A Complete Bibliography of Publications in the Journal of Mathematical Physics: 2010–2014

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

17 October 2024 Version 1.31

Title word cross-reference

(1 + 1) [1000, 1906, 294, 2457]. (1 + 2) [1493, 1654]. (1, 0) [2095]. (1/p)^{n_{q}n} [1052]. (2 + 1) [2094, 669, 1302, 718, 449, 1012, 377, 2620, 2228]. (3 + 1) [1499]. (4, 4, 0) [2306]. (β, q) [1297]. (C, +) [1885]. (D + 1) [2054, 2291]. (d + s) [2255]. (L^{2}, Γ, χ) [1885]. (N + 1) [1334, 155]. (n + 3) [490]. (N, N′) [1789]. (p, q, ζ) [500]. (p, q; α, β, ν; γ) [1113]. (q, μ) [500]. (q, N) [1659]. (R, p, q) [300]. (SÔ_{(q)}(N), Sp̂_{(q)}(N)) [1659]. + [2688]. −1 [1394]. −1/2 [977]. −a/r + br^{2} [945]. 1 [2659, 1714, 1004, 1212, 632, 2154, 694, 1952, 354, 661, 1985, 752]. 1 + 1 [2332]. 1 + 2 [2484]. 1/2 [1004, 144, 759]. 1 < α ≤ 2 [598]. 2 [518, 2225, 2329, 1, 1009, 2562, 2251, 1903, 1947, 1352, 1597, 465, 2675, 454, 891, 899, 2031]. 2 + 1 [884, 938, 217, 681, 939]. 2d [356]. 2N [1406]. 3 [287, 1875, 1951, 2313, 2009, 518, 2155, 799, 1095, 810, 2553, 2260, 2579, 2067, 1882, 2554, 1340, 2251, 1069, 2257, 2169, 1006, 1992, 2195, 2289]. 3 + 1 [1939]. 4 [2402, 874, 585, 1906, 640]. 4 + 2 [1939]. 4d [356]. 5 [80, 1560, 924, 2046]. 50 [22]. 6j [2111]. 9 [692]. ^{(1)} [381]. ^{∗} [628, 491, 236]. _{0} [1462]. _{2} [1398]. _{3} [1272]. n [2517, 1603]. A, D, E [655]. a = 4 [379]. A_{3}^{∗} [2169]. A_{1}^{(1)} [2382]. A_{2}^{(1)} [2382].

A_{2}^{(2)} [2382]. A_{4}^{(1)} [1155]. A_{n} [2568]. A_{r} [356]. AdS_{2} [1004]. α [488, 136]. α∗ [1899]. B [705, 286, 225, 710, 567]. B(m, n) [1507]. B_{∞∞}^{−1(ln)} + B_{X_{r}}^{−1+r,2/(1−r)} + L^{2} [1798]. B^{2,s} [1777]. B_{2} [412]. B_{3} [1974]. β [1431, 428, 1469, 2287, 2419, 132]. βγ [411]. C [1677, 2156]. C∗ [2213]. C^{∗} [1600, 802]. C^{1} [1391]. C^{k} [2269]. C_{3} [1974]. χ^{2} [543]. D [471, 2636, 2331, 1492, 1247, 572, 1915, 134, 1954, 945, 2050, 963]. D = 11 [1817]. D = 2 [781, 1436, 944]. D = 4 [1817]. D_{N} [782]. D_{n}^{(1)} [1777]. † [903]. ẍ + f(x)ẋ^{2} + g(x) = 0 [2719, 2718, 2707, 1836, 2325]. δ [1807, 2163, 2294]. E_{10} [1988]. E_{6} [2188, 1946]. E_{7} [97]. E_{8} [97]. ℓ [2012]. End_{u_{k}(2)}(Ω_{k}^{⊗r}) [929]. ϵ(3) [622]. f [2684]. f(R) [1025]. F^{\frac{⊗1}{2}} [2537]. G [2536, 2153, 1803, 2459]. G_{2} [2135, 412]. G_{2}(1) [1119]. G_{2(2)} [1644]. GF (p^{ℓ}) [202]. GL(3, C) [2195]. gl(m|n) [2675, 1647, 2154]. H [1621, 2039, 1088, 2496]. H^{n}(≀∫√(1|2), M) [447]. P̂ [97].

K [289, 2094, 657, 1460, 2016]. κ [791]. l [47, 1336, 2519, 2016]. L^{2} [2573, 925]. L^{p} [332, 297, 918, 1933]. L_{2} [2106]. Λ [2039]. M [447]. M_{2}(C) [696]. M_{2} ×_{W} Σ_{d−2} [1770]. M_{g} [988]. cH^{N} [1345]. CP^{1} [1781]. CP^{N} [2658, 1345]. D_{2} [1503]. E^{3} [1040]. Q/Z [826]. R^{+} [244]. R^{1|n} [184]. R^{3} × M [538]. R^{4} [1107]. R^{N} [1709, 338]. Z^{2} [1721]. Z_{2} [257, 1167]. Z_{3} [1593]. z_{4m} [933]. ⌉_{6}^{(1)} [2188]. {_{4} [248]. }_{2} [248]. }↕(m|n) [453]. ∫↕(2|1) [1388]. ∫↕(3, C) [1003]. ∫↕(n, C) [735]. ∫↕_{−1}(2) [1699]. ∫√(4, C) [1003]. N

[996, 2508, 1, 2232, 1349, 938, 427, 270, 191, 1567, 1293, 773, 115, 112, 1439, 1400, 2419, 533, 1068, 1002, 239, 550, 1281, 490, 1831, 354, 296, 175, 2577, 686]. N = (0, 4) [555]. N = (2, 2) [2492]. N = 1 [2337, 2519]. N = 2 [2012, 135, 1632, 379, 385, 1336, 356, 1273]. N = 2, 4, 8 [1247]. N = 4 [2306, 632, 1487, 2169]. N = 4, 7, 8 [1492]. N = 5 [924, 1289]. N = 6 [2155]. N = 6, 8 [1095]. N = 8 [472, 1909]. O [856]. O(3) [908]. p [1298, 2629, 2561, 220, 1101, 1664, 2576]. ∂ [403]. Φ^{2} [437]. Φ^{2k} [414]. Φ_{2}^{4} [1226]. π [1370, 2640, 281]. pp [1065, 1688]. psl(2|2) [2460]. P T [2076]. Q [1881, 1878, 840, 2388, 1297, 432, 2383, 2489, 79, 427, 281, 1196, 2172, 584, 833, 404]. q > 1 [833, 136]. q^{2} [860, 1662]. q^{±2} [860]. R [1388, 2660, 1403, 1622, 1130, 2026, 1603, 388, 1399, 2056, 2432]. R^{2} [2017, 2151]. R^{3} [2107]. R^{6} [444]. R^{d} [2153]. R^{N} [2105, 2099, 1307, 1943]. ρ [1357]. S [843]. s = −1 [1338]. S^{2} [2027, 1406]. S^{3} [850]. sh(2|2) [2055]. σ [2659, 632, 2488]. sl(2, C) [2186]. SL(2, C) [606, 411]. sl(3, C) [1945]. sl(4, C) [622]. sl(m + 1|n) [2386]. sl(n) [1153]. SL(N, C) [672]. sl(n^{k}, C) [401]. SO(10, 2) [865]. so(2, 2k + 2) [1826]. so(3, R) [2631]. so(3, R) [2058]. SO(4) [2494]. SO(5) [126, 448]. SO(8) [692]. SO(9) × SU(2) [556]. SO_{q}(N) [1944]. Sp(m) [243]. SU(1, 1) [2270, 606, 576, 2116]. SU(2) [1681, 360, 2502, 556, 581]. SU(2) ⊗ SU(2) ≈ SO(4) [2494]. SU(2, 2) [353, 2461]. SU(3) [1681, 320]. SU(6) ⊃ SU(3) ⊗ SU(2) [753]. SU(8) ⊃ SU(4) ⊗ SU(2) [753]. SU(N) [672, 781, 434, 2131, 2184, 2636, 780]. SU(N) × SU(N) [1039]. SU(r) [1571]. SU_{2,1} [900]. SU_{3} [900]. ∑_{i=1}^{d}(−∂_{i}^{2})^{s} [2050]. T [274, 2534, 2644]. T^{d} [1964]. T_{0} [1536]. T_{c} [878]. τ [2442, 2062]. θ [88]. u [2389]. U(1) [2670, 1141, 539, 273]. U(2, 2) [2387]. U(N) [360, 795]. U(Ω) ⊗ SU(r) [1571]. U_{q}( ˆ∫↕(M + 1|N + 1)) [1785]. U′_{q}(ŝl(2)) [2489]. U′_{q}(so_{3}) [771]. U_{q}(ˆ∫↕(N|1)) [1401]. U_{q}(ŝl(N|1)) [1124]. U_{q,p}(A_{N−1}^{(1)}) [619]. U_{q}[gl(2|1)] [1078]. Uq′(ŝl(2)) [1662]. V [2688, 1973]. V_{4} [411]. φ [1196, 1803]. φ^{4} [1262, 1458]. ϑ [1031]. ∨ [2569]. W [60, 408, 572]. W (2, 2) [60, 1892]. W_{∞,p} [1913]. W_{∞}^{N} [848]. W_{p,p′} [1354]. wU_{r,s}^{d}(osp(1|2n)) [2112]. X_{2} [112]. X_{m} [2118]. Y (∫↕_{N}) [1674]. Z [2276].

-actions [2153]. -adic [2561]. -affine [1000]. -algebra [60]. -algebras [628, 1095, 1167]. -ary [1068, 1006]. -average [427]. -bar [963]. -bigradings [257]. -body [2553, 1906, 2046, 175, 1567, 1293, 115, 1334, 1439]. -bounds [1933, 918]. -bridge [2562]. -categories [903, 2534, 2644]. -cohomology [1899]. -completely [488]. -component [1009, 454, 899]. -condition [2389]. -conformal [2012, 1336, 2519]. -constants [1031]. -constrained [2094]. -coordinated [1803]. -cosmology [1025]. -cosymplectic [657]. -coupled [191]. -covariant [1903]. -critical [925]. -crystal [1777]. -curvatures [2289]. -d [1352]. -deformed [1878, 1113, 2383, 404, 300, 1622]. -Derivations [1196]. -dim [640]. -dimensional [2094, 669, 2620, 490, 2402, 874, 2050, 449, 2554, 1499, 1906, 692, 2251, 294, 939, 1493, 2228, 1012, 377, 891, 572, 1915, 2054, 2291]. -dimensions [1954, 945, 2508]. -divergence [543]. -dressing [403]. -elliptic [1153]. -ensemble [428]. -ensembles [1431]. -entropy [833]. -equation [225, 710, 567]. -equivariant [1803, 320]. -exponential [500, 281, 584]. -exponentials [432]. -Fock [1196]. -fold [1992, 2195, 112, 1400]. -form [1460]. -function [2062]. -functions [2442, 1621, 2636]. -gauge [2289]. -Gaussian [1297]. -generalization [2684]. -generalized [1297]. -geometry [2135]. -graded [79, 924, 2156, 1593]. -Green [297]. -Hahn [860]. -Hermite [132]. -integrability [1391]. -invariant [2670, 243, 273]. -Jacobi [1469, 1662]. -Kampé [860]. -Korteweg [379]. -Laguerre [112]. -Laplacian [220, 1101, 1664, 2576]. -Leibniz [686, 2016]. -Lie [1068, 287, 1875, 239, 550, 2260, 2579, 490, 1831, 1069]. -matrices [1, 2026, 388, 1399, 2056, 97]. -matrix [1603, 2432, 1388, 2660, 1130]. -model [2488]. -models [2659, 632, 2287]. -module [2331, 1492, 1247]. -modules [236, 802, 2213]. -norm [2573]. -Onsager [2388]. -operation [1403]. -operators [856]. -particle [354, 270]. -perturbed [1357]. -photon [2031]. -ple [2016]. -plurality [274]. -point [354]. -positive [465]. -pseudoalgebraic [1088]. -pseudodifferential [2496]. -representable [1]. -soliton [1349, 773, 1789, 533]. -space [585]. -spaces [2039]. -sphere [1951, 810, 1281, 1882]. -spin [1298, 2459]. -stability [2106]. -stable [136]. -state [47]. -symbol [2269]. -system [411]. -systems [2569, 2536, 2153]. -tensor [1600]. -ternary [491]. -Theory [1262]. -theta [1885]. -time [1686]. -topology [1536]. -torus [2067]. -totally [2172]. -type [1807, 782, 408, 705]. -unit [427]. -valence-bond-solid [840]. -vortex [996, 1406]. -wave [2629, 2255, 2577]. -waves [1065, 1688].

/anti [649]. /anti-BRST [649].

1/2 [2416].

37 [109].

43 [942]. 45 [2295, 110]. 46 [726, 2183]. 47 [2326]. 48 [157]. 49 [2021, 2621, 262].

50 [261, 197, 263, 1252, 775, 1083, 511, 1373, 397, 198, 200, 1507]. 51 [863, 600, 510, 635, 111, 2620, 2708, 599, 509, 636, 776, 899]. 52 [1375, 2020, 1887, 774, 2626, 1129, 1128, 823, 1374, 2619, 1933]. 53 [1611, 2572, 1409, 1931, 1666, 1612, 1665, 1932, 1251, 1934, 2709, 1508, 2253, 1217]. 54 [2717, 2625, 2713, 2531, 2719, 2712, 1935, 2064, 2710, 2718, 2715, 2707, 2325, 2184, 2182, 1983, 1984]. 55 [2622, 2451, 2362, 2707, 2623, 2294, 2532]. 56 [2666, 2662, 2714, 2624]. 57 [2705, 2711]. 58 [2716].

80 [199]. 80th [1410].

= [2282].

ABC [1156]. ABCD [2633]. Abel [1163, 2627, 2205, 361, 273]. abelian [1260, 1684, 1700, 1781, 437, 1976, 2132, 1258, 271, 948, 1050, 555, 692, 761, 484]. Ablowitz [237, 2514, 1783, 1929, 1838, 1845]. above [1192, 1470, 1757]. above-threshold [1757]. Abraham [2033]. abruptly [892]. Absence [2250, 1033, 541, 1421]. Absolutely [1415, 2115, 541, 2266]. absorption [1855, 1751]. abstract [1712]. accelerated [1090, 568]. Accelerating [1182, 224]. acceleration [1753, 687, 1187, 2645]. according [1998]. accumulation [664]. acknowledgment [2254]. acoustic [1500]. Action [1732, 1722, 1355, 1681, 2603, 565, 257, 2088]. Action-angles [1732]. actions [231, 232, 1426, 2153, 2238, 2565, 2331, 1679, 1019]. active [1384]. actuators [1147]. adding [2542]. additional [1261, 816, 1706, 1635]. additive [2315, 514, 301, 438, 875]. additive-cubic [301, 438]. adiabatic [2651, 1966, 1449, 593, 558, 63]. adic [2561]. adjoint [1696, 980, 947, 1752, 2717, 1092, 2245, 765, 1762, 1978, 2176, 45]. adjointness [2303, 1229]. Adler [2293]. admitted [757]. admitting [1730, 1558]. AdS [1865, 1272, 165, 151]. AdS/CFT [165]. aether [1772]. Affine [1125, 2014, 2064, 217, 2382, 2357, 2286, 1489, 1580, 1802, 1000, 2567, 253, 2188, 1841, 1179, 2345, 1390, 1124, 1323, 381]. affinities [561]. affinizations [670]. against [839]. age [1963]. age-structured [1963]. aggregation [1519]. AGT [581]. Aharonov [846, 1165, 1261, 1165, 1676, 778, 2262, 2677, 785]. Airy [507, 1398].

Akhiezer [1810]. AKNS [1441, 752]. al [198, 1924, 182]. Al-Salam [182]. Albert [248]. Alexander [720]. Alexandrov [2564]. algebra [1354, 1707, 2286, 1113, 2388, 2539, 1925, 2525, 510, 451, 518, 549, 789, 1817, 250, 149, 689, 2383, 811, 1003, 622, 1701, 2329, 380, 2614, 1612, 1198, 1344, 1703, 935, 1699, 1756, 1981, 966, 2461, 771, 1900, 217, 950, 514, 60, 1865, 2345, 933, 99, 1988, 196, 385, 2087, 2673, 101, 739, 248, 1336, 1440, 2231, 2152, 1892, 1622, 1377, 607, 165, 1153, 2661, 1833, 1953, 2182, 213, 1914, 2157, 2459, 396, 2116, 2457]. algebra-valued [2661]. Algebraic [997, 2155, 1613, 37, 1920, 926, 905, 648, 1300, 2395, 1769, 587, 1790, 1070, 1064, 1949, 651, 1495, 1922, 742, 1993]. algebraically [557]. Algebras [1928, 2319, 2521, 2427, 627, 2012, 2110, 195, 1068, 628, 287, 1875, 2384, 1489, 2155, 630, 769, 1580, 1659, 239, 1095, 1926, 1802, 412, 2458, 550, 2260, 2579, 2660, 2188, 1701, 2329, 491, 2503, 2637, 652, 850, 79, 490, 597, 1675, 256, 2495, 1841, 2332, 300, 1801, 31, 1492, 924, 1831, 1247, 2156, 1803, 2222, 1919, 459, 1103, 1893, 1069, 1019, 2257, 198, 212, 1438, 1186, 1006, 12, 296, 586, 1399, 2056, 2432, 843, 1167, 1008, 803, 2534, 2644, 2290, 2404, 444, 1838, 634, 686, 2016, 790]. Algebro [1916, 1588, 2436, 1356]. Algebro-geometric [1916, 1588, 2436, 1356]. algebroid [2235]. algebroids [2611]. algorithm [672, 508, 600, 975, 2526, 503, 599]. Algorithmic [2144, 1958, 2405]. Algorithms [1748, 2636, 687]. Allard [1607]. alloys [277]. Almost [2698, 36, 2528, 1584, 1864, 2004, 1847]. almost-commutative [2528]. along [2646, 2321]. alternate [1663]. alternative [1434, 1611, 1051, 2005, 160, 1325]. Ambjorn [1366]. among [1246, 2237, 988]. amounts [2079]. amplifier [2270]. Amplitude [2616, 914]. Analog [2634]. Analogue [1842]. Analogues [2388]. Analysis [1236, 980, 1626, 2226, 238, 1880, 1483, 1467, 2280, 1372, 712, 1000, 583, 2011, 588, 2710, 1184, 2555, 2665, 744, 2604, 2715, 1918, 957, 2309, 1551, 1210, 2523, 299, 1813, 2716, 1296, 568, 972, 1867, 1495, 2015, 1908, 2711, 1080, 1364, 2396, 2700, 918, 1933]. Analytic [1958, 2571, 1436, 2394, 610, 369, 1566, 2126, 998, 975, 727, 2071, 1465, 418, 2498, 134, 2351, 485]. Analytical [2508, 445, 1496, 1809, 2524, 1499, 891, 180, 706, 989]. Analyticity [854, 2134]. ancilla [2688]. ancilla-free [2688]. Anderson [2462, 2052]. Ando [1861]. Angelesco [1598]. angle [1604]. angles [1732]. Angular [1037, 923, 700, 371, 1537, 796, 616]. anharmonic [624, 774, 646, 1464]. anhomomorphic [512]. Anisotropic [1772, 2435, 926, 885, 1568, 2359, 1867, 2060, 578]. anisotropy [1268, 1855]. annihilation [1619]. Announcement [1613, 2254]. annular [857, 177]. anomalies [2113, 2083, 1066, 1141, 1227]. Anomalous [2687, 180, 1444, 919, 133]. anomaly [358, 131]. ansatz [997]. antenna [2653]. anti [2646, 63, 615]. anti-BRST [649]. anti-de [2646]. anti-self-dual [63]. anti-Skyrmion [615]. anticommutation [99]. antidegradable [2469]. antiferromagnetic [2483]. antiperiodic [487, 675, 1846, 1234]. antisymmetric [428, 92, 341]. Any [47, 1225]. anyons [1324]. appearing [644]. Appell [2284]. Application [1260, 2710, 1608, 1908, 2711, 2116, 137, 1795, 2071, 56, 111, 1435, 2703, 2677, 1130, 693, 2431, 1783, 330].

Applications [1246, 1096, 2542, 1042, 1327, 1125, 467, 1926, 2320, 623, 251, 1112, 2212, 1428, 1669, 2494, 352, 99, 441, 351, 1550, 934, 1328, 565, 684, 2430, 346, 2078, 2109, 851, 444]. applied [369, 2007]. approach [926, 2522, 1594, 2312, 2073, 2144, 1221, 180, 2313, 1051, 868, 1443, 182, 2092, 255, 391, 267, 304, 2602, 2395, 208, 949, 29, 1812, 540, 2179, 744, 1349, 2653, 2627, 1055, 117, 268, 2121, 739, 176, 50, 2488, 2393, 2170, 1325, 2350, 457, 773, 2365, 2091, 2227, 1384, 580, 2609, 716, 1993, 1447]. approaches [2671, 2542, 2056]. Approximate [485, 2708, 102, 198, 12, 791, 1020]. Approximately [1814, 295]. Approximating [163]. approximation [1295, 575, 2316, 2441, 1291, 2668, 2664, 1842, 2372, 2609]. approximations [1451]. Arbitrarily [2079, 2588, 2408]. Arbitrary [620, 1818, 2510, 463, 1392, 2043, 2282, 1463, 1537, 47, 364, 1809, 270, 2589, 531, 2145, 2334, 2655]. arc [1108]. Archimedean [103, 438, 1158]. area [2676, 960, 1766]. argument [1035]. arising [866, 1356, 1046, 13, 200, 424]. Arithmetic [84, 2402]. array [1835]. arrays [480]. arrival [1249]. artificial [2465]. ary [1068, 1006]. Ashtekar [2595]. Aspects [2114, 586, 1958, 2625, 2212, 191, 351, 1839, 1326]. assignment [601]. associated [2549, 1441, 241, 411, 2383, 2188, 1885, 632, 1270, 1647, 401, 735, 2494, 492, 2148, 2194, 974, 2058, 2631, 1119, 1777, 2444, 1383, 2612, 1482, 112, 1648, 1303, 961, 2530, 378, 444]. associative [1893]. Associativity [2691]. astigmatic [2558, 364]. asymmetric [1445, 1760, 1935, 1884, 2500, 2315, 2162, 216, 1855, 2172, 279, 665, 2047]. asymmetry [1917, 878, 2067]. Asymptotic [736, 1483, 91, 1834, 1497, 2548, 1382, 639, 1184, 220, 754, 1144, 1367, 331, 1244, 2327, 219, 1151, 745, 875, 1643, 644, 2509, 786, 2027, 1832, 1609, 2111, 1638, 675, 797, 934, 1309, 207, 276, 1121, 1564, 2670, 2513, 132]. asymptotical [2582]. Asymptotically [1996, 1624, 824, 1327, 2409, 1852, 1463, 321, 1369, 1724]. Asymptotics [591, 2068, 682, 2263, 703, 1795, 2074, 2669, 2337, 536, 25, 743, 1368, 594, 1152, 870, 1194, 1915, 1872, 2496]. Atiyah [1139]. atlas [2168]. atom [1804, 1433, 1384, 538]. Atomic [630, 1246, 33, 1073, 587, 2394, 308, 636]. Atoms [2544, 1423, 733, 2715, 1067, 1813, 2716, 825]. attempt [2166]. attenuation [423]. attraction [1963, 916]. attractive [2353, 1376, 1054, 977]. Attractiveness [1466]. attractor [367]. Attractors [2000, 1940, 472, 2693, 1331, 1332, 1101, 1309, 68, 1965, 1393, 128, 419, 127, 1190]. Attributing [656]. auroral [425]. Auto [146, 1736]. Auto-Bäcklund [146, 1736]. automata [25, 2591]. automaton [1119, 1777, 1997]. Automorphic [1988]. automorphism [738]. automorphisms [209, 2474, 1188, 2038, 1932]. autonomous [2139, 1191, 1309, 2499, 1965, 1393, 1564]. autoresonance [2358]. auxiliary [1450, 927, 1383, 115, 783]. Average [631, 1022, 1475, 427]. averaged [2355]. Averaging [1291]. avoiding [1429]. axial [2688, 358, 369, 177]. axially [2478]. axiomatic [699, 571]. axis [1397]. Axisymmetric [1365, 481, 362, 2201]. azimuthal [446].

B [2456, 1992, 2195]. back [2141]. background [846, 151, 1467, 2587, 1140, 1061, 468]. Bäcklund [2620, 146, 135, 1736, 1587, 449, 532, 1799, 629, 886]. backtracking [2421]. Baker [1810]. Balaban [1998]. balance [2059, 1754, 1828, 616]. Balinskii [670]. ball [2590, 2471]. ballistic [1415, 180]. balls [971]. Banach [198, 1696, 12]. Bañados [742]. Band [664, 878, 645]. bar [963]. Bardina [799]. Bargmann [1281, 1214, 2148]. Barnes [214, 1159, 1635]. baroclinic [1512]. barothropic [2309]. barotropic [760]. barrier [530]. Barut [2584]. Barycentric [52]. Base [2197, 1119]. based [846, 2010, 1176, 2259, 2678, 2461, 2627, 1523, 394, 2244, 2121, 2058, 1150, 1217, 2586, 1886]. bases [108, 2245, 448, 630, 606, 1023, 315, 1841, 1546, 382, 1284, 84, 1058, 1399, 97, 2119]. basic [2648]. basis [2593, 2622, 861, 2193, 2407, 2301, 1073, 2071, 1885, 2148, 1078, 2112, 2394, 1897, 1950, 453, 595]. Batalin [35]. bath [646]. baths [1066]. Bavand [198]. Baxter [856, 1875, 2223, 881, 1881, 2186, 803]. Baxterized [2286]. Bayes [1224]. BBM [1937]. BC [1002, 2517, 1603]. beam [249, 1250, 2070, 340]. beams [708, 662, 1765, 2151]. Becchi [649]. Becker [922]. behavior [2670, 786, 878, 921, 1598, 1512, 458, 2256, 1609, 1938, 1305, 822, 1244, 1121, 1973, 875, 1564, 311]. behaviors [2310, 2232, 745]. behaviour [1834]. Bell [1251, 2684, 1112, 1321, 310, 1674, 1135, 532, 1420, 773]. Bell-polynomial [532, 773]. Bell-type [1251, 1135]. Belov [2361]. Bénard [940, 2480]. Bend [892]. Bend-imitating [892]. Bender [2193]. Benjamin [1937]. bent [892]. Berezin [2269]. Bergman [1117]. Bernoulli [1948, 249, 1250, 665, 260, 2051]. Berry [2267, 555]. Bertotti [2037]. Bertrand [2428, 1786]. Besov [1338, 1127, 424]. Bessel [1912, 722, 820, 2512]. Bessis [393]. beta [1980]. beta-Wishart [1980]. Bethe [997, 1976]. between [507, 1545, 406, 1543, 1666, 275, 1840, 2517, 902, 415, 1456, 2647, 394, 496, 1077, 1459, 1273, 1791, 2439, 145]. beyond [80, 1387]. BGK [589]. bi [1582, 2307, 1257, 2631, 1495, 812, 156, 285, 378, 333]. bi-Darboux [1582]. bi-Hamiltonian [2631, 1495, 156, 285, 378, 333]. bi-isotropic [812]. bi-qutrit [1257]. bi-spinors [2307]. bialgebra [2637]. bialgebras [2173, 1708, 2290]. Bianchi [275, 418, 1461, 1462, 1328, 1025]. Bianchi-I [1025]. bicharacter [2110]. bidirectional [2094]. Bifurcation [282, 702, 821]. Bifurcations [171, 2336]. big [1662, 2489]. bigraded [194, 1126]. bigradings [257]. Bihamiltonian [286]. Bilinear [2521, 888, 194, 705, 1941]. billiards [205, 1594, 264]. bimodules [1600]. bimultipliers [491]. Binary [841, 1032, 359, 184, 1321, 282, 1305]. binding [645, 827]. binomial [2549, 1472, 2137, 881]. biological [1518, 1525]. Biomixing [1518]. biorthogonal [1795, 291]. Biot [1607, 1106, 2417]. Bipartite [1679, 1545, 1572, 1766]. bipolar [2453]. biproducts [2222]. Birkhoff [2112, 521]. Birman [196]. Birth [1854]. birthday [1410]. bistable [2308]. BKP [246, 2087]. black [951, 989, 151, 2551, 1288, 1100, 520, 835, 1290, 2647, 605, 359, 2201, 474]. blackbody [2160]. Blasius [522]. Bloch [2185, 2147, 267, 986, 2590, 1792, 1343]. Block [2087, 1126]. blocks [193, 2703]. Blow [1576, 2256, 567, 2002, 2108, 2242, 1750, 1009, 1527, 1127]. Blow-up [1576, 2256, 567, 2002, 2108, 2242, 1009, 1127]. blowup [1011, 509, 143, 454, 891, 899]. bodies [2647, 684]. body [114, 2415, 993, 1392, 2043, 1639, 2136, 2553, 2395, 1608, 2121, 1906, 2046, 1567, 1293, 1867, 2641, 115, 1334, 1907, 1439, 1571, 701, 175, 955, 1613]. Bogoliubov [2713, 2158, 2006, 1403, 575, 2372]. Bogomolny [215, 970]. Bogomol’nyi [2658]. Bogomolov [2474]. Bohm [1261, 1676, 778, 2262, 2677, 785]. Bohmian [2121, 553]. Bohr [1816]. Boiti [2620, 449, 378]. Bollobás [893]. Boltzmann [2623, 2418, 1483, 332, 2106, 1911, 2298, 1842, 2613, 589]. Bona [1937]. bond [840]. Bonelli [1612, 1198]. Bonnet [1660, 741]. Boolean [547, 1167]. Borchers [1535]. borderline [2580, 1510]. Born [658, 1408, 2065, 2214, 2316, 1433]. Bose [2312, 2333, 433, 275, 483, 1608, 1796, 228, 914, 1763, 1054, 187, 680, 2641, 2372, 2423, 572]. Bose-gas [2423]. boson [2110, 2636, 105, 2537]. boson-fermion [2110, 2537]. bosonic [231, 232, 150, 1574, 1433, 2078, 1262]. bosonization [2674, 2633]. bosons [577, 1927, 862, 1883, 1948, 866, 357, 889, 1761, 2473, 434, 1039, 2258]. Bound [446, 1241, 1762, 1442, 2162, 1804, 2163, 2294, 1650, 943, 1223, 2032, 119, 2597, 2395, 2582, 162, 540, 2250, 1056, 228, 2484, 480, 1358, 1058, 123, 2480, 2700, 468]. bound-states [468]. boundaries [1496, 1749, 2559, 2564, 2243, 210]. Boundary [89, 2613, 2623, 284, 1206, 2286, 2409, 2219, 1250, 1079, 1154, 1592, 2543, 2500, 189, 1468, 1533, 1609, 1207, 2068, 859, 619, 1785, 477, 2200, 93, 1101, 497, 1521, 870, 1234, 1915, 2047, 2480, 1526, 667, 2019]. boundary-value [1521]. boundarylike [487, 675]. Bounded [2575, 1168, 2339, 2451, 536, 1942, 355, 1429, 226, 1049, 1129, 2425, 1521, 452, 1847]. Boundedness [375, 2303]. Bounding [1285]. boundness [520]. Bounds [1994, 1013, 120, 1870, 2468, 1358, 1232, 1021, 1549, 2509, 2242, 623, 940, 833, 1527, 641, 395, 983, 1725, 2124, 410, 1242, 918, 1933]. Bourin [1738]. Boussinesq [1507, 1338, 2300, 139, 1942, 379, 426, 687, 1645, 1114]. Boussinesq- [1507]. box [461, 383]. Braid [2330, 2638]. braided [2330, 2258, 2534, 2644, 2535]. brain [1432]. branching [2248]. brane [1865]. Brans [325]. Brdicka [1065]. Breaking [2200, 1504, 669, 323, 2096, 2477]. breakup [967]. Breather [766, 1930, 2490, 1306]. breathers [1646, 1464]. Bregman [199, 19]. Bresse [506]. Brezis [976]. bridge [2562]. Bright [1663]. Brillouin [2441]. Brinkman [481]. broad [280]. broken [2268]. Brownian [879, 838, 1416, 1573, 331, 1473, 851, 1123, 637]. BRST [649, 272, 649]. Brusselator [715]. BTL [816]. bubbling [1212, 1985]. Büche [481]. Buckley [1276]. Bulk [1431, 2695, 2287, 499]. bump [806, 2107]. bundle [883, 2563, 643, 2321, 484]. bundlelike [559]. bundles [290, 508, 600, 1096, 720, 503, 599]. Burgers [2710, 2711, 1986, 1317, 1395, 542, 336, 27, 934, 1746, 131, 1908]. Burghelea [2699]. buried [1844]. Büttiker [2365].

C [491, 236]. Cahn [1750, 1169]. calculation [672, 1436]. calculations [1950]. Calculus [811, 137, 1299, 495, 656, 2066, 1977, 594, 792, 131, 2435]. Caldirola [1539]. Callan [207]. Calogero [241, 782, 709, 1587, 1266, 1350, 1351]. Calogero-type [709]. Camassa [899, 326, 2454, 1009, 1938, 2140, 1105, 1799, 721, 864, 1127, 1316, 2378, 1711, 247, 420, 454]. Can [553]. cannot [2233]. Canonical [1301, 697, 2403, 2311, 492, 2564, 2683, 995, 344, 448, 1452, 373, 582, 99, 2650, 359]. Capacities [832]. capacity [314]. caps [668]. capture [837]. Caputo [2444]. CAR [514]. Carlitz [182]. carrying [2049]. Cartan [769, 1946]. Cartesian [818]. cascades [2009]. case [1192, 591, 2475, 1258, 1532, 2315, 387, 1050, 1518, 70, 2399, 1166, 1770, 1265, 817]. Casher [846, 1165]. Casimir [284, 867, 1440, 1459, 1915, 2054, 2291]. Catalan [181]. categorical [2092]. categories [1600, 1653, 903, 2289, 2389, 2534, 2644, 2535]. category [2223, 2258]. Cauchy [579, 1795, 805, 690, 2598, 1775, 710, 2425, 2412, 651, 2097, 305, 589, 2464, 2700]. Caudrey [629]. Causal [1535, 650, 2133, 319, 1870, 1188, 2038, 1932, 1586]. causality [2305]. Causally [1287]. Cause [1769]. causes [1554]. caustic [655]. Caustics [80]. cavity [1298, 1915, 308, 636]. CC [2343]. cell [2643]. cellular [25, 1119, 1777, 1997]. Center [659, 771, 835, 53, 359]. Center-of-mass [659, 53, 359]. centered [77, 1910]. Central [1567, 100, 1235, 2579, 2508, 2332, 30, 2046, 1719, 641, 175, 955]. centrally [2460]. centrifugal [2708, 102]. certain [2605, 2673, 2104, 153, 496, 362, 489, 2166, 2655]. certificates [393]. CFT [165]. CH [882]. Chain [546, 2676, 2202, 2355, 1785, 2114, 1881, 316, 1045, 664, 2224, 2516]. chains [1779, 1396, 1163, 448, 1659, 406, 2113, 970, 750, 1854, 38, 1846]. challenges [40]. Chaltikian [2361]. Chandra [2457]. Change [2287]. changing [1048, 766]. channel [2653, 2078, 2323]. channels [2079, 2588, 314, 832, 2469, 2341, 1178, 2230, 737, 1547, 1093, 2197, 1022, 1077, 2078]. chaos [2237]. chaotic [609, 477]. chaoticity [2049, 2148]. Chaplygin [1065]. character [1090]. Characterisation [1614]. Characteristic [1542, 2665, 2513, 2479, 2532, 473, 1672, 631, 280, 519]. characteristics [904, 76]. Characterization [1696, 1437, 2469, 571, 1874, 1188, 2038, 1932, 1786, 1573, 1827]. Characterizations [779, 2323]. characters [58, 2109]. charge [1677, 2610, 956, 86, 1999, 2647, 2298]. charged [926, 971, 708, 1691, 2005, 743, 1213, 1470, 1461, 474]. charges [2416, 1690, 576, 1627, 1947]. Charlier [2055]. Chasles [1687]. Chebyshev [2525]. chemical [2424, 2256, 28, 561, 2131, 2184]. Chemically [1888]. chemo [1936]. chemotaxis [1518]. Cherednik [1675]. Chern [697, 2132, 1698, 1474, 193, 2630, 2210, 877, 437, 1095, 908, 1684, 1310, 1668, 1819, 1820, 2347, 558, 1700, 1770, 2482, 1014, 2103]. Cherry [2125]. Chiral [2012, 2131, 2184, 507, 2113, 1141, 413]. Choi [737]. choice [2334]. Choquard [2328, 2573, 1215, 2221]. chord [817]. chromodynamics [1138]. circle [510, 451, 225, 1732, 1197]. circuits [2688, 2404]. Circular [1108, 1311, 1693, 1086]. Cisneros [1267]. Civita [1696, 796]. clamped [1793, 1983]. Class [1005, 65, 1218, 1047, 703, 722, 1590, 2415, 473, 876, 2122, 2463, 2380, 2474, 471, 1996, 2188, 1496, 2559, 2702, 2025, 221, 2304, 2362, 64, 69, 1787, 1294, 1029, 2107, 897, 1952, 1578, 640, 502, 2044, 1671, 2714, 2576, 1618, 127, 686]. class-preserving [2474]. classes [2580, 1659, 1452, 1469, 129, 1960, 2486, 1062, 2626, 1687, 1944, 2478]. classic [1907]. Classical [2418, 1879, 2625, 1175, 1634, 1060, 1148, 912, 2529, 963, 1103, 1839, 2026, 2048, 917, 1192, 2263, 856, 2269, 659, 747, 1620, 1767, 2364, 1853, 2329, 1840, 1212, 1104, 956, 1913, 75, 1841, 707, 838, 933, 349, 2244, 2200, 1693, 329, 1490, 159, 1830, 161, 2186, 614, 1377, 388, 1399, 2056, 2166, 1335, 1381, 1705, 1262]. Classification [1560, 2384, 2719, 1205, 486, 335, 884, 490, 2461, 2565, 2718, 1257, 385, 2707, 1593, 2446, 1836, 2325, 1402, 1822, 1771, 1590, 390, 318, 1003, 2702, 1103, 1688, 1947, 2650, 681, 1695, 2586, 2714, 2201, 1899]. Clauser [1554]. clean [1814]. Clebsch [672, 1205, 753, 1699, 453, 886]. click [1387]. Clifford [2521, 2306, 25, 2589]. clocks [1123]. cloning [314, 1104, 216]. close [525]. closed [626, 530, 2437, 1944]. closure [1477, 2005]. cluster [1073, 716]. clustering [1519]. CMV [2491]. co [2535]. coalgebra [354]. coalgebras [2427, 2644, 2535]. Coarse [1857, 554]. Coarse-grained [1857]. coated [1364]. cochain [254]. codes [647, 1285, 2402, 179]. codimension [80]. coding [544, 1767, 695]. coefficient [530, 191, 972, 1695, 890, 760, 1658]. coefficients [672, 448, 1599, 1205, 753, 1480, 1762, 2502, 2559, 1608, 1499, 487, 675, 192, 1244, 2215, 453]. Coercivity [1514]. Coherence [1432, 2571]. Coherences [1090]. Coherent [2572, 1282, 3, 2111, 2504, 942, 108, 1808, 2682, 2387, 2092, 1717, 2383, 1602, 1991, 1281, 795, 536, 2608, 904, 1078, 1619, 1086, 2270, 1719, 206, 1377, 519, 586, 1648, 1085, 780, 381, 2584, 300]. coherently [1663]. cohomologies [2025, 1427]. Cohomology [508, 600, 1096, 720, 503, 599, 184, 1878, 518, 549, 789, 1817, 1694, 129, 1732, 1899]. coin [736]. coincidence [1140]. Colbeck [2679]. cold [749]. collapse [925, 322, 1187]. Collective [523]. Collineations [1903]. collision [664]. collisional [714]. collisions [2641]. Collocation [535]. Colombeau [131]. Combinatorial [2073, 2144]. Combinatorics [354]. Combined [2542, 2543, 509, 143]. Cometary [1269]. Comment [2666, 2572, 261, 1409, 1931, 2625, 157, 2719, 197, 1322, 2710, 2620, 2708, 1887, 1083, 774, 1665, 2718, 2715, 2662, 2716, 2709, 2720, 2707, 198, 2253, 942, 2714, 2711, 262, 109, 2624, 2097]. Common [1769, 1554]. communication [159]. commutant [411]. commutation [1042, 1585, 582]. commutative [1927, 1879, 2528, 1685, 2704, 1806, 1981, 2104, 1899, 1329, 1458]. Commutativity [593]. commutators [1862]. commuting [36, 2058, 1724]. compact [1696, 1230, 1698, 1978, 2701, 2473, 2664, 1091, 2618]. comparative [238]. Comparison [2552, 2671, 1097]. compass [2168]. Compatible [643]. compensated [2247]. complementarity [37]. complementary [351, 84]. complementation [1898]. Complete [1590, 390, 1897, 1563, 648, 1347, 2238, 2450, 2053, 1846, 164, 763, 1953, 2182]. completed [2517]. completely [488, 2476, 1265, 2404]. Completeness [2412, 2165, 903, 1437, 1638]. Completion [1601]. Complex [2697, 1696, 2357, 2164, 383, 231, 232, 1395, 49, 2307, 1964, 811, 251, 122, 1835, 1723, 349, 268, 27, 5, 110, 2072, 2492, 292, 903, 2471, 2436]. Complex-time [2697]. complex-valued [27]. complexity [1023, 2288, 2715, 1813, 2716, 1573]. complicated [1796]. component [2638, 1637, 1504, 1657, 884, 1015, 1576, 1009, 717, 2087, 1081, 2178, 2007, 2578, 621, 721, 286, 1711, 454, 899, 2477, 1838]. components [1583, 2352, 593, 887, 2218]. composed [1583, 2352]. Composite [5, 110, 1110, 991, 1305, 484]. composites [1013, 1961]. Compound [1087, 2408, 1994]. compressible [2318, 1672, 1891, 1534, 1797, 2455, 2479, 2532, 760]. compression [630]. computation [1176, 269]. computational [508, 600]. computations [1748]. Computer [1395, 164]. computing [2651, 1023]. concatenation [647]. concave [2065, 1750]. concavity [755, 2468]. concentrated [1643]. concentration [2328, 1853, 1172, 2647, 2533, 2220]. concerning [607]. concrete [1885, 2148]. Concurrent [727]. condensate [1763, 2641]. condensates [1366, 433, 275, 1608, 1796, 914, 187, 680]. condensation [2333, 483, 2372]. Condensed [1613]. condition [559, 90, 2205, 487, 675, 477, 1816, 665, 2389, 1526, 1256]. Conditional [1030, 1224, 1517, 2406]. Conditionally [158, 691]. Conditions [2341, 238, 284, 2219, 1079, 1154, 1592, 2543, 1394, 1595, 2500, 1967, 702, 1547, 859, 2200, 1952, 2465, 2035, 2613, 2623, 1905, 2078, 1234, 1558, 2480, 2019]. conducting [2430]. conduction [83, 2481, 962]. conductivity [1311]. conductor [2610]. cone [1541, 2643]. cones [1614]. configuration [1879, 1640]. configurational [978]. configurations [2027, 2438, 1567, 1334, 175, 955]. confined [905, 1165, 304, 945, 2392, 1764, 2076, 2544]. confinement [1161, 1697]. confining [1208]. Confluent [896, 2526, 20, 775, 687]. Conformal [2600, 2234, 125, 1682, 2012, 2249, 1185, 2659, 289, 1747, 2538, 2387, 2411, 811, 1863, 2614, 353, 32, 1180, 2331, 217, 469, 950, 2332, 1843, 2156, 193, 1336, 2519, 2703, 212, 614, 165, 1649, 356, 67, 1833, 2157, 2290]. Conformally [1683, 90, 2503, 1874, 1180, 2513]. Congruence [1632]. congruences [1185]. conjecture [393, 432, 2705, 2706, 2633, 999, 1284, 1924, 581]. conjectured [1370]. conjugacy [2580, 1687, 1944]. Conjugate [314, 2088]. conjugated [920]. conjugates [1057]. Conne [2705]. connection [611, 394, 2178, 153, 496, 484, 361, 145]. connections [2569, 1557, 528, 1258, 1314, 1315, 792, 63, 608, 643]. connectivity [2462]. Connes [2706, 602]. Consecutive [2063]. Consensus [363]. Conservation [1278, 2252, 504, 505, 2671, 669, 1982, 368, 1240, 144, 759, 1866, 2040, 657, 1754, 1842, 2215, 2297, 1682, 2361, 752, 2191]. Conservative [1969, 326, 864]. conserved [1677, 2239, 2045]. conserving [2298]. consideration [1866]. Considerations [292]. considered [676]. Consistency [1990, 1409, 1222, 1931, 1136, 1665, 565]. Consistent [1227, 1269, 922, 890]. Constant [1692, 824, 1560, 2505, 2282, 275, 1891, 885, 1369, 568, 339, 1461, 74, 1309, 521, 1493, 2003]. Constantin [1516]. constants [2670, 976, 1031, 709]. constitutive [746, 1775, 748]. constrained [2557, 81, 2094, 57, 1790, 2127, 2488, 705, 1505, 2467]. constraining [1541]. Constraint [2011, 1070, 2280, 1214, 321, 415, 1361]. constraints [272, 676, 2013, 1183, 546, 2565, 1492, 1337, 2062, 2657, 1341, 1170]. construct [1723]. Constructing [1230, 2534, 2010, 294, 285]. Construction [1068, 240, 1557, 1221, 1073, 2503, 2346, 1885, 329, 382, 1830, 2471, 2584, 2110, 2671, 2678, 1024, 160, 2344, 1139, 613, 1871, 2180, 2513, 206, 2097]. Constructive [1087, 1226, 2292]. contact [2455]. contain [2233]. contextuality [1953, 2182]. continuation [2454]. Continuity [1021, 389, 1849, 2491, 833]. Continuous [2437, 399, 2217, 1570, 1415, 2357, 1633, 2448, 606, 541, 860, 1703, 2340, 2115, 1303, 1558, 677, 758, 2266]. Continuous-time [399]. continuum [2654, 1524, 841, 1032, 480, 280, 895]. Contractions [1865, 634, 1601, 1945]. Contragredient [1989, 2712]. contrast [2147]. contribution [2192, 1422, 1508, 2639, 1384]. control [391, 1386, 2574, 98, 1381, 964, 1525]. controls [2339, 2451]. convected [1523]. convection [2627, 940, 821, 958, 2480, 1216]. convective [1750, 1313]. Convergence [63, 1498, 1130, 2698, 784, 687, 762, 350]. Convergent [658, 1159, 1928]. convex [1431, 1322, 1, 1093, 2401, 479]. Convexity [1861, 2401]. cooling [2652]. coordinate [1935, 1884, 523, 2075, 556]. coordinated [1803]. coordinates [896, 250, 104, 440, 64, 698, 1283, 2446, 359, 1272]. core [48]. cored [1231, 2554]. corners [1604]. corrected [1296]. correcting [2402]. correction [1412, 1332, 1988]. corrections [1966]. Correlated [133, 1358]. Correlation [77, 2206, 308, 636, 840, 2079, 2705, 2706, 303, 870]. Correlations [618, 1667, 2424, 1424, 1554, 2034, 349, 1971, 2374, 961]. correlators [469]. correspondence [2537, 275, 1350, 1351]. correspondences [2110, 809]. corresponding [1040]. coset [344, 1632]. cosine [92]. Cosmic [1829]. Cosmic-ray [1829]. Cosmological [59, 1772, 1560, 2518, 275, 1823, 1209, 1911, 1102, 1461, 521]. cosmology [275, 740, 1025]. cost [2408]. cosymplectic [657]. cotangent [883, 643]. Coulomb [2717, 644, 1752, 1951, 1423, 2185, 1419, 894, 2301, 1806, 1001, 1442, 48, 2392, 1302, 1133, 1832, 1656, 1417, 1132, 1256]. Coulomb-distorted [1832]. Coulomb-like [1656]. Coulombic [2126]. Count [222, 2687, 2602]. countable [1696]. counter [916]. counter-propagating [916]. counterexample [839]. Counterexamples [1233]. counterpropagating [2151]. Counting [1034, 556, 842, 1555, 215, 1765]. Coupled [1236, 1605, 1784, 683, 894, 141, 1356, 2069, 2320, 673, 2475, 476, 2113, 1349, 2452, 598, 1162, 1870, 1310, 743, 609, 1804, 646, 299, 191, 1253, 2366, 2720, 1616, 2178, 496, 479, 2096, 1663, 1975, 864, 1127, 2131, 2184, 1774, 87, 1216, 677, 333, 758]. Coupling [2502, 378, 1909, 1314, 1315, 962, 1397, 156, 977, 1929, 190, 311, 468, 448, 311]. couplings [2426, 1464, 1076, 378, 190]. covariance [1778, 2200, 499]. Covariant [746, 1075, 1866, 2040, 224, 917, 2600, 431, 669, 1050, 1680, 1437, 1903, 453, 1740]. covariants [2521]. covering [1062, 2626]. Cox [55]. Coxeter [2064, 2014]. CPT [1677]. CPT-conserved [1677]. CQ [2343]. crack [1844, 106, 1237, 813]. cracks [595, 1364]. Cranking [268]. creation [59, 404]. Cremmer [1640]. criteria [2300, 844, 1576, 226, 1516, 1780, 1615, 2218, 801]. criterion [2547, 1340, 711, 1161]. Critical [1413, 1598, 1492, 1234, 1363, 1973, 1562, 185, 2108, 2006, 2713, 1529, 1617, 1172, 1870, 1518, 2098, 1305, 2099, 925, 1664, 1041, 2220, 1190]. criticality [1522]. crystal [1073, 1841, 711, 1119, 1777]. crystallographic [2014, 2064]. CTL [816]. cubic [909, 1964, 1499, 1799, 245, 1695, 621, 367, 2378, 301, 438]. cubic-quintic [1499, 1695]. cubics [633]. Cucker [482]. Cummings [1840]. cumulant [330]. curl [682]. current [2049, 645, 2371, 2066, 714, 212, 1661]. current-type [212]. currents [1683, 2581]. Curvature [2261, 1164, 554, 2582, 2505, 868, 1443, 2610, 1258, 2190, 885, 74, 2446, 1405]. Curvature-induced [2582]. curvatures [2289]. curve [815, 2017, 2321]. Curved [1660, 2701, 90, 2553, 2445, 2127, 1229]. curves [2313, 1859, 1685, 215, 1040, 2238, 1882, 2606, 1786, 258, 1586]. cuspidal [2188]. cuspon [2140]. cutoff [1610, 1634, 1680]. cutoffs [653, 429]. Cycle [1745]. cycles [2059]. Cyclic [1284]. cyclidic [1935, 1884]. cyclotomic [984]. cylinders [1835, 1574, 796, 2512, 366, 480]. cylindrical [733, 1500, 369, 177, 1784]. Cylindrically [325, 1561].

d [1560, 923, 2537, 2659, 2313, 2009, 799, 1714, 2225, 1212, 632, 694, 1340, 2087, 1952, 1947, 661, 2169, 1597, 1985, 752, 1352]. d’Alembert [706]. damped [577, 542, 805, 1991, 1713, 2351, 1645]. Damping [1702, 26, 1498, 964, 2576, 667]. Dancoff [2383]. Darboux [717, 1582, 669, 148, 969, 814, 418, 2251, 2566, 527, 233, 155, 927, 1493, 1383, 1400, 752, 333]. Darboux-integrable [717]. Darbouxian [704]. Dark [1606, 703, 1483]. Data [1659, 1957, 2129, 473, 784, 2598, 436, 2513, 1072, 1303, 977]. datum [2570]. Davey [1939]. Davies [2124]. death [1854]. Debye [481, 2170]. decagonal [813]. Decay [964, 2576, 579, 1589, 979, 338, 2024, 1735, 303, 2034, 1712, 1782, 340, 370, 1352, 2453, 2378]. decaying [2069]. decays [620]. decide [1995]. Decoherence [2545, 696, 266, 730, 1432]. decomposable [56, 111]. Decomposition [529, 344, 239, 52, 2643, 415, 1978, 2635, 1516, 1828, 1037, 1167, 2479, 2532]. decompositions [209, 2660, 1775, 1946, 1204]. decoupling [1551]. decreasing [1996]. deduced [2071]. defects [2027, 2552]. deficiencies [86]. deficiency [853, 1887]. deficit [796]. defined [2245, 1063, 1374]. defining [1031]. definite [2181, 2642, 2486]. definition [818, 549]. deformation [689, 513, 2153, 306, 1142, 379, 2345, 71, 1139, 2199, 347, 1922, 285]. Deformations [467, 2536, 2143, 987, 2167, 353, 1694, 1930, 1485, 16, 511, 2594, 1345, 388, 2063, 686]. deformative [2637]. Deformed [1348, 2684, 1878, 997, 1113, 2383, 2563, 2541, 935, 79, 300, 2053, 459, 1893, 1622, 2169, 404, 581]. Degasperis [1316, 305]. Degeneracies [825]. degeneracy [1101]. Degenerate [1646, 1891, 1874, 1686, 2452, 1337, 21, 1373, 2270, 1446]. degenerated [1888]. degradability [314]. degree [1394, 2602, 2237, 778, 1237, 610, 2251]. degrees [2565]. Delay [1131, 1589, 1736, 2377, 1331, 2024, 1782, 2390, 1007, 1274, 1309, 1360, 1965, 953]. delayed [2308, 953]. deleting [2542]. delocalization [2053]. delta [432, 357, 1675, 281, 394, 1197, 1955, 1054, 592]. delta-Bose [1054]. delta-potentials [1675]. Dempster [2397]. densities [184, 1181, 2419]. Density [180, 2091, 2318, 1396, 1562, 2312, 1856, 1322, 2448, 15, 179, 1199, 1938, 986, 1857, 397, 2253, 760, 132, 831, 2266]. density-dependent [2318, 760]. density-functional [1857]. dependence [710, 1105, 478, 1425]. dependent [2318, 1221, 309, 868, 1443, 157, 2601, 141, 2122, 2463, 729, 2300, 932, 1005, 2541, 2255, 2653, 2559, 1513, 76, 1796, 2075, 487, 675, 1955, 400, 1319, 2262, 1253, 9, 50, 2393, 2720, 1624, 206, 2466, 2612, 2591, 2431, 1473, 474, 1085, 2686, 1824, 760, 311]. derivation [1218, 976, 689, 1537, 2640, 1676]. Derivations [1196, 198, 12]. derivative [261, 17, 263, 1741, 1465, 1349, 2204, 71, 711]. Derivatives [2284, 495, 582, 2444, 2000]. derive [177]. derived [379, 181, 2096, 1816]. deriving [747]. descendants [835]. describing [1193]. Description [1059, 1375, 905, 90, 2305, 2264, 2417]. descriptions [1450]. Deser [405]. destruction [1146]. destructive [1087]. detailed [1828]. detector [1387]. detectors [1090]. determinant [2607, 2267, 668, 1646]. determinantal [1473, 2482]. determinants [261, 529, 17, 263, 2699, 1034]. Determination [683, 2059]. determine [1941]. determined [2459]. determining [1532, 2376]. deterministic [1965]. detour [272]. Development [587, 345, 2431, 2097]. Developments [34]. deviation [1552, 105]. deviations [889, 714, 2481, 227]. device [2256]. DeWitt [1387]. diagonal [2446, 2138]. Diagonalization [1168, 619, 1785, 694]. diagrammatic [2092]. diagrams [2560, 1262]. Diamond [149, 1003]. diatomic [120]. dichotomic [203]. Dicke [325, 1840]. Diejen [2517]. dielectric [2279, 480, 2639]. dielectrics [1470]. diffeomorphic [2552]. diffeomorphisms [836, 373]. difference [2523, 177, 687]. differences [2027]. different [2406]. differentiable [678, 1824]. Differential [2171, 1495, 184, 1120, 861, 2010, 510, 451, 1639, 1300, 876, 504, 505, 2022, 146, 152, 422, 1967, 935, 2150, 765, 1736, 48, 2564, 2204, 1213, 2523, 1866, 2040, 192, 651, 1460, 153, 496, 219, 1780, 2655, 243, 1968, 1644, 953, 2692, 1116, 1658, 273, 757]. Differential-algebraic [1495]. differential-delay [1736]. differential-difference [2523]. differentials [254, 988]. diffusion [1667, 2042, 515, 230, 1528, 180, 1963, 1595, 2369, 2370, 2627, 1331, 1295, 2243, 2656, 1828, 1498, 489, 2444, 1493, 919, 133, 1829, 1751, 2321, 68, 1965, 1393, 2000]. diffusions [1047]. Diffusive [921]. diffusivity [2300]. dihedral [2566]. dilogarithms [2510]. dilute [2539, 228]. Diluted [1235]. dim [640]. Dimension [193, 1643, 982, 2607, 2384, 2605, 1506, 1486, 1918, 1492, 1373, 2484, 1054, 414, 1150, 1217, 426, 134, 757]. Dimensional [1633, 2596, 2299, 461, 2319, 2522, 627, 1149, 577, 214, 204, 1402, 1165, 2442, 1396, 1625, 325, 2405, 1423, 1528, 277, 1467, 2625, 1011, 1395, 1286, 2525, 2021, 2505, 1013, 2638, 49, 642, 1497, 769, 880, 390, 386, 1678, 1672, 2671, 2094, 1651, 1522, 1193, 1030, 583, 2260, 1891, 403, 623, 486, 335, 1271, 1142, 669, 639, 1005, 1532, 1811, 32, 56, 111, 1684, 1211, 2620, 490, 1868, 1517, 1487, 2665, 336, 597, 959, 2402, 713, 874, 1302, 598, 2050, 1809, 2250, 92, 341, 42, 1171, 449, 1609, 1804, 1734, 1640, 2038, 181, 2554, 1499, 1906, 865, 692]. dimensional [477, 27, 813, 193, 711, 2251, 294, 1166, 1919, 1837, 1132, 2029, 1839, 2668, 792, 869, 1481, 1716, 1353, 972, 2095, 2348, 1378, 1593, 153, 370, 590, 1037, 1527, 939, 283, 1842, 2613, 2623, 1493, 1383, 2077, 2056, 1797, 2520, 572, 2435, 664, 1915, 2054, 2291, 2138, 522, 278, 1558, 2228, 745, 1012, 2455, 1914, 1526, 2218, 581, 475, 377, 2375, 2618, 891, 1658, 2457, 875, 1564, 419, 1190, 468, 1256, 937]. dimensionality [2590]. Dimensions [1678, 2510, 35, 1278, 2101, 1822, 1771, 1954, 518, 789, 1817, 1324, 1633, 709, 52, 2395, 58, 2508, 884, 787, 1537, 288, 1478, 1723, 47, 945, 274, 1387, 2495, 217, 718, 1314, 1315, 973, 21, 2501, 698, 936, 681, 527, 1649, 155, 927, 1764, 923, 276, 700, 2479, 2532, 1654, 2530, 127]. dimer [2353, 1843, 2209]. dipole [446, 20, 775, 734, 2041]. Dipoles [2597]. Dirac [2717, 1524, 2708, 774, 2294, 636, 910, 947, 1752, 1223, 2032, 1530, 151, 1591, 1760, 883, 880, 432, 2526, 304, 2198, 2598, 1556, 1853, 2067, 1004, 319, 740, 624, 1376, 1302, 1075, 1133, 405, 992, 854, 102, 791, 281, 85, 859, 1681, 692, 121, 2472, 144, 759, 584, 2216, 2488, 1325, 2348, 1367, 527, 2649, 603, 1130, 123, 2163, 474, 1816, 308, 1335, 1025, 817, 1636, 1650, 2396, 1379, 2023, 1788]. Dirac- [2294, 2163]. Dirac-attractive [1376]. Direct [1597, 1844, 415, 401, 999]. directed [2484]. direction [1855, 1072, 2645]. directional [711, 410]. Dirichlet [2042, 2207, 2050, 1314, 1315, 2068, 1242, 1256]. dirty [2201]. disclination [446]. Disconjugacy [1923]. discontinuities [1672, 2455]. discontinuity [2588, 1115]. discontinuous [2215]. Discovering [1946]. Discrete [1079, 945, 88, 161, 260, 1570, 456, 861, 2415, 1392, 2043, 880, 673, 406, 2013, 606, 2567, 2011, 1885, 853, 1887, 237, 536, 2315, 1761, 2615, 2565, 1843, 1240, 1790, 2062, 1362, 2413, 2540, 374, 2334, 1464, 766]. Discrete-time [161, 260, 1570, 2415, 1392, 2043]. discretisation [1845]. Discretising [1157]. Discretization [969, 2132, 2433, 2527]. discretized [2277]. discrimination [1549, 2509, 2548, 57, 2197, 117, 435, 316]. Discussion [1658, 1236]. disentangling [498, 295]. Disk [1053]. disorder [1149, 77]. Disordered [982, 2676, 1419, 889, 1766, 1575]. Dispersion [941, 772, 858, 2687, 2189, 1741, 1515, 423, 1507, 82]. dispersionless [2429, 562]. Dispersive [1109, 252, 2609]. displaced [2285]. displacement [1661]. dissipated [2085]. dissipation [340, 131]. Dissipative [1583, 1805, 421, 2134, 1614, 1996, 2693, 1009, 2603, 1937, 2002, 2160, 770, 721, 864, 420, 667, 424]. distance [1285, 719, 2020, 1177, 1453]. Distances [986, 2424]. distillation [2408]. Distilling [463]. distinguishability [2548, 2374]. Distinguished [947]. distinguishes [2081]. distorted [1832]. distortion [1767]. Distributed [2444, 2377, 30, 1499, 506]. Distribution [2681, 1398, 2042, 1396, 719, 2020, 584, 2485, 919, 2000, 1972]. distributions [1297, 1472, 2137, 250, 255, 1469, 1270, 1956, 85, 455, 2583, 136, 817, 1737]. distributive [2024]. disturbed [1381]. divergence [2546, 2129, 1501, 2349, 543]. divergence-free [2349]. divergences [323, 19, 199]. diverse [2306]. DLR [2423]. DNA [1125]. Do [1161, 608]. Dodd [629]. does [1387, 1220]. domain [81, 1117, 1942, 1521, 1080, 2210]. Domains [2134, 1230, 1834, 2456, 2050, 1513, 1429, 2574, 226, 963, 2425, 807, 870, 1965, 1793, 1983, 1564]. Donaldson [1476]. Donder [1140, 1061]. Dong [346]. Dorfman [2290]. Döring [922]. dot [1855]. dots [2185]. Double [2633, 1292, 2302, 1877, 2474, 1894, 2162, 1729, 2488, 480, 1045, 1203, 2146]. Double-bosonization [2633]. double-time [1877]. double-well [2302, 1894, 2162]. doubles [1346]. dressing [403]. drift [230, 2369, 2370, 1473, 371]. drift-diffusion [230]. Drifting [2318]. Drinfeld [1139, 2223, 486, 274, 2538, 2087, 2222, 2173, 2389]. driven [461, 2247, 2709, 978, 2160, 1792, 1330, 958]. drop [1692]. drums [857, 205, 1594]. dual [342, 1372, 928, 437, 908, 2411, 860, 1699, 1909, 1668, 2281, 63, 83, 1897, 1377]. dual-1 [1699]. dual-phase [83]. dualities [1882]. Duality [2517, 783, 2521, 215, 573, 293, 835, 1321, 1909, 720, 871, 459, 936, 625, 1791, 2675]. duality-invariants [835]. duals [494]. due [2647, 177]. Duffin [2021, 157, 9, 468]. Duffing [2694]. Dugdale [813]. Dullin [1504, 2002, 82, 2477]. Dunne [2193]. Dvoretzky [43]. dwell [1628]. dyadic [1673]. Dyck [888]. Dym [2178, 1273]. Dynamic [1386, 1169, 7, 726, 1250, 275, 1607, 754]. Dynamical [461, 1031, 18, 1083, 2519, 350, 2649, 985, 1221, 1386, 29, 2336, 1870, 1551, 865, 2709, 2001, 1191, 347, 2499, 170, 1330, 128, 1886]. Dynamically [1063, 1374]. Dynamics [1011, 750, 967, 1333, 2368, 1689, 2278, 479, 1147, 224, 917, 482, 834, 671, 2189, 1623, 2600, 784, 1904, 1996, 372, 2308, 44, 1541, 706, 2346, 1901, 2311, 1896, 1812, 75, 2083, 751, 2529, 373, 39, 349, 324, 436, 714, 1028, 1134, 1952, 1716, 1461, 1962, 1516, 1763, 168, 187, 1107, 1671, 1907, 410, 2686, 2467, 371]. Dziobek [1334].

E7 [545]. Eardley [1065]. Eckart [2032]. Eden [2696]. Edge [574, 1778, 2354, 1257, 1973, 132]. Editorial [1535]. Effect [1376, 284, 1090, 1805, 630, 380, 44, 1761, 36, 531, 2399, 2262, 296, 1459, 1915, 2054, 1439, 785]. Effective [943, 1223, 1258, 1835, 1445, 1677, 231, 232, 1621, 2541, 1311, 400, 2335, 206, 2686]. Effective-mass [943, 1223]. Effectiveness [40]. Effects [1855, 2641, 1692, 574, 1342, 749, 2199, 870, 1405, 953]. efficiency [1176, 1518]. efficiently [2589]. Efimov [1761]. eigenfunction [1704, 1542, 838]. Eigenfunctions [979, 2463, 264, 960, 1130]. Eigenkets [404]. eigensolutions [2021]. eigenstates [2676]. Eigenvalue [2443, 895, 2147, 222, 1469, 902, 2664, 100, 1235, 1303, 961, 132, 462]. Eigenvalues [2657, 2628, 725, 2642, 91, 937, 772, 222, 2071, 1537, 2441, 858, 973, 1007, 1274, 1440, 2334, 395, 1242, 1793, 1983]. eigenvectors [2208]. eight [2501]. eighth [2490]. Einstein [342, 1772, 126, 2333, 1561, 433, 1565, 473, 2411, 275, 2369, 2370, 740, 218, 321, 741, 483, 1608, 1796, 914, 2410, 1462, 1763, 187, 680, 2641, 2372, 123, 1922]. Eisenhart [2202]. Elastic [1773, 1561, 2450, 506]. elasticity [1775]. elastodynamics [2376]. electric [114, 824, 446, 20, 775, 734, 1670, 1784, 1774, 785, 1705]. electrodynamics [917, 1192, 995, 635, 1067, 1106, 653, 308, 636, 1705, 1361, 1122, 548]. electrohydrodynamics [424]. Electromagnetic [2279, 733, 480, 2430, 1915, 2019, 1261, 1269, 1614, 2240, 2497, 956, 1683, 1199, 2170, 1304, 2253, 2060, 1605, 1784, 1975, 1774, 2003, 650]. electromagnetism [1460]. electron [878, 662, 267, 303, 2394, 1858, 1304, 2639, 553, 1384, 892]. electron-hole [878]. electrons [2420, 1053, 2179, 1776, 1343]. electrostatic [1147]. electroweak [1366]. element [107, 2589, 1114]. elementary [1260, 2416, 1115]. elements [1707, 518, 789, 1817, 1647, 2154, 2394, 2675]. elimination [593, 1341]. Elliott [1410]. Elliptic [1789, 2666, 2510, 2318, 1599, 129, 1501, 2236, 2662, 2098, 1045, 1153, 1194, 2624, 886]. Embedded [725, 1571, 937, 4, 2295]. Embedding [253, 1742, 976, 2705, 2706, 167, 602, 1571, 742]. Embeddings [622, 422, 1003, 2224]. Emden [1539, 1289]. Emden-type [1539]. Emergence [2307, 2072]. Emergent [2310, 2550]. emerging [2161]. endowed [206]. energies [2027]. Energy [579, 2009, 2463, 2024, 1460, 2651, 980, 1626, 2226, 2032, 1589, 1423, 1948, 421, 1463, 2279, 59, 1073, 1715, 1424, 2615, 458, 355, 228, 1143, 1520, 1934, 761, 2484, 1716, 614, 1827, 67, 1233, 1418, 538, 1307, 616, 977, 372]. energy-momentum [1463, 2279, 614, 67, 616]. Engineered [2008]. enhanced [483]. Ensemble [2404, 574, 91, 1980, 428, 1433, 2091, 961, 1973]. ensembles [507, 1431, 1598, 1469, 889, 348, 41, 384, 2091, 291, 1151, 1571, 2419, 132]. Enskog [1479, 142]. entangled [1059, 1452, 1375, 750, 1487, 516, 1450, 408, 1257]. Entanglement [1545, 162, 2168, 2590, 946, 1420, 2676, 834, 2408, 463, 1659, 318, 787, 25, 1451, 307, 1679, 42, 1766, 1058, 983, 409]. entire [241]. Entropic [2049, 1094, 1312, 2237, 351]. entropies [1471, 1659, 2130, 641]. Entropy [2715, 1813, 2716, 788, 1308, 1453, 1098, 43, 1021, 1917, 2420, 2405, 2273, 676, 2571, 1178, 1714, 691, 839, 2128, 787, 571, 427, 1861, 2237, 1548, 2468, 1022, 161, 833, 2406]. environment [2484, 2161]. epsilon [2284]. epsilon-expansions [2284]. Epstein [2596]. equal [2553, 366, 1334]. equality [2035]. Equation [2379, 951, 1643, 1745, 596, 2418, 830, 2328, 725, 383, 1223, 1880, 2241, 856, 941, 2032, 515, 1530, 1562, 1409, 1222, 1931, 1990, 1621, 151, 579, 1589, 1483, 1372, 2219, 1011, 2010, 249, 852, 1250, 46, 1286, 433, 402, 1477, 1639, 1760, 1208, 2108, 981, 1497, 90, 772, 2719, 169, 954, 805, 1269, 2223, 431, 332, 1241, 1964, 504, 505, 1228, 1543, 1666, 255, 921, 152, 2380, 2242, 1595, 2598, 2301, 1030, 2500, 1556, 583, 2225, 1894, 1840, 403, 1743, 372, 20, 775, 445, 225, 1382, 251, 669, 588, 1710, 902, 1084, 2710, 1511, 439, 1646, 1930, 2490, 1715]. equation [218, 2541, 2620, 266, 730, 1089, 1406, 2708, 1290, 237, 1736, 624, 1349, 2693, 2452, 48, 1145, 774, 2627, 1133, 764, 1665, 1506, 2314, 313, 139, 1849, 2454, 392, 1749, 2559, 1809, 312, 992, 1331, 1652, 1310, 379, 2667, 2718, 102, 791, 2391, 2493, 265, 449, 731, 1938, 1207, 2068, 1313, 116, 1608, 1046, 436, 858, 1404, 2524, 1499, 1787, 71, 1213, 2681, 487, 675, 1986, 96, 399, 1955, 400, 2523, 21, 171, 477, 299, 710, 1373, 2317, 2247, 2656, 897, 121, 2472, 220, 1169, 1782, 2062, 1028, 2694, 1629, 1713, 567, 93, 974, 1289, 1937, 1081, 1799, 1215, 2498, 50, 53, 1858]. equation [1317, 934, 2002, 131, 1516, 2707, 340, 2186, 1753, 527, 370, 773, 629, 187, 2177, 206, 2018, 674, 2206, 2466, 1521, 489, 207, 8, 2326, 1851, 1597, 2099, 252, 919, 925, 133, 603, 1194, 890, 687, 2221, 367, 2227, 1418, 1836, 2325, 1279, 292, 2351, 1908, 2711, 429, 1048, 1020, 1012, 1121, 1579, 1645, 213, 1479, 551, 305, 142, 2378, 1711, 1526, 2000, 138, 301, 438, 1306, 1987, 1650, 1740, 13, 200, 227, 2003, 803, 1941, 247, 82, 1789, 977, 420, 767, 1111, 1204, 1254, 1886, 1654, 2530, 2577, 2324, 2477, 667, 1114, 1379, 2023, 592]. Equations [2299, 1640, 2666, 238, 569, 863, 2318, 1278, 2101, 2670, 1043, 861, 2420, 126, 1940, 703, 1528, 1590, 7, 726, 1395, 1248, 1561, 1637, 542, 1565, 473, 185, 1051, 90, 2689, 2089, 2531, 1852, 1685, 141, 386, 876, 1356, 2136, 1877, 2069, 2277, 1529, 2252, 2022, 481, 676, 146, 485, 2526, 1798, 1371, 1522, 1193, 1617, 2296, 573, 1939, 326, 1714, 2188, 2239, 1891, 1967, 1750, 225, 2497, 1112, 2575, 2238, 2558, 1532, 1435, 1236, 148, 298, 1465, 1517, 321, 2255, 495, 223, 1157, 2315, 336, 94, 844, 744, 1407, 938, 969, 1698, 660, 2627, 1675, 440, 1172, 1982]. equations [1276, 2514, 1501, 1942, 679, 1668, 1468, 1534, 1199, 1332, 967, 76, 1160, 1735, 838, 2236, 658, 871, 226, 1188, 1932, 2616, 1485, 717, 124, 16, 511, 2149, 1173, 2281, 1712, 2298, 188, 194, 1240, 2140, 2573, 2105, 2098, 532, 1291, 294, 1101, 2058, 2709, 192, 1369, 1607, 754, 1277, 931, 2703, 1616, 872, 1538, 681, 972, 1695, 1367, 822, 748, 2233, 443, 2178, 153, 496, 2253, 2478, 320, 2048, 1360, 2096, 1527, 384, 2376, 1842, 2613, 2623, 1663, 2658, 219, 489, 1238, 1670, 155, 927, 1493, 1383, 1709, 705, 113, 2044, 1848, 2578, 388, 509, 143, 1797, 1330, 1519, 2655]. equations [2217, 812, 864, 1127, 1273, 243, 1303, 998, 704, 1968, 1494, 745, 87, 68, 1080, 1400, 1316, 1965, 1664, 2609, 1307, 1505, 2576, 2533, 1202, 1393, 1170, 1318, 377, 1889, 2692, 616, 1943, 1507, 454, 891, 899, 361, 2479, 2532, 1116, 578, 190, 766, 760, 801, 1658, 1618, 1661, 2624, 875, 752, 533, 273, 650, 2191, 1984, 1739, 918, 1933, 2662, 1262]. equatorial [989]. equidistant [1625]. equilibria [1082, 1639, 996, 2553, 615, 1869]. equilibrium [1047, 1710, 646, 1063, 1374, 841, 1032]. Equivalence [1452, 810, 807, 302, 1668, 1995, 1754, 2073]. Equivalences [2174]. equivalent [2632, 2410]. Equivariant [422, 2496, 1803, 320]. equivariantizations [1653]. Erdahl [1322]. ergodic [2653, 116, 887]. Ergodicity [178]. Ermakov [1816]. Ermakov-type [1816]. Erratum [863, 2622, 2717, 726, 2021, 2713, 2531, 1375, 2712, 2721, 263, 1252, 775, 1612, 2183, 2020, 2295, 2621, 1932, 2626, 1373, 1129, 1934, 397, 110, 2623, 2326, 1128, 509, 2294, 2325, 636, 2182, 199, 1374, 200, 776, 899, 2619, 1984, 2707]. error [2509, 2588, 2402, 57, 435, 316, 410]. ESK [2041]. Essential [1229, 2445, 2104]. estimate [765, 1721]. Estimates [1793, 1983, 73, 1109, 2697, 1426, 2211, 799, 2448, 485, 1414, 2300, 1921, 1439, 964, 2576, 297]. estimation [2355]. estimators [1737]. Euclidean [1983, 2282, 622, 2329, 2050, 585, 1019, 1793]. Euler [2532, 2381, 249, 1250, 1672, 2704, 1714, 225, 1171, 1534, 2668, 2216, 519, 2217, 891, 2479]. evaluate [1748]. evaluating [1034]. evaluation [1912, 2073, 2144, 2082]. even [2238, 1723, 1640]. even-dimensional [1640]. Every [380, 2593, 2622]. Evolution [154, 461, 736, 386, 431, 548, 635, 148, 2527, 679, 2702, 1173, 294, 2465, 946, 1360, 2714, 45, 693, 2356, 1740, 190]. evolutionary [1301, 681]. evolutions [563, 161]. evolved [350]. Evolving [2433]. Exact [596, 2353, 782, 786, 2449, 2320, 406, 1053, 33, 1768, 1478, 2627, 1162, 734, 2667, 731, 1608, 1796, 64, 694, 69, 61, 568, 1325, 770, 919, 2003, 1259, 1218, 712, 180, 505, 1890, 1743, 2082, 2493, 1860, 2639, 134, 1507]. Exactly [49, 2334, 1218, 542, 403, 158, 691, 400, 374, 2077, 346]. Exactness [575]. Examples [160, 2306, 2461]. Exceptional [412, 2434, 158, 1923, 1758, 2612]. exchange [933, 2639]. Excitation [673]. excited [2694]. exciton [1809]. exclusion [279, 665, 2047]. Exclusivity [1898]. excursion [960]. Existence [703, 1419, 1286, 612, 141, 849, 2136, 1963, 908, 1853, 1245, 439, 2299, 1407, 2100, 1172, 638, 1332, 1057, 1067, 2298, 2107, 2694, 2668, 370, 1360, 2585, 123, 2692, 977, 2210, 1589, 971, 437, 1024, 1714, 203, 425, 1009, 1474, 1468, 2647, 2573, 1616, 2248, 1309, 362, 994, 1128, 509, 143, 426, 1121, 1393, 2375, 667, 1618, 1661, 2220]. exotic [1872]. expanding [769, 1680]. Expansion [1715, 309, 1497, 2685, 1655, 1073, 639, 319, 1900, 2691, 1832, 675, 2634, 331, 414, 1226, 330, 328, 1187, 2109, 2419, 716]. expansions [2284, 2111, 838, 2513, 1130, 311]. expectation [1246, 1224]. expected [2696]. experimentally [1370]. Explicit [1441, 15, 2306, 1435, 397, 1345, 2324, 1354, 672, 787, 1905]. Exploding [1395]. Exploitation [676, 1308]. exponent [1617, 2237]. Exponential [2651, 905, 1403, 1735, 303, 316, 506, 953, 255, 2346, 500, 92, 281, 2034, 584, 295, 2378]. Exponentially [1018, 2721, 2558]. exponentials [432, 1862, 2071]. exponentiation [1557]. exponents [2509, 2006, 2713, 842, 1234, 1190]. expression [787]. Extendability [2450]. Extended [1927, 1691, 2005, 1324, 2293, 1675, 1486, 2084, 981, 2599, 2065, 2710, 571, 1756, 1923, 1514, 2205, 194, 1028, 337, 2542, 2460, 2491, 773, 1908, 2711]. Extending [1455, 2117]. extensibility [2645]. Extension [2347, 817, 66, 547, 1856, 1112, 836, 217, 1548, 1292, 2631, 2029, 1336, 2118, 138, 496]. Extensions [873, 2386, 2601, 2559, 2035, 947, 1752, 2717, 1585, 2014, 2064, 2579, 353, 809, 1001, 1055, 2332, 2033, 1304]. exterior [2456]. External [1525, 824, 879, 332, 548, 635, 2598, 1263, 575, 1792, 2206, 1479]. extra [1633]. Extremal [2458, 907, 216, 1601, 398, 822]. Extremality [1547]. extremals [1414]. Extreme [1050, 91, 1184].

F [425]. F-region [425]. face [619]. faces [209]. factor [1053]. factorizable [883]. Factorization [1272, 1162]. factorizations [317]. factorized [784]. factors [2551, 1600, 920, 1846]. Faddeev [872, 948, 2203, 1137]. fall [460]. families [389, 1723, 1974, 1758]. family [655, 1189, 820, 1463, 2207, 2604, 2355, 352, 1681, 63, 1719, 1377, 1851, 156, 961, 1170]. far [1891, 177]. far-field [177]. Fast [2652, 269, 1966, 593, 2694, 1751]. Fate [283, 40, 1458]. features [2235, 902]. featuring [1392, 2043]. Fedosov [643]. feedback [593, 1782]. feedbacks [1589, 506]. Fefferman [1516]. Fefferman-Stein [1516]. Fériet [644, 860, 2284]. Fermat [1586]. Fermi [1524, 1569, 1715, 64, 1470, 870]. fermion [2110, 2537, 2403, 1422, 1508]. Fermionic [431, 1574, 1740, 2407, 2113, 2305, 2274, 270, 99, 375]. fermionization [698]. fermions [1927, 1702, 2189, 304, 1084, 1326, 694]. ferromagnetic [1233]. ferromagnets [2654, 957]. few [948]. Feynman [2179]. fiber [433]. fibers [191]. Fibonacci [1284]. Field [654, 1262, 114, 824, 1696, 2353, 1667, 1120, 446, 231, 232, 2189, 1467, 990, 1948, 1013, 1561, 612, 1426, 2538, 460, 2449, 699, 1269, 2571, 267, 548, 635, 1690, 2240, 2497, 122, 2274, 1005, 32, 1148, 1572, 1117, 1686, 1263, 1809, 1769, 1729, 469, 1683, 1310, 743, 604, 368, 31, 1855, 658, 1124, 920, 1843, 124, 2473, 1102, 369, 1134, 1417, 1016, 657, 176, 354, 987, 62, 177, 1461, 1323, 1325, 2160, 614, 168, 1755, 1304, 1719, 2048, 1272, 1649, 1343, 356, 115, 783, 994, 1128, 450, 1459, 1682, 1203, 2146, 1335, 2459, 2218, 2003]. field [470]. field-electron [1304]. field-particle [2497]. fields [2521, 224, 1261, 119, 697, 90, 471, 1018, 2721, 1657, 2364, 1998, 1891, 1040, 1412, 353, 2344, 2555, 2487, 2621, 221, 355, 1680, 349, 560, 1134, 1369, 178, 1389, 1328, 1225, 565, 1200, 147, 2061, 1397, 474, 2004, 10, 262, 2041]. Fierz [827, 2366]. fifth [1213, 2703, 1111]. fifth-order [1111]. filament [502]. Filiform [2257]. Filippov [634, 686]. filling [1053]. film [1652]. filter [1720]. fine [360, 401, 735]. Finite [2442, 1531, 386, 1166, 2054, 1914, 419, 1190, 35, 241, 1549, 2588, 289, 402, 1678, 1976, 52, 1053, 1023, 2232, 853, 1887, 597, 427, 1066, 2168, 400, 869, 423, 1459, 649, 1536, 745, 1824, 1488, 1114, 2442]. Finite-dimensional [1166, 419, 1190, 1678, 869]. Finite-gap [386]. finite-temperature [1023]. finitely [1773, 2334]. finiteness [1761]. Finsler [968, 74, 1586, 1922]. First [2167, 2147, 2410, 2609, 1149, 1677, 768, 689, 2083, 1698, 234, 1292, 70, 1156, 651, 919, 998, 128]. First-order [2410]. first-passage [234]. Fischer [2635]. Fisher [1297, 902, 2606, 1028, 1578, 1232, 2085, 1458]. Five [581, 1402, 789, 32, 193, 1132, 276, 700]. Five-dimensional [581, 1402, 1132]. Fixed [132, 1834, 491, 2559, 779, 1520, 1934, 1072, 1725, 1458, 744]. fixed/moving [2559]. FK [2423]. Flag [2041, 74]. Flag-dipole [2041]. Flat [97, 1808, 2409, 1874, 1180, 2359, 522]. flavor [1138]. flights [1896, 1478]. flip [1910]. flocking [482]. Floquet [2431]. flow [2522, 232, 1327, 1269, 1890, 1209, 2275, 959, 991, 2156, 1210, 711, 1693, 2481, 1418, 522, 2088, 1318, 1232]. Flows [2379, 1441, 671, 815, 1672, 1651, 1657, 1312, 288, 1577, 2604, 800, 1341, 1520, 1934, 2650, 1152, 2417, 154, 2102, 2455]. Fluctuation [1426]. Fluctuations [1384, 1745, 2371, 524, 100, 1235]. Fluid [7, 726, 1657, 1212, 1010, 2605, 1523, 1652, 2511, 1533, 1313, 991, 1961, 1516, 154, 67, 1985, 522, 1564]. Fluid-dynamic [7, 726]. fluid-particle [1212, 1985]. fluid-structure [1533]. Fluids [1509, 1193, 1496, 2456, 28, 1213, 800, 502, 1797, 1111]. flux [1018, 2721, 1220]. fluxes [2594]. foam [1142, 874]. Fock [54, 2537, 2158, 2685, 2689, 2558, 1430, 2148, 1196, 1384]. focusing [2219, 1011, 2500, 1494]. Fokas [1372, 71]. Fokker [2709, 1710, 1749, 2559, 76, 897, 1493, 919, 1330]. fold [71, 112, 1992, 2195, 1400]. foliations [1185, 1874, 1636]. forbidden [837]. force [332, 956, 177, 1479, 616]. forced [1466, 1964, 2134, 1991, 328]. forces [2046, 978]. Forchheimer [2604]. forcing [1965]. Forest [2596]. Form [1846, 834, 2403, 1532, 1901, 1052, 1501, 1199, 920, 1460, 1770, 2253, 551, 420, 1929]. Formal [1344, 537, 70, 1200]. formalism [1260, 2569, 1445, 2557, 1683, 85, 605, 1094, 1191, 413, 374]. formation [1746]. former [1905]. forms [65, 888, 769, 1452, 1239, 737, 1832, 2347, 248, 687]. Formula [2596, 951, 591, 530, 2640, 1349, 2699, 2365, 2672, 1091, 1641, 581]. formulae [1152, 140, 776]. Formulas [279, 1370, 1862, 1768, 911, 1435, 86, 2425, 1345, 233]. formulation [211, 699, 167, 259, 2710, 751, 324, 2631, 1757, 1905, 2297, 1908, 2711, 273]. formulations [83]. Foucault [365]. Foulis [1074]. foundation [2550]. Four [1974, 2331, 444, 2319, 214, 789, 390, 1059, 1452, 1375, 486, 1142, 1211, 516, 1761, 1387, 1491, 2168, 718, 1314, 1315, 972, 625, 2586, 2119]. four-dimensional [214, 390, 1142, 1211, 972]. four-identical-particle [2119]. four-level [2586]. four-manifolds [1491]. four-qubit [516]. Fourier [2167, 1532, 1118, 1981, 500, 88]. fourth [249, 1250, 1930, 21, 1373, 2703, 1864, 1851, 1121]. fourth-order [249, 1250, 1864, 1121]. Fox [1621]. Fractal [957, 2194, 2435]. Fractional [815, 1595, 1197, 1629, 1713, 1123, 137, 535, 1409, 1222, 1931, 1990, 1621, 1444, 1852, 1241, 729, 2225, 345, 1236, 2150, 1407, 2693, 1268, 598, 1665, 2702, 1199, 265, 731, 2204, 76, 2244, 796, 1787, 399, 1955, 2656, 1573, 370, 2253, 2206, 489, 1709, 2099, 2714, 1780, 2163, 2294, 1783, 2426, 851, 2533, 767, 1929, 2464, 2700, 592]. fractional-order [1268]. Frame [62, 834, 359]. frames [973, 1182]. framework [224, 2611, 1677, 2659, 2352, 2274, 679, 1438].

Fraunhofer [1118]. Fréchet [2485]. Fredholm [1611, 1434, 1494]. Free [2427, 460, 1584, 1124, 2545, 1696, 782, 2688, 1600, 868, 1443, 1448, 1496, 122, 850, 1703, 1468, 1533, 1609, 1961, 1287, 2484, 2349, 2258, 1827, 1229, 870, 1907]. Free-fall [460]. freedom [2602, 2565, 207]. freeness [1800]. Frenkel [1146]. frequencies [683, 201]. frequency [1246, 975, 2653, 675, 2414]. friction [1466, 1502, 930, 660]. Friedlander [2699]. Friedman [361]. Friedmann [1562, 275, 1680]. Frobenius [1346, 2661]. fronts [1610]. frustration [1448]. frustration-free [1448]. Fuchs [2238]. Fuchsian [1346]. fuel [2256]. Full [1555, 2507, 1765, 90, 1515]. Fully [2366, 1680, 1507]. Function [1858, 910, 644, 73, 250, 1543, 1666, 2475, 2610, 2518, 2567, 1385, 554, 357, 729, 251, 588, 639, 1035, 1662, 1491, 281, 194, 2062, 88, 584, 1325, 331, 2639, 2649, 843, 1167, 572, 1303, 532]. Functional [261, 17, 263, 495, 1763, 2607, 2286, 786, 1414, 839, 2710, 1868, 1959, 1857, 617, 1780, 1908, 2711, 438, 2692, 918, 1933]. functionals [1968, 2088]. functions [2666, 1594, 241, 840, 861, 1912, 2442, 1621, 1472, 2144, 829, 722, 894, 896, 157, 2307, 2678, 376, 1018, 2721, 2437, 2636, 1073, 58, 403, 1823, 1810, 129, 2344, 1885, 2284, 1974, 2568, 1587, 36, 1861, 300, 500, 92, 341, 2564, 2486, 838, 303, 2236, 2662, 2401, 244, 394, 1380, 1949, 9, 354, 1858, 651, 2072, 1045, 2663, 2115, 2206, 233, 1153, 538, 19, 199, 2624, 917]. functor [54]. Fundamental [2656, 2241, 2282, 1299, 2304, 2362, 1160, 62, 450]. Furry [2305]. Fusion [1653, 2286, 1530]. Future [1462]. fuzziness [2443]. fuzzy [1004, 301, 438].

Gacs [2405]. gain [1792]. Gaiotto [1612, 1198, 581]. galactic [1231, 1333]. Galilean [217, 950, 394, 165]. Galilei [2012, 1925, 1000, 932, 2264, 1336, 2519]. Gallouet [976]. Game [2469]. Game-theoretic [2469]. games [983]. Gamow [1447]. gap [2651, 386, 1084, 313, 2304, 2362, 38, 1367, 2124]. gaps [2574]. Gardner [379, 1485]. Garnier [847]. gas [2312, 751, 959, 1609, 228, 331, 1054, 870, 1997, 2591, 2423, 2479, 2532]. gasdynamics [939]. gaseous [891]. gases [2084, 1691, 2005, 1830, 2372, 572]. gasket [663]. gaslike [524]. Gaudin [997, 1840, 886]. Gauge [75, 728, 2470, 2658, 272, 1195, 1863, 2235, 912, 1355, 1632, 1310, 1819, 1820, 1140, 2690, 1141, 781, 2203, 2262, 1949, 1061, 1559, 2095, 320, 356, 994, 1128, 2131, 2184, 278, 1227, 484, 2289, 1014]. gauges [504, 948]. Gauss [741, 1660, 1657, 1118, 500]. Gaussian [2709, 574, 1297, 91, 2685, 428, 2052, 737, 2667, 1553, 960, 2583, 2091, 1330, 1571, 2419]. Gaussianity [178]. Gazeau [2608, 1924]. Gel’fand [453, 2290]. Gell [1945]. Gell-Mann [1945]. General [2285, 2621, 1782, 497, 1277, 340, 2078, 2677, 10, 262, 1098, 564, 1431, 549, 1639, 448, 1100, 1821, 2377, 51, 477, 710, 567, 1559, 568, 435, 1152, 870, 794, 2287, 1058, 890, 87, 823, 724, 301, 438, 2645, 951, 1866, 1397]. Generalised [1434, 1611, 2202, 1189, 1810, 1355, 1579]. Generalization [2158, 2205, 1088, 893, 136, 2684, 2094, 666, 1078, 692, 96, 1081, 2130, 1045, 330, 1012]. Generalizations [2249, 856, 2057, 92, 341, 2673, 2406, 2031].

Generalized [2383, 2380, 2567, 1657, 2541, 904, 1093, 871, 1674, 459, 1132, 732, 1627, 1378, 2292, 1194, 1907, 19, 199, 2338, 1958, 1259, 2302, 1043, 2611, 1966, 2042, 515, 118, 1591, 1297, 1472, 2137, 995, 2648, 2525, 1181, 805, 1504, 1976, 1706, 222, 481, 2207, 2296, 2508, 150, 2238, 911, 1435, 2434, 702, 336, 735, 1133, 2314, 2454, 2604, 1276, 791, 1547, 2197, 707, 838, 1608, 1619, 1551, 2523, 2140, 2200, 1876, 191, 974, 2399, 2359, 2046, 792, 1826, 2170, 822, 2178, 527, 2663, 1792, 1377, 2206, 233, 388, 1146, 1308, 998, 704, 2351, 2426, 305, 1615, 285, 767, 1111, 578]. generalized [762, 667, 333]. generally [2040]. generate [1185]. generated [984, 2427, 2133, 386, 1412, 2392, 1571]. Generating [1472, 769, 1968, 831, 2144, 850, 1703, 804, 272, 504]. generation [65, 25]. generators [932, 2412, 2124]. generic [839, 1952]. Geng [2317]. gently [429]. geodesic [1484]. geodesics [1577]. Geometric [2271, 876, 1812, 364, 1017, 1450, 585, 2127, 797, 2102, 1381, 2261, 815, 528, 1859, 1356, 391, 1651, 307, 2168, 1916, 2121, 1340, 188, 1786, 2594, 1716, 130, 1573, 2365, 484, 2455, 1588, 2436, 273]. geometrical [208, 2376, 327]. geometrically [81]. geometries [2306, 1234]. geometrodynamics [6, 2183]. Geometrothermodynamics [1747, 2440, 2249]. Geometry [1372, 373, 2190, 756, 761, 2343, 1120, 365, 66, 1582, 1164, 909, 2171, 2135, 1053, 2360, 2288, 2556, 417, 2643, 1729, 2608, 1679, 2690, 1097, 1203]. geostrophic [1512, 918, 1933]. Gerdjikov [1916, 1306, 915, 1204]. ghost [246]. Gibbon [629]. Gibbs [2438, 526]. Gil’ [241]. Ginibre [2356]. Ginzburg [141, 1964, 2255, 1608, 2007, 1216]. Girardello [2584]. given [2142]. giving [2461]. Glaser [2596]. glass [1298]. glasses [2698, 1910, 1572]. Global [1589, 1714, 1750, 2497, 1212, 1015, 94, 959, 1010, 2605, 1009, 1474, 1609, 859, 1911, 1028, 1101, 1027, 1616, 1461, 661, 1309, 426, 721, 864, 1121, 1216, 2375, 2630, 578, 667, 760, 1661, 424, 127, 2208, 473, 2100, 2693, 2106, 418, 509, 143, 985, 2103, 589]. globally [2528, 2097]. gluing [2699]. golden [179, 339, 2212]. Goldstone [604]. good [1950]. Gordan [672, 1205, 753, 1699, 453]. Gordon [943, 2558, 113, 596, 118, 1011, 1030, 445, 1382, 2575, 135, 2246, 1506, 523, 1173, 96, 974, 1045, 292, 551, 1987, 1464, 1654, 2530, 667]. Goryachev [1065]. Gottwald [1504, 2002, 82, 2477]. Gould [954, 1924]. Goursat [1249]. governed [2225]. governing [676, 439, 1238]. Graded [1945, 2156, 2213, 79, 2564, 924, 1485, 1593]. gradient [1426, 1312, 1365, 2481, 2578, 2085]. gradient-flow [2481]. gradients [676]. grading [735, 933]. Gradings [248, 401]. gradually [708]. grained [1857]. graining [554]. Gram [529]. granular [525, 7, 726]. Graph [647, 457, 1818, 2510, 1898, 1759, 2379, 2248, 1720]. Graphene [654, 2420, 2597, 304, 1430, 1417, 1343]. graphene-like [1343]. graphs [1415, 2687, 2462, 119, 2333, 2449, 901, 1454, 1385, 541, 526, 853, 1887, 1632, 2304, 2362, 1734, 1416, 2606, 1596, 100, 1235, 893, 1873]. Grassmann [832, 1556, 99, 192]. Grassmannian [492]. Grassmannians [2442, 1781]. gravitating [2511, 324, 1906]. gravitation [1120, 1866, 2040].

Gravitational [605, 558, 655, 460, 2214, 1962, 1325]. gravities [2041]. Graviton [1559, 1140, 1061]. Gravity [836, 815, 2409, 1467, 2171, 1228, 2037, 167, 1863, 2235, 795, 2100, 741, 1390, 1099, 166, 1770, 1353, 2234, 607, 608, 1864, 125, 484]. Gray [2289]. Gray-categories [2289]. Green [917, 2475, 481, 729, 588, 933, 651, 1325, 233, 297]. Green-functions [917]. Greybody [2551]. Gröbli [173]. Gröbner [1023]. Groenewold [1899]. Grosche [1650]. Gross [690, 1371, 1715, 1406, 967, 187]. Ground [1948, 1852, 1448, 1430, 42, 2151, 443, 2499, 1709, 1048, 2023, 2651, 1286, 2447, 1715, 1172, 638, 355, 1433, 1067, 228]. Ground-state [1448, 2651, 355]. Group [2101, 267, 1722, 1537, 1676, 679, 1652, 2180, 2357, 873, 231, 232, 1484, 1347, 1590, 1600, 2638, 883, 1000, 2678, 1998, 1205, 2268, 932, 2710, 738, 1577, 1148, 29, 1355, 1050, 1974, 1145, 874, 735, 1978, 2702, 2589, 781, 1210, 865, 2523, 2143, 2196, 1267, 2566, 1687, 2583, 2488, 2634, 2714, 1908, 2711, 1644, 2609, 716, 1041, 2644, 2535, 373]. Group-invariant [2101]. Group-theoretical [267, 1537, 1676, 679]. groupoids [2259]. groups [2580, 1125, 1584, 1000, 1717, 2385, 2567, 2014, 2064, 1981, 2568, 401, 1075, 1160, 936, 1353, 972, 258, 633, 2492, 1438, 1110]. Growth [1480, 2696, 1172, 2098, 2422, 1921, 2099, 1476, 2220]. Gudder [666]. guide [399].

h [855]. Haar [1800, 1110]. Hadamard [315, 1723, 1546, 1229]. Hahn [860, 1699]. Haldane [1396]. half [980, 1626, 2226, 284, 372, 765, 2100]. half-line [765]. half-space [372]. Hall [1053, 36]. Hamilton [1298, 1467, 2557, 2280, 612, 1337, 83, 2446, 2177, 1335, 1733]. Hamiltonian [2319, 1301, 2569, 1966, 1320, 1637, 612, 1176, 2600, 663, 2601, 2599, 2135, 1502, 884, 930, 1982, 2608, 707, 324, 1333, 1029, 172, 366, 610, 2251, 2414, 1837, 2631, 2680, 1152, 2278, 1495, 1200, 2096, 939, 359, 2044, 2102, 1482, 1630, 1791, 2004, 156, 2361, 475, 285, 1076, 767, 1929, 378, 333, 1838]. Hamiltonian-based [1176]. Hamiltonians [204, 2245, 1677, 828, 1448, 2463, 160, 1245, 2346, 1762, 269, 590, 202]. Hankel [1601]. Hans [1535]. Hans-Jürgen [1535]. happens [2201]. Hard [28, 2428, 1778, 134]. Hard-sphere [28]. Hardy [947, 2192, 546]. Hardy-type [546]. Harish [2457]. Harish-Chandra [2457]. Harmonic [984, 440, 462, 1958, 952, 577, 2572, 1113, 782, 2312, 1109, 2570, 185, 1991, 251, 1005, 1435, 1282, 2059, 2487, 312, 1253, 2029, 2216, 2720, 2398, 1037, 1085, 1920, 1379, 1055]. Harry [2178, 1273]. Hartman [1805]. Hartree [2158, 1208, 2108, 2689, 784, 1430, 1433, 93, 2533]. Hartree-type [93]. Hastings [162]. Hausdorff [756, 127]. having [1059, 1452, 1375, 366]. Heat [2607, 928, 2074, 2359, 255, 1655, 1480, 1066, 646, 991, 83, 2481, 962, 1872, 2480]. heat-conduction [2481]. heat-trace [1872]. heating [958]. heavenly [342]. Hecke [2286]. Height [1976]. Heiles [69]. Heiles-type [69]. Heisenberg [2057, 2614, 79, 2483, 101, 2350, 1689, 1892, 1622]. helicities [1657]. helicity [560, 1225, 2417, 154, 2283]. Helicoidal [885]. Helmholtz [588, 1827].

Hénon [69]. Hepp [2006, 2713]. Herglotz [1868]. Hermite [2263, 2164, 1758, 1648, 965, 132]. Hermitean [493]. Hermitian [507, 204, 2175, 1092, 828, 160, 2187, 1622, 499, 2146, 311]. heterointeractions [113]. heterostructures [2335]. Heun [775, 1951, 896, 20, 48]. Heyting [2182, 1953]. HH3 [145]. Hidden [1243, 1790, 1135, 1251, 1019, 1378]. hierarchies [815, 808, 690, 816, 2322, 2494, 2008, 16, 511, 2087, 974, 882]. hierarchy [1631, 2320, 246, 1706, 1704, 2548, 2094, 2429, 237, 1982, 1916, 2514, 379, 1606, 1046, 71, 562, 194, 1126, 1790, 2062, 2657, 1341, 337, 2631, 2152, 1495, 1921, 705, 2661, 1791, 2672, 1783, 2426, 286, 1588, 1503, 847, 2361, 377, 767, 1929, 2436, 378, 190, 752, 333, 1838, 1845]. Higgs [697, 320, 1014, 2103]. High [1626, 1439, 1307, 2147, 878, 1583, 2352, 355, 1915, 2054]. high- [878]. High-energy [1626]. high-loss [1583, 2352]. High-velocity [1439]. Higher [1966, 2689, 2013, 1651, 1684, 1180, 2693, 576, 927, 790, 65, 261, 1880, 1677, 2388, 325, 1657, 17, 263, 326, 1214, 2503, 1646, 288, 1065, 778, 973, 21, 1373, 191, 329, 698, 1881, 1846, 2095, 1191, 174, 359, 283, 1658]. higher-dimensional [1658]. Higher-order [2689, 2013, 326, 1214, 174]. highest [411, 732]. Hilbert [2522, 2357, 2405, 2193, 845, 1902, 167, 1168, 1349, 802, 2433, 1346, 349, 436, 2410, 906, 1928, 2323, 1648, 2213, 1447]. Hill [1745, 313, 913, 1279]. Hilliard [1750, 1169]. Hirota [2320, 2188, 1157, 194, 705, 1941]. Historical [2212]. History [2591, 259]. Hitchin [1139]. Hodograph [169]. Hohenberg [2616]. Hölder [1849, 2491]. hole [151, 878, 2387, 1100, 520, 605, 1287, 359, 474]. hole-free [1287]. holes [951, 989, 2551, 1288, 835, 1290, 2647, 2201]. Holm [899, 1504, 326, 2454, 1009, 1938, 2140, 1105, 1799, 2002, 721, 864, 1127, 1316, 2378, 1711, 247, 82, 420, 454, 2477]. holographic [950]. Holomorphic [290, 1183, 1071, 2437, 988, 1323]. holonomic [2310, 1337]. Holonomy [484, 1260, 1145, 2234]. Hom [195, 1068, 1926, 2223, 2222, 2173, 2389, 803, 2534, 2644, 2290, 670]. Hom-bialgebras [2173]. Hom-modules [2223]. Hom-version [670]. homeomorphic [2552]. homoclinic [328, 475]. homogeneous [1394, 1413, 1651, 2301, 740, 1844, 2121, 610, 2251, 2470, 1417, 1328, 1527]. Homogenization [569, 863, 1936, 2335]. homographic [1567]. homology [1925, 271]. homomorphisms [1831]. Homotopy [1507, 683, 1386, 1174, 1062, 2626, 299]. honeycomb [2687, 2209]. Honor [1410]. Hopf [2710, 2711, 2427, 689, 1926, 552, 702, 256, 2222, 2594, 1438, 1908, 2534, 2644]. Hopper [954, 1924]. Horava [1100]. horizon [2412]. horizons [557, 2201]. Horne [1554]. horseshoes [2277]. host [7, 726]. Householder [344, 317]. Hubbard [807, 1033]. Hulthén [1445, 2032, 118, 1760, 47]. Hunter [1015, 2314]. Hurwitz [1035, 1346, 1132]. Huygens [1987]. hybrid [749]. Hydrodynamic [2420, 1477, 884, 524]. hydrodynamical [551]. hydrodynamics [875]. hydrogen [1067, 825]. hydrogen-like [1067]. hydrogenic [2715, 1813, 2716, 538]. hyper [2306, 1540]. hyperasymptotic [1184]. hyperbolic [2302, 861, 1327, 868, 1443, 2122, 1463, 1894, 2275, 321, 2402, 969, 2162, 1869, 1988, 717, 61, 2515, 807, 1742, 1603, 1498, 1077, 1091, 292, 2097, 2576, 1650]. hyperbolic-type [861]. hyperbolicity [1027, 748]. hypercomplex [1582]. hypercontractive [1550]. hypercube [2304, 2362]. hypercubic [792]. hyperdiamond [1326]. hyperdissipative [1529]. hyperelastic [1576]. hyperelasticity [2671]. hyperelliptic [1100, 2238]. hyperfinite [1857]. hyperfunctions [255]. Hypergeometric [777, 931, 644, 2648]. hyperhamiltonian [2311]. hyperkähler [2311]. hypermatrix [307]. hypersurface [1730, 2127]. hypersurfaces [2646, 1143]. hypothesis [1543, 1666, 1036]. Hysteresis [277].

Ichikawa [1791]. ideal [2312, 1210, 2417, 2479, 2532]. Ideally [4, 2295]. identical [2119]. identically [30]. Identities [1868, 723, 1910, 2113, 2704, 2239, 1866, 2040, 88, 576, 245]. identity [204, 1380, 366, 1783, 2426, 767, 1929]. if [676]. ignition [2256]. II [989, 2226, 1441, 2670, 1021, 232, 2144, 518, 2043, 1443, 880, 2307, 1454, 2370, 1998, 1403, 1502, 1032, 730, 387, 1315, 1820, 1638, 560, 1067, 759, 2040, 1919, 1274, 1144, 1462, 2513, 1186, 891, 1886]. II. [2527, 331]. III [80, 789, 2024, 2542, 1116]. Ill [1529, 1534]. Ill-posedness [1529, 1534]. imaginary [197, 1035, 14]. imaginary-real [197, 14]. IMBq [805]. imitating [892]. immanant [2568]. impact [170]. impedances [708]. implementability [168]. Implementation [582]. implications [1517]. implies [1415, 2034]. improper [208]. improved [1645, 1114, 940]. impulse [953]. impurities [1417]. impurity [2516]. incidence [2360]. incident [1072]. including [2708, 102, 16, 511]. inclusions [1311, 2204]. Incompressible [1509, 2009, 2671, 1511, 2376]. increasing [1466]. Indecomposable [149, 2539, 385, 1914]. independence [91, 62]. Independent [2119, 2232, 30, 556]. independent-value [2232]. Index [86, 671, 1338, 588, 223, 598, 1436, 1947, 95, 1922]. indexed [1923, 1979, 2174, 2560]. indices [853, 1887, 2486]. Indirect [1423]. induced [195, 446, 1113, 2409, 879, 645, 2014, 2064, 2582, 2541, 2495, 988, 1895, 2506, 1328, 2700]. induction [1194]. inductions [2633]. inductive [293]. inelastic [639, 1479, 142]. Inequalities [2647, 2340, 1246, 947, 1471, 1297, 2181, 2273, 77, 1414, 1956, 1738, 1815, 310, 1135, 1251, 1550]. inequality [976, 1957, 2129, 2192, 2212, 1411, 1554, 1420, 2035]. inequivalence [1179]. Inequivalent [144, 759]. inertia [1904]. inextendible [1287]. Infeld [2065, 2214, 658, 1408]. infimum [1049, 1129, 452]. Infinite [1206, 336, 1857, 2056, 2405, 769, 406, 2260, 669, 733, 1811, 1435, 1868, 2377, 619, 813, 792, 821, 1755, 1360, 2323, 522, 2686, 958, 2480, 475, 2700]. Infinite-dimensional [336, 2056, 769, 2260, 1811, 1868, 792, 475]. Infinite-volume [1857]. Infinitely [2105, 245, 2597, 495, 1682, 1012]. Infinitesimal [2440]. infinity [1599, 1145, 276, 700, 1873]. Information [691, 1573, 1297, 2571, 544, 2288, 902, 738, 466, 1956, 2653, 352, 1680, 1550, 1578, 906, 1232, 2085]. information-theoretic [1680]. informational [1437]. informationally [648, 164]. Infrared [1263, 323]. inherently [1525]. inhomogeneity [1098]. Inhomogeneous [1569, 185, 93, 433, 2108, 2386, 588, 746, 1295, 2068, 2196, 191, 1687, 1792, 1438, 2533]. initial [1248, 2570, 473, 784, 2598, 1207, 2068, 436, 2309, 2498, 2513, 1521, 665, 977]. initial-boundary [1207, 2068]. inner [160, 415]. Inozemtsev [1380]. Inozemtsev-type [1380]. inspirals [1184]. inspired [66]. Instability [1082, 2275, 713, 1363]. instant [834, 553]. instantaneous [593, 201]. Instantons [1581, 555, 2501]. instead [910]. integer [840, 1220, 2435, 955]. Integrability [1248, 433, 2429, 191, 1538, 2661, 1579, 238, 1771, 1347, 249, 1637, 90, 1394, 146, 2556, 1290, 969, 2511, 1231, 1029, 610, 913, 1102, 2251, 1566, 1495, 939, 2044, 1563, 2351, 763, 1654, 758, 1391]. Integrable [2319, 2538, 87, 882, 1002, 993, 2282, 386, 2320, 884, 439, 148, 1424, 2665, 809, 969, 2514, 2008, 1313, 69, 717, 16, 511, 71, 96, 171, 188, 418, 366, 2414, 2631, 681, 153, 347, 283, 674, 2432, 621, 1482, 243, 804, 1265, 2426, 1503, 156, 1170, 285, 247, 1076, 1204, 1929, 378, 190, 1658, 1201, 1845]. Integral [768, 1591, 584, 118, 402, 380, 1100, 2337, 1270, 415, 1302, 537, 1787, 2681, 121, 2472, 1757, 1325, 1384, 1151, 637]. Integrals [1926, 2549, 35, 214, 137, 591, 1657, 2238, 1810, 1159, 512, 2587, 656, 1141, 1292, 70, 1156, 329, 2229, 923, 998, 704]. integrand [2179]. Integrated [413, 2266]. Integration [718, 1959, 1786, 2, 2577, 109]. integro [861, 1967]. integro-differential [861, 1967]. intensity [680]. interacting [1948, 319, 694, 524, 331, 1316]. Interaction [1910, 1115, 2353, 2185, 49, 2320, 357, 1212, 639, 1376, 1133, 2017, 2483, 1064, 2233, 443, 1304, 479, 2161, 1985, 2423, 2291, 650]. interactions [456, 1807, 2367, 1181, 1245, 1411, 1608, 1656, 2473, 2348, 680, 1571, 1201]. interactive [2070]. interatomic [1608]. interband [1855]. interesting [2461]. interface [898, 2581, 1952, 2465]. interference [1087, 823, 724]. interior [362]. interlacing [2086]. intermediate [1207, 385]. intermittency [1517]. Internal [958, 659, 1589, 1782, 2206]. Interpolatability [2081]. interpolated [1600]. interpolating [2091]. interpolation [976, 591, 507]. interpretation [528, 403]. Intertwiners [2489, 360]. Intertwining [155, 50]. interval [1441, 380, 392, 1027]. intervals [1275]. Intrinsic [688, 1712, 371]. introduces [1613]. Introduction [1509, 2363, 22, 1410]. Intuitionistic [301]. invariance [642, 2307, 75, 797, 614, 2195, 1654, 1361]. Invariant [747, 1250, 2299, 2414, 1553, 1064, 1688, 926, 184, 2101, 2670, 1625, 90, 513, 876, 476, 376, 2503, 1211, 2293, 751, 1959, 741, 1729, 2331, 1652, 1683, 42, 2264, 394, 728, 72, 366, 1995, 1949, 1559, 245, 330, 1150, 1217, 1864, 927, 1146, 170, 97, 1992, 1816, 243, 1227, 1899, 1644, 409, 1705, 273, 757, 2191]. Invariants [2240, 1647, 1039, 1966, 1221, 2474, 2132, 11, 1252, 835, 1005, 1684, 288, 271, 797, 2476]. Inverse [2219, 1759, 1848, 2617, 501, 2618, 2522, 2625, 2500, 2541, 2529, 854, 2391, 718, 2066, 2114, 106, 1237, 1839, 1578, 1072, 2649, 1597, 804, 1439, 955]. inversely [174, 468]. Inversion [1100, 1488, 1017, 55, 308, 636, 2672]. inverted [20, 775]. invertible [566, 2413]. Investigation [563, 1111, 2312]. inviscid [1510, 1341, 1516]. invisibility [613]. Involutive [2557]. Involutivity [124]. involving [1297, 2461, 1501, 192, 682, 2576]. ion [1433]. ionic [569, 863]. ionization [33, 1757]. ions [355]. IRF [1489]. Irreducible [1113, 289, 606, 1003, 622, 2154, 434, 1039, 354, 1167]. Irregular [2703]. irreversible [1828, 45]. isentropic [1714, 2455]. Ising [1424, 2421, 1870, 618, 1843, 1421, 1305, 570, 2138]. island [2465]. isochronicity [2205]. isokinetic [229]. Isolated [1546]. isomerism [1279]. isometric [167]. isometry [1353]. Isomonodromic [388]. isomorphic [1873]. isospectral [1441, 50]. isothermal [2511, 891]. isotonic [1259, 1056, 2398, 1377]. Isotropic [2476, 1864, 626, 740, 1478, 2376, 812]. Issue [1509, 22, 1410]. issues [565]. Iterated [2549]. iterates [410]. iteration [1382, 754]. IV [1817, 2044]. Ivanov [1916, 1306, 1204]. Iwasawa [472].

J [2666, 1611, 863, 2622, 2572, 261, 2717, 1409, 1931, 2625, 726, 600, 2451, 2021, 510, 2713, 157, 2531, 2719, 197, 1375, 1666, 2712, 1935, 2721, 2064, 635, 263, 1252, 2705, 775, 1612, 2710, 111, 2620, 2183, 2020, 2708, 1887, 1083, 2295, 774, 2621, 1665, 2718, 2362, 2715, 2662, 1932, 2626, 511, 1373, 1129, 2716, 1251, 1934, 397, 2709, 1508, 2720, 110, 2707, 198, 2253, 942, 1217, 599, 2623, 2326, 2714, 1128, 509, 2294, 2325, 2184, 636, 2182, 2711, 199, 262, 823, 1374, 200, 776, 1507, 1983, 899, 2532, 2619, 109, 2624, 1984, 1933, 1019]. Jackiw [405]. Jacobi [1298, 1467, 2557, 2280, 1031, 612, 1469, 1662, 2489, 1577, 1337, 88, 887, 2583, 1086, 245, 83, 1045, 2446, 2177, 2118, 2482, 2560, 1973, 1733]. Jacobi-theta [2482]. Jacobi-type [2118]. Jamiolkowski [737]. Jaulent [1318]. Jauregui [1370, 999]. Jaynes [1840]. Jensen [1158]. Jet [1716]. Joint [2390, 1398, 869, 279]. Jordan [198, 491, 1552, 1831, 1019, 12, 1008]. Jordanian [828]. Jost [928, 1656, 2126]. Journal [22, 1613]. Julia [283]. Jumarie [2204]. Jump [515, 1963, 1762]. Jump-diffusion [515]. jumps [1994, 841, 1032, 744, 1360]. Jupiter [1746]. Jupp [633]. Jürgen [1535].

Kac [253, 1988, 1876, 1742, 2382]. Kadomtsev [2442, 1356, 2094, 2429, 1790, 2062, 2657, 705, 2661, 286, 1505, 1789, 1886]. Kähler [2036, 2306, 1491, 193, 2470, 74, 1540]. Kampé [644, 860, 2284]. Kanai [1539]. Kappeler [2699]. Kapteyn [768]. Kármán [1735]. Kashiwara [2458]. Kaup [1783, 1929, 1838, 379, 2426]. KdV [938, 2178, 1038, 1202, 1941]. KdV6 [188, 533]. Kelvin [1936]. Kemmer [2021, 157, 9, 468]. Kent [633]. Kepler [1001, 865, 459, 1132, 539, 1019, 732, 1267, 1627, 1826, 1378]. Kernel [1587, 2074, 255, 1655, 338, 1380, 2359, 1151]. kernels [2473, 2216]. Kerov [2086]. Kerr [951, 989, 151, 990, 1288, 520, 1605, 1784]. Kerrness [1230]. killing [2600, 65, 2039, 1040, 1730, 1065, 2487, 1328, 2177, 67]. Kim [2035]. kind [768, 652, 1268, 657]. kinds [1204, 190]. kinematics [387]. Kinetic [838, 1787, 1776, 489]. kink [1045]. Kirchhoff [2450, 2107, 1671, 1307, 1943, 419, 127, 1190]. Kirchhoff-type [1307, 1943]. Kitaev [2259]. Klauder [2608]. Klein [943, 118, 1011, 1030, 1382, 2575, 2558, 2246, 1506, 1173, 113, 292, 551, 1987, 1464, 667]. Knizhnik [388]. Knot [1125]. knots [2562]. Koch [1475]. Kochen [640]. Kohno [2569]. Kolmogorov [1610, 2237, 1028]. Konno [1791]. Kontorova [1146]. Kontsevich [2564]. Koornwinder [2525]. Korteweg [2670, 383, 1372, 808, 2320, 1193, 2541, 298, 2693, 2452, 392, 379, 1485, 436, 2524, 2498, 2096, 1273, 1012, 1111, 190, 145, 333]. Korteweg-type [1193]. Kosaki [1947]. Kotera [1588]. Kovalevskaya [1065]. KP [337, 890, 2672]. Krall [1794]. Kramers [2441]. Krein [2293, 236]. Krichever [2523]. Kronig [1544]. KSGNS [236]. Kummer [1984, 1739]. Kundt [471]. Kuper [882]. Kupershmidt [285]. Kuramoto [1904, 2616, 1216]. Kuznetsov [1521].

Laakso [1219]. labeled [1707]. laced [2188, 1742]. ladder [2684, 866, 329, 1103, 2029, 2542, 501]. Ladik [237, 2514, 1845]. lagging [83]. Lagrange [706, 702, 1293, 1922]. Lagrangian [261, 263, 2253, 17, 2553, 2602, 167, 2352, 1199, 368, 1294, 1337, 2410, 2297, 1968, 1733]. Lagrangians [1775, 1865, 72]. Laguerre [1923, 1758, 245, 112]. Lamb [1638]. Lambert [572]. laminar [800]. Landau [577, 846, 1165, 141, 1964, 1602, 1005, 86, 2255, 948, 1608, 2203, 576, 26, 2007, 1848, 1216, 2003]. Landau-type [1216]. Landauer [2373, 2365]. Landsberg [968]. landscape [1386]. Lane [1539, 1289]. Langevin [2206, 133]. Laplace [564, 668, 2120, 495, 973, 2145, 2485]. Laplacian [250, 1180, 585, 2574, 220, 1876, 1101, 2628, 1664, 2576, 1242]. Laplacians [1721, 2359, 698]. Large [889, 1938, 714, 2481, 105, 227, 2666, 2462, 2415, 2211, 1212, 2232, 2236, 2662, 1515, 100, 1235, 2131, 2184, 2356, 2419, 2624, 1998]. large- [2356, 2232, 2419]. largest [1529]. lasso [1759]. Late [367]. Late-time [367]. LATKes [652]. Lattice [372, 952, 982, 2697, 697, 820, 849, 1976, 673, 1706, 1704, 1214, 1311, 139, 2562, 1368, 1358, 2209, 1138, 2169, 1997, 2591, 687, 501, 2397, 156, 1170, 2361, 128, 1789, 677, 654]. lattices [1569, 2133, 609, 2368, 792, 1850, 1464, 1618]. Laughlin [1574]. law [2676, 1778, 2671, 2123, 583, 1108, 2150, 904, 2355, 2053, 1766, 748, 1842, 2215, 522, 955]. laws [1278, 504, 505, 669, 1982, 368, 1240, 1866, 2040, 657, 1754, 2297, 1682, 2361, 752, 2191]. Lax [2620, 1516, 238, 534, 1557, 1357, 2660, 1214, 669, 449, 1266, 881, 1240, 532, 629, 2056, 2432, 413]. layer [712, 1512, 2100, 1605]. layered [2170, 2060]. layers [2445, 1341, 1362, 2544]. leading [2355, 2111]. leading-order [2355]. leads [613]. least [553]. leaves [1874, 1826]. Lee [1271, 2447, 1428, 1421]. left [2172, 1438]. left-most [2172]. Legendre [1666, 1543, 2440]. Leibniz [1217, 1925, 2701, 1150, 686, 2016]. Leibniz-like [1217, 1150]. Lemaître [275]. lemma [1635]. lemmas [1635]. Lenells [71]. Length [799, 1994, 150, 2562, 1796, 62, 450, 1920, 2683]. Length-scale [799]. lengths [1678, 280]. lens [951]. Lensing [951, 989, 655]. Lenz [564, 2120, 2145, 1378]. Leon [2620, 449]. Lesche [839]. less [977].

Level [1396, 1555, 421, 2339, 2451, 733, 2586, 790]. levels [577, 1602, 86, 1233]. Leverett [1276]. Levi [1696, 2580, 796, 1944]. Levi-Civita [1696, 796]. Levin [1967, 744]. Levinson [1626, 901, 1454, 1385, 778]. Lévy [2709, 1896, 1270, 1971, 1330, 136]. Lewandowski [2595]. Lichnerowicz [1641]. Lie [510, 2719, 2710, 2718, 2707, 2714, 2325, 2711, 1984, 2467, 926, 2319, 648, 2611, 195, 1068, 287, 1875, 2384, 712, 289, 451, 883, 149, 239, 412, 550, 1193, 811, 2260, 2579, 2563, 2660, 2188, 1003, 1701, 2329, 335, 2637, 2235, 2614, 652, 935, 1913, 490, 1756, 1647, 2154, 1981, 2461, 274, 1978, 2702, 2055, 1728, 924, 1831, 2523, 343, 1069, 2460, 936, 1440, 2488, 2257, 258, 633, 212, 2492, 1153, 2056, 2432, 453, 1833, 1836, 1908, 2109, 1644, 1946, 2675, 2157, 803, 2290, 444, 1838, 2016, 1739]. Lieb [1410, 755, 2158, 2539, 1659, 1996, 1411, 1738, 1861]. Liénard [2719, 2718, 1162, 2707, 1836, 2325]. Liénard-type [1162]. Lifschitz [1848]. Lifshitz [2448, 1100, 576, 2003]. lift [2202]. Lifting [1821]. light [2643, 1855, 798, 2327, 553]. light-cone [2643]. Lightcone [1882]. Lightlike [2646]. like [1396, 2312, 591, 966, 1773, 1067, 1656, 2501, 797, 382, 2393, 1150, 1217, 1255, 2658, 1343, 1020, 1254, 731, 2466]. likelihood [19, 199]. limit [644, 2462, 1419, 1604, 1524, 20, 775, 1634, 1710, 1032, 24, 229, 1715, 855, 2059, 1424, 593, 1376, 1263, 405, 1474, 1314, 1315, 30, 1140, 1433, 963, 1857, 2473, 558, 2335, 1498, 283, 100, 1235, 123, 653, 2054, 2356, 1526, 82, 1570]. Limitations [1104]. limiting [1972]. limits [1667, 1879, 1714, 293, 841, 1843, 962, 677, 758]. Lindblad [515, 312, 2191]. Lindbladians [2431]. Line [187, 980, 1626, 2226, 2659, 508, 600, 1096, 2475, 2070, 765, 1870, 720, 2243, 2498, 503, 599]. Line-soliton [187]. Linear [1168, 1552, 76, 988, 1116, 1852, 1264, 2279, 1414, 1743, 1967, 520, 1487, 2627, 2604, 1775, 16, 511, 192, 1462, 1323, 496, 100, 1235, 2078, 125, 620, 964, 82, 2516, 1847, 2338, 468]. Linearisable [1697]. linearizable [927]. Linearization [321, 566]. linearized [569, 863, 2409]. lines [2025]. Link [288, 2132, 1684, 271, 560, 1607, 2283]. links [1657, 2562]. Liouville [456, 1685, 1501, 1513, 116, 1296, 233, 1170]. Liouvillian [1288, 1290, 312, 2511]. Lipschitz [822, 1780]. liquid [959, 711, 1470, 2639]. liquid-gas [959]. Liu [676]. Lobachevski [1703]. Local [1778, 2448, 1024, 2527, 1468, 2053, 1135, 1251, 1152, 1014, 2404, 456, 1090, 2537, 2569, 1372, 442, 1898, 1545, 1136, 215, 1655, 546, 1554, 179, 1679, 1731, 1995, 1389, 269, 159, 1353, 748, 499, 983, 1194, 1127, 2351, 2191]. locality [2697, 40, 1268]. localizability [869]. Localization [455, 1700, 2051, 483]. Localized [264, 870, 1018, 2721, 2558, 36, 1197, 1225]. Locally [1328, 1667, 334, 1651, 1480, 853, 1887, 355, 506]. Location [2445, 1421]. LOCC [2081]. Lochner [2042]. loci [855]. log [829, 563, 617, 331, 19, 199]. log-likelihood [19, 199]. log-normal [617]. log-partition [331]. log-periodic [829, 563]. logarithm [1231, 2554]. Logarithmic [469, 1354, 950, 1815, 2394]. logarithmical [844]. Logarithmically [1510]. logical [208]. logics [512]. Long [671, 1512, 1368, 2483, 2422, 1671, 2424, 439, 2615, 1207, 814, 2540, 1971, 822, 436]. Long-range [2483, 2424, 2615, 2540].

Long-run [2422]. Long-term [671]. Long-time [1512, 1368, 1671, 822, 436]. long-wave [439, 1207]. Longo [1947]. look [2593, 2622, 2083]. Loop [1580, 2614, 1262, 414, 2157, 1818, 2510, 1748, 2312, 2638, 769, 573, 795, 1099, 2488, 1226, 607, 608, 1399, 1970]. loop-gas-like [2312]. loops [2080, 331]. Lorentz [846, 2307, 390, 1614, 874, 1267, 1687, 2634, 1014, 1654]. Lorentzian [2626, 1484, 852, 1183, 1062]. loss [1583, 2352, 1792]. loss/gain [1792]. lossless [1583, 2352]. Lotka [153, 773]. Lovelock [2037, 1770]. Low [2638, 1149, 2384, 1176, 975, 2669, 59, 56, 111, 458, 274, 675, 2354, 1919, 1711]. Low-dimensional [2638, 1149]. low-frequency [975, 675]. low-rank [1176]. Low-regularity [1711]. Low-temperature [2354]. Lower [1527, 1921, 983, 2124, 2065, 2242, 835, 228, 749, 1242, 2016]. lower-rank [835]. Lowest [627]. lubrication [177]. luminescent [2327]. lunar [913]. Lund [1046]. Lusztig [2388, 2112]. Lyapunov [2237, 45].

Macdonald [1810]. MacDowell [2234]. machines [216]. Mackey [2504]. macro [350]. macro-observables [350]. Macroscopic [223, 2084, 524, 1361, 1122]. Madelung [542]. Magnetic [1040, 741, 389, 446, 2074, 2313, 2211, 2449, 81, 267, 2279, 1018, 2721, 1657, 1073, 2192, 1412, 1690, 1005, 39, 1721, 1809, 979, 355, 116, 585, 1855, 560, 173, 1294, 594, 2105, 369, 1134, 1417, 1016, 1369, 176, 1627, 1826, 1755, 1343, 1383, 1784, 2221, 1397, 1774, 1363, 1705]. magneto [1013, 939, 875]. magneto-gasdynamics [939]. magneto-hydrodynamics [875]. magneto-transport [1013]. magnetohydrodynamic [2009, 799, 1363, 2375, 1661]. magnetohydrodynamics [661, 2218]. magnetostatic [2065]. magnification [80, 951]. Mahony [1937]. Majda [1516]. Majid [2633]. Majorana [257]. Majoranas [2080]. make [553]. Manakov [169]. Manev [1568]. manifestly [224, 917, 2040]. manifold [2505, 2337, 321, 1959, 2699]. Manifolds [1535, 2528, 2313, 1859, 876, 476, 1463, 550, 1245, 1271, 1874, 1180, 2402, 1491, 2450, 1346, 2470, 2102, 170, 2004, 2146, 297]. Manin [1139]. manipulations [532]. Mann [1945]. Mansouri [2234]. Many [2333, 2432, 2415, 993, 1392, 2043, 1639, 662, 1691, 2005, 2597, 495, 1773, 303, 2105, 1995, 1422, 1508, 245, 684, 2334, 1682, 1012, 1613]. many-body [2415, 993, 1392, 2043, 1639, 1613]. many-electron [303]. many-fermion [1422, 1508]. Many-particle [2333]. Many-poled [2432]. map [566, 1849, 609, 366, 1897, 2297, 2217]. mapping [782, 406, 1255, 1820]. Mappings [2224, 1158, 311]. maps [723, 1657, 2536, 56, 111, 1981, 1552, 256, 488, 802, 560, 881, 1166, 1919, 601, 283, 465, 804, 2404]. Marangoni [1216, 1692]. Marchenko [1778]. marginal [2340, 398]. marginally [1211]. Markov [2545, 1043, 1642, 514, 2350, 617, 1725, 543]. Markovian [2418, 515, 161, 2224]. Marr [207]. Maruyoshi [1612, 1198]. Mass [2533, 1192, 943, 1223, 1445, 1677, 1327, 659, 2108, 1051, 2122, 2463, 1184, 2541, 1290, 1762, 1999, 400, 1319, 797, 50, 53, 2335, 2393, 359, 206, 1842, 2686, 977]. mass-critical [2108]. mass-deformed [2541]. mass-like [797]. mass-ratio [1184]. mass-shell [1192]. masses [1818, 2510, 2553, 1334]. Massive [1760, 2006, 2713, 2280, 990, 949, 1461, 1459]. Massless [2475, 1225, 2067, 353, 1788]. Master [476, 515, 921, 415, 312]. Master-slave [476]. materials [223]. Math [2666, 1611, 863, 2622, 2572, 261, 2717, 1409, 1931, 2625, 726, 600, 2451, 2021, 510, 2713, 157, 2531, 2719, 197, 1375, 1666, 2712, 1935, 2721, 2064, 635, 263, 1252, 2705, 775, 1612, 2710, 111, 2620, 2183, 2020, 2708, 1887, 1083, 2295, 774, 2621, 1665, 2718, 2362, 2715, 2662, 1932, 2626, 511, 1373, 1129, 2716, 1251, 1934, 397, 2709, 1508, 2720, 110, 2707, 198, 2253, 942, 1217, 599, 2623, 2326, 2714, 1128, 509, 2294, 2325, 2184, 636, 2182, 2711, 199, 262, 823, 1374, 200, 776, 1507, 1983, 899, 2532, 2619, 109, 2624, 1984, 1933, 1662]. math-modules [1662]. Mathematical [583, 1825, 1296, 22, 1613, 2015, 595, 1364, 1256, 2239, 2316, 2376, 1575]. mathematics [1883, 51]. Mathieu [2694, 1435]. matrices [2208, 2181, 2642, 1087, 2086, 344, 1778, 1176, 1856, 1322, 315, 2563, 2705, 2706, 1, 1723, 735, 36, 2695, 2025, 493, 986, 881, 2053, 887, 1546, 1800, 842, 2026, 1724, 430, 1077, 499, 388, 1399, 2056, 97, 317, 1151, 1972, 1737, 831]. Matrix [1280, 808, 318, 237, 2154, 1476, 409, 2675, 507, 980, 1626, 2226, 1388, 529, 2167, 2312, 1471, 2073, 1795, 510, 451, 793, 2449, 1877, 2094, 631, 2660, 1108, 2354, 2212, 2232, 1647, 107, 2190, 1870, 235, 619, 1785, 2394, 1919, 698, 1007, 1274, 556, 350, 2491, 242, 1358, 570, 1603, 2126, 2035, 1493, 2432, 1130, 328, 2482, 291, 243, 356]. matroid [1299]. Matsumoto [559]. Matter [1613, 680, 1483, 1561, 1095, 652, 1461]. Maurer [769]. MaxEnt [1342]. Maximal [66, 1024, 310, 2196, 398, 757, 948, 1552, 179, 2097]. maximally [516]. maximum [1398, 2420, 2298, 410, 161]. Maxwell [697, 2106, 1474, 2253, 1792, 2630, 1082, 467, 1853, 223, 1698, 741, 743, 1199, 658, 2149, 748, 1408, 812, 1080, 2210]. Maxwellian [531, 589]. Maxwellians [2106]. Maya [2560]. Mayer [98]. MBPT [587]. McIntosh [1267]. McLennan [41]. Mean [2189, 2473, 1134, 2353, 1948, 1572, 885, 350, 1972]. Mean-field [2189, 1134, 2353]. meanders [1768]. meaning [41]. means [528, 299, 1754]. measurability [869]. measure [1098, 1177, 307, 1110, 2595]. Measurement [2272, 348, 410]. measurements [648, 2081, 52, 1060, 203, 407, 310, 788]. Measures [2299, 1224, 1814, 2194, 2247, 2473, 1553, 887, 2491, 608, 164, 1063, 1374]. Measuring [1359]. mechanical [534, 391, 1556, 2640, 2205, 2196, 2199, 2677, 2530]. mechanics [834, 830, 535, 1092, 815, 1471, 1879, 547, 1444, 460, 513, 1717, 2678, 2122, 845, 530, 1131, 306, 729, 2225, 2045, 2306, 1104, 611, 2337, 1806, 2120, 2371, 1909, 2331, 2008, 2373, 281, 1819, 1820, 537, 1860, 1219, 1492, 85, 728, 781, 1436, 2244, 2030, 51, 2075, 1247, 944, 2393, 2594, 1871, 5, 110, 374, 895, 2117, 2166, 2163, 2294, 2227, 1418, 134, 826, 1254, 2338]. media [569, 863, 1667, 671, 73, 2147, 1511, 2179, 568, 1305, 524, 2170, 1971, 770, 2327, 423, 2060, 812, 2435, 1705]. medium [7, 726, 481, 2279, 1614, 1844, 1295, 1670]. meets [2336]. Mehta [1810].

Meixner [1744]. Mellin [214, 1159]. membrane [440]. membranes [273]. memory [277, 2592, 2377, 1735, 1712, 487, 675, 2247, 497, 1671, 745]. memory-type [497]. merging [1726, 2408]. Mermin [1734]. Meromorphic [152, 1801]. metals [1459]. metaplectic [2380]. method [683, 535, 2010, 448, 883, 2006, 2713, 403, 491, 2071, 1382, 251, 1174, 2059, 2093, 2322, 1723, 1133, 1056, 718, 609, 1292, 106, 1237, 299, 2472, 1876, 294, 1837, 2007, 1272, 2287, 115, 783, 804, 1920, 817, 1020, 1650, 2618, 1507, 1114]. methods [238, 1912, 1870, 1490, 1397, 1829, 1613]. Metric [1863, 2190, 559, 2249, 1747, 239, 490, 1416, 2701, 1866, 2040, 988, 906]. metric-torsion [1866, 2040]. metrics [342, 852, 61, 1144, 362]. MHD [1615]. Mickens [1157]. microelectromechanical [1147]. microscopic [878, 1122]. microstructure [1523]. MICZ [865, 459, 732, 1378, 1826]. MICZ-Kepler [459, 732]. migration [1692]. Mikhailov [763, 915]. Mild [2052, 1360]. Mills [707, 211, 126, 2280, 1565, 652, 2275, 1314, 1315, 871, 728, 781, 1436, 2501, 944, 1716, 320]. Mimetic [2132]. mini [2061]. mini-twistors [2061]. Minimal [2339, 2451, 2108, 353, 1610, 2027, 2006, 2713, 150, 1624, 1920, 2683, 468]. minimal-nonminimal [468]. minimally [2475]. minimization [421]. minimizing [1178]. Minimum [2562, 57, 1602, 435, 410]. Minimum-error [57, 435, 410]. minimum-uncertainty [1602]. Minkowski [1128, 2475, 1211, 2038, 994, 1182, 2283]. minus [1398, 613]. Miodek [1318]. mirror [2036]. Miura [804]. Mixed [1207, 322, 851, 2271, 1967, 1812, 2255, 117, 1952, 2480, 1202, 301, 438, 468]. Mixing [1509, 221, 1815, 2034, 477, 1520, 1934, 543]. mixtures [208, 1888]. mKdV [758]. mKdV6 [533]. mobility [282, 1169]. Mode [2344, 2507, 1139, 2641, 2031]. Mode-sum [2344]. Model [800, 2651, 482, 2353, 1936, 997, 525, 840, 1002, 782, 2462, 1673, 1625, 2265, 1555, 1795, 2185, 1948, 2084, 793, 878, 1477, 49, 662, 799, 849, 2065, 1976, 2671, 908, 33, 1890, 631, 1556, 1544, 1840, 1524, 1271, 2447, 358, 1634, 1212, 1512, 835, 949, 639, 1411, 439, 120, 552, 959, 1010, 2100, 874, 2052, 2421, 1870, 638, 827, 1533, 2055, 523, 2256, 933, 2243, 619, 920, 2483, 1319, 813, 814, 418, 1470, 715, 576, 2488, 1296, 2335, 1317, 1305, 1746, 1776, 83, 978, 570, 1689, 617, 2422, 1304, 1603, 895, 962, 330, 1150, 1217, 770, 207, 1137, 2677, 1985]. model [1146, 2482, 1418, 257, 2138, 1216, 2375, 413, 589, 2031, 2210, 886]. modeling [230, 95, 1829]. modelling [1135, 1251]. models [1218, 2302, 1772, 1396, 2659, 277, 937, 642, 1510, 2259, 81, 1877, 406, 573, 1312, 2507, 2424, 1502, 1142, 930, 2541, 1960, 2232, 1424, 632, 1487, 2665, 1071, 2605, 1523, 492, 2564, 2345, 2690, 2090, 1843, 1876, 350, 1947, 1358, 2095, 2492, 1006, 356, 2287, 1671, 1540, 922, 1476, 1482, 892, 1033, 2400, 2459, 419, 127, 1190]. modes [1288, 1532, 733, 948, 364, 2203, 2650]. Modified [577, 1228, 1490, 2540, 2670, 383, 1372, 808, 1356, 1543, 1666, 298, 201, 1576, 1332, 1046, 436, 1986, 1105, 1799, 1962, 2018, 1515, 721, 1012, 1579, 1588, 1505, 2378, 1711, 1973, 247, 2477]. modular [2133, 2474, 1632]. Modularity [2036]. Modulated [914, 703]. modulation [2616]. module [411, 2331, 1492, 1247]. modules [1354, 447, 2261, 2539, 289, 2538, 888, 2223, 1003, 1662, 597, 966, 236, 802, 60, 385, 1803, 101, 2222, 2173, 732, 1892, 257, 2213, 1914, 396, 2457]. moduli [2643, 61]. Modulus [2401]. molecular [2463, 664]. molecules [120]. moment [2144, 1477, 1577, 1842, 2613, 2623]. momenta [2282, 202]. Moments [719, 2020, 1007, 1274, 2419, 1246, 2073, 2144, 662, 1691, 2005, 975, 1118, 2390]. Momentum [1070, 1463, 2279, 2536, 1403, 1537, 2592, 1938, 1918, 2127, 614, 1037, 67, 700, 616, 2338, 371]. monodromy [2176]. monoidal [1926, 2222, 2258, 2534]. monolayer [1716]. Monomer [2209, 2353]. Monomer-dimer [2209, 2353]. monopolar [371]. monopole [692, 1132]. monopoles [970, 692]. monostable [1528]. monotonic [1610]. Monotonicity [2128]. Moody [253, 1988, 1742, 2382]. Mooney [2671]. Morita [302]. Morse [2708, 1189, 517, 1001, 1055, 102, 791, 1993]. Moser [1587, 1266]. Moses [2033]. most [935, 2172]. motion [830, 2670, 1261, 2625, 2282, 709, 1100, 850, 1640, 329, 1839, 320, 1473, 701, 637, 2645]. motions [1732, 29, 1416, 1971, 1573]. Moussa [393]. moves [2527]. moving [2136, 1749, 2559, 1629]. Moyal [1623, 1686, 462, 843, 1899]. MUBs [1179]. Multi [1388, 2070, 806, 2233, 1792, 2221, 2560, 1278, 2101, 2537, 1646, 2490, 1923, 1608, 2107, 1837, 2542, 1081, 1979, 2174, 2423, 1335, 1080, 1158, 1838]. multi-body [1608]. Multi-bump [806, 2107]. multi-component [1081, 1838]. multi-dimensional [1837]. multi-dimensions [1278, 2101]. multi-Dirac [1335]. Multi-indexed [2560, 1923, 1979, 2174]. multi-Jensen [1158]. multi-local [2537]. Multi-parametric [1388]. Multi-peak [2221]. multi-rogue [1646, 2490]. Multi-soliton [1792]. multi-step [2542]. multi-symplectic [1080]. Multi-time [2233]. Multi-transmission-line-beam [2070]. multi-type [2423]. multibaker [2355]. multiband [2335]. Multibump [2098]. Multicalorons [183]. multichannel [2126]. multicomponent [138]. Multidimensional [504, 505, 1669, 1242, 455, 1740, 2577]. Multilocal [2674]. multimode [2685, 1553]. multinegaton [677]. multiparameter [2660]. multiparticle [53, 1858]. multipartite [307, 1135, 1251]. multiperiodic [677]. Multiple [1599, 1215, 1943, 137, 1748, 2509, 727, 889, 2653, 1314, 1315, 967, 316, 2163, 2294, 1116]. multiple-state [727]. multiplets [2306]. Multiplication [85]. multiplicative [482, 2308, 2422, 227]. multiplicativity [56, 111]. Multiplicity [2328]. multiplier [2474, 843]. multiplying [1295]. multipole [2019]. multipositon [677]. Multiscale [1523, 1918]. Multiscaling [280]. multishell [324]. Multisoliton [139, 677, 1081, 1799]. Multisymplectic [383]. Murakami [196]. mutation [2090]. mutation-selection [2090]. mutual [2653]. mutually [2593, 2622, 315, 1546, 1284, 84]. Myers [2551].

N [2282, 1905, 2356, 546]. N. [1905]. Nagy [666]. Nambu [195, 1068, 39, 550, 2045]. nanocomposites [280]. nanostructures [230, 95]. Nappi [1065, 1802]. Nariai [2037]. natural [683]. nature [483, 2072].

Navier [2318, 2136, 1529, 2242, 1798, 1522, 1891, 1532, 1517, 844, 1514, 226, 1340, 2478, 1527, 2578, 1797, 745, 2453, 1526, 1615, 578, 760, 801]. near [1011, 2106, 366, 1596, 142, 1041, 589]. nearest [1779]. nearest-neighbor [1779]. nearly [2414]. necessary [1558]. needed [1995, 983]. negative [2518, 1338, 223, 1143, 936, 1318, 247, 424]. negative-order [424]. neighbor [1779]. neither [418]. Nekhoroshev [1905]. Nelson [1634, 638]. nematic [1193, 711]. Neoclassical [2416]. Nernst [1670]. nested [1551]. network [1631, 695, 2281]. networks [941, 1320, 1475, 2288, 2082, 593, 363, 1034, 2368]. Neumann [386, 1356, 2081, 2238, 2486, 2504, 2628, 1405, 2063]. neutral [390, 1538, 1360, 219, 2692]. Neutron [1295]. Newell [1783, 1929, 1838, 2426]. Newman [151, 1288]. Newton [591, 1237]. Newtonian [1496, 1710, 800, 1962, 359, 1985, 1797, 123, 701, 1564]. next [2111]. Nicole [552]. Nielsen [39]. Nieto [580]. Nikiforov [1133, 1056, 1650]. Nikodým [488]. nilpotent [649]. nine [865, 1378]. nine-dimensional [865, 1378]. no [1761, 827, 57, 1067]. no-binding [827]. no-pair [1067]. no-signaling [57]. Nodal [1617, 2296, 2687]. Noether [1642, 678, 657, 614]. Nohel [1967]. noise [879, 2449, 2308, 2179, 582, 2206, 170, 227, 1525]. noises [482, 260, 2051, 875]. Non [2245, 2133, 2301, 1556, 1271, 493, 103, 555, 1029, 1105, 1944, 2594, 2033, 1781, 1622, 1255, 2044, 1220, 994, 1128, 507, 2418, 1927, 204, 2569, 1230, 515, 1625, 1879, 2528, 1431, 1685, 437, 1655, 1910, 678, 1614, 2704, 2014, 2064, 1714, 2508, 2071, 1496, 2575, 2346, 1806, 765, 1812, 1424, 1874, 1981, 1268, 2627, 756, 2421, 2332, 1313, 368, 692, 800, 1102, 2709, 1566, 1952, 761, 2139, 1462, 2270, 2104, 2176, 1191, 1309, 1377, 2499, 388, 983, 1985, 1797, 1330, 1992, 1630, 2435, 1308, 1899, 484, 1329, 1965, 1393, 438, 1158, 2618, 1458, 2479, 2532, 1564, 311, 1032]. Non-abelian [1781, 555, 437, 692, 761, 484]. Non-Archimedean [103, 438, 1158]. non-autonomous [2139, 1191, 1309, 2499, 1965, 1393, 1564]. non-backtracking [2421]. non-centered [1910]. non-central [2508]. non-classical [1377]. non-commutative [1879, 1685, 2704, 1806, 1981, 2104, 1899, 1329, 1458]. non-compact [2618]. non-convex [1431]. non-crystallographic [2014, 2064]. non-degenerate [1874, 2270]. non-differentiable [678]. non-dissipative [1614]. Non-equilibrium [1032]. Non-existence [994, 1128]. non-Gaussian [2709, 1330]. Non-geometric [2594]. Non-Grassmann [1556]. non-Hamiltonian [1630]. non-Hausdorff [756]. Non-Hermitean [493]. Non-Hermitian [1622, 507, 204, 311]. Non-homogeneous [2301]. non-ideal [2479, 2532]. Non-integer [1220, 2435]. Non-integrability [1029, 2044, 1102, 1566]. non-integrable [1424, 1313]. non-Kerrness [1230]. non-Lagrangian [368]. Non-Levi [1944]. non-linear [2627, 1462]. non-local [2569, 1655, 983]. non-locality [1268]. non-Markovian [2418, 515]. non-mixed [1952]. Non-modular [2133]. non-Newtonian [1496, 800, 1985, 1797, 1564]. Non-orthogonal [1255]. Non-polynomial [2033, 1992]. non-regular [1308]. Non-relativistic [1271, 1714, 1812, 2332]. non-resonant [2575]. Non-self-adjoint [2245, 765, 2176]. non-selfadjoint [2071]. non-separable [1625]. non-skew-symmetric [388]. non-symmetric [2346]. non-topological [994, 1128]. non-trivial [2528]. Non-uniform [1105]. Nonadditivity [43]. nonassociative [254, 2594]. nonautonomous [1331, 752]. noncentral [734, 174, 819, 1379]. nonclassical [1060, 1331, 68]. Noncolliding [1473]. Noncommutative [1402, 689, 1769, 210, 1139, 1636, 1348, 2261, 66, 460, 699, 302, 1717, 2678, 1660, 304, 306, 2120, 75, 2246, 2690, 2030, 2701, 337, 2470, 1389, 2180, 1345, 2658, 1038, 1227, 2228, 2471, 808]. noncommutativity [828, 2075]. Noncommuting [1554]. noncompact [256, 297]. nonconstant [1599, 1949, 1389]. nonconvex [817]. Nondegeneracy [2447]. nondegenerate [702, 468]. nondiagonalizable [49]. nondispersive [2028]. nondissipative [579]. Nondisturbing [407]. Nonequilibrium [24, 2371, 1418]. Nonequivalence [1698]. Nonexistence [615, 2314, 70, 2103, 393]. nonextensive [427, 281]. nonextensive-statistical-mechanics [281]. nonholonomic [611, 706, 16, 511, 71, 1337]. nonhomogeneous [644]. nonintegrability [707]. noninteracting [105]. nonisospectral [156, 1111]. Nonisotropic [477]. Nonlinear [862, 1585, 2093, 713, 2281, 2664, 1753, 2048, 2455, 1076, 1886, 2530, 2666, 952, 1643, 683, 238, 2684, 2572, 725, 671, 861, 1808, 703, 1589, 1347, 2185, 1590, 2219, 1011, 1248, 433, 185, 1208, 386, 849, 784, 2069, 2277, 2022, 152, 2500, 1853, 1840, 1743, 1174, 949, 1646, 1930, 2490, 148, 1465, 1282, 1407, 1506, 1162, 1982, 1501, 718, 2068, 2236, 2662, 1638, 717, 920, 1499, 71, 70, 477, 806, 2573, 2247, 2151, 800, 2105, 191, 1291, 2709, 749, 2668, 1490, 972, 1695, 340, 1367, 443, 153, 496, 1792, 2398, 1498, 1377, 770, 2376, 1842, 2613, 2623, 1663, 1238]. nonlinear [2077, 2118, 1764, 1709, 1851, 1597, 252, 113, 586, 925, 1975, 1408, 123, 426, 170, 1330, 2221, 367, 2655, 501, 1303, 2351, 1774, 87, 429, 1048, 1400, 1121, 964, 1123, 138, 1740, 13, 200, 377, 1507, 1464, 2516, 766, 1658, 1618, 2220, 2624, 752, 757]. nonlinearities [1616, 174, 1663, 509, 143]. nonlinearity [1643, 703, 2151, 1799, 1605, 1784, 621, 2655, 1664, 2378, 766, 2023]. nonlinearization [1214]. Nonlinearizing [16, 511]. Nonlocal [1163, 2089, 2531, 2320, 1228, 505, 1371, 1896, 540, 731, 496, 895, 2466, 505]. nonlocality [2621, 265, 10, 262, 823, 724]. Nonlocally [2700, 2010, 334, 2439]. Nonminimal [1329, 632, 468, 387]. nonmonotone [1082]. nonperiodic [2023]. nonperturbative [857, 205]. Nonrelativistic [877, 1541, 2447, 1628, 405, 469, 1474, 1056]. nonseparable [49]. nonsingular [55]. Nonstationary [475, 1675, 568, 2478, 2613, 2623]. nonsymmetric [1082]. nontraveling [294]. Nontrivial [2579, 1543, 1666]. Nonuniform [710, 588, 568]. nonzero [2219, 2588, 914, 74]. nor [418]. Nordström [474]. norm [2573]. Normal [2684, 1288, 427, 103, 617, 2321, 1151, 2459]. Normalization [2305]. normed [438, 1158]. norms [2548, 2575, 2197, 352]. Note [1611, 600, 2451, 510, 1666, 1935, 2064, 635, 2705, 111, 2362, 511, 1712, 1251, 1508, 1217, 599, 2184, 823, 1983, 2532, 1933, 393, 1449, 218, 1290, 2148, 2373, 562, 1156, 2578, 1972]. Notes [1298, 1089, 1340]. Noumi [1155]. novel [96, 1081]. Novikov [2345, 2488, 763, 78, 954, 1743, 1849, 2523, 670]. Null [1859, 995, 1614, 1775, 276, 700, 213]. null-plane [995]. number [2666, 2285, 2236, 2662, 1358, 2270, 940, 821, 395, 1363, 958, 2480, 2624]. numbers [1260, 2684, 5, 110, 903, 1692]. Numerical [2255, 1173, 1080, 1645, 672, 666, 466, 2524]. Nusselt [940].

o [2154]. observables [989, 2409, 1321, 1050, 2421, 1437, 1049, 1129, 350, 435, 84, 452]. observation [1377]. observations [1823]. Observed [2027]. observer [1821, 2015]. obstacles [1773, 1500, 595, 1364]. obstructed [215]. obtain [1507]. obtained [2420, 2644]. obtaining [238]. ocean [1468]. Octonionic [812, 2306]. octonions [242]. octopole [1705]. octopole-magnetic [1705]. odd [1478]. Off [1656, 649, 917, 1192, 1397]. off-axis [1397]. Off-shell [1656, 649, 917, 1192]. often [1387]. Old [965]. Oldroyd [2456]. Oldroyd-B [2456]. Olesen [39, 1366]. Olver [2436]. omni [1209]. omni-potential [1209]. One [910, 2348, 461, 982, 627, 204, 2607, 1396, 1528, 277, 1011, 2408, 1286, 2021, 880, 1053, 1193, 1030, 1891, 623, 639, 935, 1885, 1487, 713, 598, 1506, 42, 1609, 1804, 1681, 477, 711, 532, 1054, 590, 1842, 2613, 2623, 1764, 2578, 1797, 553, 664, 1384, 2595, 2618, 145, 468, 1256]. One-dimensional [2348, 204, 1396, 1528, 277, 1011, 1286, 2021, 880, 1193, 1891, 623, 639, 1487, 713, 598, 42, 1609, 1804, 477, 590, 1842, 2613, 2623, 1797, 664, 468, 1256]. one-parameter [1681]. one-way [2408]. ones [366, 2016]. Onsager [2388, 282, 1169]. onto [2224]. open [2367, 626, 253, 970, 1089, 2587, 1801, 587, 2047, 2224]. open-shell [587]. open-string [1801]. Opening [2574]. operation [1403]. operations [1545, 1455, 779, 210, 159]. Operator [2142, 1900, 1548, 1304, 1744, 1354, 980, 1626, 2226, 2141, 1441, 1677, 1280, 928, 182, 104, 2067, 623, 2071, 15, 1224, 1089, 765, 948, 2653, 186, 2632, 598, 2050, 2691, 1501, 2194, 31, 2401, 244, 859, 1197, 2017, 1067, 1619, 2030, 2203, 397, 50, 1881, 2393, 1106, 1755, 2115, 1038, 164, 664, 2672, 1872, 1644, 404, 140, 776, 2396, 1824, 1788]. operator-fractal [2194]. Operators [1168, 1696, 456, 1807, 905, 2684, 184, 2175, 947, 1752, 2717, 389, 1092, 1042, 1585, 856, 2269, 2074, 510, 451, 2193, 866, 1300, 880, 2089, 2531, 1357, 2448, 2380, 777, 900, 2704, 2437, 1414, 293, 666, 623, 2503, 541, 1004, 668, 1794, 935, 2150, 1052, 853, 1887, 1075, 313, 854, 979, 2304, 2362, 1196, 116, 1486, 778, 2187, 235, 963, 103, 1401, 1681, 1240, 1596, 1952, 329, 1103, 2029, 2542, 1039, 2413, 2216, 2664, 2231, 651, 2540, 1578, 375, 2348, 2186, 2176, 608, 295, 395, 2056, 2432, 1329, 1636, 2496, 915, 1242, 1488, 1847, 757, 2338, 2266]. Oppenheimer [2316]. optical [2571, 364, 191, 2076, 2617, 1618]. optically [2327]. optics [433, 1238, 2455, 13, 200, 327]. Optimal [2688, 391, 1209, 1721, 1520, 1934, 435, 1352, 252, 1834, 1451]. Optimum [2159]. OPUC [1244]. Orbifold [652, 2025]. Orbifolds [655, 2474]. orbit [1484, 894, 2567, 2708, 2093, 1974, 102, 1804]. Orbital [1643, 1851, 1343]. orbits [1466, 1898, 1679, 707, 1231, 2554, 172, 366, 913, 1102, 1568, 1267, 475, 1391]. Order [311, 1149, 241, 1966, 1677, 2167, 2388, 249, 1250, 442, 1639, 2282, 2689, 2065, 2013, 1394, 485, 152, 1651, 326, 1214, 2238, 1646, 1930, 935, 364, 1698, 1268, 1652, 2355, 2111, 560, 1213, 2483, 21, 1373, 2410, 191, 329, 1800, 754, 651, 2257, 2349, 2476, 1283, 153, 1191, 174, 1360, 359, 219, 1921, 2444, 1864, 927, 1851, 1780, 2655, 818, 1558, 1536, 1121, 2609, 1705, 128, 1111, 424, 2016]. Order-dependent [311]. ordered [2346, 1052]. ordering [2684, 1233, 45]. orderings [2141, 186, 1744, 1026, 1033]. ordinary [1639, 146, 496, 1254]. organized [1693]. orientational [818]. oriented [1267, 1586]. origin [956]. Orthogonal [186, 2580, 510, 451, 406, 810, 158, 889, 1730, 2434, 2495, 2008, 1758, 1800, 1979, 2174, 2446, 2177, 1255, 2612, 2091, 1008]. orthonormal [1885, 2148, 382]. Orthosymplectically [376]. oscillating [338, 1670]. oscillations [749, 1490]. oscillator [1259, 577, 2572, 566, 1113, 1484, 517, 1591, 1347, 1109, 2525, 2021, 2570, 542, 2387, 2383, 1991, 150, 1005, 1282, 624, 774, 2246, 312, 2495, 462, 1056, 2055, 646, 459, 2029, 1893, 2216, 1086, 1766, 2160, 174, 2398, 1037, 1622, 1377, 2077, 2118, 1764, 1085, 2351, 1920, 2464, 1055]. Oscillators [2228, 683, 782, 829, 1539, 689, 2652, 1904, 1174, 1253, 2720, 563, 2104, 1624, 479]. Oscillatory [2670, 2482, 593, 819, 1379]. Osterwalder [697]. Ostrovsky [2314]. other [1444]. output [1178, 1022]. overlap [410].

P [1092]. P-self-adjoint [1092]. packet [2686]. packets [2028, 2161]. packing [525, 179]. Painlevé [238, 1877, 237, 2322, 1157, 2315, 1145, 1046, 931, 2703, 1921, 2044, 2520, 1973, 1350, 1351]. pair [238, 2660, 2620, 449, 1057, 1266, 1067, 629]. pairs [1989, 2712, 1214, 669, 1612, 1198, 1909, 2564, 532, 1718]. Palatini [2171]. palenstrophy [1520, 1934]. para [1582, 192, 2146]. para-Grassmann [192]. para-Hermitian [2146]. para-hypercomplex [1582]. parabola [1398]. Parabolic [258, 1940, 290, 2089, 2531, 2022, 485, 2052, 1160, 21, 1373, 1888, 1283]. parabolic-type [2089, 2531]. paradox [2428]. parafields [107]. parallel [390, 2190, 1234]. parameter [772, 2601, 2458, 571, 858, 1681, 594, 1851, 1170]. parameterization [1969, 1339, 1110]. parameters [1392, 2043, 1930, 2490, 478, 818, 1085]. parametric [1388, 542, 2270, 2160, 2227]. parametrically [1466]. parametrized [1659]. Parasyuk [2006, 2713]. Parity [2268, 2493, 144, 759, 1161]. Parity-time [2268]. part [1303]. Partial [1768, 1832, 1536, 861, 163, 2010, 547, 1059, 876, 1452, 1375, 504, 505, 2022, 146, 1601, 1057, 1257, 1656, 243, 1968, 1658]. partially [2231, 1364]. particle [926, 1445, 118, 879, 2625, 2333, 2505, 2387, 626, 868, 1443, 837, 157, 2403, 2310, 59, 1312, 1542, 1100, 2497, 1212, 1703, 75, 1669, 1628, 270, 743, 1765, 2172, 2127, 1629, 1839, 9, 354, 1325, 1438, 1985, 2119, 2467]. particle-hole [2387]. particles [1415, 1002, 517, 7, 726, 2387, 2463, 1411, 838, 144, 759, 1470, 1693, 1317, 1538, 524, 2430, 2018, 468]. particular [2520]. partition [58, 331]. partitions [270]. partners [777]. passage [234, 919]. passing [2142]. passive [131]. Past [557]. Pastur [1778]. patch [2344]. Path [118, 1591, 1384, 2337, 1302, 2304, 2362, 537, 1141, 1757, 1325, 161, 2229, 2, 637, 109]. path-integral [1757]. paths [888, 2371]. pathway [1787]. patterns [1507]. Pauli [2688, 401, 735, 827, 2366, 1538, 1718]. PDE [384]. peak [2221]. Peakon [2140]. peakons [82]. Pekar [786]. Pempinelli [2620, 378, 449]. pencil [2142, 140, 776]. pencils [318]. Pendulum [129, 365, 328, 273]. penetrable [595]. Penney [1544]. pentagon [2360]. percolation [2424, 234]. Peregrine [1646, 1930, 2490]. Perelomov [2270]. Perfect [67, 1193, 1213, 613, 825]. perfectly [2430]. perimeter [2696, 1834]. Periodic [2615, 618, 707, 1231, 1313, 2554, 366, 1102, 1568, 114, 569, 863, 1466, 703, 829, 879, 1565, 1504, 2136, 1964, 2242, 1595, 1371, 1018, 2721, 225, 1465, 1311, 1835, 2250, 1009, 2493, 1368, 1486, 1918, 2574, 2281, 2606, 172, 913, 2694, 480, 563, 570, 1234, 1153, 665, 1202, 1711, 82, 2324]. periodically [703, 1964]. Periodics [1391]. permanent [409]. permittivity [1835]. Permutability [135]. permutation [2231, 409]. Perry [2551]. Persistence [2454]. Persistent [1418]. perspective [1347, 1654]. Perturbation [1725, 683, 309, 727, 491, 2071, 2575, 1174, 319, 638, 2139, 1507, 1824]. Perturbations [937, 520, 284, 2125, 574, 442, 2538, 2211, 2364, 1435, 1368, 2540]. Perturbative [1262, 2305, 2274, 1594, 658, 897, 857, 205]. Perturbed [2492, 1357, 2524, 1291, 168, 1664]. Peter [2634]. Petiau [2021, 157, 9, 468]. Petrov [471, 2201]. Petrovskii [1610]. Petviashvili [2442, 1356, 2094, 2429, 1790, 2062, 2657, 705, 2661, 286, 1505, 1789, 1886]. Pfaffians [1402]. Phase [879, 81, 2654, 900, 2243, 2309, 121, 2472, 715, 1757, 1086, 617, 2085, 637, 277, 1320, 1623, 2267, 1859, 1956, 959, 1676, 2604, 1978, 1295, 1437, 555, 1421, 2180, 83, 978, 2159, 1427, 1488]. Phase-integral [121, 2472]. Phase-space [637]. Phases [201, 364, 914, 484, 1381]. phenomena [1504, 1009, 2002, 1453]. phenomenon [2533, 2477]. Phillips [534, 1357]. photoionization [1384]. Photon [327, 2507, 2031]. Photonic [1850, 849]. Phys [2666, 1611, 863, 2622, 2572, 261, 2717, 1409, 1931, 2625, 726, 600, 2451, 2021, 510, 2713, 157, 2531, 2719, 197, 1375, 1666, 2712, 1935, 2721, 2064, 635, 263, 1252, 2705, 775, 1612, 2710, 111, 2620, 2183, 2020, 2708, 1887, 1083, 2295, 774, 2621, 1665, 2718, 2362, 2715, 2662, 1932, 2626, 511, 1373, 1129, 2716, 1251, 1934, 397, 2709, 1508, 2720, 110, 2707, 198, 2253, 942, 1217, 599, 2623, 2326, 2714, 1128, 509, 2294, 2325, 2184, 636, 2182, 2711, 199, 262, 823, 1374, 200, 776, 1507, 1983, 899, 2532, 2619, 109, 2624, 1984, 1933]. Physical [1456, 2009, 2155, 448, 2602, 403, 1451, 821, 2417]. Physics [1613, 2239, 1722, 756, 2400, 22, 1613]. Picard [2238]. picture [2350]. piecewise [1844]. Piscounov [1610]. piston [1714]. Pitaevskii [690, 1371, 1715, 1406, 967, 187]. Pizzetti [1293]. Planar [1079, 1154, 1834, 2610, 2555, 405, 1500, 2180, 2209, 1391]. Planck [2709, 1710, 1749, 2559, 76, 897, 1670, 1493, 919, 1330]. Plane [1604, 2125, 995, 1690, 1832, 462, 728, 1986, 2515, 2650, 1325, 2321, 522, 1488]. plane-wave [728]. planes [1179]. planetary [1036, 1746]. plasma [373, 94, 1010, 2605, 531, 749]. plasmas [1530, 746, 1332, 502, 773, 1397]. plasmonic [95]. plastic [1210]. plate [1234, 2291, 1793, 1983]. play [983]. ple [2016]. Plebański [1404]. plurality [274]. Plus [613, 118]. Plus-minus [613]. Pochhammer [2284]. Podolsky [995]. Pohlmeyer [1046]. Poincaré [873, 2381, 1324, 1701, 2344, 2112, 1186]. Point [1901, 2515, 224, 2027, 2719, 1193, 491, 2268, 526, 1741, 956, 1411, 2344, 744, 1773, 1869, 2718, 1693, 354, 1881, 1830, 1871, 2707, 2248, 1097, 1473, 2482, 105, 664, 1836, 2325, 1091, 1458, 1041]. Point-form [1901]. point-group [2268]. point-like [1773]. points [779, 1725]. Poisson [1083, 1911, 2427, 507, 1963, 2669, 1228, 550, 417, 702, 744, 18, 2052, 1172, 274, 1171, 2690, 881, 2298, 806, 2247, 210, 1103, 1893, 2668, 176, 1826, 2488, 1481, 1708, 1360, 1670, 1797, 2453, 977, 891, 2220]. Pol [2694, 299]. polar [250, 698]. polarization [1727]. polarized [1524, 1084]. poled [2432]. poles [458]. Pollaczek [1744]. polygonal [996]. polygons [61, 2143]. polyharmonic [1486]. polyhedra [2143]. Polyhedral [1321, 1841]. polylogarithms [984, 1958]. polymers [2484]. Polynomial [709, 1954, 1639, 2386, 11, 1252, 889, 855, 2168, 871, 610, 532, 459, 1103, 2033, 773, 603, 1992, 893, 506, 2116]. polynomials [984, 2263, 2164, 2141, 1951, 1795, 2648, 954, 182, 406, 1543, 1666, 2704, 631, 158, 1112, 1732, 860, 1699, 1923, 2434, 186, 2495, 500, 2008, 2055, 1977, 2635, 1995, 1758, 1086, 245, 1979, 2174, 2117, 2612, 1284, 1744, 2560, 97, 112, 1648, 291, 1008, 965, 1924]. Polytope [2109]. polytropic [751]. Pontryagin [1335]. Popov [948, 2203]. population [1963]. pore [134]. Porod [1118]. porous [569, 863, 671, 1511, 1971, 1670]. portraits [2309]. Pöschl [1133, 2608, 2116, 2584]. posedness [1529, 1338, 2377, 660, 1533, 1534, 822, 1515, 113, 1127, 1216, 1014, 424]. Position [2592, 50, 1398, 2122, 2463, 2596, 2541, 310, 2172, 400, 1319, 2393, 206, 2686]. position-dependent [2122, 2463, 2541, 400, 1319, 2393]. Position-momentum [2592]. positions [202]. Positive [1181, 1016, 163, 2181, 2642, 696, 1059, 1452, 1375, 160, 1224, 488, 2256, 2486, 1999, 1257, 400, 2573, 1166, 1919, 2366, 164, 465, 2404, 2220]. positive-operator-valued [164]. Positivity [2588, 697, 1897]. possessing [1939]. Possible [387, 2332, 2242, 51]. post [359, 1361]. post-Newtonian [359]. postselected [348]. potential [1415, 1223, 1445, 1530, 1621, 879, 2625, 1948, 928, 1431, 1760, 185, 1208, 157, 1241, 334, 2252, 2380, 2598, 1894, 2508, 20, 775, 1541, 1209, 2708, 540, 47, 2529, 2392, 2052, 734, 679, 2667, 102, 791, 523, 1804, 1918, 1796, 244, 2550, 1197, 2053, 610, 1629, 1417, 1839, 2498, 9, 1425, 1962, 1867, 479, 1792, 206, 2115, 2165, 2076, 2131, 2184, 1048, 1020, 297, 2618, 819, 1379, 1993, 2584, 1256]. potentials [114, 1092, 2032, 118, 1954, 708, 402, 2282, 2211, 2069, 332, 504, 2122, 1394, 2526, 1371, 2123, 2198, 583, 357, 403, 623, 158, 691, 1750, 1382, 855, 2293, 1001, 1923, 48, 945, 1017, 2615, 1675, 1133, 1055, 904, 2391, 2008, 575, 1231, 1486, 28, 2554, 1955, 2105, 2251, 1731, 1596, 2542, 2413, 50, 2443, 245, 2033, 55, 2233, 1719, 2126, 641, 2334, 1670, 927, 2612, 603, 322, 308, 636, 1151, 1650, 2116, 592]. Power [231, 232, 772, 197, 583, 741, 904, 658, 858, 1520, 1934, 1616, 14, 509, 143, 522, 955]. power-law [583, 904, 522]. power-type [1616, 509, 143]. powers [1180, 465]. Prandtl [821, 958, 2480]. Prasad [215, 970, 2658]. Pre [1902]. Pre- [1902]. predominance [1256]. Prelle [2322]. Prelle-Singer [2322]. Prepotential [1055]. prescribed [2573, 2414]. prescribing [2645]. presence [446, 879, 150, 2592, 1016, 1461, 1753, 1077, 2206]. preserving [2474, 969, 1552, 1339]. pressure [1365, 1731, 177]. pressure-gradient [1365]. Presymplectic [2066]. primes [1543, 1666]. Primitive [2299, 1468, 2413, 518, 789, 1817]. principal [2539, 1674, 484, 413]. principle [2158, 2420, 1591, 676, 2385, 40, 57, 2373, 2205, 2603, 2298, 2399, 53, 105, 1586, 1308, 1335, 1987]. Principles [1769, 689, 1731, 1044, 2619]. private [544]. probabilistic [685, 435, 330, 1150, 1217, 580]. probabilities [2509, 1976, 1175, 348, 2159, 279, 2397]. probability [1814, 2172, 2166, 817]. Probing [2037]. Problem [1975, 2428, 2522, 943, 1779, 1990, 1951, 579, 712, 421, 1248, 547, 993, 1392, 2043, 1639, 1856, 996, 805, 690, 1136, 222, 810, 267, 548, 635, 2553, 1714, 2395, 729, 2071, 2556, 1005, 1806, 1118, 266, 730, 1699, 853, 1887, 702, 1844, 1376, 1302, 462, 854, 1468, 2391, 1314, 1315, 582, 1609, 602, 338, 1207, 2068, 2066, 1346, 2441, 2205, 436, 2114, 2309, 1906, 865, 1986, 106, 1237, 710, 913, 1568, 93, 2058, 1566, 2046, 1422, 1508, 2443, 568, 651, 98, 1724, 1567, 1293, 1867, 1378, 1115, 1072, 2209, 1521, 115, 1249, 1334, 1303, 1774, 305, 175, 955, 1793, 1983]. problem [589, 882, 1391]. Problems [539, 644, 1250, 2415, 473, 89, 1599, 2013, 1890, 2301, 2225, 1052, 363, 2340, 2107, 1019, 732, 1627, 1826, 1490, 1578, 161, 895, 2327, 2649, 870, 233, 1848, 2677, 1405, 595, 1364, 2464, 2645]. procedures [1794, 1308]. Procesi [1316, 305]. Process [2230, 1398, 2042, 1032, 2172, 1473, 2482, 279, 665, 2047]. processes [1745, 1043, 1642, 2696, 526, 841, 1854, 802, 2709, 1830, 1871, 2481, 1828, 1330, 1725, 105, 543, 1308, 1381]. processing [1957, 2129]. Product [1862, 2407, 2341, 1178, 160, 2502, 2190, 415, 401, 735, 1900, 2691, 968, 1057, 2699, 1546, 1892, 430, 1058, 2595, 1899]. production [1232]. productivity [671]. products [722, 2181, 1662, 107, 63, 2177, 1928, 1972]. programme [2440]. Progress [1575]. Projecting [2321]. projection [2014, 2064, 1039]. Projective [2196, 2025, 906]. projector [2305, 2274]. projectors [2458, 492]. prolongation [528, 431, 1740]. Proof [1149, 599, 1924, 755, 2150, 1551, 999, 174, 521, 676, 503]. proofs [1635, 625]. propagating [916]. Propagation [2149, 2241, 1236, 1775, 83, 1792, 423, 1605, 1784, 1975, 1774]. propagator [2682, 1140, 1061, 1559, 1054, 919, 780]. Proper [2411, 2390, 1007, 1274, 346]. Properties [2648, 2342, 95, 1189, 2269, 1771, 1300, 880, 2252, 908, 2279, 1205, 2582, 1583, 2429, 1290, 1430, 2246, 2050, 2454, 462, 1652, 2111, 585, 1843, 2231, 2130, 375, 1377, 2115, 1038, 1764, 2423, 2217, 87, 1658, 928]. property [2208, 221, 188, 38]. protocols [907]. prototypes [1328]. provide [2481]. proving [56, 111]. Przanowski [1491]. Pseudo [866, 2276, 577, 1927, 862, 1702, 1883, 2282, 828, 160, 1874, 2667, 2140]. Pseudo- [2276]. Pseudo-bosons [866, 577, 862, 1883]. pseudo-Euclidean [2282]. pseudo-fermions [1927, 1702]. pseudo-Gaussian [2667]. pseudo-Hermitian [828, 160]. pseudo-peakon [2140]. pseudo-Riemannian [1874]. pseudoalgebraic [1088]. Pseudobosons [108, 240]. pseudodifferential [389, 979, 2496, 377]. pseudoharmonic [1086]. pseudoscalar [2526]. pseudosphere [390]. Pseudospin [819, 624, 774, 1376, 791, 1379]. PT [1092, 1357, 2396]. PT-symmetric [1357]. Publisher [1935, 2064, 2705, 111, 2184, 1983, 1933, 1611, 600, 2451, 510, 1666, 635, 2362, 511, 1251, 1508, 1217, 599, 823, 2532]. Pullback [1331, 1965, 1564, 68]. pulse [2241, 1637, 1081, 2617]. pulsed [1384]. pulsrodon [939]. Pure [1487, 928, 2685, 1545, 2403, 2037, 471, 15, 652, 397, 159, 1464]. pure-state [15, 397]. pure-states [2403]. Purely [2569]. purification [2680]. Puzzles [204]. pyramids [1150, 1217].

Q [971, 687]. Q-balls [971]. Q-difference [687]. QCA [1997]. QHWM [510, 451]. Qiang [346]. Qiao [1372, 1606, 674, 1791]. QLGA [1997]. Qu [1791]. Quadratic [1506, 739, 1114, 926, 54, 2384, 1426, 49, 2719, 2303, 2565, 2718, 988, 2707, 174, 1377, 1836, 2325, 2609, 140, 776, 1929]. quadratic-form [1929]. quadrupole [1705]. Qualitative [1867, 1290]. Quantal [1342]. Quantifying [2288]. quantisation [1070, 1949]. quantities [2239, 2045, 2355]. Quantization [2580, 254, 2505, 1981, 2433, 2527, 2581, 2028, 846, 697, 272, 1717, 1902, 2303, 306, 1302, 1016, 651, 1895, 2506, 590, 346, 1816, 643, 1922, 470]. quantization-induced [2506]. Quantized [550, 1117, 1152]. Quantum [830, 1549, 2546, 1261, 230, 119, 1623, 2570, 1477, 1444, 1767, 2369, 2370, 2364, 293, 44, 1142, 2640, 2336, 1117, 512, 464, 2402, 802, 1815, 1219, 349, 2075, 2465, 1296, 2199, 2160, 1249, 2374, 1720, 1279, 826, 2397, 1350, 1351, 1446, 2467, 926, 2521, 114, 2651, 2545, 736, 982, 1415, 1149, 834, 905, 857, 535, 205, 1594, 2271, 1021, 2509, 1092, 1165, 782, 2125, 534, 1471, 647, 1957, 1879, 2405, 1539, 1805, 2185, 2625, 1726, 1419, 2079, 2588, 2408, 2333, 421, 2339, 2451, 2697, 314, 442, 2543, 1569, 2367, 547, 1489, 542, 2698, 2387, 1181]. quantum [642, 1176, 868, 1443, 2273, 460, 772, 2449, 1580, 699, 513, 2259, 1545, 1239, 1678, 2223, 406, 52, 1717, 2678, 2122, 1053, 921, 2341, 1178, 77, 412, 845, 2385, 1023, 907, 530, 1131, 635, 1385, 1996, 306, 208, 1386, 729, 2225, 1175, 2507, 345, 15, 1060, 1224, 1814, 1285, 1541, 932, 2337, 2274, 259, 1722, 795, 738, 203, 1768, 2592, 1806, 2120, 2541, 1896, 1435, 1089, 466, 1812, 593, 2502, 387, 1686, 1854, 2529, 1177, 2230, 25, 1050, 1628, 1552, 874, 2621, 1075, 1066, 36, 2587, 771, 407, 1455, 1769, 737, 1841, 300, 1456, 42]. quantum [2025, 2008, 695, 221, 604, 30, 2373, 1819, 1820, 2197, 582, 1728, 1332, 2547, 537, 352, 2264, 838, 1860, 2345, 1734, 1855, 1078, 85, 2468, 1022, 1390, 117, 963, 1124, 1401, 268, 1067, 728, 781, 1436, 2244, 2030, 2340, 2445, 51, 2701, 779, 1803, 210, 1319, 1099, 1049, 1129, 1135, 1251, 2196, 1362, 605, 1134, 1553, 397, 688, 739, 1839, 38, 576, 946, 354, 944, 1490, 1550, 987, 2393, 1389, 2130, 2594, 1947, 269, 62, 931, 1895, 2506, 1830, 1846, 5, 110, 316, 374, 98, 1763, 83, 2350, 457, 161, 37, 1225, 168, 2248, 1593, 2117, 2398]. quantum [906, 2048, 1438, 2159, 2585, 607, 608, 1138, 2077, 2118, 1764, 346, 1997, 2591, 788, 2078, 115, 2677, 450, 693, 2166, 2423, 2323, 1725, 653, 1630, 2163, 2294, 664, 543, 2227, 2406, 308, 636, 410, 1970, 1439, 1381, 202, 1536, 10, 262, 260, 823, 724, 2051, 409, 1453, 2224, 1575, 452, 2375, 1254, 2404, 1886, 2530, 1044, 2619, 2031, 2338, 1570, 548]. quantum-dynamical [1886]. quantum-mechanical [2677, 2530]. Quantum-to-classical [1767]. quarter [1986]. quarter-plane [1986]. Quartic [1893, 566]. Quasi [1259, 1074, 1964, 1991, 1153, 1085, 989, 1218, 2263, 2175, 1402, 628, 1779, 1288, 1512, 1465, 1861, 1486, 1918, 2121, 72, 1803, 1135, 1251, 2694, 1607, 2594, 2398, 1229, 2166, 112, 817, 964, 1692, 1202, 2324, 2535, 918, 1933]. quasi-Bohmian [2121]. quasi-classical [2263]. Quasi-coherent [1991, 1085]. quasi-entropy [1861]. quasi-equatorial [989]. Quasi-exact [1259]. quasi-exactly [1218]. quasi-free [1229]. quasi-geostrophic [1512, 918, 1933]. quasi-harmonic [2398]. quasi-Hermitian [2175]. quasi-Hopf [2594]. quasi-invariant [72]. quasi-linear [964]. quasi-normal [1288]. Quasi-periodic [1964, 1153, 1465, 1486, 1918, 2694, 1202, 2324]. quasi-periodically [1964]. quasi-probability [2166, 817]. quasi-solvability [112]. quasi-spin [1402]. Quasi-splitting [1074]. quasi-static [1607]. quasi-steady [1692]. quasi-Turaev [2535]. quasi-uniform [1779]. Quasiclassical [536]. quasicrystals [813]. quasiexactly [374]. Quasifinite [1913]. quasigeostrophic [712, 370, 227]. Quasigraded [1399]. quasihomogeneous [1567, 1293]. quasilinear [2328, 1278, 1617, 2296, 1169, 2098, 682, 2576, 1889]. quasilocal [628]. Quasiperiodic [29, 1318, 298]. quasiperiodically [328]. quasirelativistic [1215]. quasistable [1901]. Quaternion [1491, 2036, 588]. Quaternionic [1168, 2357, 530, 1131, 2123, 2198, 8, 2326, 1648]. quaternions [181]. qubit [516, 2168, 986, 1995]. qubits [545, 11, 1252, 546]. quenched [1149, 77]. question [1738]. quintic [1499, 1695]. Quispel [809]. quiver [215, 320]. qutrit [1452, 1257]. qutrits [1059, 1375].

Rabi [2507, 2031]. Racah [448]. Radford [2222]. radial [1278, 1246, 861, 908, 104, 541, 1376, 121, 2472, 698, 2443, 1151]. radiating [1776]. Radiation [224, 1192, 1562, 471, 2071, 480, 2160]. Radiation-reaction [1192]. Radiative [1667, 2409, 1609, 661, 322, 1187, 1661]. Radon [488, 147]. railroad [269]. railroad-switch [269]. Randall [1074]. Random [525, 1877, 2371, 1970, 128, 2651, 736, 1415, 507, 1667, 1457, 73, 1940, 2073, 2086, 2089, 2531, 1178, 2448, 2438, 2207, 1108, 1544, 2354, 1224, 526, 2212, 162, 1768, 1478, 2179, 1721, 2695, 1295, 575, 30, 2547, 493, 349, 2053, 399, 2247, 178, 2425, 1007, 1274, 350, 2484, 242, 2422, 347, 2209, 430, 2591, 100, 1235, 105, 291, 1151, 1571, 2404, 1847, 1972, 831, 2266, 133]. randomly [1295]. randomness [1060]. range [402, 2424, 466, 2615, 1804, 1368, 2172, 2483, 2540, 2348, 1971]. ranges [666]. Rank [1136, 2633, 1176, 1059, 1452, 1375, 412, 1701, 2329, 835, 1885, 1065, 398]. Rank-inductions [2633]. ranks [1257]. Rapid [2034, 1815]. rapidly [328]. rarefaction [898]. Rate [784, 1767, 1290, 568, 2422, 2453, 964, 2576]. rates [1712, 316, 1498, 1352]. ratio [197, 1184, 179, 14]. Rational [1155, 2118, 2601, 723, 2517, 32, 1001, 1055, 2029, 1399]. rationalities [2632]. rationally [1923, 2542]. ray [2061, 1829, 1185]. Raychaudhuri [998, 704]. Rayleigh [940, 1950, 1363, 2480]. rays [798]. Razumov [1881]. reach [553]. reacting [1888]. reaction [1192, 1528, 2627, 2243, 1518, 2444, 1965, 1393, 2000]. reaction-convection-diffusion [2627]. reaction-diffusion [1528, 2243, 2444, 1965, 1393, 2000]. reactions [1518, 561]. reactive [1609]. Real [1465, 1800, 257, 2319, 2012, 910, 197, 2461, 1346, 248, 14, 1459, 1946, 470]. real-time [470]. Realigning [1457]. reality [546]. realization [1354, 2387, 1095, 428, 950, 1124, 1674, 1826, 381]. Realizations [287, 212, 1580, 2636, 2093, 1841, 1893, 2519]. reciprocal [2284, 2317, 441]. Reconstructing [1500]. Reconstruction [450, 975, 1209]. Recurrence [1979]. Recursion [1240, 793, 1038, 2672, 915]. Recursive [1794, 581]. redistribution [680]. reduced [2312, 1322, 1, 1647, 2100, 1978, 1064]. Reduction [2022, 1754, 2677, 2381, 2171, 1633, 1136, 494, 2547, 72, 1294, 2566, 788, 1824]. reductions [1163, 1193, 562, 2655, 243]. reentering [878]. Referee [2254]. reference [1182]. Refined [1949, 793, 1070]. reflected [898]. Reflection [898, 1125, 1978]. Refocusing [798]. Reformulating [46]. reformulations [2166]. refraction [898, 588]. regarding [1390]. Regge [167, 458, 656, 1046]. regime [1212, 827, 2465, 499, 1235, 1985]. region [425]. regions [2628]. Regular [1648, 2265, 250, 2518, 1338, 658, 1308, 1872, 1873]. regular-singular [1872]. Regularity [2300, 2299, 1516, 1517, 1923, 844, 2693, 1942, 226, 1340, 220, 711, 2578, 1711, 1615, 2218, 578, 801, 487]. regularization [2596, 2215]. Regularized [1141, 1510, 1276, 913, 576]. Reisner [519]. Reissner [474]. related [2357, 1092, 1396, 2010, 2696, 334, 2071, 1236, 1761, 2568, 2608, 2008, 2382, 233, 156, 2439, 882]. Relating [2406]. Relation [1543, 1666, 2271, 2687, 1771, 2369, 2370, 2360, 902, 415, 582, 1323, 1630].

Relations [1077, 241, 2286, 1042, 1585, 2388, 2272, 2057, 1741, 2592, 746, 1775, 2237, 99, 351, 988, 869, 1094, 1979, 155, 783, 2439]. relationship [1840]. relationships [747]. Relative [1098, 1869, 168, 1021, 1917, 996, 2553, 129, 1732, 2128, 787, 571, 1731, 833, 2085]. Relativistic [834, 846, 1412, 1056, 2512, 749, 1687, 8, 2326, 1397, 2283, 1259, 224, 917, 1047, 517, 659, 2189, 1286, 662, 1714, 357, 1271, 1710, 120, 1812, 746, 1587, 2332, 221, 604, 355, 1067, 739, 62, 1461, 1867, 1182, 965, 977, 2630]. relativity [1098, 2307, 1821, 51, 2072, 794, 2645]. Relaxation [423, 561, 1904]. relaying [1877]. relevant [462, 687]. relics [1483]. Remainder [2273]. remark [2484]. Remarks [2135, 844, 2035, 1615, 2218, 2175, 948, 601]. Renner [2679]. renormalisation [1403]. renormalizable [1960]. Renormalization [1041, 231, 232, 1998, 2596, 122, 29, 2200, 2001, 1226, 2609, 716]. Renormalized [357]. Rényi [43, 2129, 2128, 2130, 641, 2406]. reparameterization [611]. Repeated [2367]. Representable [628, 1, 2115]. Representation [1613, 768, 1280, 2682, 542, 432, 2678, 1614, 353, 1646, 2490, 1981, 734, 1099, 584, 1323, 1763, 607, 1026, 1558]. Representations [848, 1802, 1728, 1078, 2460, 627, 2537, 231, 232, 1113, 2638, 2386, 149, 2341, 606, 622, 360, 1159, 203, 2489, 2330, 1913, 2154, 2502, 236, 281, 1275, 1492, 1247, 2196, 556, 242, 1229, 1928, 453, 45, 693, 2166, 1970, 2675, 2224, 381]. representatives [1898]. Representing [1707, 1727]. reps [2331]. repulsion [1442, 605]. repulsive [2391, 2423]. required [2015]. requirements [1858]. rescaling [1070, 319]. reservoir [406]. residue [1329, 1636]. resolution [204, 2587, 312]. Resolvent [2211, 765]. resonance [1599, 70]. Resonances [534, 1184, 1223, 1620, 946, 2585, 693]. Resonant [2060, 1555, 2575, 540, 916]. resonant-level [1555]. Resource [2408]. resources [2588, 463]. respect [2593, 2622]. Response [2719, 2662, 2716, 2711, 2134, 76, 2649]. rest [834, 359]. rest-frame [834, 359]. restoration [2131, 2184]. Restricted [1239, 466, 1334]. restriction [2595]. restrictions [1003, 622]. result [338, 2024, 1277, 1905]. resulting [1442]. results [2353, 857, 205, 529, 1610, 2408, 1322, 1053, 2571, 33, 2456, 1761, 2377, 1514, 1390, 2348, 607, 1343, 134, 580, 965, 1041, 1788]. retarded [1425, 128]. Retraction [1507]. Return [646, 366]. reversal [2587]. Reversibility [2078, 410]. Review [29]. Revisited [2285, 319, 867, 1861, 1485, 183, 1516, 1495, 1872]. revivals [461]. rhomboidal [2046]. Riccati [1163, 266, 730, 1089]. Ricci [1327, 1903]. Riem [912]. Riemann [1666, 1346, 2522, 1543, 1890, 668, 540, 1349, 436, 568, 63, 1495, 1115]. Riemannian [365, 1164, 2313, 1271, 2337, 1874, 764, 2450, 234, 297]. Riesz [108, 2245, 2071]. rigged [1447]. rigging [415]. Rigid [1600, 1732, 368, 569, 863, 2520, 1907]. Rigidity [1999, 1344]. Rigorous [2312, 1610, 41, 2353, 2337, 1343, 2480, 1041]. ring [1165, 1382, 734, 1746, 519, 819, 1379]. ring-shaped [1382, 734, 819, 1379]. rings [857]. Riordan [893]. rise [2461]. Ritz [1950]. Rivlin [2671]. Roberts [809]. Robertson [275, 1680, 1683]. Robin [1592, 2582]. Robinson [2037, 1996]. rod [1576]. rods [2450]. Rogers [500]. rogue [1646, 2490, 1306]. role [2027, 5, 110, 2085]. Romanovski [2117]. root [1266, 181, 296]. root-type [1266]. roots [1732, 855]. ropelength [2562]. Rosen [2708, 102]. Rosochatius [847, 2063]. Rössler [1265]. Rota [1875, 2186]. Rotating [1406, 517, 660, 371]. rotation [1686, 821]. rotational [1413, 1279]. rotationally [564, 513]. rotations [2688, 973]. rotator [2125]. rotators [1734]. Rouet [649]. rounding [1149]. Routh [1294]. Routhian [72]. Ruijsenaars [494, 2517]. rule [197, 1224, 14, 346]. rules [1912, 722, 1653, 1205, 430]. run [2422, 1123]. Runge [564, 2120, 2145, 1378]. runged [501]. Rydberg [2715, 2716, 1813].

Salam [182]. same [2593, 2622]. sample [1549, 1778, 733, 499]. sandpile [1976]. Sandwiched [2129]. Santini [169]. Sasaki [559]. Sasakian [290]. satellites [1746]. satisfies [2129]. satisfy [1161]. Satsuma [2320]. saturable [2151]. Saturn [1746]. Savadkouhi [198]. Savart [1106, 2417]. Sawada [1588]. Saxon [1223, 118, 1020]. Saxon-like [1020]. Saxton [1015]. Scalar [2194, 118, 990, 2006, 2713, 2475, 122, 1133, 1680, 1102, 1461, 681, 131, 1272, 2215]. Scale [1425, 799, 2307, 1073, 992, 2514, 2072, 330, 1150, 1217, 1864]. scale-invariant [330, 1150, 1217, 1864]. scaled [1596, 570]. scales [2009, 754, 2278]. Scaling [1475, 355, 653, 747, 723, 583, 841, 1032, 1424, 1492, 2359, 1152]. Scarf [1650]. scatterer [1091]. Scattering [114, 824, 1223, 1445, 2543, 645, 2225, 743, 2358, 2229, 1873, 2522, 943, 73, 534, 2219, 894, 402, 2449, 1357, 2301, 2500, 729, 2541, 1762, 1844, 2314, 1263, 1773, 2391, 718, 1275, 1796, 1638, 1759, 106, 1237, 2170, 2348, 1072, 684, 2430, 1077, 1597, 2165, 1848, 2060, 693, 2617, 474, 501, 1439, 595, 1364, 2618, 327]. scenario [981]. scenarios [1321]. scheme [861, 55, 2677, 1080]. schemes [383, 1969, 1403, 1339]. Scherk [1640]. Schneider [2517]. Schott [956]. Schrader [697]. Schrödinger [1409, 1931, 1666, 775, 1887, 1665, 2681, 509, 200, 776, 952, 1643, 627, 980, 1626, 2226, 238, 1807, 2101, 1043, 725, 941, 1222, 1990, 1621, 703, 2219, 2074, 46, 1286, 433, 542, 185, 1051, 981, 2689, 1852, 2069, 673, 1241, 1543, 2448, 2380, 2301, 2500, 1617, 2296, 583, 2225, 2057, 1894, 1840, 403, 623, 20, 541, 2637, 902, 1646, 1930, 2490, 1465, 853, 1407, 1349, 1675, 598, 1172, 1982, 313, 1809, 1310, 1668, 2667, 2391, 2304, 2362, 265, 731, 2068, 116, 244, 1197, 2017, 920, 1499, 71, 399, 1955, 400, 806, 2673, 2151, 2105, 191, 1629, 1596, 1952, 1369, 2668, 2413, 50, 53]. Schrödinger [1858, 931, 2540, 1616, 1538, 1323, 972, 1695, 2233, 1753, 443, 1755, 1792, 206, 2115, 1663, 2466, 1238, 155, 1383, 1709, 1851, 1597, 2099, 395, 113, 925, 2297, 143, 1130, 45, 1833, 2227, 501, 87, 1048, 1020, 1914, 138, 13, 377, 140, 1889, 1824, 2464, 1254, 396, 766, 1618, 2220, 2457, 752, 592]. Schrödinger-like [731, 2466]. Schrödinger-type [2464]. Schrödinger-Virasoro [1833]. Schur [2434, 291]. Schwartz [2564]. Schwarz [1984, 933, 1739]. Schwarzschild [66]. Schwinger [498, 434, 1039]. scintillation [73]. Scott [1412]. Screenings [1401]. Scri [2513]. SDiff [1404]. sea [319]. Seba [264]. Second [1878, 166, 1639, 485, 1035, 237, 1698, 1237, 1800, 754, 2349, 153, 1360, 219, 1352, 1921, 2655]. second-order [1639, 754, 2349, 1360, 219, 2655]. sections [1613]. sector [320]. Segur [1783, 1929, 1838]. Seiberg [1685, 215]. Selberg [591]. Selberg-like [591]. select [2589]. selection [1205, 2090]. Self [342, 1752, 2717, 820, 2303, 1693, 45, 616, 2516, 454, 899, 1696, 980, 947, 1092, 2245, 437, 1269, 908, 2411, 1412, 956, 765, 1762, 1668, 1429, 28, 871, 324, 2281, 1906, 1034, 63, 2176, 1229, 890, 1519]. Self-adjoint [1752, 2717, 45, 1696, 980, 947, 1762]. Self-adjointness [2303, 1229]. self-avoiding [1429]. self-consistent [1269, 890]. Self-dual [342, 437, 908, 2411, 1668, 2281]. self-duality [871]. Self-force [616]. self-generated [1412]. self-gravitating [324, 1906]. Self-organized [1693]. self-potentials [28]. Self-similar [820, 454, 899, 1034, 1519]. Self-trapping [2516]. selfadjoint [1626, 2226, 2071]. Semi [2269, 861, 1181, 406, 1853, 813, 1490, 522, 1705]. Semi-classical [2269, 1853, 1490, 1705]. semi-densities [1181]. semi-discrete [861]. semi-infinite [406, 813, 522]. semibounded [456]. semicircle [2053]. Semiclassical [2161, 780, 2073, 2144, 879, 2682, 855, 2441, 2465, 2586, 934]. semicontinuity [68]. semidiscrete [237, 1201, 758]. semidualisation [2026]. semiduals [1353]. semifinite [313]. semifinite-gap [313]. semigroup [2545, 278]. semigroups [696, 1847]. semilinear [2101, 1599, 1173, 487, 675, 2000]. semisimple [1978]. sending [1475]. sense [656]. Sensitive [478]. Separability [270, 1678]. Separable [1168, 209, 1625, 163, 2081, 2341, 738, 1283]. Separation [1935, 1884, 896, 1846, 2446, 2177]. separatrix [2358]. sequences [2141, 1189, 2086, 1977]. sequential [296]. Series [299, 231, 232, 768, 772, 1159, 860, 658, 858, 385, 934, 2618]. Serrin [801]. Serrin-type [801]. set [456, 2593, 2622, 163, 2301, 1386, 1, 1950, 283, 817, 1536]. Sethi [911]. sets [1545, 960, 2513, 640]. setting [1911, 1888]. seven [2501]. several [1392, 2043, 11, 1252, 2239]. sextuple [251]. Shabat [46, 1355, 2441, 235, 858]. Shafer [2397]. shallow [660, 721]. Shannon [2653, 641]. Shape [2700, 1625, 277, 642, 975, 876, 2424, 1118, 2083, 2293, 234, 1359, 245]. shape-invariant [245]. shaped [1382, 734, 819, 1379]. Sharma [2436]. Sharp [976, 509, 143]. Shastry [1396]. shear [1961, 1341, 2650, 2376]. shear-free [1961]. Shearing [1187]. sheaves [2506, 519]. sheets [851, 1123]. Sheffer [1794, 1977]. shell [917, 1192, 587, 1656, 1341, 649]. shells [2054]. shift [1385, 2148]. shifted [1667]. shifts [1676]. Shimamoto [207]. shock [568]. shocks [898]. Short [2337, 1637, 439, 2150, 2148, 1804, 1368, 2149, 814, 1081, 2617]. short-range [1804, 1368]. Short-time [2337]. short-waves [439]. sided [410]. Siegel [988, 2583]. Sierpinski [663]. sigma [908, 949, 1487, 1071, 492, 2564, 933, 2492, 1540]. sigma-models [1487]. sign [1697, 1048, 766]. sign-changing [1048]. signal [1000]. signaling [57]. significance [648, 756]. signs [1663]. similar [820, 1034, 1519, 454, 899]. Similarity [1749, 897, 238, 1163, 104, 2559, 1173, 1952]. Simons [1668, 1014, 2103, 877, 697, 437, 1095, 908, 2132, 1684, 1698, 1474, 1310, 1819, 1820, 2347, 193, 558, 1700, 1770, 2482, 2630, 2210]. Simons-matter [1095]. Simple [2032, 625, 2457, 755, 289, 1584, 1701, 1743, 2461, 1287, 521, 279, 665, 2047, 1644, 686]. simplicial [1195, 720, 519]. simplicity [1183, 825]. simplified [662, 799, 1010, 568]. simply [2188, 1742]. Simulating [1429, 1194]. simulation [1645]. simulations [1395]. Sinai [2237]. sine [596, 341, 1654, 2530, 445, 135, 523, 96, 974, 1045]. sine-Gordon [596, 445, 135, 523, 96, 974]. Singer [2322]. Single [73, 2403, 1858, 842]. single-electron [1858]. Singular [284, 887, 114, 250, 1245, 623, 1511, 1017, 638, 307, 116, 244, 2677, 1872, 961, 1972, 1733]. Singularities [27, 1369, 655, 204, 685, 2687, 894, 1531, 652, 439, 1697, 177]. Singularity [2555, 1752, 2717, 2697, 957, 1161, 2076, 1303]. singularly [1291, 1596]. Sitter [2475, 1264, 2344, 218, 2646, 1140, 1727, 1143, 323, 1061, 1559, 987, 168, 125, 1987]. Sivashinski [2616]. Sivashinsky [1216]. Six [1930, 2095, 2520, 1822, 2403, 486, 1581, 1649]. Six-dimensional [2095, 2520]. Six-parameters [1930]. six-sphere [1581]. sixth [1652]. sixth-order [1652]. size [1549, 2647]. Sjöstrand [2380]. skein [888]. skew [2546, 2563, 388]. skew-symmetric [2563]. Skin [531]. Skyrmion [615]. Skyrmions [723]. slave [476]. slicings [1024]. slope [339]. slow [1966]. slow-fast [1966]. Smale [482]. Small [980, 2226, 1154, 1314, 1315, 1745, 231, 232, 2079, 473, 1904, 1604, 855, 363, 1066, 2483, 1367, 684, 2430]. Small-energy [980, 2226]. small-h [855]. small-world [363]. Smallest [1469]. Smilansky [2265]. Smooth [674, 1298, 1015, 1799, 2513, 661]. smooth-cavity [1298]. smoothness [2513]. Smorodinsky [739]. Snell [2123]. Snyder [828]. Sobolev [976, 1510, 2575, 1815, 1527, 2692]. Sobolev-type [2692]. Soft [48, 332, 2354]. Soft-core [48]. softcore [2392]. Sokolov [2087]. solid [840, 859, 2121]. solids [181, 1343]. solitary [820, 1851, 1507, 247]. Soliton [298, 1119, 1777, 1889, 1254, 1201, 928, 808, 1617, 2296, 1939, 669, 2541, 2665, 1349, 2494, 523, 1499, 188, 814, 532, 2058, 2631, 773, 187, 1792, 680, 674, 1400, 1789, 767, 533]. Soliton-like [1254]. soliton-potential [523]. Solitonic [1137, 815]. Solitons [2620, 449, 2670, 703, 552, 1606, 2149, 191, 1367, 502, 1663, 994, 1128, 2102, 1654, 2210]. Solution [624, 774, 1787, 2498, 1379, 2353, 2241, 782, 402, 1639, 1497, 662, 141, 2518, 1894, 445, 2497, 2541, 730, 2255, 94, 959, 2100, 2052, 1849, 734, 2667, 658, 1266, 173, 2656, 121, 1782, 176, 568, 773, 1072, 2327, 2586, 252, 1408, 2221, 2217, 2227, 1127, 2453, 138, 1306, 2003, 2103, 2630, 762, 1661]. Solutions [2122, 1262, 781, 489, 1829, 1664, 2666, 1218, 596, 2328, 2318, 2101, 2032, 126, 1409, 1222, 1931, 1990, 325, 579, 1589, 118, 712, 928, 1395, 1561, 612, 1565, 2108, 808, 169, 1852, 386, 81, 849, 1356, 2136, 2320, 332, 1599, 1964, 2214, 505, 908, 485, 152, 2134, 2242, 1798, 1890, 2037, 2198, 2301, 1617, 2296, 1939, 326, 1853, 1714, 1743, 2508, 1863, 1496, 1212, 2558, 439, 1646, 1930, 2490, 1478, 1406, 298, 1465, 2708, 1517, 855, 751, 47, 1015, 1010, 1576, 2605, 2452, 425, 2627, 1665, 1162, 741, 139, 1749, 2559, 1809, 1916, 1365, 1652, 1468, 102, 791, 1314, 1315, 2493]. solutions [2256, 731, 1313, 1160, 1608, 1796, 2236, 2662, 728, 69, 1499, 1656, 2281, 1911, 1955, 171, 188, 27, 299, 806, 2140, 2573, 2107, 897, 814, 144, 759, 2105, 2098, 1028, 2694, 567, 294, 1289, 1937, 1369, 1155, 1081, 1799, 1215, 931, 934, 872, 1462, 1770, 978, 972, 682, 1567, 2186, 822, 370, 661, 629, 2478, 1097, 1360, 1527, 770, 674, 1243, 1921, 1238, 2499, 1383, 2118, 1709, 1137, 2578, 1985, 1797, 922, 1519, 2520, 721, 864, 429, 1400, 1121, 213, 1588, 1307, 2378, 1202, 1711, 1318, 1650, 13, 200, 1889, 1943, 1507, 1789, 2436, 454, 891, 899, 1254, 2324, 766, 760, 801, 1618, 2023, 2220, 2624]. solutions [677, 1564, 533]. solvability [2302, 1248, 1750, 1609, 112, 278, 2031]. Solvable [1001, 1218, 2415, 542, 937, 1392, 2043, 1639, 49, 403, 158, 691, 1055, 400, 374, 2033, 2334, 2077, 346, 603]. solve [1020]. solved [1640]. Solving [2428, 2464, 861]. Some [2175, 1300, 1356, 1823, 1502, 930, 1428, 1960, 148, 2377, 601, 1550, 607, 1343, 2323, 1558, 2439, 1041, 801, 1788, 447, 1261, 574, 1954, 2273, 1322, 1245, 2148, 1728, 656, 1421, 532, 2029, 2130, 1358, 563, 565, 1719, 2478, 590, 994, 1128, 580, 377, 1201]. Sommerfeld [215, 970, 1816]. Sommerfield [2658]. Sommerfield-like [2658]. Song [1791]. sound [2071, 1236, 1352]. Source [1380, 224, 2475, 741, 2170, 145]. sources [2408, 658, 890]. Space [1168, 23, 65, 2510, 627, 284, 2537, 66, 1808, 1484, 1409, 1222, 1931, 1990, 1879, 990, 1623, 2193, 90, 845, 304, 2563, 2225, 372, 2596, 1100, 2344, 1703, 2120, 218, 1211, 2643, 2275, 1821, 1956, 1686, 597, 1050, 4, 2295, 1302, 2246, 2050, 1665, 1978, 2514, 2148, 2646, 1093, 1196, 1918, 585, 933, 1062, 2626, 798, 1437, 64, 2550, 2030, 61, 487, 865, 692, 675, 21, 1373, 688, 2276, 556, 1757, 2443, 1858, 2583, 557, 821, 161, 168, 2159, 608, 1186, 67, 994, 1128, 416, 561, 125, 2435, 2163, 2294, 1648, 1427, 1182, 1026, 1558, 2228, 1393, 2085, 637]. space [2283, 2338, 1447]. space-fractional [2225, 2163, 2294]. space-times [4]. Spacelike [1143, 2646]. Spaces [1127, 1696, 342, 2036, 2357, 447, 1940, 1560, 2405, 1181, 868, 1443, 1510, 1448, 1529, 2039, 2411, 1651, 1902, 1338, 1537, 1811, 2330, 1730, 1480, 2665, 764, 756, 2433, 1683, 968, 1219, 349, 1057, 2701, 936, 2382, 2180, 1328, 74, 1527, 843, 1928, 2323, 1922, 438, 1158, 1793, 1983, 424]. spacetime [2475, 1264, 2038, 2552, 1143, 987, 1389, 2054, 2291, 742, 1987, 2457]. spacetimes [1880, 2409, 1444, 1065, 1680, 1287, 1461, 1229, 1586]. spacing [1036]. spanning [1034]. Sparse [880, 2695, 2165]. spatial [2669, 2571, 836, 2090, 1568, 653, 2119]. Spatially [1268, 1051, 740]. spatiotemporal [461, 1796, 1499, 477]. Special [1509, 1690, 902, 22, 1303, 644, 909, 1743, 300, 557, 2115, 1110, 1507, 1410]. Specific [1066]. Specker [640]. Spectra [2392, 2413, 2266, 389, 2032, 197, 541, 536, 945, 14, 538]. Spectral [456, 204, 1189, 1779, 2067, 1052, 2050, 312, 244, 1681, 2606, 38, 2348, 2176, 1764, 2076, 1594, 1834, 772, 1685, 1385, 1414, 1017, 854, 2194, 2574, 2561, 858, 2058, 887, 2491, 2649, 870, 2124, 1720, 1303, 2496, 2464, 882, 1737, 880]. Spectroscopy [857, 205]. Spectrum [850, 1703, 1415, 1625, 1079, 2507, 1761, 867, 2445, 1362, 944, 1846, 1358, 2104, 2248, 2115, 664]. specularity [531]. speed [1610, 423, 553]. speeds [1528, 2084, 1413]. Sphere [2291, 1026, 1808, 211, 1951, 2310, 810, 1281, 1869, 1882, 28, 986, 1581, 1106, 1558, 371]. Sphere-plate [2291]. spheres [2261, 3, 2511, 1640, 942, 134]. Spherical [1807, 1264, 179, 868, 1443, 1024, 1679, 1171, 2121, 1341, 178, 960, 1272, 2054, 134, 977]. spherical-shell [1341]. Spherically [978, 65, 1880, 126, 852, 1565]. spheroidal [896]. Spin [840, 570, 1298, 1402, 2416, 1396, 230, 517, 894, 1569, 2698, 1659, 77, 1910, 1073, 970, 1524, 2503, 1142, 2082, 1084, 1572, 2708, 624, 1854, 774, 874, 102, 791, 1804, 394, 196, 1857, 1421, 2483, 144, 759, 38, 1881, 1947, 1846, 2145, 316, 1970, 2459]. spin- [144, 759]. spin-0 [517]. spin-1 [2483]. spin-cluster [1073]. spin-foam [874]. spin-orbit [2708, 102, 1804]. spin-orbit-coupled [894]. spin-polarized [1524, 1084]. Spin-spin [840]. spinless [1445, 694, 2018]. spinning [359]. Spinor [1099, 213, 2521, 910, 1320, 90, 2307, 1139, 2072, 2041]. Spinors [1822, 2307, 1025]. spins [2654, 1004]. spirals [339]. Split [1069, 256, 343, 38]. splitting [1074, 328]. spoil [1950]. Spreading [1528]. square [2625, 2529, 2399, 1839, 2686]. squared [1704]. squares [393, 1229]. squashed [2345]. squeeze [2393]. squeeze-like [2393]. Squeezed [2164]. Stability [1672, 1524, 2456, 744, 1831, 1417, 2046, 1519, 438, 1158, 247, 82, 1464, 1643, 990, 332, 255, 1030, 1967, 839, 321, 2377, 1407, 2106, 2604, 854, 1514, 487, 2368, 1277, 1144, 2650, 1462, 2350, 1689, 821, 1367, 1309, 187, 1360, 219, 1851, 1408, 1780, 985, 506, 953, 301, 2692, 442]. stabilization [89, 497, 667]. Stabilizer [1956, 1094]. stable [971, 2027, 1270, 1027, 136]. Stäckel [1065]. standard [866, 257, 1254]. Standing [2069, 1310, 1643, 1208, 1407]. Stanley [519]. star [119, 1385, 1596, 1928]. Stark [2585, 1439, 1391]. stars [891]. state [2651, 840, 1549, 2509, 2032, 1948, 1726, 2408, 2049, 2682, 1852, 849, 1448, 2092, 1241, 2548, 1717, 215, 1178, 727, 2395, 1891, 380, 2447, 15, 1811, 1715, 47, 1430, 638, 57, 791, 42, 355, 1574, 1093, 1078, 1433, 2504, 228, 1135, 1251, 397, 688, 2542, 556, 159, 435, 316, 1238, 2499, 1709, 922, 1384, 429, 1048, 381, 1692, 1650, 581, 13, 200, 2023]. state-adding [2542]. state-deleting [2542]. States [2593, 2622, 952, 209, 2164, 2572, 2271, 943, 1402, 1457, 1223, 108, 628, 1808, 446, 163, 119, 2409, 894, 1286, 1631, 2387, 2685, 1545, 1059, 1452, 1375, 2403, 2383, 2341, 2448, 2597, 2438, 900, 59, 1602, 2285, 1991, 1281, 2582, 1612, 1198, 795, 738, 1768, 1901, 1282, 1812, 1762, 536, 516, 540, 1442, 2190, 3, 2162, 2314, 1172, 2250, 1450, 300, 2608, 1456, 270, 1056, 904, 221, 514, 2547, 1804, 2111, 1486, 408, 117, 1067, 1619, 1257, 2151, 1674, 369, 1553, 1995, 1086, 480, 1358, 398, 2270, 2343, 443, 1719, 1624, 942, 206, 1377, 2091, 586, 561]. states [2374, 123, 2163, 2294, 1648, 1816, 1085, 780, 409, 1063, 1374, 2700, 2584, 468, 2266, 1447]. Static [126, 2639, 1097, 1654, 1880, 1561, 1565, 615, 2511, 1607]. Statics [39]. Stationary [952, 429, 703, 403, 658, 1765, 1028, 395, 561].

Statistical [2244, 134, 1471, 2371, 2246, 904, 2373, 281, 1871, 1418]. statistics [507, 1555, 348, 2608, 1765, 394, 2368, 100, 1235, 464]. Steady [1238, 13, 200, 2049, 1497, 849, 1672, 2669, 2604, 922, 2455, 1692, 1618]. steady-state [849, 922]. steepness [1905]. Stein [1516]. stellar [322]. step [231, 106, 2542, 665]. steplike [436]. steps [888, 2207]. Stewartson [1939]. sticky [1317]. Stieltjes [241, 1473]. Stirling [2684]. stirring [1520, 1934]. Stochastic [482, 2381, 2308, 2299, 2247, 347, 470, 1745, 2042, 876, 2136, 1963, 2652, 476, 744, 802, 2149, 714, 1291, 176, 2139, 934, 1746, 98, 2481, 962, 1360, 219, 1476, 1418, 1381, 745, 429, 1645, 1965, 953, 1216, 128, 227, 2692, 875, 1525]. stochastisation [1036]. Stockwell [201]. Stokes [1797, 2453, 2318, 1497, 2136, 1529, 2242, 1798, 1522, 1891, 1532, 1517, 844, 2025, 1514, 1513, 226, 1340, 2262, 2478, 1527, 2578, 745, 1526, 1615, 1242, 578, 760, 801]. Stone [2504]. Stora [649]. straight [2543]. strategies [1177]. Stratonovich [807]. strength [1809]. stresses [130]. stretching [1522]. strict [2548, 251]. string [137, 440, 1801, 2062, 2657, 1225]. strings [1594, 39, 985]. strip [813]. strips [2248]. Stroganov [1881]. Strong [1795, 1985, 1405, 1013, 959, 1010, 1548, 2017, 176, 2035, 1797, 1751, 311]. Strong-coupling [311]. Strongly [2315, 683, 2134, 514, 1670, 2131, 2184]. Structural [117, 1038, 1451, 280]. Structure [2231, 2629, 2249, 995, 549, 1747, 1497, 2133, 1322, 431, 811, 1, 2235, 15, 360, 929, 1552, 1533, 2493, 514, 60, 1825, 2121, 957, 1790, 397, 688, 1166, 1919, 1949, 2379, 2481, 74, 1603, 2126, 2096, 2492, 939, 384, 2376, 276, 664, 286, 2459, 156, 1740, 767, 1929, 378, 333]. structured [1963]. structures [547, 708, 1898, 2135, 2637, 2311, 18, 1083, 1982, 2250, 1837, 1088, 1482, 1335, 1899, 2361, 1838]. stu [835]. studied [1298]. Study [107, 989, 208, 2399, 164, 2287]. studying [2524]. Stueckelberg [224, 917, 2344, 605]. Sturm [456, 233]. sub [365, 1164, 1426, 1636]. sub-Dirac [1636]. sub-quadratic [1426]. sub-Riemannian [365, 1164]. subadditive [1731]. subadditivity [1548, 2035]. subalgebra [510, 451, 848, 448]. subalgebras [1913, 1728, 1946]. subcritical [1529]. subdiffusive [76]. subdomains [1106]. subdominant [855]. subentropy [2342]. subgroup [2459]. subgroups [412, 1885, 1632]. subject [1234]. subjected [1066]. submanifolds [883, 2646, 2488]. subsets [1093]. Subsingular [1892]. subspaces [2545, 1074, 162, 1992, 1648, 2397, 757]. Substituting [565]. substitution [2413]. subsystems [504, 1953, 2182]. subtraction [2006, 2713, 1403]. subwavelength [480, 1494]. Successive [1341]. Sufficient [702, 1558]. sum [1912, 722, 393, 2344, 430, 2115]. Sums [973, 984, 1958, 2549, 1748, 860, 30]. sunrise [1818, 2510]. Super [1112, 165, 175, 955, 627, 2282, 1982, 2331, 2096, 2258, 1783, 1740, 1076, 2457, 1186]. super-bosons [2258]. Super-Galilean [165]. super-integrable [2282]. Superalgebra [1095, 1647, 2154, 2055, 1078, 1124, 1401, 2519, 2460, 2018, 453, 2675, 381]. superalgebras [1878, 78, 467, 289, 253, 189, 2461, 2632, 924, 343, 1742, 1200, 670]. superbialgebras [335]. Supercharges [1540]. superconductivity [1641]. superconductors [2629, 878, 2255, 369, 502]. superconformal [58, 1909, 31, 1492, 1247, 385, 2095]. supercoset [933]. supercurves [1071]. superdoubles [486, 274]. superfield [2487]. superfluid [1194]. superfluids [1413, 967]. superGalileon [1865]. supergeometry [2564]. supergravities [1186]. supergravity [1243]. supergroups [1239]. superintegrability [2029, 2542, 174]. superintegrable [2319, 329, 1103, 1758, 1283]. Supermatrix [573, 2690]. supermultiplets [632, 1487]. supernonlinear [431]. superoscillating [2570]. Superposition [2666, 2282, 2236, 2662, 2624, 1261, 2205]. superpositions [1743]. superpotentials [2200]. superquadratic [2023]. superspace [184, 211, 376, 845, 2635]. Supersymmetric [751, 2008, 441, 2396, 2012, 1677, 1324, 1989, 2712, 777, 1112, 353, 2337, 135, 1071, 379, 2690, 728, 781, 1436, 1619, 1336, 944, 2096, 2492, 825, 2169, 2091, 1540, 2195, 2227, 1273]. supersymmetries [79]. Supersymmetry [518, 549, 789, 1817, 405, 1324, 2526, 727, 1785, 555, 1538, 112, 1992]. supertori [302]. supertrace [1783]. Supertransvectants [1694]. support [1938, 608, 2618]. supported [1596]. supremum [1049, 1129]. sure [2698]. surely [1847]. surface [1531, 1285, 540, 1470, 822, 1515, 1091, 2467]. surfaces [390, 2598, 1211, 1698, 1632, 492, 885, 1064, 339, 63]. surmises [1481]. survey [2214]. surveys [1173]. susceptibility [1343, 2138]. Suslov [2556, 1566]. Sutherland [1002, 494, 2517, 1603]. swap [2404]. Swift [2616]. switch [269]. switching [1449, 680]. Symanzik [207]. symbol [2269]. Symbolic [1912, 1120]. symbols [2284, 2111]. Symmetric [78, 2137, 164, 2459, 65, 648, 1880, 1092, 126, 325, 564, 852, 1561, 1565, 1357, 2563, 1205, 949, 2346, 1480, 764, 2587, 92, 341, 1903, 2276, 2628, 936, 2382, 2476, 978, 2478, 2076, 388, 136]. Symmetries [249, 654, 1046, 196, 2655, 2389, 1278, 1880, 659, 1163, 2719, 2320, 816, 1706, 505, 2039, 2411, 1193, 2239, 2045, 1868, 1180, 336, 401, 679, 2702, 2718, 368, 2112, 657, 2707, 2278, 614, 1593, 1243, 2714, 2677, 1836, 2325, 1012, 2439, 1570]. Symmetry [1880, 1414, 1160, 1210, 1695, 1339, 1044, 2619, 2319, 2036, 2318, 1206, 846, 564, 712, 2010, 246, 1704, 1717, 1024, 1073, 1205, 2268, 2440, 1756, 624, 1050, 774, 1376, 791, 1171, 865, 1126, 2196, 2262, 2650, 972, 1378, 347, 1138, 23, 705, 2131, 2184, 1563, 1571, 1503, 1170, 2396, 819, 1379, 730]. symmetry-based [2010]. Symplectic [1819, 1820, 1427, 2642, 848, 883, 494, 2564, 2196, 643, 2004, 1080]. synchronization [476, 363, 609]. Synchronous [2705, 2706]. system [1165, 1082, 2049, 46, 2599, 1504, 2320, 2652, 2669, 673, 2310, 406, 411, 1935, 1884, 1338, 1853, 2308, 2497, 2300, 2238, 2517, 2070, 1355, 1015, 1576, 713, 2106, 405, 1009, 1365, 1474, 1668, 2355, 718, 1514, 1171, 523, 1513, 2024, 2264, 324, 235, 268, 1638, 858, 991, 2309, 1333, 1906, 1340, 1911, 27, 806, 1156, 191, 1105, 688, 1064, 459, 1132, 1949, 2668, 1155, 1607, 1616, 1462, 682, 2278, 443, 153, 1792, 2096, 1498, 1842, 2613, 2623, 1515, 2297, 509, 143, 621, 1408, 123, 426, 1482, 721, 2617, 1265, 501, 763, 2686, 1316, 2453, 506, 964, 2063, 2218, 475, 915, 1014, 2103, 2630]. system [1886, 2530, 762, 2220, 424, 677, 758, 886]. system-reservoir [406]. Systematic [804]. systems [2319, 114, 1149, 1301, 905, 1246, 2569, 1966, 240, 1466, 534, 564, 659, 2557, 1805, 1419, 2010, 421, 1620, 1109, 2339, 2451, 2697, 1031, 1569, 2367, 709, 2601, 386, 1356, 1678, 1963, 2407, 334, 504, 505, 2252, 476, 1053, 391, 77, 2134, 33, 1036, 2536, 2153, 2011, 345, 494, 884, 1583, 2352, 706, 1089, 1236, 29, 2336, 1669, 916, 2377, 18, 1083, 1587, 1552, 427, 1066, 2587, 1162, 2433, 2527, 2168, 282, 2702, 42, 707, 1160, 303, 99, 1825, 181, 69, 717, 196, 16, 511, 1294, 1380, 2603, 1421, 1337, 70, 1029, 2151, 1169, 172, 366, 610, 2251]. systems [2414, 1470, 2709, 329, 739, 1103, 1758, 2029, 2001, 1161, 1697, 869, 2366, 2139, 1766, 1746, 5, 110, 524, 2350, 280, 95, 457, 1283, 1593, 1191, 347, 1309, 2649, 2444, 2499, 2586, 346, 2044, 2714, 2432, 1780, 170, 1330, 2520, 2195, 1630, 243, 804, 1970, 1439, 202, 1536, 2224, 1575, 953, 1307, 847, 2119, 156, 1615, 128, 1147, 2439, 285, 1044, 2619, 1525, 1201, 2016, 1733, 1984, 1739, 1446]. Szego [500].

tachyonic [2015]. Tachyons [794]. tails [2448, 2126]. Tamm [2383]. tangent [1292, 1755]. Tanzini [1612, 1198]. tapering [708]. Target [1909, 2665, 933]. Tasso [2436]. Tau [194, 532]. Tau-function [532]. Taylor [1363]. TE [1605, 1975]. technique [612, 583, 1162, 2264, 754, 138]. techniques [1298, 587]. Teitelboim [167, 742]. Teller [1133, 2608, 2116, 2584]. temperature [1023, 2354, 2300, 2366, 1459, 1915, 2054]. temperature-dependent [2300]. Tempered [1940, 2656]. Temperley [2539, 1659]. Templeton [405]. temporal [45]. temporally [2433, 2527]. tension [1531]. Tensor [1662, 1006, 465, 1822, 1771, 1600, 1631, 1178, 2279, 1960, 1835, 2502, 1376, 735, 1133, 1775, 1903, 2476, 2152, 1892, 67, 893]. tensors [2600, 481, 1730, 1065, 2349, 614, 2177]. term [671, 1589, 2065, 956, 2708, 102, 731, 477, 1782, 2398, 2466]. terms [2510, 2273, 660, 2355, 582, 711, 2394, 1888, 2520, 1965, 964, 2576, 2218]. Ternary [195, 491, 198, 12]. tessellation [1481]. test [163, 843]. Testing [159]. Tetrad [167]. Tetrads [6, 2183]. th [296]. their [2549, 2409, 467, 1248, 1163, 2387, 1926, 2320, 777, 1023, 403, 2582, 494, 353, 841, 1684, 632, 1487, 2284, 2604, 904, 1057, 1843, 1247, 1257, 1607, 350, 972, 153, 1377, 2586, 2078, 1928, 2520, 1725, 851, 444, 677, 127]. theorem [80, 1434, 1611, 1626, 43, 755, 1642, 2125, 976, 1582, 901, 1454, 676, 678, 498, 1385, 1030, 666, 1449, 1299, 2305, 1428, 2293, 1669, 488, 604, 30, 1734, 778, 1999, 2504, 234, 2679, 2262, 1687, 1689, 1293, 521, 100, 1235, 2431, 1922, 1824]. theorems [2424, 1861, 1501, 1513, 625]. theoretic [2469, 1680]. theoretical [1347, 267, 1537, 1676, 679, 1716, 2180]. theories [261, 612, 2538, 1095, 272, 17, 263, 1258, 1684, 2487, 1686, 1632, 871, 2690, 1866, 2040, 2470, 987, 435, 320, 2658, 1649, 356, 994, 1128, 1682, 2131, 2184, 1335, 278, 1227, 701, 311]. Theory [1262, 1607, 1613, 524, 1260, 114, 456, 1120, 2611, 2416, 2158, 211, 534, 325, 2073, 2144, 1125, 309, 250, 2006, 2713, 699, 1859, 1685, 437, 431, 2214, 1195, 215, 1357, 727, 2132, 306, 2354, 251, 122, 2274, 259, 738, 32, 1211, 466, 1148, 75, 912, 2336, 1117, 1430, 1769, 1729, 469, 604, 1819, 1820, 352, 609, 368, 31, 1141, 1843, 1857, 2681, 1337, 193, 1470, 657, 34, 1700, 354, 1007, 1274, 556, 936, 1550, 2139, 2199, 62, 2506, 177, 242, 374, 1323, 83, 130, 2095, 37, 614, 1304, 906, 2048, 414, 1226, 2007, 2229, 607]. theory [450, 1130, 1476, 2482, 474, 1279, 1203, 2146, 484, 2400, 1575, 1740, 2618, 1044, 2619, 470, 2019, 1733, 1873, 654]. there [1761]. Thermal [1808, 2300, 604, 83, 130, 2480, 470]. thermal-wave [83]. thermally [661, 1661]. thermo [1936]. thermo-chemo-viscoelastic [1936]. thermocapillary [1692]. Thermodynamic [229, 1419, 24]. thermodynamically [922]. thermodynamics [1691, 2005]. thermoelastic [1309]. thermoelasticity [2024, 497, 1352]. Thermosolutal [821]. thermostats [24, 229]. theta [1885, 2482]. thick [1574, 2327]. Thin [1592, 1652, 2544]. Third [1394, 1283, 152, 652, 1046, 560, 2476, 1842, 2655]. third- [2655]. Third-order [1283, 2476]. Thirring [358, 1411, 1738]. Thomas [837, 1524, 1715, 1470, 1476]. Thomas-forbidden [837]. Thompson [2212, 809, 55]. Three [1625, 2299, 1959, 2632, 1058, 190, 1643, 277, 1013, 2638, 1324, 709, 2403, 1651, 1522, 1657, 2395, 58, 335, 1517, 959, 2605, 924, 181, 173, 711, 1064, 2590, 1353, 1106, 443, 1037, 1527, 2641, 2115, 2077, 1764, 1400, 2455, 1526, 955, 2218, 2375]. three-algebras [924]. three-body [2395, 2641, 955]. Three-by-three [1058]. three-component [1657]. three-dimension [2605]. Three-dimensional [1625, 277, 1013, 1651, 1522, 1517, 959, 711, 1353, 1037, 1527, 2077, 1526, 2218, 2375]. three-dimensional-root [181]. three-dimensionality [2590]. three-fermion [2403]. Three-manifold [1959]. three-sphere [1106]. three-wave [1064]. threefolds [290]. threesome [356]. threshold [1011, 937, 673, 1442, 1757]. thresholds [509, 143]. Thue [1189]. tight [645, 973]. tight-binding [645]. tiling [1125]. Time [1621, 548, 635, 345, 932, 25, 2587, 1955, 1253, 2720, 2466, 2612, 474, 114, 824, 2510, 66, 1466, 1589, 1280, 1221, 990, 2339, 2451, 2697, 309, 2415, 1392, 2043, 157, 1531, 141, 2136, 1877, 2518, 1475, 1131, 729, 1823, 1512, 611, 2268, 2337, 1005, 2346, 218, 2255, 1686, 2377, 2653, 1628, 1302, 2559, 992, 2514, 2702, 1513, 1938, 1368, 76, 1796, 798, 2550, 487, 675, 399, 1782, 2262, 9, 350, 754, 563, 161, 822, 2278, 2233, 347, 1624, 1309, 1515, 23, 919, 1671, 2714, 1249, 2431, 367, 1473, 1586, 1182, 1085, 1026, 2686, 260, 1080, 2283, 1824, 470, 1570]. time [436]. Time-dependent [1253, 2720, 2466, 474, 1221, 309, 157, 141, 729, 1005, 2559, 1513, 76, 1796, 2262, 9, 1624, 2431, 1473, 1085, 2686]. time-domain [1080]. Time-evolution [548, 635]. time-increasing [1466]. time-ordered [2346]. time-periodic [114, 2136]. Time-reversal [2587]. time-space [2514]. time-varying [1782]. times [65, 90, 1100, 4, 2295, 64, 2276, 2390, 1007, 1274, 557, 280, 1186, 67, 416, 543, 1558]. Timoshenko [2377, 2024, 964]. Timoshenko-type [2024]. TM [1605]. Toda [2202, 1706, 1704, 1214, 1587, 2493, 1368, 2114, 562, 194, 1126, 156]. Toeplitz [2269]. Tomasevich [786]. tomography [15, 2336, 2230, 397, 2159]. tool [466]. tools [1120]. Topics [909]. topoi [1895]. Topological [442, 2277, 1326, 1189, 2659, 1320, 1631, 2348, 1593, 2417, 994, 1128]. Topologically [2280, 750]. Topologies [1895]. topology [1386, 1359, 1027, 2343, 1536]. Topos [2506, 2400, 259, 1722]. tori [1348, 2261, 1820, 1728, 2701, 366, 2414, 1146]. toric [720]. torics [519]. torsion [2306, 1866, 2040, 1540, 1025]. torsionless [1863]. torus [877, 1660, 2067, 859, 1153]. Total [1146, 1143]. totally [2172]. Trace [2181, 2607, 1471, 1480, 1738, 576, 1578, 1152, 233, 1091, 1872, 2426, 140, 776, 767, 132]. traces [778]. tracial [398]. tractability [2, 109]. traffic [1890]. trajectories [2339, 2451, 1623, 1542, 1690, 2358, 2645]. transcendents [1973]. transfer [1667, 1087, 1870, 619, 1785, 842, 1358, 1077]. transform [214, 2219, 2500, 1281, 1662, 2489, 2541, 495, 1270, 1981, 500, 718, 609, 1787, 88, 351, 2485, 2617, 501]. transformation [238, 1489, 1545, 2620, 148, 449, 2317, 814, 441, 1132, 1799, 2566, 629, 147, 752, 333]. Transformations [408, 1301, 747, 344, 146, 2311, 135, 1736, 1587, 1456, 532, 807, 1754, 527, 233, 155, 927, 1493, 1383, 1273, 649, 2683, 1400, 1644, 2289, 716, 1041, 886]. transforms [2357, 2567, 201, 2504, 95, 2061]. Transition [348, 693, 81, 2172, 617, 2516, 1122]. transitions [1149, 277, 2654, 2243, 1421, 1169, 715, 978, 280]. transitive [1050]. Translation [1275, 23, 1227, 1899]. translation-invariant [1227, 1899]. Translational [1361]. Translationally [1705, 2191, 2293, 42]. transmission [89, 1659, 544, 530, 2070, 1007, 1274]. transonic [1672]. Transport [1544, 1415, 569, 863, 230, 2073, 2144, 879, 1013, 372, 1209, 2190, 477, 1305, 2327, 2480]. transportation [2473]. transpose [163, 1059, 1452, 1375, 1768, 1257]. transposed [2231]. transvector [2503]. transverse [1784, 1774]. transverse-electric-transverse [1784, 1774]. trap [952]. trapped [1211]. trapping [1620, 2516]. traps [2312, 1413]. Traveling [2452, 171, 1937, 2378]. travelling [2277, 2627]. treatment [926, 1259, 517, 879, 180, 2316, 369]. Tree [309, 1415, 2462, 1673, 541, 2292, 1873]. trees [1707, 1034]. Tremblay [2599]. trend [1710]. Trends [1047]. Tri [1791]. Tri-Hamiltonian [1791]. triad [608]. triangulations [1870]. Tridiagonal [428, 734]. Trigonometric [494, 2517, 1675, 1650]. triple [2561, 2016]. triplectic [1582]. triplets [514]. trivial [2528, 1408]. Trkalian [147, 2061]. Truax [580]. truncated [597]. truncation [1108]. truth [1175]. Tsallis [1370, 1471, 584, 999]. Tsirelson [602, 1724]. Tu [378]. tube [1497, 2379]. Tunneling [1620]. Turaev [2535]. Turbiner [2599]. Turbulence [1509, 2009, 1512, 1523, 1194]. turbulent [800]. turns [22]. Twenty [32]. Twist [513, 689, 828, 2594]. Twisted [2110, 974, 2169, 1079, 1154, 1802, 462, 385, 101, 1892, 1503, 790]. twisting [557, 1144]. twistor [90, 1902, 1644]. twistors [342, 1581, 1649, 2061]. twists [254]. Two [1467, 403, 1690, 835, 1741, 2490, 2494, 92, 341, 1961, 71, 594, 242, 2178, 1383, 2056, 410, 522, 1204, 114, 1818, 2510, 2522, 577, 1002, 1165, 1423, 1795, 2625, 712, 1395, 421, 2525, 2339, 2451, 2505, 878, 1637, 49, 642, 1497, 1059, 1452, 1375, 1504, 1672, 1706, 2671, 412, 2458, 1614, 631, 2395, 2507, 1469, 335, 1271, 1142, 1512, 884, 750, 1005, 1532, 733, 203, 1321, 2344, 32, 571, 2665, 1015, 336, 959, 1010, 1576, 2100, 1809, 2250, 2604, 718, 1171, 1635, 1275, 1486, 1918, 1734, 2038, 117, 717, 991, 106, 27, 2087, 2140, 2151, 813, 2394, 2029, 1839]. two [2668, 1481, 1613, 1716, 1867, 1593, 527, 153, 370, 2007, 2641, 927, 1851, 621, 426, 721, 2138, 134, 278, 1558, 1571, 745, 506, 286, 1711, 1170, 2218, 2479, 2532, 2477, 1972, 2031, 875, 1564, 937]. two-band [878]. two-body [114, 1867, 1571]. Two-by-two [242]. Two-center [835]. Two-component [2178, 1637, 1504, 884, 1015, 1576, 717, 2087, 2007, 621, 721, 286, 1711, 2477]. Two-dimensional [1467, 403, 92, 341, 1383, 522, 2522, 577, 1165, 1423, 1395, 2505, 49, 642, 1497, 1672, 2671, 1271, 1005, 1532, 32, 2665, 336, 1809, 2250, 1171, 1734, 2038, 27, 813, 2029, 1481, 1716, 1593, 370, 2138, 1558, 745, 875, 1564, 937]. two-dimensions [2479, 2532]. two-electron [2394]. two-fluid [1010]. Two-fold [71]. two-layer [712, 1512]. two-level [421, 2339, 2451, 733]. two-loop [1818, 2510]. two-matrix [1795, 631]. two-mode [2507, 2641, 2031]. Two-parameter [594, 2458, 571]. two-phase [959, 2604]. two-photon [2507]. Two-point [1741, 2344]. two-qutrit [1452]. Two-sided [410]. two-Toda [1706]. type [1696, 1807, 2537, 1752, 2717, 861, 782, 1539, 289, 510, 451, 848, 2089, 2531, 2719, 1356, 2252, 777, 1601, 471, 1193, 1939, 884, 1032, 2456, 298, 2154, 969, 488, 1861, 1978, 2608, 1501, 968, 2718, 2024, 871, 408, 1266, 1787, 1380, 1126, 2087, 1790, 2107, 1876, 1135, 1251, 2542, 1155, 497, 1144, 1616, 2707, 212, 2118, 1851, 705, 843, 509, 143, 2423, 2221, 2195, 1816, 1836, 2325, 2675, 286, 1216, 1307, 2692, 1943, 1941, 1464, 2464, 2577, 667, 2584, 1992, 709, 546, 1162, 69, 93, 801]. types [1002, 2331, 1841, 1462, 994, 1128]. typicality [350]. Tyutin [649].

U [464]. U-statistics [464]. Uhrig [1551]. ultra [2149]. ultra-short [2149]. ultracold [1804]. ultradiscrete [2493, 1161, 1697]. ultrametric [2089, 2531]. ultrarelativistic [2084]. Ultrastatic [416]. ultraviolet [1680, 429]. Umbral [1977]. unambiguous [117]. unbiased [2593, 2622, 315, 1546, 1284, 84]. unbounded [1042, 1585, 1241, 2654, 293, 2346, 1513, 1965, 1824, 1564]. Uncertainties [1619]. Uncertainty [1246, 869, 1630, 2271, 1591, 2272, 2385, 1602, 2057, 2592, 2399, 351, 1094, 1624]. Unchanged [2645]. unconfined [945]. unconstrained [2297]. Unconventional [2006, 2713]. underactuated [391]. underlying [607]. Understanding [823, 724]. Unextendible [2407, 1058]. Unfinished [1535]. unidirectional [2241]. unification [51]. Unified [1218, 1113, 374, 1191, 65, 625]. uniform [2676, 1779, 460, 2300, 854, 1105, 1753, 1309, 1397, 1393, 667]. uniformly [224]. unimodular [2385]. Unique [209, 2454]. Uniqueness [1072, 922, 239, 1963, 2341, 1404, 2298, 2668, 1360, 426, 2692, 2210, 1361]. unit [1441, 1732, 427, 1882, 1825, 977]. unital [210]. Unitary [606, 507, 574, 91, 344, 591, 2259, 1584, 1108, 353, 2330, 2154, 2230, 1679, 2143, 1995, 732, 1724, 168, 384, 2091, 1110, 317, 1571, 2675, 961]. unity [2587]. Universal [1726, 1489, 544, 430, 384, 80, 216, 1440, 269, 1941]. Universality [2208, 1572, 1551, 961, 1431, 2695, 34, 499, 2287]. universe [1864]. universes [626]. unraveled [361]. unravelling [515]. Unruh [1090, 1387]. Upper [2509, 833, 641, 395, 68]. Upscaling [1312, 1971]. use [2284, 817]. used [1530]. uses [793]. using [861, 163, 772, 1545, 432, 406, 2503, 588, 2059, 1162, 281, 858, 605, 584, 2650, 1829, 410, 1924, 1946, 1650]. usual [2314]. Uvarov [1133, 1056, 1650].

vacuum [325, 473, 1863, 107, 959, 1534, 1727, 1688, 1144, 1770, 142]. Vafa [911]. Vakhnenko [1579, 2324]. Valchev [915]. valence [840]. value [1250, 1248, 2232, 1050, 1468, 1207, 2068, 2309, 93, 1521, 870]. valued [1224, 2489, 349, 27, 164, 2661]. values [1246, 2437, 1175, 307, 350, 1116, 1972]. vanishing [2651, 1498, 1048, 1526]. Variable [972, 2164, 1994, 1497, 251, 1331, 523, 1608, 1135, 1251, 191, 192, 1016, 1695, 890, 1658]. variable-coefficient [191, 890, 1658]. variable-coefficients [1608]. variable-length [1994]. variables [896, 1935, 1884, 1224, 2493, 30, 2340, 1161, 1697, 1846, 2072, 2287, 45]. Variance [1628]. Variational [2239, 1731, 2007, 683, 2158, 2013, 811, 2011, 1174, 1868, 1786, 1829, 2426]. variations [1745, 137, 2066]. varieties [1450]. varying [1589, 2079, 2588, 2408, 1051, 2433, 2527, 2017, 1782, 1244]. Vector [358, 381, 118, 290, 900, 2344, 2489, 2120, 2330, 1730, 2555, 2487, 1133, 1982, 2145, 1328, 1378, 1200, 2435, 1459, 2004, 138, 757, 468]. vector-valued [2489]. vectors [1463, 1057, 1892]. velocities [718]. Velocity [991, 1996, 2030, 711, 2578, 1439, 2218]. Verblunsky [1244]. Verma [60, 1892]. versatile [466]. version [2265, 2577, 670]. versus [2417, 1563, 1907, 361]. vertex [1354, 2110, 1396, 119, 1489, 2632, 1801, 1401, 2156, 1803, 414, 1226]. vertex-IRF [1489]. vertical [2480]. vertices [1596]. very [2544]. Veselov [954, 1743]. vessels [177]. VI [1145, 1462, 1351]. via [926, 2651, 1298, 2141, 947, 504, 727, 1414, 403, 162, 495, 1270, 1870, 379, 720, 560, 1857, 946, 354, 2026, 680, 519, 2578, 804, 616, 918, 1933]. vibrating [137]. vibrations [477]. view [685, 1097]. viewpoint [2288]. VIII [418]. Vilkovisky [35]. Villani [393]. violating [1554]. violation [846, 310]. violations [1135, 1251]. Virasoro [1612, 2537, 2637, 2614, 1198, 1344, 597, 966, 1790, 2673, 101, 1892, 705, 2157, 1833]. Virasoro-like [966]. virial [977]. viscoelastic [1936, 338, 1782, 1277, 423]. viscoelasticity [1268, 620]. viscosity [2318, 1497, 2518, 1891, 2300, 1526, 760]. viscous [1338, 959, 660, 1652, 1609, 800, 252, 522, 2375]. Vlasov [1530, 1082, 2669, 1032, 1577, 713, 2106, 1474, 1911, 176, 1462, 977, 2630]. Vlasov-type [1032]. Voigt [1936]. Volterra [744, 153, 773, 677]. volume [1721, 1857, 1114]. Voronoi [1481]. Vortex [2629, 2080, 1522, 1107, 925, 996, 2571, 1406, 61, 369, 502, 985]. vortexlike [425]. vortices [877, 437, 39, 1698, 1869, 967, 173, 761, 2515, 1781, 371]. vorticity [2252, 934]. vorticity-type [2252]. Vries [1372, 2670, 383, 808, 2320, 2541, 298, 2693, 2452, 392, 379, 1485, 436, 2524, 2498, 2096, 1273, 1012, 1111, 190, 145, 333]. vs [2155].

Waals [1425]. Wadati [1791]. Wagner [1734]. Wake [708]. walk [399, 133]. Walker [275, 1683, 1680]. walks [736, 982, 2089, 2531, 663, 2207, 2421, 1429, 2425, 2591, 1570]. wall [81, 2210]. Wallis [2640]. Wang [763]. Wannier [2721, 1018, 36, 2663, 2585]. Ward [2113]. warped [559, 1560, 968, 2699, 2177]. Watanabe [1669]. water [1531, 2620, 660, 449, 721]. Wave [829, 852, 2084, 157, 623, 916, 764, 235, 1952, 9, 2216, 684, 538, 2686, 2477, 2629, 2028, 1667, 1278, 1880, 579, 1589, 1590, 894, 1248, 820, 626, 90, 2307, 1504, 2277, 1595, 104, 403, 898, 2558, 639, 439, 218, 2255, 2452, 2627, 1832, 1775, 1207, 1796, 1188, 1932, 728, 1173, 1712, 487, 675, 171, 477, 2247, 1782, 1713, 1291, 294, 1064, 1937, 568, 1277, 2540, 2072, 83, 1325, 443, 680, 1072, 2430, 2161, 8, 2326, 252, 1605, 1784, 1975, 367, 1418, 1774, 429, 1121, 213, 1405, 2378, 1306, 2577]. wave-diffusion [1595]. wave-packets [2161]. wave-particle [626]. wavefunction [2119]. waveguide [1079, 1154, 2543, 1784]. waveguides [1592, 1604, 2582, 2250, 892]. wavelet [2357, 609, 382]. wavelet-like [382]. waves [1643, 73, 1208, 1531, 2069, 2277, 1264, 520, 439, 1646, 2490, 2620, 1065, 1407, 1310, 718, 449, 1656, 2281, 1500, 814, 914, 1713, 1688, 1115, 2376, 1515, 1851, 247]. waves-short [814]. way [2408, 1941]. Weak [1042, 1585, 1798, 1756, 962, 141, 2136, 1015, 94, 2100, 1325, 340, 1115, 801, 2535, 311]. weak-coupling [311]. weakly [1589, 2022, 2693, 598, 1009, 2002, 1358]. webs [810]. Wedge [1535, 1389, 1755]. Wedge-local [1389]. Wegner [2448, 1721]. Weierstrass [1937]. Weight [597, 627, 411, 2148, 732, 1244, 1892]. weighted [184, 1940, 1475, 1101]. weights [2292, 961]. Weitzenböck [611, 1641]. Weizsäcker [1524]. Well [1338, 660, 822, 1515, 113, 2302, 671, 1894, 1435, 2377, 2162, 1533, 1629, 2399, 1482, 1127, 2686, 1216, 1014, 2464, 2700, 424, 1534]. Well-posedness [1338, 660, 822, 1515, 2377, 1533, 1216, 1014, 424]. wells [772, 2465, 2163, 2294]. Wentzel [2441]. Wenzl [196]. Wess [358, 2345, 2488, 2492]. Weyl [2611, 1707, 1822, 1771, 2682, 2567, 2150, 1052, 2083, 1974, 2568, 735, 594, 2634, 2396]. Weyl-ordered [1052]. which [2552]. white [482, 2179, 582]. Whitham [1837]. Whittaker [1612, 1198, 966, 313, 101, 396]. whole [1393]. whose [1639, 67]. Wick [1611, 1434, 1686, 1229]. width [1154]. Wiedemann [1284]. Wigert [1473]. Wigner [2208, 1477, 2678, 2053, 2473, 455, 1481, 1763, 590, 1037, 1255, 1737]. Wigner-like [1255]. Wilson [1458]. Winternitz [2599, 739]. wireless [1877]. Wishart [1980, 493, 1972]. within [1445, 605, 1858, 565, 1438, 2396, 2516, 1254]. without [830, 1620, 1633, 1541, 2587, 821, 1828, 1627]. witnesses [1451]. Witt [1878, 1344, 2112, 2062, 396]. Witten [1685, 1802, 911, 1065, 2345, 1436, 1876, 2488, 2492]. WKB [2668]. Woods [1223, 118, 1020]. Work [1535, 1220, 2085]. world [2659, 363, 2168]. world-line [2659]. wormhole [741]. Wronskian [2490, 138].

xo [2155]. xo-algebras [2155]. XOR [983]. Xue [2317]. XXX [1846]. XXZ [1881]. XY [2676].

Yamada [1669, 1155]. Yang [126, 1565, 2223, 881, 803, 856, 211, 2280, 652, 1428, 2275, 1314, 1315, 707, 871, 728, 781, 1436, 692, 1421, 2501, 1132, 944, 1716, 2186, 320]. Yangian [1388, 1674]. Yano [65, 1730]. years [32]. Yetter [2389, 2538, 2223, 2222, 2173]. YM [2169]. Young [2187]. Yukawa [943, 713].

Zaboronsky [2564]. Zakharov [46, 1355, 1332, 2441, 235, 858, 1521, 762]. Zamolodchikov [388]. Zanelli [742]. Zeeman [1391]. Zeldovich [1746]. Zeno [44]. Zero [1459, 1018, 2721, 948, 2172, 2203, 1139, 2366, 2348, 414, 1343, 977]. zero-field [1343]. zero-range [2348]. zeros [2648, 1543, 1666, 1421]. Zeta [668, 1594, 1543, 1666, 1035, 2699, 1116, 1116]. zeta-determinants [2699]. zeta-values [1116]. Zetlin [453]. Zhu [790]. zigzag [501]. zigzagrunged [501]. Zimmermann [2713, 2006]. Zumino [2345, 2488, 2492]. Zwanziger [1267].

References

Erdahl:1972:CSS

[1] R. M. Erdahl. The convex structure of the set of N-representable reduced 2-matrices. Journal of Mathematical Physics, 13(10):1608–1621, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

      tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i10/p1608_s1. See comment [1322].

Wasilkowski:1996:TPI

[2] Grzegorz W. Wasilkowski and Henryk Woźniakowski. On tractability of path integration. Journal of Mathematical Physics, 37(4):2071–2086, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [109].

Hall:2002:CSS

[3] Brian C. Hall and Jeffrey J. Mitchell. Coherent states on spheres. Journal of Mathematical Physics, 43(3):1211–1236, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [942].

Haesen:2004:IES

[4] Stefan Haesen and Leopold Verstraelen. Ideally embedded space-times. Journal of Mathematical Physics, 45(4):1497–1510, April 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i4/p1497_

s1. See erratum [2295].

Niestegge:2004:CSR

[5] Gerd Niestegge. Composite systems and the role of the complex numbers in quantum mechanics. Journal of Mathematical Physics, 45(12):4714– 4725, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4714_s1. See erratum [110].

Garat:2005:TG

[6] Alcides Garat. Tetrads in geometrodynamics. Journal of Mathematical Physics, 46(10):102502, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p102502_s1. See [2183].

Bisi:2005:FDE

[7] M. Bisi and G. Spiga. Fluid-dynamic equations for granular particles in a host medium. Journal of Mathematical Physics, 46(11):113301, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113301_s1. See erratum [726].

Schwartz:2006:RQW

[8] Charles Schwartz. Relativistic quaternionic wave equation. Journal of Mathematical Physics, 47(12):122301, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p122301_s1. See

erratum [2326].

Merad:2007:WFD

[9] M. Merad and S. Bensaid. Wave functions for a Duffin–Kemmer– Petiau particle in a time-dependent potential. Journal of Mathematical Physics, 48(7):073515, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073515_s1. See comment [157].

Wang:2008:GNQ

[10] Hai-Jun Wang. General nonlocality in quantum fields. Journal of Mathematical Physics, 49(3):033513, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i3/p033513_s1. See comment [262]

and erratum [2621].

Dokovic:2009:PIS

[11] D. Ž. – Doković and A. Osterloh. On polynomial invariants of several qubits. Journal of Mathematical Physics, 50(3):033509, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p033509_s1.

Savadkouhi:2009:ATJ

[12] M. Bavand Savadkouhi, M. Eshaghi Gordji, J. M. Rassias, and N. Ghobadipour. Approximate ternary Jordan derivations on Banach ternary algebras. Journal of Mathematical Physics, 50(4):042303, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042303_s1. See comment [198].

Yang:2009:SSS

[13] Yisong Yang and Ruifeng Zhang. Steady state solutions for nonlinear Schrödinger equation arising in optics. Journal of Mathematical Physics, 50(5):053501, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p053501_s1. See erratum [200].

Qian:2009:IRR

[14] Min-Ping Qian and Jian-Sheng Xie. On the imaginary-real ratio rule of power spectra. Journal of Mathematical Physics, 50(6):063301, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p063301_s1. See comment [197].

Fan:2009:EPS

[15] Hong yi Fan and Cui hong Lv. Explicit pure-state density operator structure for quantum tomography. Journal of Mathematical Physics, 50(10): 102108, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102108_s1. See [397].

Kundu:2009:NLE

[16] Anjan Kundu. Nonlinearizing linear equations to integrable systems including new hierarchies with nonholonomic deformations. Journal of Mathematical Physics, 50(10):102702, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p102702_s1. See

publisher’s note [511].

DiCriscienzo:2009:FDH

[17] Roberto Di Criscienzo and Sergio Zerbini. Functional determinants in higher derivative Lagrangian theories. Journal of Mathematical Physics, 50(10):103517, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p103517_s1. See comment [261] and er-

ratum [263].

Gurses:2009:DSP

[18] Metin Gürses, Gusein Sh. Guseinov, and Kostyantyn Zheltukhin. Dynamical systems and Poisson structures. Journal of Mathematical Physics, 50(11):112703, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p112703_s1. See comment [1083].

Wada:2009:GLL

[19] Tatsuaki Wada. Generalized log-likelihood functions and Bregman divergences. Journal of Mathematical Physics, 50(11):113301, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p113301_s1. See erratum [199].

El-Jaick:2009:LCH

[20] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. A limit of the confluent Heun equation and the Schrödinger equation for an inverted potential and for an electric dipole. Journal of Mathematical Physics, 50(12):123511, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p123511_s1. See erratum [775].

Li:2009:FOD

[21] Junjie Li. On a fourth order degenerate parabolic equation in higher space dimensions. Journal of Mathematical Physics, 50(12):123524, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123524_s1. See erratum [1373].

Nachtergaele:2010:ISI

[22] Bruno L. Z. Nachtergaele. Introduction to special issue: Journal of Mathematical Physics turns 50. Journal of Mathematical Physics, 51(1): 015101, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i1/p015101_s1.

Schwarz:2010:STT

[23] A. Schwarz. Space and time from translation symmetry. Journal of Mathematical Physics, 51(1):015201, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i1/p015201_s1.

Gallavotti:2010:NTT

[24] G. Gallavotti and E. Presutti. Nonequilibrium, thermostats, and thermodynamic limit. Journal of Mathematical Physics, 51(1):015202, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i1/p015202_s1.

Gutschow:2010:TAE

[25] Johannes Gütschow, Sonja Uphoff, Reinhard F. Werner, and Zoltán Zimborás. Time asymptotics and entanglement generation of Clifford quantum cellular automata. Journal of Mathematical Physics, 51(1): 015203, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i1/p015203_s1.

Mouhot:2010:LD

[26] C. Mouhot and C. Villani. Landau damping. Journal of Mathematical Physics, 51(1):015204, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i1/p015204_s1.

Li:2010:SCV

[27] Dong Li and Yakov G. Sinai. Singularities of complex-valued solutions of the two-dimensional Burgers system. Journal of Mathematical Physics, 51(1):015205, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i1/p015205_s1.

Kiessling:2010:HSF

[28] M. K.-H. Kiessling and J. K. Percus. Hard-sphere fluids with chemical self-potentials. Journal of Mathematical Physics, 51(1):015206, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i1/p015206_s1.

Gentile:2010:QMD

[29] Guido Gentile. Quasiperiodic motions in dynamical systems: Review of a renormalization group approach. Journal of Mathematical Physics, 51(1): 015207, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i1/p015207_s1.

Jaksic:2010:QCL

[30] V. Jaksić, Y. Pautrat, and C.-A. Pillet. A quantum central limit theorem for sums of independent identically distributed random variables. Journal of Mathematical Physics, 51(1):015208, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i1/p015208_s1.

Kawahigashi:2010:OAS

[31] Yasuyuki Kawahigashi. From operator algebras to superconformal field theory. Journal of Mathematical Physics, 51(1):015209, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i1/
p015209_s1.

Fuchs:2010:TFY

[32] Jürgen Fuchs, Ingo Runkel, and Christoph Schweigert. Twenty five years of two-dimensional rational conformal field theory. Journal of Mathematical Physics, 51(1):015210, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i1/p015210_s1.

Costin:2010:ERI

[33] O. Costin, J. L. Lebowitz, C. Stucchio, and S. Tanveer. Exact results for ionization of model atomic systems. Journal of Mathematical Physics, 51(1):015211, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i1/p015211_s1.

Mastropietro:2010:DTU

[34] Vieri Mastropietro. Developments in the theory of universality. Journal of Mathematical Physics, 51(1):015212, January 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i1/p015212_s1.

Albert:2010:BVI

[35] C. Albert, B. Bleile, and J. Fröhlich. Batalin–Vilkovisky integrals in finite dimensions. Journal of Mathematical Physics, 51(1):015213, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i1/p015213_s1.

Hastings:2010:ACM

[36] Matthew B. Hastings and Terry A. Loring. Almost commuting matrices, localized Wannier functions, and the quantum Hall effect. Journal of Mathematical Physics, 51(1):015214, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i1/p015214_s1.

Petz:2010:ACQ

[37] Dénes Petz. Algebraic complementarity in quantum theory. Journal of Mathematical Physics, 51(1):015215, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i1/p015215_s1.

Matsui:2010:SGS

[38] Taku Matsui. Spectral gap, and split property in quantum spin chains. Journal of Mathematical Physics, 51(1):015216, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i1/p015216_s1.

Gustafson:2010:SDM

[39] S. Gustafson, I. M. Sigal, and T. Tzaneteas. Statics and dynamics of magnetic vortices and of Nielsen–Olesen (Nambu) strings. Journal of Mathematical Physics, 51(1):015217, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i1/p015217_s1.

Doplicher:2010:PLE

[40] Sergio Doplicher. The principle of locality: Effectiveness, fate, and challenges. Journal of Mathematical Physics, 51(1):015218, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i1/
p015218_s1.

Maes:2010:RMM

[41] Christian Maes and Karel Netocný. Rigorous meaning of McLennan ensembles. Journal of Mathematical Physics, 51(1):015219, January 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i1/p015219_s1.

Irani:2010:GSE

[42] Sandy Irani. Ground state entanglement in one-dimensional translationally invariant quantum systems. Journal of Mathematical Physics, 51(2):022101, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022101_s1.

Aubrun:2010:NRE

[43] Guillaume Aubrun, Stanislaw Szarek, and Elisabeth Werner. Nonadditivity of Rényi entropy and Dvoretzky’s theorem. Journal of Mathematical Physics, 51(2):022102, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p022102_s1.

Facchi:2010:QZE

[44] Paolo Facchi and Marilena Ligabò. Quantum Zeno effect and dynamics. Journal of Mathematical Physics, 51(2):022103, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/p022103_s1.

Strauss:2010:SAL

[45] Y. Strauss. Self-adjoint Lyapunov variables, temporal ordering, and irreversible representations of Schrödinger evolution. Journal of Mathematical Physics, 51(2):022104, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p022104_s1.

Boonserm:2010:RSE

[46] Petarpa Boonserm and Matt Visser. Reformulating the Schrödinger equation as a Shabat–Zakharov system. Journal of Mathematical Physics, 51(2):022105, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022105_s1.

Gu:2010:SSH

[47] Xiao-Yan Gu and Jian-Qiang Sun. Any l-state solutions of the Hulthén potential in arbitrary dimensions. Journal of Mathematical Physics, 51(2):022106, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022106_s1.

Hall:2010:SCC

[48] Richard L. Hall, Nasser Saad, and K. D. Sen. Soft-core Coulomb potentials and Heun’s differential equation. Journal of Mathematical Physics, 51(2):022107, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022107_s1.

Cannata:2010:ESN

[49] F. Cannata, M. V. Ioffe, and D. N. Nishnianidze. Exactly solvable nonseparable and nondiagonalizable two-dimensional model with quadratic complex interaction. Journal of Mathematical Physics, 51(2):022108, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p022108_s1.

Midya:2010:PDM

[50] Bikashkali Midya, B. Roy, and R. Roychoudhury. Position dependent mass Schrödinger equation and isospectral potentials: Intertwining operator approach. Journal of Mathematical Physics, 51(2):022109, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p022109_s1.

Kryukov:2010:PMU

[51] A. Kryukov. A possible mathematics for the unification of quantum mechanics and general relativity. Journal of Mathematical Physics, 51(2):022110, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022110_s1.

Chiribella:2010:BDQ

[52] Giulio Chiribella, Giacomo Mauro D’Ariano, and Dirk Schlingemann. Barycentric decomposition of quantum measurements in finite dimensions. Journal of Mathematical Physics, 51(2):022111, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p022111_s1.

Mohlenkamp:2010:CMP

[53] Martin J. Mohlenkamp. A center-of-mass principle for the multiparticle Schrödinger equation. Journal of Mathematical Physics, 51(2):022112, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p022112_s1.

Accardi:2010:QFF

[54] Luigi Accardi and Ameur Dhahri. The quadratic Fock functor. Journal of Mathematical Physics, 51(2):022113, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i2/p022113_s1.

Palmai:2010:NPC

[55] Tamás Pálmai and Barnabás Apagyi. On nonsingular potentials of Cox–Thompson inversion scheme. Journal of Mathematical Physics, 51(2):022114, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022114_s1.

Fukuda:2010:ADM

[56] Motohisa Fukuda. An application of decomposable maps in proving multiplicativity of low dimensional maps. Journal of Mathematical Physics, 51(2):022201, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022201_s1. See publisher’s note [111].

Hwang:2010:MES

[57] Won-Young Hwang and Joonwoo Bae. Minimum-error state discrimination constrained by the no-signaling principle. Journal of Mathematical Physics, 51(2):022202, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022202_s1.

Dolan:2010:SCP

[58] F. A. Dolan. On superconformal characters and partition functions in three dimensions. Journal of Mathematical Physics, 51(2):022301, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p022301_s1.

Degner:2010:CPC

[59] Andreas Degner and Rainer Verch. Cosmological particle creation in states of low energy. Journal of Mathematical Physics, 51(2):022302, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p022302_s1.

Jiang:2010:SVM

[60] Wei Jiang and Yufeng Pei. On the structure of Verma modules over the W -algebra W (2, 2). Journal of Mathematical Physics, 51(2):022303, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p022303_s1.

Krusch:2010:EMS

[61] S. Krusch and J. M. Speight. Exact moduli space metrics for hyperbolic vortex polygons. Journal of Mathematical Physics, 51(2):022304, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p022304_s1.

Nagamachi:2010:FIF

[62] S. Nagamachi and E. Brüning. Frame independence of the fundamental length in relativistic quantum field theory. Journal of Mathematical Physics, 51(2):022305, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022305_s1.

Nishinou:2010:CAF

[63] Takeo Nishinou. Convergence of adiabatic family of anti-self-dual connections on products of Riemann surfaces. Journal of Mathematical Physics, 51(2):022306, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022306_s1.

Klein:2010:EFC

[64] David Klein and Peter Collas. Exact Fermi coordinates for a class of space–times. Journal of Mathematical Physics, 51(2):022501, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p022501_s1.

Acik:2010:KYF

[65] Ö. Açik, Ü. Ertem, M. Önder, and A. Verçin. Killing–Yano forms of a class of spherically symmetric space–times: a unified generation of higher forms. Journal of Mathematical Physics, 51(2):022502, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p022502_s1.

Arraut:2010:MES

[66] I. Arraut, D. Batic, and M. Nowakowski. Maximal extension of the Schwarzschild space–time inspired by noncommutative geometry. Journal of Mathematical Physics, 51(2):022503, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/p022503_s1.

Sharma:2010:PFS

[67] Ramesh Sharma and Amalendu Ghosh. Perfect fluid space–times whose energy-momentum tensor is conformal Killing. Journal of Mathematical Physics, 51(2):022504, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022504_s1.

Wang:2010:USP

[68] Yonghai Wang and Yuming Qin. Upper semicontinuity of pullback attractors for nonclassical diffusion equations. Journal of Mathematical Physics, 51(2):022701, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022701_s1.

Kostov:2010:ESC

[69] N. A. Kostov, V. S. Gerdjikov, and V. Mioc. Exact solutions for a class of integrable Hénon–Heiles-type systems. Journal of Mathematical Physics, 51(2):022702, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022702_s1.

Li:2010:NFF

[70] Wenlei Li, Zhiguo Xu, and Shaoyun Shi. Nonexistence of formal first integrals for nonlinear systems under the case of resonance. Journal of Mathematical Physics, 51(2):022703, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i2/p022703_s1.

Kundu:2010:TFI

[71] Anjan Kundu. Two-fold integrable hierarchy of nonholonomic deformation of the derivative nonlinear Schrödinger and the Lenells–Fokas equation. Journal of Mathematical Physics, 51(2):022901, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p022901_s1.

Langerock:2010:RRQ

[72] B. Langerock, F. Cantrijn, and J. Vankerschaver. Routhian reduction for quasi-invariant Lagrangians. Journal of Mathematical Physics, 51(2):022902, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022902_s1.

Bal:2010:SSE

[73] Guillaume Bal, Ian Langmore, and Olivier Pinaud. Single scattering estimates for the scintillation function of waves in random media. Journal of Mathematical Physics, 51(2):022903, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p022903_s1.

Peyghan:2010:KSF

[74] E. Peyghan and A. Tayebi. A Kähler structure on Finsler spaces with nonzero constant flag curvature. Journal of Mathematical Physics, 51(2):022904, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p022904_s1.

Gitman:2010:GIC

[75] D. M. Gitman and V. G. Kupriyanov. Gauge invariance and classical dynamics of noncommutative particle theory. Journal of Mathematical Physics, 51(2):022905, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p022905_s1.

Kang:2010:LRC

[76] Yan-Mei Kang and Yao-Lin Jiang. Linear response characteristics in time-dependent subdiffusive fractional Fokker–Planck equations. Journal of Mathematical Physics, 51(2):023301, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023301_s1.

Contucci:2010:CIQ

[77] Pierluigi Contucci and Joel L. Lebowitz. Correlation inequalities for quantum spin systems with quenched centered disorder. Journal of Mathematical Physics, 51(2):023302, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p023302_s1.

Ayadi:2010:SNS

[78] Imen Ayadi and Saïd Benayadi. Symmetric Novikov superalgebras. Journal of Mathematical Physics, 51(2):023501, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023501_s1.

Geloun:2010:GHA

[79] Joseph Ben Geloun and Mahouton Norbert Hounkonnou. q-graded Heisenberg algebras and deformed supersymmetries. Journal of Mathematical Physics, 51(2):023502, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p023502_s1.

Aazami:2010:UMT

[80] A. B. Aazami and A. O. Petters. A universal magnification theorem. III. Caustics beyond codimension 5. Journal of Mathematical Physics, 51(2):023503, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p023503_s1.

Chen:2010:PTS

[81] Shouxin Chen and Yisong Yang. Phase transition solutions in geometrically constrained magnetic domain wall models. Journal of Mathematical Physics, 51(2):023504, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p023504_s1.

Yin:2010:SPL

[82] Jiuli Yin and Lixin Tian. Stability of peakons and linear dispersion limit for the periodic Dullin–Gottwald–Holm equation. Journal of Mathematical Physics, 51(2):023505, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i2/p023505_s1.

Ordonez-Miranda:2010:HJQ

[83] J. Ordonez-Miranda, Miguel A. Zambrano-Arjona, and J. J. Alvarado-Gil. Hamilton–Jacobi and quantum theory formulations of thermal-wave propagation under the dual-phase lagging model of heat conduction. Journal of Mathematical Physics, 51(2):023506, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023506_s1.

Sheppeard:2010:AMU

[84] M. D. Sheppeard. Arithmetic, mutually unbiased bases and complementary observables. Journal of Mathematical Physics, 51(2):023507, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023507_s1.

Kim:2010:MDD

[85] Namhoon Kim. Multiplication of distributions and Dirac formalism of quantum mechanics. Journal of Mathematical Physics, 51(2):023508, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p023508_s1.

Goffeng:2010:IFC

[86] Magnus Goffeng. Index formulas and charge deficiencies on the Landau levels. Journal of Mathematical Physics, 51(2):023509, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023509_s1.

Wang:2010:IPG

[87] Deng-Shan Wang, Da-Jun Zhang, and Jianke Yang. Integrable properties of the general coupled nonlinear Schrödinger equations. Journal of Mathematical Physics, 51(2):023510, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023510_s1.

Malekar:2010:DFT

[88] R. A. Malekar and H. Bhate. Discrete Fourier transform and Jacobi θ function identities. Journal of Mathematical Physics, 51(2):023511, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p023511_s1.

Cardoso:2010:BST

[89] Fernando Cardoso and Georgi Vodev. Boundary stabilization of transmission problems. Journal of Mathematical Physics, 51(2):023512, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023512_s1.

Cardoso:2010:FDI

[90] J. G. Cardoso. A full description of the integrability condition for the twistor equation in curved space–times and new wave equations for conformally invariant spinor fields. Journal of Mathematical Physics, 51(2):023513, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p023513_s1.

Bornemann:2010:AIE

[91] Folkmar Bornemann. Asymptotic independence of the extreme eigenvalues of Gaussian unitary ensemble. Journal of Mathematical Physics, 51(2):023514, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p023514_s1.

Hrivnak:2010:TDSa

[92] Jirí Hrivnák and Jirí Patera. Two-dimensional symmetric and antisymmetric generalizations of exponential and cosine functions. Journal of Mathematical Physics, 51(2):023515, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i2/p023515_s1.

Ma:2010:IBV

[93] Li Ma and Pei Cao. Inhomogeneous boundary value problem for Hartreetype equation. Journal of Mathematical Physics, 51(2):023516, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023516_s1.

Guo:2010:GWS

[94] Boling Guo and Daiwen Huang. Global weak solution for a equations in plasma. Journal of Mathematical Physics, 51(2):023517, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023517_s1.

Passian:2010:PIT

[95] A. Passian, S. Koucheckian, S. B. Yakubovich, and T. Thundat. Properties of index transforms in modeling of nanostructures and plasmonic systems. Journal of Mathematical Physics, 51(2):023518, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p023518_s1.

Lenells:2010:NIG

[96] J. Lenells and A. S. Fokas. On a novel integrable generalization of the sine-Gordon equation. Journal of Mathematical Physics, 51(2):023519, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p023519_s1.

Talamini:2010:FBI

[97] Vittorino Talamini. Flat bases of invariant polynomials and P̂-matrices of E_{7} and E_{8}. Journal of Mathematical Physics, 51(2):023520, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023520_s1.

Ogundiran:2010:MPQ

[98] M. O. Ogundiran and E. O. Ayoola. Mayer problem for quantum stochastic control. Journal of Mathematical Physics, 51(2):023521, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i2/p023521_s1.

Keyl:2010:AGC

[99] Michael Keyl and Dirk-M. Schlingemann. The algebra of Grassmann canonical anticommutation relations and its applications to fermionic systems. Journal of Mathematical Physics, 51(2):023522, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p023522_s1.

Shcherbina:2010:CLT

[100] Mariya Shcherbina and Brunello Tirozzi. Central limit theorem for fluctuations of linear eigenvalue statistics of large random graphs. Journal of Mathematical Physics, 51(2):023523, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023523_s1.

Liu:2010:WMT

[101] Dong Liu, Yuezhu Wu, and Linsheng Zhu. Whittaker modules for the twisted Heisenberg–Virasoro algebra. Journal of Mathematical Physics, 51(2):023524, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p023524_s1.

Ikhdair:2010:ASD

[102] Sameer M. Ikhdair. Approximate solutions of the Dirac equation for the Rosen–Morse potential including the spin-orbit centrifugal term. Journal of Mathematical Physics, 51(2):023525, February 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023525_s1. See

comment [2708].

Kochubei:2010:NAN

[103] Anatoly N. Kochubei. Non-Archimedean normal operators. Journal of Mathematical Physics, 51(2):023526, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i2/p023526_s1.

Donninger:2010:RWO

[104] Roland Donninger. The radial wave operator in similarity coordinates. Journal of Mathematical Physics, 51(2):023527, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/p023527_s1.

Tamura:2010:LDP

[105] Hiroshi Tamura and Valentin A. Zagrebnov. Large deviation principle for noninteracting boson random point processes. Journal of Mathematical Physics, 51(2):023528, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i2/p023528_s1.

Lee:2010:TSM

[106] Kuo-Ming Lee. A two step method in inverse scattering problem for a crack. Journal of Mathematical Physics, 51(2):023529, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p023529_s1.

Greenberg:2010:SVM

[107] O. W. Greenberg and A. K. Mishra. Study of the vacuum matrix element of products of parafields. Journal of Mathematical Physics, 51(2):023530, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i2/p023530_s1.

Bagarello:2010:PRB

[108] F. Bagarello. Pseudobosons, Riesz bases, and coherent states. Journal of Mathematical Physics, 51(2):023531, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i2/p023531_s1.

Zhang:2010:CTP

[109] Shun Zhang. Comment on “On tractability of path integration”, [J. Math. Phys. 37, 2071 (1996)]. Journal of Mathematical Physics, 51(2):024101, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p024101_s1. See [2].

Niestegge:2010:ECS

[110] Gerd Niestegge. Erratum: “Composite systems and the role of the complex numbers in quantum mechanics”, [J. Math. Phys. 45, 4714 (2004)]. Journal of Mathematical Physics, 51(2):029901, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p029901_s1. See [5].

Fukuda:2010:PNA

[111] Motohisa Fukuda. Publisher’s note: “An application of decomposable maps in proving multiplicativity of low dimensional maps”, [J. Math. Phys. 51, 022201 (2010)]. Journal of Mathematical Physics, 51(2):029902, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p029902_s1. See [56].

Tanaka:2010:FSQ

[112] Toshiaki Tanaka. N-fold supersymmetry and quasi-solvability associated with X_{2}-Laguerre polynomials. Journal of Mathematical Physics, 51(3): 032101, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032101_s1.

Shi:2010:WPN

[113] Qihong Shi, Shu Wang, Yong Li, and Changyou Wang. Well posedness for the nonlinear Klein–Gordon–Schrödinger equations with heterointeractions. Journal of Mathematical Physics, 51(3):032102, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i3/p032102_s1.

Adachi:2010:STT

[114] Tadayoshi Adachi, Toshiyuki Kimura, and Yoshimasa Shimizu. Scattering theory for two-body quantum systems with singular potentials in a time-periodic electric field. Journal of Mathematical Physics, 51(3): 032103, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032103_s1.

Silvestre-Brac:2010:QBP

[115] Bernard Silvestre-Brac, Claude Semay, Fabien Buisseret, and Fabian Brau. The quantum N-body problem and the auxiliary field method. Journal of Mathematical Physics, 51(3):032104, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/p032104_s1.

Kang:2010:LES

[116] Yang Kang and Abel Klein. The Liouville equation for singular ergodic magnetic Schrödinger operators. Journal of Mathematical Physics, 51(3): 032105, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032105_s1.

Kleinmann:2010:SAU

[117] M. Kleinmann, H. Kampermann, and D. Bruß. Structural approach to unambiguous discrimination of two mixed quantum states. Journal of Mathematical Physics, 51(3):032201, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i3/p032201_s1.

Benamira:2010:PIS

[118] F. Benamira, L. Guechi, S. Mameri, and M. A. Sadoun. Path integral solutions for Klein–Gordon particle in vector plus scalar generalized Hulthén and Woods–Saxon potentials. Journal of Mathematical Physics, 51(3):032301, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p032301_s1.

Bellazzini:2010:QFS

[119] B. Bellazzini, M. Mintchev, and P. Sorba. Quantum fields on star graphs with bound states at the vertex. Journal of Mathematical Physics, 51(3): 032302, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032302_s1.

Gilka:2010:BDM

[120] Natalie Gilka. Bounds on diatomic molecules in a relativistic model. Journal of Mathematical Physics, 51(3):032303, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/p032303_s1.

Linnaeus:2010:PIS

[121] Staffan Linnæus. Phase-integral solution of the radial Dirac equation. Journal of Mathematical Physics, 51(3):032304, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/p032304_s1.

Ferrari:2010:RCS

[122] Ruggero Ferrari. On the renormalization of the complex scalar free field theory. Journal of Mathematical Physics, 51(3):032305, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p032305_s1.

Stuart:2010:ENL

[123] David Stuart. Existence and Newtonian limit of nonlinear bound states in the Einstein–Dirac system. Journal of Mathematical Physics, 51(3): 032501, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032501_s1.

Kruglikov:2010:IFE

[124] Boris Kruglikov. Involutivity of field equations. Journal of Mathematical Physics, 51(3):032502, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p032502_s1.

Takook:2010:CLG

[125] M. V. Takook, M. R. Tanhayi, and S. Fatemi. Conformal linear gravity in de Sitter space. Journal of Mathematical Physics, 51(3):032503, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i3/p032503_s1.

Bartnik:2010:SSS

[126] Robert Bartnik, Mark Fisher, and Todd A. Oliynyk. Static spherically symmetric solutions of the SO(5) Einstein Yang–Mills equations. Journal of Mathematical Physics, 51(3):032504, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i3/p032504_s1.

Zhijian:2010:GAT

[127] Yang Zhijian. Global attractors and their Hausdorff dimensions for a class of Kirchhoff models. Journal of Mathematical Physics, 51(3): 032701, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p032701_s1.

Yan:2010:RAF

[128] Weiping Yan, Yong Li, and Shuguan Ji. Random attractors for first order stochastic retarded lattice dynamical systems. Journal of Mathematical Physics, 51(3):032702, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p032702_s1.

Francoise:2010:PEF

[129] J.-P. Françoise, P. L. Garrido, and G. Gallavotti. Pendulum, elliptic functions, and relative cohomology classes. Journal of Mathematical Physics, 51(3):032901, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p032901_s1.

Ozakin:2010:GTT

[130] Arkadas Ozakin and Arash Yavari. A geometric theory of thermal stresses. Journal of Mathematical Physics, 51(3):032902, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p032902_s1.

Ohkitani:2010:BEP

[131] Koji Ohkitani and Mark Dowker. Burgers equation with a passive scalar: Dissipation anomaly and Colombeau calculus. Journal of Mathematical Physics, 51(3):033101, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033101_s1.

Zhou:2010:FTH

[132] Da-Sheng Zhou, Dang-Zheng Liu, and Tao Qian. Fixed trace β-Hermite ensembles: Asymptotic eigenvalue density and the edge of the density. Journal of Mathematical Physics, 51(3):033301, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/p033301_s1.

Sogo:2010:CAD

[133] Kiyoshi Sogo, Yoshiaki Kishikawa, Shuhei Ohnishi, Takenori Yamamoto, Susumu Fujiwara, and Keiko M. Aoki. Correlated anomalous diffusion: Random walk and Langevin equation. Journal of Mathematical Physics, 51(3):033302, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033302_s1.

Urrutia:2010:SMT

[134] Ignacio Urrutia and Leszek Szybisz. Statistical mechanics of two hard spheres in a spherical pore, exact analytic results in D dimension. Journal of Mathematical Physics, 51(3):033303, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i3/p033303_s1.

Gomes:2010:PBT

[135] J. F. Gomes, L. H. Ymai, and A. H. Zimerman. Permutability of Bäcklund transformations for N = 2 supersymmetric sine-Gordon. Journal of Mathematical Physics, 51(3):033501, March 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i3/p033501_s1.

Umarov:2010:GSS

[136] Sabir Umarov, Constantino Tsallis, Murray Gell-Mann, and Stanly Steinberg. Generalization of symmetric α-stable Lévy distributions for q > 1. Journal of Mathematical Physics, 51(3):033502, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p033502_s1.

Almeida:2010:FCV

[137] Ricardo Almeida, Agnieszka B. Malinowska, and Delfim F. M. Torres. A fractional calculus of variations for multiple integrals with application to vibrating string. Journal of Mathematical Physics, 51(3):033503, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i3/p033503_s1.

Xu:2010:EWT

[138] Tao Xu and Bo Tian. An extension of the Wronskian technique for the multicomponent Wronskian solution to the vector nonlinear Schrödinger equation. Journal of Mathematical Physics, 51(3):033504, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p033504_s1.

Hietarinta:2010:MSL

[139] Jarmo Hietarinta and Da jun Zhang. Multisoliton solutions to the lattice Boussinesq equation. Journal of Mathematical Physics, 51(3): 033505, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p033505_s1.

Yang:2010:NTF

[140] Chuan Fu Yang. New trace formulae for a quadratic pencil of the Schrödinger operator. Journal of Mathematical Physics, 51(3):033506, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p033506_s1. See erratum [776].

Chen:2010:EWS

[141] Shuhong Chen and Boling Guo. Existence of the weak solution of coupled time-dependent Ginzburg–Landau equations. Journal of Mathematical Physics, 51(3):033507, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033507_s1.

Wu:2010:IEE

[142] Zhigang Wu. On the inelastic Enskog equation near vacuum. Journal of Mathematical Physics, 51(3):033508, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i3/p033508_s1.

Song:2010:STG

[143] Xianfa Song. Sharp thresholds of global existence and blowup for a system of Schrödinger equations with combined power-type nonlinearities. Journal of Mathematical Physics, 51(3):033509, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/

p033509_s1. See erratum [509].

Liu:2010:ISD

[144] C. J. Liu. Inequivalent solutions for Dirac spin-1/2 particles under conservation of parity I. Journal of Mathematical Physics, 51(3):033510, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i3/p033510_s1.

Zhao:2010:CBH

[145] Jun xiao Zhao and Robert Conte. A connection between HH3 and Korteweg–de Vries with one source. Journal of Mathematical Physics, 51(3):033511, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033511_s1.

Conde:2010:ABT

[146] Juan M. Conde, Pilar R. Gordoa, and Andrew Pickering. Auto-Bäcklund transformations and integrability of ordinary and partial differential equations. Journal of Mathematical Physics, 51(3):033512, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p033512_s1.

Saygili:2010:TFR

[147] K. Saygili. Trkalian fields and Radon transformation. Journal of Mathematical Physics, 51(3):033513, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i3/p033513_s1.

Geng:2010:SNI

[148] Xianguo Geng and Guoliang He. Some new integrable nonlinear evolution equations and Darboux transformation. Journal of Mathematical Physics, 51(3):033514, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033514_s1.

Casati:2010:IRD

[149] Paolo Casati, Stefania Minniti, and Valentina Salari. Indecomposable representations of the Diamond Lie algebra. Journal of Mathematical Physics, 51(3):033515, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033515_s1.

Falek:2010:GBO

[150] M. Falek and M. Merad. A generalized bosonic oscillator in the presence of a minimal length. Journal of Mathematical Physics, 51(3): 033516, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p033516_s1.

Belgiorno:2010:DEK

[151] Francesco Belgiorno and Sergio L. Cacciatori. The Dirac equation in Kerr–Newman–Ads black hole background. Journal of Mathematical Physics, 51(3):033517, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p033517_s1.

Conte:2010:MST

[152] Robert Conte and Tuen-Wai Ng. Meromorphic solutions of a third order nonlinear differential equation. Journal of Mathematical Physics, 51(3): 033518, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p033518_s1.

Pradeep:2010:CNI

[153] R. Gladwin Pradeep, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. On certain new integrable second order nonlinear differential equations and their connection with two dimensional Lotka–Volterra system. Journal of Mathematical Physics, 51(3):033519, March 2010.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p033519_s1.

Scofield:2010:EHF

[154] D. F. Scofield and Pablo Huq. Evolution of helicity in fluid flows. Journal of Mathematical Physics, 51(3):033520, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i3/p033520_s1.

Schulze-Halberg:2010:IRD

[155] Axel Schulze-Halberg. Intertwining relations and Darboux transformations for Schrödinger equations in (n + 1) dimensions. Journal of Mathematical Physics, 51(3):033521, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i3/p033521_s1.

Xu:2010:ICF

[156] Xi-Xiang Xu. An integrable coupling family of the Toda lattice systems, its bi-Hamiltonian structure, and a related nonisospectral integrable lattice family. Journal of Mathematical Physics, 51(3):033522, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i3/
p033522_s1.

Castro:2010:CWF

[157] L. B. Castro and A. S. de Castro. Comment on “Wave functions for a Duffin–Kemmer–Petiau particle in a time-dependent potential”, [J. Math. Phys. 48, 073515 (2007)]. Journal of Mathematical Physics, 51(3): 034101, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p034101_s1. See [9].

Dutta:2010:CES

[158] D. Dutta and P. Roy. Conditionally exactly solvable potentials and exceptional orthogonal polynomials. Journal of Mathematical Physics, 51(4): 042101, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p042101_s1.

Nathanson:2010:TPS

[159] Michael Nathanson. Testing for a pure state with local operations and classical communication. Journal of Mathematical Physics, 51(4):042102, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p042102_s1.

Das:2010:ACP

[160] Ashok Das and L. Greenwood. An alternative construction of the positive inner product for pseudo-Hermitian Hamiltonians: Examples. Journal of Mathematical Physics, 51(4):042103, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p042103_s1.

Pavon:2010:DTC

[161] Michele Pavon and Francesco Ticozzi. Discrete-time classical and quantum Markovian evolutions: Maximum entropy problems on path space. Journal of Mathematical Physics, 51(4):042104, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i4/p042104_s1.

Fukuda:2010:ERS

[162] Motohisa Fukuda and Christopher King. Entanglement of random subspaces via the Hastings bound. Journal of Mathematical Physics, 51(4): 042201, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p042201_s1.

Beigi:2010:ASS

[163] Salman Beigi and Peter W. Shor. Approximating the set of separable states using the positive partial transpose test. Journal of Mathematical Physics, 51(4):042202, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p042202_s1.

Scott:2010:SIC

[164] A. J. Scott and M. Grassl. Symmetric informationally complete positiveoperator-valued measures: a new computer study. Journal of Mathematical Physics, 51(4):042203, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i4/p042203_s1.

Sakaguchi:2010:SGC

[165] Makoto Sakaguchi. Super-Galilean conformal algebra in AdS/CFT. Journal of Mathematical Physics, 51(4):042301, April 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i4/p042301_s1.

Nash:2010:SG

[166] Patrick L. Nash. Second gravity. Journal of Mathematical Physics, 51(4): 042501, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p042501_s1.

Estabrook:2010:HLI

[167] Frank B. Estabrook. The Hilbert Lagrangian and isometric embedding: Tetrad formulation of Regge–Teitelboim gravity. Journal of Mathematical Physics, 51(4):042502, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i4/p042502_s1.

Poon:2010:RUI

[168] Gary Kaiman Poon. Relative unitary implementability of perturbed quantum field dynamics on de Sitter space. Journal of Mathematical Physics, 51(4):042503, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p042503_s1.

Chang:2010:HSM

[169] Jen-Hsu Chang and Yu-Tung Chen. Hodograph solutions for the Manakov–Santini equation. Journal of Mathematical Physics, 51(4): 042701, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p042701_s1.

Sun:2010:INI

[170] Xu Sun, Jinqiao Duan, and Xiaofan Li. An impact of noise on invariant manifolds in nonlinear dynamical systems. Journal of Mathematical Physics, 51(4):042702, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p042702_s1.

Li:2010:BTW

[171] Jibin Li and Zhijun Qiao. Bifurcations of traveling wave solutions for an integrable equation. Journal of Mathematical Physics, 51(4):042703, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p042703_s1.

Llibre:2010:POH

[172] Jaume Llibre and Ana Rodrigues. On the periodic orbits of Hamiltonian systems. Journal of Mathematical Physics, 51(4):042704, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p042704_s1.

Komineas:2010:GST

[173] Stavros Komineas and Nikos Papanicolaou. Gröbli solution for three magnetic vortices. Journal of Mathematical Physics, 51(4):042705, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p042705_s1.

Ranada:2010:NPH

[174] Manuel F. Rañada, Miguel A. Rodríguez, and Mariano Santander. A new proof of the higher-order superintegrability of a noncentral oscillator with inversely quadratic nonlinearities. Journal of Mathematical Physics, 51(4):042901, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p042901_s1.

Xie:2010:SCC

[175] Zhifu Xie. Super central configurations of the n-body problem. Journal of Mathematical Physics, 51(4):042902, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p042902_s1.

Mendes:2010:PVS

[176] R. Vilela Mendes. Poisson–Vlasov in a strong magnetic field: a stochastic solution approach. Journal of Mathematical Physics, 51(4):043101, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p043101_s1.

Navardi:2010:NLT

[177] Shahin Navardi and Sukalyan Bhattacharya. A new lubrication theory to derive far-field axial pressure difference due to force singularities in cylindrical or annular vessels. Journal of Mathematical Physics, 51(4): 043102, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p043102_s1.

Marinucci:2010:EGS

[178] Domenico Marinucci and Giovanni Peccati. Ergodicity and Gaussianity for spherical random fields. Journal of Mathematical Physics, 51(4): 043301, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p043301_s1.

Hopkins:2010:SCM

[179] Adam B. Hopkins, Frank H. Stillinger, and Salvatore Torquato. Spherical codes, maximal local packing density, and the golden ratio. Journal of Mathematical Physics, 51(4):043302, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043302_s1.

Bologna:2010:DAB

[180] Mauro Bologna, Gianluca Ascolani, and Paolo Grigolini. Density approach to ballistic anomalous diffusion: an exact analytical treatment. Journal of Mathematical Physics, 51(4):043303, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i4/p043303_s1.

Koca:2010:CSD

[181] Mehmet Koca, Nazife Ozdes Koca, and Ramazan Koç. Catalan solids derived from three-dimensional-root systems and quaternions. Journal of Mathematical Physics, 51(4):043501, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043501_s1.

Chen:2010:OAA

[182] William Y. C. Chen, Husam L. Saad, and Lisa H. Sun. An operator approach to the Al-Salam–Carlitz polynomials. Journal of Mathematical Physics, 51(4):043502, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p043502_s1.

Nakamula:2010:MR

[183] Atsushi Nakamula and Jun Sakaguchi. Multicalorons revisited. Journal of Mathematical Physics, 51(4):043503, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043503_s1.

Ammar:2010:BID

[184] Mabrouk Ben Ammar, Nizar Ben Fraj, and Salem Omri. The binary invariant differential operators on weighted densities on the superspace R^{1|n} and cohomology. Journal of Mathematical Physics, 51(4):043504, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p043504_s1.

Cao:2010:ICN

[185] Daomin Cao and Pigong Han. Inhomogeneous critical nonlinear Schrödinger equations with a harmonic potential. Journal of Mathematical Physics, 51(4):043505, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i4/p043505_s1.

Hamdi:2010:OPO

[186] Adel Hamdi and Jiang Zeng. Orthogonal polynomials and operator orderings. Journal of Mathematical Physics, 51(4):043506, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i4/
p043506_s1.

Radha:2010:LSD

[187] R. Radha, V. Ramesh Kumar, and Miki Wadati. Line-soliton dynamics and stability of Bose–Einstein condensates in (2 + 1) Gross–Pitaevskii equation. Journal of Mathematical Physics, 51(4):043507, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i4/
p043507_s1.

Li:2010:GPS

[188] Jibin Li and Yi Zhang. The geometric property of soliton solutions for the integrable KdV6 equations. Journal of Mathematical Physics, 51(4): 043508, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i4/p043508_s1.

Doikou:2010:BS

[189] Anastasia Doikou. On boundary superalgebras. Journal of Mathematical Physics, 51(4):043509, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p043509_s1.

Zhang:2010:TKC

[190] Yufeng Zhang and Honwah Tam. Three kinds of coupling integrable couplings of the Korteweg–de Vries hierarchy of evolution equations. Journal of Mathematical Physics, 51(4):043510, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043510_s1.

Lu:2010:IAO

[191] Xing Lü, Juan Li, Hai-Qiang Zhang, Tao Xu, Li-Li Li, and Bo Tian. Integrability aspects with optical solitons of a generalized variable-coefficient N-coupled higher order nonlinear Schrödinger system from inhomogeneous optical fibers. Journal of Mathematical Physics, 51(4):043511, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p043511_s1.

Mansour:2010:LDE

[192] Toufik Mansour and Matthias Schork. On linear differential equations with variable coefficients involving a para-Grassmann variable. Journal of Mathematical Physics, 51(4):043512, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043512_s1.

Liu:2010:DCB

[193] Haitao Liu. Dimension of conformal blocks in five dimensional Kähler– Chern–Simons theory. Journal of Mathematical Physics, 51(4):043513, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p043513_s1.

Li:2010:TFH

[194] Chuanzhong Li, Jingsong He, Ke Wu, and Yi Cheng. Tau function and Hirota bilinear equations for the extended bigraded Toda hierarchy. Journal of Mathematical Physics, 51(4):043514, April 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i4/p043514_s1.

Arnlind:2010:THN

[195] Joakim Arnlind, Abdenacer Makhlouf, and Sergei Silvestrov. Ternary Hom–Nambu–Lie algebras induced by Hom–Lie algebras. Journal of Mathematical Physics, 51(4):043515, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p043515_s1.

Kulish:2010:SSS

[196] P. P. Kulish, N. Manojlović, and Z. Nagy. Symmetries of spin systems and Birman–Wenzl–Murakami algebra. Journal of Mathematical Physics, 51(4):043516, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p043516_s1.

Chen:2010:CIR

[197] Yong Chen. Comment on “On the imaginary-real ratio rule of power spectra”, [J. Math. Phys. 50, 063301 (2009)]. Journal of Mathematical Physics, 51(4):044101, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p044101_s1. See [14].

Park:2010:CAT

[198] Choonkil Park and M. Eshaghi Gordji. Comment on “Approximate ternary Jordan derivations on Banach ternary algebras”, [Bavand Savadkouhi et al. J. Math. Phys. 50, 042303 (2009)]. Journal of Mathematical Physics, 51(4):044102, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p044102_s1. See [12].

Wada:2010:EGL

[199] Tatsuaki Wada. Erratum: “Generalized log-likelihood functions and Bregman divergences”, [J. Math. Phys. 80, 113301 (2009)]. Journal of Mathematical Physics, 51(4):049901, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p049901_s1. See [19].

Yang:2010:ESS

[200] Yisong Yang and Ruifeng Zhang. Erratum: “Steady state solutions for nonlinear Schrödinger equation arising in optics”, [J. Math. Phys. 50, 053501 (2009)]. Journal of Mathematical Physics, 51(4):049902, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p049902_s1. See [13].

Guo:2010:PMS

[201] Qiang Guo, Shahla Molahajloo, and M. W. Wong. Phases of modified Stockwell transforms and instantaneous frequencies. Journal of Mathematical Physics, 51(5):052101, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p052101_s1.

Vourdas:2010:HQS

[202] A. Vourdas. Hamiltonians of quantum systems with positions and momenta in GF (p^{ℓ}). Journal of Mathematical Physics, 51(5):052102, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052102_s1.

Fritz:2010:EQR

[203] Tobias Fritz. On the existence of quantum representations for two dichotomic measurements. Journal of Mathematical Physics, 51(5):052103, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052103_s1.

Andrianov:2010:SSN

[204] A. A. Andrianov, F. Cannata, and A. V. Sokolov. Spectral singularities for non-Hermitian one-dimensional Hamiltonians: Puzzles with resolution of identity. Journal of Mathematical Physics, 51(5):052104, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052104_s1.

Amore:2010:SDQ

[205] Paolo Amore. Spectroscopy of drums and quantum billiards: Perturbative and nonperturbative results. Journal of Mathematical Physics, 51 (5):052105, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052105_s1.

Ruby:2010:CCS

[206] V. Chithiika Ruby and M. Senthilvelan. On the construction of coherent states of position dependent mass Schrödinger equation endowed with effective potential. Journal of Mathematical Physics, 51(5):052106, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052106_s1.

Scarfone:2010:CSE

[207] Leonard M. Scarfone. Callan–Symanzik equation and asymptotic freedom in the Marr–Shimamoto model. Journal of Mathematical Physics, 51 (5):052107, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052107_s1.

Domenech:2010:QLG

[208] Graciela Domenech, Federico Holik, and César Massri. A quantum logical and geometrical approach to the study of improper mixtures. Journal of Mathematical Physics, 51(5):052108, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p052108_s1.

Alfsen:2010:UDF

[209] Erik Alfsen and Fred Shultz. Unique decompositions, faces, and automorphisms of separable states. Journal of Mathematical Physics, 51(5): 052201, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052201_s1.

Lim:2010:NPB

[210] Bunrith Jacques Lim. Noncommutative Poisson boundaries of unital quantum operations. Journal of Mathematical Physics, 51(5):052202, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052202_s1.

Banerjee:2010:SFY

[211] Rabin Banerjee and Shinichi Deguchi. A superspace formulation of Yang– Mills theory on sphere. Journal of Mathematical Physics, 51(5):052301, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052301_s1.

Pei:2010:RCC

[212] Yufeng Pei and Chengming Bai. Realizations of conformal current-type Lie algebras. Journal of Mathematical Physics, 51(5):052302, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/
p052302_s1.

Wehbe:2010:SAN

[213] Mohammad Wehbe. Spinor algebra and null solutions of the wave equation. Journal of Mathematical Physics, 51(5):052303, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/p052303_s1.

Allendes:2010:NFD

[214] Pedro Allendes, Natanael Guerrero, Igor Kondrashuk, and Eduardo A. Notte Cuello. New four-dimensional integrals by Mellin–Barnes transform. Journal of Mathematical Physics, 51(5):052304, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/
p052304_s1.

Chuang:2010:BPS

[215] Wu-Yen Chuang and Guang Pan. Bogomolny–Prasad–Sommerfeld state counting in local obstructed curves from quiver theory and Seiberg duality. Journal of Mathematical Physics, 51(5):052305, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/p052305_s1.

Jiang:2010:EAU

[216] Mingming Jiang and Sixia Yu. Extremal asymmetric universal cloning machines. Journal of Mathematical Physics, 51(5):052306, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/
p052306_s1.

Hosseiny:2010:AEG

[217] Ali Hosseiny and Shahin Rouhani. Affine extension of Galilean conformal algebra in 2 + 1 dimensions. Journal of Mathematical Physics, 51(5): 052307, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052307_s1.

Galstian:2010:NWE

[218] Anahit Galstian, Tamotu Kinoshita, and Karen Yagdjian. A note on wave equation in Einstein and de Sitter space–time. Journal of Mathematical Physics, 51(5):052501, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p052501_s1.

Sakthivel:2010:ASS

[219] R. Sakthivel, Yong Ren, and Hyunsoo Kim. Asymptotic stability of second-order neutral stochastic differential equations. Journal of Mathematical Physics, 51(5):052701, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p052701_s1.

Liu:2010:ARL

[220] Yuewei Liu, Lu Yang, and Chengkui Zhong. Asymptotic regularity for p-Laplacian equation. Journal of Mathematical Physics, 51(5):052702, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052702_s1.

Jakel:2010:MPC

[221] Christian D. Jäkel, Heide Narnhofer, and Walter F. Wreszinski. On the mixing property for a class of states of relativistic quantum fields. Journal of Mathematical Physics, 51(5):052703, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p052703_s1.

Chugunova:2010:CEG

[222] Marina Chugunova and Dmitry Pelinovsky. Count of eigenvalues in the generalized eigenvalue problem. Journal of Mathematical Physics, 51(5): 052901, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052901_s1.

Gralak:2010:MME

[223] B. Gralak and A. Tip. Macroscopic Maxwell’s equations and negative index materials. Journal of Mathematical Physics, 51(5):052902, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p052902_s1.

Aharonovich:2010:RFU

[224] I. Aharonovich and L. P. Horwitz. Radiation fields of a uniformly accelerating point source in the framework of Stueckelberg’s manifestly covariant relativistic dynamics. Journal of Mathematical Physics, 51(5): 052903, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p052903_s1.

Escher:2010:PEE

[225] Joachim Escher and Jörg Seiler. The periodic b-equation and Euler equations on the circle. Journal of Mathematical Physics, 51(5):053101, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053101_s1.

Kim:2010:RCN

[226] Jae-Myoung Kim. On regularity criteria of the Navier–Stokes equations in bounded domains. Journal of Mathematical Physics, 51(5):053102, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053102_s1.

Yang:2010:LDS

[227] Desheng Yang and Jinqiao Duan. Large deviations for the stochastic quasigeostrophic equation with multiplicative noise. Journal of Mathematical Physics, 51(5):053301, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p053301_s1.

Lee:2010:LBG

[228] Ji Oon Lee and Jun Yin. A lower bound on the ground state energy of dilute Bose gas. Journal of Mathematical Physics, 51(5):053302, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053302_s1.

Gallavotti:2010:TLI

[229] G. Gallavotti and E. Presutti. Thermodynamic limit for isokinetic thermostats. Journal of Mathematical Physics, 51(5):053303, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/
p053303_s1.

Barletti:2010:QDD

[230] Luigi Barletti and Florian Méhats. Quantum drift-diffusion modeling of spin transport in nanostructures. Journal of Mathematical Physics, 51 (5):053304, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053304_s1.

Balaban:2010:PSRa

[231] Tadeusz Balaban, Joel Feldman, Horst Knörrer, and Eugene Trubowitz. Power series representations for complex bosonic effective actions. I. A small field renormalization group step. Journal of Mathematical Physics, 51(5):053305, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053305_s1.

Balaban:2010:PSRb

[232] Tadeusz Balaban, Joel Feldman, Horst Knörrer, and Eugene Trubowitz. Power series representations for complex bosonic effective actions. II. A small field renormalization group flow. Journal of Mathematical Physics, 51(5):053306, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053306_s1.

Schulze-Halberg:2010:GFT

[233] Axel Schulze-Halberg. Green’s functions and trace formulas for generalized Sturm–Liouville problems related by Darboux transformations. Journal of Mathematical Physics, 51(5):053501, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/p053501_s1.

LaGatta:2010:STR

[234] T. LaGatta and J. Wehr. A shape theorem for Riemannian first-passage percolation. Journal of Mathematical Physics, 51(5):053502, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i5/
p053502_s1.

Klaus:2010:WOM

[235] Martin Klaus and Cornelis van der Mee. Wave operators for the matrix Zakharov–Shabat system. Journal of Mathematical Physics, 51(5): 053503, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053503_s1.

Heo:2010:KRK

[236] Jaeseong Heo, Jang Pyo Hong, and Un Cig Ji. On KSGNS representations on Krein C^{∗}-modules. Journal of Mathematical Physics, 51(5): 053504, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053504_s1.

Gordoa:2010:MSA

[237] Pilar R. Gordoa, Andrew Pickering, and Zuo-Nong Zhu. Matrix semidiscrete Ablowitz–Ladik equation hierarchy and a matrix discrete second Painlevé equation. Journal of Mathematical Physics, 51(5):053505, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053505_s1.

AlKhawaja:2010:CAP

[238] U. Al Khawaja. A comparative analysis of Painlevé, Lax pair, and similarity transformation methods in obtaining the integrability conditions of nonlinear Schrödinger equations. Journal of Mathematical Physics, 51 (5):053506, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053506_s1.

Chen:2010:DML

[239] Zhiqi Chen and Ke Liang. The decomposition of metric n-Lie algebras and the uniqueness. Journal of Mathematical Physics, 51(5):053507, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053507_s1.

Bagarello:2010:CPS

[240] F. Bagarello. Construction of pseudobosons systems. Journal of Mathematical Physics, 51(5):053508, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p053508_s1.

Anghel:2010:SCG

[241] N. Anghel. Stieltjes–Calogero–Gil’ relations associated to entire functions of finite order. Journal of Mathematical Physics, 51(5):053509, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053509_s1.

Nieminen:2010:TTR

[242] John M. Nieminen. Two-by-two random matrix theory with matrix representations of octonions. Journal of Mathematical Physics, 51(5): 053510, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053510_s1.

Tsuchida:2010:NRI

[243] Takayuki Tsuchida. New reductions of integrable matrix partial differential equations: Sp(m)-invariant systems. Journal of Mathematical Physics, 51(5):053511, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053511_s1.

Kirsten:2010:SFS

[244] Klaus Kirsten and Paul Loya. Spectral functions for the Schrödinger operator on R^{+} with a singular potential. Journal of Mathematical Physics, 51(5):053512, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053512_s1.

Odake:2010:IMS

[245] S. Odake and R. Sasaki. Infinitely many shape-invariant potentials and cubic identities of the Laguerre and Jacobi polynomials. Journal of Mathematical Physics, 51(5):053513, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p053513_s1.

Cheng:2010:GSB

[246] Jipeng Cheng, Jingsong He, and Sen Hu. The “ghost” symmetry of the BKP hierarchy. Journal of Mathematical Physics, 51(5):053514, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053514_s1.

Yin:2010:SNS

[247] Jiuli Yin, Lixin Tian, and Xinghua Fan. Stability of negative solitary waves for an integrable modified Camassa–Holm equation. Journal of Mathematical Physics, 51(5):053515, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p053515_s1.

Martin:2010:GRF

[248] A. J. Calderón Martín, C. Draper, and C. Martín González. Gradings on the real forms of the Albert algebra, of }_{2}, and of {_{4}. Journal of Mathematical Physics, 51(5):053516, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i5/p053516_s1.

Bokhari:2010:SIF

[249] Ashfaque H. Bokhari, F. M. Mahomed, and F. D. Zaman. Symmetries and integrability of a fourth-order Euler–Bernoulli beam equation. Journal of Mathematical Physics, 51(5):053517, May 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i5/p053517_s1.

Cantelaube:2010:LPC

[250] Y. C. Cantelaube and A. L. Khelif. Laplacian in polar coordinates, regular singular function algebra, and theory of distributions. Journal of Mathematical Physics, 51(5):053518, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p053518_s1.

Fan:2010:STC

[251] Tian You Fan, Zhi Yi Tang, Lian He Li, and Wu Li. The strict theory of complex variable function method for sextuple harmonic equation and applications. Journal of Mathematical Physics, 51(5):053519, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053519_s1.

Shen:2010:OSV

[252] Chunyu Shen and Anna Gao. Optimal solution for the viscous nonlinear dispersive wave equation. Journal of Mathematical Physics, 51(5): 053520, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p053520_s1.

Daboul:2010:EOS

[253] Claudia Daboul, Jamil Daboul, and Marc de Montigny. Embedding of “open” superalgebras into affine Kac–Moody superalgebras. Journal of Mathematical Physics, 51(5):053521, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p053521_s1.

Beggs:2010:QCT

[254] E. J. Beggs and S. Majid. Quantization by cochain twists and nonassociative differentials. Journal of Mathematical Physics, 51(5):053522, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053522_s1.

Chung:2010:HKA

[255] Jae-Young Chung. A heat kernel approach to the stability of exponential equation in distributions and hyperfunctions. Journal of Mathematical Physics, 51(5):053523, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053523_s1.

Hasebe:2010:SAN

[256] Kazuki Hasebe. The split algebras and noncompact Hopf maps. Journal of Mathematical Physics, 51(5):053524, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p053524_s1.

Tolksdorf:2010:RBM

[257] Jürgen Tolksdorf. Real Z_{2}-bigradings, Majorana modules, and the standard model action. Journal of Mathematical Physics, 51(5):053525, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053525_s1.

Pauley:2010:PCL

[258] Michael Pauley. Parabolic curves in Lie groups. Journal of Mathematical Physics, 51(5):053526, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i5/p053526_s1.

Flori:2010:TFH

[259] C. Flori. A topos formulation of history quantum theory. Journal of Mathematical Physics, 51(5):053527, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i5/p053527_s1.

Wang:2010:DTQ

[260] Caishi Wang, Huifang Chai, and Yanchun Lu. Discrete-time quantum Bernoulli noises. Journal of Mathematical Physics, 51(5):053528, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p053528_s1.

Andrzejewski:2010:CFD

[261] Krzysztof Andrzejewski, Joanna Gonera, and Paweł Maślanka. Comment on “Functional determinants in higher derivative Lagrangian theories”, [J. Math. Phys. 50, 103517 (2009)]. Journal of Mathematical Physics, 51 (5):054101, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p054101_s1. See [17].

Wang:2010:CGN

[262] Hai-Jun Wang. Comment on “General nonlocality in quantum fields”, [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 51(5): 054102, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p054102_s1. See [10].

DiCriscienzo:2010:EFD

[263] Roberto Di Criscienzo and Sergio Zerbini. Erratum: “Functional determinants in higher derivative Lagrangian theories”, [J. Math. Phys. 50, 103517 (2009)]. Journal of Mathematical Physics, 51(5):059901, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p059901_s1. See [17].

Keating:2010:LES

[264] J. P. Keating, J. Marklof, and B. Winn. Localized eigenfunctions in Seba billiards. Journal of Mathematical Physics, 51(6):062101, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p062101_s1.

Jeng:2010:NFS

[265] M. Jeng, S.-L.-Y. Xu, E. Hawkins, and J. M. Schwarz. On the nonlocality of the fractional Schrödinger equation. Journal of Mathematical Physics, 51(6):062102, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p062102_s1.

Gardas:2010:REP

[266] Bartlomiej Gardas. Riccati equation and the problem of decoherence. Journal of Mathematical Physics, 51(6):062103, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p062103_s1.

Cosic:2010:GTA

[267] Marko Ćosić. Group-theoretical approach to Bloch electron in magnetic field problem. Journal of Mathematical Physics, 51(6):062104, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p062104_s1.

Kolomietz:2010:CAC

[268] V. M. Kolomietz and S. V. Radionov. Cranking approach for a complex quantum system. Journal of Mathematical Physics, 51(6):062105, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p062105_s1.

Nagaj:2010:FUQ

[269] Daniel Nagaj. Fast universal quantum computation with railroad-switch local Hamiltonians. Journal of Mathematical Physics, 51(6):062201, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p062201_s1.

Ichikawa:2010:SPF

[270] Tsubasa Ichikawa, Toshihiko Sasaki, and Izumi Tsutsui. Separability of N-particle fermionic states for arbitrary partitions. Journal of Mathematical Physics, 51(6):062202, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i6/p062202_s1.

Guadagnini:2010:ALI

[271] Enore Guadagnini and Francesco Mancarella. Abelian link invariants and homology. Journal of Mathematical Physics, 51(6):062301, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p062301_s1.

Cherney:2010:BDQ

[272] D. Cherney, E. Latini, and A. Waldron. BRST detour quantization: Generating gauge theories from constraints. Journal of Mathematical Physics, 51(6):062302, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p062302_s1.

Zheltukhin:2010:IMG

[273] A. A. Zheltukhin and M. Trzetrzelewski. U(1)-invariant membranes: The geometric formulation, Abel, and pendulum differential equations. Journal of Mathematical Physics, 51(6):062303, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p062303_s1.

Hlavaty:2010:DSP

[274] Ladislav Hlavatý, Vojtech Stepán, and Jan Vysoký. Drinfel’d superdoubles and Poisson–Lie T -plurality in low dimensions. Journal of Mathematical Physics, 51(6):062304, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i6/p062304_s1.

DAmbroise:2010:DCB

[275] Jennie D’Ambroise and Floyd L. Williams. A dynamic correspondence between Bose–Einstein condensates and Friedmann–Lemaître– Robertson–Walker and Bianchi I cosmology with a cosmological constant. Journal of Mathematical Physics, 51(6):062501, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p062501_s1.

Tanabe:2010:ASN

[276] Kentaro Tanabe, Norihiro Tanahashi, and Tetsuya Shiromizu. On asymptotic structure at null infinity in five dimensions. Journal of Mathematical Physics, 51(6):062502, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i6/p062502_s1.

Berti:2010:HPT

[277] V. Berti, M. Fabrizio, and D. Grandi. Hysteresis and phase transitions for one-dimensional and three-dimensional models in shape memory alloys. Journal of Mathematical Physics, 51(6):062901, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p062901_s1.

Varga:2010:STD

[278] Péter Varga. On the solvability of two dimensional semigroup gauge theories. Journal of Mathematical Physics, 51(6):063301, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063301_s1.

Tracy:2010:FJP

[279] Craig A. Tracy and Harold Widom. Formulas for joint probabilities for the asymmetric simple exclusion process. Journal of Mathematical Physics, 51(6):063302, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p063302_s1.

Pankavich:2010:MSB

[280] S. Pankavich and P. Ortoleva. Multiscaling for systems with a broad continuum of characteristic lengths and times: Structural transitions in nanocomposites. Journal of Mathematical Physics, 51(6):063303, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p063303_s1.

Jauregui:2010:NRD

[281] M. Jauregui and C. Tsallis. New representations of π and Dirac delta using the nonextensive-statistical-mechanics q-exponential function. Journal of Mathematical Physics, 51(6):063304, June 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063304_s1.

Hsia:2010:BBS

[282] Chun-Hsiung Hsia. Bifurcation of binary systems with the Onsager mobility. Journal of Mathematical Physics, 51(6):063305, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/
p063305_s1.

Saito:2010:FJS

[283] Satoru Saito and Noriko Saitoh. Fate of the Julia set of higher dimensional maps in the integrable limit. Journal of Mathematical Physics, 51 (6):063501, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i6/p063501_s1.

Albeverio:2010:SPB

[284] S. Albeverio, G. Cognola, M. Spreafico, and S. Zerbini. Singular perturbations with boundary conditions and the Casimir effect in the half space. Journal of Mathematical Physics, 51(6):063502, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063502_s1.

Yao:2010:GKD

[285] Yuqin Yao and Yunbo Zeng. The generalized Kupershmidt deformation for constructing new integrable systems from integrable bi-Hamiltonian systems. Journal of Mathematical Physics, 51(6):063503, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/
p063503_s1.

Wu:2010:BST

[286] Chao-Zhong Wu and Dingdian Xu. Bihamiltonian structure of the twocomponent Kadomtsev–Petviashvili hierarchy of type B. Journal of Mathematical Physics, 51(6):063504, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i6/p063504_s1.

Bai:2010:RLA

[287] Ruipu Bai, Chengming Bai, and Jinxiu Wang. Realizations of 3-Lie algebras. Journal of Mathematical Physics, 51(6):063505, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/
p063505_s1.

Garcia-Compean:2010:LIF

[288] Hugo García-Compeán and Roberto Santos-Silva. Link invariants for flows in higher dimensions. Journal of Mathematical Physics, 51(6): 063506, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i6/p063506_s1.

Boyallian:2010:IMF

[289] Carina Boyallian, Victor G. Kac, and José I. Liberati. Irreducible modules over finite simple Lie conformal superalgebras of type K. Journal of Mathematical Physics, 51(6):063507, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i6/p063507_s1.

Biswas:2010:HVB

[290] Indranil Biswas. Holomorphic vector bundles on Sasakian threefolds and parabolic vector bundles. Journal of Mathematical Physics, 51(6): 063508, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i6/p063508_s1.

Tierz:2010:SPB

[291] Miguel Tierz. Schur polynomials and biorthogonal random matrix ensembles. Journal of Mathematical Physics, 51(6):063509, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063509_s1.

Ulrych:2010:CHC

[292] S. Ulrych. Considerations on the hyperbolic complex Klein–Gordon equation. Journal of Mathematical Physics, 51(6):063510, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063510_s1.

Dosi:2010:QDU

[293] Anar Dosi. Quantum duality, unbounded operators, and inductive limits. Journal of Mathematical Physics, 51(6):063511, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063511_s1.

Ma:2010:MCN

[294] Yulan Ma and Bangqing Li. A method for constructing nontraveling wave solutions for (1 + 1)-dimensional evolution equations. Journal of Mathematical Physics, 51(6):063512, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i6/p063512_s1.

Scholz:2010:ADE

[295] Daniel Scholz, Volodymyr G. Voronov, and Michael Weyrauch. Approximately disentangling exponential operators. Journal of Mathematical Physics, 51(6):063513, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p063513_s1.

Shen:2010:TRS

[296] Jun Shen and Junde Wu. The n th root of sequential effect algebras. Journal of Mathematical Physics, 51(6):063514, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/p063514_s1.

Xiao:2010:GPE

[297] Jie Xiao. L^{p}-Green potential estimates on noncompact Riemannian manifolds. Journal of Mathematical Physics, 51(6):063515, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/
p063515_s1.

Geng:2010:SSQ

[298] Xianguo Geng and Bo Xue. Soliton solutions and quasiperiodic solutions of modified Korteweg–de Vries type equations. Journal of Mathematical Physics, 51(6):063516, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p063516_s1. See [2317].

Li:2010:SSC

[299] Yajie Li, Ben T. Nohara, and Shijun Liao. Series solutions of coupled van der Pol equation by means of homotopy analysis method. Journal of Mathematical Physics, 51(6):063517, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i6/p063517_s1.

Hounkonnou:2010:DQA

[300] Mahouton Norbert Hounkonnou and Joseph Désiré Bukweli Kyemba. (R, p, q)-deformed quantum algebras: Coherent states and special functions. Journal of Mathematical Physics, 51(6):063518, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i6/
p063518_s1.

Xu:2010:IFS

[301] Tian Zhou Xu, John Michael Rassias, and Wan Xin Xu. Intuitionistic fuzzy stability of a general mixed additive-cubic equation. Journal of Mathematical Physics, 51(6):063519, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i6/p063519_s1.

Chang-Young:2010:MEN

[302] Ee Chang-Young, Hoil Kim, and Hiroaki Nakajima. Morita equivalence of noncommutative supertori. Journal of Mathematical Physics, 51(6): 063520, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i6/p063520_s1.

Kashima:2010:EDC

[303] Yohei Kashima. Exponential decay of correlation functions in manyelectron systems. Journal of Mathematical Physics, 51(6):063521, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i6/p063521_s1.

Dayi:2010:NSA

[304] Ömer F. Dayi and Ahmed Jellal. A noncommutative space approach to confined Dirac fermions in graphene. Journal of Mathematical Physics, 51(6):063522, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i6/p063522_s1.

Wu:2010:CPG

[305] Xinglong Wu and Zhaoyang Yin. On the Cauchy problem for a generalized Degasperis–Procesi equation. Journal of Mathematical Physics, 51 (6):063523, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i6/p063523_s1.

Dias:2010:DQT

[306] Nuno Costa Dias, Maurice de Gosson, Franz Luef, and João Nuno Prata. A deformation quantization theory for noncommutative quantum mechanics. Journal of Mathematical Physics, 51(7):072101, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p072101_s1.

Hilling:2010:GMM

[307] Joseph J. Hilling and Anthony Sudbery. The geometric measure of multipartite entanglement and the singular values of a hypermatrix. Journal of Mathematical Physics, 51(7):072102, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p072102_s1.

Trisetyarso:2010:CDP

[308] Agung Trisetyarso. Correlation of Dirac potentials and atomic inversion in cavity quantum electrodynamics. Journal of Mathematical Physics, 51 (7):072103, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p072103_s1. See erratum [636].

Brouder:2010:TET

[309] Christian Brouder, Ângela Mestre, and Frédéric Patras. Tree expansion in time-dependent perturbation theory. Journal of Mathematical Physics, 51(7):072104, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p072104_s1.

Kiukas:2010:MVB

[310] J. Kiukas and R. F. Werner. Maximal violation of Bell inequalities by position measurements. Journal of Mathematical Physics, 51(7):072105, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p072105_s1.

Zinn-Justin:2010:ODM

[311] Jean Zinn-Justin and Ulrich D. Jentschura. Order-dependent mappings: Strong-coupling behavior from weak-coupling expansions in non-Hermitian theories. Journal of Mathematical Physics, 51(7):072106, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p072106_s1.

Honda:2010:SRL

[312] Daigo Honda, Hiromichi Nakazato, and Motoyuki Yoshida. Spectral resolution of the Liouvillian of the Lindblad master equation for a harmonic oscillator. Journal of Mathematical Physics, 51(7):072107, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p072107_s1.

Hemery:2010:WHE

[313] A. D. Hemery and A. P. Veselov. Whittaker–Hill equation and semifinitegap Schrödinger operators. Journal of Mathematical Physics, 51(7): 072108, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p072108_s1.

Bradler:2010:CDQ

[314] Kamil Brádler, Nicolas Dutil, Patrick Hayden, and Abubakr Muhammad. Conjugate degradability and the quantum capacity of cloning channels. Journal of Mathematical Physics, 51(7):072201, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i7/p072201_s1.

Dita:2010:HMM

[315] P. Dita. Hadamard matrices from mutually unbiased bases. Journal of Mathematical Physics, 51(7):072202, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p072202_s1.

Nussbaum:2010:EER

[316] Michael Nussbaum and Arleta Szkola. Exponential error rates in multiple state discrimination on a quantum spin chain. Journal of Mathematical Physics, 51(7):072203, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p072203_s1.

Urias:2010:HFU

[317] Jesús Urías. Householder factorizations of unitary matrices. Journal of Mathematical Physics, 51(7):072204, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p072204_s1.

Chitambar:2010:MPE

[318] Eric Chitambar, Carl A. Miller, and Yaoyun Shi. Matrix pencils and entanglement classification. Journal of Mathematical Physics, 51(7): 072205, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p072205_s1.

Finster:2010:CPE

[319] Felix Finster and Andreas Grotz. The causal perturbation expansion revisited: Rescaling the interacting Dirac sea. Journal of Mathematical Physics, 51(7):072301, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p072301_s1.

Rahn:2010:YME

[320] Thorsten Rahn. Yang–Mills equations of motion for the Higgs sector of SU(3)-equivariant quiver gauge theories. Journal of Mathematical Physics, 51(7):072302, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p072302_s1.

Gicquaud:2010:LSE

[321] Romain Gicquaud. Linearization stability of the Einstein constraint equations on an asymptotically hyperbolic manifold. Journal of Mathematical Physics, 51(7):072501, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i7/p072501_s1.

Thirukkanesh:2010:MPR

[322] S. Thirukkanesh and S. D. Maharaj. Mixed potentials in radiative stellar collapse. Journal of Mathematical Physics, 51(7):072502, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i7/
p072502_s1.

Miao:2010:SBT

[323] S. P. Miao, N. C. Tsamis, and R. P. Woodard. de Sitter breaking through infrared divergences. Journal of Mathematical Physics, 51(7): 072503, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p072503_s1.

Kijowski:2010:HFD

[324] J. Kijowski, G. Magli, and D. Malafarina. The Hamiltonian formulation for the dynamics of a multishell self-gravitating system. Journal of Mathematical Physics, 51(7):072504, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p072504_s1.

Baykal:2010:CSV

[325] Ahmet Baykal, Dilek K. Çiftci, and Özgür Delice. Cylindrically symmetric vacuum solutions in higher dimensional Brans–Dicke theory. Journal of Mathematical Physics, 51(7):072505, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p072505_s1.

Ding:2010:CSH

[326] Danping Ding and Peng Lv. Conservative solutions for higher-order Camassa–Holm equations. Journal of Mathematical Physics, 51(7): 072701, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p072701_s1.

vonWaldenfels:2010:PSG

[327] Wilhelm von Waldenfels. Photon scattering in geometrical optics. Journal of Mathematical Physics, 51(7):072901, July 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i7/p072901_s1.

Stenlund:2010:EHS

[328] Mikko Stenlund. An expansion of the homoclinic splitting matrix for the rapidly, quasiperiodically, forced pendulum. Journal of Mathematical Physics, 51(7):072902, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p072902_s1.

Marquette:2010:CCS

[329] Ian Marquette. Construction of classical superintegrable systems with higher order integrals of motion from ladder operators. Journal of Mathematical Physics, 51(7):072903, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i7/p072903_s1.

Rodriguez:2010:GCE

[330] A. Rodríguez and C. Tsallis. A generalization of the cumulant expansion. Application to a scale-invariant probabilistic model. Journal of Mathematical Physics, 51(7):073301, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i7/p073301_s1.

Poghosyan:2010:AEL

[331] Suren Poghosyan. Asymptotic expansion of the log-partition function for a gas of interacting Brownian loops. II. Journal of Mathematical Physics, 51(7):073302, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p073302_s1.

Cheng:2010:SSB

[332] Chi Honn Cheng. L^{p} stability of solutions of Boltzmann equation with external force in soft potentials. Journal of Mathematical Physics, 51(7): 073303, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p073303_s1.

Zhaqilao:2010:GCK

[333] Zhaqilao and Sirendaoreji. A generalized coupled Korteweg–de Vries hierarchy, bi-Hamiltonian structure, and Darboux transformation. Journal of Mathematical Physics, 51(7):073501, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p073501_s1.

Cheviakov:2010:LNR

[334] Alexei F. Cheviakov and George W. Bluman. On locally and nonlocally related potential systems. Journal of Mathematical Physics, 51(7): 073502, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p073502_s1.

Eghbali:2010:CTT

[335] A. Eghbali, A. Rezaei-Aghdam, and F. Heidarpour. Classification of two and three dimensional Lie superbialgebras. Journal of Mathematical Physics, 51(7):073503, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p073503_s1.

Gungor:2010:IDS

[336] F. Güngör. Infinite-dimensional symmetries of two-dimensional generalized Burgers equations. Journal of Mathematical Physics, 51(7):073504, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p073504_s1.

Ma:2010:ENK

[337] Wen-Xiu Ma. An extended noncommutative KP hierarchy. Journal of Mathematical Physics, 51(7):073505, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p073505_s1.

Kafini:2010:DRV

[338] Mohammad Kafini and Nasser eddine Tatar. A decay result to a viscoelastic problem in R^{n} with an oscillating kernel. Journal of Mathematical Physics, 51(7):073506, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i7/p073506_s1.

Munteanu:2010:GSC

[339] Marian Ioan Munteanu. From golden spirals to constant slope surfaces. Journal of Mathematical Physics, 51(7):073507, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i7/p073507_s1.

Park:2010:GDN

[340] Jong Yeoul Park and Sun Hye Park. General decay for a nonlinear beam equation with weak dissipation. Journal of Mathematical Physics, 51(7): 073508, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i7/p073508_s1.

Hrivnak:2010:TDSb

[341] Jirí Hrivnák, Lenka Motlochová, and Jirí Patera. Two-dimensional symmetric and antisymmetric generalizations of sine functions. Journal of Mathematical Physics, 51(7):073509, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i7/p073509_s1.

Alexandrov:2010:SDE

[342] Sergei Alexandrov, Boris Pioline, and Stefan Vandoren. Self-dual Einstein spaces, heavenly metrics, and twistors. Journal of Mathematical Physics, 51(7):073510, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i7/p073510_s1.

Martin:2010:SLS

[343] Antonio J. Calderón Martín and José M. Sánchez Delgado. On split Lie superalgebras. Journal of Mathematical Physics, 51(7):073511, July 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i7/p073511_s1.

Cabrera:2010:CCD

[344] Renan Cabrera, Traci Strohecker, and Herschel Rabitz. The canonical coset decomposition of unitary matrices through Householder transformations. Journal of Mathematical Physics, 51(8):082101, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p082101_s1.

Ertik:2010:TFD

[345] Hüseyin Ertik, Dogan Demirhan, Hüseyin Sirin, and Fevzi Büyükkiliç. Time fractional development of quantum systems. Journal of Mathematical Physics, 51(8):082102, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i8/p082102_s1.

Serrano:2010:QDP

[346] F. A. Serrano, Xiao-Yan Gu, and Shi-Hai Dong. Qiang–Dong proper quantization rule and its applications to exactly solvable quantum systems. Journal of Mathematical Physics, 51(8):082103, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p082103_s1.

Privault:2010:SDI

[347] Nicolas Privault and Jean-Claude Zambrini. Stochastic deformation of integrable dynamical systems and random time symmetry. Journal of Mathematical Physics, 51(8):082104, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p082104_s1.

Fritz:2010:TPM

[348] Tobias Fritz. Transition probabilities and measurement statistics of postselected ensembles. Journal of Mathematical Physics, 51(8):082105, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p082105_s1.

Khrennikov:2010:QCD

[349] Andrei Khrennikov. Quantum correlations and dynamics from classical random fields valued in complex Hilbert spaces. Journal of Mathematical Physics, 51(8):082106, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082106_s1.

Miranda:2010:DTC

[350] P. A. Vidal Miranda and G. Mahler. Dynamical typicality: Convergence of time evolved macro-observables to their mean values in random matrix models. Journal of Mathematical Physics, 51(8):082107, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p082107_s1.

Mandayam:2010:TCA

[351] Prabha Mandayam, Stephanie Wehner, and Niranjan Balachandran. A transform of complementary aspects with applications to entropic uncertainty relations. Journal of Mathematical Physics, 51(8):082201, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p082201_s1.

Johnston:2010:FNA

[352] Nathaniel Johnston and David W. Kribs. A family of norms with applications in quantum information theory. Journal of Mathematical Physics, 51(8):082202, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082202_s1.

Fernando:2010:MUR

[353] Sudarshan Fernando and Murat Günaydin. Minimal unitary representation of SU(2, 2) and its deformations as massless conformal fields and their supersymmetric extensions. Journal of Mathematical Physics, 51(8):082301, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082301_s1.

Mestre:2010:CPI

[354] Ângela Mestre. Combinatorics of 1-particle irreducible n-point functions via coalgebra in quantum field theory. Journal of Mathematical Physics, 51(8):082302, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082302_s1.

Jakubassa-Amundsen:2010:SGS

[355] D. H. Jakubassa-Amundsen. Scaling of the ground-state energy of relativistic ions in high locally bounded magnetic fields. Journal of Mathematical Physics, 51(8):082303, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i8/p082303_s1.

Schiappa:2010:TMM

[356] Ricardo Schiappa and Niclas Wyllard. An A_{r} threesome: Matrix models, 2d conformal field theories, and 4d N = 2 gauge theories. Journal of Mathematical Physics, 51(8):082304, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p082304_s1.

Dogan:2010:RIR

[357] Çağlar Doğan and O. Teoman Turgut. Renormalized interaction of relativistic bosons with delta function potentials. Journal of Mathematical Physics, 51(8):082305, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082305_s1.

Falco:2010:VAA

[358] Pierluigi Falco. Vector and axial anomaly in the Thirring–Wess model. Journal of Mathematical Physics, 51(8):082306, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/p082306_s1.

Rothe:2010:BSB

[359] Tilman J. Rothe and Gerhard Schäfer. Binary spinning black hole Hamiltonian in canonical center-of-mass and rest-frame coordinates through higher post-Newtonian order. Journal of Mathematical Physics, 51(8): 082501, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p082501_s1.

Freidel:2010:FSI

[360] Laurent Freidel and Etera R. Livine. The fine structure of SU(2) intertwiners from U(N) representations. Journal of Mathematical Physics, 51(8):082502, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082502_s1.

Yurov:2010:FVA

[361] A. V. Yurov and V. A. Yurov. Friedman versus Abel equations: a connection unraveled. Journal of Mathematical Physics, 51(8):082503, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p082503_s1.

Santacruz:2010:ECA

[362] C. Angulo Santacruz, D. Batic, and M. Nowakowski. On the existence of certain axisymmetric interior metrics. Journal of Mathematical Physics, 51(8):082504, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082504_s1.

Gu:2010:CSP

[363] Lei Gu, Xiao-Dong Zhang, and Qing Zhou. Consensus and synchronization problems on small-world networks. Journal of Mathematical Physics, 51(8):082701, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082701_s1.

Habraken:2010:GPA

[364] Steven J. M. Habraken and Gerard Nienhuis. Geometric phases in astigmatic optical modes of arbitrary order. Journal of Mathematical Physics, 51(8):082702, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p082702_s1.

Anzaldo-Meneses:2010:FPS

[365] A. Anzaldo-Meneses and F. Monroy-Pérez. Foucault pendulum and sub-Riemannian geometry. Journal of Mathematical Physics, 51(8): 082703, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p082703_s1.

Llibre:2010:POI

[366] Jaume Llibre, Ricardo Miranda Martins, and Marco Antonio Teixeira. Periodic orbits, invariant tori, and cylinders of Hamiltonian systems near integrable ones having a return map equal to the identity. Journal of Mathematical Physics, 51(8):082704, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p082704_s1.

Szpak:2010:LTA

[367] Nikodem Szpak. Late-time attractor for the cubic nonlinear wave equation. Journal of Mathematical Physics, 51(8):082901, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p082901_s1.

Kaparulin:2010:RSC

[368] D. S. Kaparulin, S. L. Lyakhovich, and A. A. Sharapov. Rigid symmetries and conservation laws in non-Lagrangian field theory. Journal of Mathematical Physics, 51(8):082902, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p082902_s1.

Ludu:2010:ATV

[369] A. Ludu, J. Van Deun, M. V. Milosević, A. Cuyt, and F. M. Peeters. Analytic treatment of vortex states in cylindrical superconductors in applied axial magnetic field. Journal of Mathematical Physics, 51(8): 082903, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p082903_s1.

Pu:2010:EDS

[370] Xueke Pu and Boling Guo. Existence and decay of solutions to the two-dimensional fractional quasigeostrophic equation. Journal of Mathematical Physics, 51(8):083101, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i8/p083101_s1.

vanderToorn:2010:AMD

[371] Ramses van der Toorn and Joseph T. F. Zimmerman. Angular momentum dynamics and the intrinsic drift of monopolar vortices on a rotating sphere. Journal of Mathematical Physics, 51(8):083102, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083102_s1.

Dudnikova:2010:LDH

[372] T. V. Dudnikova. Lattice dynamics in the half-space. Energy transport equation. Journal of Mathematical Physics, 51(8):083301, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083301_s1.

Gumral:2010:GPD

[373] Hasan Gümral. Geometry of plasma dynamics. I. Group of canonical diffeomorphisms. Journal of Mathematical Physics, 51(8):083501, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083501_s1.

Odake:2010:UTE

[374] Satoru Odake and Ryu Sasaki. Unified theory of exactly and quasiexactly solvable “discrete” quantum mechanics. I. Formalism. Journal of Mathematical Physics, 51(8):083502, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p083502_s1.

Otte:2010:BPF

[375] Peter Otte. Boundedness properties of fermionic operators. Journal of Mathematical Physics, 51(8):083503, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p083503_s1.

Coulembier:2010:OIF

[376] K. Coulembier, H. De Bie, and F. Sommen. Orthosymplectically invariant functions in superspace. Journal of Mathematical Physics, 51(8): 083504, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p083504_s1.

Yang:2010:NSH

[377] Xiao Yang and Dianlou Du. From nonlinear Schrödinger hierarchy to some (2+1)-dimensional nonlinear pseudodifferential equations. Journal of Mathematical Physics, 51(8):083505, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p083505_s1.

Zhang:2010:CIC

[378] Yufeng Zhang and Engui Fan. Coupling integrable couplings and bi-Hamiltonian structure associated with the Boiti–Pempinelli–Tu hierarchy. Journal of Mathematical Physics, 51(8):083506, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p083506_s1.

Hussin:2010:SKV

[379] V. Hussin, A. V. Kiselev, A. O. Krutov, and T. Wolf. N = 2 supersymmetric a = 4-Korteweg–de Vries hierarchy derived via Gardner’s deformation of Kaup–Boussinesq equation. Journal of Mathematical Physics, 51(8):083507, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p083507_s1.

Dvurecenskij:2010:ESI

[380] Anatolij Dvurecenskij. Every state on interval effect algebra is integral. Journal of Mathematical Physics, 51(8):083508, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083508_s1.

Wu:2010:VCS

[381] Yuezhu Wu. Vector coherent state realization of representations of the affine superalgebra A^{(1)} (0,1). Journal of Mathematical Physics, 51(8): 083509, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p083509_s1.

Morillas:2010:COW

[382] Patricia Mariela Morillas. Construction of orthonormal wavelet-like bases. Journal of Mathematical Physics, 51(8):083510, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p083510_s1.

Aydin:2010:MBS

[383] A. Aydin and B. Karasözen. Multisymplectic box schemes for the complex modified Korteweg–de Vries equation. Journal of Mathematical Physics, 51(8):083511, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p083511_s1.

Rumanov:2010:USU

[384] Igor Rumanov. Universal structure and universal equations (PDE) for unitary ensembles. Journal of Mathematical Physics, 51(8):083512, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083512_s1.

Li:2010:CIM

[385] Junbo Li, Yucai Su, and Linsheng Zhu. Classification of indecomposable modules of the intermediate series over the twisted N = 2 superconformal algebra. Journal of Mathematical Physics, 51(8):083513, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083513_s1.

Chen:2010:FGS

[386] Jinbing Chen. Finite-gap solutions of 2 + 1 dimensional integrable nonlinear evolution equations generated by the Neumann systems. Journal of Mathematical Physics, 51(8):083514, August 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083514_s1.

Gromov:2010:PQK

[387] N. A. Gromov. Possible quantum kinematics. II. Nonminimal case. Journal of Mathematical Physics, 51(8):083515, August 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083515_s1.

Skrypnyk:2010:IDG

[388] T. Skrypnyk. Isomonodromic deformations, generalized Knizhnik– Zamolodchikov equations and non-skew-symmetric classical r-matrices. Journal of Mathematical Physics, 51(8):083516, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083516_s1.

Athmouni:2010:CSF

[389] Nassim Athmouni, Marius Măntoiu, and Radu Purice. On the continuity of spectra for families of magnetic pseudodifferential operators. Journal of Mathematical Physics, 51(8):083517, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i8/p083517_s1.

Chen:2010:CCP

[390] Bang-Yen Chen. Complete classification of parallel Lorentz surfaces in four-dimensional neutral pseudosphere. Journal of Mathematical Physics, 51(8):083518, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i8/p083518_s1.

Colombo:2010:OCU

[391] Leonardo Colombo, David Martín De Diego, and Marcela Zuccalli. Optimal control of underactuated mechanical systems: a geometric approach. Journal of Mathematical Physics, 51(8):083519, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083519_s1.

Hitzazis:2010:KVE

[392] Iasonas Hitzazis and Dimitri Tsoubelis. The Korteweg–de Vries equation on the interval. Journal of Mathematical Physics, 51(8):083520, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p083520_s1.

Cafuta:2010:NNS

[393] Kristijan Cafuta, Igor Klep, and Janez Povh. A note on the nonexistence of sum of squares certificates for the Bessis–Moussa–Villani conjecture. Journal of Mathematical Physics, 51(8):083521, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/p083521_s1.

Kobayashi:2010:CBS

[394] Masanori Kobayashi, Marc de Montigny, and Faqir C. Khanna. A connection between spin and statistics based on Galilean invariant delta functions. Journal of Mathematical Physics, 51(8):083522, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i8/
p083522_s1.

Shi:2010:UBN

[395] Dai Shi. Upper bounds on the number of eigenvalues of stationary Schrödinger operators. Journal of Mathematical Physics, 51(8):083523, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p083523_s1.

Zhang:2010:WMS

[396] Xiufu Zhang, Shaobin Tan, and Haifeng Lian. Whittaker modules for the Schrödinger–Witt algebra. Journal of Mathematical Physics, 51(8): 083524, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i8/p083524_s1.

Lv:2010:EEP

[397] Cui hong Lv and Hong yi Fan. Erratum: “Explicit pure-state density operator structure for quantum tomography”, [J. Math. Phys. 50, 102108 (2009)]. Journal of Mathematical Physics, 51(8):089901, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p089901_s1. See [15].

Ohno:2010:MRE

[398] Hiromichi Ohno. Maximal rank of extremal marginal tracial states. Journal of Mathematical Physics, 51(9):092101, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092101_s1.

Lenzi:2010:CTR

[399] E. K. Lenzi, H. V. Ribeiro, H. Mukai, and R. S. Mendes. Continuoustime random walk as a guide to fractional Schrödinger equation. Journal of Mathematical Physics, 51(9):092102, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092102_s1.

Levai:2010:ESS

[400] G. Lévai and O. Özer. An exactly solvable Schrödinger equation with finite positive position-dependent effective mass. Journal of Mathematical Physics, 51(9):092103, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i9/p092103_s1.

Han:2010:SFG

[401] Gang Han. The symmetries of the fine gradings of sl(n^{k}, C) associated with direct product of Pauli groups. Journal of Mathematical Physics, 51(9):092104, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p092104_s1.

Cacciari:2010:SIE

[402] Ilaria Cacciari, Marco Lantieri, and Paolo Moretti. On the solution of the integral equation of scattering for finite range potentials. Journal of Mathematical Physics, 51(9):092105, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092105_s1.

Dubrovsky:2010:TDS

[403] V. G. Dubrovsky, A. V. Topovsky, and M. Yu. Basalaev. Twodimensional stationary Schrödinger equation via the ∂-dressing method: New exactly solvable potentials, wave functions, and their physical interpretation. Journal of Mathematical Physics, 51(9):092106, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p092106_s1.

Xu:2010:EDC

[404] Ye jun Xu, Jun Song, Hong yi Fan, and Qiu yu Liu. Eigenkets of the q-deformed creation operator. Journal of Mathematical Physics, 51(9):092107, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p092107_s1.

Horvathy:2010:SPD

[405] Peter A. Horváthy, Mikhail S. Plyushchay, and Mauricio Valenzuela. Supersymmetry of the planar Dirac–Deser–Jackiw–Templeton system and of its nonrelativistic limit. Journal of Mathematical Physics, 51 (9):092108, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p092108_s1.

Chin:2010:EMB

[406] Alex W. Chin, Ángel Rivas, Susana F. Huelga, and Martin B. Plenio. Exact mapping between system-reservoir quantum models and semi-infinite discrete chains using orthogonal polynomials. Journal of Mathematical Physics, 51(9):092109, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i9/p092109_s1.

Heinosaari:2010:NQM

[407] Teiko Heinosaari and Michael M. Wolf. Nondisturbing quantum measurements. Journal of Mathematical Physics, 51(9):092201, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p092201_s1.

Kintas:2010:TTE

[408] S. Kintas and S. Turgut. Transformations of W -type entangled states. Journal of Mathematical Physics, 51(9):092202, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/
p092202_s1.

Wei:2010:MPQ

[409] Tzu-Chieh Wei and Simone Severini. Matrix permanent and quantum entanglement of permutation invariant states. Journal of Mathematical Physics, 51(9):092203, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i9/p092203_s1.

Tyson:2010:TSB

[410] Jon Tyson. Two-sided bounds on minimum-error quantum measurement, on the reversibility of quantum dynamics, and on maximum overlap using directional iterates. Journal of Mathematical Physics, 51(9):092204, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p092204_s1.

Chu:2010:CSA

[411] Yan-Jun Chu, Fang Huang, and Zhu-Jun Zheng. A commutant of βγsystem associated to the highest weight module V_{4} of sl(2, C). Journal of Mathematical Physics, 51(9):092301, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092301_s1.

Coquereaux:2010:EQS

[412] R. Coquereaux, R. Rais, and E. H. Tahri. Exceptional quantum subgroups for the rank two Lie algebras B_{2} and G_{2}. Journal of Mathematical Physics, 51(9):092302, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i9/p092302_s1.

Yilmaz:2010:ILF

[413] Nejat T. Yilmaz. Integrated Lax formalism for principal chiral model. Journal of Mathematical Physics, 51(9):092303, September 2010. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/
p092303_s1.

Rivasseau:2010:LVE

[414] Vincent Rivasseau and Zhituo Wang. Loop vertex expansion for Φ^{2k} theory in zero dimension. Journal of Mathematical Physics, 51(9):092304, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p092304_s1.

Han:2010:RBR

[415] Muxin Han and Thomas Thiemann. On the relation between rigging inner product and master constraint direct integral decomposition. Journal of Mathematical Physics, 51(9):092501, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092501_s1.

Sonego:2010:UST

[416] Sebastiano Sonego. Ultrastatic space–times. Journal of Mathematical Physics, 51(9):092502, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i9/p092502_s1.

Foth:2010:PGS

[417] Philip Foth and McKenzie Lamb. The Poisson geometry of SU(1,1). Journal of Mathematical Physics, 51(9):092701, September 2010. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/
p092701_s1.

Llibre:2010:BVM

[418] Jaume Llibre and Clàudia Valls. The Bianchi VIII model is neither global analytic nor Darboux integrable. Journal of Mathematical Physics, 51(9):092702, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p092702_s1.

Zhijian:2010:FDA

[419] Yang Zhijian. Finite-dimensional attractors for the Kirchhoff models. Journal of Mathematical Physics, 51(9):092703, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/
p092703_s1.

Yu:2010:DFC

[420] Shengqi Yu and Mingxin Wang. On a dissipative form of Camassa–Holm equation. Journal of Mathematical Physics, 51(9):092704, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p092704_s1.

Bonnard:2010:EMP

[421] B. Bonnard, O. Cots, N. Shcherbakova, and D. Sugny. The energy minimization problem for two-level dissipative quantum systems. Journal of Mathematical Physics, 51(9):092705, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p092705_s1.

Cross:2010:EDE

[422] Daniel J. Cross and R. Gilmore. Equivariant differential embeddings. Journal of Mathematical Physics, 51(9):092706, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/
p092706_s1.

Seredynska:2010:RDA

[423] M. Seredyńska and Andrzej Hanyga. Relaxation, dispersion, attenuation, and finite propagation speed in viscoelastic media. Journal of Mathematical Physics, 51(9):092901, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i9/p092901_s1.

Zhao:2010:GWP

[424] Jihong Zhao, Chao Deng, and Shangbin Cui. Global well-posedness of a dissipative system arising in electrohydrodynamics in negative-order Besov spaces. Journal of Mathematical Physics, 51(9):093101, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093101_s1.

Hamza:2010:EVS

[425] A. M. Hamza. On the existence of vortexlike solutions in the auroral F-region. Journal of Mathematical Physics, 51(9):093102, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093102_s1.

Sulaiman:2010:GEU

[426] Samira Sulaiman. Global existence and uniqueness for a nonlinear Boussinesq system in dimension two. Journal of Mathematical Physics, 51(9):093103, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093103_s1.

Hasegawa:2010:EFU

[427] Hideo Hasegawa. The entropy in finite N-unit nonextensive systems: The normal average and q-average. Journal of Mathematical Physics, 51(9):093301, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093301_s1.

Dumitriu:2010:TRA

[428] Ioana Dumitriu and Peter J. Forrester. Tridiagonal realization of the antisymmetric Gaussian β-ensemble. Journal of Mathematical Physics, 51(9):093302, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093302_s1.

Wang:2010:SSS

[429] Yao Wang and Lawrence E. Thomas. Stationary state solutions for a gently stochastic nonlinear wave equation with ultraviolet cutoffs. Journal of Mathematical Physics, 51(9):093303, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/p093303_s1.

Rogers:2010:USP

[430] Tim Rogers. Universal sum and product rules for random matrices. Journal of Mathematical Physics, 51(9):093304, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v51/i9/p093304_s1.

Cheng:2010:FCP

[431] Ji-Peng Cheng, Shi-Kun Wang, Ke Wu, and Wei-Zhong Zhao. Fermionic covariant prolongation structure theory for supernonlinear evolution equation. Journal of Mathematical Physics, 51(9):093501, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093501_s1.

Chevreuil:2010:CAD

[432] A. Chevreuil, A. Plastino, and C. Vignat. On a conjecture about Dirac’s delta representation using q-exponentials. Journal of Mathematical Physics, 51(9):093502, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i9/p093502_s1.

Brugarino:2010:IIN

[433] Tommaso Brugarino and Michele Sciacca. Integrability of an inhomogeneous nonlinear Schrödinger equation in Bose–Einstein condensates and fiber optics. Journal of Mathematical Physics, 51(9):093503, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093503_s1.

Mathur:2010:ISB

[434] Manu Mathur, Indrakshi Raychowdhury, and Ramesh Anishetty. SU(N) irreducible Schwinger bosons. Journal of Mathematical Physics, 51 (9):093504, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093504_s1.

Nuida:2010:OOM

[435] Koji Nuida, Gen Kimura, and Takayuki Miyadera. Optimal observables for minimum-error state discrimination in general probabilistic theories. Journal of Mathematical Physics, 51(9):093505, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093505_s1.

Kotlyarov:2010:RHP

[436] Vladimir Kotlyarov and Alexander Minakov. Riemann–Hilbert problem to the modified Korteweg–de Vries equation: Long-time dynamics of the steplike initial data. Journal of Mathematical Physics, 51(9):093506, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p093506_s1.

Chen:2010:ESD

[437] Shouxin Chen and Ying Wang. The existence of self-dual vortices in a non-Abelian Φ^{2} Chern–Simons theory. Journal of Mathematical Physics, 51(9):093507, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093507_s1.

Xu:2010:SGM

[438] Tian Zhou Xu, John Michael Rassias, and Wan Xin Xu. Stability of a general mixed additive-cubic functional equation in non-Archimedean fuzzy normed spaces. Journal of Mathematical Physics, 51(9):093508, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p093508_s1.

Fu:2010:ESS

[439] Yuxia Fu and Zhaoyang Yin. Existence and singularities of solutions to an integrable equation governing short-waves in a long-wave model. Journal of Mathematical Physics, 51(9):093509, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p093509_s1.

He:2010:HCS

[440] Chun-Lei He and Shou-Jun Huang. Harmonic coordinates in the string and membrane equations. Journal of Mathematical Physics, 51 (9):093510, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093510_s1.

Liu:2010:SRT

[441] Q. P. Liu, Ziemowit Popowicz, and Kai Tian. Supersymmetric reciprocal transformation and its applications. Journal of Mathematical Physics, 51(9):093511, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093511_s1.

Bravyi:2010:TQO

[442] Sergey Bravyi, Matthew B. Hastings, and Spyridon Michalakis. Topological quantum order: Stability under local perturbations. Journal of Mathematical Physics, 51(9):093512, September 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i9/p093512_s1.

Pomponio:2010:GSS

[443] A. Pomponio. Ground states for a system of nonlinear Schrödinger equations with three wave interaction. Journal of Mathematical Physics, 51(9):093513, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093513_s1.

Zhang:2010:FLA

[444] Yufeng Zhang and Honwah Tam. Four Lie algebras associated with R^{6} and their applications. Journal of Mathematical Physics, 51(9):093514, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p093514_s1.

Eleuch:2010:ASS

[445] H. Eleuch and Y. V. Rostovtsev. Analytical solution to sine-Gordon equation. Journal of Mathematical Physics, 51(9):093515, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093515_s1.

Bakke:2010:BSI

[446] K. Bakke. Bound states for an induced electric dipole in the presence of an azimuthal magnetic field and a disclination. Journal of Mathematical Physics, 51(9):093516, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i9/p093516_s1.

Arnal:2010:SSM

[447] Didier Arnal, Mabrouk Ben Ammar, and Bechir Dali. The spaces H^{n}(≀∫√(1|2), M) for some modules M. Journal of Mathematical Physics,

51(9):093517, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093517_s1.

Caprio:2010:RMG

[448] M. A. Caprio, K. D. Sviratcheva, and A. E. McCoy. Racah’s method for general subalgebra chains: Coupling coefficients of SO(5) in canonical and physical bases. Journal of Mathematical Physics, 51(9):093518, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i9/p093518_s1.

Jiang:2010:SBT

[449] Yan Jiang, Bo Tian, Wen-Jun Liu, Min Li, Pan Wang, and Kun Sun. Solitons, Bäcklund transformation, and Lax pair for the (2 + 1)-dimensional Boiti–Leon–Pempinelli equation for the water waves. Journal of Mathematical Physics, 51(9):093519, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i9/p093519_s1. See [2620].

Soloviev:2010:RQF

[450] M. A. Soloviev. Reconstruction in quantum field theory with a fundamental length. Journal of Mathematical Physics, 51(9):093520, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093520_s1.

Boyallian:2010:QOT

[451] Carina Boyallian and Vanesa Meinardi. QHWM of the orthogonal type Lie subalgebra of the Lie algebra of matrix differential operators on the circle. Journal of Mathematical Physics, 51(9):093521, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v51/i9/p093521_s1. See publisher’s note [510].

Xu:2010:IBQ

[452] Xiao-Ming Xu, Hong-Ke Du, and Xiaochun Fang. On the infimum of bounded quantum observables. Journal of Mathematical Physics, 51(9):093522, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093522_s1.

Stoilova:2010:GZB

[453] N. I. Stoilova and J. Van der Jeugt. Gel’fand–Zetlin basis and Clebsch– Gordan coefficients for covariant representations of the Lie superalgebra }↕(m|n). Journal of Mathematical Physics, 51(9):093523, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093523_s1.

Yuen:2010:SSB

[454] Manwai Yuen. Self-similar blowup solutions to the 2-component Camassa–Holm equations. Journal of Mathematical Physics, 51(9): 093524, September 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i9/p093524_s1. See erratum [899].

Lieb:2010:LMW

[455] Elliott H. Lieb and Yaron Ostrover. Localization of multidimensional Wigner distributions. Journal of Mathematical Physics, 51(10):102101, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p102101_s1.

Albeverio:2010:STS

[456] Sergio Albeverio, Aleksey Kostenko, and Mark Malamud. Spectral theory of semibounded Sturm–Liouville operators with local interactions on a discrete set. Journal of Mathematical Physics, 51(10):102102, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p102102_s1.

Pavicic:2010:GAQ

[457] Mladen Pavicić, Brendan D. McKay, Norman D. Megill, and Kresimir Fresl. Graph approach to quantum systems. Journal of Mathematical Physics, 51(10):102103, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p102103_s1.

Hiscox:2010:LEB

[458] A. Hiscox, B. M. Brown, and M. Marletta. On the low energy behavior of Regge poles. Journal of Mathematical Physics, 51(10):102104, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p102104_s1.

Marquette:2010:GMK

[459] Ian Marquette. Generalized MICZ-Kepler system, duality, polynomial, and deformed oscillator algebras. Journal of Mathematical Physics, 51(10):102105, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i10/p102105_s1.

Castello-Branco:2010:FFU

[460] K. H. C. Castello-Branco and A. G. Martins. Free-fall in a uniform gravitational field in noncommutative quantum mechanics. Journal of Mathematical Physics, 51(10):102106, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p102106_s1.

Abbas:2010:DRS

[461] Tasawar Abbas and Farhan Saif. Dynamical revivals in spatiotemporal evolution of driven one dimensional box. Journal of Mathematical Physics, 51(10):102107, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p102107_s1.

Hounkonnou:2010:HOT

[462] Mahouton Norbert Hounkonnou and Dine Ousmane Samary. Harmonic oscillator in twisted Moyal plane: Eigenvalue problem and relevant properties. Journal of Mathematical Physics, 51(10):102108, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p102108_s1.

Buscemi:2010:DEA

[463] Francesco Buscemi and Nilanjana Datta. Distilling entanglement from arbitrary resources. Journal of Mathematical Physics, 51(10):102201, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p102201_s1.

Guta:2010:QUS

[464] Madalin Guta and Cristina Butucea. Quantum U-statistics. Journal of Mathematical Physics, 51(10):102202, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p102202_s1.

Stormer:2010:TPP

[465] Erling Størmer. Tensor powers of 2-positive maps. Journal of Mathematical Physics, 51(10):102203, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p102203_s1.

Gawron:2010:RNR

[466] Piotr Gawron, Zbigniew Puchala, Jaroslaw Adam Miszczak, Lukasz Skowronek, and Karol Zyczkowski. Restricted numerical range: a versatile tool in the theory of quantum information. Journal of Mathematical Physics, 51(10):102204, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p102204_s1.

Bonanos:2010:DMS

[467] Sotirios Bonanos, Joaquim Gomis, Kiyoshi Kamimura, and Jerzy Lukierski. Deformations of Maxwell superalgebras and their applications. Journal of Mathematical Physics, 51(10):102301, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p102301_s1.

deCastro:2010:DKP

[468] A. S. de Castro. On Duffin–Kemmer–Petiau particles with a mixed minimal-nonminimal vector coupling and the nondegenerate boundstates for the one-dimensional inversely linear background. Journal of Mathematical Physics, 51(10):102302, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p102302_s1.

Hosseiny:2010:LCN

[469] Ali Hosseiny and Shahin Rouhani. Logarithmic correlators in nonrelativistic conformal field theory. Journal of Mathematical Physics, 51(10): 102303, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p102303_s1.

deAguiar:2010:SQR

[470] T. C. de Aguiar, N. F. Svaiter, and G. Menezes. Stochastic quantization of real-time thermal field theory. Journal of Mathematical Physics, 51(10):102304, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i10/p102304_s1.

DeGroote:2010:PTP

[471] Liselotte De Groote, Norbert Van den Bergh, and Lode Wylleman. Petrov type D pure radiation fields of Kundt’s class. Journal of Mathematical Physics, 51(10):102501, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p102501_s1.

Cacciatori:2010:IA

[472] Sergio L. Cacciatori, Bianca L. Cerchiai, and Alessio Marrani. Iwasawa N = 8 attractors. Journal of Mathematical Physics, 51(10):102502, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p102502_s1.

Caciotta:2010:CGC

[473] Giulio Caciotta and Francesco Nicolò. On a class of global characteristic problems for the Einstein vacuum equations with small initial data. Journal of Mathematical Physics, 51(10):102503, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p102503_s1.

Thierry:2010:TDS

[474] Daudé Thierry. Time-dependent scattering theory for charged Dirac fields on a Reissner–Nordström black hole. Journal of Mathematical Physics, 51(10):102504, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p102504_s1.

Yang:2010:NHO

[475] Minbo Yang. Nonstationary homoclinic orbits for an infinite-dimensional Hamiltonian system. Journal of Mathematical Physics, 51(10):102701, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p102701_s1.

Chueshov:2010:MSS

[476] Igor Chueshov and Björn Schmalfuß. Master-slave synchronization and invariant manifolds for coupled stochastic systems. Journal of Mathematical Physics, 51(10):102702, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p102702_s1.

Li:2010:NSC

[477] Liangliang Li, Yuanlong Chen, and Yu Huang. Nonisotropic spatiotemporal chaotic vibrations of the one-dimensional wave equation with a mixing transport term and general nonlinear boundary condition. Journal of Mathematical Physics, 51(10):102703, October 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i10/p102703_s1.

Misiurewicz:2010:SDP

[478] Michal Misiurewicz. Sensitive dependence on parameters. Journal of Mathematical Physics, 51(10):102704, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p102704_s1.

Qin:2010:DCO

[479] Wen-Xin Qin and Chun-Lan Xu. Dynamics of coupled oscillators with convex interaction potential. Journal of Mathematical Physics, 51(10): 102705, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p102705_s1.

Ndangali:2010:EBS

[480] Rémy F. Ndangali and Sergei V. Shabanov. Electromagnetic bound states in the radiation continuum for periodic double arrays of subwavelength dielectric cylinders. Journal of Mathematical Physics, 51(10): 102901, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p102901_s1.

Cichocki:2010:GTD

[481] Bogdan Cichocki and Maria L. Ekiel-Jezewska. Green tensors for Debye– Büche–Brinkman equations generalized for axisymmetric medium. Journal of Mathematical Physics, 51(10):103101, October 2010. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p103101_s1.

Ahn:2010:SFD

[482] Shin Mi Ahn and Seung-Yeal Ha. Stochastic flocking dynamics of the Cucker–Smale model with multiplicative white noises. Journal of Mathematical Physics, 51(10):103301, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p103301_s1.

Jaeck:2010:NBE

[483] Thomas Jaeck, Joseph V. Pulé, and Valentin A. Zagrebnov. On the nature of Bose–Einstein condensation enhanced by localization. Journal of Mathematical Physics, 51(10):103302, October 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i10/p103302_s1.

Viennot:2010:HPC

[484] David Viennot. Holonomy of a principal composite bundle connection, non-Abelian geometric phases, and gauge theory of gravity. Journal of Mathematical Physics, 51(10):103501, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p103501_s1.

Constantinescu:2010:ASS

[485] Radu Constantinescu, Nick Costanzino, Anna L. Mazzucato, and Victor Nistor. Approximate solutions to second order parabolic equations. I: Analytic estimates. Journal of Mathematical Physics, 51(10):103502, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p103502_s1.

Eghbali:2010:CFS

[486] A. Eghbali, A. Rezaei-Aghdam, and F. Heidarpour. Classification of four and six dimensional Drinfel’d superdoubles. Journal of Mathematical Physics, 51(10):103503, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p103503_s1.

Le:2010:SWE

[487] Út V. Lê. A semilinear wave equation with space–time dependent coefficients and a memory boundarylike antiperiodic condition: Regularity and stability. Journal of Mathematical Physics, 51(10):103504, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p103504_s1.

Heo:2010:RNT

[488] Jaeseong Heo and Un Cig Ji. Radon–Nikodým type theorem for αcompletely positive maps. Journal of Mathematical Physics, 51(10): 103505, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p103505_s1.

Saxena:2010:SCF

[489] R. K. Saxena, A. M. Mathai, and H. J. Haubold. Solutions of certain fractional kinetic equations and a fractional diffusion equation. Journal of Mathematical Physics, 51(10):103506, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p103506_s1.

Geng:2010:CDM

[490] Qiaozhi Geng, Mingming Ren, and Zhiqi Chen. Classification of (n + 3)- dimensional metric n-Lie algebras. Journal of Mathematical Physics, 51(10):103507, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i10/p103507_s1.

Ebadian:2010:FPM

[491] A. Ebadian, N. Ghobadipour, and M. Eshaghi Gordji. A fixed point method for perturbation of bimultipliers and Jordan bimultipliers in C^{∗}- ternary algebras. Journal of Mathematical Physics, 51(10):103508, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p103508_s1.

Hussin:2010:CSA

[492] V. Hussin, I. Yurdusen, and W. J. Zakrzewski. Canonical surfaces associated with projectors in Grassmannian sigma models. Journal of Mathematical Physics, 51(10):103509, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p103509_s1.

Kanzieper:2010:NHW

[493] Eugene Kanzieper and Navinder Singh. Non-Hermitean Wishart random matrices (I). Journal of Mathematical Physics, 51(10):103510, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p103510_s1.

Feher:2010:TSS

[494] L. Fehér and V. Ayadi. Trigonometric Sutherland systems and their Ruijsenaars duals from symplectic reduction. Journal of Mathematical Physics, 51(10):103511, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p103511_s1.

Gorka:2010:FCL

[495] P. Górka, H. Prado, and E. G. Reyes. Functional calculus via Laplace transform and equations with infinitely many derivatives. Journal of Mathematical Physics, 51(10):103512, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p103512_s1.

Pradeep:2010:NCB

[496] R. Gladwin Pradeep, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. A nonlocal connection between certain linear and nonlinear ordinary differential equations: Extension to coupled equations. Journal of Mathematical Physics, 51(10):103513, October 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i10/p103513_s1.

Messaoudi:2010:GBS

[497] Salim A. Messaoudi and Aisha Al-Shehri. General boundary stabilization of memory-type thermoelasticity. Journal of Mathematical Physics, 51(10):103514, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i10/p103514_s1.

Cross:2010:SDT

[498] Daniel J. Cross and Robert Gilmore. A Schwinger disentangling theorem. Journal of Mathematical Physics, 51(10):103515, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p103515_s1.

Shcherbina:2010:UBL

[499] Tatyana Shcherbina. On universality of bulk local regime of the Hermitian sample covariance matrices. Journal of Mathematical Physics, 51(10):103516, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i10/p103516_s1.

Hounkonnou:2010:EFR

[500] Mahouton Norbert Hounkonnou and Elvis Benzo Ngompe Nkouankam. (q, μ) and (p, q, ζ)-exponential functions: Rogers–Szegő polynomials and Fourier–Gauss transform. Journal of Mathematical Physics, 51(10): 103517, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p103517_s1.

Vakhnenko:2010:IST

[501] Oleksiy O. Vakhnenko. Inverse scattering transform for the nonlinear Schrödinger system on a zigzag-runged ladder lattice. Journal of Mathematical Physics, 51(10):103518, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p103518_s1.

Perline:2010:CVF

[502] Ron Perline. A class of vortex filament solitons in fluids, plasmas, and superconductors. Journal of Mathematical Physics, 51(10):103519, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i10/p103519_s1.

Roschy:2010:CLB

[503] Helmut Roschy and Thorsten Rahn. Cohomology of line bundles: Proof of the algorithm. Journal of Mathematical Physics, 51(10):103520, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v51/i10/p103520_s1. See publisher’s note [599].

Cheviakov:2010:MPDa

[504] Alexei F. Cheviakov and George W. Bluman. Multidimensional partial differential equation systems: Generating new systems via conservation laws, potentials, gauges, subsystems. Journal of Mathematical Physics, 51(10):103521, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p103521_s1.

Cheviakov:2010:MPDb

[505] Alexei F. Cheviakov and George W. Bluman. Multidimensional partial differential equation systems: Nonlocal symmetries, nonlocal conservation laws, exact solutions. Journal of Mathematical Physics, 51(10): 103522, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i10/p103522_s1.

Wehbe:2010:EPS

[506] Ali Wehbe and Wael Youssef. Exponential and polynomial stability of an elastic Bresse system with two locally distributed feedbacks. Journal of Mathematical Physics, 51(10):103523, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p103523_s1.

Akemann:2010:IBA

[507] G. Akemann and M. Bender. Interpolation between Airy and Poisson statistics for unitary chiral non-Hermitian random matrix ensembles. Journal of Mathematical Physics, 51(10):103524, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i10/
p103524_s1.

Blumenhagen:2010:CLB

[508] Ralph Blumenhagen, Benjamin Jurke, Thorsten Rahn, and Helmut Roschy. Cohomology of line bundles: a computational algorithm. Journal of Mathematical Physics, 51(10):103525, October 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i10/p103525_s1. See

publisher’s note [600].

Song:2010:EST

[509] Xianfa Song. Erratum to “Sharp thresholds of global existence and blowup for a system of Schrödinger equations with combined power-type nonlinearities”, [J. Math. Phys. 51, 033509 (2010)]. Journal of Mathematical Physics, 51(10):109901, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i10/p109901_s1. See [143].

Boyallian:2010:PNQ

[510] Carina Boyallian and Vanesa Meinardi. Publisher’s note: “QHWM of the orthogonal type Lie subalgebra of the Lie algebra of matrix differential operators on the circle”, [J. Math. Phys. 51, 093521 (2010)]. Journal of Mathematical Physics, 51(10):109902, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i10/p109902_s1. See [451].

Kundu:2010:PNN

[511] Anjan Kundu. Publisher’s note: “Nonlinearizing linear equations to integrable systems including new hierarchies with nonholonomic deformations”, [J. Math. Phys. 50, 102702 (2009)]. Journal of Mathematical Physics, 51(10):109903, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p109903_s1. See [16].

Gudder:2010:QIA

[512] Stan Gudder. Quantum integrals and anhomomorphic logics. Journal of Mathematical Physics, 51(11):112101, November 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i11/p112101_s1.

Chakraborty:2010:TDR

[513] B. Chakraborty, Z. Kuznetsova, and F. Toppan. Twist deformation of rotationally invariant quantum mechanics. Journal of Mathematical Physics, 51(11):112102, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i11/p112102_s1.

Jencova:2010:SSA

[514] Anna Jencová. The structure of strongly additive states and Markov triplets on the CAR algebra. Journal of Mathematical Physics, 51 (11):112103, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112103_s1.

Barchielli:2010:JDU

[515] A. Barchielli and C. Pellegrini. Jump-diffusion unravelling of a non-Markovian generalized Lindblad master equation. Journal of Mathematical Physics, 51(11):112104, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i11/p112104_s1.

Gour:2010:AME

[516] Gilad Gour and Nolan R. Wallach. All maximally entangled four-qubit states. Journal of Mathematical Physics, 51(11):112201, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i11/
p112201_s1.

Bayrak:2010:RTS

[517] O. Bayrak, A. Soylu, and I. Boztosun. The relativistic treatment of spin-0 particles under the rotating Morse oscillator. Journal of Mathematical Physics, 51(11):112301, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i11/p112301_s1.

Brandt:2010:SACa

[518] Friedemann Brandt. Supersymmetry algebra cohomology. II. Primitive elements in 2 and 3 dimensions. Journal of Mathematical Physics, 51(11):112303, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112303_s1.

Schenck:2010:ECC

[519] Hal Schenck. Euler characteristic of coherent sheaves on simplicial torics via the Stanley–Reisner ring. Journal of Mathematical Physics, 51(11):112304, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112304_s1.

Eskin:2010:PKB

[520] G. Eskin. Perturbations of the Kerr black hole and the boundness of linear waves. Journal of Mathematical Physics, 51(11):112501, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i11/p112501_s1.

Schleich:2010:SPB

[521] Kristin Schleich and Donald M. Witt. A simple proof of Birkhoff’s theorem for cosmological constant. Journal of Mathematical Physics, 51(11):112502, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112502_s1.

VanGorder:2010:TDB

[522] Robert A. Van Gorder. Two-dimensional Blasius viscous flow of a powerlaw fluid over a semi-infinite flat plane. Journal of Mathematical Physics, 51(11):112901, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112901_s1.

Javidan:2010:CCV

[523] Kurosh Javidan. Collective coordinate variable for soliton-potential system in sine-Gordon model. Journal of Mathematical Physics, 51 (11):112902, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p112902_s1.

Nikolayenko:2010:TMF

[524] S. O. Nikolayenko and Yu. V. Slyusarenko. Theory of macroscopic fluctuations in systems of particles, interacting with hydrodynamic and gaslike media. Journal of Mathematical Physics, 51(11):113301, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i11/
p113301_s1.

Aristoff:2010:RCP

[525] David Aristoff and Charles Radin. Random close packing in a granular model. Journal of Mathematical Physics, 51(11):113302, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i11/p113302_s1.

Ferrari:2010:GRG

[526] Pablo A. Ferrari, Eugene A. Pechersky, Valentin V. Sisko, and Anatoly A. Yambartsev. Gibbs random graphs on point processes. Journal of Mathematical Physics, 51(11):113303, November 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i11/p113303_s1.

Pozdeeva:2010:DTG

[527] Ekaterina Pozdeeva and Axel Schulze-Halberg. Darboux transformations for a generalized Dirac equation in two dimensions. Journal of Mathematical Physics, 51(11):113501, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i11/p113501_s1.

Bracken:2010:GIP

[528] Paul Bracken. A geometric interpretation of prolongation by means of connections. Journal of Mathematical Physics, 51(11):113502, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i11/p113502_s1.

Banica:2010:DRG

[529] Teodor Banica and Stephen Curran. Decomposition results for Gram matrix determinants. Journal of Mathematical Physics, 51(11):113503, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i11/p113503_s1.

DeLeo:2010:CFB

[530] Stefano De Leo, Gisele Ducati, Vinicius Leonardi, and Kenia Pereira. A closed formula for the barrier transmission coefficient in quaternionic quantum mechanics. Journal of Mathematical Physics, 51(11):113504, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i11/p113504_s1.

Latyshev:2010:SEA

[531] A. V. Latyshev and A. A. Yushkanov. Skin effect with arbitrary specularity in Maxwellian plasma. Journal of Mathematical Physics, 51 (11):113505, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p113505_s1.

Lu:2010:BPM

[532] Xing Lü, Bo Tian, Kun Sun, and Pan Wang. Bell-polynomial manipulations on the Bäcklund transformations and Lax pairs for some soliton equations with one Tau-function. Journal of Mathematical Physics, 51(11):113506, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i11/p113506_s1.

Zhaqilao:2010:SSK

[533] Zhaqilao and Sirendaoreji. N-soliton solutions of the KdV6 and mKdV6 equations. Journal of Mathematical Physics, 51(11):113507, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i11/p113507_s1.

Baumgartel:2010:RQM

[534] Hellmut Baumgärtel. Resonances of quantum mechanical scattering systems and Lax–Phillips scattering theory. Journal of Mathematical Physics, 51(11):113508, November 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i11/p113508_s1.

Amore:2010:CMF

[535] Paolo Amore, Francisco M. Fernández, Christoph P. Hofmann, and Ricardo A. Sáenz. Collocation method for fractional quantum mechanics. Journal of Mathematical Physics, 51(12):122101, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122101_s1.

Gorska:2010:QAC

[536] K. Górska, K. A. Penson, A. Horzela, G. H. E. Duchamp, P. Blasiak, and A. I. Solomon. Quasiclassical asymptotics and coherent states for bounded discrete spectra. Journal of Mathematical Physics, 51 (12):122102, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p122102_s1.

Johnson-Freyd:2010:FPI

[537] Theo Johnson-Freyd. The formal path integral and quantum mechanics. Journal of Mathematical Physics, 51(12):122103, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122103_s1.

VanGorder:2010:WFE

[538] Robert A. Van Gorder. Wave functions and energy spectra for the hydrogenic atom in R^{3} × M. Journal of Mathematical Physics, 51(12):122104, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122104_s1.

Meng:2010:KP

[539] Guowu Meng. The U(1)–Kepler problems. Journal of Mathematical Physics, 51(12):122105, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p122105_s1.

Grama:2010:RSA

[540] Cornelia Grama, N. Grama, and I. Zamfirescu. Riemann surface approach to bound and resonant states for a nonlocal potential. Journal of Mathematical Physics, 51(12):122106, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p122106_s1.

Exner:2010:AAC

[541] Pavel Exner and Jirí Lipovský. On the absence of absolutely continuous spectra for Schrödinger operators on radial tree graphs. Journal of Mathematical Physics, 51(12):122107, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p122107_s1.

Buyukasik:2010:MRD

[542] Sirin A. Büyükasik and Oktay K. Pashaev. Madelung representation of damped parametric quantum oscillator and exactly solvable Schrödinger– Burgers equations. Journal of Mathematical Physics, 51(12):122108, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122108_s1.

Temme:2010:DMT

[543] K. Temme, M. J. Kastoryano, M. B. Ruskai, M. M. Wolf, and F. Verstraete. The χ^{2}-divergence and mixing times of quantum Markov processes. Journal of Mathematical Physics, 51(12):122201, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122201_s1.

Datta:2010:UCT

[544] Nilanjana Datta and Min-Hsiu Hsieh. Universal coding for transmission of private information. Journal of Mathematical Physics, 51(12):122202, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122202_s1.

Cerchiai:2010:QE

[545] Bianca Letizia Cerchiai and Bert van Geemen. From qubits to E7. Journal of Mathematical Physics, 51(12):122203, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122203_s1.

Ghosh:2010:CHT

[546] Sibasish Ghosh and Shasanka Mohan Roy. Chain of Hardy-type local reality constraints for n qubits. Journal of Mathematical Physics, 51(12):122204, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p122204_s1.

Budroni:2010:EPP

[547] Costantino Budroni and Giovanni Morchio. The extension problem for partial Boolean structures in quantum mechanics. Journal of Mathematical Physics, 51(12):122205, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v51/i12/p122205_s1.

Deckert:2010:TEE

[548] D.-A. Deckert, D. Dürr, F. Merkl, and M. Schottenloher. Time-evolution of the external field problem in Quantum Electrodynamics. Journal of Mathematical Physics, 51(12):122301, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p122301_s1. See

publisher’s note [635].

Brandt:2010:SACb

[549] Friedemann Brandt. Supersymmetry algebra cohomology. I. Definition and general structure. Journal of Mathematical Physics, 51(12):122302, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122302_s1.

DeBellis:2010:QNP

[550] Joshua DeBellis, Christian Sämann, and Richard J. Szabo. Quantized Nambu–Poisson manifolds and n-Lie algebras. Journal of Mathematical Physics, 51(12):122303, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p122303_s1.

Wong:2010:KGE

[551] C. Y. Wong. Klein–Gordon equation in hydrodynamical form. Journal of Mathematical Physics, 51(12):122304, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p122304_s1.

Gillard:2010:HSN

[552] Mike Gillard and Paul Sutcliffe. Hopf solitons in the Nicole model. Journal of Mathematical Physics, 51(12):122305, December 2010. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122305_s1.

Tausk:2010:CWM

[553] Daniel V. Tausk and Roderich Tumulka. Can we make a Bohmian electron reach the speed of light, at least for one instant? Journal of Mathematical Physics, 51(12):122306, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i12/p122306_s1.

Diaz-Marin:2010:CFC

[554] Homero Díaz-Marín and José A. Zapata. Curvature function and coarse graining. Journal of Mathematical Physics, 51(12):122307, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p122307_s1.

Laia:2010:NAB

[555] João N. Laia. Non-Abelian Berry phase, instantons, and N = (0, 4) supersymmetry. Journal of Mathematical Physics, 51(12):122308, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p122308_s1.

Michishita:2010:CRC

[556] Yoji Michishita. Counting SO(9) × SU(2) representations in coordinate independent state space of SU(2) matrix theory. Journal of Mathematical Physics, 51(12):122309, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p122309_s1.

Natorf:2010:PHT

[557] Włodzimierz Natorf. Past horizons in twisting algebraically special space–times. Journal of Mathematical Physics, 51(12):122501, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p122501_s1.

McLellan:2010:GCS

[558] Brendan McLellan. Gravitational Chern–Simons and the adiabatic limit. Journal of Mathematical Physics, 51(12):122502, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p122502_s1.

Alipour-Fakhri:2010:WSM

[559] Y. Alipour-Fakhri and M. M. Rezaii. The warped Sasaki–Matsumoto metric and bundlelike condition. Journal of Mathematical Physics, 51 (12):122701, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p122701_s1.

Komendarczyk:2010:TOH

[560] R. Komendarczyk. The third order helicity of magnetic fields via link maps. II. Journal of Mathematical Physics, 51(12):122702, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p122702_s1.

Sonnino:2010:RCR

[561] Giorgio Sonnino. Relaxation of chemical reactions to stationary states in the chemical affinities space. Journal of Mathematical Physics, 51 (12):122703, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p122703_s1.

Lee:2010:NRD

[562] Niann-Chern Lee, Hsin-Fu Shen, and Ming-Hsien Tu. A note on reductions of the dispersionless Toda hierarchy. Journal of Mathematical Physics, 51(12):122704, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p122704_s1.

Ozeren:2010:ITE

[563] Serhat F. Özeren. Investigation of the time evolutions of some logperiodic oscillators. Journal of Mathematical Physics, 51(12):122901, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122901_s1.

Ben-Yaacov:2010:LRL

[564] Uri Ben-Ya’acov. Laplace–Runge–Lenz symmetry in general rotationally symmetric systems. Journal of Mathematical Physics, 51(12):122902, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122902_s1.

Pons:2010:SFW

[565] Josep M. Pons. Substituting fields within the action: Consistency issues and some applications. Journal of Mathematical Physics, 51(12):122903, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p122903_s1.

Anderson:2010:ILM

[566] Robert L. Anderson. An invertible linearization map for the quartic oscillator. Journal of Mathematical Physics, 51(12):122904, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p122904_s1.

Lv:2010:BSG

[567] Guangying Lv and Mingxin Wang. Blow-up solutions of the general bequation. Journal of Mathematical Physics, 51(12):123101, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p123101_s1.

Mori:2010:ESG

[568] Koichi Mori. Exact solution to the general Riemann problem in nonuniform and nonstationary media: a simplified analysis of a shock wave accelerated at a constant rate. Journal of Mathematical Physics, 51 (12):123102, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123102_s1.

Allaire:2010:HLI

[569] Grégoire Allaire, Andro Mikelić, and Andrey Piatnitski. Homogenization of the linearized ionic transport equations in rigid periodic porous media. Journal of Mathematical Physics, 51(12):123103, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/

p123103_s1. See erratum [863].

Palmer:2010:SMS

[570] John Palmer and Grethe Hystad. Spin matrix for the scaled periodic Ising model. Journal of Mathematical Physics, 51(12):123301, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p123301_s1.

Furuichi:2010:ACT

[571] Shigeru Furuichi. An axiomatic characterization of a two-parameter extended relative entropy. Journal of Mathematical Physics, 51(12):123302, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p123302_s1.

Tanguay:2010:DBG

[572] J. Tanguay, M. Gil, D. J. Jeffrey, and S. R. Valluri. D-dimensional Bose gases and the Lambert W function. Journal of Mathematical Physics, 51(12):123303, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123303_s1.

Desrosiers:2010:SML

[573] Patrick Desrosiers and Bertrand Eynard. Supermatrix models, loop equations, and duality. Journal of Mathematical Physics, 51(12):123304, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p123304_s1.

Bassler:2010:EES

[574] K. E. Bassler, P. J. Forrester, and N. E. Frankel. Edge effects in some perturbations of the Gaussian unitary ensemble. Journal of Mathematical Physics, 51(12):123305, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p123305_s1.

Jaeck:2010:EBA

[575] Thomas Jaeck and Valentin A. Zagrebnov. Exactness of the Bogoliubov approximation in random external potentials. Journal of Mathematical Physics, 51(12):123306, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p123306_s1.

Melikyan:2010:HCR

[576] A. Melikyan, A. Pinzul, and G. Weber. Higher charges and regularized quantum trace identities in su(1, 1) Landau–Lifshitz model. Journal of Mathematical Physics, 51(12):123501, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p123501_s1.

Ali:2010:MLL

[577] S. Twareque Ali, F. Bagarello, and Jean Pierre Gazeau. Modified Landau levels, damped harmonic oscillator, and two-dimensional pseudobosons. Journal of Mathematical Physics, 51(12):123502, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123502_s1.

Zhang:2010:GRG

[578] Ting Zhang. Global regularity for generalized anisotropic Navier–Stokes equations. Journal of Mathematical Physics, 51(12):123503, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p123503_s1.

Benaissa:2010:EDS

[579] Abbes Benaissa and Sohbi Benazzouz. Energy decay of solutions to the Cauchy problem for a nondissipative wave equation. Journal of Mathematical Physics, 51(12):123504, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i12/p123504_s1.

Vignat:2010:PAS

[580] C. Vignat. A probabilistic approach to some results by Nieto and Truax. Journal of Mathematical Physics, 51(12):123505, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123505_s1.

Yanagida:2010:FDA

[581] Shintarou Yanagida. Five-dimensional SU(2) AGT conjecture and recursive formula of deformed Gaiotto state. Journal of Mathematical Physics, 51(12):123506, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123506_s1.

Ji:2010:IPC

[582] Un Cig Ji and Nobuaki Obata. Implementation problem for the canonical commutation relation in terms of quantum white noise derivatives. Journal of Mathematical Physics, 51(12):123507, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123507_s1.

Ding:2010:MAD

[583] Zhonghai Ding, Goong Chen, and Chang-Shou Lin. Mathematical analysis of the dimensional scaling technique for the Schrödinger equation with power-law potentials. Journal of Mathematical Physics, 51(12):123508, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p123508_s1.

Mamode:2010:IRD

[584] Malik Mamode. Integral representation of Dirac distribution using the Tsallis q-exponential function. Journal of Mathematical Physics, 51(12):123509, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123509_s1.

Kazmierowski:2010:GPM

[585] Dominique Kazmierowski, Azzouz Zinoun, and Ahmed Intissar. Geometric properties of the magnetic Laplacian on the Euclidean 4space. Journal of Mathematical Physics, 51(12):123510, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123510_s1.

Shreecharan:2010:ACS

[586] T. Shreecharan and K. V. S. Shiv Chaitanya. Aspects of coherent states of nonlinear algebras. Journal of Mathematical Physics, 51(12):123511, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p123511_s1.

Jursenas:2010:DAT

[587] Rytis Jursenas and Gintaras Merkelis. Development of algebraic techniques for the atomic open-shell MBPT(3). Journal of Mathematical Physics, 51(12):123512, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p123512_s1.

Fathi:2010:GFI

[588] S. M. Saberi Fathi. Green function of the inhomogeneous Helmholtz equation with nonuniform refraction index, using quaternion analysis. Journal of Mathematical Physics, 51(12):123513, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123513_s1.

Yun:2010:CPB

[589] Seok-Bae Yun. Cauchy problem for the Boltzmann–BGK model near a global Maxwellian. Journal of Mathematical Physics, 51(12):123514, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i12/p123514_s1.

Regniers:2010:WQS

[590] G. Regniers and J. Van der Jeugt. Wigner quantization of some one-dimensional Hamiltonians. Journal of Mathematical Physics, 51 (12):123515, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123515_s1.

Carre:2010:ASL

[591] Christophe Carré, Matthieu Deneufchatel, Jean-Gabriel Luque, and Pierpaolo Vivo. Asymptotics of Selberg-like integrals: The unitary case and Newton’s interpolation formula. Journal of Mathematical Physics, 51(12):123516, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123516_s1.

deOliveira:2010:FSE

[592] Edmundo Capelas de Oliveira, Felix Silva Costa, and Jayme Vaz, Jr. The fractional Schrödinger equation for delta potentials. Journal of Mathematical Physics, 51(12):123517, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i12/p123517_s1.

Gough:2010:CAE

[593] John E. Gough, Hendra I. Nurdin, and Sebastian Wildfeuer. Commutativity of the adiabatic elimination limit of fast oscillatory components and the instantaneous feedback limit in quantum feedback networks. Journal of Mathematical Physics, 51(12):123518, December 2010. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v51/i12/p123518_s1.

Lein:2010:TPA

[594] Max Lein. Two-parameter asymptotics in magnetic Weyl calculus. Journal of Mathematical Physics, 51(12):123519, December 2010.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i12/
p123519_s1.

Yan:2010:MBS

[595] Guozheng Yan and Mao Yao. Mathematical basis of scattering problems from penetrable obstacles and cracks. Journal of Mathematical Physics, 51(12):123520, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123520_s1.

Aktosun:2010:ESS

[596] Tuncay Aktosun, Francesco Demontis, and Cornelis van der Mee. Exact solutions to the sine-Gordon equation. Journal of Mathematical Physics, 51(12):123521, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i12/p123521_s1.

Guo:2010:WMF

[597] Xiangqian Guo and Xuewen Liu. Weight modules with a finite dimensional weight space over truncated Virasoro algebras. Journal of Mathematical Physics, 51(12):123522, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i12/p123522_s1.

Hatzinikitas:2010:WCF

[598] Agapitos N. Hatzinikitas. The weakly coupled fractional one-dimensional Schrödinger operator with index 1 < α ≤ 2. Journal of Mathematical Physics, 51(12):123523, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i12/p123523_s1.

Roschy:2010:PNC

[599] Helmut Roschy and Thorsten Rahn. Publisher’s note: “Cohomology of line bundles: Proof of the algorithm”, [J. Math. Phys. 51, 103520 (2010)]. Journal of Mathematical Physics, 51(12):129901, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i12/p129901_s1. See [503].

Blumenhagen:2010:PNC

[600] Ralph Blumenhagen, Benjamin Jurke, Thorsten Rahn, and Helmut Roschy. Publisher’s note: “Cohomology of line bundles: A computational algorithm”, [J. Math. Phys. 51, 103525 (2010)]. Journal of Mathematical Physics, 51(12):129902, December 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i12/p129902_s1. See [508].

Masillo:2011:SRA

[601] F. Masillo, G. Scolarici, and L. Solombrino. Some remarks on assignment maps. Journal of Mathematical Physics, 52(1):012101, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p012101_s1.

Junge:2011:CEP

[602] M. Junge, M. Navascues, C. Palazuelos, D. Perez-Garcia, V. B. Scholz, and R. F. Werner. Connes’ embedding problem and Tsirelson’s problem. Journal of Mathematical Physics, 52(1):012102, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/p012102_s1.

Stachowiak:2011:SDE

[603] Tomasz Stachowiak. On solvable Dirac equation with polynomial potentials. Journal of Mathematical Physics, 52(1):012301, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/
p012301_s1.

Jakel:2011:GTT

[604] Christian D. Jäkel and Walter F. Wreszinski. A Goldstone theorem in thermal relativistic quantum field theory. Journal of Mathematical Physics, 52(1):012302, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012302_s1.

Ludwin:2011:GRW

[605] D. M. Ludwin and L. P. Horwitz. Gravitational repulsion within a black hole using the Stueckelberg quantum formalism. Journal of Mathematical Physics, 52(1):012303, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012303_s1.

Conrady:2011:UIR

[606] Florian Conrady and Jeff Hnybida. Unitary irreducible representations of SL(2, C) in discrete and continuous SU(1, 1) bases. Journal of Mathematical Physics, 52(1):012501, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i1/p012501_s1.

Sahlmann:2011:SRC

[607] Hanno Sahlmann. Some results concerning the representation theory of the algebra underlying loop quantum gravity. Journal of Mathematical Physics, 52(1):012502, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012502_s1.

Sahlmann:2011:WDM

[608] Hanno Sahlmann. When do measures on the space of connections support the triad operators of loop quantum gravity? Journal of Mathematical Physics, 52(1):012503, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012503_s1.

Juang:2011:TWT

[609] Jonq Juang and Chin-Lung Li. The theory of wavelet transform method on chaotic synchronization of coupled map lattices. Journal of Mathematical Physics, 52(1):012701, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i1/p012701_s1.

Llibre:2011:AIH

[610] Jaume Llibre, Adam Mahdi, and Claudia Valls. Analytic integrability of Hamiltonian systems with a homogeneous polynomial potential of degree 4. Journal of Mathematical Physics, 52(1):012702, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p012702_s1.

Fernandez:2011:WCT

[611] O. E. Fernandez and A. M. Bloch. The Weitzenböck connection and time reparameterization in nonholonomic mechanics. Journal of Mathematical Physics, 52(1):012901, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012901_s1.

Bruno:2011:ESH

[612] Danilo Bruno. Existence of solutions for Hamiltonian field theories by the Hamilton–Jacobi technique. Journal of Mathematical Physics, 52(1):

012902, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i1/p012902_s1.

Nacher:2011:PMC

[613] J. C. Nacher and T. Ochiai. Plus-minus construction leads to perfect invisibility. Journal of Mathematical Physics, 52(1):012903, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p012903_s1.

Pons:2011:NSE

[614] Josep M. Pons. Noether symmetries, energy-momentum tensors, and conformal invariance in classical field theory. Journal of Mathematical Physics, 52(1):012904, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p012904_s1.

Gibbons:2011:NSS

[615] G. W. Gibbons, C. M. Warnick, and W. W. Wong. Nonexistence of Skyrmion–Skyrmion and Skyrmion–anti-Skyrmion static equilibria. Journal of Mathematical Physics, 52(1):012905, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/p012905_s1.

Yaremko:2011:SFE

[616] Yurij Yaremko. Self-force via energy-momentum and angular momentum balance equations. Journal of Mathematical Physics, 52(1):012906, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p012906_s1.

Pirjol:2011:PTL

[617] Dan Pirjol. Phase transition in a log-normal Markov functional model. Journal of Mathematical Physics, 52(1):013301, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/p013301_s1.

Hystad:2011:PIC

[618] Grethe Hystad. Periodic Ising correlations. Journal of Mathematical Physics, 52(1):013302, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p013302_s1.

Kojima:2011:DIT

[619] Takeo Kojima. Diagonalization of infinite transfer matrix of boundary U_{q,p}(A_{N−1}^{(1)}) face model. Journal of Mathematical Physics, 52(1): 013501, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i1/p013501_s1.

Tatar:2011:ADL

[620] Nasser eddine Tatar. Arbitrary decays in linear viscoelasticity. Journal of Mathematical Physics, 52(1):013502, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i1/p013502_s1.

Song:2011:NIT

[621] Junfeng Song, Changzheng Qu, and Zhijun Qiao. A new integrable two-component system with cubic nonlinearity. Journal of Mathematical Physics, 52(1):013503, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i1/p013503_s1.

Douglas:2011:EEA

[622] Andrew Douglas and Joe Repka. Embeddings of the Euclidean algebra ϵ(3) into sl(4, C) and restrictions of irreducible representations of sl(4, C). Journal of Mathematical Physics, 52(1):013504, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/p013504_s1.

Duchene:2011:WOB

[623] Vincent Duchêne, Jeremy L. Marzuola, and Michael I. Weinstein. Wave operator bounds for one-dimensional Schrödinger operators with singular potentials and applications. Journal of Mathematical Physics, 52(1): 013505, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i1/p013505_s1.

Goudarzi:2011:SDE

[624] H. Goudarzi, M. Sohbati, and S. Zarrin. Solution of Dirac equation with spin and pseudospin symmetry for an anharmonic oscillator. Journal of Mathematical Physics, 52(1):013506, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i1/p013506_s1. See comment

Rowe:2011:SUP

[625] D. J. Rowe, J. Repka, and M. J. Carvalho. Simple unified proofs of four duality theorems. Journal of Mathematical Physics, 52(1):013507, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p013507_s1.

Campos:2011:WPC

[626] L. M. B. C. Campos. On a wave-particle in closed and open isotropic universes. Journal of Mathematical Physics, 52(1):013508, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/
p013508_s1.

Aizawa:2011:LWR

[627] N. Aizawa. Lowest weight representations of super Schrödinger algebras in one dimensional space. Journal of Mathematical Physics, 52(1): 013509, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i1/p013509_s1.

Bagarello:2011:RSQ

[628] F. Bagarello, C. Trapani, and S. Triolo. Representable states on quasilocal quasi^{∗}-algebras. Journal of Mathematical Physics, 52(1):013510, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p013510_s1.

Qu:2011:BTL

[629] Qi-Xing Qu, Bo Tian, Kun Sun, and Yan Jiang. Bäcklund transformation, Lax pair, and solutions for the Caudrey–Dodd–Gibbon equation. Journal of Mathematical Physics, 52(1):013511, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/p013511_s1.

Caragheorgheopol:2011:AEA

[630] Dan Caragheorgheopol and Josef Tkadlec. Atomic effect algebras with compression bases. Journal of Mathematical Physics, 52(1):013512, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p013512_s1.

Delvaux:2011:ACP

[631] Steven Delvaux. Average characteristic polynomials in the two-matrix model. Journal of Mathematical Physics, 52(1):013513, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i1/
p013513_s1.

Gonzales:2011:NST

[632] Marcelo Gonzales, Sadi Khodaee, and Francesco Toppan. On nonminimal N = 4 supermultiplets in 1D and their associated σ-models. Journal of Mathematical Physics, 52(1):013514, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i1/p013514_s1.

Pauley:2011:JKC

[633] Michael Pauley. Jupp and Kent’s cubics in Lie groups. Journal of Mathematical Physics, 52(1):013515, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i1/p013515_s1.

deAzcarraga:2011:CFA

[634] José A. de Azcárraga, José M. Izquierdo, and Moisés Picón. Contractions of Filippov algebras. Journal of Mathematical Physics, 52(1): 013516, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i1/p013516_s1.

Deckert:2011:PNT

[635] D.-A. Deckert, D. Dürr, F. Merkl, and M. Schottenloher. Publisher’s note: “Time-evolution of the external field problem in quantum electrodynamics”, [J. Math. Phys. 51, 122301 (2010)]. Journal of Mathematical Physics, 52(1):019901, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i1/p019901_s1. See [548].

Trisetyarso:2011:ECD

[636] Agung Trisetyarso. Erratum: “Correlation of Dirac potentials and atomic inversion in cavity quantum electrodynamics”, [J. Math. Phys. 51, 072103 (2010)]. Journal of Mathematical Physics, 52(1):019902, January 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i1/p019902_s1. See [308].

Yamashita:2011:PSP

[637] Hideyasu Yamashita. Phase-space path integral and Brownian motion. Journal of Mathematical Physics, 52(2):022101, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/p022101_s1.

Hidaka:2011:EGS

[638] Takeru Hidaka. Existence of a ground state for the Nelson model with a singular perturbation. Journal of Mathematical Physics, 52(2):022102, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022102_s1.

Finco:2011:AEW

[639] Domenico Finco and Alessandro Teta. Asymptotic expansion for the wave function in a one-dimensional model of inelastic interaction. Journal of Mathematical Physics, 52(2):022103, February 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i2/p022103_s1.

Pavicic:2011:NCD

[640] Mladen Pavicić, Norman D. Megill, P. K. Aravind, and Mordecai Waegell. New class of 4-dim Kochen–Specker sets. Journal of Mathematical Physics, 52(2):022104, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p022104_s1.

Sanchez-Moreno:2011:UBS

[641] P. Sánchez-Moreno, S. Zozor, and J. S. Dehesa. Upper bounds on Shannon and Rényi entropies for central potentials. Journal of Mathematical Physics, 52(2):022105, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p022105_s1.

Cannata:2011:NTD

[642] F. Cannata, M. V. Ioffe, and D. N. Nishnianidze. New two-dimensional quantum models with shape invariance. Journal of Mathematical Physics, 52(2):022106, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p022106_s1.

Tosiek:2011:CSC

[643] Jaromir Tosiek. Compatible symplectic connections on a cotangent bundle and the Fedosov quantization. Journal of Mathematical Physics, 52(2):022107, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p022107_s1.

Ancarani:2011:SAL

[644] L. U. Ancarani and G. Gasaneo. A special asymptotic limit of a Kampé de Fériet hypergeometric function appearing in nonhomogeneous Coulomb problems. Journal of Mathematical Physics, 52(2):022108, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p022108_s1.

Bruneau:2011:SIC

[645] L. Bruneau, S. De Bièvre, and C.-A. Pillet. Scattering induced current in a tight-binding band. Journal of Mathematical Physics, 52(2):022109, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022109_s1.

Konenberg:2011:REA

[646] Martin Könenberg. Return to equilibrium for an anharmonic oscillator coupled to a heat bath. Journal of Mathematical Physics, 52(2):022110, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022110_s1.

Beigi:2011:GCQ

[647] Salman Beigi, Isaac Chuang, Markus Grassl, Peter Shor, and Bei Zeng. Graph concatenation for quantum codes. Journal of Mathematical Physics, 52(2):022201, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p022201_s1.

Appleby:2011:LAS

[648] D. M. Appleby, Steven T. Flammia, and Christopher A. Fuchs. The Lie algebraic significance of symmetric informationally complete measurements. Journal of Mathematical Physics, 52(2):022202, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p022202_s1.

Upadhyay:2011:SNF

[649] Sudhaker Upadhyay, Sumit Kumar Rai, and Bhabani Prasad Mandal. Off-shell nilpotent finite Becchi–Rouet–Stora–Tyutin (BRST)/anti-BRST transformations. Journal of Mathematical Physics, 52(2):022301, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022301_s1.

Zinoviev:2011:CEI

[650] Yury M. Zinoviev. Causal electromagnetic interaction equations. Journal of Mathematical Physics, 52(2):022302, February 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i2/p022302_s1.

Muhlhoff:2011:CPG

[651] Rainer Mühlhoff. Cauchy problem and Green’s functions for first order differential operators and algebraic quantization. Journal of Mathematical Physics, 52(2):022303, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p022303_s1.

Friedmann:2011:OSL

[652] Tamar Friedmann. Orbifold singularities, Lie algebras of the third kind (LATKes), and pure Yang–Mills with matter. Journal of Mathematical Physics, 52(2):022304, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p022304_s1.

Takaesu:2011:SLQ

[653] Toshimitsu Takaesu. Scaling limit of quantum electrodynamics with spatial cutoffs. Journal of Mathematical Physics, 52(2):022305, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p022305_s1.

Drissi:2011:GLF

[654] L. B. Drissi, E. H. Saidi, and M. Bousmina. Graphene, Lattice Field Theory and symmetries. Journal of Mathematical Physics, 52(2):022306, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i2/p022306_s1.

Aazami:2011:OFC

[655] A. B. Aazami, A. O. Petters, and J. M. Rabin. Orbifolds, the A, D, E family of caustic singularities, and gravitational lensing. Journal of Mathematical Physics, 52(2):022501, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p022501_s1.

Khatsymovsky:2011:ASS

[656] V. M. Khatsymovsky. Attributing sense to some integrals in Regge calculus. Journal of Mathematical Physics, 52(2):022502, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/
p022502_s1.

Marrero:2011:KNS

[657] Juan Carlos Marrero, Narciso Román-Roy, Modesto Salgado, and Silvia Vilariño. On a kind of Noether symmetries and conservation laws in k-cosymplectic field theory. Journal of Mathematical Physics, 52(2):022901, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p022901_s1.

Kiessling:2011:CPP

[658] M. K.-H. Kiessling. Convergent perturbative power series solution of the stationary Maxwell–Born–Infeld field equations with regular sources. Journal of Mathematical Physics, 52(2):022902, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/p022902_s1.

Ben-Yaacov:2011:CMI

[659] Uri Ben-Ya’acov. Center-of-mass and internal symmetries in classical relativistic systems. Journal of Mathematical Physics, 52(2):022903, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p022903_s1.

Hao:2011:WPV

[660] Chengchun Hao. Well-posedness for the viscous rotating shallow water equations with friction terms. Journal of Mathematical Physics, 52(2):023101, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023101_s1.

Qin:2011:GSS

[661] Yuming Qin and Guili Hu. Global smooth solutions for 1D thermally radiative magnetohydrodynamics. Journal of Mathematical Physics, 52(2):023102, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023102_s1.

Carrisi:2011:MMM

[662] Maria Cristina Carrisi and Sebastiano Pennisi. The model with many moments for relativistic electron beams: a simplified solution. Journal of Mathematical Physics, 52(2):023103, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i2/p023103_s1.

Chang:2011:HWS

[663] Shu-Chiuan Chang and Lung-Chi Chen. Hamiltonian walks on the Sierpinski gasket. Journal of Mathematical Physics, 52(2):023301, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023301_s1.

Tay:2011:BSA

[664] B. A. Tay, Kazuki Kanki, Satoshi Tanaka, and Tomio Petrosky. Band structure and accumulation point in the spectrum of quantum collision operator in a one-dimensional molecular chain. Journal of Mathematical Physics, 52(2):023302, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p023302_s1.

Tracy:2011:ASE

[665] Craig A. Tracy and Harold Widom. On asymmetric simple exclusion process with periodic step Bernoulli initial condition. Journal of Mathematical Physics, 52(2):023303, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p023303_s1.

Du:2011:GGN

[666] Hong-Ke Du, Yan-Ni Dou, Hai-Yan Zhang, and Ling-Min Shen. A generalization of Gudder–Nagy’s theorem with numerical ranges of operators. Journal of Mathematical Physics, 52(2):023501, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/p023501_s1.

Zhang:2011:GEU

[667] Zai yun Zhang, Zhen hai Liu, Xiu jin Miao, and Yue zhong Chen. Global existence and uniform stabilization of a generalized dissipative Klein– Gordon equation type with boundary damping. Journal of Mathematical Physics, 52(2):023502, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p023502_s1.

Flachi:2011:ZDL

[668] Antonino Flachi and Guglielmo Fucci. Zeta determinant for Laplace operators on Riemann caps. Journal of Mathematical Physics, 52 (2):023503, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023503_s1.

Fan:2011:DCL

[669] Engui Fan and Kwok Wing Chow. Darboux covariant Lax pairs and infinite conservation laws of the (2 + 1)-dimensional breaking soliton equation. Journal of Mathematical Physics, 52(2):023504, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/
p023504_s1.

Zhang:2011:HVA

[670] Runxuan Zhang, Dongping Hou, and Chengming Bai. A Hom-version of the affinizations of Balinskii–Novikov and Novikov superalgebras. Journal of Mathematical Physics, 52(2):023505, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/p023505_s1.

Aulisa:2011:LTD

[671] Eugenio Aulisa, Lidia Bloshanskaya, and Akif Ibragimov. Long-term dynamics for well productivity index for nonlinear flows in porous media. Journal of Mathematical Physics, 52(2):023506, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/p023506_s1.

Alex:2011:NAE

[672] Arne Alex, Matthias Kalus, Alan Huckleberry, and Jan von Delft. A numerical algorithm for the explicit calculation of SU(N) and SL(N, C) Clebsch–Gordan coefficients. Journal of Mathematical Physics, 52 (2):023507, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023507_s1.

Cheng:2011:ETD

[673] Ming Cheng. Excitation threshold for the discrete coupled Schrödinger lattice system. Journal of Mathematical Physics, 52(2):023508, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023508_s1.

Sakovich:2011:SSS

[674] Sergei Sakovich. Smooth soliton solutions of a new integrable equation by Qiao. Journal of Mathematical Physics, 52(2):023509, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023509_s1.

Le:2011:SWE

[675] Út V. Lê. A semilinear wave equation with space–time dependent coefficients and a memory boundarylike antiperiodic condition: a lowfrequency asymptotic expansion. Journal of Mathematical Physics, 52(2):023510, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023510_s1.

Cimmelli:2011:EEP

[676] V. A. Cimmelli, F. Oliveri, and V. Triani. Exploitation of the entropy principle: Proof of Liu theorem if the gradients of the governing equations are considered as constraints. Journal of Mathematical Physics, 52(2):023511, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p023511_s1.

Zhao:2011:MMM

[677] Hai qiong Zhao and Zuo nong Zhu. Multisoliton, multipositon, multinegaton, and multiperiodic solutions of a coupled Volterra lattice system and their continuous limits. Journal of Mathematical Physics, 52(2):023512, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023512_s1.

Cresson:2011:NDN

[678] Jacky Cresson and Isabelle Greff. A non-differentiable Noether’s theorem. Journal of Mathematical Physics, 52(2):023513, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i2/
p023513_s1.

Huang:2011:GTF

[679] Qing Huang, Changzheng Qu, and Renat Zhdanov. Group-theoretical framework for potential symmetries of evolution equations. Journal of Mathematical Physics, 52(2):023514, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i2/p023514_s1.

Rajendran:2011:MWS

[680] S. Rajendran, M. Lakshmanan, and P. Muruganandam. Matter wave switching in Bose–Einstein condensates via intensity redistribution soliton interactions. Journal of Mathematical Physics, 52(2):023515, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023515_s1.

Novikov:2011:CSE

[681] V. S. Novikov and E. V. Ferapontov. On the classification of scalar evolutionary integrable equations in 2+1 dimensions. Journal of Mathematical Physics, 52(2):023516, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p023516_s1.

Pan:2011:ASQ

[682] Xing-Bin Pan. Asymptotics of solutions of a quasilinear system involving curl. Journal of Mathematical Physics, 52(2):023517, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023517_s1.

Akbarzade:2011:DNF

[683] M. Akbarzade and J. Langari. Determination of natural frequencies by coupled method of homotopy perturbation and variational method for strongly nonlinear oscillators. Journal of Mathematical Physics, 52(2):023518, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023518_s1.

Ramm:2011:WSM

[684] A. G. Ramm and A. Rona. Wave scattering by many small bodies and applications. Journal of Mathematical Physics, 52(2):023519, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i2/p023519_s1.

Bailleul:2011:PVS

[685] Ismaël Bailleul. A probabilistic view on singularities. Journal of Mathematical Physics, 52(2):023520, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i2/p023520_s1.

deAzcarraga:2011:CLD

[686] José A. de Azcárraga and José M. Izquierdo. On a class of n-Leibniz deformations of the simple Filippov algebras. Journal of Mathematical Physics, 52(2):023521, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i2/p023521_s1.

Sun:2011:QDC

[687] Jian-Qing Sun, Yi He, Xing-Biao Hu, and Hon-Wah Tam. Q-difference and confluent forms of the lattice Boussinesq equation and the relevant convergence acceleration algorithms. Journal of Mathematical Physics, 52(2):023522, February 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i2/p023522_s1.

Ma:2011:ISS

[688] Zhi-Hao Ma and Sen Zhu. Intrinsic structure of state space of a quantum system. Journal of Mathematical Physics, 52(3):032101, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032101_s1.

Castro:2011:NOH

[689] P. G. Castro, B. Chakraborty, R. Kullock, and F. Toppan. Noncommutative oscillators from a Hopf algebra twist deformation. A first principles derivation. Journal of Mathematical Physics, 52(3):032102, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032102_s1.

Chen:2011:CPG

[690] Zeqian Chen and Chuangye Liu. On the Cauchy problem for Gross– Pitaevskii hierarchies. Journal of Mathematical Physics, 52(3):032103, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i3/p032103_s1.

Dutta:2011:IEC

[691] D. Dutta and P. Roy. Information entropy of conditionally exactly solvable potentials. Journal of Mathematical Physics, 52(3):032104, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032104_s1.

Le:2011:NAM

[692] Van-Hoang Le and Thanh-Son Nguyen. A non-Abelian SO(8) monopole as generalization of Dirac–Yang monopoles for a 9-dimensional space. Journal of Mathematical Physics, 52(3):032105, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/p032105_s1.

Strauss:2011:TRQ

[693] Y. Strauss. Transition representations of quantum evolution with application to scattering resonances. Journal of Mathematical Physics, 52(3): 032106, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i3/p032106_s1.

Kohler:2011:EDI

[694] Heiner Kohler. Exact diagonalization of 1D interacting spinless Fermions. Journal of Mathematical Physics, 52(3):032107, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/p032107_s1.

Jain:2011:QNC

[695] Avinash Jain, Massimo Franceschetti, and David A. Meyer. On quantum network coding. Journal of Mathematical Physics, 52(3):032201, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032201_s1.

Carbone:2011:DPS

[696] Raffaella Carbone, Emanuela Sasso, and Veronica Umanità. Decoherence for positive semigroups on M_{2}(C). Journal of Mathematical Physics, 52(3):032202, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i3/p032202_s1.

Bowman:2011:CQL

[697] Daniel A. Bowman and John L. Challifour. Canonical quantization of lattice Higgs–Maxwell–Chern–Simons fields: Osterwalder–Schrader positivity. Journal of Mathematical Physics, 52(3):032301, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/
p032301_s1.

Masuku:2011:LPM

[698] Mthokozisi Masuku and João P. Rodrigues. Laplacians in polar matrix coordinates and radial fermionization in higher dimensions. Journal of Mathematical Physics, 52(3):032302, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p032302_s1.

Chaichian:2011:TAF

[699] M. Chaichian, M. N. Mnatsakanova, K. Nishijima, A. Tureanu, and Yu. S. Vernov. Toward an axiomatic formulation of noncommutative quantum field theory. Journal of Mathematical Physics, 52(3): 032303, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i3/p032303_s1.

Tanabe:2011:AMN

[700] Kentaro Tanabe, Norihiro Tanahashi, and Tetsuya Shiromizu. Angular momentum at null infinity in five dimensions. Journal of Mathematical Physics, 52(3):032501, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i3/p032501_s1.

Weatherall:2011:MBN

[701] James Owen Weatherall. The motion of a body in Newtonian theories. Journal of Mathematical Physics, 52(3):032502, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/p032502_s1.

Guirao:2011:SCN

[702] Juan L. G. Guirao and Juan A. Vera. Sufficient conditions for a nondegenerate Hopf bifurcation in a generalized Lagrange–Poisson problem. Journal of Mathematical Physics, 52(3):032701, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/p032701_s1.

Belmonte-Beitia:2011:EDS

[703] J. Belmonte-Beitia and J. Cuevas. Existence of dark solitons in a class of stationary nonlinear Schrödinger equations with periodically modulated nonlinearity and periodic asymptotics. Journal of Mathematical Physics, 52(3):032702, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i3/p032702_s1.

Valls:2011:DIG

[704] Claudia Valls. Darbouxian integrals for generalized Raychaudhuri equations. Journal of Mathematical Physics, 52(3):032703, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/
p032703_s1.

Shen:2011:CTK

[705] Hsin-Fu Shen and Ming-Hsien Tu. On the constrained B-type Kadomtsev–Petviashvili hierarchy: Hirota bilinear equations and Virasoro symmetry. Journal of Mathematical Physics, 52(3):032704, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032704_s1.

Flannery:2011:DLA

[706] M. R. Flannery. d’Alembert–Lagrange analytical dynamics for nonholonomic systems. Journal of Mathematical Physics, 52(3):032705, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032705_s1.

Jimenez-Lara:2011:PON

[707] Lidia Jiménez-Lara and Jaume Llibre. Periodic orbits and nonintegrability of generalized classical Yang–Mills Hamiltonian systems. Journal of Mathematical Physics, 52(3):032901, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p032901_s1.

Burton:2011:WPI

[708] D. A. Burton, D. C. Christie, J. D. A. Smith, and R. W. Tucker. Wake potentials and impedances of charged beams in gradually tapering structures. Journal of Mathematical Physics, 52(3):032902, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p032902_s1.

Chanu:2011:PCM

[709] Claudia Chanu, Luca Degiovanni, and Giovanni Rastelli. Polynomial constants of motion for Calogero-type systems in three dimensions. Journal of Mathematical Physics, 52(3):032903, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p032903_s1.

Li:2011:NDC

[710] Yan Li. Nonuniform dependence for the Cauchy problem of the general b-equation. Journal of Mathematical Physics, 52(3):033101, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p033101_s1.

Liu:2011:RCT

[711] Qiao Liu, Jihong Zhao, and Shangbin Cui. A regularity criterion for the three-dimensional nematic liquid crystal flow in terms of one directional derivative of the velocity. Journal of Mathematical Physics, 52(3): 033102, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i3/p033102_s1.

Bihlo:2011:LSA

[712] Alexander Bihlo and Roman O. Popovych. Lie symmetry analysis and exact solutions of the quasigeostrophic two-layer problem. Journal of Mathematical Physics, 52(3):033103, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p033103_s1.

Ha:2011:NIO

[713] Seung-Yeal Ha, Taeyoung Ha, Chi-Ok Hwang, and Ho Lee. Nonlinear instability of the one-dimensional Vlasov–Yukawa system. Journal of Mathematical Physics, 52(3):033301, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p033301_s1.

Lefevere:2011:LDC

[714] Raphaël Lefevere, Mauro Mariani, and Lorenzo Zambotti. Large deviations of the current in stochastic collisional dynamics. Journal of Mathematical Physics, 52(3):033302, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i3/p033302_s1.

Ma:2011:PTB

[715] Tian Ma and Shouhong Wang. Phase transitions for the Brusselator model. Journal of Mathematical Physics, 52(3):033501, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/
p033501_s1.

Yin:2011:CEA

[716] Mei Yin. A cluster expansion approach to renormalization group transformations. Journal of Mathematical Physics, 52(3):033502, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/
p033502_s1.

Kostrigina:2011:DIT

[717] O. S. Kostrigina and A. V. Zhiber. Darboux-integrable two-component nonlinear hyperbolic systems of equations. Journal of Mathematical Physics, 52(3):033503, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i3/p033503_s1.

Ismailov:2011:INS

[718] Mansur I. Ismailov. Integration of nonlinear system of four waves with two velocities in (2 + 1) dimensions by the inverse scattering transform method. Journal of Mathematical Physics, 52(3):033504, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i3/p033504_s1.

Garcia-Pelayo:2011:MDD

[719] Ricardo García-Pelayo. Moments of the distribution of distance. Journal of Mathematical Physics, 52(3):033505, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p033505_s1.

Jow:2011:CTL

[720] Shin-Yao Jow. Cohomology of toric line bundles via simplicial Alexander duality. Journal of Mathematical Physics, 52(3):033506, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/
p033506_s1.

Tan:2011:GDS

[721] Wenke Tan and Zhaoyang Yin. Global dissipative solutions of a modified two-component Camassa–Holm shallow water system. Journal of Mathematical Physics, 52(3):033507, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i3/p033507_s1.

Bevilacqua:2011:NCS

[722] G. Bevilacqua, V. Biancalana, Y. Dancheva, T. Mansour, and L. Moi. A new class of sum rules for products of Bessel functions. Journal of Mathematical Physics, 52(3):033508, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p033508_s1.

Charalampidis:2011:SRM

[723] E. G. Charalampidis, T. A. Ioannidou, and N. S. Manton. Skyrmions, rational maps, and scaling identities. Journal of Mathematical Physics, 52(3):033509, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i3/p033509_s1.

Wang:2011:UQI

[724] Hai-Jun Wang. Understanding quantum interference in general nonlocality. Journal of Mathematical Physics, 52(3):033510, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i3/

p033510_s1. See publisher’s note [823].

Asad:2011:EEN

[725] R. Asad and G. Simpson. Embedded eigenvalues and the nonlinear Schrödinger equation. Journal of Mathematical Physics, 52(3): 033511, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i3/p033511_s1.

Bisi:2011:EFD

[726] M. Bisi and G. Spiga. Erratum: “Fluid-dynamic equations for granular particles in a host medium”, [J. Math. Phys. 46, 113301 (2005)]. Journal of Mathematical Physics, 52(3):039901, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p039901_s1. See [7].

Dhatt:2011:CMS

[727] Sharmistha Dhatt and Kamal Bhattacharyya. Concurrent multiple-state analytic perturbation theory via supersymmetry. Journal of Mathematical Physics, 52(4):042101, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i4/p042101_s1.

Korcyl:2011:GIP

[728] Piotr Korcyl. Gauge invariant plane-wave solutions in supersymmetric Yang–Mills quantum mechanics. Journal of Mathematical Physics, 52(4): 042102, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p042102_s1.

Dong:2011:GFT

[729] Jianping Dong. Green’s function for the time-dependent scattering problem in the fractional quantum mechanics. Journal of Mathematical Physics, 52(4):042103, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042103_s1.

Gardas:2011:REP

[730] Bartlomiej Gardas. Riccati equation and the problem of decoherence II: Symmetry and the solution of the Riccati equation. Journal of Mathematical Physics, 52(4):042104, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i4/p042104_s1.

Jiang:2011:ESF

[731] Xiaoyun Jiang, Haitao Qi, and Mingyu Xu. Exact solutions of fractional Schrödinger-like equation with a nonlocal term. Journal of Mathematical Physics, 52(4):042105, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042105_s1. See comment [1409].

Meng:2011:GMK

[732] Guowu Meng and Ruibin Zhang. Generalized MICZ-Kepler problems and unitary highest weight modules. Journal of Mathematical Physics, 52(4): 042106, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p042106_s1.

Friedberg:2011:EMI

[733] Richard Friedberg and Jamal T. Manassah. Electromagnetic modes of an infinite cylindrical sample of two-level atoms. Journal of Mathematical Physics, 52(4):042107, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042107_s1.

Huang-Fu:2011:ESN

[734] Guo-Qing Huang-Fu and Min-Cang Zhang. Exact solution for a noncentral electric dipole ring-shaped potential in the tridiagonal representation. Journal of Mathematical Physics, 52(4):042108, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/
p042108_s1.

Han:2011:WGF

[735] Gang Han. The Weyl group of the fine grading of ∫↕(n, C) associated with tensor product of generalized Pauli matrices. Journal of Mathematical Physics, 52(4):042109, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042109_s1.

Ahlbrecht:2011:AEQ

[736] A. Ahlbrecht, H. Vogts, A. H. Werner, and R. F. Werner. Asymptotic evolution of quantum walks with random coin. Journal of Mathematical Physics, 52(4):042201, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042201_s1.

Holevo:2011:CJF

[737] A. S. Holevo. The Choi–Jamiolkowski forms of quantum Gaussian channels. Journal of Mathematical Physics, 52(4):042202, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/
p042202_s1.

Friedland:2011:AGS

[738] Shmuel Friedland, Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. The automorphism group of separable states in quantum information theory. Journal of Mathematical Physics, 52(4):042203, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/p042203_s1.

Marquette:2011:QAA

[739] Ian Marquette. Quadratic algebra approach to relativistic quantum Smorodinsky–Winternitz systems. Journal of Mathematical Physics, 52(4):042301, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042301_s1.

Finster:2011:SHI

[740] Felix Finster and Christian Hainzl. A spatially homogeneous and isotropic Einstein–Dirac cosmology. Journal of Mathematical Physics, 52(4):042501, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042501_s1.

Hendi:2011:MWS

[741] S. H. Hendi. Magnetic wormhole solutions in Einstein–Gauss–Bonnet gravity with power Maxwell invariant source. Journal of Mathematical Physics, 52(4):042502, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042502_s1.

Willison:2011:BTZ

[742] Steven Willison. The Bañados, Teitelboim, and Zanelli spacetime as an algebraic embedding. Journal of Mathematical Physics, 52(4):042503, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p042503_s1.

Imaykin:2011:SAC

[743] Valery Imaykin, Alexander Komech, and Herbert Spohn. Scattering asymptotics for a charged particle coupled to the Maxwell field. Journal of Mathematical Physics, 52(4):042701, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p042701_s1.

Guo:2011:SAS

[744] Lifang Guo and Quanxin Zhu. Stability analysis for stochastic Volterra– Levin equations with Poisson jumps: Fixed point approach. Journal of Mathematical Physics, 52(4):042702, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p042702_s1.

Wan:2011:ABS

[745] Li Wan and Qinghua Zhou. Asymptotic behaviors of stochastic twodimensional Navier–Stokes equations with finite memory. Journal of Mathematical Physics, 52(4):042703, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p042703_s1.

Gratus:2011:CCR

[746] J. Gratus and R. W. Tucker. Covariant constitutive relations and relativistic inhomogeneous plasmas. Journal of Mathematical Physics, 52(4): 042901, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p042901_s1.

Bludman:2011:IRD

[747] Sidney Bludman and Dallas C. Kennedy. Invariant relationships deriving from classical scaling transformations. Journal of Mathematical Physics, 52(4):042902, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p042902_s1.

Perlick:2011:HME

[748] Volker Perlick. On the hyperbolicity of Maxwell’s equations with a local constitutive law. Journal of Mathematical Physics, 52(4):042903, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p042903_s1.

Maity:2011:REN

[749] Chandan Maity, Nikhil Chakrabarti, and Sudip Sengupta. Relativistic effects on nonlinear lower hybrid oscillations in cold plasma. Journal of Mathematical Physics, 52(4):043101, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p043101_s1.

Ferrari:2011:DTT

[750] F. Ferrari, J. Paturej, M. Pia̧tek, and T. A. Vilgis. Dynamics of two topologically entangled chains. Journal of Mathematical Physics, 52(4): 043301, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p043301_s1.

Grundland:2011:SFP

[751] A. M. Grundland and A. J. Hariton. Supersymmetric formulation of polytropic gas dynamics and its invariant solutions. Journal of Mathematical Physics, 52(4):043501, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i4/p043501_s1.

Zhao:2011:AHD

[752] Dun Zhao, Yu-Juan Zhang, Wei-Wei Lou, and Hong-Gang Luo. AKNS hierarchy, Darboux transformation and conservation laws of the 1D nonautonomous nonlinear Schrödinger equations. Journal of Mathematical Physics, 52(4):043502, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i4/p043502_s1.

Garcia-Recio:2011:CGC

[753] C. Garcia-Recio and L. L. Salcedo. SU(6) ⊃ SU(3) ⊗ SU(2) and SU(8) ⊃ SU(4) ⊗ SU(2) Clebsch–Gordan coefficients. Journal of Mathematical Physics, 52(4):043503, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043503_s1.

Misir:2011:AIT

[754] Adil Misir and Suleyman Ogrekci. Asymptotic iteration technique for second-order dynamic equations on time scales. Journal of Mathematical Physics, 52(4):043504, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043504_s1.

Aujla:2011:SPL

[755] Jaspal Singh Aujla. A simple proof of Lieb concavity theorem. Journal of Mathematical Physics, 52(4):043505, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p043505_s1.

Heller:2011:GNH

[756] Michael Heller, Leszek Pysiak, and Wieslaw Sasin. Geometry of non-Hausdorff spaces and its significance for physics. Journal of Mathematical Physics, 52(4):043506, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043506_s1.

Zhu:2011:MDI

[757] Chunrong Zhu and Changzheng Qu. Maximal dimension of invariant subspaces admitted by nonlinear vector differential operators. Journal of Mathematical Physics, 52(4):043507, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p043507_s1.

Zhu:2011:CLI

[758] Zuo nong Zhu, Hai qiong Zhao, and Xiao nan Wu. On the continuous limits and integrability of a new coupled semidiscrete mKdV system. Journal of Mathematical Physics, 52(4):043508, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/p043508_s1.

Liu:2011:ISD

[759] C. J. Liu. Inequivalent solutions for Dirac spin-1/2 particles under conservation of parity II. Journal of Mathematical Physics, 52(4):043509, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p043509_s1.

Zhang:2011:GSC

[760] Ting Zhang. Global solutions of compressible barotropic Navier–Stokes equations with a density-dependent viscosity coefficient. Journal of Mathematical Physics, 52(4):043510, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p043510_s1.

Manton:2011:GEN

[761] Nicholas S. Manton and Norman A. Rink. Geometry and energy of non-Abelian vortices. Journal of Mathematical Physics, 52(4):043511, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p043511_s1.

Zhang:2011:CSG

[762] Jingjun Zhang and Yanfeng Guo. On the convergence of the solution for a generalized Zakharov system. Journal of Mathematical Physics, 52(4): 043512, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p043512_s1.

Vojcak:2011:CIM

[763] Petr Vojcák. On complete integrability of the Mikhailov–Novikov–Wang system. Journal of Mathematical Physics, 52(4):043513, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/
p043513_s1.

Hassani:2011:WER

[764] A. Hassani. Wave equation on Riemannian symmetric spaces. Journal of Mathematical Physics, 52(4):043514, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p043514_s1.

Gil:2011:ERN

[765] M. I. Gil. An estimate for the resolvent of a non-self-adjoint differential operator on the half-line. Journal of Mathematical Physics, 52(4): 043515, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p043515_s1.

Zhang:2011:BSD

[766] Guoping Zhang. Breather solutions of the discrete nonlinear Schrödinger equations with sign changing nonlinearity. Journal of Mathematical Physics, 52(4):043516, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043516_s1.

Yu:2011:GFT

[767] Fajun Yu. A generalized fractional trace identity and Hamiltonian structure of the fractional soliton equation hierarchy. Journal of Mathematical Physics, 52(4):043517, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043517_s1.

Baricz:2011:IRF

[768] Árpád Baricz, Dragana Jankov, and Tibor K. Pogány. Integral representation of first kind Kapteyn series. Journal of Mathematical Physics, 52(4):043518, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043518_s1.

Caroca:2011:GID

[769] R. Caroca, N. Merino, P. Salgado, and O. Valdivia. Generating infinitedimensional algebras from loop algebras by expanding Maurer–Cartan forms. Journal of Mathematical Physics, 52(4):043519, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/
p043519_s1.

Ruggieri:2011:ESN

[770] Marianna Ruggieri and Antonino Valenti. Exact solutions for a nonlinear model of dissipative media. Journal of Mathematical Physics, 52(4): 043520, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i4/p043520_s1.

Havlicek:2011:CQA

[771] Miloslav Havlícek and Severin Posta. Center of quantum algebra U′_{q}(so_{3}). Journal of Mathematical Physics, 52(4):043521, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i4/p043521_s1.

Castillo-Perez:2011:DEE

[772] Raúl Castillo-Pérez, Vladislav V. Kravchenko, Héctor Oviedo-Galdeano, and Vladimir S. Rabinovich. Dispersion equation and eigenvalues for quantum wells using spectral parameter power series. Journal of Mathematical Physics, 52(4):043522, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i4/p043522_s1.

Qin:2011:BPA

[773] Bo Qin, Bo Tian, Li-Cai Liu, Ming Wang, Zhi-Qiang Lin, and Wen-Jun Liu. Bell-polynomial approach and N-soliton solution for the extended Lotka–Volterra equation in plasmas. Journal of Mathematical Physics, 52(4):043523, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i4/p043523_s1.

Hamzavi:2011:CSD

[774] M. Hamzavi. Comment on “Solution of Dirac equation with spin and pseudospin symmetry for an anharmonic oscillator”, [J. Math. Phys. 52, 013506 (2011)]. Journal of Mathematical Physics, 52(4):044101, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p044101_s1. See [624].

El-Jaick:2011:ELC

[775] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. Erratum: “A limit of the confluent Heun equation and the Schrödinger equation for an inverted potential and for an electric dipole”, [J. Math. Phys. 50, 123511 (2009)]. Journal of Mathematical Physics, 52(4):049901, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p049901_s1. See [20].

Yang:2011:ENT

[776] Chuan Fu Yang. Erratum: “New trace formulae for a quadratic pencil of the Schrödinger operator”, [J. Math. Phys. 51, 033506 (2010)]. Journal of Mathematical Physics, 52(4):049902, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i4/p049902_s1. See [140].

Cotfas:2011:HTO

[777] Nicolae Cotfas and Liviu Adrian Cotfas. Hypergeometric type operators and their supersymmetric partners. Journal of Mathematical Physics, 52 (5):052101, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052101_s1.

Kellendonk:2011:LTH

[778] Johannes Kellendonk, Konstantin Pankrashkin, and Serge Richard. Levinson’s theorem and higher degree traces for Aharonov–Bohm operators. Journal of Mathematical Physics, 52(5):052102, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p052102_s1.

Li:2011:CFP

[779] Yuan Li. Characterizations of fixed points of quantum operations. Journal of Mathematical Physics, 52(5):052103, May 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i5/p052103_s1.

Viscondi:2011:SPC

[780] Thiago F. Viscondi and Marcus A. M. de Aguiar. Semiclassical propagator for SU(n) coherent states. Journal of Mathematical Physics, 52(5): 052104, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052104_s1.

Korcyl:2011:SSY

[781] Piotr Korcyl. Solutions of D = 2 supersymmetric Yang–Mills quantum mechanics with SU(N) gauge group. Journal of Mathematical Physics, 52 (5):052105, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052105_s1.

Banerjee:2011:EST

[782] Pratyay Banerjee and B. Basu-Mallick. Exact solution of D_{N}-type quantum Calogero model through a mapping to free harmonic oscillators. Journal of Mathematical Physics, 52(5):052106, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p052106_s1.

Silvestre-Brac:2011:DRA

[783] Bernard Silvestre-Brac and Claude Semay. Duality relations in the auxiliary field method. Journal of Mathematical Physics, 52(5):052107, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i5/p052107_s1.

Chen:2011:RCN

[784] Li Chen and Ji Oon Lee. Rate of convergence in nonlinear Hartree dynamics with factorized initial data. Journal of Mathematical Physics, 52 (5):052108, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052108_s1.

Weder:2011:EAB

[785] Ricardo Weder. The electric Aharonov–Bohm effect. Journal of Mathematical Physics, 52(5):052109, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p052109_s1.

Benguria:2011:EAB

[786] Rafael D. Benguria and Gonzalo A. Bley. Exact asymptotic behavior of the Pekar–Tomasevich functional. Journal of Mathematical Physics, 52 (5):052110, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052110_s1.

Friedland:2011:EER

[787] Shmuel Friedland and Gilad Gour. An explicit expression for the relative entropy of entanglement in all dimensions. Journal of Mathematical Physics, 52(5):052201, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i5/p052201_s1.

Shirokov:2011:ERQ

[788] M. E. Shirokov. Entropy reduction of quantum measurements. Journal of Mathematical Physics, 52(5):052202, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p052202_s1.

Brandt:2011:SAC

[789] Friedemann Brandt. Supersymmetry algebra cohomology. III. Primitive elements in four and five dimensions. Journal of Mathematical Physics, 52 (5):052301, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052301_s1.

vanEkeren:2011:HLT

[790] Jethro van Ekeren. Higher level twisted Zhu algebras. Journal of Mathematical Physics, 52(5):052302, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p052302_s1.

Ikhdair:2011:ASS

[791] Sameer M. Ikhdair. An approximate κ state solutions of the Dirac equation for the generalized Morse potential under spin and pseudospin symmetry. Journal of Mathematical Physics, 52(5):052303, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/
p052303_s1.

Mendes:2011:IDC

[792] R. Vilela Mendes. An infinite-dimensional calculus for generalized connections on hypercubic lattices. Journal of Mathematical Physics, 52(5): 052304, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052304_s1.

Brini:2011:URM

[793] Andrea Brini, Marcos Mariño, and Sébastien Stevan. The uses of the refined matrix model recursion. Journal of Mathematical Physics, 52(5): 052305, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p052305_s1.

Schwartz:2011:TGR

[794] Charles Schwartz. Tachyons in general relativity. Journal of Mathematical Physics, 52(5):052501, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p052501_s1.

Freidel:2011:CSL

[795] Laurent Freidel and Etera R. Livine. U(N) coherent states for loop quantum gravity. Journal of Mathematical Physics, 52(5):052502, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i5/p052502_s1.

Krisch:2011:LCC

[796] J. P. Krisch and E. N. Glass. Levi-Civita cylinders with fractional angular deficit. Journal of Mathematical Physics, 52(5):052503, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/
p052503_s1.

Michel:2011:GIM

[797] B. Michel. Geometric invariance of mass-like asymptotic invariants. Journal of Mathematical Physics, 52(5):052504, May 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i5/p052504_s1.

Kinlaw:2011:RLR

[798] Paul A. Kinlaw. Refocusing of light rays in space–time. Journal of Mathematical Physics, 52(5):052505, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p052505_s1.

Catania:2011:LSE

[799] Davide Catania. Length-scale estimates for the 3D simplified Bardina magnetohydrodynamic model. Journal of Mathematical Physics, 52(5): 053101, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p053101_s1.

Litvinov:2011:MLT

[800] W. G. Litvinov. Model for laminar and turbulent flows of viscous and nonlinear viscous non-Newtonian fluids. Journal of Mathematical Physics, 52(5):053102, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i5/p053102_s1.

Zhang:2011:SST

[801] Zujin Zhang, Zheng an Yao, Ming Lu, and Lidiao Ni. Some Serrin-type regularity criteria for weak solutions to the Navier–Stokes equations. Journal of Mathematical Physics, 52(5):053103, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p053103_s1.

Heo:2011:QSP

[802] Jaeseong Heo and Un Cig Ji. Quantum stochastic processes for maps on Hilbert C^{∗}-modules. Journal of Mathematical Physics, 52(5):053501, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i5/p053501_s1.

Yau:2011:HYB

[803] Donald Yau. The Hom–Yang–Baxter equation and Hom–Lie algebras. Journal of Mathematical Physics, 52(5):053502, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p053502_s1.

Tsuchida:2011:SMG

[804] Takayuki Tsuchida. Systematic method of generating new integrable systems via inverse Miura maps. Journal of Mathematical Physics, 52 (5):053503, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p053503_s1.

Chen:2011:CPD

[805] Guowang Chen, Weifang Rui, and Xiangying Chen. Cauchy problem for a damped generalized IMBq equation. Journal of Mathematical Physics, 52 (5):053504, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p053504_s1.

Li:2011:MBS

[806] Gongbao Li, Shuangjie Peng, and Chunhua Wang. Multi-bump solutions for the nonlinear Schrödinger–Poisson system. Journal of Mathematical Physics, 52(5):053505, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i5/p053505_s1.

Muller-Hill:2011:EDH

[807] J. Müller-Hill and M. R. Zirnbauer. Equivalence of domains for hyperbolic Hubbard–Stratonovich transformations. Journal of Mathematical Physics, 52(5):053506, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i5/p053506_s1.

Carillo:2011:MKV

[808] Sandra Carillo and Cornelia Schiebold. Matrix Korteweg–de Vries and modified Korteweg–de Vries hierarchies: Noncommutative soliton solutions. Journal of Mathematical Physics, 52(5):053507, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/
p053507_s1.

Grammaticos:2011:QRT

[809] B. Grammaticos, A. Ramani, K. M. Tamizhmani, and R. Willox. On Quispel–Roberts–Thompson extensions and integrable correspondences. Journal of Mathematical Physics, 52(5):053508, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p053508_s1.

Cochran:2011:EPO

[810] Caroline M. Cochran, Raymond G. McLenaghan, and Roman G. Smirnov. Equivalence problem for the orthogonal webs on the 3-sphere. Journal of Mathematical Physics, 52(5):053509, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/p053509_s1.

DeSole:2011:CSL

[811] Alberto De Sole, Pedram Hekmati, and Victor G. Kac. Calculus structure on the Lie conformal algebra complex and the variational complex. Journal of Mathematical Physics, 52(5):053510, May 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i5/p053510_s1.

Tanisli:2011:OME

[812] M. Tanisli and M. E. Kansu. Octonionic Maxwell’s equations for biisotropic media. Journal of Mathematical Physics, 52(5):053511, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i5/p053511_s1.

Ling-yun:2011:DMS

[813] Xie Ling-yun and Fan Tian-you. The Dugdale model for a semi-infinite crack in a strip of two-dimensional decagonal quasicrystals. Journal of Mathematical Physics, 52(5):053512, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p053512_s1.

Ling:2011:LWS

[814] Liming Ling and Q. P. Liu. A long waves-short waves model: Darboux transformation and soliton solutions. Journal of Mathematical Physics, 52(5):053513, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i5/p053513_s1.

Baleanu:2011:FCF

[815] Dumitru Baleanu and Sergiu I. Vacaru. Fractional curve flows and solitonic hierarchies in gravity and geometric mechanics. Journal of Mathematical Physics, 52(5):053514, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p053514_s1.

Cheng:2011:ASB

[816] Jipeng Cheng, Kelei Tian, and Jingsong He. The additional symmetries for the BTL and CTL hierarchies. Journal of Mathematical Physics, 52 (5):053515, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i5/p053515_s1.

Vlasov:2011:EDC

[817] Alexander Yu. Vlasov. Extension of Dirac’s chord method to the case of a nonconvex set by the use of quasi-probability distributions. Journal of Mathematical Physics, 52(5):053516, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p053516_s1.

Turzi:2011:CDO

[818] Stefano S. Turzi. On the Cartesian definition of orientational order parameters. Journal of Mathematical Physics, 52(5):053517, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i5/
p053517_s1.

Zhang:2011:PSN

[819] Min-Cang Zhang and Guo-Qing Huang-Fu. Pseudospin symmetry for a new oscillatory ring-shaped noncentral potential. Journal of Mathematical Physics, 52(5):053518, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p053518_s1.

Cai:2011:SSS

[820] Zebin Cai, Jianchu Liang, Xiongping Xia, Yue Jiang, Haiqin Jin, and Lin Yi. Self-similar solitary wave family in Bessel lattice. Journal of Mathematical Physics, 52(5):053519, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i5/p053519_s1.

Park:2011:TCI

[821] Jungho Park. Thermosolutal convection at infinite Prandtl number with or without rotation: Bifurcation and stability in physical space. Journal of Mathematical Physics, 52(5):053701, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p053701_s1.

Peng:2011:WPL

[822] Yue-Jun Peng and Yong-Fu Yang. Well-posedness and long-time behavior of Lipschitz solutions to generalized extremal surface equations. Journal of Mathematical Physics, 52(5):053702, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p053702_s1.

Wang:2011:PNU

[823] Hai-Jun Wang. Publisher’s note: “Understanding quantum interference in general nonlocality”, [J. Math. Phys. 52, 033510 (2011)]. Journal of Mathematical Physics, 52(5):059901, May 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i5/p059901_s1. See [724].

Adachi:2011:SEE

[824] Tadayoshi Adachi and Atsuhide Ishida. Scattering in an external electric field asymptotically constant in time. Journal of Mathematical Physics, 52(6):062101, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p062101_s1.

Rube:2011:SPA

[825] Tomas Rube and Jay G. Wacker. The simplicity of perfect atoms: Degeneracies in supersymmetric hydrogen. Journal of Mathematical Physics, 52 (6):062102, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p062102_s1.

Vourdas:2011:QMM

[826] A. Vourdas. Quantum mechanics on Q/Z. Journal of Mathematical Physics, 52(6):062103, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p062103_s1.

Hiroshima:2011:NBR

[827] Fumio Hiroshima, Herbert Spohn, and Akito Suzuki. The no-binding regime of the Pauli–Fierz model. Journal of Mathematical Physics, 52 (6):062104, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p062104_s1.

Castro:2011:SNP

[828] P. G. Castro, R. Kullock, and F. Toppan. Snyder noncommutativity and pseudo-Hermitian Hamiltonians from a Jordanian twist. Journal of Mathematical Physics, 52(6):062105, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p062105_s1.

Bessa:2011:WFL

[829] V. Bessa and I. Guedes. Wave functions of log-periodic oscillators. Journal of Mathematical Physics, 52(6):062106, June 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i6/p062106_s1.

Alhaidari:2011:QME

[830] A. D. Alhaidari. Quantum mechanics without an equation of motion. Journal of Mathematical Physics, 52(6):062107, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/p062107_s1.

Zyczkowski:2011:GRD

[831] Karol Zyczkowski, Karol A. Penson, Ion Nechita, and Benoît Collins. Generating random density matrices. Journal of Mathematical Physics, 52(6):062201, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p062201_s1.

Bradler:2011:CGC

[832] Kamil Brádler, Tomas Jochym-O’Connor, and Rocío Jáuregui. Capacities of Grassmann channels. Journal of Mathematical Physics, 52(6): 062202, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p062202_s1.

Rastegin:2011:UCB

[833] Alexey E. Rastegin. Upper continuity bounds on the relative q-entropy for q > 1. Journal of Mathematical Physics, 52(6):062203, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/
p062203_s1.

Alba:2011:RQM

[834] David Alba, Horace W. Crater, and Luca Lusanna. Relativistic quantum mechanics and relativistic entanglement in the rest-frame instant form of dynamics. Journal of Mathematical Physics, 52(6):062301, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p062301_s1.

Ferrara:2011:TCB

[835] Sergio Ferrara, Alessio Marrani, Emanuele Orazi, Raymond Stora, and Armen Yeranyan. Two-center black holes duality-invariants for stu model and its lower-rank descendants. Journal of Mathematical Physics, 52(6): 062302, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p062302_s1.

Farkas:2011:GES

[836] Szilard Farkas and Emil J. Martinec. Gravity from the extension of spatial diffeomorphisms. Journal of Mathematical Physics, 52(6):062501, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p062501_s1.

Carter:2011:TFP

[837] John H. Carter and Michael Lieber. Thomas-forbidden particle capture. Journal of Mathematical Physics, 52(6):062901, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/p062901_s1.

Kanki:2011:KEC

[838] Kazuki Kanki, Satoshi Tanaka, and Tomio Petrosky. Kinetic equations for classical and quantum Brownian particles and eigenfunction expansions as generalized functions. Journal of Mathematical Physics, 52(6): 063301, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p063301_s1.

ElKaabouchi:2011:CAL

[839] A. El Kaabouchi, C. J. Ou, J. C. Chen, G. Z. Su, and Q. A. Wang. A counterexample against the Lesche stability of a generic entropy functional. Journal of Mathematical Physics, 52(6):063302, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/
p063302_s1.

Arita:2011:SSC

[840] Chikashi Arita and Kohei Motegi. Spin-spin correlation functions of the q-valence-bond-solid state of an integer spin model. Journal of Mathematical Physics, 52(6):063303, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i6/p063303_s1.

Finkelshtein:2011:BJCa

[841] Dmitri L. Finkelshtein, Yuri G. Kondratiev, Oleksandr V. Kutoviy, and Eugene Lytvynov. Binary jumps in continuum. I. Equilibrium processes and their scaling limits. Journal of Mathematical Physics, 52(6):063304, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p063304_s1.

Molinari:2011:CES

[842] Luca Guido Molinari and Giuseppe Lacagnina. Counting the exponents of single transfer matrices. Journal of Mathematical Physics, 52(6): 063501, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p063501_s1.

Soloviev:2011:MMA

[843] M. A. Soloviev. Moyal multiplier algebras of the test function spaces of type S. Journal of Mathematical Physics, 52(6):063502, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/
p063502_s1.

Guo:2011:RLR

[844] Zhengguang Guo and Sadek Gala. Remarks on logarithmical regularity criteria for the Navier–Stokes equations. Journal of Mathematical Physics, 52(6):063503, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p063503_s1.

Coulembier:2011:HSQ

[845] K. Coulembier and H. De Bie. Hilbert space for quantum mechanics on superspace. Journal of Mathematical Physics, 52(6):063504, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p063504_s1.

Bakke:2011:RLA

[846] K. Bakke, H. Belich, and E. O. Silva. Relativistic Landau–Aharonov– Casher quantization based on the Lorentz symmetry violation background. Journal of Mathematical Physics, 52(6):063505, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/
p063505_s1.

Xia:2011:HGR

[847] Baoqiang Xia. A hierarchy of Garnier–Rosochatius systems. Journal of Mathematical Physics, 52(6):063506, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p063506_s1.

Boyallian:2011:RST

[848] Carina Boyallian and Vanesa Meinardi. Representations of a symplectic type subalgebra of W_{∞}^{N}. Journal of Mathematical Physics, 52(6): 063507, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p063507_s1.

Chen:2011:ESS

[849] Shouxin Chen and Yuqiong Lei. Existence of steady-state solutions in a nonlinear photonic lattice model. Journal of Mathematical Physics, 52 (6):063508, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p063508_s1.

Gadella:2011:SGA

[850] M. Gadella, J. Negro, L. M. Nieto, G. P. Pronko, and M. Santander. Spectrum generating algebras for the free motion in S^{3}. Journal of Mathematical Physics, 52(6):063509, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i6/p063509_s1.

Wu:2011:MFB

[851] Dongsheng Wu. Mixed fractional Brownian sheets and their applications. Journal of Mathematical Physics, 52(6):063510, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/p063510_s1.

Bokhari:2011:WES

[852] Ashfaque H. Bokhari, Ahmad Y. Al-Dweik, A. H. Kara, M. Karim, and F. D. Zaman. Wave equation on spherically symmetric Lorentzian metrics. Journal of Mathematical Physics, 52(6):063511, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i6/p063511_s1.

Golenia:2011:PDI

[853] Sylvain Golénia and Christoph Schumacher. The problem of deficiency indices for discrete Schrödinger operators on locally finite graphs. Journal of Mathematical Physics, 52(6):063512, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p063512_s1.

Hryniv:2011:AUS

[854] Rostyslav O. Hryniv. Analyticity and uniform stability in the inverse spectral problem for Dirac operators. Journal of Mathematical Physics, 52(6):063513, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p063513_s1.

Giller:2011:SSH

[855] Stefan Giller. The semiclassical small-h limit of loci of roots of subdominant solutions for polynomial potentials. Journal of Mathematical Physics, 52(6):063514, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p063514_s1.

Bai:2011:GCY

[856] Chengming Bai, Li Guo, and Xiang Ni. Generalizations of the classical Yang–Baxter equation and O-operators. Journal of Mathematical Physics, 52(6):063515, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p063515_s1.

Alvarado:2011:SAD

[857] Carlos Alvarado and Paolo Amore. Spectroscopy of annular drums and quantum rings: Perturbative and nonperturbative results. Journal of Mathematical Physics, 52(6):063516, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p063516_s1.

Kravchenko:2011:DEE

[858] Vladislav V. Kravchenko and Ulises Velasco-García. Dispersion equation and eigenvalues for the Zakharov–Shabat system using spectral parameter power series. Journal of Mathematical Physics, 52(6):063517, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p063517_s1.

Klimek:2011:GBC

[859] Slawomir Klimek and Matt McBride. Global boundary conditions for a Dirac operator on the solid torus. Journal of Mathematical Physics, 52 (6):063518, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i6/p063518_s1.

Gade:2011:KFS

[860] R. M. Gade. q^{2}-Kampé de Fériet series and sums of continuous dual q^{±2}-Hahn polynomials. Journal of Mathematical Physics, 52(6):063519, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i6/p063519_s1.

Avazzadeh:2011:SDS

[861] Z. Avazzadeh, Z. Beygi Rizi, F. M. Maalek Ghaini, and G. B. Loghmani. A semi-discrete scheme for solving nonlinear hyperbolic-type partial integro-differential equations using radial basis functions. Journal of Mathematical Physics, 52(6):063520, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p063520_s1.

Bagarello:2011:NPB

[862] F. Bagarello. Nonlinear pseudo-bosons. Journal of Mathematical Physics, 52(6):063521, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i6/p063521_s1.

Allaire:2011:EHL

[863] Grégoire Allaire, Andro Mikelić, and Andrey Piatnitski. Erratum: “Homogenization of the linearized ionic transport equations in rigid periodic porous media”, [J. Math. Phys. 51, 123103 (2010)]. Journal of Mathematical Physics, 52(6):063701, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i6/p063701_s1. See [569].

Tian:2011:GCD

[864] Lixin Tian, Yujuan Wang, and Jiangbo Zhou. Global conservative and dissipative solutions of a coupled Camassa–Holm equations. Journal of Mathematical Physics, 52(6):063702, June 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i6/p063702_s1.

Le:2011:DSG

[865] Van-Hoang Le, Thanh-Tu Phan, and Cat-Tuong Truong. On the SO(10, 2) dynamical symmetry group of the MICZ–Kepler problem in a nine-dimensional space. Journal of Mathematical Physics, 52(7):072101, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p072101_s1.

Calabrese:2011:PBA

[866] Fabio F. G. Calabrese. Pseudo-bosons arising from standard ladder operators. Journal of Mathematical Physics, 52(7):072102, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/
p072102_s1.

Herdeiro:2011:CSR

[867] Carlos A. R. Herdeiro, Marco O. P. Sampaio, and Jaime E. Santos. The Casimir spectrum revisited. Journal of Mathematical Physics, 52(7): 072103, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p072103_s1.

Carinena:2011:QFP

[868] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. The quantum free particle on spherical and hyperbolic spaces: a curvature dependent approach. Journal of Mathematical Physics, 52(7):072104, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p072104_s1.

Miyadera:2011:URJ

[869] Takayuki Miyadera. Uncertainty relations for joint localizability and joint measurability in finite-dimensional systems. Journal of Mathematical Physics, 52(7):072105, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p072105_s1.

Schmidt:2011:LSA

[870] Bernd Schmidt. Localized spectral asymptotics for boundary value problems and correlation effects in the free Fermi gas in general domains. Journal of Mathematical Physics, 52(7):072106, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/p072106_s1.

Kihara:2011:GSD

[871] Hironobu Kihara. Generalized self-duality equations of polynomial type in Yang–Mills theories. Journal of Mathematical Physics, 52(7):072301, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p072301_s1.

Niemi:2011:SFN

[872] Antti J. Niemi and Andrzej Wereszczyński. On solutions to the “Faddeev–Niemi” equations. Journal of Mathematical Physics, 52(7): 072302, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p072302_s1.

Antoniadis:2011:EPG

[873] Ignatios Antoniadis, Lars Brink, and George Savvidy. Extensions of the Poincaré group. Journal of Mathematical Physics, 52(7):072303, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p072303_s1.

Han:2011:DSF

[874] Muxin Han. 4-dimensional spin-foam model with quantum Lorentz group. Journal of Mathematical Physics, 52(7):072501, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/
p072501_s1.

Zhao:2011:ABT

[875] Wenqiang Zhao and Yangrong Li. Asymptotic behavior of twodimensional stochastic magneto-hydrodynamics equations with additive noises. Journal of Mathematical Physics, 52(7):072701, July 2011.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/
p072701_s1.

Chen:2011:GSI

[876] Guanggan Chen, Jinqiao Duan, and Jian Zhang. Geometric shape of invariant manifolds for a class of stochastic partial differential equations. Journal of Mathematical Physics, 52(7):072702, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/p072702_s1.

Akerblom:2011:NCS

[877] N. Akerblom, G. Cornelissen, G. Stavenga, and J. W. van Holten. Nonrelativistic Chern–Simons vortices on the torus. Journal of Mathematical Physics, 52(7):072901, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p072901_s1.

Bru:2011:MTB

[878] J.-B. Bru, W. de Siqueira Pedra, and A.-S. Dömel. A microscopic twoband model for the electron-hole asymmetry in high-T_{c} superconductors and reentering behavior. Journal of Mathematical Physics, 52(7): 073301, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p073301_s1.

Bhattacharya:2011:PIT

[879] Satyabrata Bhattacharya, Sudip Chattopadhyay, Pinaki Chaudhury, and Jyotipratim Ray Chaudhuri. Phase induced transport of a Brownian particle in a periodic potential in the presence of an external noise: a semiclassical treatment. Journal of Mathematical Physics, 52(7):073302, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073302_s1.

Carvalho:2011:SOD

[880] S. L. Carvalho, C. R. de Oliveira, and R. A. Prado. Sparse onedimensional discrete Dirac operators II: Spectral properties. Journal of Mathematical Physics, 52(7):073501, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i7/p073501_s1.

Kouloukas:2011:PYB

[881] Theodoros E. Kouloukas and Vassilios G. Papageorgiou. Poisson Yang– Baxter maps with binomial Lax matrices. Journal of Mathematical Physics, 52(7):073502, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p073502_s1.

Zhang:2011:IHR

[882] Ling Zhang and Dafeng Zuo. Integrable hierarchies related to the Kuper– CH spectral problem. Journal of Mathematical Physics, 52(7):073503, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073503_s1.

Capriotti:2011:DMS

[883] S. Capriotti and H. Montani. Dirac method and symplectic submanifolds in the cotangent bundle of a factorizable Lie group. Journal of Mathematical Physics, 52(7):073504, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i7/p073504_s1.

Ferapontov:2011:CIT

[884] E. V. Ferapontov, A. V. Odesskii, and N. M. Stoilov. Classification of integrable two-component Hamiltonian systems of hydrodynamic type in 2 + 1 dimensions. Journal of Mathematical Physics, 52(7):073505, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073505_s1.

Kuhns:2011:HSC

[885] Chad Kuhns and Bennett Palmer. Helicoidal surfaces with constant anisotropic mean curvature. Journal of Mathematical Physics, 52(7): 073506, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p073506_s1.

Zullo:2011:BTE

[886] Federico Zullo. Bäcklund transformations for the elliptic Gaudin model and a Clebsch system. Journal of Mathematical Physics, 52(7):073507, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073507_s1.

Marx:2011:SCS

[887] C. A. Marx. Singular components of spectral measures for ergodic Jacobi matrices. Journal of Mathematical Physics, 52(7):073508, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073508_s1.

Cai:2011:BFS

[888] Xuanting Cai and Toufik Mansour. Bilinear forms on skein modules and steps in Dyck paths. Journal of Mathematical Physics, 52(7):073509, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073509_s1.

Eichelsbacher:2011:LDD

[889] Peter Eichelsbacher, Jens Sommerauer, and Michael Stolz. Large deviations for disordered bosons and multiple orthogonal polynomial ensembles. Journal of Mathematical Physics, 52(7):073510, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/p073510_s1.

Sun:2011:GVC

[890] Ye-Peng Sun and Hon-Wah Tam. On the general variable-coefficient KP equation with self-consistent sources. Journal of Mathematical Physics, 52(7):073511, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p073511_s1.

Yuen:2011:ABS

[891] Manwai Yuen. Analytical blowup solutions to the 2-dimensional isothermal Euler–Poisson equations of gaseous stars II. Journal of Mathematical Physics, 52(7):073512, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p073512_s1.

Vakhnenko:2011:BIM

[892] Oleksiy O. Vakhnenko. Bend-imitating models of abruptly bent electron waveguides. Journal of Mathematical Physics, 52(7):073513, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073513_s1.

Tanasa:2011:GBR

[893] Adrian Tanasa. Generalization of the Bollobás–Riordan polynomial for tensor graphs. Journal of Mathematical Physics, 52(7):073514, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i7/p073514_s1.

Bogdanski:2011:CSS

[894] P. Bogdanski and H. Ouerdane. Coulomb singularities in scattering wave functions of spin-orbit-coupled states. Journal of Mathematical Physics, 52(7):073515, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p073515_s1.

Qi:2011:EPM

[895] Jiangang Qi and Shaozhu Chen. Eigenvalue problems of the model from nonlocal continuum mechanics. Journal of Mathematical Physics, 52(7): 073516, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p073516_s1.

Boyack:2011:CHF

[896] Rufus Boyack and John Lekner. Confluent Heun functions and separation of variables in spheroidal coordinates. Journal of Mathematical Physics, 52(7):073517, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i7/p073517_s1.

Lin:2011:SSC

[897] Wen-Tsan Lin and Choon-Lin Ho. Similarity solutions of a class of perturbative Fokker–Planck equation. Journal of Mathematical Physics, 52 (7):073701, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i7/p073701_s1.

Fang:2011:RRS

[898] Beixiang Fang, Ya-Guang Wang, and Hairong Yuan. Reflection and refraction of shocks on an interface with a reflected rarefaction wave. Journal of Mathematical Physics, 52(7):073702, July 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i7/p073702_s1.

Yuen:2011:ESS

[899] Manwai Yuen. Erratum: “Self-similar blowup solutions to the 2component Camassa–Holm equations”, [J. Math. Phys. 51, 093524 (2010)]. Journal of Mathematical Physics, 52(7):079901, July 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i7/
p079901_s1. See [454].

Daoud:2011:POP

[900] M. Daoud and M. R. Kibler. Phase operators, phase states and vector phase states for SU_{3} and SU_{2,1}. Journal of Mathematical Physics, 52(8): 082101, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p082101_s1.

Childs:2011:LTG

[901] Andrew M. Childs and D. J. Strouse. Levinson’s theorem for graphs. Journal of Mathematical Physics, 52(8):082102, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/p082102_s1.

Flego:2011:SFR

[902] S. P. Flego, A. Plastino, and A. R. Plastino. Special features of the relation between Fisher information and Schrödinger eigenvalue equation. Journal of Mathematical Physics, 52(8):082103, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/p082103_s1.

Vicary:2011:CCC

[903] Jamie Vicary. Completeness of †-categories and the complex numbers. Journal of Mathematical Physics, 52(8):082104, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/p082104_s1.

Iqbal:2011:GCS

[904] Shahid Iqbal and Farhan Saif. Generalized coherent states and their statistical characteristics in power-law potentials. Journal of Mathematical Physics, 52(8):082105, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p082105_s1.

Aleixo:2011:EOA

[905] A. N. F. Aleixo and A. B. Balantekin. Exponential operators and the algebraic description of quantum confined systems. Journal of Mathematical Physics, 52(8):082106, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i8/p082106_s1.

Reeb:2011:HPM

[906] David Reeb, Michael J. Kastoryano, and Michael M. Wolf. Hilbert’s projective metric in quantum information theory. Journal of Mathematical Physics, 52(8):082201, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p082201_s1.

DAriano:2011:EQP

[907] Giacomo Mauro D’Ariano, Paolo Perinotti, and Michal Sedlák. Extremal quantum protocols. Journal of Mathematical Physics, 52(8):082202, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p082202_s1.

Choe:2011:EPR

[908] Kwangseok Choe and Jongmin Han. Existence and properties of radial solutions in the self-dual Chern–Simons O(3) sigma model. Journal of Mathematical Physics, 52(8):082301, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p082301_s1.

Bellucci:2011:TCS

[909] Stefano Bellucci, Alessio Marrani, and Raju Roychowdhury. Topics in cubic special geometry. Journal of Mathematical Physics, 52(8):082302, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p082302_s1.

Akhmeteli:2011:ORF

[910] Andrey Akhmeteli. One real function instead of the Dirac spinor function. Journal of Mathematical Physics, 52(8):082303, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/
p082303_s1.

Fullwood:2011:GSV

[911] James Fullwood. On generalized Sethi–Vafa–Witten formulas. Journal of Mathematical Physics, 52(8):082304, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p082304_s1.

Gomes:2011:CGT

[912] Henrique de A. Gomes. Classical gauge theory in Riem. Journal of Mathematical Physics, 52(8):082501, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p082501_s1.

Llibre:2011:POI

[913] Jaume Llibre and Luci Any Roberto. On the periodic orbits and the integrability of the regularized Hill lunar problem. Journal of Mathematical Physics, 52(8):082701, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i8/p082701_s1.

Liu:2011:MAW

[914] Qihuai Liu and Dingbian Qian. Modulated amplitude waves with nonzero phases in Bose–Einstein condensates. Journal of Mathematical Physics, 52(8):082702, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p082702_s1.

Yanovski:2011:ROG

[915] A. B. Yanovski. On the recursion operators for the Gerdjikov, Mikhailov, and Valchev system. Journal of Mathematical Physics, 52(8):082703, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p082703_s1.

Grenier:2011:WAR

[916] M. Grenier, H.-R. Jauslin, C. Klein, and V. B. Matveev. Wave attraction in resonant counter-propagating wave systems. Journal of Mathematical Physics, 52(8):082704, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p082704_s1.

Aharonovich:2011:GFC

[917] I. Aharonovich and L. P. Horwitz. On the Green-functions of the classical off-shell electrodynamics under the manifestly covariant relativistic dynamics of Stueckelberg. Journal of Mathematical Physics, 52(8): 082901, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p082901_s1.

delaLlave:2011:BQG

[918] Rafael de la Llave and Enrico Valdinoci. L^{p}-bounds for quasi-geostrophic equations via functional analysis. Journal of Mathematical Physics, 52(8):083101, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p083101_s1.

Silva:2011:EPF

[919] A. T. Silva, E. K. Lenzi, L. R. Evangelista, M. K. Lenzi, H. V. Ribeiro, and A. A. Tateishi. Exact propagator for a Fokker–Planck equation, first passage time distribution, and anomalous diffusion. Journal of Mathematical Physics, 52(8):083301, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i8/p083301_s1.

Kozlowski:2011:FFC

[920] K. K. Kozlowski. On form factors of the conjugated field in the nonlinear Schrödinger model. Journal of Mathematical Physics, 52(8):083302, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083302_s1.

Clark:2011:DBQ

[921] Jeremy Clark, W. De Roeck, and Christian Maes. Diffusive behavior from a quantum master equation. Journal of Mathematical Physics, 52(8): 083303, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p083303_s1.

Ssemaganda:2011:USS

[922] Vincent Ssemaganda, Katharina Holstein, and Gerald Warnecke. Uniqueness of steady-state solutions for thermodynamically consistent Becker– Döring models. Journal of Mathematical Physics, 52(8):083304, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083304_s1.

Somogyi:2011:AID

[923] Gábor Somogyi. Angular integrals in d dimensions. Journal of Mathematical Physics, 52(8):083501, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i8/p083501_s1.

Kim:2011:TAG

[924] Sung-Soo Kim and Jakob Palmkvist. N = 5 three-algebras and 5-graded Lie superalgebras. Journal of Mathematical Physics, 52(8):083502, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083502_s1.

Simpson:2011:VCC

[925] G. Simpson and I. Zwiers. Vortex collapse for the L^{2}-critical nonlinear Schrödinger equation. Journal of Mathematical Physics, 52(8): 083503, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p083503_s1.

Abdalla:2011:LAA

[926] M. Sebawe Abdalla and P. G. L. Leach. Lie algebraic approach and quantum treatment of an anisotropic charged particle via the quadratic invariant. Journal of Mathematical Physics, 52(8):083504, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/
p083504_s1.

Schulze-Halberg:2011:HOD

[927] Axel Schulze-Halberg. Higher order Darboux transformations in two dimensions, linearizable auxiliary equations, and invariant potentials. Journal of Mathematical Physics, 52(8):083505, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/p083505_s1.

Boiti:2011:HOP

[928] M. Boiti, F. Pempinelli, and A. K. Pogrebkov. Heat operator with pure soliton potential: Properties of Jost and dual Jost solutions. Journal of Mathematical Physics, 52(8):083506, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p083506_s1.

Fu:2011:SEK

[929] Qiang Fu and Qunguang Yang. On the structure of End_{u_{k}(2)}(Ω_{k}^{⊗r}). Journal of Mathematical Physics, 52(8):083507, August 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i8/p083507_s1.

Frohlich:2011:SHM

[930] Jürg Fröhlich, Zhou Gang, and Avy Soffer. Some Hamiltonian models of friction. Journal of Mathematical Physics, 52(8):083508, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083508_s1.

Nagoya:2011:HSS

[931] H. Nagoya. Hypergeometric solutions to Schrödinger equations for the quantum Painlevé equations. Journal of Mathematical Physics, 52(8): 083509, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p083509_s1.

Filippelli:2011:TDQ

[932] G. Filippelli. Time dependent quantum generators for the Galilei group. Journal of Mathematical Physics, 52(8):083510, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/p083510_s1.

Ke:2011:CEA

[933] Sanmin Ke, Wenli Yang, Chun Wang, Kexia Jiang, and Kangjie Shi. The classical exchange algebra of a Green–Schwarz sigma model on supercoset target space with z_{4m} grading. Journal of Mathematical Physics, 52(8): 083511, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p083511_s1.

Neate:2011:SBE

[934] Andrew Neate, Scott Reasons, and Aubrey Truman. The stochastic Burgers equation with vorticity: Semiclassical asymptotic series solutions with applications. Journal of Mathematical Physics, 52(8):083512, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083512_s1.

Gao:2011:DLA

[935] Shoulan Gao, Dong Liu, and Yufeng Pei. The deformed Lie algebra of differential operators of order at most one. Journal of Mathematical Physics, 52(8):083513, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i8/p083513_s1.

Mkrtchyan:2011:DND

[936] Ruben L. Mkrtchyan and Alexander P. Veselov. On duality and negative dimensions in the theory of Lie groups and symmetric spaces. Journal of Mathematical Physics, 52(8):083514, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p083514_s1.

Cacciapuoti:2011:PEE

[937] Claudio Cacciapuoti, Raffaele Carlone, and Rodolfo Figari. Perturbations of eigenvalues embedded at threshold: Two-dimensional solvable models. Journal of Mathematical Physics, 52(8):083515, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i8/
p083515_s1.

Gurses:2011:KE

[938] Metin Gürses and Asli Pekcan. 2 + 1 KdV(N ) equations. Journal of Mathematical Physics, 52(8):083516, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i8/p083516_s1.

Rogers:2011:PDM

[939] C. Rogers and W. K. Schief. The pulsrodon in 2+1-dimensional magnetogasdynamics: Hamiltonian structure and integrability. Journal of Mathematical Physics, 52(8):083701, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i8/p083701_s1.

Otto:2011:RBC

[940] Felix Otto and Christian Seis. Rayleigh–Bénard convection: Improved bounds on the Nusselt number. Journal of Mathematical Physics, 52(8): 083702, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p083702_s1.

Banica:2011:DSE

[941] Valeria Banica and Liviu I. Ignat. Dispersion for the Schrödinger equation on networks. Journal of Mathematical Physics, 52(8):083703, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i8/p083703_s1.

Rabeie:2011:CCS

[942] A. Rabeie and Z. Jalilian. Comment on “Coherent states on spheres”, [J. Math. Phys. 43, 1211 (2002)]. Journal of Mathematical Physics, 52(8): 084101, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p084101_s1. See [3].

Arda:2011:EMK

[943] Altug Arda and Ramazan Sever. Effective-mass Klein–Gordon–Yukawa problem for bound and scattering states. Journal of Mathematical Physics, 52(9):092101, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i9/p092101_s1.

Michishita:2011:SSY

[944] Yoji Michishita. On the spectrum of D = 2 supersymmetric Yang–Mills quantum mechanics. Journal of Mathematical Physics, 52(9):092102, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i9/p092102_s1.

Hall:2011:DSC

[945] Richard L. Hall, Nasser Saad, and K. D. Sen. Discrete spectra for confined and unconfined −a/r + br^{2} potentials in d-dimensions. Journal of Mathematical Physics, 52(9):092103, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p092103_s1.

Merkli:2011:EEQ

[946] Marco Merkli. Entanglement evolution via quantum resonances. Journal of Mathematical Physics, 52(9):092201, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p092201_s1.

Arrizabalaga:2011:DSA

[947] Naiara Arrizabalaga. Distinguished self-adjoint extensions of Dirac operators via Hardy–Dirac inequalities. Journal of Mathematical Physics, 52(9):092301, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p092301_s1.

Guimaraes:2011:FRZ

[948] M. S. Guimaraes and S. P. Sorella. A few remarks on the zero modes of the Faddeev–Popov operator in the Landau and maximal Abelian gauges. Journal of Mathematical Physics, 52(9):092302, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i9/
p092302_s1.

Ferrari:2011:SAM

[949] Ruggero Ferrari. A symmetric approach to the massive nonlinear sigma model. Journal of Mathematical Physics, 52(9):092303, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i9/
p092303_s1.

Hosseiny:2011:HRL

[950] Ali Hosseiny and Ali Naseh. On holographic realization of logarithmic Galilean conformal algebra. Journal of Mathematical Physics, 52 (9):092501, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p092501_s1.

Aazami:2011:LKBa

[951] Amir B. Aazami, Charles R. Keeton, and A. O. Petters. Lensing by Kerr black holes. I. General lens equation and magnification formula. Journal of Mathematical Physics, 52(9):092502, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p092502_s1.

Achilleos:2011:SSN

[952] V. Achilleos, G. Theocharis, P. G. Kevrekidis, N. I. Karachalios, F. K. Diakonos, and D. J. Frantzeskakis. Stationary states of a nonlinear Schrödinger lattice with a harmonic trap. Journal of Mathematical Physics, 52(9):092701, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i9/p092701_s1.

Wu:2011:ESS

[953] Xiaotai Wu, Litan Yan, Wenbing Zhang, and Yang Tang. Exponential stability of stochastic differential delay systems with delayed impulse effects. Journal of Mathematical Physics, 52(9):092702, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p092702_s1.

Chang:2011:GHP

[954] Jen-Hsu Chang. The Gould–Hopper polynomials in the Novikov–Veselov equation. Journal of Mathematical Physics, 52(9):092703, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p092703_s1.

Xie:2011:SCC

[955] Zhifu Xie, Kenyaita Hodge, Krystolyn Henderson, and Michael Westbrook. Super central configurations of the three-body problem under the inverse integer power law. Journal of Mathematical Physics, 52(9):092901, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p092901_s1.

Ferris:2011:OST

[956] Michael R. Ferris and Jonathan Gratus. The origin of the Schott term in the electromagnetic self force of a classical point charge. Journal of Mathematical Physics, 52(9):092902, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p092902_s1.

Kuetche:2011:FSF

[957] Victor K. Kuetche, Thomas B. Bouetou, and Timoleon C. Kofane. Fractal structure of ferromagnets: The singularity structure analysis. Journal of Mathematical Physics, 52(9):092903, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p092903_s1.

Whitehead:2011:IHD

[958] Jared P. Whitehead and Charles R. Doering. Internal heating driven convection at infinite Prandtl number. Journal of Mathematical Physics, 52(9):093101, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093101_s1.

Guo:2011:GSSa

[959] Zhenhua Guo, Jing Yang, and Lei Yao. Global strong solution for a threedimensional viscous liquid-gas two-phase flow model with vacuum. Journal of Mathematical Physics, 52(9):093102, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v52/i9/p093102_s1.

Marinucci:2011:AES

[960] Domenico Marinucci and Igor Wigman. On the area of excursion sets of spherical Gaussian eigenfunctions. Journal of Mathematical Physics, 52(9):093301, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093301_s1.

Xu:2011:UEC

[961] Shuai-Xia Xu, Yu-Qiu Zhao, and Jian-Rong Zhou. Universality for eigenvalue correlations from the unitary ensemble associated with a family of singular weights. Journal of Mathematical Physics, 52(9):093302, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i9/p093302_s1.

Redig:2011:WCL

[962] Frank Redig and Kiamars Vafayi. Weak coupling limits in a stochastic model of heat conduction. Journal of Mathematical Physics, 52 (9):093303, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093303_s1.

Klimek:2011:CLB

[963] Slawomir Klimek and Matt McBride. Classical limit of the d-bar operators on quantum domains. Journal of Mathematical Physics, 52 (9):093501, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093501_s1.

Wu:2011:DRE

[964] Yuhu Wu and Xiaoping Xue. Decay rate estimates for the quasi-linear Timoshenko system with nonlinear control and damping terms. Journal of Mathematical Physics, 52(9):093502, September 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i9/p093502_s1.

Vignat:2011:ONR

[965] C. Vignat. Old and new results about relativistic Hermite polynomials. Journal of Mathematical Physics, 52(9):093503, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i9/
p093503_s1.

Guo:2011:WMV

[966] Xiangqian Guo and Xuewen Liu. Whittaker modules over Virasorolike algebra. Journal of Mathematical Physics, 52(9):093504, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p093504_s1.

Jonsson:2011:DBM

[967] B. L. G. Jonsson, Yu. N. Ovchinnikov, I. M. Sigal, and F. S. T. Ting. Dynamics of breakup of multiple vortices in Gross–Pitaevskii equations of superfluids. Journal of Mathematical Physics, 52(9):093505, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p093505_s1.

Hushmandi:2011:WPF

[968] Ataabak B. Hushmandi and Morteza M. Rezaii. On warped product Finsler spaces of Landsberg type. Journal of Mathematical Physics, 52(9):093506, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093506_s1.

Habibullin:2011:DHT

[969] Ismagil Habibullin, Natalya Zheltukhina, and Alfia Sakieva. Discretization of hyperbolic type Darboux integrable equations preserving integrability. Journal of Mathematical Physics, 52(9):093507, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p093507_s1.

Doikou:2011:BPS

[970] Anastasia Doikou and Theodora Ioannidou. Bogomolny–Prasad– Sommerfeld monopoles and open spin chains. Journal of Mathematical Physics, 52(9):093508, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i9/p093508_s1.

Benci:2011:ESC

[971] Vieri Benci and Donato Fortunato. On the existence of stable charged Qballs. Journal of Mathematical Physics, 52(9):093701, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i9/
p093701_s1.

Ozemir:2011:VCN

[972] C. Özemir and F. Güngör. Variable coefficient nonlinear Schrödinger equations with four-dimensional symmetry groups and analysis of their solutions. Journal of Mathematical Physics, 52(9):093702, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i9/p093702_s1.

Laugesen:2011:SLE

[973] R. S. Laugesen and B. A. Siudeja. Sums of Laplace eigenvalues: Rotations and tight frames in higher dimensions. Journal of Mathematical Physics, 52(9):093703, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i9/p093703_s1.

Ma:2011:THA

[974] Hui Ma and Derchyi Wu. Twisted hierarchies associated with the generalized sine-Gordon equation. Journal of Mathematical Physics, 52(9):093704, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093704_s1.

Charalambopoulos:2011:AAS

[975] Antonios Charalambopoulos. An analytic algorithm for shape reconstruction from low-frequency moments. Journal of Mathematical Physics, 52(9):093705, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093705_s1.

Bartuccelli:2011:SCS

[976] Michele V. Bartuccelli and John D. Gibbon. Sharp constants in the Sobolev embedding theorem and a derivation of the Brezis–Gallouet interpolation inequality. Journal of Mathematical Physics, 52(9):093706, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i9/p093706_s1.

Young:2011:ESI

[977] Brent Young. Existence of spherical initial data with unit mass, zero energy, and virial less than −1/2 for the relativistic Vlasov–Poisson equation with attractive coupling. Journal of Mathematical Physics, 52(9):093707, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i9/p093707_s1.

Ou:2011:SSS

[978] Yaobin Ou and Peicheng Zhu. Spherically symmetric solutions to a model for phase transitions driven by configurational forces. Journal of Mathematical Physics, 52(9):093708, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i9/p093708_s1.

Iftimie:2011:EDM

[979] Viorel Iftimie and Radu Purice. Eigenfunctions decay for magnetic pseudodifferential operators. Journal of Mathematical Physics, 52(9):093709, September 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i9/p093709_s1.

Aktosun:2011:SEA

[980] Tuncay Aktosun, Martin Klaus, and Ricardo Weder. Small-energy analysis for the self-adjoint matrix Schrödinger operator on the half line. Journal of Mathematical Physics, 52(10):102101, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i10/
p102101_s1.

Capak:2011:ESS

[981] M. Çapak, Y. Cançelik, Ö. L. Ünsal, S. Atay, and B. Gönül. An extended scenario for the Schrödinger equation. Journal of Mathematical Physics, 52(10):102102, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i10/p102102_s1.

Ahlbrecht:2011:DQW

[982] Andre Ahlbrecht, Volkher B. Scholz, and Albert H. Werner. Disordered quantum walks in one lattice dimension. Journal of Mathematical Physics, 52(10):102201, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p102201_s1.

Slofstra:2011:LBE

[983] William Slofstra. Lower bounds on the entanglement needed to play XOR non-local games. Journal of Mathematical Physics, 52(10):102202, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p102202_s1.

Ablinger:2011:HSP

[984] Jakob Ablinger, Johannes Blümlein, and Carsten Schneider. Harmonic sums and polylogarithms generated by cyclotomic polynomials. Journal of Mathematical Physics, 52(10):102301, October 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i10/p102301_s1.

Tho:2011:DSG

[985] Nguyen Vien Tho and To Ba Ha. Dynamical stability of global vortex strings. Journal of Mathematical Physics, 52(10):102302, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p102302_s1.

Kim:2011:DQD

[986] Sejong Kim. Distances of qubit density matrices on Bloch sphere. Journal of Mathematical Physics, 52(10):102303, October 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i10/p102303_s1.

Morfa-Morales:2011:DQF

[987] Eric Morfa-Morales. Deformations of quantum field theories on de Sitter spacetime. Journal of Mathematical Physics, 52(10):102304, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p102304_s1.

Matone:2011:LRA

[988] Marco Matone and Roberto Volpato. Linear relations among holomorphic quadratic differentials and induced Siegel’s metric on M_{g}. Journal of Mathematical Physics, 52(10):102305, October 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i10/p102305_s1.

Aazami:2011:LKBb

[989] Amir B. Aazami, Charles R. Keeton, and A. O. Petters. Lensing by Kerr black holes. II: Analytical study of quasi-equatorial lensing observables. Journal of Mathematical Physics, 52(10):102501, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i10/
p102501_s1.

Beyer:2011:SMS

[990] Horst Reinhard Beyer. On the stability of the massive scalar field in Kerr space–time. Journal of Mathematical Physics, 52(10):102502, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p102502_s1.

Krisch:2011:VHF

[991] J. P. Krisch and E. N. Glass. Velocity and heat flow in a composite two fluid system. Journal of Mathematical Physics, 52(10):102503, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p102503_s1.

Hovhannisyan:2011:DET

[992] Gro Hovhannisyan. On Dirac equation on a time scale. Journal of Mathematical Physics, 52(10):102701, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p102701_s1.

Calogero:2011:IMB

[993] F. Calogero. An integrable many-body problem. Journal of Mathematical Physics, 52(10):102702, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i10/p102702_s1.

Smolyakov:2011:NEN

[994] Mikhail N. Smolyakov. Non-existence of non-topological solitons in some types of gauge field theories in Minkowski space. Journal of Mathematical Physics, 52(10):102901, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i10/p102901_s1. See erratum [1128].

Bertin:2011:CSP

[995] M. C. Bertin, B. M. Pimentel, and G. E. R. Zambrano. The canonical structure of Podolsky’s generalized electrodynamics on the nullplane. Journal of Mathematical Physics, 52(10):102902, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i10/
p102902_s1.

Celli:2011:PRE

[996] M. Celli, E. A. Lacomba, and E. Pérez-Chavela. On polygonal relative equilibria in the N-vortex problem. Journal of Mathematical Physics, 52(10):103101, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i10/p103101_s1.

Antonio:2011:ABA

[997] N. Cirilo António, N. Manojlović, and A. Stolin. Algebraic Bethe ansatz for deformed Gaudin model. Journal of Mathematical Physics, 52(10): 103501, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i10/p103501_s1.

Valls:2011:AFI

[998] Claudia Valls. Analytic first integrals for generalized Raychaudhuri equations. Journal of Mathematical Physics, 52(10):103502, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i10/
p103502_s1.

Plastino:2011:DPJ

[999] A. Plastino and M. C. Rocca. A direct proof of Jauregui–Tsallis’ conjecture. Journal of Mathematical Physics, 52(10):103503, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i10/
p103503_s1.

Chowdhury:2011:AGS

[1000] S. Hasibul Hassan Chowdhury and S. Twareque Ali. All the groups of signal analysis from the (1 + 1)-affine Galilei group. Journal of Mathematical Physics, 52(10):103504, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p103504_s1.

Grandati:2011:SRE

[1001] Yves Grandati. Solvable rational extensions of the Morse and Kepler– Coulomb potentials. Journal of Mathematical Physics, 52(10):103505, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i10/p103505_s1.

Ayadi:2011:IBS

[1002] V. Ayadi and L. Fehér. An integrable BC(n ) Sutherland model with two types of particles. Journal of Mathematical Physics, 52(10):103506, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p103506_s1.

Douglas:2011:CED

[1003] Andrew Douglas, Wainright Joseph, and Joe Repka. A classification of the embeddings of the Diamond Lie algebra into ∫↕(3, C) and ∫√(4, C),

and restrictions of irreducible modules. Journal of Mathematical Physics, 52(10):103507, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i10/p103507_s1.

Fakhri:2011:DOF

[1004] H. Fakhri and M. Lotfizadeh. Dirac operators on the fuzzy AdS_{2} with the spins 1/2 and 1. Journal of Mathematical Physics, 52(10):103508, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p103508_s1.

Fiore:2011:CIT

[1005] Gaetano Fiore and Laure Gouba. Class of invariants for the twodimensional time-dependent Landau problem and harmonic oscillator in a magnetic field. Journal of Mathematical Physics, 52(10):103509, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p103509_s1.

Sasakura:2011:TMA

[1006] Naoki Sasakura. Tensor models and 3-ary algebras. Journal of Mathematical Physics, 52(10):103510, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p103510_s1.

Mezzadri:2011:MTE

[1007] F. Mezzadri and N. J. Simm. Moments of the transmission eigenvalues, proper delay times, and random matrix theory. I. Journal of Mathematical Physics, 52(10):103511, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p103511_s1.

Tsujimoto:2011:JAO

[1008] Satoshi Tsujimoto, Luc Vinet, and Alexei Zhedanov. Jordan algebras and orthogonal polynomials. Journal of Mathematical Physics, 52(10): 103512, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2011:GEB

[1009] Qiaoyi Hu. Global existence and blow-up phenomena for a weakly dissipative periodic 2-component Camassa–Holm system. Journal of Mathematical Physics, 52(10):103701, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p103701_s1.

Guo:2011:GSSb

[1010] Boling Guo and Daiwen Huang. Global strong solutions of a simplified two-fluid model in plasma. Journal of Mathematical Physics, 52(10): 103702, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i10/p103702_s1.

Bizon:2011:DNT

[1011] Piotr Bizoń, Tadeusz Chmaj, and Nikodem Szpak. Dynamics near the threshold for blowup in the one-dimensional focusing nonlinear Klein– Gordon equation. Journal of Mathematical Physics, 52(10):103703, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p103703_s1.

Wang:2011:NDG

[1012] Jianyong Wang, Hanwei Hu, and Jun Yu. A new (2 + 1)-dimensional generalization of the modified Korteweg–de Vries equation and its infinitely many symmetries. Journal of Mathematical Physics, 52(10): 103704, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i10/p103704_s1.

Briane:2011:BSF

[1013] Marc Briane and Graeme W. Milton. Bounds on strong field magnetotransport in three-dimensional composites. Journal of Mathematical Physics, 52(10):103705, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i10/p103705_s1.

Yuan:2011:LWP

[1014] Jianjun Yuan. Local well-posedness of Chern–Simons–Higgs system in the Lorentz gauge. Journal of Mathematical Physics, 52(10):103706, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i10/p103706_s1.

Guan:2011:GWS

[1015] Chunxia Guan and Zhaoyang Yin. Global weak solutions and smooth solutions for a two-component Hunter–Saxton system. Journal of Mathematical Physics, 52(10):103707, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i10/p103707_s1.

Mantoiu:2011:PQP

[1016] Marius Măntoiu, Radu Purice, and Serge Richard. Positive quantization in the presence of a variable magnetic field. Journal of Mathematical Physics, 52(11):112101, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i11/p112101_s1.

Hall:2011:GSI

[1017] Richard L. Hall and Wolfgang Lucha. Geometric spectral inversion for singular potentials. Journal of Mathematical Physics, 52(11):112102, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i11/p112102_s1.

DeNittis:2011:ELW

[1018] G. De Nittis and M. Lein. Exponentially localized Wannier functions in periodic zero flux magnetic fields. Journal of Mathematical Physics, 52(11):112103, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p112103_s1. See erratum [2721].

Meng:2011:EJA

[1019] Guowu Meng. Euclidean Jordan algebras, hidden actions, and J–Kepler problems. Journal of Mathematical Physics, 52(11):112104, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p112104_s1.

Wang:2011:NAMa

[1020] Guozhong Wang, Hui Li, Yifeng Shen, and Biqin Dong. New approximate method to solve the Schrödinger equation with a Woods–Saxon-like potential. Journal of Mathematical Physics, 52(11):112105, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p112105_s1.

Audenaert:2011:CBQ

[1021] Koenraad M. R. Audenaert and Jens Eisert. Continuity bounds on the quantum relative entropy — II. Journal of Mathematical Physics, 52(11):112201, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p112201_s1.

King:2011:AOE

[1022] Christopher King and David K. Moser. Average output entropy for quantum channels. Journal of Mathematical Physics, 52(11):112202, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p112202_s1.

Crompton:2011:GBF

[1023] P. R. Crompton. Gröbner bases for finite-temperature quantum computing and their complexity. Journal of Mathematical Physics, 52 (11):112203, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p112203_s1.

Cordero-Carrion:2011:MSS

[1024] Isabel Cordero-Carrión, José María Ibáñez, and Juan Antonio Morales-Lladosa. Maximal slicings in spherical symmetry: Local existence and construction. Journal of Mathematical Physics, 52(11):112501, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p112501_s1.

Vignolo:2011:DSB

[1025] Stefano Vignolo, Luca Fabbri, and Roberto Cianci. Dirac spinors in Bianchi-I f(R)-cosmology with torsion. Journal of Mathematical Physics, 52(11):112502, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p112502_s1.

Vatandoost:2011:SOR

[1026] M. Vatandoost and Y. Bahrampour. Sphere orderings representation of space–time. Journal of Mathematical Physics, 52(11):112503, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p112503_s1.

Navarro:2011:GHS

[1027] J. J. Benavides Navarro and E. Minguzzi. Global hyperbolicity is stable in the interval topology. Journal of Mathematical Physics, 52(11):112504, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i11/p112504_s1.

Llibre:2011:GDS

[1028] Jaume Llibre, Marcelo Messias, and Paulo R. da Silva. Global dynamics of stationary solutions of the extended Fisher–Kolmogorov equation. Journal of Mathematical Physics, 52(11):112701, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i11/
p112701_s1.

Li:2011:NIC

[1029] Wenlei Li, Shaoyun Shi, and Bing Liu. Non-integrability of a class of Hamiltonian systems. Journal of Mathematical Physics, 52(11):112702, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i11/p112702_s1.

Demirkaya:2011:CST

[1030] Aslihan Demirkaya and Milena Stanislavova. Conditional stability theorem for the one dimensional Klein–Gordon equation. Journal of Mathematical Physics, 52(11):112703, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i11/p112703_s1.

Brezhnev:2011:DSD

[1031] Yurii V. Brezhnev, Simon L. Lyakhovich, and Alexey A. Sharapov. Dynamical systems defining Jacobi’s ϑ-constants. Journal of Mathematical Physics, 52(11):112704, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i11/p112704_s1.

Finkelshtein:2011:BJCb

[1032] Dmitri Finkelshtein, Yuri Kondratiev, Oleksandr Kutoviy, and Eugene Lytvynov. Binary jumps in continuum. II. Non-equilibrium process and a Vlasov-type scaling limit. Journal of Mathematical Physics, 52 (11):113301, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p113301_s1.

Wojtkiewicz:2011:AOH

[1033] Jacek Wojtkiewicz. Absence of orderings in Hubbard models. Journal of Mathematical Physics, 52(11):113302, November 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i11/p113302_s1.

Lin:2011:CST

[1034] Yuan Lin, Bin Wu, Zhongzhi Zhang, and Guanrong Chen. Counting spanning trees in self-similar networks by evaluating determinants. Journal of Mathematical Physics, 52(11):113303, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i11/
p113303_s1.

Fucci:2011:HZF

[1035] Guglielmo Fucci. On the Hurwitz zeta function of imaginary second argument. Journal of Mathematical Physics, 52(11):113501, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p113501_s1.

Cresson:2011:SHS

[1036] Jacky Cresson. The stochastisation hypothesis and the spacing of planetary systems. Journal of Mathematical Physics, 52(11):113502, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p113502_s1.

Regniers:2011:AMD

[1037] G. Regniers and J. Van der Jeugt. Angular momentum decomposition of the three-dimensional Wigner harmonic oscillator. Journal of Mathematical Physics, 52(11):113503, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i11/p113503_s1.

Schiebold:2011:SPN

[1038] C. Schiebold. Structural properties of the noncommutative KdV recursion operator. Journal of Mathematical Physics, 52(11):113504, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p113504_s1.

Mathur:2011:IPO

[1039] Manu Mathur, Indrakshi Raychowdhury, and T. P. Sreeraj. Invariants, projection operators and SU(N) × SU(N) irreducible Schwinger bosons. Journal of Mathematical Physics, 52(11):113505, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i11/p113505_s1.

Druta-Romaniuc:2011:MCC

[1040] Simona Luiza Druţă-Romaniuc and Marian Ioan Munteanu. Magnetic curves corresponding to Killing magnetic fields in E^{3}. Journal of Mathematical Physics, 52(11):113506, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i11/p113506_s1.

Yin:2011:RGT

[1041] Mei Yin. Renormalization group transformations near the critical point: Some rigorous results. Journal of Mathematical Physics, 52(11):113507, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i11/p113507_s1.

Bagarello:2011:WCR

[1042] Fabio Bagarello, Atsushi Inoue, and Camillo Trapani. Weak commutation relations of unbounded operators and applications. Journal of Mathematical Physics, 52(11):113508, November 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i11/p113508_s1.

Andrisani:2011:MPG

[1043] Andrea Andrisani and Nicola Cufaro Petroni. Markov processes and generalized Schrödinger equations. Journal of Mathematical Physics, 52(11):113509, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p113509_s1.

Zeier:2011:SPQ

[1044] Robert Zeier and Thomas Schulte-Herbrüggen. Symmetry principles in quantum systems theory. Journal of Mathematical Physics, 52 (11):113510, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p113510_s1. See erratum [2619].

Pawellek:2011:GJE

[1045] Michael Pawellek. On a generalization of Jacobi’s elliptic functions and the double sine–Gordon kink chain. Journal of Mathematical Physics, 52(11):113701, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i11/p113701_s1.

Kikuchi:2011:STP

[1046] Tetsuya Kikuchi. Symmetries in the third Painlevé equation arising from the modified Pohlmeyer–Lund–Regge hierarchy. Journal of Mathematical Physics, 52(11):113702, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i11/p113702_s1.

Angst:2011:TEC

[1047] Jürgen Angst. Trends to equilibrium for a class of relativistic diffusions. Journal of Mathematical Physics, 52(11):113703, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i11/
p113703_s1.

Wang:2011:GSN

[1048] Zhengping Wang and Huan-Song Zhou. Ground state for nonlinear Schrödinger equation with sign-changing and vanishing potential. Journal of Mathematical Physics, 52(11):113704, November 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i11/
p113704_s1.

Long:2011:SIB

[1049] Long Long and Shifang Zhang. On the supremum and infimum of bounded quantum observables. Journal of Mathematical Physics, 52 (12):122101, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122101_s1. See erratum [1129].

Haapasalo:2011:ECQ

[1050] Erkka Theodor Haapasalo and Juha-Pekka Pellonpää. Extreme covariant quantum observables in the case of an Abelian symmetry group and a transitive value space. Journal of Mathematical Physics, 52(12):122102, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i12/p122102_s1.

Capak:2011:AAS

[1051] M. Çapak and B. Gönül. An alternative approach to Schrödinger equations with a spatially varying mass. Journal of Mathematical Physics, 52(12):122103, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122103_s1.

Gianfreda:2011:SPW

[1052] M. Gianfreda and G. Landolfi. Spectral problems for the Weyl-ordered form of operators (1/p)^{n_{q}n}. Journal of Mathematical Physics, 52 (12):122104, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122104_s1.

Ciftja:2011:ERF

[1053] Orion Ciftja. Exact results for finite quantum Hall systems of electrons at filling factor one: Disk geometry. Journal of Mathematical Physics, 52(12):122105, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122105_s1.

Prolhac:2011:PAD

[1054] Sylvain Prolhac and Herbert Spohn. The propagator of the attractive delta-Bose gas in one dimension. Journal of Mathematical Physics, 52(12):122106, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122106_s1.

Ho:2011:PAS

[1055] C.-L. Ho. Prepotential approach to solvable rational extensions of Harmonic Oscillator and Morse potentials. Journal of Mathematical Physics, 52(12):122107, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122107_s1.

Ikhdair:2011:RNB

[1056] Sameer M. Ikhdair and Ramazan Sever. Relativistic and nonrelativistic bound states of the isotonic oscillator by Nikiforov–Uvarov method. Journal of Mathematical Physics, 52(12):122108, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i12/p122108_s1.

Kiem:2011:EPV

[1057] Young-Hoon Kiem, Seung-Hyeok Kye, and Jungseob Lee. Existence of product vectors and their partial conjugates in a pair of spaces. Journal of Mathematical Physics, 52(12):122201, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p122201_s1.

Skowronek:2011:TTB

[1058] Łukasz Skowronek. Three-by-three bound entanglement with general unextendible product bases. Journal of Mathematical Physics, 52 (12):122202, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122202_s1.

Chen:2011:DRF

[1059] Lin Chen and Dragomir Ž. – Doković. Description of rank four entangled states of two qutrits having positive partial transpose. Journal of Mathematical Physics, 52(12):122203, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p122203_s1. See

erratum [1375].

Farenick:2011:CNR

[1060] Douglas Farenick, Sarah Plosker, and Jerrod Smith. Classical and nonclassical randomness in quantum measurements. Journal of Mathematical Physics, 52(12):122204, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v52/i12/p122204_s1.

Miao:2011:GPD

[1061] S. P. Miao, N. C. Tsamis, and R. P. Woodard. The graviton propagator in de Donder gauge on de Sitter background. Journal of Mathematical Physics, 52(12):122301, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i12/p122301_s1.

Kim:2011:LCS

[1062] Do-Hyung Kim. Lorentzian covering space and homotopy classes. Journal of Mathematical Physics, 52(12):122501, December 2011. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i12/

p122501_s1. See erratum [2626].

Werner:2011:DDM

[1063] Ivan Werner. Dynamically defined measures and equilibrium states. Journal of Mathematical Physics, 52(12):122701, December 2011. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i12/

p122701_s1. See erratum [1374].

Mahdi:2011:IAS

[1064] Adam Mahdi and Claudia Valls. Invariant algebraic surfaces for the reduced three-wave interaction system. Journal of Mathematical Physics, 52(12):122702, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122702_s1.

Gibbons:2011:GCK

[1065] G. W. Gibbons and C. Rugina. Goryachev–Chaplygin, Kovalevskaya, and Brdička–Eardley–Nappi–Witten pp-waves spacetimes with higher rank Stäckel–Killing tensors. Journal of Mathematical Physics, 52 (12):122901, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p122901_s1.

Hasegawa:2011:SHA

[1066] Hideo Hasegawa. Specific heat anomalies of small quantum systems subjected to finite baths. Journal of Mathematical Physics, 52(12):123301, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i12/p123301_s1.

Konenberg:2011:EGS

[1067] Martin Könenberg, Oliver Matte, and Edgardo Stockmeyer. Existence of ground states of hydrogen-like atoms in relativistic quantum electrodynamics. II. The no-pair operator. Journal of Mathematical Physics, 52(12):123501, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123501_s1.

Arnlind:2011:CLA

[1068] Joakim Arnlind, Abdenacer Makhlouf, and Sergei Silvestrov. Construction of n-Lie algebras and n-ary Hom–Nambu–Lie algebras. Journal of Mathematical Physics, 52(12):123502, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p123502_s1.

Martin:2011:SLA

[1069] Antonio J. Calderón Martín and M. Forero Piulestán. Split 3-Lie algebras. Journal of Mathematical Physics, 52(12):123503, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i12/
p123503_s1.

Louko:2011:CRR

[1070] Jorma Louko and Eric Martínez-Pascual. Constraint rescaling in refined algebraic quantisation: Momentum constraint. Journal of Mathematical Physics, 52(12):123504, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i12/p123504_s1.

Groeger:2011:HSS

[1071] Josua Groeger. Holomorphic supercurves and supersymmetric sigma models. Journal of Mathematical Physics, 52(12):123505, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i12/p123505_s1.

Ramm:2011:USI

[1072] A. G. Ramm. Uniqueness of the solution to inverse scattering problem with scattering data at a fixed direction of the incident wave. Journal of Mathematical Physics, 52(12):123506, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p123506_s1.

Dietermann:2011:CBF

[1073] F. Dietermann, R. Singer, and M. Fähnle. Construction of basis functions with crystal symmetry for the spin-cluster expansion of the magnetic energy on the atomic scale. Journal of Mathematical Physics, 52(12):123507, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123507_s1.

Buhagiar:2011:QSS

[1074] D. Buhagiar, E. Chetcuti, and A. Dvurečenskij. Quasi-splitting subspaces and Foulis–Randall subspaces. Journal of Mathematical Physics, 52(12):123508, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123508_s1.

Harju:2011:CDO

[1075] Antti J. Harju. Covariant Dirac operators on quantum groups. Journal of Mathematical Physics, 52(12):123509, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p123509_s1.

You:2011:NSI

[1076] Fucai You. Nonlinear super integrable Hamiltonian couplings. Journal of Mathematical Physics, 52(12):123510, December 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v52/i12/p123510_s1.

Sadel:2011:RBT

[1077] Christian Sadel. Relations between transfer and scattering matrices in the presence of hyperbolic channels. Journal of Mathematical Physics, 52(12):123511, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123511_s1.

Kien:2011:RQS

[1078] Nguyen Cong Kien, Nguyen Anh Ky, Le Ba Nam, and Nguyen Thi Hong Van. Representations of quantum superalgebra U_{q}[gl(2|1)] in a coherent state basis and generalization. Journal of Mathematical Physics, 52(12):123512, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123512_s1.

Borisov:2011:PWT

[1079] Denis Borisov and Giuseppe Cardone. Planar waveguide with “twisted” boundary conditions: Discrete spectrum. Journal of Mathematical Physics, 52(12):123513, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i12/p123513_s1.

Wang:2011:NAMb

[1080] Yushun Wang, Juan Jiang, and Wenjun Cai. Numerical analysis of a multi-symplectic scheme for the time-domain Maxwell’s equations. Journal of Mathematical Physics, 52(12):123701, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/i12/
p123701_s1.

Matsuno:2011:NMC

[1081] Yoshimasa Matsuno. A novel multi-component generalization of the short pulse equation and its multisoliton solutions. Journal of Mathematical Physics, 52(12):123702, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v52/i12/p123702_s1.

Ben-Artzi:2011:INN

[1082] Jonathan Ben-Artzi. Instability of nonsymmetric nonmonotone equilibria of the Vlasov–Maxwell system. Journal of Mathematical Physics, 52(12):123703, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p123703_s1.

Haas:2011:CDS

[1083] F. Haas. Comment on “Dynamical systems and Poisson structures” [J. Math. Phys. 50, 112703 (2009)]. Journal of Mathematical Physics, 52(12):124101, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p124101_s1. See [18].

Freiji:2012:GES

[1084] Abraham Freiji, Christian Hainzl, and Robert Seiringer. The gap equation for spin-polarized fermions. Journal of Mathematical Physics, 53(1): 012101, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p012101_s1.

Unal:2012:QCS

[1085] Nuri Ünal. Quasi-coherent states for harmonic oscillator with timedependent parameters. Journal of Mathematical Physics, 53(1):012102, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p012102_s1.

Mouayn:2012:PCS

[1086] Zouhaïr Mouayn. Phase coherent states with circular Jacobi polynomials for the pseudoharmonic oscillator. Journal of Mathematical Physics, 53(1):012103, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p012103_s1.

Boonserm:2012:CTM

[1087] Petarpa Boonserm and Matt Visser. Compound transfer matrices: Constructive and destructive interference. Journal of Mathematical Physics, 53(1):012104, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i1/p012104_s1.

Sun:2012:GPS

[1088] Qinxiu Sun. Generalization of H-pseudoalgebraic structures. Journal of Mathematical Physics, 53(1):012105, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i1/p012105_s1.

Gardas:2012:NRO

[1089] Bartlomiej Gardas and Zbigniew Puchala. Notes on the Riccati operator equation in open quantum systems. Journal of Mathematical Physics, 53(1):012106, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p012106_s1.

Anastopoulos:2012:CAD

[1090] Charis Anastopoulos and Ntina Savvidou. Coherences of accelerated detectors and the local character of the Unruh effect. Journal of Mathematical Physics, 53(1):012107, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i1/p012107_s1.

Ueberschar:2012:TFP

[1091] Henrik Ueberschär. The trace formula for a point scatterer on a compact hyperbolic surface. Journal of Mathematical Physics, 53(1):012108, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p012108_s1.

Azizov:2012:PSH

[1092] Tomas Ya. Azizov and Carsten Trunk. PT symmetric, Hermitian and P-self-adjoint operators related to potentials in PT quantum mechanics. Journal of Mathematical Physics, 53(1):012109, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i1/p012109_s1.

Jencova:2012:GCC

[1093] Anna Jencová. Generalized channels: Channels for convex subsets of the state space. Journal of Mathematical Physics, 53(1):012201, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p012201_s1.

Niekamp:2012:EUR

[1094] Sönke Niekamp, Matthias Kleinmann, and Otfried Gühne. Entropic uncertainty relations and the stabilizer formalism. Journal of Mathematical Physics, 53(1):012202, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p012202_s1.

Chen:2012:SRA

[1095] Fa-Min Chen. Superalgebra realization of the 3-algebras in N = 6, 8 Chern–Simons-matter theories. Journal of Mathematical Physics, 53(1): 012301, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p012301_s1.

Blumenhagen:2012:CLB

[1096] Ralph Blumenhagen, Benjamin Jurke, Thorsten Rahn, and Helmut Roschy. Cohomology of line bundles: Applications. Journal of Mathematical Physics, 53(1):012302, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i1/p012302_s1.

Reiris:2012:SSP

[1097] Martin Reiris. Static solutions from the point of view of comparison geometry. Journal of Mathematical Physics, 53(1):012501, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i1/
p012501_s1.

Akerblom:2012:REM

[1098] Nikolas Akerblom and Gunther Cornelissen. Relative entropy as a measure of inhomogeneity in general relativity. Journal of Mathematical Physics, 53(1):012502, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p012502_s1.

Livine:2012:SRL

[1099] Etera Livine and Johannes Tambornino. Spinor representation for loop quantum gravity. Journal of Mathematical Physics, 53(1):012503, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p012503_s1.

Enolski:2012:IGH

[1100] Victor Enolski, Betti Hartmann, Valeria Kagramanova, Jutta Kunz, Claus Lämmerzahl, and Parinya Sirimachan. Inversion of a general hyperelliptic integral and particle motion in Horava–Lifshitz black hole space–times. Journal of Mathematical Physics, 53(1):012504, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p012504_s1.

Ma:2012:GAW

[1101] Shan Ma and Hongtao Li. Global attractors for weighted p-Laplacian equations with boundary degeneracy. Journal of Mathematical Physics, 53(1):012701, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i1/p012701_s1.

Llibre:2012:PON

[1102] Jaume Llibre and Claudio Vidal. Periodic orbits and non-integrability in a cosmological scalar field. Journal of Mathematical Physics, 53(1): 012702, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p012702_s1.

Marquette:2012:CLO

[1103] Ian Marquette. Classical ladder operators, polynomial Poisson algebras, and classification of superintegrable systems. Journal of Mathematical Physics, 53(1):012901, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p012901_s1.

Fenyes:2012:LCC

[1104] Aaron Fenyes. Limitations on cloning in classical mechanics. Journal of Mathematical Physics, 53(1):012902, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i1/p012902_s1.

Lv:2012:NUD

[1105] Guangying Lv and Mingxin Wang. Non-uniform dependence for a modified Camassa–Holm system. Journal of Mathematical Physics, 53(1): 013101, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p013101_s1.

Parsley:2012:BSO

[1106] Jason Parsley. The Biot–Savart operator and electrodynamics on subdomains of the three-sphere. Journal of Mathematical Physics, 53(1): 013102, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p013102_s1.

Shashikanth:2012:VD

[1107] Banavara N. Shashikanth. Vortex dynamics in R^{4}. Journal of Mathematical Physics, 53(1):013103, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i1/p013103_s1.

Dong:2012:CLA

[1108] Zhishan Dong, Tiefeng Jiang, and Danning Li. Circular law and arc law for truncation of random unitary matrix. Journal of Mathematical Physics, 53(1):013301, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013301_s1.

Borovyk:2012:DEH

[1109] Vita Borovyk and Robert Sims. Dispersive estimates for harmonic oscillator systems. Journal of Mathematical Physics, 53(1):013302, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p013302_s1.

Spengler:2012:CPH

[1110] Christoph Spengler, Marcus Huber, and Beatrix C. Hiesmayr. Composite parameterization and Haar measure for all unitary and special unitary groups. Journal of Mathematical Physics, 53(1):013501, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p013501_s1.

Yu:2012:INF

[1111] Xin Yu, Yi-Tian Gao, Zhi-Yuan Sun, and Ying Liu. Investigation on a nonisospectral fifth-order Korteweg–de Vries equation generalized from fluids. Journal of Mathematical Physics, 53(1):013502, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i1/
p013502_s1.

Fan:2012:SEB

[1112] Engui Fan and Y. C. Hon. Super extension of Bell polynomials with applications to supersymmetric equations. Journal of Mathematical Physics, 53(1):013503, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013503_s1.

Baloitcha:2012:UDO

[1113] E. Baloitcha, M. N. Hounkonnou, and E. B. Ngompe Nkouankam. Unified (p, q; α, β, ν; γ)-deformed oscillator algebra: Irreducible representations and induced deformed harmonic oscillator. Journal of Mathematical Physics, 53(1):013504, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i1/p013504_s1.

Zhang:2012:QFV

[1114] Zhiyue Zhang and Fuqiang Lu. Quadratic finite volume element method for the improved Boussinesq equation. Journal of Mathematical Physics, 53(1):013505, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i1/p013505_s1.

Radha:2012:IWD

[1115] R. Radha and V. D. Sharma. Interaction of a weak discontinuity with elementary waves of Riemann problem. Journal of Mathematical Physics, 53(1):013506, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013506_s1.

Zakrzewski:2012:LDE

[1116] Michal Zakrzewski and Henryk Żoła̧dek. Linear differential equations and multiple zeta-values. III. Zeta(3). Journal of Mathematical Physics, 53(1):013507, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i1/p013507_s1.

Grosse:2012:QFT

[1117] H. Grosse, P. Presnajder, and Zhituo Wang. Quantum field theory on quantized Bergman domain. Journal of Mathematical Physics, 53(1): 013508, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p013508_s1.

Gallatin:2012:FGF

[1118] Gregg M. Gallatin. Fourier, Gauss, Fraunhofer, Porod and the shape from moments problem. Journal of Mathematical Physics, 53(1):013509, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p013509_s1.

Misra:2012:SCA

[1119] Kailash C. Misra, Masato Okado, and Evan A. Wilson. Soliton cellular automaton associated with G_{2}(1) crystal base. Journal of Mathematical Physics, 53(1):013510, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013510_s1.

Anderson:2012:NST

[1120] I. M. Anderson and C. G. Torre. New symbolic tools for differential geometry, gravitation, and field theory. Journal of Mathematical Physics, 53(1):013511, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013511_s1.

Wang:2012:GEA

[1121] Yu-Zhu Wang and Yin-Xia Wang. Global existence and asymptotic behavior of solutions to a nonlinear wave equation of fourth-order. Journal of Mathematical Physics, 53(1):013512, January 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i1/p013512_s1.

deLange:2012:TMM

[1122] O. L. de Lange, R. E. Raab, and A. Welter. On the transition from microscopic to macroscopic electrodynamics. Journal of Mathematical Physics, 53(1):013513, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013513_s1.

Wu:2012:FBS

[1123] Dongsheng Wu. Fractional Brownian sheets run with nonlinear clocks. Journal of Mathematical Physics, 53(1):013514, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i1/p013514_s1.

Kojima:2012:FFR

[1124] Takeo Kojima. Free field realization of quantum affine superalgebra U_{q}(ŝl(N|1)). Journal of Mathematical Physics, 53(1):013515, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p013515_s1.

Bodner:2012:ARG

[1125] M. Bodner, J. Patera, and M. Peterson. Affine reflection groups for tiling applications: Knot theory and DNA. Journal of Mathematical Physics, 53(1):013516, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i1/p013516_s1.

Li:2012:BTS

[1126] Chuanzhong Li, Jingsong He, and Yucai Su. Block type symmetry of bigraded Toda hierarchy. Journal of Mathematical Physics, 53(1): 013517, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p013517_s1.

Tian:2012:LWP

[1127] Lixin Tian, Wanfeng Yan, and Guilong Gui. On the local well posedness and blow-up solution of a coupled Camassa–Holm equations in Besov spaces. Journal of Mathematical Physics, 53(1):013701, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i1/p013701_s1.

Smolyakov:2012:ENE

[1128] Mikhail N. Smolyakov. Erratum: “Non-existence of non-topological solitons in some types of gauge field theories in Minkowski space” [J. Math. Phys. 52, 102901 (2011)]. Journal of Mathematical Physics, 53(1): 019901, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i1/p019901_s1. See [994].

Long:2012:ESI

[1129] Long Long and Shifang Zhang. Erratum: “On the supremum and infimum of bounded quantum observables” [J. Math. Phys. 52, 122101 (2011)]. Journal of Mathematical Physics, 53(1):019902, January 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i1/

p019902_s1. See [1049].

Stasinska:2012:CES

[1130] Julia Stasińska. Convergence of expansions in Schrödinger and Dirac eigenfunctions, with an application to the R-matrix theory. Journal of Mathematical Physics, 53(2):022101, February 2012. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i2/p022101_s1.

DeLeo:2012:DTQ

[1131] Stefano De Leo and Gisele Ducati. Delay time in quaternionic quantum mechanics. Journal of Mathematical Physics, 53(2):022102, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p022102_s1.

Marquette:2012:GFD

[1132] Ian Marquette. Generalized five-dimensional Kepler system, Yang– Coulomb monopole, and Hurwitz transformation. Journal of Mathematical Physics, 53(2):022103, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i2/p022103_s1.

Hassanabadi:2012:DEG

[1133] H. Hassanabadi, E. Maghsoodi, S. Zarrinkamar, and H. Rahimov. Dirac equation for generalized Pöschl–Teller scalar and vector potentials and a Coulomb tensor interaction by Nikiforov–Uvarov method. Journal of Mathematical Physics, 53(2):022104, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i2/p022104_s1.

Luhrmann:2012:MFQ

[1134] Jonas Lührmann. Mean-field quantum dynamics with magnetic fields. Journal of Mathematical Physics, 53(2):022105, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/p022105_s1.

Loubenets:2012:LQH

[1135] Elena R. Loubenets. Local quasi hidden variable modelling and violations of Bell-type inequalities by a multipartite quantum state. Journal of Mathematical Physics, 53(2):022201, February 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i2/p022201_s1. See

note [1251].

Chen:2012:RRL

[1136] Jianxin Chen, Zhengfeng Ji, Alexander Klyachko, David W. Kribs, and Bei Zeng. Rank reduction for the local consistency problem. Journal of Mathematical Physics, 53(2):022202, February 2012. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i2/p022202_s1.

Shi:2012:SSF

[1137] Chang-Guang Shi and Minoru Hirayama. Solitonic solutions of Faddeev model. Journal of Mathematical Physics, 53(2):022301, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p022301_s1.

Saidi:2012:FSL

[1138] El Hassan Saidi. On flavor symmetry in lattice quantum chromodynamics. Journal of Mathematical Physics, 53(2):022302, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/
p022302_s1.

Maeda:2012:NDS

[1139] Yoshiaki Maeda and Akifumi Sako. Noncommutative deformation of spinor zero mode and Atiyah–Drinfeld–Hitchin–Manin construction. Journal of Mathematical Physics, 53(2):022303, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/p022303_s1.

Kahya:2012:CLG

[1140] E. O. Kahya, S. P. Miao, and R. P. Woodard. The coincidence limit of the graviton propagator in de donder gauge on de Sitter background. Journal of Mathematical Physics, 53(2):022304, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/p022304_s1.

Kopper:2012:RPI

[1141] Christoph Kopper and Benjamin Lévêque. Regularized path integrals and anomalies: U(1) chiral gauge theory. Journal of Mathematical Physics, 53(2):022305, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p022305_s1.

Fairbairn:2012:QDT

[1142] Winston J. Fairbairn and Catherine Meusburger. Quantum deformation of two four-dimensional spin foam models. Journal of Mathematical Physics, 53(2):022501, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i2/p022501_s1.

Liang:2012:SHN

[1143] Zhuobin Liang and Xiao Zhang. Spacelike hypersurfaces with negative total energy in de Sitter spacetime. Journal of Mathematical Physics, 53(2):022502, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p022502_s1.

Natorf:2012:ASV

[1144] Włodzimierz Natorf. Asymptotic stability of vacuum twisting type II metrics. Journal of Mathematical Physics, 53(2):022503, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/
p022503_s1.

Hamed:2012:HGI

[1145] Bassem Ben Hamed, Lubomir Gavrilov, and Martine Klughertz. The holonomy group at infinity of the Painlevé VI equation. Journal of Mathematical Physics, 53(2):022701, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i2/p022701_s1.

Su:2012:TDI

[1146] Xifeng Su and Lin Wang. Total destruction of invariant tori for the generalized Frenkel–Kontorova model. Journal of Mathematical Physics, 53(2):022702, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p022702_s1.

Yang:2012:DEM

[1147] Yisong Yang, Ruifeng Zhang, and Le Zhao. Dynamics of electrostatic microelectromechanical systems actuators. Journal of Mathematical Physics, 53(2):022703, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p022703_s1.

Geloun:2012:CGF

[1148] Joseph Ben Geloun. Classical group field theory. Journal of Mathematical Physics, 53(2):022901, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p022901_s1.

Aizenman:2012:PRQ

[1149] Michael Aizenman, Rafael L. Greenblatt, and Joel L. Lebowitz. Proof of rounding by quenched disorder of first order transitions in lowdimensional quantum systems. Journal of Mathematical Physics, 53 (2):023301, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023301_s1.

Rodriguez:2012:DSI

[1150] A. Rodríguez and C. Tsallis. A dimension scale-invariant probabilistic model based on Leibniz-like pyramids. Journal of Mathematical Physics, 53(2):023302, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023302_s1. See note [1217].

Veneziani:2012:AIK

[1151] Alexei M. Veneziani, Tiago Pereira, and Domingos H. U. Marchetti. Asymptotic integral kernel for ensembles of random normal matrices with radial potentials. Journal of Mathematical Physics, 53(2):023303, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023303_s1.

Paoletti:2012:LTF

[1152] Roberto Paoletti. Local trace formulae and scaling asymptotics for general quantized Hamiltonian flows. Journal of Mathematical Physics, 53(2):023501, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023501_s1.

Skrypnyk:2012:QPF

[1153] T. Skrypnyk. Quasi-periodic functions on the torus and sl(n)-elliptic Lie algebra. Journal of Mathematical Physics, 53(2):023502, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023502_s1.

Borisov:2012:PWT

[1154] Denis Borisov and Giuseppe Cardone. Planar waveguide with “twisted” boundary conditions: Small width. Journal of Mathematical Physics, 53(2):023503, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023503_s1.

Matsuda:2012:RSN

[1155] Kazuhide Matsuda. Rational solutions of the Noumi and Yamada system of type A_{4}^{(1)}. Journal of Mathematical Physics, 53(2):023504, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023504_s1.

Llibre:2012:NFI

[1156] Jaume Llibre and Clàudia Valls. A note on the first integrals of the ABC system. Journal of Mathematical Physics, 53(2):023505, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023505_s1.

Grammaticos:2012:DPE

[1157] B. Grammaticos, A. Ramani, J. Satsuma, and R. Willox. Discretising the Painlevé equations à la Hirota–Mickens. Journal of Mathematical Physics, 53(2):023506, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023506_s1.

Xu:2012:SMJ

[1158] Tian Zhou Xu. Stability of multi-Jensen mappings in non-Archimedean normed spaces. Journal of Mathematical Physics, 53(2):023507, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023507_s1.

Friot:2012:CSR

[1159] Samuel Friot and David Greynat. On convergent series representations of Mellin — Barnes integrals. Journal of Mathematical Physics, 53(2):023508, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023508_s1.

Kang:2012:SGF

[1160] Jing Kang and Changzheng Qu. Symmetry groups and fundamental solutions for systems of parabolic equations. Journal of Mathematical Physics, 53(2):023509, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023509_s1.

Mimura:2012:DUS

[1161] N. Mimura, S. Isojima, M. Murata, J. Satsuma, A. Ramani, and B. Grammaticos. Do ultradiscrete systems with parity variables satisfy the singularity confinement criterion? Journal of Mathematical Physics, 53(2):023510, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023510_s1.

Hazra:2012:ESC

[1162] Tamaghna Hazra, V. K. Chandrasekar, R. Gladwin Pradeep, and M. Lakshmanan. Exact solutions of coupled Liénard-type nonlinear systems using factorization technique. Journal of Mathematical Physics, 53(2):023511, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023511_s1.

Bruzon:2012:NSR

[1163] M. S. Bruzon, M. L. Gandarias, and M. Senthilvelan. Nonlocal symmetries of Riccati and Abel chains and their similarity reductions. Journal of Mathematical Physics, 53(2):023512, February 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i2/p023512_s1.

Bejancu:2012:CSR

[1164] Aurel Bejancu. Curvature in sub-Riemannian geometry. Journal of Mathematical Physics, 53(2):023513, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i2/p023513_s1.

Bakke:2012:ACS

[1165] K. Bakke and C. Furtado. On the Aharonov–Casher system and the Landau–Aharonov–Casher system confined to a two-dimensional quantum ring. Journal of Mathematical Physics, 53(2):023514, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023514_s1.

Majewski:2012:SPM

[1166] Wladyslaw A. Majewski. On the structure of positive maps: Finitedimensional case. Journal of Mathematical Physics, 53(2):023515, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023515_s1.

Takane:2012:ABF

[1167] Martha Takane and Federico Zertuche. mathbbz_{2}-algebras in the Boolean function irreducible decomposition. Journal of Mathematical Physics, 53(2):023516, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p023516_s1.

Feng:2012:DBL

[1168] Youling Feng, Yang Cao, and Haijun Wang. Diagonalization of bounded linear operators on separable quaternionic Hilbert space. Journal of Mathematical Physics, 53(2):023517, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i2/p023517_s1.

Liu:2012:DTQ

[1169] Honghu Liu, Taylan Sengul, and Shouhong Wang. Dynamic transitions for quasilinear systems and Cahn–Hilliard equation with Onsager mobility. Journal of Mathematical Physics, 53(2):023518, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i2/
p023518_s1.

Xu:2012:FLI

[1170] Xi-Xiang Xu. A family of Liouville integrable lattice equations with a parameter and its two symmetry constraints. Journal of Mathematical Physics, 53(2):023519, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023519_s1.

Jang:2012:TDE

[1171] Juhi Jang. The two-dimensional Euler–Poisson system with spherical symmetry. Journal of Mathematical Physics, 53(2):023701, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i2/p023701_s1.

He:2012:ECG

[1172] Xiaoming He and Wenming Zou. Existence and concentration of ground states for Schrödinger–Poisson equations with critical growth. Journal of Mathematical Physics, 53(2):023702, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i2/p023702_s1.

Kycia:2012:SES

[1173] Radoslaw A. Kycia. On similarity in the evolution of semilinear wave and Klein–Gordon equations: Numerical surveys. Journal of Mathematical Physics, 53(2):023703, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i2/p023703_s1.

Fernandez:2012:VHP

[1174] Francisco M. Fernández. On the variational homotopy perturbation method for nonlinear oscillators. Journal of Mathematical Physics, 53(2):024101, February 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i2/p024101_s1.

Doring:2012:CQP

[1175] Andreas Döring and Chris J. Isham. Classical and quantum probabilities as truth values. Journal of Mathematical Physics, 53(3):032101, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p032101_s1.

Cao:2012:EHB

[1176] Zhenwei Cao and Alexander Elgart. On the efficiency of Hamiltonianbased quantum computation for low-rank matrices. Journal of Mathematical Physics, 53(3):032201, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i3/p032201_s1.

Gutoski:2012:MDQ

[1177] Gus Gutoski. On a measure of distance for quantum strategies. Journal of Mathematical Physics, 53(3):032202, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p032202_s1.

Collins:2012:TSM

[1178] Benoît Collins, Motohisa Fukuda, and Ion Nechita. Towards a state minimizing the output entropy of a tensor product of random quantum channels. Journal of Mathematical Physics, 53(3):032203, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p032203_s1.

Kantor:2012:MIA

[1179] W. M. Kantor. MUBs inequivalence and affine planes. Journal of Mathematical Physics, 53(3):032204, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i3/p032204_s1.

Gover:2012:HSC

[1180] A. Rod Gover and Josef Silhan. Higher symmetries of the conformal powers of the Laplacian on conformally flat manifolds. Journal of Mathematical Physics, 53(3):032301, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i3/p032301_s1.

Canarutto:2012:PSG

[1181] Daniel Canarutto. Positive spaces, generalized semi-densities, and quantum interactions. Journal of Mathematical Physics, 53(3):032302, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p032302_s1.

Turyshev:2012:ARR

[1182] Slava G. Turyshev, Olivier L. Minazzoli, and Viktor T. Toth. Accelerating relativistic reference frames in Minkowski space–time. Journal of Mathematical Physics, 53(3):032501, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p032501_s1.

Dupuis:2012:HLS

[1183] Maïté Dupuis, Laurent Freidel, Etera R. Livine, and Simone Speziale. Holomorphic Lorentzian simplicity constraints. Journal of Mathematical Physics, 53(3):032502, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p032502_s1.

Gair:2012:REM

[1184] Jonathan Gair, Nicolás Yunes, and Carl M. Bender. Resonances in extreme mass-ratio inspirals: Asymptotic and hyperasymptotic analysis. Journal of Mathematical Physics, 53(3):032503, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p032503_s1.

Baird:2012:RCG

[1185] Paul Baird and Mohammad Wehbe. Ray congruences that generate conformal foliations. Journal of Mathematical Physics, 53(3):032504, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p032504_s1.

Santi:2012:SPA

[1186] A. Santi and A. Spiro. Super–Poincarè algebras, space–times, and supergravities. II. Journal of Mathematical Physics, 53(3):032505, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p032505_s1.

Thirukkanesh:2012:SRC

[1187] S. Thirukkanesh, S. S. Rajah, and S. D. Maharaj. Shearing radiative collapse with expansion and acceleration. Journal of Mathematical Physics, 53(3):032506, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p032506_s1.

Kim:2012:CCA

[1188] Do-Hyung Kim. A characterization of causal automorphisms by wave equations. Journal of Mathematical Physics, 53(3):032507, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/

p032507_s1. See erratum [1932].

Baake:2012:STP

[1189] Michael Baake, Franz Gähler, and Uwe Grimm. Spectral and topological properties of a family of generalised Thue–Morse sequences. Journal of Mathematical Physics, 53(3):032701, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p032701_s1.

Zhijian:2012:FDA

[1190] Yang Zhijian. Finite-dimensional attractors for the Kirchhoff models with critical exponents. Journal of Mathematical Physics, 53(3): 032702, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p032702_s1.

Prieto-Martinez:2012:UFH

[1191] Pedro Daniel Prieto-Martínez and Narciso Román-Roy. Unified formalism for higher order non-autonomous dynamical systems. Journal of Mathematical Physics, 53(3):032901, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p032901_s1.

Aharonovich:2012:RRC

[1192] I. Aharonovich and L. P. Horwitz. Radiation-reaction in classical off-shell electrodynamics. I. The above mass-shell case. Journal of Mathematical Physics, 53(3):032902, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p032902_s1.

DeMatteis:2012:LPS

[1193] Giovanni De Matteis and Luigi Martina. Lie point symmetries and reductions of one-dimensional equations describing perfect Korteweg-type nematic fluids. Journal of Mathematical Physics, 53(3):033101, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p033101_s1.

Strong:2012:GLI

[1194] Scott A. Strong and Lincoln D. Carr. Generalized local induction equation, elliptic asymptotics, and simulating superfluid turbulence. Journal of Mathematical Physics, 53(3):033102, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p033102_s1.

Christiansen:2012:SGT

[1195] Snorre H. Christiansen and Tore G. Halvorsen. A simplicial gauge theory. Journal of Mathematical Physics, 53(3):033501, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033501_s1.

Ji:2012:DOF

[1196] Un Cig Ji and Jae Hee Kim. φ-derivations on operators on q-Fock space. Journal of Mathematical Physics, 53(3):033502, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033502_s1.

Kondej:2012:FSO

[1197] Sylwia Kondej and Jayme Vaz, Jr. Fractional Schrödinger operator with delta potential localized on circle. Journal of Mathematical Physics, 53(3):033503, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p033503_s1.

Felinska:2012:WPV

[1198] Ewa Felińska, Zbigniew Jaskólski, and Michal Kosztolowicz. Whittaker pairs for the Virasoro algebra and the Gaiotto–Bonelli–Maruyoshi– Tanzini states. Journal of Mathematical Physics, 53(3):033504, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v53/i3/p033504_s1. See erratum [1612].

Jaradat:2012:MEE

[1199] E. K. Jaradat, R. S. Hijjawi, and J. M. Khalifeh. Maxwell’s equations and electromagnetic Lagrangian density in fractional form. Journal of Mathematical Physics, 53(3):033505, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i3/p033505_s1. See comment

[2253].

Ren:2012:HSF

[1200] Li Ren, Qiang Mu, and Yongzheng Zhang. Hamiltonian superalgebras of formal vector fields. Journal of Mathematical Physics, 53(3): 033506, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p033506_s1.

Zhou:2012:SIS

[1201] Tong Zhou and Zuo nong Zhu. Soliton interactions in some semidiscrete integrable systems. Journal of Mathematical Physics, 53(3): 033507, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p033507_s1.

Xia:2012:QPS

[1202] Baoqiang Xia. The quasi-periodic solutions of mixed KdV equations. Journal of Mathematical Physics, 53(3):033508, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033508_s1.

Vaisman:2012:GDF

[1203] Izu Vaisman. On the geometry of double field theory. Journal of Mathematical Physics, 53(3):033509, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i3/p033509_s1.

Yu:2012:TKN

[1204] Jing Yu, Jingsong He, and Jingwei Han. Two kinds of new integrable decompositions of the Gerdjikov–Ivanov equation. Journal of Mathematical Physics, 53(3):033510, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p033510_s1.

Doma:2012:CSR

[1205] S. B. Doma and M. M. Hammad. Classification, selection rules, and symmetry properties of the Clebsch–Gordan coefficients of symmetric group. Journal of Mathematical Physics, 53(3):033511, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033511_s1.

Arponen:2012:ISB

[1206] Heikki Arponen. Infinite symmetry on the boundary of SL (3)/ SO (3). Journal of Mathematical Physics, 53(3):033512, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033512_s1.

Kaikina:2012:MIB

[1207] Elena I. Kaikina. Mixed initial-boundary value problem for intermediate long-wave equation. Journal of Mathematical Physics, 53(3): 033701, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p033701_s1.

Cao:2012:SWN

[1208] Pei Cao, Jing Wang, and Wenming Zou. On the standing waves for nonlinear Hartree equation with confining potential. Journal of Mathematical Physics, 53(3):033702, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i3/p033702_s1.

Frisch:2012:OTO

[1209] Uriel Frisch, Olga Podvigina, Barbara Villone, and Vladislav Zheligovsky. Optimal transport by omni-potential flow and cosmological reconstruction. Journal of Mathematical Physics, 53(3):033703, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i3/p033703_s1.

Lamothe:2012:SGA

[1210] Vincent Lamothe. Symmetry group analysis of an ideal plastic flow. Journal of Mathematical Physics, 53(3):033704, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033704_s1.

Ganchev:2012:ITM

[1211] Georgi Ganchev and Velichka Milousheva. An invariant theory of marginally trapped surfaces in the four-dimensional Minkowski space. Journal of Mathematical Physics, 53(3):033705, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033705_s1.

Fang:2012:GCL

[1212] Daoyuan Fang, Ruizhao Zi, and Ting Zhang. Global classical large solutions to a 1D fluid-particle interaction model: The bubbling regime. Journal of Mathematical Physics, 53(3):033706, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i3/p033706_s1.

Kweyama:2012:FOD

[1213] M. C. Kweyama, K. S. Govinder, and S. D. Maharaj. A fifth order differential equation for charged perfect fluids. Journal of Mathematical Physics, 53(3):033707, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p033707_s1.

Dong:2012:NLP

[1214] Huanhe Dong, Jie Su, Fangjiao Yi, and Tongqian Zhang. New Lax pairs of the Toda lattice and the nonlinearization under a higherorder Bargmann constraint. Journal of Mathematical Physics, 53(3):

033708, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p033708_s1.

Melgaard:2012:MSQ

[1215] M. Melgaard and F. Zongo. Multiple solutions of the quasirelativistic Choquard equation. Journal of Mathematical Physics, 53(3):033709, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p033709_s1.

Wu:2012:GWP

[1216] Wei Wu, Shangbin Cui, and Jinqiao Duan. Global well-posedness of a stochastic coupled Kuramoto–Sivashinsky and Ginzburg–Landautype model for the Marangoni convection. Journal of Mathematical Physics, 53(3):033710, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i3/p033710_s1.

Rodriguez:2012:PND

[1217] A. Rodríguez and C. Tsallis. Publisher’s note: “A dimension scaleinvariant probabilistic model based on Leibniz-like pyramids” [J. Math. Phys. 53, 023302 (2012)]. Journal of Mathematical Physics, 53(3): 039901, March 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i3/p039901_s1. See [1150].

Agboola:2012:UDE

[1218] Davids Agboola and Yao-Zhong Zhang. Unified derivation of exact solutions for a class of quasi-exactly solvable models. Journal of Mathematical Physics, 53(4):042101, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p042101_s1.

Kauffman:2012:QML

[1219] Christopher J. Kauffman, Robert M. Kesler, Amanda G. Parshall, Evelyn A. Stamey, and Benjamin A. Steinhurst. Quantum mechanics on Laakso spaces. Journal of Mathematical Physics, 53(4):042102, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p042102_s1.

Smilga:2012:NIF

[1220] A. V. Smilga. Non-integer flux: Why it does not work. Journal of Mathematical Physics, 53(4):042103, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i4/p042103_s1.

Bertin:2012:CTD

[1221] M. C. Bertin, B. M. Pimentel, and J. A. Ramirez. Construction of timedependent dynamical invariants: a new approach. Journal of Mathematical Physics, 53(4):042104, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i4/p042104_s1.

Bayin:2012:CSS

[1222] Selçuk S. Bayin. On the consistency of the solutions of the space fractional Schrödinger equation. Journal of Mathematical Physics, 53(4): 042105, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i4/p042105_s1. See comments [1665].

Aydogdu:2012:EMD

[1223] Oktay Aydogdu, Altug Arda, and Ramazan Sever. Effective-mass Dirac equation for Woods–Saxon potential: Scattering, bound states, and resonances. Journal of Mathematical Physics, 53(4):042106, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/
p042106_s1.

Farenick:2012:CEB

[1224] Douglas Farenick and Michael J. Kozdron. Conditional expectation and Bayes’ rule for quantum random variables and positive operator valued measures. Journal of Mathematical Physics, 53(4):042201, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p042201_s1.

Plaschke:2012:MSL

[1225] M. Plaschke and J. Yngvason. Massless, string localized quantum fields for any helicity. Journal of Mathematical Physics, 53(4):042301, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p042301_s1.

Rivasseau:2012:CRT

[1226] Vincent Rivasseau and Zhituo Wang. Constructive renormalization for Φ_{2}^{4} theory with loop vertex expansion. Journal of Mathematical Physics, 53(4):042302, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p042302_s1.

Varshovi:2012:CAT

[1227] Amir Abbass Varshovi. Consistent anomalies in translation-invariant noncommutative gauge theories. Journal of Mathematical Physics, 53(4): 042303, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i4/p042303_s1.

Chicone:2012:NGM

[1228] C. Chicone and B. Mashhoon. Nonlocal gravity: Modified Poisson’s equation. Journal of Mathematical Physics, 53(4):042501, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/
p042501_s1.

Sanders:2012:ESA

[1229] Ko Sanders. Essential self-adjointness of Wick squares in quasi-free Hadamard representations on curved spacetimes. Journal of Mathematical Physics, 53(4):042502, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i4/p042502_s1.

Backdahl:2012:CNK

[1230] Thomas Bäckdahl and Juan A. Valiente Kroon. Constructing “non-Kerrness” on compact domains. Journal of Mathematical Physics, 53(4): 042503, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i4/p042503_s1.

Jimenez-Lara:2012:CLG

[1231] Lidia Jiménez-Lara and Jaume Llibre. The cored and logarithm galactic potentials: Periodic orbits and integrability. Journal of Mathematical Physics, 53(4):042901, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p042901_s1.

Yamano:2012:BFI

[1232] Takuya Yamano. Bounds for Fisher information and its production under flow. Journal of Mathematical Physics, 53(4):043301, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/
p043301_s1.

Spitzer:2012:CFO

[1233] Wolfgang Spitzer, Shannon Starr, and Lam Tran. Counterexamples to ferromagnetic ordering of energy levels. Journal of Mathematical Physics, 53(4):043302, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p043302_s1.

Silva:2012:CEP

[1234] José B. da Silva, Jr. and Marcelo M. Leite. Critical exponents from parallel plate geometries subject to periodic and antiperiodic boundary conditions. Journal of Mathematical Physics, 53(4):043303, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043303_s1.

Shcherbina:2012:CLT

[1235] Mariya Shcherbina and Brunello Tirozzi. Central limit theorem for fluctuations of linear eigenvalue statistics of large random graphs: Diluted regime. Journal of Mathematical Physics, 53(4):043501, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043501_s1.

Garra:2012:CSF

[1236] Roberto Garra and Federico Polito. Coupled systems of fractional equations related to sound propagation: Analysis and discussion. Journal of Mathematical Physics, 53(4):043502, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i4/p043502_s1.

Lee:2012:SDN

[1237] Kuo-Ming Lee. A second degree Newton method in inverse scattering problem for a crack. Journal of Mathematical Physics, 53(4):043503, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043503_s1.

Schechter:2012:SSS

[1238] Martin Schechter. Steady state solutions for Schrödinger equations governing nonlinear optics. Journal of Mathematical Physics, 53(4):043504, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043504_s1.

Chen:2012:RFQ

[1239] Jialei Chen and Shilin Yang. Restricted forms of quantum supergroups. Journal of Mathematical Physics, 53(4):043505, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/p043505_s1.

Li:2012:ROC

[1240] Hong-Min Li, Biao Li, and Yu-Qi Li. Recursion operators and conservation laws for discrete Lax equations. Journal of Mathematical Physics, 53(4):043506, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p043506_s1.

Cheng:2012:BSF

[1241] Ming Cheng. Bound state for the fractional Schrödinger equation with unbounded potential. Journal of Mathematical Physics, 53(4):043507, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043507_s1.

Yolcu:2012:MLB

[1242] Selma Yildirim Yolcu and Türkay Yolcu. Multidimensional lower bounds for the eigenvalues of Stokes and Dirichlet Laplacian operators. Journal of Mathematical Physics, 53(4):043508, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i4/p043508_s1.

Santillan:2012:HSS

[1243] Osvaldo P. Santillan. Hidden symmetries and supergravity solutions. Journal of Mathematical Physics, 53(4):043509, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/p043509_s1.

Poplavskyi:2012:ABV

[1244] M. Poplavskyi. Asymptotic behavior of the Verblunsky coefficients for the OPUC with a varying weight. Journal of Mathematical Physics, 53(4):043510, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p043510_s1.

Dogan:2012:EHS

[1245] Çaglar Doğan, Fatih Erman, and O. Teoman Turgut. Existence of Hamiltonians for some singular interactions on manifolds. Journal of Mathematical Physics, 53(4):043511, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i4/p043511_s1.

Angulo:2012:UIA

[1246] J. C. Angulo, P. A. Bouvrie, and J. Antolín. Uncertainty inequalities among frequency moments and radial expectation values: Applications to atomic systems. Journal of Mathematical Physics, 53(4):043512, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043512_s1.

Kuznetsova:2012:MRS

[1247] Zhanna Kuznetsova and Francesco Toppan. D-module representations of N = 2, 4, 8 superconformal algebras and their superconformal mechanics. Journal of Mathematical Physics, 53(4):043513, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i4/p043513_s1.

Borgna:2012:INW

[1248] Juan Pablo Borgna, Antonio Degasperis, Mariano Fernando De Leo, and Diego Rial. Integrability of nonlinear wave equations and solvability of their initial value problem. Journal of Mathematical Physics, 53(4): 043701, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i4/p043701_s1.

Sombillo:2012:QTA

[1249] Denny Lane B. Sombillo and Eric A. Galapon. Quantum time of arrival Goursat problem. Journal of Mathematical Physics, 53(4):043702, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043702_s1.

Bokhari:2012:IBV

[1250] Ashfaque H. Bokhari, F. M. Mahomed, and F. D. Zaman. Invariant boundary value problems for a fourth-order dynamic Euler–Bernoulli beam equation. Journal of Mathematical Physics, 53(4):043703, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i4/p043703_s1.

Loubenets:2012:PNL

[1251] Elena R. Loubenets. Publisher’s note: “Local quasi hidden variable modelling and violations of Bell-type inequalities by a multipartite quantum state” [J. Math. Phys. 53, 022201 (2012)]. Journal of Mathematical Physics, 53(4):049901, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i4/p049901_s1. See [1135].

Dokovic:2012:EPI

[1252] Dragomir Ž. – Doković and Andreas Osterloh. Erratum: “On polynomial invariants of several qubits” [J. Math. Phys. 50, 033509 (2009)]. Journal of Mathematical Physics, 53(4):049902, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i4/p049902_s1. See [11].

Macedo:2012:TDC

[1253] D. X. Macedo and I. Guedes. Time-dependent coupled harmonic oscillators. Journal of Mathematical Physics, 53(5):052101, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/

p052101_s1. See [2720].

Zamboni-Rached:2012:SLS

[1254] Michel Zamboni-Rached and Erasmo Recami. Soliton-like solutions to the ordinary Schrödinger equation within standard quantum mechanics. Journal of Mathematical Physics, 53(5):052102, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/p052102_s1.

Sainz:2012:NOW

[1255] Isabel Sainz, Luis Roa, and A. B. Klimov. Non-orthogonal Wigner-like mapping. Journal of Mathematical Physics, 53(5):052103, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/
p052103_s1.

deOliveira:2012:MPD

[1256] César R. de Oliveira and Alessandra A. Verri. Mathematical predominance of Dirichlet condition for the one-dimensional Coulomb potential. Journal of Mathematical Physics, 53(5):052104, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/p052104_s1.

Kye:2012:CBQ

[1257] Seung-Hyeok Kye and Hiroyuki Osaka. Classification of bi-qutrit positive partial transpose entangled edge states by their ranks. Journal of Mathematical Physics, 53(5):052201, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i5/p052201_s1.

Diaz-Marin:2012:ETC

[1258] Homero G. Díaz-Marín and José A. Zapata. Effective theories of connections and curvature: Abelian case. Journal of Mathematical Physics, 53 (5):052301, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p052301_s1.

Agboola:2012:QET

[1259] D. Agboola. Quasi-exact treatment of the relativistic generalized isotonic oscillator. Journal of Mathematical Physics, 53(5):052302, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i5/p052302_s1.

Abe:2012:AAH

[1260] Yasuhiro Abe. Application of abelian holonomy formalism to the elementary theory of numbers. Journal of Mathematical Physics, 53(5): 052303, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p052303_s1.

Bagrov:2012:QMS

[1261] V. G. Bagrov, D. M. Gitman, and A. D. Levin. Quantum motion in superposition of Aharonov–Bohm with some additional electromagnetic fields. Journal of Mathematical Physics, 53(5):052304, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/p052304_s1.

Finster:2012:BLD

[1262] Felix Finster and Jürgen Tolksdorf. Bosonic Loop Diagrams as perturbative solutions of the Classical Field Equations in φ^{4}-theory. Journal of Mathematical Physics, 53(5):052305, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i5/p052305_s1.

Herdegen:2012:ILE

[1263] Andrzej Herdegen. Infrared limit in external field scattering. Journal of Mathematical Physics, 53(5):052306, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i5/p052306_s1.

Costa:2012:SLW

[1264] João L. Costa, Artur Alho, and José Natário. Spherical linear waves in de Sitter spacetime. Journal of Mathematical Physics, 53(5):052501, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i5/p052501_s1.

Tudoran:2012:CIC

[1265] Razvan M. Tudoran and Anania Gîrban. On the completely integrable case of the Rössler system. Journal of Mathematical Physics, 53(5): 052701, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p052701_s1.

Kluck:2012:CMS

[1266] Timo Kluck. On the Calogero–Moser solution by root-type Lax pair. Journal of Mathematical Physics, 53(5):052702, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/p052702_s1.

Meng:2012:LGO

[1267] Guowu Meng. Lorentz group and oriented McIntosh–Cisneros– Zwanziger–Kepler orbits. Journal of Mathematical Physics, 53(5): 052901, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p052901_s1.

Hanyga:2012:SFO

[1268] A. Hanyga and M. Seredyńska. Spatially fractional-order viscoelasticity, non-locality, and a new kind of anisotropy. Journal of Mathematical Physics, 53(5):052902, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i5/p052902_s1.

Chen:2012:CFE

[1269] Zili Chen and Xianwen Zhang. Cometary flow equation with a selfconsistent electromagnetic field. Journal of Mathematical Physics, 53(5): 053301, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p053301_s1.

Gorska:2012:LSD

[1270] K. Górska and K. A. Penson. Lévy stable distributions via associated integral transform. Journal of Mathematical Physics, 53(5):053302, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i5/p053302_s1.

Erman:2012:NRL

[1271] Fatih Erman and O. Teoman Turgut. Non-relativistic Lee model in twodimensional Riemannian manifolds. Journal of Mathematical Physics, 53 (5):053501, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p053501_s1.

Sadeghi:2012:FMS

[1272] J. Sadeghi and M. Rostami. Factorization method in scalar field on AdS_{3} in spherical coordinates. Journal of Mathematical Physics, 53(5): 053502, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p053502_s1.

Tian:2012:TBS

[1273] Kai Tian and Q. P. Liu. The transformations between N = 2 supersymmetric Korteweg–de Vries and Harry Dym equations. Journal of Mathematical Physics, 53(5):053503, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i5/p053503_s1.

Mezzadri:2012:MTE

[1274] F. Mezzadri and N. J. Simm. Moments of the transmission eigenvalues, proper delay times and random matrix theory II. Journal of Mathematical Physics, 53(5):053504, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i5/p053504_s1.

Jorgensen:2012:TRS

[1275] Palle Jorgensen, Steen Pedersen, and Feng Tian. Translation representations and scattering by two intervals. Journal of Mathematical Physics, 53(5):053505, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i5/p053505_s1.

Hong:2012:GBL

[1276] John Meng-Kai Hong, Jiahong Wu, and Juan-Ming Yuan. The generalized Buckley–Leverett and the regularized Buckley–Leverett equations. Journal of Mathematical Physics, 53(5):053701, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i5/p053701_s1.

Mustafa:2012:GSR

[1277] Muhammad I. Mustafa and Salim A. Messaoudi. General stability result for viscoelastic wave equations. Journal of Mathematical Physics, 53(5): 053702, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i5/p053702_s1.

Anco:2012:CLS

[1278] Stephen C. Anco, Steven A. MacNaughton, and Thomas Wolf. Conservation laws and symmetries of quasilinear radial wave equations in multi-dimensions. Journal of Mathematical Physics, 53(5):053703, May 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i5/p053703_s1.

Ugulava:2012:QTR

[1279] A. Ugulava, Z. Toklikishvili, S. Chkhaidze, R. Abramishvili, and L. Chotorlishvili. Quantum theory of rotational isomerism and Hill equation. Journal of Mathematical Physics, 53(6):062101, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p062101_s1.

Bender:2012:MRT

[1280] Carl M. Bender and Mariagiovanna Gianfreda. Matrix representation of the time operator. Journal of Mathematical Physics, 53(6):062102, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p062102_s1.

Diaz-Ortiz:2012:BTC

[1281] Erik I. Díaz-Ortiz and Carlos Villegas-Blas. On a Bargmann transform and coherent states for the n-sphere. Journal of Mathematical Physics, 53 (6):062103, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p062103_s1.

Ghosh:2012:CSN

[1282] Subir Ghosh. Coherent states for the nonlinear harmonic oscillator. Journal of Mathematical Physics, 53(6):062104, June 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i6/p062104_s1. See

comment [2572].

Popper:2012:TOS

[1283] I. Popper, S. Post, and P. Winternitz. Third-order superintegrable systems separable in parabolic coordinates. Journal of Mathematical Physics, 53(6):062105, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p062105_s1.

Seyfarth:2012:CMU

[1284] Ulrich Seyfarth and Kedar S. Ranade. Cyclic mutually unbiased bases, Fibonacci polynomials and Wiedemann’s conjecture. Journal of Mathematical Physics, 53(6):062201, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i6/p062201_s1.

Fetaya:2012:BDQ

[1285] Ethan Fetaya. Bounding the distance of quantum surface codes. Journal of Mathematical Physics, 53(6):062202, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i6/p062202_s1.

Borgna:2012:EGS

[1286] Juan P. Borgna and Diego F. Rial. Existence of ground states for a onedimensional relativistic Schrödinger equation. Journal of Mathematical Physics, 53(6):062301, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p062301_s1.

Minguzzi:2012:CSI

[1287] E. Minguzzi. Causally simple inextendible spacetimes are hole-free. Journal of Mathematical Physics, 53(6):062501, June 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i6/p062501_s1.

Couch:2012:LQN

[1288] W. E. Couch and C. L. Holder. Liouvillian quasi-normal modes of Kerr– Newman black holes. Journal of Mathematical Physics, 53(6):062502, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p062502_s1.

Mach:2012:ASL

[1289] Patryk Mach. All solutions of the n = 5 Lane–Emden equation. Journal of Mathematical Physics, 53(6):062503, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i6/p062503_s1.

Gine:2012:NLI

[1290] Jaume Giné and Daniel Peralta-Salas. A note on the Liouvillian integrability and the qualitative properties of the mass rate equation for black holes. Journal of Mathematical Physics, 53(6):062701, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/
p062701_s1.

Lv:2012:AAS

[1291] Yan Lv and A. J. Roberts. Averaging approximation to singularly perturbed nonlinear stochastic wave equations. Journal of Mathematical Physics, 53(6):062702, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p062702_s1.

Lazrag:2012:DTE

[1292] Lanouar Lazrag. Double tangent extension method and first integrals. Journal of Mathematical Physics, 53(6):062703, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p062703_s1.

Paraschiv:2012:LPT

[1293] Victor Paraschiv. Lagrange–Pizzetti theorem for the quasihomogeneous N-body problem. Journal of Mathematical Physics, 53(6):062901, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p062901_s1.

Langerock:2012:RRC

[1294] B. Langerock, E. García-Toraño Andrés, and F. Cantrijn. Routh reduction and the class of magnetic Lagrangian systems. Journal of Mathematical Physics, 53(6):062902, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i6/p062902_s1.

Imre:2012:NDR

[1295] Kaya Imre and A. Ziya Akcasu. Neutron diffusion in a randomly inhomogeneous multiplying medium with random phase approximation. Journal of Mathematical Physics, 53(6):063301, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i6/p063301_s1.

Morandi:2012:QCL

[1296] O. Morandi. Quantum corrected Liouville model: Mathematical analysis. Journal of Mathematical Physics, 53(6):063302, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p063302_s1.

Bercher:2012:GFI

[1297] J.-F. Bercher. On a (β, q)-generalized Fisher information and inequalities involving q-Gaussian distributions. Journal of Mathematical Physics, 53 (6):063303, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p063303_s1.

Agliari:2012:NSG

[1298] Elena Agliari, Adriano Barra, Raffaella Burioni, and Aldo Di Biasio. Notes on the p-spin glass studied via Hamilton–Jacobi and smoothcavity techniques. Journal of Mathematical Physics, 53(6):063304, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063304_s1.

Faris:2012:FTC

[1299] William G. Faris. The fundamental theorem of calculus for a matroid. Journal of Mathematical Physics, 53(6):063305, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p063305_s1.

Carpentier:2012:SAP

[1300] Sylvain Carpentier, Alberto De Sole, and Victor G. Kac. Some algebraic properties of differential operators. Journal of Mathematical Physics, 53 (6):063501, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p063501_s1.

Al-Ashhab:2012:CTH

[1301] Samer Al-Ashhab. Canonical transformations and Hamiltonian evolutionary systems. Journal of Mathematical Physics, 53(6):063502, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063502_s1.

Haouat:2012:DCP

[1302] S. Haouat and L. Chetouani. On Dirac–Coulomb problem in (2 + 1) dimensional space–time and path integral quantization. Journal of Mathematical Physics, 53(6):063503, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i6/p063503_s1.

Vakhnenko:2012:SSF

[1303] V. O. Vakhnenko and E. J. Parkes. Special singularity function for continuous part of the spectral data in the associated eigenvalue problem for nonlinear equations. Journal of Mathematical Physics, 53(6):063504, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063504_s1.

Popov:2012:OET

[1304] I. Yu. Popov. Operator extensions theory model for electromagnetic fieldelectron interaction. Journal of Mathematical Physics, 53(6):063505, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063505_s1.

Murphy:2012:IMC

[1305] N. B. Murphy and K. M. Golden. The Ising model and critical behavior of transport in binary composite media. Journal of Mathematical Physics, 53(6):063506, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p063506_s1.

Xu:2012:RWB

[1306] Shuwei Xu and Jingsong He. The rogue wave and breather solution of the Gerdjikov–Ivanov equation. Journal of Mathematical Physics, 53(6): 063507, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p063507_s1.

Wu:2012:HES

[1307] Xian Wu. High energy solutions of systems of Kirchhoff-type equations in R^{N}. Journal of Mathematical Physics, 53(6):063508, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p063508_s1.

Triani:2012:EPN

[1308] V. Triani and V. A. Cimmelli. Entropy principle, non-regular processes, and generalized exploitation procedures. Journal of Mathematical Physics, 53(6):063509, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p063509_s1.

Qin:2012:GEA

[1309] Yuming Qin, Jia Ren, and Tianhui Wei. Global existence, asymptotic stability, and uniform attractors for non-autonomous thermoelastic systems with constant time delay. Journal of Mathematical Physics, 53(6): 063701, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p063701_s1.

Huh:2012:SWS

[1310] Hyungjin Huh. Standing waves of the Schrödinger equation coupled with the Chern–Simons gauge field. Journal of Mathematical Physics, 53(6):

063702, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i6/p063702_s1.

Godin:2012:ECP

[1311] Yuri A. Godin. The effective conductivity of a periodic lattice of circular inclusions. Journal of Mathematical Physics, 53(6):063703, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063703_s1.

Dirr:2012:UPM

[1312] Nicolas Dirr, Vaios Laschos, and Johannes Zimmer. Upscaling from particle models to entropic gradient flows. Journal of Mathematical Physics, 53(6):063704, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p063704_s1.

Kamenov:2012:PSN

[1313] Ognyan Y. Kamenov. Periodic solutions of the non-integrable convective fluid equation. Journal of Mathematical Physics, 53(6):063705, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063705_s1.

Isobe:2012:SCLa

[1314] Takeshi Isobe and Antonella Marini. Small coupling limit and multiple solutions to the Dirichlet problem for Yang–Mills connections in four dimensions. I. Journal of Mathematical Physics, 53(6):063706, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063706_s1.

Isobe:2012:SCLb

[1315] Takeshi Isobe and Antonella Marini. Small coupling limit and multiple solutions to the Dirichlet problem for Yang–Mills connections in four dimensions. II. Journal of Mathematical Physics, 53(6):063707, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i6/p063707_s1.

Wang:2012:ISC

[1316] Mingxin Wang and Shengqi Yu. An interacting system of the Camassa– Holm and Degasperis–Procesi equations. Journal of Mathematical Physics, 53(6):063708, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p063708_s1.

Moutsinga:2012:BES

[1317] Octave Moutsinga. Burgers’ equation and the sticky particles model. Journal of Mathematical Physics, 53(6):063709, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/p063709_s1.

Xue:2012:QSJ

[1318] Bo Xue, Xianguo Geng, and Fang Li. Quasiperiodic solutions of Jaulent–Miodek equations with a negative flow. Journal of Mathematical Physics, 53(6):063710, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i6/p063710_s1.

Lima:2012:YAP

[1319] Jonas R. F. Lima, M. Vieira, C. Furtado, F. Moraes, and Cleverson Filgueiras. Yet another position-dependent mass quantum model. Journal of Mathematical Physics, 53(7):072101, July 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i7/p072101_s1.

Bonzom:2012:NHT

[1320] Valentin Bonzom and Etera R. Livine. A new Hamiltonian for the topological BF phase with spinor networks. Journal of Mathematical Physics, 53(7):072201, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p072201_s1.

Fritz:2012:PDB

[1321] Tobias Fritz. Polyhedral duality in Bell scenarios with two binary observables. Journal of Mathematical Physics, 53(7):072202, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p072202_s1.

Chen:2012:CSR

[1322] Jianxin Chen, Zhengfeng Ji, Mary Beth Ruskai, Bei Zeng, and Duan-Lu Zhou. Comment on some results of Erdahl and the convex structure of reduced density matrices. Journal of Mathematical Physics, 53(7): 072203, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p072203_s1. See [1].

Oeckl:2012:SRR

[1323] Robert Oeckl. The Schrödinger representation and its relation to the holomorphic representation in linear and affine field theory. Journal of Mathematical Physics, 53(7):072301, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p072301_s1.

Chaichian:2012:EPS

[1324] M. Chaichian, A. Tureanu, and R. B. Zhang. Extended Poincaré supersymmetry in three dimensions and supersymmetric anyons. Journal of Mathematical Physics, 53(7):072302, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p072302_s1.

Ould-Lahoucine:2012:EGF

[1325] H. K. Ould-Lahoucine and L. Chetouani. Exact Green function for a Dirac particle in a weak gravitational plane wave field. Alternative path integral approach. Journal of Mathematical Physics, 53(7):072303, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p072303_s1.

Saidi:2012:TAF

[1326] E. H. Saidi, O. Fassi-Fehri, and M. Bousmina. Topological aspects of fermions on hyperdiamond. Journal of Mathematical Physics, 53(7): 072304, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p072304_s1.

Balehowsky:2012:RFA

[1327] T. Balehowsky and E. Woolgar. The Ricci flow of asymptotically hyperbolic mass and applications. Journal of Mathematical Physics, 53(7): 072501, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p072501_s1.

Papadopoulos:2012:LHS

[1328] G. O. Papadopoulos and Th. Grammenos. Locally homogeneous spaces, induced Killing vector fields and applications to Bianchi prototypes. Journal of Mathematical Physics, 53(7):072502, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p072502_s1.

Wang:2012:NON

[1329] Jian Wang and Yong Wang. Nonminimal operators and non-commutative residue. Journal of Mathematical Physics, 53(7):072503, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p072503_s1.

Sun:2012:FPE

[1330] Xu Sun and Jinqiao Duan. Fokker–Planck equations for nonlinear dynamical systems driven by non-Gaussian Lévy processes. Journal of Mathematical Physics, 53(7):072701, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p072701_s1. See comment

[2709].

Hu:2012:PAN

[1331] Zhiyuan Hu and Yejuan Wang. Pullback attractors for a nonautonomous nonclassical diffusion equation with variable delay. Journal of Mathematical Physics, 53(7):072702, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p072702_s1.

Jin:2012:EAM

[1332] Lingyu Jin, Shaomei Fang, and Boling Guo. Existence of attractors for modified Zakharov equations for plasmas with a quantum correction. Journal of Mathematical Physics, 53(7):072703, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p072703_s1.

Lacomba:2012:DGH

[1333] Ernesto A. Lacomba and Jaume Llibre. Dynamics of a galactic Hamiltonian system. Journal of Mathematical Physics, 53(7):072901, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p072901_s1.

Tsai:2012:DCR

[1334] Ya lun Tsai. Dziobek configurations of the restricted (N + 1)-body problem with equal masses. Journal of Mathematical Physics, 53(7): 072902, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p072902_s1.

Vankerschaver:2012:HPP

[1335] J. Vankerschaver, H. Yoshimura, and M. Leok. The Hamilton–Pontryagin principle and multi-Dirac structures for classical field theories. Journal of Mathematical Physics, 53(7):072903, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p072903_s1.

Masterov:2012:SEC

[1336] Ivan Masterov. N = 2 supersymmetric extension of l-conformal Galilei algebra. Journal of Mathematical Physics, 53(7):072904, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p072904_s1.

Leok:2012:HJT

[1337] Melvin Leok, Tomoki Ohsawa, and Diana Sosa. Hamilton–Jacobi theory for degenerate Lagrangian systems with holonomic and nonholonomic constraints. Journal of Mathematical Physics, 53(7):072905, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p072905_s1.

Deng:2012:WPV

[1338] Chao Deng and Shangbin Cui. Well-posedness of the viscous Boussinesq system in Besov spaces of negative regular index s = −1. Journal of Mathematical Physics, 53(7):073101, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p073101_s1.

Popovych:2012:SPP

[1339] Roman O. Popovych and Alexander Bihlo. Symmetry preserving parameterization schemes. Journal of Mathematical Physics, 53(7):073102, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073102_s1.

Lee:2012:NGR

[1340] Jihoon Lee. Notes on the geometric regularity criterion of 3D Navier– Stokes system. Journal of Mathematical Physics, 53(7):073103, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073103_s1.

Livermore:2012:SES

[1341] Philip W. Livermore and Rainer Hollerbach. Successive elimination of shear layers by a hierarchy of constraints in inviscid spherical-shell flows. Journal of Mathematical Physics, 53(7):073104, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073104_s1.

Holik:2012:QEM

[1342] F. Holik and A. Plastino. Quantal effects and MaxEnt. Journal of Mathematical Physics, 53(7):073301, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073301_s1.

Savoie:2012:ZFO

[1343] Baptiste Savoie. On the zero-field orbital magnetic susceptibility of Bloch electrons in graphene-like solids: Some rigorous results. Journal of Mathematical Physics, 53(7):073302, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073302_s1.

Fialowski:2012:FRW

[1344] Alice Fialowski. Formal rigidity of the Witt and Virasoro algebra. Journal of Mathematical Physics, 53(7):073501, July 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073501_s1.

Sako:2012:EFN

[1345] Akifumi Sako, Toshiya Suzuki, and Hiroshi Umetsu. Explicit formulas for noncommutative deformations of cP^{N} and cH^{N}. Journal of Mathematical Physics, 53(7):073502, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073502_s1.

Khreibani:2012:FRH

[1346] H. Khreibani and V. Shramchenko. Fuchsian Riemann–Hilbert problem for “real doubles” of Hurwitz Frobenius manifolds. Journal of Mathematical Physics, 53(7):073503, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073503_s1.

Bhuvaneswari:2012:CIN

[1347] A. Bhuvaneswari, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. On the complete integrability of a nonlinear oscillator from group theoretical perspective. Journal of Mathematical Physics, 53(7): 073504, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p073504_s1.

Arnlind:2012:DNT

[1348] Joakim Arnlind and Harald Grosse. Deformed noncommutative tori. Journal of Mathematical Physics, 53(7):073505, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073505_s1.

Guo:2012:RHA

[1349] Boling Guo and Liming Ling. Riemann–Hilbert approach and N-soliton formula for coupled derivative Schrödinger equation. Journal of Mathematical Physics, 53(7):073506, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073506_s1.

Zabrodin:2012:QPCa

[1350] A. Zabrodin and A. Zotov. Quantum Painlevé–Calogero correspondence. Journal of Mathematical Physics, 53(7):073507, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073507_s1.

Zabrodin:2012:QPCb

[1351] A. Zabrodin and A. Zotov. Quantum Painlevé–Calogero correspondence for Painlevé VI. Journal of Mathematical Physics, 53(7):073508, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073508_s1.

Sare:2012:ORD

[1352] Hugo D. Fernández Sare and Jaime E. Muñoz Rivera. Optimal rates of decay in 2-d thermoelasticity with second sound. Journal of Mathematical Physics, 53(7):073509, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073509_s1.

Osei:2012:SLI

[1353] Prince K. Osei and Bernd J. Schroers. On the semiduals of local isometry groups in three-dimensional gravity. Journal of Mathematical Physics, 53 (7):073510, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p073510_s1.

Adamovic:2012:ERL

[1354] Drazen Adamović and Antun Milas. An explicit realization of logarithmic modules for the vertex operator algebra W_{p,p′}. Journal of Mathematical Physics, 53(7):073511, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p073511_s1.

Grahovski:2012:GZS

[1355] Georgi G. Grahovski. The generalised Zakharov–Shabat system and the gauge group action. Journal of Mathematical Physics, 53(7):073512, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073512_s1.

Chen:2012:SAG

[1356] Jinbing Chen. Some algebro-geometric solutions for the coupled modified Kadomtsev–Petviashvili equations arising from the Neumann type systems. Journal of Mathematical Physics, 53(7):073513, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p073513_s1.

Cojuhari:2012:LPS

[1357] Petru A. Cojuhari and Sergii Kuzhel. Lax–Phillips scattering theory for PT-symmetric ρ-perturbed operators. Journal of Mathematical Physics, 53(7):073514, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p073514_s1.

OCarroll:2012:BNB

[1358] Michael O’Carroll. Bounds on the number of bound states in the transfer matrix spectrum for some weakly correlated lattice models. Journal of Mathematical Physics, 53(7):073515, July 2012. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p073515_s1.

MacPherson:2012:MST

[1359] Robert MacPherson and Benjamin Schweinhart. Measuring shape with topology. Journal of Mathematical Physics, 53(7):073516, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p073516_s1.

Ren:2012:EUS

[1360] Yong Ren and R. Sakthivel. Existence, uniqueness, and stability of mild solutions for second-order neutral stochastic evolution equations with infinite delay and Poisson jumps. Journal of Mathematical Physics, 53 (7):073517, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p073517_s1.

deLange:2012:TIP

[1361] O. L. de Lange, R. E. Raab, and A. Welter. Translational invariance, the Post constraint and uniqueness in macroscopic electrodynamics. Journal of Mathematical Physics, 53(7):073518, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p073518_s1.

Lu:2012:DSQ

[1362] Zhiqin Lu and Julie Rowlett. On the discrete spectrum of quantum layers. Journal of Mathematical Physics, 53(7):073519, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073519_s1.

Wang:2012:CMN

[1363] Yanjin Wang. Critical magnetic number in the magnetohydrodynamic Rayleigh–Taylor instability. Journal of Mathematical Physics, 53(7): 073701, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p073701_s1.

Yan:2012:MAS

[1364] Guozheng Yan, Jianguo Ye, and Jun Guo. Mathematical analysis of scattering problems for partially coated obstacles and cracks. Journal of Mathematical Physics, 53(7):073702, July 2012. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p073702_s1.

Hu:2012:ASP

[1365] Yanbo Hu. Axisymmetric solutions of the pressure-gradient system. Journal of Mathematical Physics, 53(7):073703, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/p073703_s1.

Bartolucci:2012:AOE

[1366] Daniele Bartolucci and Francesca De Marchis. On the Ambjorn–Olesen electroweak condensates. Journal of Mathematical Physics, 53(7):073704, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073704_s1.

Pelinovsky:2012:ASS

[1367] Dmitry E. Pelinovsky and Atanas Stefanov. Asymptotic stability of small gap solitons in nonlinear Dirac equations. Journal of Mathematical Physics, 53(7):073705, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p073705_s1.

Kamvissis:2012:LTA

[1368] Spyridon Kamvissis and Gerald Teschl. Long-time asymptotics of the periodic Toda lattice under short-range perturbations. Journal of Mathematical Physics, 53(7):073706, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i7/p073706_s1.

Mao:2012:SSS

[1369] Shikuan Mao. Singularities for solutions to Schrödinger equations with asymptotically constant magnetic fields. Journal of Mathematical Physics, 53(7):073707, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p073707_s1.

Amdeberhan:2012:FEC

[1370] Tewodros Amdeberhan, David Borwein, Jonathan M. Borwein, and Armin Straub. On formulas for π experimentally conjectured by Jauregui–Tsallis. Journal of Mathematical Physics, 53(7):073708, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073708_s1.

Curtis:2012:NGP

[1371] Christopher W. Curtis. On nonlocal Gross–Pitaevskii equations with periodic potentials. Journal of Mathematical Physics, 53(7):073709, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i7/p073709_s1.

Bies:2012:DMK

[1372] Piotr Michal Bies, Przemyslaw Górka, and Enrique G. Reyes. The dual modified Korteweg–de Vries–Fokas–Qiao equation: Geometry and local analysis. Journal of Mathematical Physics, 53(7):073710, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i7/
p073710_s1.

Li:2012:EFO

[1373] Junjie Li. Erratum: “On a fourth order degenerate parabolic equation in higher space dimension” [J. Math. Phys. 50, 123524 (2009)]. Journal of Mathematical Physics, 53(7):079901, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p079901_s1. See [21].

Werner:2012:EDD

[1374] Ivan Werner. Erratum: “Dynamically defined measures and equilibrium states” [J. Math. Phys. 52, 122701 (2011)]. Journal of Mathematical Physics, 53(7):079902, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i7/p079902_s1. See [1063].

Chen:2012:EDR

[1375] Lin Chen and Dragomir Ž. – Doković. Erratum: “Description of rank four entangled states of two qutrits having positive partial transpose” [J. Math. Phys. 52, 122203 (2011)]. Journal of Mathematical Physics, 53(7): 079903, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p079903_s1. See [1059].

Hamzavi:2012:ETI

[1376] M. Hamzavi, M. Eshghi, and S. M. Ikhdair. Effect of tensor interaction in the Dirac-attractive radial problem under pseudospin symmetry limit. Journal of Mathematical Physics, 53(8):082101, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p082101_s1.

Ruby:2012:OQA

[1377] V. Chithiika Ruby and M. Senthilvelan. An observation of quadratic algebra, dual family of nonlinear coherent states and their non-classical properties, in the generalized isotonic oscillator. Journal of Mathematical Physics, 53(8):082102, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082102_s1.

Phan:2012:GRL

[1378] Ngoc-Hung Phan and Van-Hoang Le. Generalized Runge–Lenz vector and a hidden symmetry of the nine-dimensional MICZ–Kepler problem. Journal of Mathematical Physics, 53(8):082103, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p082103_s1.

Zhang:2012:SDE

[1379] Min-Cang Zhang and Zhi-Cun Liu. Solution of the Dirac equation with pseudospin symmetry for a new harmonic oscillatory ring-shaped noncentral potential. Journal of Mathematical Physics, 53(8):082104, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p082104_s1.

Langmann:2012:SIK

[1380] Edwin Langmann and Kouichi Takemura. Source identity and kernel functions for Inozemtsev-type systems. Journal of Mathematical Physics, 53(8):082105, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082105_s1.

Viennot:2012:GPQ

[1381] David Viennot. Geometric phases in quantum control disturbed by classical stochastic processes. Journal of Mathematical Physics, 53(8): 082106, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i8/p082106_s1.

Falaye:2012:KGE

[1382] Babatunde James Falaye. The Klein–Gordon equation with ring-shaped potentials: Asymptotic iteration method. Journal of Mathematical Physics, 53(8):082107, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082107_s1.

Schulze-Halberg:2012:TDM

[1383] Axel Schulze-Halberg. Two-dimensional magnetic Schrödinger equations, Darboux transformations and solutions of associated auxiliary equations. Journal of Mathematical Physics, 53(8):082108, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p082108_s1.

Thrapsaniotis:2012:PIA

[1384] E. G. Thrapsaniotis. Path integral approach to the photoionization of one active electron atom by a pulsed Fock state: Fluctuations contribution. Journal of Mathematical Physics, 53(8):082109, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p082109_s1.

Demirel:2012:SSF

[1385] Semra Demirel. The spectral shift function and Levinson’s theorem for quantum star graphs. Journal of Mathematical Physics, 53(8): 082110, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i8/p082110_s1.

Dominy:2012:DHL

[1386] Jason Dominy and Herschel Rabitz. Dynamic homotopy and landscape dynamical set topology in quantum control. Journal of Mathematical Physics, 53(8):082201, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082201_s1.

Hodgkinson:2012:HOD

[1387] Lee Hodgkinson and Jorma Louko. How often does the Unruh–DeWitt detector click beyond four dimensions? Journal of Mathematical Physics, 53(8):082301, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082301_s1.

Babichenko:2012:MPM

[1388] Andrei Babichenko and Alessandro Torrielli. Multi-parametric R-matrix for the ∫↕(2|1) Yangian. Journal of Mathematical Physics, 53(8):082302, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i8/p082302_s1.

Much:2012:WLQ

[1389] A. Much. Wedge-local quantum fields on a nonconstant noncommutative spacetime. Journal of Mathematical Physics, 53(8):082303, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p082303_s1.

Klauder:2012:RRR

[1390] John R. Klauder. Recent results regarding affine quantum gravity. Journal of Mathematical Physics, 53(8):082501, August 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i8/p082501_s1.

deBustos:2012:POI

[1391] María T. de Bustos, Juan L. G. Guirao, Juan A. Vera, and Jesús Vigo-Aguiar. Periodics orbits and C^{1}-integrability in the planar Stark–Zeeman problem. Journal of Mathematical Physics, 53(8):082701, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/
p082701_s1.

Calogero:2012:NSD

[1392] F. Calogero and F. Leyvraz. New solvable discrete-time many-body problem featuring several arbitrary parameters. Journal of Mathematical Physics, 53(8):082702, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p082702_s1.

Xie:2012:EUA

[1393] Yongqin Xie, Kaixuan Zhu, and Chunyou Sun. The existence of uniform attractors for non-autonomous reaction-diffusion equations on the whole space. Journal of Mathematical Physics, 53(8):082703, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p082703_s1.

Combot:2012:TOI

[1394] Thierry Combot and Christoph Koutschan. Third order integrability conditions for homogeneous potentials of degree −1. Journal of Mathematical Physics, 53(8):082704, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i8/p082704_s1.

Boldrighini:2012:ESC

[1395] C. Boldrighini, S. Frigio, and P. Maponi. Exploding solutions of the complex two-dimensional Burgers equations: Computer simulations. Journal of Mathematical Physics, 53(8):083101, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v53/i8/p083101_s1. See pub-

lisher’s note [1933].

Banerjee:2012:LDD

[1396] Pratyay Banerjee and B. Basu-Mallick. Level density distribution for onedimensional vertex models related to Haldane–Shastry like spin chains. Journal of Mathematical Physics, 53(8):083301, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p083301_s1.

Tautz:2012:RPU

[1397] R. C. Tautz and I. Lerche. Relativistic plasmas in uniform magnetic fields. I. General off-axis coupling methods. Journal of Mathematical Physics, 53(8):083302, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p083302_s1.

Baik:2012:JDM

[1398] Jinho Baik, Karl Liechty, and Grégory Schehr. On the joint distribution of the maximum and its position of the Airy_{2} process minus a parabola. Journal of Mathematical Physics, 53(8):083303, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p083303_s1.

Skrypnyk:2012:QBL

[1399] T. Skrypnyk. Quasigraded bases in loop algebras and classical rational r-matrices. Journal of Mathematical Physics, 53(8):083501, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p083501_s1.

Wang:2012:FDT

[1400] Fan-Fan Wang and Ai-Hua Chen. N-fold Darboux transformations and soliton solutions of three nonlinear equations. Journal of Mathematical Physics, 53(8):083502, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p083502_s1.

Kojima:2012:SVO

[1401] Takeo Kojima. Screenings and vertex operators of quantum superalgebra U_{q}(ˆ∫↕(N|1)). Journal of Mathematical Physics, 53(8):083503, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p083503_s1.

Artamonov:2012:NPC

[1402] Dmitry Artamonov and Valentina Goloubeva. Noncommutative Pfaffians and classification of states of five-dimensional quasi-spin. Journal of Mathematical Physics, 53(8):083504, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i8/p083504_s1.

Ebrahimi-Fard:2012:ERI

[1403] Kurusch Ebrahimi-Fard and Frédéric Patras. Exponential renormalisation. II. Bogoliubov’s R-operation and momentum subtraction schemes. Journal of Mathematical Physics, 53(8):083505, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/p083505_s1.

Kruglikov:2012:SUP

[1404] Boris Kruglikov and Oleg Morozov. SDiff(2) and uniqueness of the Plebański equation. Journal of Mathematical Physics, 53(8):083506, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i8/p083506_s1.

Willatzen:2012:SCE

[1405] M. Willatzen, A. Pors, and J. Gravesen. Strong curvature effects in Neumann wave problems. Journal of Mathematical Physics, 53(8): 083507, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i8/p083507_s1.

Gelantalis:2012:RVS

[1406] Michael Gelantalis and Peter Sternberg. Rotating 2N-vortex solutions to the Gross–Pitaevskii equation on S^{2}. Journal of Mathematical Physics, 53(8):083701, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p083701_s1.

Guo:2012:ESS

[1407] Boling Guo and Daiwen Huang. Existence and stability of standing waves for nonlinear fractional Schrödinger equations. Journal of Mathematical Physics, 53(8):083702, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i8/p083702_s1.

Speck:2012:NST

[1408] Jared Speck. The nonlinear stability of the trivial solution to the Maxwell–Born–Infeld system. Journal of Mathematical Physics, 53(8): 083703, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i8/p083703_s1.

Bayin:2012:CCS

[1409] Selçuk S. Bayin. Comment on “On the consistency of the solutions of the space fractional Schrödinger equation” [J. Math. Phys. 53, 042105 (2012)]. Journal of Mathematical Physics, 53(8):084101, August 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i8/

p084101_s1. See [1222, 1665].

Nachtergaele:2012:ISI

[1410] Bruno Nachtergaele, Jan Philip Solovej, and Horng-Tzer Yau. Introduction to Special Issue: In Honor of Elliott Lieb’s 80th birthday. Journal of Mathematical Physics, 53(9):095101, September 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i9/p095101_s1.

Frank:2012:LTI

[1411] Rupert L. Frank and Robert Seiringer. Lieb–Thirring inequality for a model of particles with point interactions. Journal of Mathematical Physics, 53(9):095201, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i9/p095201_s1.

Erdos:2012:RSC

[1412] László Erdős, Søren Fournais, and Jan Philip Solovej. Relativistic Scott correction in self-generated magnetic fields. Journal of Mathematical Physics, 53(9):095202, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i9/p095202_s1.

Correggi:2012:CRS

[1413] M. Correggi, F. Pinsker, N. Rougerie, and J. Yngvason. Critical rotational speeds for superfluids in homogeneous traps. Journal of Mathematical Physics, 53(9):095203, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i9/p095203_s1.

Dolbeault:2012:SEF

[1414] Jean Dolbeault, Maria J. Esteban, and Michael Loss. Symmetry of extremals of functional inequalities via spectral estimates for linear operators. Journal of Mathematical Physics, 53(9):095204, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/
p095204_s1.

Aizenman:2012:ACS

[1415] Michael Aizenman and Simone Warzel. Absolutely continuous spectrum implies ballistic transport for quantum particles in a random potential on tree graphs. Journal of Mathematical Physics, 53(9):095205, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i9/p095205_s1.

Kostrykin:2012:BMM

[1416] Vadim Kostrykin, Jürgen Potthoff, and Robert Schrader. Brownian motions on metric graphs. Journal of Mathematical Physics, 53(9):095206, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i9/p095206_s1.

Maier:2012:SIC

[1417] Thomas Maier and Heinz Siedentop. Stability of impurities with Coulomb potential in graphene with homogeneous magnetic field. Journal of Mathematical Physics, 53(9):095207, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/p095207_s1.

Thomas:2012:PEF

[1418] Lawrence E. Thomas. Persistent energy flow for a stochastic wave equation model in nonequilibrium statistical mechanics. Journal of Mathematical Physics, 53(9):095208, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i9/p095208_s1.

Blanc:2012:ETL

[1419] Xavier Blanc and Mathieu Lewin. Existence of the thermodynamic limit for disordered quantum Coulomb systems. Journal of Mathematical Physics, 53(9):095209, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i9/p095209_s1.

Narnhofer:2012:EBI

[1420] Heide Narnhofer and Walter Thirring. Entanglement, Bell inequality and all that. Journal of Mathematical Physics, 53(9):095210, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i9/p095210_s1.

Lebowitz:2012:LLY

[1421] Joel L. Lebowitz, David Ruelle, and Eugene R. Speer. Location of the Lee–Yang zeros and absence of phase transitions in some Ising spin systems. Journal of Mathematical Physics, 53(9):095211, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/
p095211_s1.

Mattis:2012:CMF

[1422] Daniel C. Mattis. A contribution to the many-fermion problem. Journal of Mathematical Physics, 53(9):095212, September 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i9/p095212_s1. See

note [1508].

Benguria:2012:ICE

[1423] Rafael D. Benguria and Matej Tusek. Indirect Coulomb energy for two-dimensional atoms. Journal of Mathematical Physics, 53(9):095213, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i9/p095213_s1.

Giuliani:2012:SLE

[1424] Alessandro Giuliani, Rafael L. Greenblatt, and Vieri Mastropietro. The scaling limit of the energy correlations in non-integrable Ising models. Journal of Mathematical Physics, 53(9):095214, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/
p095214_s1.

Miyao:2012:SDR

[1425] Tadahiro Miyao and Herbert Spohn. Scale dependence of the retarded van der Waals potential. Journal of Mathematical Physics, 53(9):095215, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i9/p095215_s1.

Brydges:2012:FES

[1426] David Brydges and Thomas Spencer. Fluctuation estimates for subquadratic gradient field actions. Journal of Mathematical Physics, 53 (9):095216, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i9/p095216_s1.

Tsai:2012:SCP

[1427] Chung-Jun Tsai, Li-Sheng Tseng, and Shing-Tung Yau. Symplectic cohomologies on phase space. Journal of Mathematical Physics, 53(9):095217, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i9/p095217_s1.

Frohlich:2012:SAL

[1428] Jürg Fröhlich and Pierre-François Rodriguez. Some applications of the Lee–Yang theorem. Journal of Mathematical Physics, 53(9):095218, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i9/p095218_s1.

Kennedy:2012:SSA

[1429] Tom Kennedy. Simulating self-avoiding walks in bounded domains. Journal of Mathematical Physics, 53(9):095219, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/p095219_s1.

Hainzl:2012:GSP

[1430] Christian Hainzl, Mathieu Lewin, and Christof Sparber. Ground state properties of graphene in Hartree–Fock theory. Journal of Mathematical Physics, 53(9):095220, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i9/p095220_s1.

Bourgade:2012:BUG

[1431] Paul Bourgade, László Erdős, and Horng-Tzer Yau. Bulk universality of general β-ensembles with non-convex potential. Journal of Mathematical Physics, 53(9):095221, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i9/p095221_s1.

Hepp:2012:CDB

[1432] K. Hepp. Coherence and decoherence in the brain. Journal of Mathematical Physics, 53(9):095222, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i9/p095222_s1.

Kiessling:2012:HLB

[1433] Michael K.-H. Kiessling. The Hartree limit of Born’s ensemble for the ground state of a bosonic atom or ion. Journal of Mathematical Physics, 53(9):095223, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i9/p095223_s1.

Abraham:2012:GWT

[1434] D. B. Abraham. Generalised Wick theorem and the Fredholm alternative. Journal of Mathematical Physics, 53(9):095224, September 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i9/

p095224_s1. See note [1611].

Garcia-Ravelo:2012:EFG

[1435] J. García-Ravelo, A. Schulze-Halberg, and A. L. Trujillo. Explicit formulas for generalized harmonic perturbations of the infinite quantum well with an application to Mathieu equations. Journal of Mathematical Physics, 53(10):102101, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102101_s1.

Korcyl:2012:ACW

[1436] Piotr Korcyl. Analytic calculation of Witten index in D = 2 supersymmetric Yang–Mills quantum mechanics. Journal of Mathematical Physics, 53(10):102102, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p102102_s1.

Kiukas:2012:CIC

[1437] J. Kiukas, P. Lahti, J. Schultz, and R. F. Werner. Characterization of informational completeness for covariant phase space observables. Journal of Mathematical Physics, 53(10):102103, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p102103_s1.

Rodriguez-Romo:2012:LHA

[1438] Suemi Rodríguez-Romo. On left Hopf algebras within the framework of inhomogeneous quantum groups for particle algebras. Journal of Mathematical Physics, 53(10):102104, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102104_s1.

Valencia:2012:HVE

[1439] Gerardo Daniel Valencia and Ricardo Weder. High-velocity estimates and inverse scattering for quantum N-body systems with Stark effect. Journal of Mathematical Physics, 53(10):102105, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p102105_s1.

Mkrtchyan:2012:CEU

[1440] R. L. Mkrtchyan, A. N. Sergeev, and A. P. Veselov. Casimir eigenvalues for universal Lie algebra. Journal of Mathematical Physics, 53(10): 102106, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p102106_s1.

Amour:2012:EIF

[1441] Laurent Amour. Explicit isospectral flows associated to the AKNS operator on the unit interval. II. Journal of Mathematical Physics, 53(10): 102107, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p102107_s1.

Gridnev:2012:BST

[1442] Dmitry K. Gridnev. Bound states at threshold resulting from Coulomb repulsion. Journal of Mathematical Physics, 53(10):102108, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102108_s1.

Carinena:2012:QFP

[1443] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. The quantum free particle on spherical and hyperbolic spaces: a curvature dependent approach. II. Journal of Mathematical Physics, 53(10):102109, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102109_s1.

Calcagni:2012:QMF

[1444] Gianluca Calcagni, Giuseppe Nardelli, and Marco Scalisi. Quantum mechanics in fractional and other anomalous spacetimes. Journal of Mathematical Physics, 53(10):102110, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102110_s1.

Aydogdu:2012:SSP

[1445] Oktay Aydogdu, Altug Arda, and Ramazan Sever. Scattering of a spinless particle by an asymmetric Hulthén potential within the effective mass formalism. Journal of Mathematical Physics, 53(10):102111, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102111_s1.

deMicheli:2012:QDS

[1446] Fiorenza de Micheli and Jorge Zanelli. Quantum degenerate systems. Journal of Mathematical Physics, 53(10):102112, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p102112_s1.

delaMadrid:2012:RHS

[1447] Rafael de la Madrid. The rigged Hilbert space approach to the Gamow states. Journal of Mathematical Physics, 53(10):102113, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p102113_s1.

Chen:2012:GSS

[1448] Jianxin Chen, Zhengfeng Ji, David Kribs, Zhaohui Wei, and Bei Zeng. Ground-state spaces of frustration-free Hamiltonians. Journal of Mathematical Physics, 53(10):102201, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102201_s1.

Elgart:2012:NSA

[1449] Alexander Elgart and George A. Hagedorn. A note on the switching adiabatic theorem. Journal of Mathematical Physics, 53(10):102202, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102202_s1.

Holweck:2012:GDE

[1450] Frédéric Holweck, Jean-Gabriel Luque, and Jean-Yves Thibon. Geometric descriptions of entangled states by auxiliary varieties. Journal of Mathematical Physics, 53(10):102203, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p102203_s1.

Ha:2012:SPA

[1451] Kil-Chan Ha and Seung-Hyeok Kye. The structural physical approximations and optimal entanglement witnesses. Journal of Mathematical Physics, 53(10):102204, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p102204_s1.

Chen:2012:ECC

[1452] Lin Chen and Dragomir Ž. – Doković. Equivalence classes and canonical forms for two-qutrit entangled states of rank four having positive partial transpose. Journal of Mathematical Physics, 53(10):102205, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102205_s1.

Weis:2012:EDN

[1453] Stephan Weis and Andreas Knauf. Entropy distance: New quantum phenomena. Journal of Mathematical Physics, 53(10):102206, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102206_s1.

Childs:2012:LTG

[1454] Andrew M. Childs and David Gosset. Levinson’s theorem for graphs II. Journal of Mathematical Physics, 53(10):102207, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p102207_s1.

Heinosaari:2012:EQO

[1455] Teiko Heinosaari, Maria A. Jivulescu, David Reeb, and Michael M. Wolf. Extending quantum operations. Journal of Mathematical Physics, 53(10):102208, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p102208_s1.

Huang:2012:PTB

[1456] Zejun Huang, Chi-Kwong Li, Edward Poon, and Nung-Sing Sze. Physical transformations between quantum states. Journal of Mathematical Physics, 53(10):102209, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p102209_s1.

Aubrun:2012:RRS

[1457] Guillaume Aubrun and Ion Nechita. Realigning random states. Journal of Mathematical Physics, 53(10):102210, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p102210_s1.

Ydri:2012:FWF

[1458] Badis Ydri and Adel Bouchareb. The fate of the Wilson–Fisher fixed point in non-commutative φ^{4}. Journal of Mathematical Physics, 53(10): 102301, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p102301_s1.

Teo:2012:ZFT

[1459] L. P. Teo. Zero and finite temperature Casimir effect of massive vector field between real metals. Journal of Mathematical Physics, 53(10): 102302, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p102302_s1.

Navarro:2012:EED

[1460] J. Navarro and J. B. Sancho. Energy and electromagnetism of a differential k-form. Journal of Mathematical Physics, 53(10):102501, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p102501_s1.

Noutchegueme:2012:GDR

[1461] Norbert Noutchegueme and Alexis Nangue. Global dynamics for a relativistic charged matter in the presence of a massive scalar field and the presence of a cosmological constant on Bianchi spacetimes. Journal of Mathematical Physics, 53(10):102502, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p102502_s1.

Nungesser:2012:FNL

[1462] Ernesto Nungesser. Future non-linear stability for solutions of the Einstein–Vlasov system of Bianchi types II and VI_{0}. Journal of Mathematical Physics, 53(10):102503, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102503_s1.

Cortier:2012:FAH

[1463] Julien Cortier. A family of asymptotically hyperbolic manifolds with arbitrary energy-momentum vectors. Journal of Mathematical Physics, 53(10):102504, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p102504_s1.

Yoshimura:2012:SDB

[1464] Kazuyuki Yoshimura. Stability of discrete breathers in nonlinear Klein– Gordon type lattices with pure anharmonic couplings. Journal of Mathematical Physics, 53(10):102701, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102701_s1.

Geng:2012:RAQ

[1465] Jiansheng Geng and Jian Wu. Real analytic quasi-periodic solutions for the derivative nonlinear Schrödinger equations. Journal of Mathematical Physics, 53(10):102702, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p102702_s1.

Bartuccelli:2012:APO

[1466] Michele Bartuccelli, Jonathan Deane, and Guido Gentile. Attractiveness of periodic orbits in parametrically forced systems with time-increasing friction. Journal of Mathematical Physics, 53(10):102703, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p102703_s1.

Bertin:2012:TDB

[1467] M. C. Bertin, B. M. Pimentel, and C. E. Valcárcel. Two-dimensional background field gravity: a Hamilton–Jacobi analysis. Journal of Mathematical Physics, 53(10):102901, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p102901_s1.

Ignatova:2012:LES

[1468] Mihaela Ignatova, Igor Kukavica, and Mohammed Ziane. Local existence of solutions to the free boundary value problem for the primitive equations of the ocean. Journal of Mathematical Physics, 53(10): 103101, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103101_s1.

Dumitriu:2012:SED

[1469] Ioana Dumitriu. Smallest eigenvalue distributions for two classes of β-Jacobi ensembles. Journal of Mathematical Physics, 53(10):103301, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p103301_s1.

Lytvtnenko:2012:TFM

[1470] D. M. Lytvtnenko, Yu. V. Slyusarenko, and A. I. Kirdin. The Thomas– Fermi model in the theory of systems of charged particles above the surface of liquid dielectrics. Journal of Mathematical Physics, 53(10): 103302, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103302_s1.

Bebiano:2012:TEM

[1471] N. Bebiano, J. da Providência, and J. P. da Providência. Tsallis entropies and matrix trace inequalities in quantum statistical mechanics. Journal of Mathematical Physics, 53(10):103303, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p103303_s1.

Bergeron:2012:GFG

[1472] H. Bergeron, E. M. F. Curado, J. P. Gazeau, and Ligia M. C. S. Rodrigues. Generating functions for generalized binomial distributions. Journal of Mathematical Physics, 53(10):103304, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103304_s1.

Takahashi:2012:NBM

[1473] Yuta Takahashi and Makoto Katori. Noncolliding Brownian motion with drift and time-dependent Stieltjes–Wigert determinantal point process. Journal of Mathematical Physics, 53(10):103305, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103305_s1.

Huh:2012:GEN

[1474] Hyungjin Huh and Ho Lee. Global existence and nonrelativistic limit for the Vlasov–Maxwell–Chern–Simons system. Journal of Mathematical Physics, 53(10):103306, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p103306_s1.

Dai:2012:SAS

[1475] Meifeng Dai and Jie Liu. Scaling of average sending time on weighted Koch networks. Journal of Mathematical Physics, 53(10):103501, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p103501_s1.

Szabo:2012:MMS

[1476] Richard J. Szabo and Miguel Tierz. Matrix models and stochastic growth in Donaldson–Thomas theory. Journal of Mathematical Physics, 53(10): 103502, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103502_s1.

Cai:2012:QHM

[1477] Zhenning Cai, Yuwei Fan, Ruo Li, Tiao Lu, and Yanli Wang. Quantum hydrodynamic model by moment closure of Wigner equation. Journal of Mathematical Physics, 53(10):103503, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i10/p103503_s1.

Garcia-Pelayo:2012:ESI

[1478] Ricardo García-Pelayo. Exact solutions for isotropic random flights in odd dimensions. Journal of Mathematical Physics, 53(10):103504, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p103504_s1.

Wei:2012:IEE

[1479] Jinbo Wei and Xianwen Zhang. The inelastic Enskog equation with external force. Journal of Mathematical Physics, 53(10):103505, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p103505_s1.

Gilkey:2012:GHT

[1480] P. Gilkey and R. J. Miatello. Growth of heat trace coefficients for locally symmetric spaces. Journal of Mathematical Physics, 53(10):103506, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i10/p103506_s1.

Muche:2012:WST

[1481] Lutz Muche and John M. Nieminen. Wigner surmises and the twodimensional Poisson–Voronoi tessellation. Journal of Mathematical Physics, 53(10):103507, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p103507_s1.

Tam:2012:ISA

[1482] Hon-Wah Tam and Yufeng Zhang. An integrable system and associated integrable models as well as Hamiltonian structures. Journal of Mathematical Physics, 53(10):103508, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p103508_s1.

Bender:2012:AAB

[1483] Carl M. Bender and Sarben Sarkar. Asymptotic analysis of the Boltzmann equation for dark matter relics. Journal of Mathematical Physics, 53(10):103509, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103509_s1.

Batat:2012:LOG

[1484] W. Batat, P. M. Gadea, and J. A. Oubiña. The Lorentzian oscillator group as a geodesic orbit space. Journal of Mathematical Physics, 53(10): 103510, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103510_s1.

Kiselev:2012:GDG

[1485] A. V. Kiselev and A. O. Krutov. Gardner’s deformations of the graded Korteweg–de Vries equations revisited. Journal of Mathematical Physics, 53(10):103511, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103511_s1.

Karpeshina:2012:ESP

[1486] Yulia Karpeshina and Roman Shterenberg. Extended states for polyharmonic operators with quasi-periodic potentials in dimension two. Journal of Mathematical Physics, 53(10):103512, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p103512_s1.

Gonzales:2012:PEL

[1487] Marcelo Gonzales, Kevin Iga, Sadi Khodaee, and Francesco Toppan. Pure and entangled N = 4 linear supermultiplets and their onedimensional sigma-models. Journal of Mathematical Physics, 53(10): 103513, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103513_s1.

Zak:2012:IOF

[1488] J. Zak. Inversion operators in finite phase plane. Journal of Mathematical Physics, 53(10):103514, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p103514_s1.

Buffenoir:2012:UVI

[1489] E. Buffenoir, Ph. Roche, and V. Terras. Universal vertex-IRF transformation for quantum affine algebras. Journal of Mathematical Physics, 53(10):103515, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103515_s1.

Moncrief:2012:MSC

[1490] Vincent Moncrief, Antonella Marini, and Rachel Maitra. Modified semiclassical methods for nonlinear quantum oscillations problems. Journal of Mathematical Physics, 53(10):103516, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i10/p103516_s1.

Hogner:2012:QKF

[1491] Moritz Högner. Quaternion–Kähler four-manifolds and Przanowski’s function. Journal of Mathematical Physics, 53(10):103517, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103517_s1.

Khodaee:2012:CSD

[1492] Sadi Khodaee and Francesco Toppan. Critical scaling dimension of Dmodule representations of N = 4, 7, 8 superconformal algebras and constraints on superconformal mechanics. Journal of Mathematical Physics, 53(10):103518, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103518_s1.

Schulze-Halberg:2012:DTD

[1493] Axel Schulze-Halberg. Darboux transformations for (1 + 2)-dimensional Fokker–Planck equations with constant diffusion matrix. Journal of Mathematical Physics, 53(10):103519, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i10/p103519_s1.

Velazquez-Arcos:2012:FES

[1494] J. M. Velázquez-Arcos. Fredholm’s equations for subwavelength focusing. Journal of Mathematical Physics, 53(10):103520, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103520_s1.

Prykarpatsky:2012:DAB

[1495] Yarema A. Prykarpatsky, Orest D. Artemovych, Maxim V. Pavlov, and Anatoliy K. Prykarpatsky. Differential-algebraic and bi-Hamiltonian integrability analysis of the Riemann hierarchy revisited. Journal of Mathematical Physics, 53(10):103521, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p103521_s1.

Fang:2012:ASC

[1496] Li Fang and Zhenhua Guo. Analytical solutions to a class of non-Newtonian fluids with free boundaries. Journal of Mathematical Physics, 53(10):103701, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103701_s1.

Cardone:2012:AES

[1497] G. Cardone, R. Fares, and G. P. Panasenko. Asymptotic expansion of the solution of the steady Stokes equation with variable viscosity in a twodimensional tube structure. Journal of Mathematical Physics, 53(10): 103702, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103702_s1.

Ruan:2012:CRV

[1498] Lizhi Ruan and Haiyan Yin. Convergence rates of vanishing diffusion limit on nonlinear hyperbolic system with damping and diffusion. Journal of Mathematical Physics, 53(10):103703, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p103703_s1.

Kumar:2012:ASS

[1499] Hitender Kumar, Anand Malik, and Fakir Chand. Analytical spatiotemporal soliton solutions to (3 + 1)-dimensional cubic-quintic nonlinear Schrödinger equation with distributed coefficients. Journal of Mathematical Physics, 53(10):103704, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p103704_s1.

Li:2012:RAO

[1500] Jingzhi Li, Hongyu Liu, Hongpeng Sun, and Jun Zou. Reconstructing acoustic obstacles by planar and cylindrical waves. Journal of Mathematical Physics, 53(10):103705, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i10/p103705_s1.

Hsini:2012:LTT

[1501] M. Hsini. Liouville type theorems for nonlinear elliptic equations involving operator in divergence form. Journal of Mathematical Physics, 53(10):103706, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i10/p103706_s1.

Egli:2012:SHM

[1502] Daniel Egli and Zhou Gang. Some Hamiltonian models of friction II. Journal of Mathematical Physics, 53(10):103707, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103707_s1.

Wu:2012:TIH

[1503] Derchyi Wu. A twisted integrable hierarchy with D_{2} symmetry. Journal of Mathematical Physics, 53(10):103708, October 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i10/p103708_s1.

Chen:2012:WBP

[1504] Caixia Chen and Yonghua Yan. On the wave breaking phenomena for the generalized periodic two-component Dullin–Gottwald–Holm system. Journal of Mathematical Physics, 53(10):103709, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103709_s1.

Wu:2012:MCK

[1505] Zhiwei Wu. On the modified constrained Kadomtsev–Petviashvili equations. Journal of Mathematical Physics, 53(10):103710, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p103710_s1.

Hayashi:2012:QNK

[1506] Nakao Hayashi and Pavel I. Naumkin. Quadratic nonlinear Klein–Gordon equation in one dimension. Journal of Mathematical Physics, 53(10):

103711, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i10/p103711_s1.

Yildirim:2012:RHP

[1507] Ahmet Yildirim. Retraction: “Homotopy perturbation method to obtain exact special solutions with solitary patterns for Boussinesq-B(m, n) equations with fully nonlinear dispersion” [J. Math. Phys. 50, 023510 (2009)]. Journal of Mathematical Physics, 53(10):109901, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i10/
p109901_s1.

Mattis:2012:PNC

[1508] Daniel C. Mattis. Publisher’s note: “A contribution to the many-fermion problem” [J. Math. Phys. 53, 095212 (2012)]. Journal of Mathematical Physics, 53(10):109902, October 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i10/p109902_s1. See [1422].

Doering:2012:ISI

[1509] Charles R. Doering, Igor Kukavica, and Edriss S. Titi. Introduction to special issue: Incompressible fluids, turbulence and mixing. Journal of Mathematical Physics, 53(11):115501, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115501_s1.

Chae:2012:LRI

[1510] Dongho Chae and Jiahong Wu. Logarithmically regularized inviscid models in borderline Sobolev spaces. Journal of Mathematical Physics, 53(11):115601, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115601_s1.

Friedlander:2012:SIP

[1511] Susan Friedlander, Francisco Gancedo, Weiran Sun, and Vlad Vicol. On a singular incompressible porous media equation. Journal of Mathematical Physics, 53(11):115602, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i11/p115602_s1.

Farhat:2012:LTB

[1512] Aseel Farhat, R. Lee Panetta, Edriss S. Titi, and Mohammed Ziane. Long-time behavior of a two-layer model of baroclinic quasi-geostrophic turbulence. Journal of Mathematical Physics, 53(11):115603, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i11/p115603_s1.

Jia:2012:LTU

[1513] Hao Jia, Gregory Seregin, and Vladimír Sverák. Liouville theorems in unbounded domains for the time-dependent Stokes system. Journal of Mathematical Physics, 53(11):115604, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115604_s1.

Iyer:2012:CSR

[1514] Gautam Iyer, Robert L. Pego, and Arghir Zarnescu. Coercivity and stability results for an extended Navier–Stokes system. Journal of Mathematical Physics, 53(11):115605, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i11/p115605_s1.

Saut:2012:WPL

[1515] Jean-Claude Saut and Li Xu. Well-posedness on large time for a modified full dispersion system of surface waves. Journal of Mathematical Physics, 53(11):115606, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115606_s1.

Ohkitani:2012:FSD

[1516] Koji Ohkitani. The Fefferman-Stein decomposition for the Constantin– Lax–Majda equation: Regularity criteria for inviscid fluid dynamics revisited. Journal of Mathematical Physics, 53(11):115607, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i11/
p115607_s1.

Gibbon:2012:CRS

[1517] J. D. Gibbon. Conditional regularity of solutions of the three-dimensional Navier–Stokes equations and implications for intermittency. Journal of Mathematical Physics, 53(11):115608, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115608_s1.

Kiselev:2012:BCE

[1518] Alexander Kiselev and Lenya Ryzhik. Biomixing by chemotaxis and efficiency of biological reactions: The critical reaction case. Journal of Mathematical Physics, 53(11):115609, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115609_s1.

Sun:2012:SCS

[1519] Hui Sun, David Uminsky, and Andrea L. Bertozzi. Stability and clustering of self-similar solutions of aggregation equations. Journal of Mathematical Physics, 53(11):115610, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i11/p115610_s1.

Lunasin:2012:OMO

[1520] Evelyn Lunasin, Zhi Lin, Alexei Novikov, Anna Mazzucato, and Charles R. Doering. Optimal mixing and optimal stirring for fixed energy, fixed power, or fixed palenstrophy flows. Journal of Mathematical Physics, 53(11):115611, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i11/p115611_s1. See erratum [1934].

Saut:2012:IBV

[1521] Jean-Claude Saut, Roger Temam, and Chuntian Wang. An initial and boundary-value problem for the Zakharov–Kuznetsov equation in a bounded domain. Journal of Mathematical Physics, 53(11):115612, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i11/p115612_s1.

Dascaliuc:2012:VSC

[1522] R. Dascaliuc and Z. Grujić. Vortex stretching and criticality for the threedimensional Navier–Stokes equations. Journal of Mathematical Physics, 53(11):115613, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115613_s1.

Holm:2012:MTM

[1523] Darryl D. Holm and Cesare Tronci. Multiscale turbulence models based on convected fluid microstructure. Journal of Mathematical Physics, 53(11):115614, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115614_s1.

E:2012:SCL

[1524] Weinan E and Jianfeng Lu. Stability and the continuum limit of the spin-polarized Thomas–Fermi–Dirac–von Weizsäcker model. Journal of Mathematical Physics, 53(11):115615, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115615_s1.

Zheng:2012:ENC

[1525] Likun Zheng, Meng Chen, and Qing Nie. External noise control in inherently stochastic biological systems. Journal of Mathematical Physics, 53(11):115616, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115616_s1.

Xiao:2012:NBC

[1526] Yuelong Xiao and Zhouping Xin. A new boundary condition for the three-dimensional Navier–Stokes equation and the vanishing viscosity limit. Journal of Mathematical Physics, 53(11):115617, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i11/
p115617_s1.

Robinson:2012:LBB

[1527] James C. Robinson, Witold Sadowski, and Ricardo P. Silva. Lower bounds on blow up solutions of the three-dimensional Navier–Stokes equations in homogeneous Sobolev spaces. Journal of Mathematical Physics, 53(11):115618, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i11/p115618_s1.

Berestycki:2012:SSO

[1528] Henri Berestycki and Grégoire Nadin. Spreading speeds for onedimensional monostable reaction-diffusion equations. Journal of Mathematical Physics, 53(11):115619, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i11/p115619_s1.

Cheskidov:2012:IPS

[1529] A. Cheskidov and R. Shvydkoy. Ill-posedness for subcritical hyperdissipative Navier–Stokes equations in the largest critical spaces. Journal of Mathematical Physics, 53(11):115620, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115620_s1.

Bardos:2012:VED

[1530] Claude Bardos and Anne Nouri. A Vlasov equation with Dirac potential used in fusion plasmas. Journal of Mathematical Physics, 53(11):115621, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i11/p115621_s1.

Castro:2012:FTS

[1531] Angel Castro, Diego Córdoba, Charles Fefferman, Francisco Gancedo, and Javier Gómez-Serrano. Finite time singularities for water waves with surface tension. Journal of Mathematical Physics, 53(11):115622, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i11/p115622_s1.

Foias:2012:DFT

[1532] Ciprian Foias, Michael S. Jolly, Rostyslav Kravchenko, and Edriss S. Titi. A determining form for the two-dimensional Navier–Stokes equations: The Fourier modes case. Journal of Mathematical Physics, 53 (11):115623, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115623_s1.

Ignatova:2012:WPF

[1533] Mihaela Ignatova, Igor Kukavica, Irena Lasiecka, and Amjad Tuffaha. On well-posedness for a free boundary fluid-structure model. Journal of Mathematical Physics, 53(11):115624, November 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i11/p115624_s1.

Jang:2012:WIP

[1534] Juhi Jang and Nader Masmoudi. Well and ill-posedness for compressible Euler equations with vacuum. Journal of Mathematical Physics, 53(11):115625, November 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i11/p115625_s1.

Hollands:2012:EWC

[1535] Stefan Hollands and Karl-Henning Rehren. Editorial: Wedge causal manifolds — an unfinished work of Hans-Jürgen Borchers. Journal of Mathematical Physics, 53(12):120401, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p120401_s1.

Vourdas:2012:POT

[1536] A. Vourdas. Partial order and a T_{0}-topology in a set of finite quantum systems. Journal of Mathematical Physics, 53(12):122101, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p122101_s1.

Friedmann:2012:GTD

[1537] Tamar Friedmann and C. R. Hagen. Group-theoretical derivation of angular momentum eigenvalues in spaces of arbitrary dimensions. Journal of Mathematical Physics, 53(12):122102, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122102_s1.

Nikitin:2012:ISS

[1538] A. G. Nikitin. Integrability and supersymmetry of Schrödinger–Pauli equations for neutral particles. Journal of Mathematical Physics, 53 (12):122103, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122103_s1.

Bessa:2012:QLE

[1539] V. H. L. Bessa and I. Guedes. The quantum Lane–Emden-type Kanai– Caldirola oscillators. Journal of Mathematical Physics, 53(12):122104, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p122104_s1.

Smilga:2012:SHK

[1540] A. V. Smilga. Supercharges in the hyper–Kähler with torsion supersymmetric sigma models. Journal of Mathematical Physics, 53(12):122105, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p122105_s1.

Filgueiras:2012:NQD

[1541] C. Filgueiras, E. O. Silva, and F. M. Andrade. Nonrelativistic quantum dynamics on a cone with and without a constraining potential. Journal of Mathematical Physics, 53(12):122106, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122106_s1.

Elci:2012:CPT

[1542] A. Elçi. Characteristic particle trajectories for an eigenfunction. Journal of Mathematical Physics, 53(12):122107, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122107_s1.

Choi:2012:RBP

[1543] Seongsoo Choi, J. W. Chung, and Kwang S. Kim. Relation between primes and nontrivial zeros in the Riemann hypothesis; Legendre polynomials, modified zeta function and Schrödinger equation. Journal of Mathematical Physics, 53(12):122108, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122108_s1. See

note [1666].

Drabkin:2012:TRK

[1544] Maxim Drabkin, Werner Kirsch, and Hermann Schulz-Baldes. Transport in the random Kronig–Penney model. Journal of Mathematical Physics, 53(12):122109, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122109_s1.

Chau:2012:ETB

[1545] H. F. Chau, Chi-Hang Fred Fung, Chi-Kwong Li, Edward Poon, and Nung-Sing Sze. Entanglement transformation between sets of bipartite pure quantum states using local operations. Journal of Mathematical Physics, 53(12):122201, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i12/p122201_s1.

McNulty:2012:IHM

[1546] Daniel McNulty and Stefan Weigert. Isolated Hadamard matrices from mutually unbiased product bases. Journal of Mathematical Physics, 53(12):122202, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122202_s1.

Jencova:2012:ECG

[1547] Anna Jencová. Extremality conditions for generalized channels. Journal of Mathematical Physics, 53(12):122203, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122203_s1.

Kim:2012:OES

[1548] Isaac H. Kim. Operator extension of strong subadditivity of entropy. Journal of Mathematical Physics, 53(12):122204, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/

p122204_s1. See remarks [2035].

Audenaert:2012:QSD

[1549] Koenraad M. R. Audenaert, Milán Mosonyi, and Frank Verstraete. Quantum state discrimination bounds for finite sample size. Journal of Mathematical Physics, 53(12):122205, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122205_s1.

Montanaro:2012:SAH

[1550] Ashley Montanaro. Some applications of hypercontractive inequalities in quantum information theory. Journal of Mathematical Physics, 53 (12):122206, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122206_s1.

Kuo:2012:UPA

[1551] Wan-Jung Kuo, Gregory Quiroz, Gerardo Andres Paz-Silva, and Daniel A. Lidar. Universality proof and analysis of generalized nested Uhrig dynamical decoupling. Journal of Mathematical Physics, 53 (12):122207, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122207_s1.

Hamhalter:2012:LMP

[1552] Jan Hamhalter. Linear maps preserving maximal deviation and the Jordan structure of quantum systems. Journal of Mathematical Physics, 53(12):122208, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122208_s1.

Lupo:2012:IMM

[1553] C. Lupo, S. Mancini, A. De Pasquale, P. Facchi, G. Florio, and S. Pascazio. Invariant measures on multimode quantum Gaussian states. Journal of Mathematical Physics, 53(12):122209, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p122209_s1.

Hofer-Szabo:2012:NLC

[1554] Gábor Hofer-Szabó and Péter Vecsernyés. Noncommuting local common causes for correlations violating the Clauser–Horne inequality. Journal of Mathematical Physics, 53(12):122301, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122301_s1.

Bernard:2012:FCS

[1555] Denis Bernard and Benjamin Doyon. Full counting statistics in the resonant-level model. Journal of Mathematical Physics, 53(12):122302, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p122302_s1.

Deriglazov:2012:NGM

[1556] A. A. Deriglazov, B. F. Rizzuti, G. P. Zamudio, and P. S. Castro. Non-Grassmann mechanical model of the Dirac equation. Journal of Mathematical Physics, 53(12):122303, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i12/p122303_s1.

Beisert:2012:CLC

[1557] Niklas Beisert and Florian Lücker. Construction of Lax connections by exponentiation. Journal of Mathematical Physics, 53(12):122304, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p122304_s1.

Vatandoost:2012:SNS

[1558] M. Vatandoost and Y. Bahrampour. Some necessary and sufficient conditions for admitting a continuous sphere order representation of two-dimensional space–times. Journal of Mathematical Physics, 53 (12):122501, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122501_s1.

Mora:2012:GPG

[1559] P. J. Mora, N. C. Tsamis, and R. P. Woodard. Graviton propagator in a general invariant gauge on de Sitter. Journal of Mathematical Physics, 53(12):122502, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122502_s1.

Bejancu:2012:CDW

[1560] Aurel Bejancu, Constantin Calin, and Hani Reda Farran. Classification of 5 d warped spaces with cosmological constant. Journal of Mathematical Physics, 53(12):122503, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i12/p122503_s1.

Brito:2012:CSS

[1561] I. Brito, J. Carot, F. C. Mena, and E. G. L. R. Vaz. Cylindrically symmetric static solutions of the Einstein field equations for elastic matter. Journal of Mathematical Physics, 53(12):122504, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p122504_s1.

Baumgartel:2012:CRD

[1562] Hellmut Baumgärtel. On a critical radiation density in the Friedmann equation. Journal of Mathematical Physics, 53(12):122505, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p122505_s1.

Tudoran:2012:CIV

[1563] Razvan M. Tudoran. Complete integrability versus symmetry. Journal of Mathematical Physics, 53(12):122701, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122701_s1.

Zhao:2012:PAB

[1564] Caidi Zhao. Pullback asymptotic behavior of solutions for a nonautonomous non-Newtonian fluid on two-dimensional unbounded domains. Journal of Mathematical Physics, 53(12):122702, December 2012.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p122702_s1.

Buzzi:2012:PSS

[1565] Claudio A. Buzzi and Jaume Llibre. On the periodic solutions of the static, spherically symmetric Einstein–Yang–Mills equations. Journal of Mathematical Physics, 53(12):122703, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p122703_s1.

Mahdi:2012:ANI

[1566] Adam Mahdi and Claudia Valls. Analytic non-integrability of the Suslov problem. Journal of Mathematical Physics, 53(12):122901, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p122901_s1.

Paraschiv:2012:CCH

[1567] Victor Paraschiv. Central configurations and homographic solutions for the quasihomogeneous N-body problem. Journal of Mathematical Physics, 53(12):122902, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i12/p122902_s1.

Llibre:2012:POS

[1568] Jaume Llibre and Ammar Makhlouf. Periodic orbits of the spatial anisotropic Manev problem. Journal of Mathematical Physics, 53 (12):122903, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p122903_s1.

Bru:2012:IFQ

[1569] J.-B. Bru and W. de Siqueira Pedra. Inhomogeneous Fermi and quantum spin systems on lattices. Journal of Mathematical Physics, 53(12):123301, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p123301_s1.

diMolfetta:2012:DTQ

[1570] G. di Molfetta and F. Debbasch. Discrete-time quantum walks: Continuous limit and symmetries. Journal of Mathematical Physics, 53 (12):123302, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123302_s1.

Vyas:2012:EGU

[1571] Manan Vyas and V. K. B. Kota. Embedded Gaussian unitary ensembles with U(Ω) ⊗ SU(r) embedding generated by random two-body interactions with SU(r) symmetry. Journal of Mathematical Physics, 53(12):123303, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123303_s1.

Genovese:2012:UBM

[1572] Giuseppe Genovese. Universality in bipartite mean field spin glasses. Journal of Mathematical Physics, 53(12):123304, December 2012. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123304_s1.

Peng:2012:IGC

[1573] Linyu Peng, Huafei Sun, and Guoquan Xu. Information geometric characterization of the complexity of fractional Brownian motions. Journal of Mathematical Physics, 53(12):123305, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123305_s1.

Jansen:2012:FBL

[1574] Sabine Jansen. Fermionic and bosonic Laughlin state on thick cylinders. Journal of Mathematical Physics, 53(12):123306, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123306_s1.

Wreszinski:2012:PMT

[1575] Walter F. Wreszinski. Progress in the mathematical theory of quantum disordered systems. Journal of Mathematical Physics, 53(12):123307, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p123307_s1.

Guo:2012:BCS

[1576] Zhengguang Guo, Liangbing Jin, and Lidiao Ni. Blow-up criteria of solutions for a modified two-component hyperelastic rod system. Journal of Mathematical Physics, 53(12):123501, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123501_s1.

Gay-Balmaz:2012:VMF

[1577] François Gay-Balmaz and Cesare Tronci. Vlasov moment flows and geodesics on the Jacobi group. Journal of Mathematical Physics, 53 (12):123502, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123502_s1.

Nordebo:2012:FII

[1578] S. Nordebo, M. Gustafsson, A. Khrennikov, B. Nilsson, and J. Toft. Fisher information for inverse problems and trace class operators. Journal of Mathematical Physics, 53(12):123503, December 2012. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123503_s1.

Wang:2012:IMG

[1579] Yunhu Wang and Yong Chen. Integrability of the modified generalised Vakhnenko equation. Journal of Mathematical Physics, 53(12):123504, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p123504_s1.

Cautis:2012:LRQ

[1580] Sabin Cautis and Anthony Licata. Loop realizations of quantum affine algebras. Journal of Mathematical Physics, 53(12):123505, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p123505_s1.

Lechtenfeld:2012:ISS

[1581] Olaf Lechtenfeld and Alexander D. Popov. Instantons on the six-sphere and twistors. Journal of Mathematical Physics, 53(12):123506, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p123506_s1.

Batalin:2012:TBD

[1582] Igor A. Batalin and Klaus Bering. A triplectic bi-Darboux theorem and para-hypercomplex geometry. Journal of Mathematical Physics, 53(12):123507, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123507_s1.

Figotin:2012:DPS

[1583] Alexander Figotin and Aaron Welters. Dissipative properties of systems composed of high-loss and lossless components. Journal of Mathematical Physics, 53(12):123508, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v53/i12/p123508_s1.

Chirvasitu:2012:FUG

[1584] Alexandru Chirvasitu. Free unitary groups are (almost) simple. Journal of Mathematical Physics, 53(12):123509, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123509_s1.

Bagarello:2012:WCR

[1585] Fabio Bagarello, Atsushi Inoue, and Camillo Trapani. Weak commutation relations of unbounded operators: Nonlinear extensions. Journal of Mathematical Physics, 53(12):123510, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123510_s1.

Torrome:2012:FPC

[1586] Ricardo Gallego Torromé, Paolo Piccione, and Henrique Vitório. On Fermat’s principle for causal curves in time oriented Finsler spacetimes. Journal of Mathematical Physics, 53(12):123511, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123511_s1.

Hallnas:2012:KFB

[1587] Martin Hallnäs and Simon Ruijsenaars. Kernel functions and Bäcklund transformations for relativistic Calogero–Moser and Toda systems. Journal of Mathematical Physics, 53(12):123512, December 2012. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123512_s1.

Wu:2012:AGS

[1588] Lihua Wu, Guoliang He, and Xianguo Geng. Algebro-geometric solutions to the modified Sawada–Kotera hierarchy. Journal of Mathematical Physics, 53(12):123513, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i12/p123513_s1.

Benaissa:2012:GEE

[1589] Abbes Benaissa, Abdelkader Benaissa, and Salim. A. Messaoudi. Global existence and energy decay of solutions for the wave equation with a time varying delay term in the weakly nonlinear internal feedbacks. Journal of Mathematical Physics, 53(12):123514, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123514_s1.

Bihlo:2012:CGC

[1590] Alexander Bihlo, Elsa Dos Santos Cardoso-Bihlo, and Roman O. Popovych. Complete group classification of a class of nonlinear wave equations. Journal of Mathematical Physics, 53(12):123515, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p123515_s1.

Benzair:2012:PID

[1591] H. Benzair, T. Boudjedaa, and M. Merad. Path integral for Dirac oscillator with generalized uncertainty principle. Journal of Mathematical Physics, 53(12):123516, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v53/i12/p123516_s1.

Bouchitte:2012:TWR

[1592] Guy Bouchitté, Luísa Mascarenhas, and Luís Trabucho. Thin waveguides with Robin boundary conditions. Journal of Mathematical Physics, 53(12):123517, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123517_s1.

Pouladsaz:2012:CTD

[1593] D. Pouladsaz. Classification of two-dimensional quantum systems with Z_{3}-graded topological symmetries. Journal of Mathematical Physics, 53(12):123518, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123518_s1.

Amore:2012:PAS

[1594] Paolo Amore. A perturbative approach to the spectral zeta functions of strings, drums, and quantum billiards. Journal of Mathematical Physics, 53(12):123519, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123519_s1.

Costa:2012:FWD

[1595] F. S. Costa and E. Capelas de Oliveira. Fractional wave-diffusion equation with periodic conditions. Journal of Mathematical Physics, 53(12):123520, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123520_s1.

Manko:2012:SOS

[1596] S. S. Man’ko. Schrödinger operators on star graphs with singularly scaled potentials supported near the vertices. Journal of Mathematical Physics, 53(12):123521, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123521_s1.

Serov:2012:DIS

[1597] Valery Serov, Markus Harju, and Georgios Fotopoulos. Direct and inverse scattering for nonlinear Schrödinger equation in 2D. Journal of Mathematical Physics, 53(12):123522, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123522_s1.

Deschout:2012:CBA

[1598] K. Deschout and A. B. J. Kuijlaars. Critical behavior in Angelesco ensembles. Journal of Mathematical Physics, 53(12):123523, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123523_s1.

Cheng:2012:MSS

[1599] Rong Cheng, Jianhua Hu, and Zhiyong Wang. Multiple solutions for semilinear resonance elliptic problems with nonconstant coefficients at infinity. Journal of Mathematical Physics, 53(12):123524, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123524_s1.

Brothier:2012:RTC

[1600] Arnaud Brothier, Michael Hartglass, and David Penneys. Rigid C^{∗}-tensor categories of bimodules over interpolated free group factors. Journal of Mathematical Physics, 53(12):123525, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123525_s1.

Curto:2012:CHP

[1601] Raúl E. Curto, Sang Hoon Lee, and Jasang Yoon. Completion of Hankel partial contractions of extremal type. Journal of Mathematical Physics, 53(12):123526, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123526_s1.

Dehghani:2012:MUC

[1602] A. Dehghani, H. Fakhri, and B. Mojaveri. The minimum-uncertainty coherent states for Landau levels. Journal of Mathematical Physics, 53(12):123527, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123527_s1.

Pusztai:2012:MSH

[1603] B. G. Pusztai. On the r-matrix structure of the hyperbolic BC_{n} Sutherland model. Journal of Mathematical Physics, 53(12):123528, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p123528_s1.

Dauge:2012:PWC

[1604] Monique Dauge and Nicolas Raymond. Plane waveguides with corners in the small angle limit. Journal of Mathematical Physics, 53(12):123529, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p123529_s1.

Smirnov:2012:CET

[1605] Yury G. Smirnov and Dmitry V. Valovik. Coupled electromagnetic TE– TM wave propagation in a layer with Kerr nonlinearity. Journal of Mathematical Physics, 53(12):123530, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123530_s1.

Ivanov:2012:DSQ

[1606] Rossen I. Ivanov and Tony Lyons. Dark solitons of the Qiao’s hierarchy. Journal of Mathematical Physics, 53(12):123701, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i12/
p123701_s1.

Mikelic:2012:TDB

[1607] Andro Mikelić and Mary F. Wheeler. Theory of the dynamic Biot– Allard equations and their link to the quasi-static Biot system. Journal of Mathematical Physics, 53(12):123702, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p123702_s1.

Kengne:2012:ESG

[1608] E. Kengne, A. Lakhssassi, R. Vaillancourt, and Wu-Ming Liu. Exact solutions for generalized variable-coefficients Ginzburg–Landau equation: Application to Bose–Einstein condensates with multi-body interatomic interactions. Journal of Mathematical Physics, 53(12):123703, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v53/i12/p123703_s1.

Jiang:2012:GSA

[1609] Jie Jiang and Songmu Zheng. Global solvability and asymptotic behavior of a free boundary problem for the one-dimensional viscous radiative and reactive gas. Journal of Mathematical Physics, 53(12):123704, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i12/p123704_s1.

Benguria:2012:RRM

[1610] Rafael D. Benguria, M. Cristina Depassier, and Michael Loss. Rigorous results for the minimal speed of Kolmogorov–Petrovskii–Piscounov monotonic fronts with a cutoff. Journal of Mathematical Physics, 53 (12):123705, December 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v53/i12/p123705_s1.

Abraham:2012:PNG

[1611] D. B. Abraham. Publisher’s note: “Generalised Wick theorem and the Fredholm alternative” [ J. Math. Phys. 53, 095224 (2012)]. Journal of Mathematical Physics, 53(12):129901, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p129901_s1. See

[1434].

Felinska:2012:EWP

[1612] Ewa Felińska, Zbigniew Jaskólski, and Michal Kosztolowicz. Erratum: “Whittaker pairs for the Virasoro algebra and the Gaiotto–Bonelli– Maruyoshi–Tanzini states” [J. Math. Phys. 53, 033504 (2012)]. Journal of Mathematical Physics, 53(12):129902, December 2012. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v53/i12/p129902_s1. See

[1198].

Nachtergaele:2013:AJM

[1613] Bruno Nachtergaele. Announcement: Journal of Mathematical Physics introduces two new sections: Many-Body and Condensed Matter Physics and Representation Theory and Algebraic Methods. Journal of Mathematical Physics, 54(1):010201, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i1/p010201_s1.

Dahl:2013:CRN

[1614] Matias F. Dahl. Characterisation and representation of non-dissipative electromagnetic medium with two Lorentz null cones. Journal of Mathematical Physics, 54(1):011501, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i1/p011501_s1.

Yamazaki:2013:RRC

[1615] Kazuo Yamazaki. Remarks on the regularity criteria of generalized MHD and Navier–Stokes systems. Journal of Mathematical Physics, 54(1): 011502, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p011502_s1.

Nguyen:2013:GEC

[1616] Nghiem V. Nguyen, Rushun Tian, Bernard Deconinck, and Natalie Sheils. Global existence for a coupled system of Schrödinger equations with power-type nonlinearities. Journal of Mathematical Physics, 54(1): 011503, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p011503_s1.

Deng:2013:NSS

[1617] Yinbin Deng, Shuangjie Peng, and Jixiu Wang. Nodal soliton solutions for quasilinear Schrödinger equations with critical exponent. Journal of Mathematical Physics, 54(1):011504, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i1/p011504_s1.

Zhang:2013:ESS

[1618] Ruifeng Zhang, Hong Wang, and Rentao Liu. The existence of steady solutions for a class of Schrödinger equations in nonlinear optical lattices. Journal of Mathematical Physics, 54(1):011505, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p011505_s1.

Kornbluth:2013:UCS

[1619] Mordechai Kornbluth and Fredy Zypman. Uncertainties of coherent states for a generalized supersymmetric annihilation operator. Journal of Mathematical Physics, 54(1):012101, January 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i1/p012101_s1.

Borisov:2013:TRS

[1620] D. Borisov, P. Exner, and A. Golovina. Tunneling resonances in systems without a classical trapping. Journal of Mathematical Physics, 54(1): 012102, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p012102_s1.

Bayin:2013:TFS

[1621] Selçuk S. Bayin. Time fractional Schrödinger equation: Fox’s Hfunctions and the effective potential. Journal of Mathematical Physics, 54(1):012103, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i1/p012103_s1.

Roychoudhury:2013:NHO

[1622] R. Roychoudhury, B. Roy, and P. P. Dube. Non-Hermitian oscillator and R-deformed Heisenberg algebra. Journal of Mathematical Physics, 54(1): 012104, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p012104_s1.

Braunss:2013:QDP

[1623] G. Braunss. Quantum dynamics in phase space: Moyal trajectories 2. Journal of Mathematical Physics, 54(1):012105, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p012105_s1.

Punosevac:2013:AMU

[1624] Predrag Punosevac and Sam L. Robinson. Asymptotically minimal uncertainty states for time-dependent oscillators. Journal of Mathematical Physics, 54(1):012106, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i1/p012106_s1.

Bardavelidze:2013:TDS

[1625] M. S. Bardavelidze, F. Cannata, M. V. Ioffe, and D. N. Nishnianidze. Three-dimensional shape invariant non-separable model with equidistant spectrum. Journal of Mathematical Physics, 54(1):012107, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p012107_s1.

Aktosun:2013:HEA

[1626] Tuncay Aktosun and Ricardo Weder. High-energy analysis and Levinson’s theorem for the selfadjoint matrix Schrödinger operator on the half line. Journal of Mathematical Physics, 54(1):012108, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p012108_s1.

Meng:2013:GKP

[1627] Guowu Meng. Generalized Kepler problems. I. Without magnetic charges. Journal of Mathematical Physics, 54(1):012109, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p012109_s1.

Hahne:2013:VQD

[1628] G. E. Hahne. Variance of the quantum dwell time for a nonrelativistic particle. Journal of Mathematical Physics, 54(1):012110, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p012110_s1.

Luchko:2013:FSE

[1629] Yuri Luchko. Fractional Schrödinger equation for a particle moving in a potential well. Journal of Mathematical Physics, 54(1):012111, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p012111_s1.

Tarasov:2013:URN

[1630] Vasily E. Tarasov. Uncertainty relation for non-Hamiltonian quantum systems. Journal of Mathematical Physics, 54(1):012112, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p012112_s1.

Buerschaper:2013:HTT

[1631] Oliver Buerschaper, Juan Martín Mombelli, Matthias Christandl, and Miguel Aguado. A hierarchy of topological tensor network states. Journal of Mathematical Physics, 54(1):012201, January 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i1/p012201_s1.

He:2013:GTC

[1632] Yang-Hui He and John McKay. N = 2 gauge theories: Congruence subgroups, coset graphs, and modular surfaces. Journal of Mathematical Physics, 54(1):012301, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i1/p012301_s1.

Chamseddine:2013:DRC

[1633] Ali H. Chamseddine, J. Fröhlich, B. Schubnel, and D. Wyler. Dimensional reduction without continuous extra dimensions. Journal of Mathematical Physics, 54(1):012302, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i1/p012302_s1.

Falconi:2013:CLN

[1634] Marco Falconi. Classical limit of the Nelson model with cutoff. Journal of Mathematical Physics, 54(1):012303, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i1/p012303_s1.

Jantzen:2013:NPT

[1635] Bernd Jantzen. New proofs for the two Barnes lemmas and an additional lemma. Journal of Mathematical Physics, 54(1):012304, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p012304_s1.

Wang:2013:NRS

[1636] Jian Wang and Yong Wang. Noncommutative residue and sub-Dirac operators for foliations. Journal of Mathematical Physics, 54(1):012501, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p012501_s1.

Brunelli:2013:HIT

[1637] J. C. Brunelli and S. Sakovich. Hamiltonian integrability of twocomponent short pulse equations. Journal of Mathematical Physics, 54(1):012701, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i1/p012701_s1.

Komech:2013:ACS

[1638] A. I. Komech and A. E. Merzon. On asymptotic completeness of scattering in the nonlinear Lamb system, II. Journal of Mathematical Physics, 54(1):012702, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i1/p012702_s1.

Calogero:2013:SMB

[1639] Francesco Calogero. A solvable many-body problem, its equilibria, and a second-order ordinary differential equation whose general solution is polynomial. Journal of Mathematical Physics, 54(1):012703, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p012703_s1.

Kihara:2013:EMS

[1640] Hironobu Kihara. Equations of motion solved by the Cremmer–Scherk configuration on even-dimensional spheres. Journal of Mathematical Physics, 54(1):012901, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i1/p012901_s1.

Vargas-Paredes:2013:LWF

[1641] Alfredo A. Vargas-Paredes, Mauro M. Doria, and José Abdala Helayël Neto. The Lichnerowicz–Weitzenböck formula and superconductivity. Journal of Mathematical Physics, 54(1):013101, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013101_s1.

Baez:2013:NTM

[1642] John C. Baez and Brendan Fong. A Noether theorem for Markov processes. Journal of Mathematical Physics, 54(1):013301, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p013301_s1.

Adami:2013:OAS

[1643] Riccardo Adami, Diego Noja, and Cecilia Ortoleva. Orbital and asymptotic stability for standing waves of a nonlinear Schrödinger equation with concentrated nonlinearity in dimension three. Journal of Mathematical Physics, 54(1):013501, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i1/p013501_s1.

Wang:2013:TTI

[1644] Wei Wang. On twistor transformations and invariant differential operator of simple Lie group G_{2(2)}. Journal of Mathematical Physics, 54(1): 013502, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p013502_s1.

Wang:2013:NSS

[1645] Quanxiang Wang, Zhiyue Zhang, and Quanyong Zhu. Numerical simulation of the stochastic damped improved Boussinesq equation. Journal of Mathematical Physics, 54(1):013503, January 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013503_s1.

Gaillard:2013:DDR

[1646] Pierre Gaillard. Degenerate determinant representation of solutions of the nonlinear Schrödinger equation, higher order Peregrine breathers and multi-rogue waves. Journal of Mathematical Physics, 54(1):013504, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p013504_s1.

Gould:2013:IRM

[1647] Mark D. Gould, Phillip S. Isaac, and Jason L. Werry. Invariants and reduced matrix elements associated with the Lie superalgebra gl(m|n). Journal of Mathematical Physics, 54(1):013505, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013505_s1.

Thirulogasanthar:2013:RSQ

[1648] K. Thirulogasanthar and S. Twareque Ali. Regular subspaces of a quaternionic Hilbert space from quaternionic Hermite polynomials and associated coherent states. Journal of Mathematical Physics, 54(1):013506, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p013506_s1.

Samann:2013:TCF

[1649] Christian Sämann and Martin Wolf. On twistors and conformal field theories from six dimensions. Journal of Mathematical Physics, 54(1): 013507, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p013507_s1.

Yahya:2013:BSS

[1650] W. A. Yahya and K. J. Oyewumi. Bound state solutions of the Dirac equation for the trigonometric and hyperbolic Scarf–Grosche potentials using the Nikiforov–Uvarov method. Journal of Mathematical Physics, 54(1):013508, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i1/p013508_s1.

Das:2013:HOG

[1651] Sanjit Das, Kartik Prabhu, and Sayan Kar. Higher order geometric flows on three-dimensional locally homogeneous spaces. Journal of Mathematical Physics, 54(1):013509, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i1/p013509_s1.

Huang:2013:GPI

[1652] Ding jiang Huang, Qin min Yang, and Shuigeng Zhou. Group properties and invariant solutions of a sixth-order thin film equation in viscous fluid. Journal of Mathematical Physics, 54(1):013510, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013510_s1.

Burciu:2013:FRE

[1653] Sebastian Burciu and Sonia Natale. Fusion rules of equivariantizations of fusion categories. Journal of Mathematical Physics, 54(1):013511, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p013511_s1.

Zarmi:2013:SSL

[1654] Yair Zarmi. Static solitons, Lorentz invariance, and a new perspective on the integrability of the sine Gordon equation in (1 + 2) dimensions. Journal of Mathematical Physics, 54(1):013512, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013512_s1.

Codello:2013:NLH

[1655] Alessandro Codello and Omar Zanusso. On the non-local heat kernel expansion. Journal of Mathematical Physics, 54(1):013513, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p013513_s1.

Laha:2013:SJS

[1656] U. Laha and J. Bhoi. Off-shell Jost solutions for Coulomb and Coulomblike interactions in all partial waves. Journal of Mathematical Physics, 54(1):013514, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i1/p013514_s1.

DeTurck:2013:GGM

[1657] Dennis DeTurck, Herman Gluck, Rafal Komendarczyk, Paul Melvin, Clayton Shonkwiler, and David Shea Vela-Vick. Generalized Gauss maps and integrals for three-component links: Toward higher helicities for magnetic fields and fluid flows. Journal of Mathematical Physics, 54(1): 013515, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p013515_s1.

Zhang:2013:DIP

[1658] Yufeng Zhang and Honwah Tam. Discussion on integrable properties for higher-dimensional variable-coefficient nonlinear partial differential equations. Journal of Mathematical Physics, 54(1):013516, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p013516_s1.

Chakrabarti:2013:NCS

[1659] Amitabha Chakrabarti, Anirban Chakraborti, and Esteban Guevara Hidalgo. New classes of spin chains from (SÔ_{(q)}(N), Sp̂_{(q)}(N)) Temperley– Lieb algebras: Data transmission and (q, N) parametrized entanglement entropies. Journal of Mathematical Physics, 54(1):013517, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p013517_s1.

Dabrowski:2013:CNT

[1660] Ludwik Dabrowski and Andrzej Sitarz. Curved noncommutative torus and Gauss–Bonnet. Journal of Mathematical Physics, 54(1):013518, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p013518_s1.

Zhang:2013:GES

[1661] Jingjun Zhang and Boling Guo. Global existence of solution for thermally radiative magnetohydrodynamic equations with the displacement current. Journal of Mathematical Physics, 54(1):013519, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/
p013519_s1.

Gade:2013:TPM

[1662] R. M. Gade. Tensor products of Uq′(ŝl(2)) math-modules and the big q^{2}-Jacobi function transform. Journal of Mathematical Physics, 54(1): 013520, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i1/p013520_s1.

Sakkaravarthi:2013:BSC

[1663] K. Sakkaravarthi and T. Kanna. Bright solitons in coherently coupled nonlinear Schrödinger equations with alternate signs of nonlinearities. Journal of Mathematical Physics, 54(1):013701, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i1/p013701_s1.

Wang:2013:SPL

[1664] Chunhua Wang and Jiangtao Wang. Solutions of perturbed p-Laplacian equations with critical nonlinearity. Journal of Mathematical Physics, 54(1):013702, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i1/p013702_s1.

Hawkins:2013:CCS

[1665] E. Hawkins and J. M. Schwarz. Comment on “On the consistency of solutions of the space fractional Schrödinger equation” [J. Math. Phys. 53, 042105 (2012)]. Journal of Mathematical Physics, 54(1):014101, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i1/p014101_s1. See [1222, 1409].

Choi:2013:PNR

[1666] Seongsoo Choi, J. W. Chung, and Kwang S. Kim. Publisher’s note: “Relation between primes and nontrivial zeros in the Riemann hypothesis; Legendre polynomials, modified zeta function and Schrödinger equation” [J. Math. Phys. 53, 122108 (2012)]. Journal of Mathematical Physics, 54(1):019901, January 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i1/p019901_s1. See [1543].

Ammari:2013:RTD

[1667] Habib Ammari, Emmanuel Bossy, Josselin Garnier, Wenjia Jing, and Laurent Seppecher. Radiative transfer and diffusion limits for wave field correlations in locally shifted random media. Journal of Mathematical Physics, 54(2):021501, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p021501_s1.

Huh:2013:ECS

[1668] Hyungjin Huh and Jinmyoung Seok. The equivalence of the Chern– Simons–Schrödinger equations and its self-dual system. Journal of Mathematical Physics, 54(2):021502, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i2/p021502_s1.

Graczyk:2013:MYW

[1669] Piotr Graczyk and Jacek Malecki. Multidimensional Yamada–Watanabe theorem and its applications to particle systems. Journal of Mathematical Physics, 54(2):021503, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p021503_s1.

Schmuck:2013:NPM

[1670] M. Schmuck. New porous medium Poisson–Nernst–Planck equations for strongly oscillating electric potentials. Journal of Mathematical Physics, 54(2):021504, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p021504_s1.

Silva:2013:LTD

[1671] Marcio Antonio Jorge Silva and To Fu Ma. Long-time dynamics for a class of Kirchhoff models with memory. Journal of Mathematical Physics, 54(2):021505, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p021505_s1.

Chen:2013:STC

[1672] Gui-Qiang Chen, Vaibhav Kukreja, and Hairong Yuan. Stability of transonic characteristic discontinuities in two-dimensional steady compressible Euler flows. Journal of Mathematical Physics, 54(2):021506, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p021506_s1.

Barbato:2013:DMT

[1673] David Barbato, Luigi Amedeo Bianchi, Franco Flandoli, and Francesco Morandin. A dyadic model on a tree. Journal of Mathematical Physics, 54(2):021507, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p021507_s1.

Liu:2013:GBS

[1674] Ming Liu, Chengming Bai, Mo-Lin Ge, and Naihuan Jing. Generalized Bell states and principal realization of the Yangian Y (∫↕_{N}). Journal of Mathematical Physics, 54(2):021701, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i2/p021701_s1.

Hartwig:2013:ETC

[1675] J. T. Hartwig and J. V. Stokman. Extended trigonometric Cherednik algebras and nonstationary Schrödinger equations with delta-potentials.

Journal of Mathematical Physics, 54(2):021702, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/p021702_s1.

Hagen:2013:GTD

[1676] C. R. Hagen. Group-theoretical derivation of Aharonov–Bohm phase shifts. Journal of Mathematical Physics, 54(2):021703, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/
p021703_s1.

Bagchi:2013:CCE

[1677] B. Bagchi, A. Banerjee, and A. Ganguly. CPT-conserved effective mass Hamiltonians through first and higher order charge operator C in a supersymmetric framework. Journal of Mathematical Physics, 54(2):022101, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i2/p022101_s1.

Chen:2013:DLS

[1678] Lin Chen and Dragomir Ž. – Doković. Dimensions, lengths, and separability in finite-dimensional quantum systems. Journal of Mathematical Physics, 54(2):022201, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p022201_s1.

Huckleberry:2013:BES

[1679] Alan Huckleberry, Marek Kuś, and Adam Sawicki. Bipartite entanglement, spherical actions, and geometry of local unitary orbits. Journal of Mathematical Physics, 54(2):022202, February 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i2/p022202_s1.

Kempf:2013:FCI

[1680] A. Kempf, A. Chatwin-Davies, and R. T. W. Martin. A fully covariant information-theoretic ultraviolet cutoff for scalar fields in expanding Friedmann Robertson Walker spacetimes. Journal of Mathematical Physics, 54(2):022301, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p022301_s1.

Lai:2013:SAO

[1681] Alan Lai and Kevin Teh. Spectral action for a one-parameter family of Dirac operators on SU(2) and SU(3). Journal of Mathematical Physics, 54(2):022302, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p022302_s1. See erratum [2021].

Todorov:2013:CFT

[1682] Ivan Todorov. Conformal field theories with infinitely many conservation laws. Journal of Mathematical Physics, 54(2):022303, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p022303_s1.

Huguet:2013:CIF

[1683] E. Huguet and J. Renaud. Conformally invariant formalism for the electromagnetic field with currents in Robertson–Walker spaces. Journal of Mathematical Physics, 54(2):022304, February 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i2/p022304_s1.

Gallot:2013:HDA

[1684] L. Gallot, E. Pilon, and F. Thuillier. Higher dimensional abelian Chern– Simons theories and their link invariants. Journal of Mathematical Physics, 54(2):022305, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p022305_s1.

Chekhov:2013:SWE

[1685] Leonid Chekhov, Bertrand Eynard, and Sylvain Ribault. Seiberg–Witten equations and non-commutative spectral curves in Liouville theory. Journal of Mathematical Physics, 54(2):022306, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/p022306_s1.

Grosse:2013:WRQ

[1686] Harald Grosse, Gandalf Lechner, Thomas Ludwig, and Rainer Verch. Wick rotation for quantum field theories on degenerate Moyal space(- time). Journal of Mathematical Physics, 54(2):022307, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/
p022307_s1.

Minguzzi:2013:RCT

[1687] E. Minguzzi. Relativistic Chasles’ theorem and the conjugacy classes of the inhomogeneous Lorentz group. Journal of Mathematical Physics, 54(2):022501, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p022501_s1.

Milson:2013:ICV

[1688] R. Milson, D. McNutt, and A. Coley. Invariant classification of vacuum pp-waves. Journal of Mathematical Physics, 54(2):022502, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p022502_s1.

Pantelidis:2013:DHM

[1689] Leonidas Pantelidis. Dynamics of the Heisenberg model and a theorem on stability. Journal of Mathematical Physics, 54(2):022701, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p022701_s1.

Escobar-Ruiz:2013:TCP

[1690] M. A. Escobar-Ruiz and A. V. Turbiner. Two charges on a plane in a magnetic field: Special trajectories. Journal of Mathematical Physics, 54(2):022901, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p022901_s1.

Carrisi:2013:ETCa

[1691] M. C. Carrisi and S. Pennisi. Extended thermodynamics of charged gases with many moments. Journal of Mathematical Physics, 54(2):023101, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i2/p023101_s1.

Wu:2013:EMN

[1692] Zuo-Bing Wu and Wen-Rui Hu. Effects of Marangoni numbers on thermocapillary drop migration: Constant for quasi-steady state? Journal of Mathematical Physics, 54(2):023102, February 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i2/p023102_s1.

Malyshev:2013:SOC

[1693] V. A. Malyshev. Self-organized circular flow of classical point particles. Journal of Mathematical Physics, 54(2):023301, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/p023301_s1.

Fraj:2013:SCD

[1694] Nizar Ben Fraj, Ismail Laraiedh, and Salem Omri. Supertransvectants, cohomology, and deformations. Journal of Mathematical Physics, 54(2):023501, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p023501_s1.

Ozemir:2013:SCV

[1695] C. Özemir and F. Güngör. Symmetry classification of variable coefficient cubic-quintic nonlinear Schrödinger equations. Journal of Mathematical Physics, 54(2):023502, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p023502_s1.

Aguayo:2013:CCS

[1696] José Aguayo, Miguel Nova, and Khodr Shamseddine. Characterization of compact and self-adjoint operators on free Banach spaces of countable type over the complex Levi-Civita field. Journal of Mathematical Physics, 54(2):023503, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p023503_s1.

Mimura:2013:LUS

[1697] N. Mimura, J. Satsuma, A. Ramani, and B. Grammaticos. Linearisable ultradiscrete systems with sign variables and the confinement of singularities. Journal of Mathematical Physics, 54(2):023504, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p023504_s1.

Han:2013:MCS

[1698] Jongmin Han and Seongtag Kim. Maxwell–Chern–Simons vortices on compact surfaces: Nonequivalence of the first and the second order equations. Journal of Mathematical Physics, 54(2):023505, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i2/p023505_s1.

Genest:2013:ADH

[1699] Vincent X. Genest, Luc Vinet, and Alexei Zhedanov. The algebra of dual-1 Hahn polynomials and the Clebsch–Gordan problem of ∫↕_{−1}(2). Journal of Mathematical Physics, 54(2):023506, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i2/p023506_s1.

McLellan:2013:LAC

[1700] B. D. K. McLellan. Localization in abelian Chern–Simons theory. Journal of Mathematical Physics, 54(2):023507, February 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i2/p023507_s1.

Douglas:2013:PAR

[1701] Andrew Douglas, Hubert de Guise, and Joe Repka. The Poincaré algebra in rank 3 simple Lie algebras. Journal of Mathematical Physics, 54(2):023508, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p023508_s1.

Bagarello:2013:DPF

[1702] F. Bagarello. Damping and pseudo-fermions. Journal of Mathematical Physics, 54(2):023509, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p023509_s1.

Gadella:2013:SGA

[1703] M. Gadella, J. Negro, G. P. Pronko, and M. Santander. Spectrum generating algebra for the continuous spectrum of a free particle in Lobachevski space. Journal of Mathematical Physics, 54(2):023510, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i2/p023510_s1.

Cheng:2013:SES

[1704] Jipeng Cheng and Jingsong He. On the squared eigenfunction symmetry of the Toda lattice hierarchy. Journal of Mathematical Physics, 54(2):023511, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p023511_s1.

Welter:2013:TIS

[1705] A. Welter, R. E. Raab, and O. L. de Lange. Translationally invariant semi-classical electrodynamics of magnetic media to electric octopole-magnetic quadrupole order. Journal of Mathematical Physics, 54(2):023512, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i2/p023512_s1.

Cheng:2013:GAS

[1706] Jipeng Cheng, Ye Tian, Zhaowen Yan, and Jingsong He. The generalized additional symmetries of the two-Toda lattice hierarchy. Journal of Mathematical Physics, 54(2):023513, February 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v54/i2/p023513_s1.

Asakly:2013:REW

[1707] Walaa Asakly, Toufik Mansour, and Matthias Schork. Representing elements of the Weyl algebra by labeled trees. Journal of Mathematical Physics, 54(2):023514, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v54/i2/p023514_s1.

Ni:2013:PB

[1708] Xiang Ni and Chengming Bai. Poisson bialgebras. Journal of Mathematical Physics, 54(2):023515, February 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i2/p023515_s1.

Secchi:2013:GSS

[1709] Simone Secchi. Ground state solutions for nonlinear fractional Schrödinger equations in R^{N}. Journal of Mathematical Physics, 54(3): 031501, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p031501_s1.

Felix:2013:NLT

[1710] José Antonio Alcántara Félix and Simone Calogero. Newtonian limit and trend to equilibrium for the relativistic Fokker–Planck equation. Journal of Mathematical Physics, 54(3):031502, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xia:2013:LRS

[1711] Li meng Xia, Lixin Tian, and Caixia Shen. Low-regularity solutions of the periodic modified two-component Camassa–Holm equation. Journal of Mathematical Physics, 54(3):031503, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lasiecka:2013:NID

[1712] Irena Lasiecka, Salim A. Messaoudi, and Muhammad I. Mustafa. Note on intrinsic decay rates for abstract wave equations with memory. Journal of Mathematical Physics, 54(3):031504, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luchko:2013:FWE

[1713] Yuri Luchko. Fractional wave equation and damped waves. Journal of Mathematical Physics, 54(3):031505, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2013:GEN

[1714] Min Ding and Yachun Li. Global existence and non-relativistic global limits of entropy solutions to the 1D piston problem for the isentropic relativistic Euler equations. Journal of Mathematical Physics, 54(3): 031506, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gallo:2013:EEG

[1715] Clément Gallo. Expansion of the energy of the ground state of the Gross– Pitaevskii equation in the Thomas–Fermi limit. Journal of Mathematical Physics, 54(3):031507, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neagu:2013:JTY

[1716] M. Neagu, N. G. Krylova, and H. V. Grushevskaya. Jet theoretical Yang– Mills energy in the geometric dynamics of two-dimensional monolayer. Journal of Mathematical Physics, 54(3):031508, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chowdhury:2013:SGN

[1717] S. Hasibul Hassan Chowdhury and S. Twareque Ali. The symmetry groups of noncommutative quantum mechanics and coherent state quantization. Journal of Mathematical Physics, 54(3):032101, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shkarin:2013:PP

[1718] Stanislav Shkarin. On Pauli pairs. Journal of Mathematical Physics, 54 (3):032102, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popov:2013:ANF

[1719] Dusan Popov, Vjekoslav Sajfert, Nicolina Pop, and Viorel Chiritoiu. About a new family of coherent states for some SU(1,1) central field potentials. Journal of Mathematical Physics, 54(3):032103, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turek:2013:QGQ

[1720] Ondrej Turek and Taksu Cheon. Quantum graph as a quantum spectral filter. Journal of Mathematical Physics, 54(3):032104, March 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hasler:2013:OVW

[1721] David Hasler and Daniel Luckett. Optimal volume Wegner estimate for random magnetic Laplacians on Z^{2}. Journal of Mathematical Physics, 54(3):032105, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flori:2013:GAT

[1722] C. Flori. Group action in topos quantum physics. Journal of Mathematical Physics, 54(3):032106, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goyeneche:2013:NMC

[1723] D. Goyeneche. A new method to construct families of complex Hadamard matrices in even dimensions. Journal of Mathematical Physics, 54(3): 032201, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ozawa:2013:TPA

[1724] Narutaka Ozawa. Tsirelson’s problem and asymptotically commuting unitary matrices. Journal of Mathematical Physics, 54(3):032202, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szehr:2013:PBQ

[1725] Oleg Szehr and Michael M. Wolf. Perturbation bounds for quantum Markov processes and their fixed points. Journal of Mathematical Physics, 54(3):032203, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjelakovic:2013:UQS

[1726] I. Bjelaković, H. Boche, and G. Janßen. Universal quantum state merging. Journal of Mathematical Physics, 54(3):032204, March 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leonard:2013:RVP

[1727] Katie E. Leonard, Tomislav Prokopec, and Richard P. Woodard. Representing the vacuum polarization on de Sitter. Journal of Mathematical Physics, 54(3):032301, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i3/p032301_s1.

Jiang:2013:RSQ

[1728] Jingjing Jiang and Song Wang. Representations of some quantum tori Lie subalgebras. Journal of Mathematical Physics, 54(3):032302, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohm:2013:TIG

[1729] Olaf Hohm and Barton Zwiebach. Towards an invariant geometry of double field theory. Journal of Mathematical Physics, 54(3):032303, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garfinkle:2013:KYT

[1730] David Garfinkle and E. N. Glass. Killing–Yano tensors in spaces admitting a hypersurface orthogonal Killing vector. Journal of Mathematical Physics, 54(3):032501, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2013:VPR

[1731] Xianfeng Ma and Ercai Chen. Variational principles for relative local pressure with subadditive potentials. Journal of Mathematical Physics, 54(3):032701, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Francoise:2013:RMA

[1732] J.-P. Françoise, P. L. Garrido, and G. Gallavotti. Rigid motions: Actionangles, relative cohomology and polynomials with roots on the unit circle. Journal of Mathematical Physics, 54(3):032901, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2013:HJT

[1733] Manuel de León, Juan Carlos Marrero, David Martín de Diego, and Miguel Vaquero. On the Hamilton–Jacobi theory for singular Lagrangian systems. Journal of Mathematical Physics, 54(3):032902, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kelbert:2013:QMW

[1734] Mark Kelbert and Yurii Suhov. A quantum Mermin–Wagner theorem for quantum rotators on two-dimensional graphs. Journal of Mathematical Physics, 54(3):033301, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i3/p033301_s1.

Kang:2013:EDK

[1735] Jum-Ran Kang. Exponential decay for a von Kármán equations with memory. Journal of Mathematical Physics, 54(3):033501, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/
p033501_s1.

Gordoa:2013:ABT

[1736] Pilar R. Gordoa and Andrew Pickering. Auto-Bäcklund transformations for a differential-delay equation. Journal of Mathematical Physics, 54(3): 033502, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p033502_s1.

Zhou:2013:NES

[1737] Wang Zhou. New estimators of spectral distributions of Wigner matrices. Journal of Mathematical Physics, 54(3):033503, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/p033503_s1.

Hayajneh:2013:LTT

[1738] Saja Hayajneh and Fuad Kittaneh. Lieb–Thirring trace inequalities and a question of Bourin. Journal of Mathematical Physics, 54(3):

033504, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p033504_s1.

deLucas:2013:LSK

[1739] J. de Lucas and C. Sardón. On Lie systems and Kummer–Schwarz equations. Journal of Mathematical Physics, 54(3):033505, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/

p033505_s1. See erratum [1984, 1739].

Yan:2013:FCP

[1740] Zhao-Wen Yan, Min-Li Li, Ke Wu, and Wei-Zhong Zhao. Fermionic covariant prolongation structure theory for multidimensional super nonlinear evolution equation. Journal of Mathematical Physics, 54(3): 033506, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p033506_s1.

Ferreira:2013:TPD

[1741] Erasmo Ferreira and Javier Sesma. Two-point derivative dispersion relations. Journal of Mathematical Physics, 54(3):033507, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/
p033507_s1.

Nayak:2013:ESL

[1742] Saudamini Nayak and K. C. Pati. Embedding of simply laced hyperbolic Kac–Moody superalgebras. Journal of Mathematical Physics, 54(3): 033508, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p033508_s1.

Dubrovsky:2013:ASN

[1743] V. G. Dubrovsky and A. V. Topovsky. About simple nonlinear and linear superpositions of special exact solutions of Veselov–Novikov equation. Journal of Mathematical Physics, 54(3):033509, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/p033509_s1.

Shibukawa:2013:OOM

[1744] Genki Shibukawa. Operator orderings and Meixner–Pollaczek polynomials. Journal of Mathematical Physics, 54(3):033510, March 2013.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i3/
p033510_s1.

Adams:2013:HES

[1745] Fred C. Adams and Anthony M. Bloch. Hill’s equation with small fluctuations: Cycle to cycle variations and stochastic processes. Journal of Mathematical Physics, 54(3):033511, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i3/p033511_s1.

Neate:2013:SBZ

[1746] Andrew Neate and Aubrey Truman. A stochastic Burgers–Zeldovich model for the formation of planetary ring systems and the satellites of Jupiter and Saturn. Journal of Mathematical Physics, 54(3):033512, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i3/p033512_s1.

Bravetti:2013:CMS

[1747] Alessandro Bravetti, Cesar S. Lopez-Monsalvo, Francisco Nettel, and Hernando Quevedo. The conformal metric structure of geometrothermodynamics. Journal of Mathematical Physics, 54(3):033513, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i3/p033513_s1.

Anzai:2013:AEM

[1748] C. Anzai and Y. Sumino. Algorithms to evaluate multiple sums for loop computations. Journal of Mathematical Physics, 54(3):033514, March 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i3/p033514_s1.

Ho:2013:SSF

[1749] C.-L. Ho. Similarity solutions of Fokker–Planck equation with moving boundaries. Journal of Mathematical Physics, 54(4):041501, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p041501_s1.

Eden:2013:GSB

[1750] A. Eden, V. K. Kalantarov, and S. V. Zelik. Global solvability and blow up for the convective Cahn–Hilliard equations with concave potentials. Journal of Mathematical Physics, 54(4):041502, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p041502_s1.

Tersenov:2013:FDS

[1751] Alkis S. Tersenov. On the fast diffusion with strong absorption. Journal of Mathematical Physics, 54(4):041503, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p041503_s1.

Arrizabalaga:2013:SAE

[1752] Naiara Arrizabalaga, Javier Duoandikoetxea, and Luis Vega. Self-adjoint extensions of Dirac operators with Coulomb type singularity. Journal of Mathematical Physics, 54(4):041504, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p041504_s1. See erratum

[2717].

Plastino:2013:NSE

[1753] A. R. Plastino and C. Tsallis. Nonlinear Schrödinger equation in the presence of uniform acceleration. Journal of Mathematical Physics, 54(4): 041505, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i4/p041505_s1.

Oliveri:2013:RBL

[1754] F. Oliveri and M. P. Speciale. Reduction of balance laws to conservation laws by means of equivalence transformations. Journal of Mathematical Physics, 54(4):041506, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p041506_s1.

Popoff:2013:SOI

[1755] Nicolas Popoff. The Schrödinger operator on an infinite wedge with a tangent magnetic field. Journal of Mathematical Physics, 54(4):041507, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p041507_s1.

Goenner:2013:WLS

[1756] Hubert Goenner. Weak Lie symmetry and extended Lie algebra. Journal of Mathematical Physics, 54(4):041701, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p041701_s1.

Milosevic:2013:PSP

[1757] D. B. Milosević. Phase space path-integral formulation of the abovethreshold ionization. Journal of Mathematical Physics, 54(4):042101, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p042101_s1.

Marquette:2013:NFS

[1758] Ian Marquette and Christiane Quesne. New families of superintegrable systems from Hermite and Laguerre exceptional orthogonal polynomials. Journal of Mathematical Physics, 54(4):042102, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p042102_s1.

Kurasov:2013:ISL

[1759] Pavel Kurasov. Inverse scattering for lasso graph. Journal of Mathematical Physics, 54(4):042103, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i4/p042103_s1.

Candemir:2013:MDE

[1760] N. Candemir and O. Bayrak. Massive Dirac equation in asymmetric Hulthén potential. Journal of Mathematical Physics, 54(4):042104, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p042104_s1.

Gridnev:2013:WTN

[1761] Dmitry K. Gridnev. Why there is no Efimov effect for four bosons and related results on the finiteness of the discrete spectrum. Journal of Mathematical Physics, 54(4):042105, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p042105_s1. See comment

Gonzalez-Diaz:2013:BSS

[1762] L. A. González-Díaz and S. Díaz-Solórzano. Bound states and scattering coefficients of self-adjoint Hamiltonians with a mass jump. Journal of Mathematical Physics, 54(4):042106, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p042106_s1.

Opanchuk:2013:FWR

[1763] B. Opanchuk and P. D. Drummond. Functional Wigner representation of quantum dynamics of Bose — Einstein condensate. Journal of Mathematical Physics, 54(4):042107, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i4/p042107_s1.

Schulze-Halberg:2013:SPC

[1764] Axel Schulze-Halberg and Christopher R. Gordon. Spectral properties of a confined nonlinear quantum oscillator in one and three dimensions. Journal of Mathematical Physics, 54(4):042108, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p042108_s1.

Kiukas:2013:FCS

[1765] J. Kiukas, A. Ruschhaupt, and R. F. Werner. Full counting statistics of stationary particle beams. Journal of Mathematical Physics, 54(4): 042109, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i4/p042109_s1.

Nachtergaele:2013:ALB

[1766] Bruno Nachtergaele, Robert Sims, and Günter Stolz. An area law for the bipartite entanglement of disordered oscillator systems. Journal of Mathematical Physics, 54(4):042110, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p042110_s1.

Datta:2013:QCR

[1767] Nilanjana Datta, Min-Hsiu Hsieh, Mark M. Wilde, and Andreas Winter. Quantum-to-classical rate distortion coding. Journal of Mathematical Physics, 54(4):042201, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p042201_s1.

Fukuda:2013:PTR

[1768] Motohisa Fukuda and Piotr Śniady. Partial transpose of random quantum states: Exact formulas and meanders. Journal of Mathematical Physics, 54(4):042202, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p042202_s1.

Hofer-Szabo:2013:NCC

[1769] Gábor Hofer-Szabó and Péter Vecsernyés. Noncommutative common cause principles in algebraic quantum field theory. Journal of Mathematical Physics, 54(4):042301, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i4/p042301_s1.

Oliva:2013:ASF

[1770] Julio Oliva. All the solutions of the form M_{2} × _{W} Σ_{d−2} for Lovelock gravity in vacuum in the Chern–Simons case. Journal of Mathematical Physics, 54(4):042501, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p042501_s1.

Batista:2013:WTC

[1771] Carlos Batista. On the Weyl tensor classification in all dimensions and its relation with integrability properties. Journal of Mathematical Physics, 54(4):042502, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p042502_s1.

Alhulaimi:2013:AEA

[1772] B. Alhulaimi, A. Coley, and P. Sandin. Anisotropic Einstein–aether cosmological models. Journal of Mathematical Physics, 54(4):042503, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p042503_s1.

Hu:2013:ESF

[1773] Guanghui Hu and Mourad Sini. Elastic scattering by finitely many pointlike obstacles. Journal of Mathematical Physics, 54(4):042901, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p042901_s1.

Valovik:2013:PNC

[1774] Dmitry V. Valovik. On the problem of nonlinear coupled electromagnetic transverse-electric-transverse magnetic wave propagation. Journal of Mathematical Physics, 54(4):042902, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p042902_s1.

Itin:2013:CTL

[1775] Yakov Itin and Friedrich W. Hehl. The constitutive tensor of linear elasticity: Its decompositions, Cauchy relations, null Lagrangians, and wave propagation. Journal of Mathematical Physics, 54(4):042903, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p042903_s1.

Noble:2013:KMR

[1776] A. Noble, D. A. Burton, J. Gratus, and D. A. Jaroszynski. A kinetic model of radiating electrons. Journal of Mathematical Physics, 54(4): 043101, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i4/p043101_s1.

Misra:2013:SCA

[1777] Kailash C. Misra and Evan A. Wilson. Soliton cellular automaton associated with D_{n}^{(1)}-crystal B^{2,s}. Journal of Mathematical Physics, 54(4): 043301, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i4/p043301_s1.

Cacciapuoti:2013:LMP

[1778] Claudio Cacciapuoti, Anna Maltsev, and Benjamin Schlein. Local Marchenko–Pastur law at the hard edge of sample covariance matrices. Journal of Mathematical Physics, 54(4):043302, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p043302_s1.

Banchi:2013:SPQ

[1779] Leonardo Banchi and Ruggero Vaia. Spectral problem for quasi-uniform nearest-neighbor chains. Journal of Mathematical Physics, 54(4):043501, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p043501_s1.

Stamova:2013:LSC

[1780] Ivanka Stamova and Gani Stamov. Lipschitz stability criteria for functional differential systems of fractional order. Journal of Mathematical Physics, 54(4):043502, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p043502_s1.

Rink:2013:NAV

[1781] Norman A. Rink. Non-abelian vortices on CP^{1} and Grassmannians. Journal of Mathematical Physics, 54(4):043503, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p043503_s1.

Liu:2013:GDS

[1782] Wenjun Liu. General decay of the solution for a viscoelastic wave equation with a time-varying delay term in the internal feedback. Journal of Mathematical Physics, 54(4):043504, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p043504_s1.

Wang:2013:FSI

[1783] Hui Wang and Tie-Cheng Xia. The fractional supertrace identity and its application to the super Ablowitz — Kaup — Newell — Segur hierarchy. Journal of Mathematical Physics, 54(4):043505, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/p043505_s1.

Smirnov:2013:CET

[1784] Yury G. Smirnov and Dmitry V. Valovik. Coupled electromagnetic transverse-electric-transverse magnetic wave propagation in a cylindrical waveguide with Kerr nonlinearity. Journal of Mathematical Physics, 54(4):043506, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p043506_s1.

Kojima:2013:DTM

[1785] Takeo Kojima. Diagonalization of transfer matrix of supersymmetry U_{q}( ˆ∫↕(M + 1|N + 1)) chain with a boundary. Journal of Mathematical Physics, 54(4):043507, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p043507_s1.

Lucas:2013:VCG

[1786] Pascual Lucas and José Antonio Ortega-Yagües. A variational characterization and geometric integration for Bertrand curves. Journal of Mathematical Physics, 54(4):043508, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i4/p043508_s1.

Kumar:2013:SFK

[1787] Dilip Kumar. Solution of fractional kinetic equation by a class of integral transform of pathway type. Journal of Mathematical Physics, 54(4): 043509, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v54/i4/p043509_s1.

Zhong:2013:SNR

[1788] Y. Zhong and G. L. Gao. Some new results about the massless Dirac operator. Journal of Mathematical Physics, 54(4):043510, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v54/i4/
p043510_s1.

Yoo-Kong:2013:ESS

[1789] Sikarin Yoo-Kong and Frank Nijhoff. Elliptic (N, N′)-soliton solutions of the lattice Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 54(4):043511, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p043511_s1.

Li:2013:VTA

[1790] Maohua Li, Chuanzhong Li, Keilei Tian, Jingsong He, and Yi Cheng. Virasoro type algebraic structure hidden in the constrained discrete Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 54(4):043512, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v54/i4/p043512_s1.

Tian:2013:THD

[1791] Kai Tian and Q. P. Liu. Tri-Hamiltonian duality between the Wadati– Konno–Ichikawa hierarchy and the Song–Qu–Qiao hierarchy. Journal of Mathematical Physics, 54(4):043513, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v54/i4/p043513_s1.

Rajan:2013:MSP

[1792] M. S. Mani Rajan and A. Mahalingam. Multi-soliton propagation in a generalized inhomogeneous nonlinear Schrödinger–Maxwell–Bloch system with loss/gain driven by an external potential. Journal of Mathematical Physics, 54(4):043514, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v54/i4/p043514_s1.

Yolcu:2013:EEC

[1793] Selma Yildirim Yolcu and Türkay Yolcu. Estimates on the eigenvalues of the clamped plate problem on domains in Euclidean spaces. Journal of Mathematical Physics, 54(4):043515, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v54/i4/p043515_s1. See pub-

lisher’s note [1983].

Fordy:2013:RPK

[1794] Allan P. Fordy and Michael J. Scott. Recursive procedures for Krall– Sheffer operators. Journal of Mathematical Physics, 54(4):043516, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p043516_s1.

Bertola:2013:SAC

[1795] M. Bertola, M. Gekhtman, and J. Szmigielski. Strong asymptotics for Cauchy biorthogonal polynomials with application to the Cauchy twomatrix model. Journal of Mathematical Physics, 54(4):043517, April 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v54/i4/p043517_s1.

Kengne:2013:EWS

[1796] E. Kengne, A. Lakhssassi, R. Vaillancourt, and Wu-Ming Liu. Exact wave solutions for Bose–Einstein condensates with time-dependent scattering length and spatiotemporal complicated potential. Journal of Mathematical Physics, 54(5):051501, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2013:SSO

[1797] Yukun Song, Hongjun Yuan, and Yang Chen. On the strong solutions of one-dimensional Navier–Stokes–Poisson equations for compressible non-Newtonian fluids. Journal of Mathematical Physics, 54(5):051502, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2013:WSN

[1798] Shangbin Cui. Weak solutions for the Navier–Stokes equations with B_{∞∞}^{−1(ln)} + B_{X_{r}}^{−1+r,2/(1−r)} + L^{2}. Journal of Mathematical Physics, 54(5): 051503, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:2013:BTS

[1799] Yoshimasa Matsuno. Bäcklund transformation and smooth multisoliton solutions for a modified Camassa–Holm equation with cubic nonlinearity. Journal of Mathematical Physics, 54(5):051504, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mingo:2013:RSO

[1800] James A. Mingo and Mihai Popa. Real second order freeness and Haar orthogonal matrices. Journal of Mathematical Physics, 54(5):051701, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2013:MOS

[1801] Yi-Zhi Huang. Meromorphic open-string vertex algebras. Journal of Mathematical Physics, 54(5):051702, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:RTA

[1802] Xue Chen, Cuipo Jiang, and Qifen Jiang. Representations of the twisted affine Nappi–Witten algebras. Journal of Mathematical Physics, 54(5): 051703, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2013:ECQ

[1803] Haisheng Li. G-equivariant φ-coordinated quasi modules for quantum vertex algebras. Journal of Mathematical Physics, 54(5):051704, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jursenas:2013:BSS

[1804] Rytis Jursenas and Julius Ruseckas. Bound states of the spin-orbit coupled ultracold atom in a one-dimensional short-range potential. Journal of Mathematical Physics, 54(5):051901, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhattacharya:2013:HED

[1805] Samyadeb Bhattacharya and Sisir Roy. Hartman effect and dissipative quantum systems. Journal of Mathematical Physics, 54(5):052101, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galikova:2013:CPN

[1806] Veronika Gáliková and Peter Presnajder. Coulomb problem in noncommutative quantum mechanics. Journal of Mathematical Physics, 54 (5):052102, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albeverio:2013:SSO

[1807] Sergio Albeverio, Aleksey Kostenko, Mark Malamud, and Hagen Neidhardt. Spherical Schrödinger operators with δ-type interactions. Journal of Mathematical Physics, 54(5):052103, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagheri:2013:TNC

[1808] H. Bagheri and A. Mahdifar. Thermal nonlinear coherent states on a flat space and on a sphere. Journal of Mathematical Physics, 54(5): 052104, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoang-Do:2013:ASS

[1809] Ngoc-Tram Hoang-Do, Van-Hung Hoang, and Van-Hoang Le. Analytical solutions of the Schrödinger equation for a two-dimensional exciton in magnetic field of arbitrary strength. Journal of Mathematical Physics, 54(5):052105, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigin:2013:BAF

[1810] M. V. Feigin, M. A. Hallnäs, and A. P. Veselov. Baker–Akhiezer functions and generalised Macdonald–Mehta integrals. Journal of Mathematical Physics, 54(5):052106, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fritz:2013:IDS

[1811] Tobias Fritz. On infinite-dimensional state spaces. Journal of Mathematical Physics, 54(5):052107, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gimeno:2013:GAN

[1812] Vicent Gimeno and Jose M. Sotoca. Geometric approach to nonrelativistic quantum dynamics of mixed states. Journal of Mathematical Physics, 54(5):052108, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lopez-Rosa:2013:ECA

[1813] S. López-Rosa, I. V. Toranzo, P. Sánchez-Moreno, and J. S. Dehesa. Entropy and complexity analysis of hydrogenic Rydberg atoms. Journal of Mathematical Physics, 54(5):052109, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2715, 2716].

Farenick:2013:ACQ

[1814] Douglas Farenick, Remus Floricel, and Sarah Plosker. Approximately clean quantum probability measures. Journal of Mathematical Physics, 54(5):052201, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kastoryano:2013:QLS

[1815] Michael J. Kastoryano and Kristan Temme. Quantum logarithmic Sobolev inequalities and rapid mixing. Journal of Mathematical Physics, 54(5):052202, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thylwe:2013:BSQ

[1816] Karl-Erik Thylwe and Patrick McCabe. Bohr–Sommerfeld quantization condition for Dirac states derived from an Ermakov-type invariant. Journal of Mathematical Physics, 54(5):052301, May 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandt:2013:SAC

[1817] Friedemann Brandt. Supersymmetry algebra cohomology. IV. Primitive elements in all dimensions from D = 4 to D = 11. Journal of Mathematical Physics, 54(5):052302, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2013:TLS

[1818] Luise Adams, Christian Bogner, and Stefan Weinzierl. The two-loop sunrise graph with arbitrary masses. Journal of Mathematical Physics, 54(5):052303, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jeffrey:2013:SQMa

[1819] Lisa C. Jeffrey. Symplectic quantum mechanics and Chern–Simons gauge theory. I. Journal of Mathematical Physics, 54(5):052304, May 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jeffrey:2013:SQMb

[1820] Lisa C. Jeffrey. Symplectic quantum mechanics and Chern–Simons gauge theory. II. Mapping tori of tori. Journal of Mathematical Physics, 54(5): 052305, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gielen:2013:LGR

[1821] Steffen Gielen and Derek K. Wise. Lifting general relativity to observer space. Journal of Mathematical Physics, 54(5):052501, May 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Batista:2013:SWT

[1822] Carlos Batista and Bruno Carneiro da Cunha. Spinors and the Weyl tensor classification in six dimensions. Journal of Mathematical Physics, 54(5):052502, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ebrahimi:2013:SOC

[1823] N. Ebrahimi. Some observations on cosmological time functions. Journal of Mathematical Physics, 54(5):052503, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2013:RTT

[1824] Xiaoping Yuan and Kangkang Zhang. A reduction theorem for time dependent Schrödinger operator with finite differentiable unbounded perturbation. Journal of Mathematical Physics, 54(5):052701, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kitano:2013:MSU

[1825] Masao Kitano. Mathematical structure of unit systems. Journal of Mathematical Physics, 54(5):052901, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meng:2013:PRM

[1826] Guowu Meng. The Poisson realization of so(2, 2k +2) on magnetic leaves and generalized Micz–Kepler problems. Journal of Mathematical Physics, 54(5):052902, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sanami:2013:CHF

[1827] Abolfazl Sanami. A characterization of the Helmholtz free energy. Journal of Mathematical Physics, 54(5):053301, May 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qian:2013:DID

[1828] Hong Qian. A decomposition of irreversible diffusion processes without detailed balance. Journal of Mathematical Physics, 54(5):053302, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tautz:2013:CRD

[1829] R. C. Tautz and I. Lerche. Cosmic-ray diffusion modeling: Solutions using variational methods. Journal of Mathematical Physics, 54(5):053303, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nehring:2013:CPPa

[1830] Benjamin Nehring. Construction of point processes for classical and quantum gases. Journal of Mathematical Physics, 54(5):053304, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2013:SLH

[1831] Seong Sik Kim, John Michael Rassias, Yeol Je Cho, and Soo Hawn Kim. Stability of n-Lie homomorphisms and Jordan n-Lie homomorphisms on n-Lie algebras. Journal of Mathematical Physics, 54(5):053501, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hornyak:2013:CDP

[1832] I. Hornyak and A. T. Kruppa. Coulomb-distorted plane wave: Partial wave expansion and asymptotic forms. Journal of Mathematical Physics, 54(5):053502, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Su:2013:SVL

[1833] Yucai Su and Lamei Yuan. Schrödinger-Virasoro Lie conformal algebra. Journal of Mathematical Physics, 54(5):053503, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bucur:2013:ABO

[1834] Dorin Bucur and Pedro Freitas. Asymptotic behaviour of optimal spectral planar domains with fixed perimeter. Journal of Mathematical Physics, 54(5):053504, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Godin:2013:ECP

[1835] Yuri A. Godin. Effective complex permittivity tensor of a periodic array of cylinders. Journal of Mathematical Physics, 54(5):053505, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tiwari:2013:CLP

[1836] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0. Journal of Mathematical Physics, 54(5): 053506, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2325], comments [2707, 2718], and response [2719].

Maltsev:2013:MDH

[1837] A. Ya. Maltsev. The multi-dimensional Hamiltonian structures in the Whitham method. Journal of Mathematical Physics, 54(5):053507, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:2013:LAH

[1838] Xiao ying Zhu and Da jun Zhang. Lie algebras and Hamiltonian structures of multi-component Ablowitz–Kaup–Newell–Segur hierarchy. Journal of Mathematical Physics, 54(5):053508, May 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-y-Romero:2013:TDM

[1839] R. P. Martínez y Romero, H. N. Núñez-Yépez, and A. L. Salas-Brito. The two dimensional motion of a particle in an inverse square potential: Classical and quantum aspects. Journal of Mathematical Physics, 54(5):

053509, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2625].

Du:2013:RBC

[1840] Dianlou Du and Xue Geng. On the relationship between the classical Dicke–Jaynes–Cummings–Gaudin model and the nonlinear Schrödinger equation. Journal of Mathematical Physics, 54(5):053510, May 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoshino:2013:PRC

[1841] A. Hoshino. Polyhedral realizations of crystal bases for quantum algebras of classical affine types. Journal of Mathematical Physics, 54(5):053511, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakabekov:2013:AMC

[1842] A. Sakabekov and E. Auzhani. Analogue of the mass conservation law for third approximation of one-dimensional nonlinear Boltzmann’s moment system equations. Journal of Mathematical Physics, 54(5):053512, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kriz:2013:DFT

[1843] Igor Kriz, Martin Loebl, and Petr Somberg. On discrete field theory properties of the dimer and Ising models and their conformal field theory limits. Journal of Mathematical Physics, 54(5):053513, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2013:DSP

[1844] Jun Guo, Guozheng Yan, and Jianguo Ye. The direct scattering problem for a crack buried in a piecewise homogeneous medium. Journal of Mathematical Physics, 54(5):053514, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zullo:2013:IDA

[1845] Federico Zullo. On an integrable discretisation of the Ablowitz–Ladik hierarchy. Journal of Mathematical Physics, 54(5):053515, May 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Niccoli:2013:FFC

[1846] G. Niccoli. Form factors and complete spectrum of XXX antiperiodic higher spin chains by quantum separation of variables. Journal of Mathematical Physics, 54(5):053516, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:ASB

[1847] Xia Zhang and Ming Liu. On almost surely bounded semigroups of random linear operators. Journal of Mathematical Physics, 54(5):053517, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shin:2013:ISP

[1848] Jaemin Shin. Inverse scattering problems for the Landau–Lifschitz equations. Journal of Mathematical Physics, 54(5):053701, May 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2013:HCS

[1849] A. Alexandrou Himonas and John Holmes. Hölder continuity of the solution map for the Novikov equation. Journal of Mathematical Physics, 54(6):061501, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schechter:2013:PL

[1850] Martin Schechter. Photonic lattices. Journal of Mathematical Physics, 54(6):061502, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Segata:2013:OST

[1851] Jun ichi Segata. Orbital stability of a two parameter family of solitary waves for a fourth order nonlinear Schrödinger type equation. Journal of Mathematical Physics, 54(6):061503, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2013:GSS

[1852] Xiaojun Chang. Ground state solutions of asymptotically linear fractional Schrödinger equations. Journal of Mathematical Physics, 54(6): 061504, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2013:ECS

[1853] Yanheng Ding, Juncheng Wei, and Tian Xu. Existence and concentration of semi-classical solutions for a nonlinear Maxwell–Dirac system. Journal of Mathematical Physics, 54(6):061505, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grunbaum:2013:BDP

[1854] F. Alberto Grünbaum, Luc Vinet, and Alexei Zhedanov. Birth and death processes and quantum spin chains. Journal of Mathematical Physics, 54(6):062101, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khordad:2013:EMF

[1855] R. Khordad. Effects of the magnetic field direction and anisotropy on the interband light absorption of an asymmetric quantum dot. Journal of Mathematical Physics, 54(6):062102, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carlen:2013:EPD

[1856] Eric A. Carlen, Joel L. Lebowitz, and Elliott H. Lieb. On an extension problem for density matrices. Journal of Mathematical Physics, 54(6): 062103, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lammert:2013:CGS

[1857] Paul E. Lammert. Coarse-grained spin density-functional theory: Infinite-volume limit via the hyperfinite. Journal of Mathematical Physics, 54(6):062104, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohlenkamp:2013:FSR

[1858] Martin J. Mohlenkamp. Function space requirements for the singleelectron functions within the multiparticle Schrödinger equation. Journal of Mathematical Physics, 54(6):062105, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaturvedi:2013:NPC

[1859] S. Chaturvedi, E. Ercolessi, G. Morandi, A. Ibort, G. Marmo, N. Mukunda, and R. Simon. Null phase curves and manifolds in geometric phase theory. Journal of Mathematical Physics, 54(6):062106, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapustin:2013:QME

[1860] Anton Kapustin. Is quantum mechanics exact? Journal of Mathematical Physics, 54(6):062107, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiai:2013:CQE

[1861] Fumio Hiai and Dénes Petz. Convexity of quasi-entropy type functions: Lieb’s and Ando’s convexity theorems revisited. Journal of Mathematical Physics, 54(6):062201, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Childs:2013:PFE

[1862] Andrew M. Childs and Nathan Wiebe. Product formulas for exponentials of commutators. Journal of Mathematical Physics, 54(6):062202, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fabbri:2013:MST

[1863] Luca Fabbri. Metric solutions in torsionless gauge for vacuum conformal gravity. Journal of Mathematical Physics, 54(6):062501, June 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:2013:IUA

[1864] Hans-Jürgen Schmidt and Douglas Singleton. Isotropic universe with almost scale-invariant fourth-order gravity. Journal of Mathematical Physics, 54(6):062502, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamimura:2013:CAB

[1865] Kiyoshi Kamimura and Seiji Onda. Contractions of AdS brane algebra and superGalileon Lagrangians. Journal of Mathematical Physics, 54(6): 062503, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lompay:2013:CDIa

[1866] Robert R. Lompay and Alexander N. Petrov. Covariant differential identities and conservation laws in metric-torsion theories of gravitation. I. General consideration. Journal of Mathematical Physics, 54(6):062504, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pasca:2013:QAA

[1867] Daniel Pasca and Cláudia Valls. Qualitative analysis of the anisotropic two-body problem with relativistic potential. Journal of Mathematical Physics, 54(6):062701, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgieva:2013:IID

[1868] Bogdana Georgieva and Theodore Bodurov. Identities from infinitedimensional symmetries of Herglotz variational functional. Journal of Mathematical Physics, 54(6):062901, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:2013:REP

[1869] Seungsu Hwang and Sun-Chul Kim. Relative equilibria of point vortices on the hyperbolic sphere. Journal of Mathematical Physics, 54(6): 063101, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez:2013:BCL

[1870] J. C. Hernandez, Y. Suhov, A. Yambartsev, and S. Zohren. Bounds on the critical line via transfer matrix methods for an Ising model coupled to causal dynamical triangulations. Journal of Mathematical Physics, 54(6):063301, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nehring:2013:CPPb

[1871] Benjamin Nehring, Suren Poghosyan, and Hans Zessin. On the construction of point processes in statistical mechanics. Journal of Mathematical Physics, 54(6):063302, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vertman:2013:EHT

[1872] Boris Vertman. The exotic heat-trace asymptotics of a regular-singular operator revisited. Journal of Mathematical Physics, 54(6):063501, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deVerdiere:2013:STG

[1873] Yves Colin de Verdière and Françoise Truc. Scattering theory for graphs isomorphic to a regular tree at infinity. Journal of Mathematical Physics, 54(6):063502, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gomez-Lobo:2013:CND

[1874] Alfonso García-Parrado Gómez-Lobo. On the characterization of nondegenerate foliations of pseudo-Riemannian manifolds with conformally flat leaves. Journal of Mathematical Physics, 54(6):063503, June 2013.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bai:2013:RBL

[1875] RuiPu Bai, Li Guo, Jiaqian Li, and Yong Wu. Rota–Baxter 3-Lie algebras. Journal of Mathematical Physics, 54(6):063504, June 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lo:2013:WLM

[1876] Assane Lo. Witten Laplacian method for generalized models of Kac type. Journal of Mathematical Physics, 54(6):063505, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:RMM

[1877] Yang Chen, Nazmus S. Haq, and Matthew R. McKay. Random matrix models, double-time Painlevé equations, and wireless relaying. Journal of Mathematical Physics, 54(6):063506, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ammar:2013:SCD

[1878] Faouzi Ammar, Abdenacer Makhlouf, and Nejib Saadaoui. Second cohomology of q-deformed Witt superalgebras. Journal of Mathematical Physics, 54(6):063507, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benatti:2013:CLQ

[1879] Fabio Benatti and Laure Gouba. Classical limits of quantum mechanics on a non-commutative configuration space. Journal of Mathematical Physics, 54(6):063508, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Azad:2013:SAW

[1880] H. Azad, Ahmad Y. Al-Dweik, R. Ghanam, and M. T. Mustafa. Symmetry analysis of wave equation on static spherically symmetric spacetimes with higher symmetries. Journal of Mathematical Physics, 54(6):063509, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motegi:2013:BOH

[1881] Kohei Motegi. On Baxter’s Q operator of the higher spin XXZ chain at the Razumov–Stroganov point. Journal of Mathematical Physics, 54(6):

063510, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Izumiya:2013:LDC

[1882] Shyuichi Izumiya, Yang Jiang, and Takami Sato. Lightcone dualities for curves in the lightcone unit 3-sphere. Journal of Mathematical Physics, 54(6):063511, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2013:MMP

[1883] F. Bagarello. More mathematics for pseudo-bosons. Journal of Mathematical Physics, 54(6):063512, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1887].

Cohl:2013:SVA

[1884] H. S. Cohl and H. Volkmer. Separation of variables in an asymmetric cyclidic coordinate system. Journal of Mathematical Physics, 54(6):063513, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [1520].

Ghanmi:2013:CCO

[1885] Allal Ghanmi and Ahmed Intissar. Construction of concrete orthonormal basis for (L^{2}, Γ, χ)-theta functions associated to discrete subgroups of rank one in (C, +). Journal of Mathematical Physics, 54(6):063514, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zarmi:2013:NQD

[1886] Yair Zarmi. Nonlinear quantum-dynamical system based on the Kadomtsev–Petviashvili II equation. Journal of Mathematical Physics, 54(6):063515, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golenia:2013:CPD

[1887] Sylvain Golénia and Christoph Schumacher. Comment on “The problem of deficiency indices for discrete Schrödinger operators on locally finite graphs” [J. Math. Phys. 52, 063512 (2011)]. Journal of Mathematical Physics, 54(6):064101, June 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1883].

Mucha:2013:CRM

[1888] P. B. Mucha, M. Pokorný, and E. Zatorska. Chemically reacting mixtures in terms of degenerated parabolic setting. Journal of Mathematical Physics, 54(7):071501, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2013:SSQ

[1889] Jun Yang, Youjun Wang, and Ahamed Adam Abdelgadir. Soliton solutions for quasilinear Schrödinger equations. Journal of Mathematical Physics, 54(7):071502, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Curro:2013:RPE

[1890] C. Curró and N. Manganaro. Riemann problems and exact solutions to a traffic flow model. Journal of Mathematical Physics, 54(7):071503, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dou:2013:ODC

[1891] Changsheng Dou and Quansen Jiu. On one-dimensional compressible Navier–Stokes equations with degenerate viscosity and constant state at far fields. Journal of Mathematical Physics, 54(7):071504, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radobolja:2013:SVV

[1892] Gordan Radobolja. Subsingular vectors in Verma modules, and tensor product of weight modules over the twisted Heisenberg–Virasoro algebra and W (2, 2) algebra. Journal of Mathematical Physics, 54(7):071701, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marquette:2013:QPA

[1893] Ian Marquette. Quartic Poisson algebras and quartic associative algebras and realizations as deformed oscillator algebras. Journal of Mathematical Physics, 54(7):071702, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Downing:2013:SSE

[1894] C. A. Downing. On a solution of the Schrödinger equation with a hyperbolic double-well potential. Journal of Mathematical Physics, 54(7): 072101, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakayama:2013:TQT

[1895] Kunji Nakayama. Topologies on quantum topoi induced by quantization. Journal of Mathematical Physics, 54(7):072102, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garbaczewski:2013:LFN

[1896] Piotr Garbaczewski and Vladimir Stephanovich. Lévy flights and nonlocal quantum dynamics. Journal of Mathematical Physics, 54(7):072103, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paulsen:2013:CPM

[1897] Vern I. Paulsen and Fred Shultz. Complete positivity of the map from a basis to its dual basis. Journal of Mathematical Physics, 54(7):072201, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cabello:2013:ESG

[1898] Adán Cabello, Matthew G. Parker, Giannicola Scarpa, and Simone Severini. Exclusivity structures and graph representatives of local complementation orbits. Journal of Mathematical Physics, 54(7):072202, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varshovi:2013:GMP

[1899] Amir Abbass Varshovi. Groenewold–Moyal product, α∗-cohomology, and classification of translation-invariant non-commutative structures. Journal of Mathematical Physics, 54(7):072301, July 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holland:2013:OPE

[1900] Jan Holland and Stefan Hollands. Operator product expansion algebra. Journal of Mathematical Physics, 54(7):072302, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gadella:2013:PFD

[1901] M. Gadella, F. Gómez-Cubillo, L. Rodriguez, and S. Wickramasekara. Point-form dynamics of quasistable states. Journal of Mathematical Physics, 54(7):072303, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deguchi:2013:PHS

[1902] Shinichi Deguchi and Jun ichi Note. (pre-)Hilbert spaces in twistor quantization. Journal of Mathematical Physics, 54(7):072304, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llosa:2013:CSC

[1903] Josep Llosa. Collineations of a symmetric 2-covariant tensor: Ricci collineations. Journal of Mathematical Physics, 54(7):072501, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2013:RDK

[1904] Young-Pil Choi, Seung-Yeal Ha, and Se Eun Noh. On the relaxation dynamics of the Kuramoto oscillators with small inertia. Journal of Mathematical Physics, 54(7):072701, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schirinzi:2013:FNE

[1905] Gabriella Schirinzi and Massimiliano Guzzo. On the formulation of new explicit conditions for steepness from a former result of N. N. Nekhoroshev. Journal of Mathematical Physics, 54(7):072702, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lauritzen:2013:BPD

[1906] A. Lauritzen, P. Gustainis, and R. B. Mann. The 4-body problem in a (1 + 1)-dimensional self-gravitating system. Journal of Mathematical Physics, 54(7):072703, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tudoran:2013:FRB

[1907] Razvan M. Tudoran. The free rigid body dynamics: Generalized versus classic. Journal of Mathematical Physics, 54(7):072704, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Waclawczyk:2013:AEL

[1908] Marta Waclawczyk and Martin Oberlack. Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation. Journal of Mathematical Physics, 54(7):072901, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2710] and response [2711].

Gonzales:2013:TDS

[1909] Marcelo Gonzales, Sadi Khodaee, Olaf Lechtenfeld, and Francesco Toppan. Target duality in N = 8 superconformal mechanics and the coupling of dual pairs. Journal of Mathematical Physics, 54(7):072902, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contucci:2013:IFI

[1910] Pierluigi Contucci, Cristian Giardinà, and Claudio Giberti. Interaction flip identities for non-centered spin glasses. Journal of Mathematical Physics, 54(7):073301, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2013:GSV

[1911] Ho Lee. Global solutions of the Vlasov–Poisson–Boltzmann system in a cosmological setting. Journal of Mathematical Physics, 54(7):073302, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babusci:2013:SME

[1912] D. Babusci, G. Dattoli, K. Górska, and K. A. Penson. Symbolic methods for the evaluation of sum rules of Bessel functions. Journal of Mathematical Physics, 54(7):073501, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia:2013:QRC

[1913] José I. García and José I. Liberati. Quasifinite representations of classical Lie subalgebras of W_{∞,p}. Journal of Mathematical Physics, 54(7): 073502, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2013:FDI

[1914] Yuezhu Wu. Finite dimensional indecomposable modules for Schrödinger algebra. Journal of Mathematical Physics, 54(7):073503, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teo:2013:ECE

[1915] L. P. Teo. Electromagnetic Casimir effect on the boundary of a Ddimensional cavity and the high temperature asymptotics. Journal of Mathematical Physics, 54(7):073504, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hou:2013:AGS

[1916] Yu Hou, Engui Fan, and Peng Zhao. Algebro-geometric solutions for the Gerdjikov–Ivanov hierarchy. Journal of Mathematical Physics, 54(7): 073505, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Audenaert:2013:ARE

[1917] Koenraad M. R. Audenaert. On the asymmetry of the relative entropy. Journal of Mathematical Physics, 54(7):073506, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karpeshina:2013:MAM

[1918] Yulia Karpeshina and Roman Shterenberg. Multiscale analysis in momentum space for quasi-periodic potential in dimension two. Journal of Mathematical Physics, 54(7):073507, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majewski:2013:SPM

[1919] Wladyslaw A. Majewski and Tomasz I. Tylec. On the structure of positive maps. II. Low dimensional matrix algebras. Journal of Mathematical Physics, 54(7):073508, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valtancoli:2013:AMH

[1920] P. Valtancoli. Algebraic method for the harmonic oscillator with a minimal length. Journal of Mathematical Physics, 54(7):073509, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sasaki:2013:LEG

[1921] Yoshikatsu Sasaki. Lower estimates of the growth order of solutions to the second Painlevé hierarchy. Journal of Mathematical Physics, 54(7): 073510, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vacaru:2013:AIT

[1922] Sergiu I. Vacaru. The algebraic index theorem and deformation quantization of Lagrange–Finsler and Einstein spaces. Journal of Mathematical Physics, 54(7):073511, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grandati:2013:DRM

[1923] Y. Grandati and C. Quesne. Disconjugacy, regularity of multi-indexed rationally extended potentials, and Laguerre exceptional polynomials. Journal of Mathematical Physics, 54(7):073512, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vignat:2013:PCG

[1924] C. Vignat and O. Lévêque. Proof of a conjecture by Gazeau et al. using the Gould–Hopper polynomials. Journal of Mathematical Physics, 54(7): 073513, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biyogmam:2013:LHG

[1925] Guy Roger Biyogmam. Leibniz homology of the Galilei algebra. Journal of Mathematical Physics, 54(7):073514, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:IMH

[1926] Yuanyuan Chen, Zhongwei Wang, and Liangyun Zhang. Integrals for monoidal Hom–Hopf algebras and their applications. Journal of Mathematical Physics, 54(7):073515, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ali:2013:EPF

[1927] S. Twareque Ali, F. Bagarello, and Jean Pierre Gazeau. Extended pseudo-fermions from non commutative bosons. Journal of Mathematical Physics, 54(7):073516, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soloviev:2013:ACS

[1928] M. A. Soloviev. Algebras with convergent star products and their representations in Hilbert spaces. Journal of Mathematical Physics, 54(7): 073517, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yue:2013:FQF

[1929] Chao Yue and Tiecheng Xia. The fractional quadratic-form identity and Hamiltonian structure of an integrable coupling of the fractional Ablowitz–Kaup–Newell–Segur hierarchy. Journal of Mathematical Physics, 54(7):073518, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaillard:2013:SPD

[1930] Pierre Gaillard. Six-parameters deformations of fourth order Peregrine breather solutions of the nonlinear Schrödinger equation. Journal of Mathematical Physics, 54(7):073519, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bayin:2013:CCS

[1931] Selçuk S. Bayin. Comment on “On the consistency of the solutions of the space fractional Schrödinger equation” [J. Math. Phys. 53, 042105 (2012)]. Journal of Mathematical Physics, 54(7):074101, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1761].

Kim:2013:ECC

[1932] Min Kyu Kim and Do-Hyung Kim. Erratum: “A characterization of causal automorphisms by wave equations” [J. Math. Phys. 53, 032507 (2012)]. Journal of Mathematical Physics, 54(7):079901, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1188].

delaLlave:2013:PNB

[1933] Rafael de la Llave and Enrico Valdinoci. Publisher’s note: “L^{p}-bounds for quasi-geostrophic equations via functional analysis” [J. Math. Phys. 52, 083101 (2011)]. Journal of Mathematical Physics, 54(7):079902, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1395].

Lunasin:2013:EOM

[1934] Evelyn Lunasin, Zhi Lin, Alexei Novikov, Anna Mazzucato, and Charles R. Doering. Erratum: “Optimal mixing and optimal stirring for fixed energy, fixed power, or fixed palenstrophy flows” [J. Math. Phys. 53, 115611 (2012)]. Journal of Mathematical Physics, 54(7):079903, July 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1520].

Cohl:2013:PNS

[1935] H. S. Cohl and H. Volkmer. Publisher’s note: “Separation of variables in an asymmetric cyclidic coordinate system” [J. Math. Phys. 54, 063513 (2013)]. Journal of Mathematical Physics, 54(7):079904, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1884].

Amosov:2013:HTC

[1936] Andrey Amosov, Ilya Kostin, Grigory Panasenko, and Valery P. Smyshlyaev. Homogenization of a thermo-chemo-viscoelastic Kelvin– Voigt model. Journal of Mathematical Physics, 54(8):081501, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mancas:2013:WTW

[1937] Stefan C. Mancas, Greg Spradlin, and Harihar Khanal. Weierstrass traveling wave solutions for dissipative Benjamin, Bona, and Mahony (BBM) equation. Journal of Mathematical Physics, 54(8):081502, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2013:LTB

[1938] Zaihong Jiang, Yong Zhou, and Mingxuan Zhu. Large time behavior for the support of momentum density of the Camassa–Holm equation. Journal of Mathematical Physics, 54(8):081503, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakos:2013:DST

[1939] M. Dimakos and A. S. Fokas. Davey–Stewartson type equations in 4 + 2 and 3 + 1 possessing soliton solutions. Journal of Mathematical Physics, 54(8):081504, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bates:2013:TRA

[1940] Peter W. Bates, Kening Lu, and Bixiang Wang. Tempered random attractors for parabolic equations in weighted spaces. Journal of Mathematical Physics, 54(8):081505, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ye:2013:UWD

[1941] Yi-Chao Ye and Zi-Xiang Zhou. A universal way to determine Hirota’s bilinear equation of KdV type. Journal of Mathematical Physics, 54 (8):081506, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2013:RBE

[1942] Weiwei Hu, Igor Kukavica, and Mohammed Ziane. On the regularity for the Boussinesq equations in a bounded domain. Journal of Mathematical Physics, 54(8):081507, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ye:2013:MSK

[1943] Yiwei Ye and Chun-Lei Tang. Multiple solutions for Kirchhoff-type equations in R^{N}. Journal of Mathematical Physics, 54(8):081508, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mudrov:2013:NLC

[1944] Andrey Mudrov. Non-Levi closed conjugacy classes of SO_{q}(N). Journal of Mathematical Physics, 54(8):081701, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hrivnak:2013:GCG

[1945] Jirí Hrivnák and Petr Novotný. Graded contractions of the Gell-Mann graded sl(3, C). Journal of Mathematical Physics, 54(8):081702, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wangberg:2013:DRL

[1946] Aaron Wangberg and Tevian Dray. Discovering real Lie subalgebras of e_{6} using Cartan decompositions. Journal of Mathematical Physics, 54 (8):081703, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naaijkens:2013:KLI

[1947] Pieter Naaijkens. Kosaki–Longo index and classification of charges in 2D quantum spin models. Journal of Mathematical Physics, 54(8):081901, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bishop:2013:GSE

[1948] M. Bishop and J. Wehr. Ground state energy of mean field model of interacting bosons in Bernoulli potential. Journal of Mathematical Physics, 54(8):081902, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-Pascual:2013:RAQ

[1949] Eric Martínez-Pascual. Refined algebraic quantisation in a system with nonconstant gauge invariant structure functions. Journal of Mathematical Physics, 54(8):082101, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pupyshev:2013:HSG

[1950] Vladimir I. Pupyshev and H. E. Montgomery, Jr. How to spoil a good basis set for Rayleigh–Ritz calculations. Journal of Mathematical Physics, 54(8):082102, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bellucci:2013:CPS

[1951] Stefano Bellucci and Vahagn Yeghikyan. The Coulomb problem on a 3-sphere and Heun polynomials. Journal of Mathematical Physics, 54 (8):082103, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantile:2013:WOS

[1952] Andrea Mantile. Wave operators, similarity and dynamics for a class of Schrödinger operators with generic non-mixed interface conditions in 1D. Journal of Mathematical Physics, 54(8):082104, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vourdas:2013:CHA

[1953] A. Vourdas. The complete Heyting algebra of subsystems and contextuality. Journal of Mathematical Physics, 54(8):082105, August 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2182].

Brandon:2013:SPP

[1954] David Brandon, Nasser Saad, and Shi-Hai Dong. On some polynomial potentials in d-dimensions. Journal of Mathematical Physics, 54(8):082106, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenzi:2013:TDS

[1955] E. K. Lenzi, H. V. Ribeiro, M. A. F. dos Santos, R. Rossato, and R. S. Mendes. Time dependent solutions for a fractional Schrödinger equation with delta potentials. Journal of Mathematical Physics, 54(8):082107, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gross:2013:SII

[1956] David Gross and Michael Walter. Stabilizer information inequalities from phase space distributions. Journal of Mathematical Physics, 54 (8):082201, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigi:2013:NQD

[1957] Salman Beigi. A new quantum data processing inequality. Journal of Mathematical Physics, 54(8):082202, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablinger:2013:AAA

[1958] Jakob Ablinger, Johannes Blümlein, and Carsten Schneider. Analytic and algorithmic aspects of generalized harmonic sums and polylogarithms. Journal of Mathematical Physics, 54(8):082301, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guadagnini:2013:TMI

[1959] Enore Guadagnini and Frank Thuillier. Three-manifold invariant from functional integration. Journal of Mathematical Physics, 54(8):082302, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geloun:2013:SCR

[1960] Joseph Ben Geloun and Etera R. Livine. Some classes of renormalizable tensor models. Journal of Mathematical Physics, 54(8):082303, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krisch:2013:TFS

[1961] J. P. Krisch and E. N. Glass. Two fluid shear-free composites. Journal of Mathematical Physics, 54(8):082501, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2013:GPM

[1962] Manuel Núñez. On the gravitational potential of modified Newtonian dynamics. Journal of Mathematical Physics, 54(8):082502, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:EUA

[1963] Huabin Chen. Existence and uniqueness, attraction for stochastic agestructured population systems with diffusion and Poisson jump. Journal of Mathematical Physics, 54(8):082701, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2013:QPS

[1964] Hongyu Cheng and Jianguo Si. Quasi-periodic solutions for the quasiperiodically forced cubic complex Ginzburg–Landau equation on T^{d}. Journal of Mathematical Physics, 54(8):082702, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2013:PAR

[1965] Jingyu Wang and Yejuan Wang. Pullback attractors for reactiondiffusion delay equations on unbounded domains with non-autonomous deterministic and stochastic forcing terms. Journal of Mathematical Physics, 54(8):082703, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avendano-Camacho:2013:HOC

[1966] M. Avendaño-Camacho, J. A. Vallejo, and Yu. Vorobiev. Higher order corrections to adiabatic invariants of generalized slow-fast Hamiltonian systems. Journal of Mathematical Physics, 54(8):082704, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dung:2013:NSC

[1967] Nguyen Tien Dung. New stability conditions for mixed linear Levin– Nohel integro-differential equations. Journal of Mathematical Physics, 54(8):082705, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vankerschaver:2013:GFL

[1968] Joris Vankerschaver, Cuicui Liao, and Melvin Leok. Generating functionals and Lagrangian partial differential equations. Journal of Mathematical Physics, 54(8):082901, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bihlo:2013:CPS

[1969] Alexander Bihlo and George Bluman. Conservative parameterization schemes. Journal of Mathematical Physics, 54(8):083101, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ueltschi:2013:RLR

[1970] Daniel Ueltschi. Random loop representations for quantum spin systems. Journal of Mathematical Physics, 54(8):083301, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:2013:ULM

[1971] Moongyu Park. Upscaling Lévy motions in porous media with long range correlations. Journal of Mathematical Physics, 54(8):083302, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:NLM

[1972] Lun Zhang. A note on the limiting mean distribution of singular values for products of two Wishart random matrices. Journal of Mathematical Physics, 54(8):083303, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2013:CEB

[1973] Shuai-Xia Xu and Yu-Qiu Zhao. Critical edge behavior in the modified Jacobi ensemble and the Painlevé V transcendents. Journal of Mathematical Physics, 54(8):083304, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hakova:2013:FFW

[1974] Lenka Háková, Jirí Hrivnák, and Jirí Patera. Four families of Weyl group orbit functions of B_{3} and C_{3}. Journal of Mathematical Physics, 54(8): 083501, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:2013:PNC

[1975] Yury G. Smirnov and Dmitry V. Valovik. Problem of nonlinear coupled electromagnetic TE–TE wave propagation. Journal of Mathematical Physics, 54(8):083502, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:HPA

[1976] Haiyan Chen and Fuji Zhang. Height probabilities in the Abelian sandpile model on the generalized finite Bethe lattice. Journal of Mathematical Physics, 54(8):083503, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2013:UCS

[1977] Taekyun Kim, Dae San Kim, Toufik Mansour, Seog-Hoon Rim, and Matthias Schork. Umbral calculus and Sheffer sequences of polynomials. Journal of Mathematical Physics, 54(8):083504, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hofmann:2013:RTD

[1978] M. Hofmann, G. Rudolph, and M. Schmidt. On the reflection type decomposition of the adjoint reduced phase space of a compact semisimple Lie group. Journal of Mathematical Physics, 54(8):083505, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2013:RRM

[1979] Satoru Odake. Recurrence relations of the multi-indexed orthogonal polynomials. Journal of Mathematical Physics, 54(8):083506, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubbs:2013:BWE

[1980] Alexander Dubbs, Alan Edelman, Plamen Koev, and Praveen Venkataramana. The beta-Wishart ensemble. Journal of Mathematical Physics, 54 (8):083507, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guedes:2013:QMA

[1981] Carlos Guedes, Daniele Oriti, and Matti Raasakka. Quantization maps, algebra representation, and non-commutative Fourier transform for Lie groups. Journal of Mathematical Physics, 54(8):083508, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2013:SHV

[1982] Guoliang He, Yunyun Zhai, and Xianguo Geng. A super hierarchy of the vector nonlinear Schrödinger equations and Hamiltonian structures, conservation laws. Journal of Mathematical Physics, 54(8):083509, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yolcu:2013:PNE

[1983] Selma Yildirim Yolcu and Türkay Yolcu. Publisher’s note: “Estimates on the eigenvalues of the clamped plate problem on domains in Euclidean spaces” [J. Math. Phys. 54, 043515 (2013)]. Journal of Mathematical Physics, 54(8):089901, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1793].

deLucas:2013:ELS

[1984] J. de Lucas and C. Sardón. Erratum: “On Lie systems and Kummer– Schwarz equations” [J. Math. Phys. 54, 033505 (2013)]. Journal of Mathematical Physics, 54(8):089902, August 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1739, 2020].

Song:2013:SSF

[1985] Yukun Song, Hongjun Yuan, Yang Chen, and Zhidong Guo. Strong solutions for a 1D fluid-particle interaction non-Newtonian model: The bubbling regime. Journal of Mathematical Physics, 54(9):091501, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leach:2013:QPP

[1986] J. A. Leach. A quarter-plane problem for the modified Burgers’ equation. Journal of Mathematical Physics, 54(9):091502, September 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yagdjian:2013:HPK

[1987] Karen Yagdjian. Huygens’ principle for the Klein–Gordon equation in the de Sitter spacetime. Journal of Mathematical Physics, 54(9):091503, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2013:ACH

[1988] Henry H. Kim and Kyu-Hwan Lee. Automorphic correction of the hyperbolic Kac–Moody algebra E_{10}. Journal of Mathematical Physics, 54(9): 091701, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chuah:2013:CSP

[1989] Meng-Kiat Chuah. Contragredient supersymmetric pairs. Journal of Mathematical Physics, 54(9):091702, September 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2712].

Bayin:2013:CPS

[1990] Selçuk S. Bayin. Consistency problem of the solutions of the space fractional Schrödinger equation. Journal of Mathematical Physics, 54(9): 092101, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dernek:2013:QCS

[1991] Mustafa Dernek and Nuri Ünal. Quasi-coherent states for damped and forced harmonic oscillator. Journal of Mathematical Physics, 54(9):

092102, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanaka:2013:TBF

[1992] Toshiaki Tanaka. Type B 3-fold supersymmetry and non-polynomial invariant subspaces. Journal of Mathematical Physics, 54(9):092103, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:AAM

[1993] Hong-Biao Zhang, Lin Yang, and San-Xing Guo. so (3) algebraic approach to the Morse potential. Journal of Mathematical Physics, 54(9): 092104, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boonserm:2013:BVL

[1994] Petarpa Boonserm and Matt Visser. Bounds on variable-length compound jumps. Journal of Mathematical Physics, 54(9):092105, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maciazek:2013:HMI

[1995] Tomasz Maciazek, Michal Oszmaniec, and Adam Sawicki. How many invariant polynomials are needed to decide local unitary equivalence of qubit states? Journal of Mathematical Physics, 54(9):092201, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Descamps:2013:ADL

[1996] Benoît Descamps. Asymptotically decreasing Lieb–Robinson velocity for a class of dissipative quantum dynamics. Journal of Mathematical Physics, 54(9):092202, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shakeel:2013:WQC

[1997] Asif Shakeel and Peter J. Love. When is a quantum cellular automaton (QCA) a quantum lattice gas automaton (QLGA)? Journal of Mathematical Physics, 54(9):092203, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimock:2013:RGA

[1998] J. Dimock. The renormalization group according to Balaban. II. Large fields. Journal of Mathematical Physics, 54(9):092301, September 2013.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khuri:2013:RPM

[1999] Marcus A. Khuri and Gilbert Weinstein. Rigidity in the positive mass theorem with charge. Journal of Mathematical Physics, 54(9):092501, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xie:2013:ASR

[2000] Yongqin Xie, Qingsong Li, Chuangxia Huang, and Yingjun Jiang. Attractors for the semilinear reaction-diffusion equation with distribution derivatives. Journal of Mathematical Physics, 54(9):092701, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Menous:2013:DSR

[2001] Frédéric Menous. From dynamical systems to renormalization. Journal of Mathematical Physics, 54(9):092702, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Novruzov:2013:BPW

[2002] Emil Novruzov. Blow-up phenomena for the weakly dissipative Dullin– Gottwald–Holm equation. Journal of Mathematical Physics, 54(9): 092703, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yaremko:2013:ESL

[2003] Yurij Yaremko. Exact solution to the Landau–Lifshitz equation in a constant electromagnetic field. Journal of Mathematical Physics, 54(9): 092901, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaisman:2013:HVF

[2004] Izu Vaisman. Hamiltonian vector fields on almost symplectic manifolds. Journal of Mathematical Physics, 54(9):092902, September 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carrisi:2013:ETCb

[2005] M. C. Carrisi and S. Pennisi. Extended thermodynamics of charged gases with many moments: an alternative closure. Journal of Mathematical Physics, 54(9):093101, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:2013:UMS

[2006] Paulo R. S. Carvalho and Marcelo M. Leite. Unconventional minimal subtraction and Bogoliubov–Parasyuk–Hepp–Zimmermann method: Massive scalar theory and critical exponents. Journal of Mathematical Physics, 54(9):093301, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2713].

Romaguera:2013:VMA

[2007] Antonio R. de C. Romaguera and K. J. S. Silva. Variational method applied to two-component Ginzburg–Landau theory. Journal of Mathematical Physics, 54(9):093501, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iyela:2013:SQM

[2008] Daddy Balondo Iyela, Jan Govaerts, and M. Norbert Hounkonnou. Supersymmetric quantum mechanics: Engineered hierarchies of integrable potentials and related orthogonal polynomials. Journal of Mathematical Physics, 54(9):093502, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bradshaw:2013:ECP

[2009] Z. Bradshaw and Z. Grujić. Energy cascades in physical scales of 3D incompressible magnetohydrodynamic turbulence. Journal of Mathematical Physics, 54(9):093503, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bluman:2013:SBM

[2010] George W. Bluman and Zhengzheng Yang. A symmetry-based method for constructing nonlocally related partial differential equation systems. Journal of Mathematical Physics, 54(9):093504, September 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dittrich:2013:CAV

[2011] Bianca Dittrich and Philipp A. Höhn. Constraint analysis for variational discrete systems. Journal of Mathematical Physics, 54(9):093505, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2013:CRS

[2012] N. Aizawa, Z. Kuznetsova, and F. Toppan. Chiral and real N = 2 supersymmetric ℓ-conformal Galilei algebras. Journal of Mathematical Physics, 54(9):093506, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colombo:2013:HOD

[2013] Leonardo Colombo, David Martín de Diego, and Marcela Zuccalli. Higher-order discrete variational problems with constraints. Journal of Mathematical Physics, 54(9):093507, September 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dechant:2013:AEN

[2014] Pierre-Philippe Dechant, Céline Boehm, and Reidun Twarock. Affine extensions of non-crystallographic Coxeter groups induced by projection. Journal of Mathematical Physics, 54(9):093508, September 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [2064].

Unterberger:2013:MAR

[2015] André Unterberger. Mathematical analysis as required by a tachyonic observer. Journal of Mathematical Physics, 54(9):093509, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAzcarraga:2013:LAL

[2016] J. A. de Azcárraga and J. M. Izquierdo. k-Leibniz algebras from lower order ones: From Lie triple to Lie l-ple systems. Journal of Mathematical Physics, 54(9):093510, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kondej:2013:SOS

[2017] Sylwia Kondej. Schrödinger operator with a strong varying interaction on a curve in R^{2}. Journal of Mathematical Physics, 54(9):093511, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sadeghi:2013:MES

[2018] J. Sadeghi, M. Rostami, and Z. Sadeghi. The modified equation for spinless particles and superalgebra. Journal of Mathematical Physics, 54 (9):093512, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLange:2013:EBC

[2019] O. L. de Lange and R. E. Raab. Electromagnetic boundary conditions in multipole theory. Journal of Mathematical Physics, 54(9):093513, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia-Pelayo:2013:EMD

[2020] Ricardo García-Pelayo. Erratum: “Moments of the distribution of distance” [J. Math. Phys. 52, 033505 (2011)]. Journal of Mathematical Physics, 54(9):099901, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1984, 1739].

Boumali:2013:EEO

[2021] Abdelmalek Boumali. Erratum: “On the eigensolutions of onedimensional Duffin–Kemmer–Petiau oscillator” [J. Math. Phys. 49, 022302 (2008)]. Journal of Mathematical Physics, 54(9):099902, September 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1681].

Chiba:2013:RWN

[2022] Hayato Chiba. Reduction of weakly nonlinear parabolic partial differential equations. Journal of Mathematical Physics, 54(10):101501, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:GSS

[2023] Jian Zhang, Xianhua Tang, and Wen Zhang. Ground state solutions for nonperiodic Dirac equation with superquadratic nonlinearity. Journal of Mathematical Physics, 54(10):101502, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kafini:2013:EDR

[2024] Mohammad Kafini, Salim A. Messaoudi, and Muhammad I. Mustafa. Energy decay result in a Timoshenko-type system of thermoelasticity of type III with distributive delay. Journal of Mathematical Physics, 54 (10):101503, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwaki:2013:SMQ

[2025] Kohei Iwaki and Atsushi Takahashi. Stokes matrices for the quantum cohomologies of a class of orbifold projective lines. Journal of Mathematical Physics, 54(10):101701, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osei:2013:CMS

[2026] Prince K. Osei and Bernd J. Schroers. Classical r-matrices via semidualisation. Journal of Mathematical Physics, 54(10):101702, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calef:2013:OAD

[2027] M. Calef, W. Griffiths, A. Schulz, C. Fichtl, and D. Hardin. Observed asymptotic differences in energies of stable and minimal point configurations on S^{2} and the role of defects. Journal of Mathematical Physics, 54 (10):101901, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Altaisky:2013:QNW

[2028] M. V. Altaisky and N. E. Kaputkina. On quantization of nondispersive wave packets. Journal of Mathematical Physics, 54(10):102101, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marquette:2013:NLO

[2029] Ian Marquette and Christiane Quesne. New ladder operators for a rational extension of the harmonic oscillator and superintegrability of some two-dimensional systems. Journal of Mathematical Physics, 54 (10):102102, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovacik:2013:VOQ

[2030] Samuel Kovácik and Peter Presnajder. The velocity operator in quantum mechanics in noncommutative space. Journal of Mathematical Physics, 54(10):102103, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:SQR

[2031] Yao-Zhong Zhang. On the solvability of the quantum Rabi model and its 2-photon and two-mode generalizations. Journal of Mathematical Physics, 54(10):102104, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barakat:2013:SBS

[2032] T. Barakat and M. Sebawe Abdalla. Simple bound state energy spectra solutions of Dirac equation for Hulthén and Eckart potentials. Journal of Mathematical Physics, 54(10):102105, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2013:NPE

[2033] Satoru Odake and Ryu Sasaki. Non-polynomial extensions of solvable potentials à la Abraham–Moses. Journal of Mathematical Physics, 54 (10):102106, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kastoryano:2013:RMI

[2034] Michael J. Kastoryano and Jens Eisert. Rapid mixing implies exponential decay of correlations. Journal of Mathematical Physics, 54(10):102201, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruskai:2013:RKS

[2035] Mary Beth Ruskai. Remarks on Kim’s strong subadditivity matrix inequality: Extensions and equality conditions. Journal of Mathematical Physics, 54(10):102202, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1548].

Alexandrov:2013:MQK

[2036] Sergei Alexandrov and Sibasish Banerjee. Modularity, quaternion–Kähler spaces, and mirror symmetry. Journal of Mathematical Physics, 54(10): 102301, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dadhich:2013:PPL

[2037] Naresh Dadhich and Josep M. Pons. Probing pure Lovelock gravity by Nariai and Bertotti–Robinson solutions. Journal of Mathematical Physics, 54(10):102501, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2013:CCA

[2038] Do-Hyung Kim. A characterization of causal automorphisms on twodimensional Minkowski spacetime. Journal of Mathematical Physics, 54 (10):102502, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chudecki:2013:KSS

[2039] Adam Chudecki and Maciej Przanowski. Killing symmetries in H-spaces with Λ. Journal of Mathematical Physics, 54(10):102503, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lompay:2013:CDIb

[2040] Robert R. Lompay and Alexander N. Petrov. Covariant differential identities and conservation laws in metric-torsion theories of gravitation. II. Manifestly generally covariant theories. Journal of Mathematical Physics, 54(10):102504, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daRocha:2013:FDS

[2041] Roldão da Rocha, Luca Fabbri, J. M. Hoff da Silva, R. T. Cavalcanti, and J. A. Silva-Neto. Flag-dipole spinor fields in ESK gravities. Journal of Mathematical Physics, 54(10):102505, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bakosi:2013:SDP

[2042] J. Bakosi and J. R. Ristorcelli. A stochastic diffusion process for Lochner’s generalized Dirichlet distribution. Journal of Mathematical Physics, 54(10):102701, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2013:NSD

[2043] F. Calogero and F. Leyvraz. New solvable discrete-time many-body problem featuring several arbitrary parameters. II. Journal of Mathematical Physics, 54(10):102702, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2013:NIC

[2044] Shaoyun Shi and Wenlei Li. Non-integrability of a class of Painlevé IV equations as Hamiltonian systems. Journal of Mathematical Physics, 54 (10):102703, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fecko:2013:SCQ

[2045] M. Fecko. On symmetries and conserved quantities in Nambu mechanics. Journal of Mathematical Physics, 54(10):102901, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marchesin:2013:SRB

[2046] Marcelo Marchesin and Claudio Vidal. Stability in a rhomboidal 5body problem with generalized central forces. Journal of Mathematical Physics, 54(10):102902, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tracy:2013:ASE

[2047] Craig A. Tracy and Harold Widom. The asymmetric simple exclusion process with an open boundary. Journal of Mathematical Physics, 54 (10):103301, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rego-Monteiro:2013:NQE

[2048] M. A. Rego-Monteiro and F. D. Nobre. Nonlinear quantum equations: Classical field theory. Journal of Mathematical Physics, 54(10):103302, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonetto:2013:ECS

[2049] F. Bonetto and M. Loss. Entropic chaoticity for the steady state of a current carrying system. Journal of Mathematical Physics, 54(10): 103303, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hatzinikitas:2013:SPD

[2050] Agapitos N. Hatzinikitas. Spectral properties of the Dirichlet operator ∑_{i=1}^{d}(−∂_{i}^{2})^{s} on domains in d-dimensional Euclidean space. Journal of Mathematical Physics, 54(10):103501, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2013:LQB

[2051] Caishi Wang and Jihong Zhang. Localization of quantum Bernoulli noises. Journal of Mathematical Physics, 54(10):103502, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2013:MSP

[2052] Yuecai Han and Liwei Zhang. Mild solution to parabolic Anderson model in Gaussian and Poisson potential. Journal of Mathematical Physics, 54 (10):103503, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2013:LDS

[2053] Ji Oon Lee and Kevin Schnelli. Local deformed semicircle law and complete delocalization for Wigner matrices with random potential. Journal of Mathematical Physics, 54(10):103504, October 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teo:2013:FTC

[2054] L. P. Teo. Finite temperature Casimir effect on spherical shells in (D + 1)-dimensional spacetime and its high temperature limit. Journal of Mathematical Physics, 54(10):103505, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarov:2013:OML

[2055] E. I. Jafarov and J. Van der Jeugt. The oscillator model for the Lie superalgebra sh(2|2) and Charlier polynomials. Journal of Mathematical Physics, 54(10):103506, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2013:IDL

[2056] T. Skrypnyk. Infinite-dimensional Lie algebras, classical r-matrices, and Lax operators: Two approaches. Journal of Mathematical Physics, 54 (10):103507, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dou:2013:GHS

[2057] Yan-Ni Dou and Hong-Ke Du. Generalizations of the Heisenberg and Schrödinger uncertainty relations. Journal of Mathematical Physics, 54 (10):103508, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2013:SPB

[2058] Wen-Xiu Ma. A spectral problem based on so(3, R) and its associated commuting soliton equations. Journal of Mathematical Physics, 54(10): 103509, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gine:2013:DLC

[2059] Jaume Giné. On the determination of the limit cycles using the harmonic balance method. Journal of Mathematical Physics, 54(10):103510, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shipman:2013:RES

[2060] Stephen P. Shipman and Aaron T. Welters. Resonant electromagnetic scattering in anisotropic layered media. Journal of Mathematical Physics, 54(10):103511, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saygili:2013:TFR

[2061] K. Saygili. Trkalian fields: ray transforms and mini-twistors. Journal of Mathematical Physics, 54(10):103512, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2013:SEF

[2062] Shaowei Liu and Wen-Xiu Ma. The string equation and the τ-function Witt constraints for the discrete Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 54(10):103513, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xia:2013:CRD

[2063] Baoqiang Xia and Ruguang Zhou. Consecutive Rosochatius deformations of the Neumann system. Journal of Mathematical Physics, 54(10): 103514, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dechant:2013:PNA

[2064] Pierre-Philippe Dechant, Céline Boehm, and Reidun Twarock. Publisher’s note: “Affine extensions of non-crystallographic Coxeter groups induced by projection” [J. Math. Phys. 54, 093508 (2013)]. Journal of Mathematical Physics, 54(10):109901, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2014].

Chen:2013:EMB

[2065] Jun Chen and Xing-Bin Pan. An extended magnetostatic Born–Infeld model with a concave lower order term. Journal of Mathematical Physics, 54(11):111501, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khavkine:2013:PCI

[2066] Igor Khavkine. Presymplectic current and the inverse problem of the calculus of variations. Journal of Mathematical Physics, 54(11):111502, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Downes:2013:SAM

[2067] Robert J. Downes, Michael Levitin, and Dmitri Vassiliev. Spectral asymmetry of the massless Dirac operator on a 3-torus. Journal of Mathematical Physics, 54(11):111503, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaikina:2013:AID

[2068] Elena I. Kaikina. Asymptotics for inhomogeneous Dirichlet initialboundary value problem for the nonlinear Schrödinger equation. Journal of Mathematical Physics, 54(11):111504, November 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:SWC

[2069] Zhijie Chen and Wenming Zou. Standing waves for coupled nonlinear Schrödinger equations with decaying potentials. Journal of Mathematical Physics, 54(11):111505, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2013:MTL

[2070] Alexander Figotin and Guillermo Reyes. Multi-transmission-line-beam interactive system. Journal of Mathematical Physics, 54(11):111901, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ellouz:2013:RBE

[2071] Hanen Ellouz, Ines Feki, and Aref Jeribi. On a Riesz basis of exponentials related to the eigenvalues of an analytic operator and application to a non-selfadjoint problem deduced from a perturbation method for sound radiation. Journal of Mathematical Physics, 54(11):112101, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nottale:2013:ECS

[2072] Laurent Nottale and Marie-Noëlle Célérier. Emergence of complex and spinor wave functions in scale relativity. I. Nature of scale variables. Journal of Mathematical Physics, 54(11):112102, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berkolaiko:2013:CTSa

[2073] G. Berkolaiko and J. Kuipers. Combinatorial theory of the semiclassical evaluation of transport moments. I. Equivalence with the random matrix approach. Journal of Mathematical Physics, 54(11):112103, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolte:2013:HKA

[2074] Jens Bolte and Stefan Keppeler. Heat kernel asymptotics for magnetic Schrödinger operators. Journal of Mathematical Physics, 54(11):112104, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kupriyanov:2013:QMC

[2075] V. G. Kupriyanov. Quantum mechanics with coordinate dependent noncommutativity. Journal of Mathematical Physics, 54(11):112105, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sinha:2013:SSC

[2076] Anjana Sinha and R. Roychoudhury. Spectral singularity in confined P T symmetric optical potential. Journal of Mathematical Physics, 54(11): 112106, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2013:EST

[2077] A. Schulze-Halberg and J. R. Morris. An exactly solvable threedimensional nonlinear quantum oscillator. Journal of Mathematical Physics, 54(11):112107, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirokov:2013:RQC

[2078] M. E. Shirokov. Reversibility of a quantum channel: General conditions and their applications to bosonic linear channels. Journal of Mathematical Physics, 54(11):112201, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2013:ASA

[2079] H. Boche and J. Nötzel. Arbitrarily small amounts of correlation for arbitrarily varying quantum channels. Journal of Mathematical Physics, 54(11):112202, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chesi:2013:VLM

[2080] Stefano Chesi, Arthur Jaffe, Daniel Loss, and Fabio L. Pedrocchi. Vortex loops and Majoranas. Journal of Mathematical Physics, 54(11):112203, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Childs:2013:IDL

[2081] Andrew M. Childs, Debbie Leung, Laura Mancinska, and Maris Ozols. Interpolatability distinguishes LOCC from separable von Neumann measurements. Journal of Mathematical Physics, 54(11):112204, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freidel:2013:EES

[2082] Laurent Freidel and Jeff Hnybida. On the exact evaluation of spin networks. Journal of Mathematical Physics, 54(11):112301, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gomes:2013:FLW

[2083] Henrique Gomes. A first look at Weyl anomalies in shape dynamics. Journal of Mathematical Physics, 54(11):112302, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borghero:2013:WSM

[2084] F. Borghero, F. Demontis, and S. Pennisi. Wave speeds in the macroscopic extended model for ultrarelativistic gases. Journal of Mathematical Physics, 54(11):113101, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamano:2013:PSG

[2085] Takuya Yamano. Phase space gradient of dissipated work and information: A role of relative Fisher information. Journal of Mathematical Physics, 54(11):113301, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bufetov:2013:KIS

[2086] Alexey Bufetov. Kerov’s interlacing sequences and random matrices. Journal of Mathematical Physics, 54(11):113302, November 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2013:BAT

[2087] Chuanzhong Li and Jingsong He. Block algebra in two-component BKP and D type Drinfeld–Sokolov hierarchies. Journal of Mathematical Physics, 54(11):113501, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Venturi:2013:CFA

[2088] Daniele Venturi. Conjugate flow action functionals. Journal of Mathematical Physics, 54(11):113502, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon-Cortes:2013:NOP

[2089] L. F. Chacón-Cortes and W. A. Zúñiga-Galindo. Nonlocal operators, parabolic-type equations, and ultrametric random walks. Journal of Mathematical Physics, 54(11):113503, November 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kondratiev:2013:SMS

[2090] Yuri Kondratiev, Oleksandr Kutoviy, Robert Minlos, and Sergey Pirogov. On spatial mutation-selection models. Journal of Mathematical Physics, 54(11):113504, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shamis:2013:DSG

[2091] Mira Shamis. Density of states for Gaussian unitary ensemble, Gaussian orthogonal ensemble, and interpolating ensembles through supersymmetric approach. Journal of Mathematical Physics, 54(11):113505, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:DAC

[2092] Wei Chen and Bing-Sheng Lin. A diagrammatic approach to the categorical coherent state. Journal of Mathematical Physics, 54(11):113506, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonera:2013:NRO

[2093] Joanna Gonera. Nonlinear realizations and the orbit method. Journal of Mathematical Physics, 54(11):113507, November 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chvartatskyi:2013:NBG

[2094] O. I. Chvartatskyi and Yu. M. Sydorenko. A new bidirectional generalization of (2+1)-dimensional matrix k-constrained Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 54(11):113508, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:2013:SDS

[2095] Sam Palmer and Christian Sämann. Six-dimensional (1, 0) superconformal models and higher gauge theory. Journal of Mathematical Physics, 54(11):113509, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Restuccia:2013:HSC

[2096] A. Restuccia and A. Sotomayor. The Hamiltonian structure of a coupled system derived from a supersymmetric breaking of super Korteweg– de Vries equations. Journal of Mathematical Physics, 54(11):113510, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wong:2013:CCM

[2097] Willie Wai-Yeung Wong. A comment on the construction of the maximal globally hyperbolic Cauchy development. Journal of Mathematical Physics, 54(11):113511, November 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2013:MSQ

[2098] Jiaquan Liu, Zhi-Qiang Wang, and Xian Wu. Multibump solutions for quasilinear elliptic equations with critical growth. Journal of Mathematical Physics, 54(12):121501, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shang:2013:FSE

[2099] Xudong Shang, Jihui Zhang, and Yang Yang. On fractional Schrödinger equation in R^{N} with critical growth. Journal of Mathematical Physics, 54(12):121502, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2013:EGW

[2100] Zhenhua Guo, Zilai Li, and Lei Yao. Existence of global weak solution for a reduced gravity two and a half layer model. Journal of Mathematical Physics, 54(12):121503, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2013:GIS

[2101] Stephen C. Anco and Wei Feng. Group-invariant solutions of semilinear Schrödinger equations in multi-dimensions. Journal of Mathematical Physics, 54(12):121504, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:2013:GSH

[2102] Chong Song, Xiaowei Sun, and Youde Wang. Geometric solitons of Hamiltonian flows on manifolds. Journal of Mathematical Physics, 54 (12):121505, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2013:NGS

[2103] Jianjun Yuan. Nonexistence of global solution to Chern–Simons–Higgs system. Journal of Mathematical Physics, 54(12):121506, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parmeggiani:2013:ESC

[2104] Alberto Parmeggiani and Alberto Venni. On the essential spectrum of certain non-commutative oscillators. Journal of Mathematical Physics, 54(12):121507, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2013:IMS

[2105] Weiming Liu and Chunhua Wang. Infinitely many solutions for the nonlinear Schrödinger equations with magnetic potentials in R^{N}. Journal of Mathematical Physics, 54(12):121508, December 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ha:2013:SVM

[2106] Seung-Yeal Ha, Qinghua Xiao, Linjie Xiong, and Huijiang Zhao. L_{2}-stability of the Vlasov–Maxwell–Boltzmann system near global Maxwellians. Journal of Mathematical Physics, 54(12):121509, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2013:EMB

[2107] Sihua Liang and Shaoyun Shi. Existence of multi-bump solutions for a class of Kirchhoff type problems in R^{3}. Journal of Mathematical Physics, 54(12):121510, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2013:MBS

[2108] Daomin Cao and Yiming Su. Minimal blow-up solutions of mass-critical inhomogeneous Hartree equation. Journal of Mathematical Physics, 54(12):121511, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walton:2013:PEL

[2109] Mark A. Walton. Polytope expansion of Lie characters and applications. Journal of Mathematical Physics, 54(12):121701, December 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anguelova:2013:TVA

[2110] Iana I. Anguelova. Twisted vertex algebras, bicharacter construction and boson-fermion correspondences. Journal of Mathematical Physics, 54(12):121702, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaminski:2013:CSS

[2111] Wojciech Kamiński and Sebastian Steinhaus. Coherent states, 6j symbols and properties of the next to leading order asymptotic expansions. Journal of Mathematical Physics, 54(12):121703, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2013:LSP

[2112] Junli Liu and Shilin Yang. Lusztig symmetries and Poincaré–Birkhoff– Witt basis for wU_{r,s}^{d}(osp(1|2n)). Journal of Mathematical Physics, 54 (12):121704, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Costa:2013:WIC

[2113] L. C. Costa, A. Ferraz, and Vieri Mastropietro. Ward identities and chiral anomalies for coupled fermionic chains. Journal of Mathematical Physics, 54(12):121901, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlowski:2013:AIP

[2114] K. K. Kozlowski. Aspects of the inverse problem for the Toda chain. Journal of Mathematical Physics, 54(12):121902, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Safronov:2013:ACS

[2115] Oleg Safronov. Absolutely continuous spectrum of the Schrödinger operator with a potential representable as a sum of three functions with special properties. Journal of Mathematical Physics, 54(12):122101, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2013:APA

[2116] Hong-Biao Zhang and Lu Lu. Application of polynomial su(1, 1) algebra to Pöschl–Teller potentials. Journal of Mathematical Physics, 54(12): 122102, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:2013:ERP

[2117] C. Quesne. Extending Romanovski polynomials in quantum mechanics. Journal of Mathematical Physics, 54(12):122103, December 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2013:REJ

[2118] Axel Schulze-Halberg and Barnana Roy. Rational extension and Jacobitype X_{m} solutions of a quantum nonlinear oscillator. Journal of Mathematical Physics, 54(12):122104, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xiao:2013:IBS

[2119] Shuyuan Xiao, Zhixuan Deng, and Hong Chen. Independent bases on the spatial wavefunction of four-identical-particle systems. Journal of Mathematical Physics, 54(12):122105, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galikova:2013:LRL

[2120] Veronika Gáliková, Samuel Kovácik, and Peter Presnajder. Laplace– Runge–Lenz vector in quantum mechanics in noncommutative space. Journal of Mathematical Physics, 54(12):122106, December 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koupaei:2013:QBA

[2121] Jalaledin Yousefi Koupaei and Mehdi Golshani. A quasi-Bohmian approach for a homogeneous spherical solid body based on its geometric structure. Journal of Mathematical Physics, 54(12):122107, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christiansen:2013:SPD

[2122] H. R. Christiansen and M. S. Cunha. Solutions to position-dependent mass quantum mechanics for a new class of hyperbolic potentials. Journal of Mathematical Physics, 54(12):122108, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:2013:SLQ

[2123] Stefano De Leo and Gisele C. Ducati. The Snell law for quaternionic potentials. Journal of Mathematical Physics, 54(12):122109, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Temme:2013:LBS

[2124] Kristan Temme. Lower bounds to the spectral gap of Davies generators. Journal of Mathematical Physics, 54(12):122110, December 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barone:2013:QCT

[2125] Fiorella Barone and Sandro Graffi. A quantum Cherry theorem for perturbations of the plane rotator. Journal of Mathematical Physics, 54(12): 122111, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rakityansky:2013:ASM

[2126] S. A. Rakityansky and N. Elander. Analytic structure of the multichannel Jost matrix for potentials with Coulombic tails. Journal of Mathematical Physics, 54(12):122112, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2013:GMP

[2127] Q. H. Liu. Geometric momentum for a particle constrained on a curved hypersurface. Journal of Mathematical Physics, 54(12):122113, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frank:2013:MRR

[2128] Rupert L. Frank and Elliott H. Lieb. Monotonicity of a relative Rényi entropy. Journal of Mathematical Physics, 54(12):122201, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beigi:2013:SRD

[2129] Salman Beigi. Sandwiched Rényi divergence satisfies data processing inequality. Journal of Mathematical Physics, 54(12):122202, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muller-Lennert:2013:QRE

[2130] Martin Müller-Lennert, Frédéric Dupuis, Oleg Szehr, Serge Fehr, and Marco Tomamichel. On quantum Rényi entropies: A new generalization and some properties. Journal of Mathematical Physics, 54(12):122203, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tomboulis:2013:CSR

[2131] E. T. Tomboulis. Chiral symmetry restoration at large chemical potential in strongly coupled SU(N) gauge theories. Journal of Mathematical Physics, 54(12):122301, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2184].

DiBartolo:2013:MDA

[2132] Cayetano Di Bartolo, Javier Grau, and Lorenzo Leal. Mimetic discretization of the Abelian Chern–Simons theory and link invariants. Journal of Mathematical Physics, 54(12):122302, December 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cegla:2013:NML

[2133] Wojciech Cegla and Bernard Jancewicz. Non-modular lattices generated by the causal structure. Journal of Mathematical Physics, 54(12):122501, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corsi:2013:DAR

[2134] Livia Corsi, Roberto Feola, and Guido Gentile. Domains of analyticity for response solutions in strongly dissipative forced systems. Journal of Mathematical Physics, 54(12):122701, December 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cho:2013:RHS

[2135] Hyunjoo Cho, Sema Salur, and A. J. Todd. Remarks on Hamiltonian structures in G_{2}-geometry. Journal of Mathematical Physics, 54(12): 122901, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2013:ETP

[2136] Feng Chen and Yuecai Han. Existence of time-periodic weak solutions to the stochastic Navier–Stokes equations around a moving body. Journal of Mathematical Physics, 54(12):123101, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:2013:SGB

[2137] H. Bergeron, E. M. F. Curado, J. P. Gazeau, and Ligia M. C. S. Rodrigues. Symmetric generalized binomial distributions. Journal of Mathematical Physics, 54(12):123301, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tracy:2013:DST

[2138] Craig A. Tracy and Harold Widom. On the diagonal susceptibility of the two-dimensional Ising model. Journal of Mathematical Physics, 54(12): 123302, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moon:2013:SPT

[2139] W. Moon and J. S. Wettlaufer. A stochastic perturbation theory for nonautonomous systems. Journal of Mathematical Physics, 54(12):123303, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2013:PPP

[2140] Jibin Li and Zhijun Qiao. Peakon, pseudo-peakon, and cuspon solutions for two generalized Camassa–Holm equations. Journal of Mathematical Physics, 54(12):123501, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amdeberhan:2013:SPB

[2141] Tewodros Amdeberhan, Valerio De Angelis, Atul Dixit, Victor H. Moll, and Christophe Vignat. From sequences to polynomials and back, via operator orderings. Journal of Mathematical Physics, 54(12):123502, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biggs:2013:OPP

[2142] A. Biggs and H. M. Khudaverdian. Operator pencil passing through a given operator. Journal of Mathematical Physics, 54(12):123503, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Livine:2013:DPP

[2143] Etera R. Livine. Deformations of polyhedra and polygons by the unitary group. Journal of Mathematical Physics, 54(12):123504, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berkolaiko:2013:CTSb

[2144] G. Berkolaiko and J. Kuipers. Combinatorial theory of the semiclassical evaluation of transport moments II: Algorithmic approach for moment generating functions. Journal of Mathematical Physics, 54(12):123505, December 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikitin:2013:LRL

[2145] A. G. Nikitin. Laplace–Runge–Lenz vector for arbitrary spin. Journal of Mathematical Physics, 54(12):123506, December 2013. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaisman:2013:TDF

[2146] Izu Vaisman. Towards a double field theory on para-Hermitian manifolds. Journal of Mathematical Physics, 54(12):123507, December 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Briane:2014:FBE

[2147] Marc Briane and Muthusamy Vanninathan. First Bloch eigenvalue in high contrast media. Journal of Mathematical Physics, 55(1):011501, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Intissar:2014:SNC

[2148] Abdelkader Intissar. A short note on the chaoticity of a weight shift on concrete orthonormal basis associated to some Fock–Bargmann space. Journal of Mathematical Physics, 55(1):011502, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurt:2014:PUS

[2149] Levent Kurt and Tobias Schäfer. Propagation of ultra-short solitons in stochastic Maxwell’s equations. Journal of Mathematical Physics, 55 (1):011503, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geisinger:2014:SPW

[2150] Leander Geisinger. A short proof of Weyl’s law for fractional differential operators. Journal of Mathematical Physics, 55(1):011504, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2014:GSN

[2151] Tai-Chia Lin, Milivoj R. Belić, Milan S. Petrović, and Goong Chen. Ground states of nonlinear Schrödinger systems with saturable nonlinearity in R^{2} for two counterpropagating beams. Journal of Mathematical Physics, 55(1):011505, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmkvist:2014:THA

[2152] Jakob Palmkvist. The tensor hierarchy algebra. Journal of Mathematical Physics, 55(1):011701, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dherin:2014:SDA

[2153] Benoit Dherin and Igor Mencattini. G-systems and deformation of Gactions on R^{d}. Journal of Mathematical Physics, 55(1):011702, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:2014:MET

[2154] Mark D. Gould, Phillip S. Isaac, and Jason L. Werry. Matrix elements for type 1 o unitary irreducible representations of the Lie superalgebra gl(m|n). Journal of Mathematical Physics, 55(1):011703, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cantarini:2014:AVP

[2155] Nicoletta Cantarini and Victor G. Kac. Algebraic vs physical N = 6 3 xo-algebras. Journal of Mathematical Physics, 55(1):011704, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laber:2014:GVA

[2156] Rob Laber and Geoffrey Mason. C-graded vertex algebras and conformal flow. Journal of Mathematical Physics, 55(1):011705, January 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2014:LVL

[2157] Henan Wu, Qiufan Chen, and Xiaoqing Yue. Loop Virasoro Lie conformal algebra. Journal of Mathematical Physics, 55(1):011706, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bach:2014:GLV

[2158] Volker Bach, Sébastien Breteaux, Hans Konrad Knörr, and Edmund Menge. Generalization of Lieb’s variational principle to Bogoliubov– Hartree–Fock theory. Journal of Mathematical Physics, 55(1):012101, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roy:2014:OPS

[2159] Arunabha S. Roy and S. M. Roy. Optimum phase space probabilities from quantum tomography. Journal of Mathematical Physics, 55(1):012102, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pachon:2014:QDD

[2160] Leonardo A. Pachón and Paul Brumer. Quantum driven dissipative parametric oscillator in a blackbody radiation field. Journal of Mathematical Physics, 55(1):012103, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Recchia:2014:SWP

[2161] Carla Recchia and Alessandro Teta. Semiclassical wave-packets emerging from interaction with an environment. Journal of Mathematical Physics, 55(1):012104, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hartmann:2014:BSH

[2162] R. R. Hartmann. Bound states in a hyperbolic asymmetric double-well. Journal of Mathematical Physics, 55(1):012105, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tare:2014:BSM

[2163] Jeffrey D. Tare and Jose Perico H. Esguerra. Bound states for multiple Dirac-δ wells in space-fractional quantum mechanics. Journal of Mathematical Physics, 55(1):012106, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2294].

Ali:2014:SSH

[2164] S. Twareque Ali, K. Górska, A. Horzela, and F. H. Szafraniec. Squeezed states and Hermite polynomials in a complex variable. Journal of Mathematical Physics, 55(1):012107, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2014:CSP

[2165] Zhongwei Shen. Completeness for sparse potential scattering. Journal of Mathematical Physics, 55(1):012108, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stulpe:2014:ACC

[2166] Werner Stulpe. From the attempt of certain classical reformulations of quantum mechanics to quasi-probability representations. Journal of Mathematical Physics, 55(1):012109, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banica:2014:FOD

[2167] Teodor Banica. First order deformations of the Fourier matrix. Journal of Mathematical Physics, 55(1):012201, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holweck:2014:EFQ

[2168] Frédéric Holweck, Jean-Gabriel Luque, and Jean-Yves Thibon. Entanglement of four qubit systems: A geometric atlas with polynomial compass I (the finite world). Journal of Mathematical Physics, 55(1):012202, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saidi:2014:TSY

[2169] El Hassan Saidi. Twisted 3D N = 4 supersymmetric YM on deformed A_{3}^{∗} lattice. Journal of Mathematical Physics, 55(1):012301, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

ONeil:2014:GDS

[2170] Michael O’Neil. A generalized Debye source approach to electromagnetic scattering in layered media. Journal of Mathematical Physics, 55(1): 012901, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Capriotti:2014:DGP

[2171] S. Capriotti. Differential geometry, Palatini gravity and reduction. Journal of Mathematical Physics, 55(1):012902, January 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korhonen:2014:TPP

[2172] Marko Korhonen and Eunghyun Lee. The transition probability and the probability for the left-most particle’s position of the q-totally asymmetric zero range process. Journal of Mathematical Physics, 55(1):013301, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Makhlouf:2014:YDM

[2173] Abdenacer Makhlouf and Florin Panaite. Yetter–Drinfeld modules for Hom-bialgebras. Journal of Mathematical Physics, 55(1):013501, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odake:2014:EMI

[2174] Satoru Odake. Equivalences of the multi-indexed orthogonal polynomials. Journal of Mathematical Physics, 55(1):013502, January 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoine:2014:SRQ

[2175] Jean-Pierre Antoine and Camillo Trapani. Some remarks on quasi-Hermitian operators. Journal of Mathematical Physics, 55(1):013503, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Phan:2014:SMN

[2176] Quang Sang Phan. Spectral monodromy of non-self-adjoint operators. Journal of Mathematical Physics, 55(1):013504, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajaratnam:2014:KTW

[2177] Krishan Rajaratnam and Raymond G. McLenaghan. Killing tensors, warped products and the orthogonal separation of the Hamilton–Jacobi equation. Journal of Mathematical Physics, 55(1):013505, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popowicz:2014:TCC

[2178] Ziemowit Popowicz. Two-component coupled KdV equations and its connection with the generalized Harry Dym equations. Journal of Mathematical Physics, 55(1):013506, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grothaus:2014:WNA

[2179] M. Grothaus, F. Riemann, and H. P. Suryawan. A white noise approach to the Feynman integrand for electrons in random media. Journal of Mathematical Physics, 55(1):013507, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ngendakumana:2014:GTC

[2180] Ancille Ngendakumana, Joachim Nzotungicimpaye, and Leonard Todjihoundé. Group theoretical construction of planar noncommutative phase spaces. Journal of Mathematical Physics, 55(1):013508, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhatia:2014:TIP

[2181] Rajendra Bhatia. Trace inequalities for products of positive definite matrices. Journal of Mathematical Physics, 55(1):013509, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vourdas:2014:ECH

[2182] A. Vourdas. Erratum: “The complete Heyting algebra of subsystems and contextuality” [J. Math. Phys. 54, 082105 (2013)]. Journal of Mathematical Physics, 55(1):019901, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1953].

Garat:2014:ETG

[2183] Alcides Garat. Erratum: “Tetrads in geometrodynamics” [J. Math. Phys. 46, 102502 (2005)]. Journal of Mathematical Physics, 55(1):019902, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [6].

Tomboulis:2014:PNC

[2184] E. T. Tomboulis. Publisher’s note: “Chiral symmetry restoration at large chemical potential 2 in strongly coupled SU(N) gauge theories” [J. Math. Phys. 54, 122301 (2013)]. Journal of Mathematical Physics, 55 (1):019903, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2131].

Bidegaray-Fesquet:2014:NBM

[2185] Brigitte Bidegaray-Fesquet and Kole Keita. A nonlinear Bloch model for Coulomb interaction in quantum dots. Journal of Mathematical Physics, 55(2):021501, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pei:2014:RBO

[2186] Jun Pei, Chengming Bai, and Li Guo. Rota–Baxter operators on sl(2, C) and solutions of the classical Yang–Baxter equation. Journal of Mathematical Physics, 55(2):021701, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keppeler:2014:HYO

[2187] Stefan Keppeler and Malin Sjödahl. Hermitian Young operators. Journal of Mathematical Physics, 55(2):021702, February 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dodd:2014:HEA

[2188] R. K. Dodd. Hirota equations associated with simply laced affine Lie algebras: The cuspidal class E_{6} of ⌉_{6}^{(1)}. Journal of Mathematical Physics, 55 (2):021703, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benedikter:2014:MFD

[2189] Niels Benedikter, Marcello Porta, and Benjamin Schlein. Mean-field dynamics of fermions with relativistic dispersion. Journal of Mathematical Physics, 55(2):021901, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haegeman:2014:GMP

[2190] Jutho Haegeman, Michaël Mariën, Tobias J. Osborne, and Frank Verstraete. Geometry of matrix product states: Metric, parallel transport, and curvature. Journal of Mathematical Physics, 55(2):021902, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Znidaric:2014:TIC

[2191] Marko Znidaric, Giuliano Benenti, and Giulio Casati. Translationally invariant conservation laws of local Lindblad equations. Journal of Mathematical Physics, 55(2):021903, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ekholm:2014:MCH

[2192] T. Ekholm and F. Portmann. A magnetic contribution to the Hardy inequality. Journal of Mathematical Physics, 55(2):022101, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bunao:2014:BDB

[2193] Joseph Bunao and Eric A. Galapon. The Bender–Dunne basis operators as Hilbert space operators. Journal of Mathematical Physics, 55 (2):022102, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:2014:SSM

[2194] Palle E. T. Jorgensen, Keri A. Kornelson, and Karen L. Shuman. Scalar spectral measures associated with an operator-fractal. Journal of Mathematical Physics, 55(2):022103, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanaka:2014:ITB

[2195] Toshiaki Tanaka. GL(3, C) invariance of type B 3-fold supersymmetric systems. Journal of Mathematical Physics, 55(2):022104, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:2014:MQM

[2196] Stephen G. Low. Maximal quantum mechanical symmetry: Projective representations of the inhomogeneous symplectic group. Journal of Mathematical Physics, 55(2):022105, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jencova:2014:BND

[2197] A. Jencová. Base norms and discrimination of generalized quantum channels. Journal of Mathematical Physics, 55(2):022201, February 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:2014:DSQ

[2198] Stefano De Leo and Sergio Giardino. Dirac solutions for quaternionic potentials. Journal of Mathematical Physics, 55(2):022301, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Much:2014:QME

[2199] A. Much. Quantum mechanical effects from deformation theory. Journal of Mathematical Physics, 55(2):022302, February 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Livshits:2014:BGC

[2200] Gideon I. Livshits. Breaking generalized covariance, classical renormalization, and boundary conditions from superpotentials. Journal of Mathematical Physics, 55(2):022501, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanatarov:2014:WHP

[2201] I. V. Tanatarov and O. B. Zaslavskii. What happens to Petrov classification, on horizons of axisymmetric dirty black holes. Journal of Mathematical Physics, 55(2):022502, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cariglia:2014:GEL

[2202] Marco Cariglia and Gary Gibbons. Generalised Eisenhart lift of the Toda chain. Journal of Mathematical Physics, 55(2):022701, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landim:2014:ZMF

[2203] R. R. Landim, L. C. Q. Vilar, O. S. Ventura, and V. E. R. Lemes. On the zero modes of the Faddeev–Popov operator in the Landau gauge. Journal of Mathematical Physics, 55(2):022901, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamocki:2014:FDI

[2204] Rafal Kamocki and Cezary Obczyński. On fractional differential inclusions with the Jumarie derivative. Journal of Mathematical Physics, 55 (2):022902, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kinugawa:2014:GAM

[2205] Tohru Kinugawa. Generalization of Abel’s mechanical problem: The extended isochronicity condition and the superposition principle. Journal of Mathematical Physics, 55(2):022903, February 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sandev:2014:CFF

[2206] Trifce Sandev, Ralf Metzler, and Zivorad Tomovski. Correlation functions for the fractional generalized Langevin equation in the presence of internal and external noise. Journal of Mathematical Physics, 55(2): 023301, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeGregorio:2014:FRW

[2207] Alessandro De Gregorio. A family of random walks with generalized Dirichlet steps. Journal of Mathematical Physics, 55(2):023302, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bao:2014:UGP

[2208] Zhigang Bao, Guangming Pan, and Wang Zhou. Universality for a global property of the eigenvectors of Wigner matrices. Journal of Mathematical Physics, 55(2):023303, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ren:2014:MDP

[2209] Haizhen Ren, Fuji Zhang, and Jianguo Qian. Monomer-dimer problem on random planar honeycomb lattice. Journal of Mathematical Physics, 55 (2):023304, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2014:EUD

[2210] Ruifeng Zhang and Fangfang Li. Existence and uniqueness of domain wall solitons in a Maxwell–Chern–Simons model. Journal of Mathematical Physics, 55(2):023501, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardoso:2014:REP

[2211] Fernando Cardoso, Claudio Cuevas, and Georgi Vodev. Resolvent estimates for perturbations by large magnetic potentials. Journal of Mathematical Physics, 55(2):023502, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2014:GTI

[2212] Peter J. Forrester and Colin J. Thompson. The Golden–Thompson inequality: Historical aspects and random matrix applications. Journal of Mathematical Physics, 55(2):023503, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:GHM

[2213] Chunxiang Wang. Graded Hilbert C∗-modules. Journal of Mathematical Physics, 55(2):023504, February 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chern:2014:SSG

[2214] Jann-Long Chern and Sze-Guang Yang. A survey of solutions in a gravitational Born–Infeld theory. Journal of Mathematical Physics, 55(3): 031501, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shen:2014:RSC

[2215] Chun Shen. On a regularization of a scalar conservation law with discontinuous coefficients. Journal of Mathematical Physics, 55(3):031502, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohameden:2014:WKD

[2216] Ahmedou Yahya Ould Mohameden and Mohamed Vall Ould Moustapha. Wave kernels for the Dirac, Euler operators and the harmonic oscillator. Journal of Mathematical Physics, 55(3):031503, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tang:2014:CPS

[2217] Hao Tang and Zhengrong Liu. Continuous properties of the solution map for the Euler equations. Journal of Mathematical Physics, 55(3):031504, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamazaki:2014:RRC

[2218] Kazuo Yamazaki. Remarks on the regularity criteria of three-dimensional magnetohydrodynamics system in terms of two velocity field components. Journal of Mathematical Physics, 55(3):031505, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biondini:2014:IST

[2219] Gino Biondini and Gregor Kovacic. Inverse scattering transform for the focusing nonlinear Schrödinger equation with nonzero boundary conditions. Journal of Mathematical Physics, 55(3):031506, March 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2014:ECP

[2220] Jianjun Zhang. The existence and concentration of positive solutions for a nonlinear Schrödinger–Poisson system with critical growth. Journal of Mathematical Physics, 55(3):031507, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2014:MPS

[2221] Xiaomei Sun and Yimin Zhang. Multi-peak solution for nonlinear magnetic Choquard type equation. Journal of Mathematical Physics, 55(3): 031508, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2014:RBY

[2222] Ling Liu and Bingliang Shen. Radford’s biproducts and Yetter–Drinfeld modules for monoidal Hom–Hopf algebras. Journal of Mathematical Physics, 55(3):031701, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2014:CYD

[2223] Yuanyuan Chen and Liangyun Zhang. The category of Yetter–Drinfel’d Hom-modules and the quantum Hom–Yang–Baxter equation. Journal of Mathematical Physics, 55(3):031702, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woods:2014:MOQ

[2224] M. P. Woods, R. Groux, A. W. Chin, S. F. Huelga, and M. B. Plenio. Mappings of open quantum systems onto chain representations and Markovian embeddings. Journal of Mathematical Physics, 55(3):032101, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2014:SPF

[2225] Jianping Dong. Scattering problems in the fractional quantum mechanics governed by the 2D space-fractional Schrödinger equation. Journal of Mathematical Physics, 55(3):032102, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2014:SEA

[2226] Tuncay Aktosun, Martin Klaus, and Ricardo Weder. Small-energy analysis for the selfadjoint matrix Schrödinger operator on the half line. II. Journal of Mathematical Physics, 55(3):032103, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tezcan:2014:PAS

[2227] Cevdet Tezcan and Ramazan Sever. A parametric approach to supersymmetric quantum mechanics in the solution of Schrödinger equation. Journal of Mathematical Physics, 55(3):032104, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vega:2014:ODN

[2228] F. Vega. Oscillators in a (2 + 1)-dimensional noncommutative space. Journal of Mathematical Physics, 55(3):032105, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosenfelder:2014:STP

[2229] R. Rosenfelder. Scattering theory with path integrals. Journal of Mathematical Physics, 55(3):032106, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gutoski:2014:PTU

[2230] Gus Gutoski and Nathaniel Johnston. Process tomography for unitary quantum channels. Journal of Mathematical Physics, 55(3):032201, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mozrzymas:2014:SPA

[2231] Marek Mozrzymas, Michal Horodecki, and Michal Studziński. Structure and properties of the algebra of partially transposed permutation operators. Journal of Mathematical Physics, 55(3):032202, March 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geloun:2014:FLB

[2232] Joseph Ben Geloun and John R. Klauder. The finite and large-N behaviors of independent-value matrix models. Journal of Mathematical Physics, 55(3):032301, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petrat:2014:MTS

[2233] Sören Petrat and Roderich Tumulka. Multi-time Schrödinger equations cannot contain interaction potentials. Journal of Mathematical Physics, 55(3):032302, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reid:2014:CHM

[2234] James A. Reid and Charles H.-T. Wang. Conformal holonomy in MacDowell–Mansouri gravity. Journal of Mathematical Physics, 55(3): 032501, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fabi:2014:GFG

[2235] S. Fabi, B. Harms, and S. Hou. On the gauge features of gravity on a Lie algebroid structure. Journal of Mathematical Physics, 55(3):032502, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:2014:SEF

[2236] Avinash Khare and Avadh Saxena. Superposition of elliptic functions as solutions for a large number of nonlinear equations. Journal of Mathematical Physics, 55(3):032701, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2624].

Kamizawa:2014:RAE

[2237] T. Kamizawa, T. Hara, and M. Ohya. On relations among the entropic chaos degree, the Kolmogorov–Sinai entropy and the Lyapunov exponent. Journal of Mathematical Physics, 55(3):032702, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fedorov:2014:PFE

[2238] Yuri Fedorov and Chara Pantazi. The Picard–Fuchs equations for complete hyperelliptic integrals of even order curves, and the actions of the generalized Neumann system. Journal of Mathematical Physics, 55(3): 032703, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Donchev:2014:VSC

[2239] Veliko Donchev. Variational symmetries, conserved quantities and identities for several equations of mathematical physics. Journal of Mathematical Physics, 55(3):032901, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Escobar:2014:IEF

[2240] C. A. Escobar and L. F. Urrutia. Invariants of the electromagnetic field. Journal of Mathematical Physics, 55(3):032902, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babushkin:2014:FSU

[2241] I. Babushkin and L. Bergé. The fundamental solution of the unidirectional pulse propagation equation. Journal of Mathematical Physics, 55 (3):032903, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cortissoz:2014:LBP

[2242] Jean C. Cortissoz, Julio A. Montero, and Carlos E. Pinilla. On lower bounds for possible blow-up solutions to the periodic Navier–Stokes equation. Journal of Mathematical Physics, 55(3):033101, March 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khorrami:2014:PTR

[2243] Mohammad Khorrami and Amir Aghamohammadi. Phase transitions in a reaction-diffusion model on a line with boundaries. Journal of Mathematical Physics, 55(3):033301, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korichi:2014:SMB

[2244] Z. Korichi and M. T. Meftah. Statistical mechanics based on fractional classical and quantum mechanics. Journal of Mathematical Physics, 55 (3):033302, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2014:NSA

[2245] F. Bagarello, A. Inoue, and C. Trapani. Non-self-adjoint Hamiltonians defined by Riesz bases. Journal of Mathematical Physics, 55(3):033501, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hassanabadi:2014:SPK

[2246] H. Hassanabadi, S. S. Hosseini, A. Boumali, and S. Zarrinkamar. The statistical properties of Klein–Gordon oscillator in noncommutative space. Journal of Mathematical Physics, 55(3):033502, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liang:2014:SNW

[2247] Fei Liang and Hongjun Gao. Stochastic nonlinear wave equation with memory driven by compensated Poisson random measures. Journal of Mathematical Physics, 55(3):033503, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popov:2014:EPS

[2248] I. Yu. Popov, A. N. Skorynina, and I. V. Blinova. On the existence of point spectrum for branching strips quantum graph. Journal of Mathematical Physics, 55(3):033504, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Azreg-Ainou:2014:CMS

[2249] Mustapha Azreg-Aïnou. On “The conformal metric structure of geometrothermodynamics”: Generalizations. Journal of Mathematical Physics, 55(3):033505, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoang:2014:ABS

[2250] Vu Hoang and Maria Radosz. Absence of bound states for waveguides in two-dimensional periodic structures. Journal of Mathematical Physics, 55(3):033506, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2014:DID

[2251] Jaume Llibre and Claudia Valls. Darboux integrability of 2-dimensional Hamiltonian systems with homogeneous potentials of degree 3. Journal of Mathematical Physics, 55(3):033507, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheviakov:2014:CPP

[2252] Alexei F. Cheviakov. Conservation properties and potential systems of vorticity-type equations. Journal of Mathematical Physics, 55(3):033508, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rabei:2014:CME

[2253] Eqab M. Rabei, A. Al-Jamel, H. Widyan, and D. Baleanu. Comment on “Maxwell’s equations and electromagnetic Lagrangian density in fractional form” [J. Math. Phys. 53, 033505 (2012)]. Journal of Mathematical Physics, 55(3):034101, March 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1199].

Nachtergaele:2014:ARA

[2254] Bruno Nachtergaele. Announcement: Referee acknowledgment for 2013. Journal of Mathematical Physics, 55(4):040201, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goncalves:2014:NST

[2255] W. C. Gonçalves, E. Sardella, V. F. Becerra, M. V. Milosević, and F. M. Peeters. Numerical solution of the time dependent Ginzburg–Landau equations for mixed (d + s)-wave superconductors. Journal of Mathematical Physics, 55(4):041501, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2014:BBP

[2256] Yunfeng Jia, Jianhua Wu, and Hong-Kun Xu. Blow-up behavior of positive solutions for a chemical fuel ignition device model. Journal of Mathematical Physics, 55(4):041502, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:2014:FLA

[2257] R. M. Navarro. Filiform Lie algebras of order 3. Journal of Mathematical Physics, 55(4):041701, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Runkel:2014:BMC

[2258] Ingo Runkel. A braided monoidal category for free super-bosons. Journal of Mathematical Physics, 55(4):041702, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2014:KMB

[2259] Liang Chang. Kitaev models based on unitary quantum groupoids. Journal of Mathematical Physics, 55(4):041703, April 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2014:IDL

[2260] Lu Ding, Xiaoyu Jia, and Wei Zhang. On infinite-dimensional 3-Lie algebras. Journal of Mathematical Physics, 55(4):041704, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnlind:2014:CGM

[2261] Joakim Arnlind. Curvature and geometric modules of noncommutative spheres and tori. Journal of Mathematical Physics, 55(4):041705, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Macdougall:2014:STG

[2262] James Macdougall and Douglas Singleton. Stokes’ theorem, gauge symmetry and the time-dependent Aharonov–Bohm effect. Journal of Mathematical Physics, 55(4):042101, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ali:2014:HPQ

[2263] S. Twareque Ali and Miroslav Englis. Hermite polynomials and quasiclassical asymptotics. Journal of Mathematical Physics, 55(4):042102, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamuntavicius:2014:GIT

[2264] Gintautas P. Kamuntavicius. Galilei invariant technique for quantum system description. Journal of Mathematical Physics, 55(4):042103, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barseghyan:2014:RVS

[2265] Diana Barseghyan and Pavel Exner. A regular version of Smilansky model. Journal of Mathematical Physics, 55(4):042104, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delRio:2014:SRO

[2266] Rafael del Rio. Spectra of random operators with absolutely continuous integrated density of states. Journal of Mathematical Physics, 55(4): 042105, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braverman:2014:BPP

[2267] Maxim Braverman. The Berry phase and the phase of the determinant. Journal of Mathematical Physics, 55(4):042106, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2014:PTS

[2268] Francisco M. Fernández and Javier Garcia. Parity-time symmetry broken by point-group symmetry. Journal of Mathematical Physics, 55(4): 042107, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barron:2014:SCP

[2269] Tatyana Barron, Xiaonan Ma, George Marinescu, and Martin Pinsonnault. Semi-classical properties of Berezin–Toeplitz operators with C^{k}- symbol. Journal of Mathematical Physics, 55(4):042108, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ojeda-Guillen:2014:PNC

[2270] D. Ojeda-Guillén, R. D. Mota, and V. D. Granados. The SU(1, 1) Perelomov number coherent states and the non-degenerate parametric amplifier. Journal of Mathematical Physics, 55(4):042109, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2014:GUR

[2271] Ole Andersson and Hoshang Heydari. Geometric uncertainty relation for mixed quantum states. Journal of Mathematical Physics, 55(4):042110, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Busch:2014:MUR

[2272] Paul Busch, Pekka Lahti, and Reinhard F. Werner. Measurement uncertainty relations. Journal of Mathematical Physics, 55(4):042111, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carlen:2014:RTS

[2273] Eric A. Carlen and Elliott H. Lieb. Remainder terms for some quantum entropy inequalities. Journal of Mathematical Physics, 55(4):042201, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2014:PQF

[2274] Felix Finster. Perturbative quantum field theory in the framework of the fermionic projector. Journal of Mathematical Physics, 55(4):042301, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gegenberg:2014:IHS

[2275] Jack Gegenberg, Andrew C. Day, Haitao Liu, and Sanjeev S. Seahra. An instability of hyperbolic space under the Yang–Mills flow. Journal of Mathematical Physics, 55(4):042501, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantica:2014:PSS

[2276] Carlo Alberto Mantica and Young Jin Suh. Pseudo-Z symmetric space– times. Journal of Mathematical Physics, 55(4):042502, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2014:THT

[2277] Yi-Chiuan Chen, Shyan-Shiou Chen, and Juan-Ming Yuan. Topological horseshoes in travelling waves of discretized nonlinear wave equations. Journal of Mathematical Physics, 55(4):042701, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peng:2014:DSH

[2278] Keke Peng and Yiping Luo. Dynamics symmetries of Hamiltonian system on time scales. Journal of Mathematical Physics, 55(4):042702, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crenshaw:2014:EME

[2279] Michael E. Crenshaw. Electromagnetic momentum and the energymomentum tensor in a linear medium with magnetic and dielectric properties. Journal of Mathematical Physics, 55(4):042901, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertin:2014:TMY

[2280] M. C. Bertin, B. M. Pimentel, C. E. Valcárcel, and G. E. R. Zambrano. Topologically massive Yang–Mills: A Hamilton–Jacobi constraint analysis. Journal of Mathematical Physics, 55(4):042902, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laptev:2014:NPW

[2281] Denis V. Laptev and Mikhail M. Bogdan. Nonlinear periodic waves solutions of the nonlinear self-dual network equations. Journal of Mathematical Physics, 55(4):042903, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campoamor-Stursberg:2014:SSI

[2282] R. Campoamor-Stursberg. Superposition of super-integrable pseudo-Euclidean potentials in n = 2 with a fundamental constant of motion of arbitrary order in the momenta. Journal of Mathematical Physics, 55(4):042904, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoshida:2014:RHL

[2283] Z. Yoshida, Y. Kawazura, and T. Yokoyama. Relativistic helicity and link in Minkowski space–time. Journal of Mathematical Physics, 55(4): 043101, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Greynat:2014:DPR

[2284] David Greynat, Javier Sesma, and Grégory Vulvert. Derivatives of the Pochhammer and reciprocal Pochhammer symbols and their use in epsilon-expansions of Appell and Kampé de Fériet functions. Journal of Mathematical Physics, 55(4):043501, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dehghani:2014:GDS

[2285] A. Dehghani. General displaced SU (1, 1) number states: Revisited. Journal of Mathematical Physics, 55(4):043502, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babichenko:2014:BFF

[2286] A. Babichenko and V. Regelskis. On boundary fusion and functional relations in the Baxterized affine Hecke algebra. Journal of Mathematical Physics, 55(4):043503, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shcherbina:2014:CVM

[2287] M. Shcherbina. Change of variables as a method to study general βmodels: Bulk universality. Journal of Mathematical Physics, 55(4): 043504, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felice:2014:QNC

[2288] Domenico Felice, Stefano Mancini, and Marco Pettini. Quantifying networks complexity from information geometry viewpoint. Journal of Mathematical Physics, 55(4):043505, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:GTC

[2289] Wei Wang. On 3-gauge transformations, 3-curvatures, and Graycategories. Journal of Mathematical Physics, 55(4):043506, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2014:HGD

[2290] Lamei Yuan. Hom Gel’fand–Dorfman bialgebras and Hom–Lie conformal algebras. Journal of Mathematical Physics, 55(4):043507, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teo:2014:SPC

[2291] L. P. Teo. Sphere-plate Casimir interaction in (D+1)-dimensional spacetime. Journal of Mathematical Physics, 55(4):043508, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Rivasseau:2014:GCT

[2292] Vincent Rivasseau and Adrian Tanasa. Generalized constructive tree weights. Journal of Mathematical Physics, 55(4):043509, April 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gomez-Ullate:2014:EKA

[2293] David Gomez-Ullate, Yves Grandati, and Robert Milson. Extended Krein–Adler theorem for the translationally shape invariant potentials. Journal of Mathematical Physics, 55(4):043510, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tare:2014:EBS

[2294] Jeffrey D. Tare and Jose Perico H. Esguerra. Erratum: “Bound states for multiple Dirac-δ wells in space-fractional quantum mechanics” [J. Math. Phys. 55, 012106 (2014)]. Journal of Mathematical Physics, 55 (4):049901, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2163].

Haesen:2014:EIE

[2295] Stefan Haesen and Leopold Verstraelen. Erratum: “Ideally embedded space–times” [J. Math. Phys. 45, 1497 (2004)]. Journal of Mathematical Physics, 55(4):049902, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4].

Deng:2014:NSS

[2296] Yinbin Deng, Shuangjie Peng, and Jixiu Wang. Nodal soliton solutions for generalized quasilinear Schrödinger equations. Journal of Mathematical Physics, 55(5):051501, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smith:2014:ULF

[2297] Paul Smith. An unconstrained Lagrangian formulation and conservation laws for the Schrödinger map system. Journal of Mathematical Physics, 55(5):051502, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2014:CCP

[2298] Chiun-Chang Lee. The charge conserving Poisson–Boltzmann equations: Existence, uniqueness, and maximum principle. Journal of Mathematical Physics, 55(5):051503, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glatt-Holtz:2014:ERI

[2299] Nathan Glatt-Holtz, Igor Kukavica, Vlad Vicol, and Mohammed Ziane. Existence and regularity of invariant measures for the three dimensional stochastic primitive equations. Journal of Mathematical Physics, 55(5): 051504, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2014:RCU

[2300] Jishan Fan, Fucai Li, and Gen Nakamura. Regularity criteria and uniform estimates for the Boussinesq system with temperature-dependent viscosity and thermal diffusivity. Journal of Mathematical Physics, 55(5): 051505, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelPunta:2014:NHS

[2301] J. A. Del Punta, M. J. Ambrosio, G. Gasaneo, S. A. Zaytsev, and L. U. Ancarani. Non-homogeneous solutions of a Coulomb Schrödinger equation as basis set for scattering problems. Journal of Mathematical Physics, 55(5):052101, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agboola:2014:SGH

[2302] Davids Agboola. On the solvability of the generalized hyperbolic doublewell models. Journal of Mathematical Physics, 55(5):052102, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dhahri:2014:SAB

[2303] Ameur Dhahri. Self-adjointness and boundedness in quadratic quantization. Journal of Mathematical Physics, 55(5):052103, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jarret:2014:FGC

[2304] Michael Jarret and Stephen P. Jordan. The fundamental gap for a class of Schrödinger operators on path and hypercube graphs. Journal of Mathematical Physics, 55(5):052104, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [2362].

Finster:2014:PDF

[2305] Felix Finster and Jürgen Tolksdorf. Perturbative description of the fermionic projector: Normalization, causality, and Furry’s theorem.

Journal of Mathematical Physics, 55(5):052301, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fedoruk:2014:MDM

[2306] Sergey Fedoruk, Evgeny Ivanov, and Andrei Smilga. N = 4 mechanics with diverse (4, 4, 0) multiplets: Explicit examples of hyper–Kähler with torsion, Clifford Kähler with torsion, and octonionic Kähler with torsion geometries. Journal of Mathematical Physics, 55(5):052302, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celerier:2014:ECS

[2307] Marie-Noëlle Célérier and Laurent Nottale. Emergence of complex and spinor wave functions in scale relativity. II. Lorentz invariance and bispinors. Journal of Mathematical Physics, 55(5):052303, May 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dung:2014:SDD

[2308] Nguyen Tien Dung. Stochastic dynamics of a delayed bistable system with multiplicative noise. Journal of Mathematical Physics, 55(5): 052701, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuetche:2014:PPA

[2309] Victor Kamgang Kuetche, Saliou Youssoufa, and Timoleon Crepin Kofane. Phase portraits analysis of a barothropic system: The initial value problem. Journal of Mathematical Physics, 55(5):052702, May 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chi:2014:EBH

[2310] Dongpyo Chi, Sun-Ho Choi, and Seung-Yeal Ha. Emergent behaviors of a holonomic particle system on a sphere. Journal of Mathematical Physics, 55(5):052703, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaeta:2014:CTH

[2311] G. Gaeta and M. A. Rodríguez. Canonical transformations for hyperkähler structures and hyperhamiltonian dynamics. Journal of Mathematical Physics, 55(5):052901, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beau:2014:RIR

[2312] Mathieu Beau and Baptiste Savoie. Rigorous investigation of the reduced density matrix for the ideal Bose gas in harmonic traps by a loop-gas-like approach. Journal of Mathematical Physics, 55(5):053301, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bozkurt:2014:NAM

[2313] Zehra Bozkurt, Ismail Gök, Yusuf Yayli, and F. Nejat Ekmekci. A new approach for magnetic curves in 3D Riemannian manifolds. Journal of Mathematical Physics, 55(5):053501, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayashi:2014:NUS

[2314] Nakao Hayashi, Pavel I. Naumkin, and Tomoyuki Niizato. Nonexistence of the usual scattering states for the generalized Ostrovsky–Hunter equation. Journal of Mathematical Physics, 55(5):053502, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grammaticos:2014:SAD

[2315] B. Grammaticos, A. Ramani, K. M. Tamizhmani, T. Tamizhmani, and J. Satsuma. Strongly asymmetric discrete Painlevé equations: The additive case. Journal of Mathematical Physics, 55(5):053503, May 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jecko:2014:MTB

[2316] Thierry Jecko. On the mathematical treatment of the Born– Oppenheimer approximation. Journal of Mathematical Physics, 55(5): 053504, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2014:RTG

[2317] Nianhua Li and Xiaoxing Niu. A reciprocal transformation for the Geng– Xue equation. Journal of Mathematical Physics, 55(5):053505, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [298].

An:2014:DSE

[2318] Hongli An and Manwai Yuen. Drifting solutions with elliptic symmetry for the compressible Navier–Stokes equations with density-dependent viscosity. Journal of Mathematical Physics, 55(5):053506, May 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abedi-Fardad:2014:ISH

[2319] J. Abedi-Fardad, A. Rezaei-Aghdam, and Gh. Haghighatdoost. Integrable and superintegrable Hamiltonian systems with four dimensional real Lie algebras as symmetry of the systems. Journal of Mathematical Physics, 55(5):053507, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2014:NSH

[2320] Junchao Chen, Xiangpeng Xin, and Yong Chen. Nonlocal symmetries of the Hirota–Satsuma coupled Korteweg–de Vries system and their applications: Exact interaction solutions and integrable hierarchy. Journal of Mathematical Physics, 55(5):053508, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valero-Valdes:2014:PDA

[2321] Carlos Valero-Valdés and Rafael Herrera-Guzmán. Projecting diffusion along the normal bundle of a plane curve. Journal of Mathematical Physics, 55(5):053509, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordoa:2014:PSM

[2322] P. R. Gordoa, A. Pickering, and M. Senthilvelan. The prelle-Singer method and Painlevé hierarchies. Journal of Mathematical Physics, 55 (5):053510, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2014:SCQ

[2323] Xiu-Hong Sun and Yuan Li. Some characterizations of quantum channel in infinite Hilbert spaces. Journal of Mathematical Physics, 55(5): 053511, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhai:2014:EQP

[2324] Yunyun Zhai, Xianguo Geng, and Guoliang He. Explicit quasi-periodic solutions of the Vakhnenko equation. Journal of Mathematical Physics, 55(5):053512, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tiwari:2014:ECL

[2325] Ajey K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Erratum: “Classification of Lie point symmetries for quadratic Liénard type equation ẍ+f(x)ẋ^{2}+g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 55(5):059901, May 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1836] and comment [2707].

Schwartz:2014:ERQ

[2326] Charles Schwartz. Erratum: “Relativistic quaternionic wave equation” [J. Math. Phys. 47, 122301 (2006)]. Journal of Mathematical Physics, 55(5):059902, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [8].

Sahin-Biryol:2014:ASL

[2327] Derya Sahin-Biryol and Boaz Ilan. Asymptotic solution of light transport problems in optically thick luminescent media. Journal of Mathematical Physics, 55(6):061501, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2014:MCS

[2328] Claudianor O. Alves and Minbo Yang. Multiplicity and concentration of solutions for a quasilinear Choquard equation. Journal of Mathematical Physics, 55(6):061502, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Douglas:2014:EAR

[2329] Andrew Douglas, Joe Repka, and Wainwright Joseph. The Euclidean algebra in rank 2 classical Lie algebras. Journal of Mathematical Physics, 55(6):061701, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galindo:2014:BRU

[2330] César Galindo and Eric C. Rowell. Braid representations from unitary braided vector spaces. Journal of Mathematical Physics, 55(6):061702, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holanda:2014:FTS

[2331] N. L. Holanda and F. Toppan. Four types of (super)conformal mechanics: D-module reps and invariant actions. Journal of Mathematical Physics, 55(6):061703, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hosseiny:2014:PCE

[2332] Ali Hosseiny. Possible central extensions of non-relativistic conformal algebras in 1 + 1. Journal of Mathematical Physics, 55(6):061704, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolte:2014:MPQ

[2333] Jens Bolte and Joachim Kerner. Many-particle quantum graphs and Bose–Einstein condensation. Journal of Mathematical Physics, 55(6): 061901, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sasaki:2014:ESP

[2334] Ryu Sasaki. Exactly solvable potentials with finitely many discrete eigenvalues of arbitrary choice. Journal of Mathematical Physics, 55 (6):062101, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morandi:2014:HLM

[2335] O. Morandi. Homogenization limit for a multiband effective mass model in heterostructures. Journal of Mathematical Physics, 55(6):062102, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goyeneche:2014:QTM

[2336] D. Goyeneche and A. C. de la Torre. Quantum tomography meets dynamical systems and bifurcations theory. Journal of Mathematical Physics, 55(6):062103, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fine:2014:STA

[2337] Dana S. Fine and Stephen Sawin. Short-time asymptotics of a rigorous path integral for N = 1 supersymmetric quantum mechanics on a Riemannian manifold. Journal of Mathematical Physics, 55(6):062104, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daCosta:2014:GSL

[2338] Bruno G. da Costa and Ernesto P. Borges. Generalized space and linear momentum operators in quantum mechanics. Journal of Mathematical Physics, 55(6):062105, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boscain:2014:MTT

[2339] Ugo Boscain, Fredrik Grönberg, Ruixing Long, and Herschel Rabitz. Minimal time trajectories for two-level quantum systems with two bounded controls. Journal of Mathematical Physics, 55(6):062106, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [2451].

Krbek:2014:IQM

[2340] Michael Krbek, Tomás Tyc, and Jan Vlach. Inequalities for quantum marginal problems with continuous variables. Journal of Mathematical Physics, 55(6):062201, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cohen:2014:CUP

[2341] Scott M. Cohen. Conditions for uniqueness of product representations for separable quantum channels and separable quantum states. Journal of Mathematical Physics, 55(6):062202, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Datta:2014:PS

[2342] Nilanjana Datta, Tony Dorlas, Richard Jozsa, and Fabio Benatti. Properties of subentropy. Journal of Mathematical Physics, 55(6):062203, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oszmaniec:2014:GTC

[2343] Michal Oszmaniec, Piotr Suwara, and Adam Sawicki. Geometry and topology of CC and CQ states. Journal of Mathematical Physics, 55(6): 062204, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frob:2014:MSC

[2344] Markus B. Fröb and Atsushi Higuchi. Mode-sum construction of the two-point functions for the Stueckelberg vector fields in the Poincaré patch of de Sitter space. Journal of Mathematical Physics, 55(6):062301, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawaguchi:2014:DQA

[2345] Io Kawaguchi and Kentaroh Yoshida. A deformation of quantum affine algebra in squashed Wess–Zumino–Novikov–Witten models. Journal of Mathematical Physics, 55(6):062302, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Futakuchi:2014:CDT

[2346] Shinichiro Futakuchi and Kouta Usui. Construction of dynamics and time-ordered exponential for unbounded non-symmetric Hamiltonians. Journal of Mathematical Physics, 55(6):062303, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konitopoulos:2014:ECS

[2347] Spyros Konitopoulos and George Savvidy. Extension of Chern–Simons forms. Journal of Mathematical Physics, 55(6):062304, June 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pankrashkin:2014:ODD

[2348] Konstantin Pankrashkin and Serge Richard. One-dimensional Dirac operators with zero-range interactions: Spectral, scattering, and topological results. Journal of Mathematical Physics, 55(6):062305, June 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:2014:SOD

[2349] José Navarro. On second-order, divergence-free tensors. Journal of Mathematical Physics, 55(6):062501, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2014:HPA

[2350] Yu Pan, Hadis Amini, Zibo Miao, John Gough, Valery Ugrinovskii, and Matthew R. James. Heisenberg picture approach to the stability of quantum Markov systems. Journal of Mathematical Physics, 55(6):062701, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valls:2014:LAI

[2351] C. Valls. On the local analytic integrability for a generalized damped nonlinear oscillator equation. Journal of Mathematical Physics, 55(6): 062901, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figotin:2014:LFS

[2352] Alexander Figotin and Aaron Welters. Lagrangian framework for systems composed of high-loss and lossless components. Journal of Mathematical Physics, 55(6):062902, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alberici:2014:MFM

[2353] D. Alberici, P. Contucci, and E. Mingione. A mean-field monomer-dimer model with attractive interaction: Exact solution and rigorous results. Journal of Mathematical Physics, 55(6):063301, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Edelman:2014:LTR

[2354] Alan Edelman, Per-Olof Persson, and Brian D. Sutton. Low-temperature random matrix theory at the soft edge. Journal of Mathematical Physics, 55(6):063302, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ishida:2014:LOE

[2355] Hideshi Ishida. A law of order estimation and leading-order terms for a family of averaged quantities on a multibaker chain system. Journal of Mathematical Physics, 55(6):063303, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tribe:2014:GEL

[2356] Roger Tribe and Oleg Zaboronski. The Ginibre evolution in the large- n limit. Journal of Mathematical Physics, 55(6):063304, June 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ali:2014:QAG

[2357] S. Twareque Ali and K. Thirulogasanthar. The quaternionic affine group and related continuous wavelet transforms on complex and quaternionic Hilbert spaces. Journal of Mathematical Physics, 55(6):063501, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiselev:2014:STS

[2358] Oleg Kiselev and Nikolai Tarkhanov. Scattering of trajectories at a separatrix under autoresonance. Journal of Mathematical Physics, 55(6): 063502, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mamiya:2014:HKF

[2359] A. Mamiya and A. Pinzul. Heat kernel for flat generalized Laplacians with anisotropic scaling. Journal of Mathematical Physics, 55(6):063503, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:2014:PRI

[2360] Adam Doliwa and Sergey M. Sergeev. The pentagon relation and incidence geometry. Journal of Mathematical Physics, 55(6):063504, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xu:2014:HSC

[2361] Songhe Xu and Bo Xue. Hamiltonian structures and conservation laws of the Belov–Chaltikian lattice hierarchy. Journal of Mathematical Physics, 55(6):063505, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jarret:2014:PNF

[2362] Michael Jarret and Stephen P. Jordan. Publisher’s note: “The fundamental gap for a class of Schrödinger operators on path and hypercube graphs” [J. Math. Phys. 55, 052104 (2014)]. Journal of Mathematical Physics, 55(6):069901, June 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2304].

Jaksic:2014:I

[2363] Vojkan Jaksic, Claude-Alain Pillet, and Robert Seiringer. Introduction. Journal of Mathematical Physics, 55(7):075101, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Derezinski:2014:QFC

[2364] Jan Dereziński. Quantum fields with classical perturbations. Journal of Mathematical Physics, 55(7):075201, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saad:2014:GAL

[2365] R. Ben Sâad and C.-A. Pillet. A geometric approach to the Landauer– Büttiker formula. Journal of Mathematical Physics, 55(7):075202, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moller:2014:FCP

[2366] Jacob Schach Møller. Fully coupled Pauli–Fierz systems at zero and positive temperature. Journal of Mathematical Physics, 55(7):075203, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruneau:2014:RIO

[2367] Laurent Bruneau, Alain Joye, and Marco Merkli. Repeated interactions in open quantum systems. Journal of Mathematical Physics, 55(7): 075204, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Livi:2014:DSS

[2368] Roberto Livi. Dynamics, stability, and statistics on lattices and networks. Journal of Mathematical Physics, 55(7):075205, July 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeRoeck:2014:QDDa

[2369] Wojciech De Roeck, Jürg Fröhlich, and Kevin Schnelli. Quantum diffusion with drift and the Einstein relation. I. Journal of Mathematical Physics, 55(7):075206, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeRoeck:2014:QDDb

[2370] Wojciech De Roeck, Jürg Fröhlich, and Kevin Schnelli. Quantum diffusion with drift and the Einstein relation. II. Journal of Mathematical Physics, 55(7):075207, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaspard:2014:RPC

[2371] Pierre Gaspard. Random paths and current fluctuations in nonequilibrium statistical mechanics. Journal of Mathematical Physics, 55(7): 075208, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seiringer:2014:BGB

[2372] Robert Seiringer. Bose gases, Bose–Einstein condensation, and the Bogoliubov approximation. Journal of Mathematical Physics, 55(7):075209, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaksic:2014:NLP

[2373] Vojkan Jaksić and Claude-Alain Pillet. A note on the Landauer principle in quantum statistical mechanics. Journal of Mathematical Physics, 55 (7):075210, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spehner:2014:QCD

[2374] Dominique Spehner. Quantum correlations and distinguishability of quantum states. Journal of Mathematical Physics, 55(7):075211, July 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2014:GET

[2375] Jianwei Yang and Qiangchang Ju. Global existence of the threedimensional viscous quantum magnetohydrodynamic model. Journal of Mathematical Physics, 55(8):081501, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saccomandi:2014:MGS

[2376] Giuseppe Saccomandi and Raffaele Vitolo. On the mathematical and geometrical structure of the determining equations for shear waves in nonlinear isotropic incompressible elastodynamics. Journal of Mathematical Physics, 55(8):081502, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guesmia:2014:SWP

[2377] Aissa Guesmia. Some well-posedness and general stability results in Timoshenko systems with infinite memory and distributed time delay. Journal of Mathematical Physics, 55(8):081503, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2014:EDS

[2378] Xinglong Wu and Boling Guo. The exponential decay of solutions and traveling wave solutions for a modified Camassa–Holm equation with cubic nonlinearity. Journal of Mathematical Physics, 55(8):081504, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Panasenko:2014:FTS

[2379] Grigory Panasenko and Konstantin Pileckas. Flows in a tube structure: Equation on the graph. Journal of Mathematical Physics, 55(8):081505, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cordero:2014:GMO

[2380] Elena Cordero, Karlheinz Gröchenig, Fabio Nicola, and Luigi Rodino. Generalized metaplectic operators and the Schrödinger equation with a potential in the Sjöstrand class. Journal of Mathematical Physics, 55 (8):081506, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnaudon:2014:SEP

[2381] Marc Arnaudon, Xin Chen, and Ana Bela Cruzeiro. Stochastic Euler– Poincaré reduction. Journal of Mathematical Physics, 55(8):081507, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nayak:2014:AKM

[2382] Saudamini Nayak and K. C. Pati. Affine Kac–Moody symmetric spaces related with A_{1}^{(1)}, A_{2}^{(1)}, A_{2}^{(2)}. Journal of Mathematical Physics, 55(8): 081701, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:2014:GDT

[2383] Won Sang Chung, Mahouton Norbert Hounkonnou, and Sama Arjika. Generalized q-deformed Tamm–Dancoff oscillator algebra and associated coherent states. Journal of Mathematical Physics, 55(8):081702, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benayadi:2014:CQL

[2384] Saïd Benayadi and Alberto Elduque. Classification of quadratic Lie algebras of low dimension. Journal of Mathematical Physics, 55(8):081703, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crann:2014:UPU

[2385] Jason Crann and Mehrdad Kalantar. An uncertainty principle for unimodular quantum groups. Journal of Mathematical Physics, 55(8): 081704, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2014:EIP

[2386] Bintao Cao, Li Luo, and Ke Ou. Extensions of inhomogeneous polynomial representations for sl(m + 1|n). Journal of Mathematical Physics, 55(8):081705, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calixto:2014:ORC

[2387] M. Calixto and E. Pérez-Romero. On the oscillator realization of conformal U(2, 2) quantum particles and their particle-hole coherent states. Journal of Mathematical Physics, 55(8):081706, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baseilhac:2014:ALH

[2388] P. Baseilhac and T. T. Vu. Analogues of Lusztig’s higher order relations for the q-Onsager algebra. Journal of Mathematical Physics, 55 (8):081707, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:SCH

[2389] Shengxiang Wang and Shuangjian Guo. Symmetries and the u-condition in Hom–Yetter–Drinfeld categories. Journal of Mathematical Physics, 55 (8):081708, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-Arguello:2014:JMP

[2390] Angel M. Martínez-Argüello, Moisés Martínez-Mares, and Julio C. García. Joint moments of proper delay times. Journal of Mathematical Physics, 55(8):081901, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ishida:2014:ISP

[2391] Atsuhide Ishida. On inverse scattering problem for the Schrödinger equation with repulsive potentials. Journal of Mathematical Physics, 55(8): 082101, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2014:SGC

[2392] Richard L. Hall and Nasser Saad. Spectra generated by a confined softcore Coulomb potential. Journal of Mathematical Physics, 55(8):082102, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moya-Cessa:2014:SLO

[2393] Héctor M. Moya-Cessa, Francisco Soto-Eguibar, and Demetrios N. Christodoulides. A squeeze-like operator approach to position-dependent mass in quantum mechanics. Journal of Mathematical Physics, 55(8): 082103, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liverts:2014:AME

[2394] Evgeny Z. Liverts and Nir Barnea. Analytic matrix elements for the twoelectron atomic basis with logarithmic terms. Journal of Mathematical Physics, 55(8):082104, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dmitrasinovic:2014:AAT

[2395] V. Dmitrasinović and Igor Salom. SO(4) algebraic approach to the threebody bound state problem in two dimensions. Journal of Mathematical Physics, 55(8):082105, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yesiltas:2014:SAD

[2396] Özlem Yesiltas. Supersymmetric analysis of the Dirac–Weyl operator within PT symmetry. Journal of Mathematical Physics, 55(8):082106, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vourdas:2014:QPD

[2397] A. Vourdas. Quantum probabilities as Dempster–Shafer probabilities in the lattice of subspaces. Journal of Mathematical Physics, 55(8):082107, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2014:QQH

[2398] Manuel F. Rañada. A quantum quasi-harmonic nonlinear oscillator with an isotonic term. Journal of Mathematical Physics, 55(8):082108, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2014:EGU

[2399] Meng-Sen Ma and Ren Zhao. The effect of generalized uncertainty principle on square well, a case study. Journal of Mathematical Physics, 55 (8):082109, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wolters:2014:TMP

[2400] Sander Wolters. Topos models for physics and topos theory. Journal of Mathematical Physics, 55(8):082110, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2014:MCO

[2401] Isaac H. Kim. Modulus of convexity for operator convex functions. Journal of Mathematical Physics, 55(8):082201, August 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guth:2014:QEC

[2402] Larry Guth and Alexander Lubotzky. Quantum error correcting codes and 4-dimensional arithmetic hyperbolic manifolds. Journal of Mathematical Physics, 55(8):082202, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2014:CFT

[2403] Lin Chen, Dragomir Ž. – Doković, Markus Grassl, and Bei Zeng. Canonical form of three-fermion pure-states with six single particle states. Journal of Mathematical Physics, 55(8):082203, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zanardi:2014:LRQ

[2404] Paolo Zanardi. Local random quantum circuits: Ensemble completely positive maps and swap algebras. Journal of Mathematical Physics, 55 (8):082204, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benatti:2014:GQA

[2405] Fabio Benatti, Samad Khabbazi Oskouei, and Ahmad Shafiei Deh Abad. Gacs quantum algorithmic entropy in infinite dimensional Hilbert spaces. Journal of Mathematical Physics, 55(8):082205, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tomamichel:2014:RDQ

[2406] Marco Tomamichel, Mario Berta, and Masahito Hayashi. Relating different quantum generalizations of the conditional Rényi entropy. Journal of Mathematical Physics, 55(8):082206, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2014:UPB

[2407] Jianxin Chen, Lin Chen, and Bei Zeng. Unextendible product basis for fermionic systems. Journal of Mathematical Physics, 55(8):082207, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2014:RCR

[2408] H. Boche and G. Janßen. Resource cost results for one-way entanglement distillation and state merging of compound and arbitrarily varying quantum sources. Journal of Mathematical Physics, 55(8):082208, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benini:2014:ROL

[2409] Marco Benini, Claudio Dappiaggi, and Simone Murro. Radiative observables for linearized gravity on asymptotically flat spacetimes and their boundary induced states. Journal of Mathematical Physics, 55(8): 082301, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lopez:2014:FOE

[2410] M. Castrillón López, J. Muñoz Masqué, and E. Rosado María. Firstorder equivalent to Einstein–Hilbert Lagrangian. Journal of Mathematical Physics, 55(8):082501, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chudecki:2014:PCS

[2411] Adam Chudecki and Michal Dobrski. Proper conformal symmetries in self-dual Einstein spaces. Journal of Mathematical Physics, 55(8):082502, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Minguzzi:2014:CCH

[2412] E. Minguzzi. Completeness of Cauchy horizon generators. Journal of Mathematical Physics, 55(8):082503, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mei:2014:SDS

[2413] May Mei. Spectra of discrete Schrödinger operators with primitive invertible substitution potentials. Journal of Mathematical Physics, 55 (8):082701, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2014:ITP

[2414] Xuezhu Lu and Junxiang Xu. Invariant tori with prescribed frequency for nearly integrable Hamiltonian systems. Journal of Mathematical Physics, 55(8):082702, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruschi:2014:LCS

[2415] Mario Bruschi, Francesco Calogero, and François Leyvraz. A large class of solvable discrete-time many-body problems. Journal of Mathematical Physics, 55(8):082703, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babin:2014:NTE

[2416] Anatoli Babin and Alexander Figotin. Neoclassical theory of elementary charges with spin of 1/2. Journal of Mathematical Physics, 55(8):082901, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahihi:2014:BSH

[2417] Taliya Sahihi and Homayoon Eshraghi. Biot–Savart helicity versus physical helicity: A topological description of ideal flows. Journal of Mathematical Physics, 55(8):083101, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alexanian:2014:CNM

[2418] Moorad Alexanian. Classical non-Markovian Boltzmann equation. Journal of Mathematical Physics, 55(8):083301, August 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Witte:2014:MGE

[2419] N. S. Witte and P. J. Forrester. Moments of the Gaussian β ensembles and the large-N expansion of the densities. Journal of Mathematical Physics, 55(8):083302, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barletti:2014:HEE

[2420] Luigi Barletti. Hydrodynamic equations for electrons in graphene obtained from the maximum entropy principle. Journal of Mathematical Physics, 55(8):083303, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helmuth:2014:IMO

[2421] Tyler Helmuth. Ising model observables and non-backtracking walks. Journal of Mathematical Physics, 55(8):083304, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pirjol:2014:LRG

[2422] Dan Pirjol. Long-run growth rate in a random multiplicative model. Journal of Mathematical Physics, 55(8):083305, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suhov:2014:FDP

[2423] Y. Suhov and I. Stuhl. FK–DLR properties of a quantum multi-type Bose-gas with a repulsive interaction. Journal of Mathematical Physics, 55(8):083306, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drewitz:2014:CDS

[2424] Alexander Drewitz, Balázs Ráth, and Artëm Sapozhnikov. On chemical distances and shape theorems in percolation models with long-range correlations. Journal of Mathematical Physics, 55(8):083307, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazzolo:2014:CFR

[2425] Alain Mazzolo, Clélia de Mulatier, and Andrea Zoia. Cauchy’s formulas for random walks in bounded domains. Journal of Mathematical Physics, 55(8):083308, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:2014:GFT

[2426] Han-Yu Wei and Tie-Cheng Xia. The generalized fractional trace variational identity and fractional integrable couplings of Kaup–Newell hierarchy. Journal of Mathematical Physics, 55(8):083501, August 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agore:2014:FPH

[2427] A. L. Agore. Free Poisson Hopf algebras generated by coalgebras. Journal of Mathematical Physics, 55(8):083502, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aerts:2014:SHP

[2428] Diederik Aerts and Massimiliano Sassoli de Bianchi. Solving the hard problem of Bertrand’s paradox. Journal of Mathematical Physics, 55 (8):083503, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fu:2014:IPD

[2429] Wei Fu, R. Ilangovane, K. M. Tamizhmani, and Da jun Zhang. Integrability properties of the dispersionless Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 55(8):083504, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramm:2014:EWS

[2430] A. G. Ramm. Electromagnetic wave scattering by small perfectly conducting particles and applications. Journal of Mathematical Physics, 55 (8):083505, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szczygielski:2014:AFT

[2431] Krzysztof Szczygielski. On the application of Floquet theorem in development of time-dependent Lindbladians. Journal of Mathematical Physics, 55(8):083506, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2014:MPM

[2432] T. Skrypnyk. “Many-poled” r-matrix Lie algebras, Lax operators, and integrable systems. Journal of Mathematical Physics, 55(8):083507, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohn:2014:QSTa

[2433] Philipp A. Höhn. Quantization of systems with temporally varying discretization. I. Evolving Hilbert spaces. Journal of Mathematical Physics, 55(8):083508, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grandati:2014:EOP

[2434] Yves Grandati. Exceptional orthogonal polynomials and generalized Schur polynomials. Journal of Mathematical Physics, 55(8):083509, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tarasov:2014:AFM

[2435] Vasily E. Tarasov. Anisotropic fractal media by vector calculus in noninteger dimensional space. Journal of Mathematical Physics, 55(8): 083510, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yue:2014:AGS

[2436] Chao Yue and Tiecheng Xia. Algebro-geometric solutions for the complex Sharma–Tasso–Olver hierarchy. Journal of Mathematical Physics, 55(8): 083511, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Derezinski:2014:CHF

[2437] Jan Dereziński and Michal Wrochna. Continuous and holomorphic functions with values in closed operators. Journal of Mathematical Physics, 55(8):083512, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daletskii:2014:GSR

[2438] Alexei Daletskii, Yuri Kondratiev, Yuri Kozitsky, and Tanja Pasurek. Gibbs states on random configurations. Journal of Mathematical Physics, 55(8):083513, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2014:SRB

[2439] Zhengzheng Yang and Alexei F. Cheviakov. Some relations between symmetries of nonlocally related systems. Journal of Mathematical Physics, 55(8):083514, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia-Pelaez:2014:ILS

[2440] D. García-Peláez and C. S. López-Monsalvo. Infinitesimal Legendre symmetry in the geometrothermodynamics programme. Journal of Mathematical Physics, 55(8):083515, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2014:WKB

[2441] Yeongjoh Kim, Long Lee, and Gregory Lyng. The Wentzel–Kramers– Brillouin approximation of semiclassical eigenvalues of the Zakharov– Shabat problem. Journal of Mathematical Physics, 55(8):083516, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balogh:2014:FDK

[2442] F. Balogh, T. Fonseca, and J. Harnad. Finite dimensional Kadomtsev– Petviashvili τ-functions. I. Finite Grassmannians. Journal of Mathematical Physics, 55(8):083517, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mirahmadi:2014:EPR

[2443] Marjan-S. Mirahmadi and Amir H. Fatollahi. Eigenvalue problem for radial potentials in space with SU(2) fuzziness. Journal of Mathematical Physics, 55(8):083518, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saxena:2014:DOR

[2444] R. K. Saxena, A. M. Mathai, and H. J. Haubold. Distributed order reaction-diffusion systems associated with Caputo derivatives. Journal of Mathematical Physics, 55(8):083519, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krejcirik:2014:LES

[2445] David Krejcirík and Zhiqin Lu. Location of the essential spectrum in curved quantum layers. Journal of Mathematical Physics, 55(8):083520, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajaratnam:2014:CHJ

[2446] Krishan Rajaratnam and Raymond G. McLenaghan. Classification of Hamilton–Jacobi separation in orthogonal coordinates with diagonal curvature. Journal of Mathematical Physics, 55(8):083521, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Erman:2014:NGS

[2447] Fatih Erman, Berkin Malkoç, and O. Teoman Turgut. Nondegeneracy of the ground state for nonrelativistic Lee model. Journal of Mathematical Physics, 55(8):083522, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Combes:2014:LWL

[2448] Jean-Michel Combes, François Germinet, and Abel Klein. Local Wegner and Lifshitz tails estimates for the density of states for continuous random Schrödinger operators. Journal of Mathematical Physics, 55(8): 083523, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caudrelier:2014:ESM

[2449] Vincent Caudrelier, Mihail Mintchev, and Eric Ragoucy. Exact scattering matrix of graphs in magnetic field and quantum noise. Journal of Mathematical Physics, 55(8):083524, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawakubo:2014:EKE

[2450] Satoshi Kawakubo. Extendability of Kirchhoff elastic rods in complete Riemannian manifolds. Journal of Mathematical Physics, 55(8):083525, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boscain:2014:PNM

[2451] Ugo Boscain, Fredrik Grönberg, Ruixing Long, and Herschel Rabitz. Publisher’s note: “Minimal time trajectories for two-level quantum systems with two bounded controls” [J. Math. Phys. 55, 062106 (2014)]. Journal of Mathematical Physics, 55(8):089901, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2339].

Gurses:2014:TWS

[2452] Metin Gürses and Asli Pekcan. Traveling wave solutions of degenerate coupled Korteweg–de Vries equation. Journal of Mathematical Physics, 55(9):091501, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:DRS

[2453] Weike Wang and Xin Xu. The decay rate of solution for the bipolar Navier–Stokes–Poisson system. Journal of Mathematical Physics, 55(9): 091502, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Himonas:2014:PPU

[2454] A. Alexandrou Himonas and Ryan C. Thompson. Persistence properties and unique continuation for a generalized Camassa–Holm equation. Journal of Mathematical Physics, 55(9):091503, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:NGO

[2455] Ya-Guang Wang and Fang Yu. Nonlinear geometric optics for contact discontinuities in three dimensional compressible isentropic steady flows.

Journal of Mathematical Physics, 55(9):091504, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geissert:2014:SRF

[2456] Matthias Geissert, Matthias Hieber, and Nguyen Thieu Huy. Stability results for fluids of Oldroyd-B type on exterior domains. Journal of Mathematical Physics, 55(9):091505, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2014:SHC

[2457] Xiufu Zhang, Yan an Cai, and Yongjie Wang. Simple Harish-Chandra modules over super Schrödinger algebra in (1+1) dimensional spacetime. Journal of Mathematical Physics, 55(9):091701, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2014:EPT

[2458] Weideng Cui. Extremal projectors of two-parameter Kashiwara algebras. Journal of Mathematical Physics, 55(9):091702, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xin:2014:SSF

[2459] Qiaoling Xin and Lining Jiang. Symmetric structure of field algebra of Gspin models determined by a normal subgroup. Journal of Mathematical Physics, 55(9):091703, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsumoto:2014:RCE

[2460] Takuya Matsumoto and Alexander Molev. Representations of centrally extended Lie superalgebra psl(2|2). Journal of Mathematical Physics, 55 (9):091704, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guzzo:2014:CRL

[2461] H. Guzzo, Jr., I. Hernández, and O. A. Sánchez-Valenzuela. Classification of real Lie superalgebras based on a simple Lie algebra, giving rise to interesting examples involving su(2, 2). Journal of Mathematical Physics, 55(9):091705, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bapst:2014:LCL

[2462] Victor Bapst. The large connectivity limit of the Anderson model on tree graphs. Journal of Mathematical Physics, 55(9):092101, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christiansen:2014:EEP

[2463] H. R. Christiansen and M. S. Cunha. Energy eigenfunctions for positiondependent mass particles in a new class of molecular Hamiltonians. Journal of Mathematical Physics, 55(9):092102, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zaba:2014:SFS

[2464] Mariusz Zaba and Piotr Garbaczewski. Solving fractional Schrödingertype spectral problems: Cauchy oscillator and Cauchy well. Journal of Mathematical Physics, 55(9):092103, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantile:2014:QER

[2465] Andrea Mantile. Quantum evolution in the regime of quantum wells in a semiclassical island with artificial interface conditions. Journal of Mathematical Physics, 55(9):092104, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sandev:2014:TDS

[2466] Trifce Sandev, Irina Petreska, and Ervin K. Lenzi. Time-dependent Schrödinger-like equation with nonlocal term. Journal of Mathematical Physics, 55(9):092105, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:2014:QDP

[2467] Gustavo de Oliveira. Quantum dynamics of a particle constrained to lie on a surface. Journal of Mathematical Physics, 55(9):092106, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2014:BCQ

[2468] Isaac Kim and Mary Beth Ruskai. Bounds on the concavity of quantum entropy. Journal of Mathematical Physics, 55(9):092201, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buscemi:2014:GTC

[2469] Francesco Buscemi, Nilanjana Datta, and Sergii Strelchuk. Gametheoretic characterization of antidegradable channels. Journal of Mathematical Physics, 55(9):092202, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maeda:2014:GTN

[2470] Yoshiaki Maeda, Akifumi Sako, Toshiya Suzuki, and Hiroshi Umetsu. Gauge theories in noncommutative homogeneous Kähler manifolds. Journal of Mathematical Physics, 55(9):092301, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:CNC

[2471] Zhituo Wang. Construction of the noncommutative complex ball. Journal of Mathematical Physics, 55(9):092302, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Linnaeus:2014:PIM

[2472] Staffan Linnæus. Phase-integral method for the radial Dirac equation. Journal of Mathematical Physics, 55(9):092303, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liard:2014:MFL

[2473] Quentin Liard and Boris Pawilowski. Mean field limit for bosons with compact kernels interactions by Wigner measures transportation. Journal of Mathematical Physics, 55(9):092304, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davydov:2014:BMD

[2474] Alexei Davydov. Bogomolov multiplier, double class-preserving automorphisms, and modular invariants for orbifolds. Journal of Mathematical Physics, 55(9):092305, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chu:2014:LSM

[2475] Yi-Zen Chu. A line source in Minkowski for the de Sitter spacetime scalar Green’s function: Massless minimally coupled case. Journal of Mathematical Physics, 55(9):092501, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Olive:2014:IIC

[2476] M. Olive and N. Auffray. Isotropic invariants of a completely symmetric third-order tensor. Journal of Mathematical Physics, 55(9):092901, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhai:2014:WBP

[2477] Panpan Zhai, Zhengguang Guo, and Weiming Wang. Wave breaking phenomenon for a modified two-component Dullin–Gottwald–Holm equation. Journal of Mathematical Physics, 55(9):093101, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rabinowitch:2014:SCN

[2478] Alexander S. Rabinowitch. On some classes of nonstationary axially symmetric solutions to the Navier–Stokes equations. Journal of Mathematical Physics, 55(9):093102, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zafar:2014:CDC

[2479] M. Zafar and V. D. Sharma. Characteristic decomposition of compressible Euler equations for a non-ideal gas in two-dimensions. Journal of Mathematical Physics, 55(9):093103, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [2532].

Whitehead:2014:RBV

[2480] Jared P. Whitehead and Ralf W. Wittenberg. A rigorous bound on the vertical transport of heat in Rayleigh–Bénard convection at infinite Prandtl number with mixed thermal boundary conditions. Journal of Mathematical Physics, 55(9):093104, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peletier:2014:LDS

[2481] Mark A. Peletier, Frank Redig, and Kiamars Vafayi. Large deviations in stochastic heat-conduction processes provide a gradient-flow structure for heat conduction. Journal of Mathematical Physics, 55(9):093301, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takahashi:2014:OMM

[2482] Yuta Takahashi and Makoto Katori. Oscillatory matrix model in Chern– Simons theory and Jacobi-theta determinantal point process. Journal of Mathematical Physics, 55(9):093302, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lees:2014:LRO

[2483] Benjamin Lees. Long-range order for the spin-1 Heisenberg model with a small antiferromagnetic interaction. Journal of Mathematical Physics, 55 (9):093303, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakashima:2014:RBF

[2484] Makoto Nakashima. A remark on the bound for the free energy of directed polymers in random environment in 1 + 2 dimension. Journal of Mathematical Physics, 55(9):093304, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penson:2014:LTF

[2485] K. A. Penson and K. Górska. On the Laplace transform of the Fréchet distribution. Journal of Mathematical Physics, 55(9):093501, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:2014:NIC

[2486] Palle Jorgensen and Feng Tian. Von Neumann indices and classes of positive definite functions. Journal of Mathematical Physics, 55(9):093502, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Groeger:2014:KVF

[2487] Josua Groeger. Killing vector fields and harmonic superfield theories. Journal of Mathematical Physics, 55(9):093503, September 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Montani:2014:DAC

[2488] H. Montani and M. Zuccalli. Dirac approach to constrained submanifolds in a double loop group: From Wess–Zumino–Novikov–Witten to Poisson–Lie σ-model. Journal of Mathematical Physics, 55(9):093504, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gade:2014:IRV

[2489] R. M. Gade. Intertwiners of U′_{q}(ŝl(2))–representations and the vectorvalued big q Jacobi transform. Journal of Mathematical Physics, 55(9):

093505, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaillard:2014:TPW

[2490] Pierre Gaillard. Two parameters Wronskian representation of solutions of nonlinear Schrödinger equation, eighth Peregrine breather and multirogue waves. Journal of Mathematical Physics, 55(9):093506, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Munger:2014:HCS

[2491] Paul E. Munger and Darren C. Ong. The Hölder continuity of spectral measures of an extended CMV matrix. Journal of Mathematical Physics, 55(9):093507, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rezaei-Aghdam:2014:PWZ

[2492] A. Rezaei-Aghdam and M. Sephid. Perturbed Wess–Zumino–Witten models and N = (2, 2) supersymmetric sigma models on Lie groups with complex structure. Journal of Mathematical Physics, 55(9):093508, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isojima:2014:ESP

[2493] Shin Isojima. On exact solutions with periodic structure of the ultradiscrete Toda equation with parity variables. Journal of Mathematical Physics, 55(9):093509, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2014:TSH

[2494] Baiying He, Liangyun Chen, and Lili Ma. Two soliton hierarchies associated with SO(4) and the applications of SU(2)⊗SU(2) ≈ SO(4). Journal of Mathematical Physics, 55(9):093510, September 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Honnouvo:2014:DOA

[2495] G. Honnouvo and K. Thirulogasanthar. On the dimensions of oscillator algebras induced by orthogonal polynomials. Journal of Mathematical Physics, 55(9):093511, September 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weich:2014:ESA

[2496] Tobias Weich. Equivariant spectral asymptotics for h-pseudodifferential operators. Journal of Mathematical Physics, 55(10):101501, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falconi:2014:GSE

[2497] Marco Falconi. Global solution of the electromagnetic field-particle system of equations. Journal of Mathematical Physics, 55(10):101502, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melnikov:2014:SKV

[2498] Andrey Melnikov. Solution of the Korteweg–de Vries equation on the line with analytic initial potential. Journal of Mathematical Physics, 55 (10):101503, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schechter:2014:GSS

[2499] Martin Schechter. Ground state solutions for non-autonomous dynamical systems. Journal of Mathematical Physics, 55(10):101504, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Demontis:2014:IST

[2500] F. Demontis, B. Prinari, C. van der Mee, and F. Vitale. The inverse scattering transform for the focusing nonlinear Schrödinger equation with asymmetric boundary conditions. Journal of Mathematical Physics, 55 (10):101505, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loginov:2014:YML

[2501] E. K. Loginov and E. D. Loginova. Yang–Mills like instantons in eight and seven dimensions. Journal of Mathematical Physics, 55(10):101701, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Groenevelt:2014:CCT

[2502] Wolter Groenevelt. Coupling coefficients for tensor product representations of quantum SU(2). Journal of Mathematical Physics, 55(10): 101702, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eelbode:2014:CCI

[2503] D. Eelbode and T. Raeymaekers. Construction of conformally invariant higher spin operators using transvector algebras. Journal of Mathematical Physics, 55(10):101703, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirwin:2014:CST

[2504] William D. Kirwin, José M. Mourão, and João P. Nunes. Coherent state transforms and the Mackey–Stone–von Neumann theorem. Journal of Mathematical Physics, 55(10):102101, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bracken:2014:QPT

[2505] Paul Bracken. Quantization of a particle on a two-dimensional manifold of constant curvature. Journal of Mathematical Physics, 55(10):102102, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakayama:2014:TQT

[2506] Kunji Nakayama. Topos quantum theory on quantization-induced sheaves. Journal of Mathematical Physics, 55(10):102103, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dossa:2014:FST

[2507] Anselme F. Dossa and Gabriel Y. H. Avossevou. Full spectrum of the two-photon and the two-mode quantum Rabi models. Journal of Mathematical Physics, 55(10):102104, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durmus:2014:ASG

[2508] Ays.en Durmus. and Aysel Özfidan. Analytical solutions of a generalized non-central potential in N-dimensions. Journal of Mathematical Physics, 55(10):102105, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Audenaert:2014:UBE

[2509] Koenraad M. R. Audenaert and Milán Mosonyi. Upper bounds on the error probabilities and asymptotic error exponents in quantum multiple state discrimination. Journal of Mathematical Physics, 55(10):102201, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2014:TLS

[2510] Luise Adams, Christian Bogner, and Stefan Weinzierl. The two-loop sunrise graph in two space–time dimensions with arbitrary masses in terms of elliptic dilogarithms. Journal of Mathematical Physics, 55(10): 102301, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iacono:2014:LIG

[2511] Roberto Iacono and Jaume Llibre. Liouvillian integrability of gravitating static isothermal fluid spheres. Journal of Mathematical Physics, 55(10): 102501, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krisch:2014:RBC

[2512] J. P. Krisch and E. N. Glass. Relativistic Bessel cylinders. Journal of Mathematical Physics, 55(10):102502, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paetz:2014:CID

[2513] Tim-Torben Paetz. Characteristic initial data and smoothness of Scri. II. Asymptotic expansions and construction of conformally smooth data sets. Journal of Mathematical Physics, 55(10):102503, October 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hovhannisyan:2014:ALH

[2514] Gro Hovhannisyan. Ablowitz–Ladik hierarchy of integrable equations on a time-space scale. Journal of Mathematical Physics, 55(10):102701, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Montaldi:2014:PVH

[2515] James Montaldi and Citlalitl Nava-Gaxiola. Point vortices on the hyperbolic plane. Journal of Mathematical Physics, 55(10):102702, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yue:2014:STT

[2516] Haitian Yue, Mario I. Molina, Panayotis G. Kevrekidis, and Nikos I. Karachalios. Self-trapping transition for a nonlinear impurity within a linear chain. Journal of Mathematical Physics, 55(10):102703, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feher:2014:DBT

[2517] L. Fehér and T. F. Görbe. Duality between the trigonometric BC_{n} Sutherland system and a completed rational Ruijsenaars–Schneider–van Diejen system. Journal of Mathematical Physics, 55(10):102704, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cui:2014:NRC

[2518] Xiaojun Cui and Liang Jin. The negative of regular cosmological time function is a viscosity solution. Journal of Mathematical Physics, 55 (10):102705, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masterov:2014:DRC

[2519] Ivan Masterov. Dynamical realizations of N = 1 l-conformal Galilei superalgebra. Journal of Mathematical Physics, 55(10):102901, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:2014:SDP

[2520] Takao Suzuki. Six-dimensional Painlevé systems and their particular solutions in terms of rigid systems. Journal of Mathematical Physics, 55 (10):102902, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablamowicz:2014:BCS

[2521] Rafal Ablamowicz, Icaro Gonçalves, and Roldão da Rocha. Bilinear covariants and spinor fields duality in quantum Clifford algebras. Journal of Mathematical Physics, 55(10):103501, October 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agaltsov:2014:RHP

[2522] A. D. Agaltsov and R. G. Novikov. Riemann–Hilbert problem approach for two-dimensional flow inverse scattering. Journal of Mathematical Physics, 55(10):103502, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levi:2014:LGA

[2523] Decio Levi, Eugenio Ricca, Zora Thomova, and Pavel Winternitz. Lie group analysis of a generalized Krichever–Novikov differential-difference equation. Journal of Mathematical Physics, 55(10):103503, October 2014.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kudryashov:2014:ANS

[2524] Nikolai A. Kudryashov and Dmitry I. Sinelshchikov. Analytical and numerical studying of the perturbed Korteweg–de Vries equation. Journal of Mathematical Physics, 55(10):103504, October 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borzov:2014:ATD

[2525] V. V. Borzov and E. V. Damaskinsky. The algebra of two dimensional generalized Chebyshev–Koornwinder oscillator. Journal of Mathematical Physics, 55(10):103505, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contreras-Astorga:2014:CSA

[2526] Alonso Contreras-Astorga and Axel Schulze-Halberg. The confluent supersymmetry algorithm for Dirac equations with pseudoscalar potentials. Journal of Mathematical Physics, 55(10):103506, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohn:2014:QSTb

[2527] Philipp A. Höhn. Quantization of systems with temporally varying discretization. II. local evolution moves. Journal of Mathematical Physics, 55(10):103507, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boeijink:2014:GNT

[2528] Jord Boeijink and Koen van den Dungen. On globally non-trivial almost-commutative manifolds. Journal of Mathematical Physics, 55 (10):103508, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guillaumin-Espana:2014:CQD

[2529] Elisa Guillaumín-España, H. N. Núñez-Yépez, and A. L. Salas-Brito. Classical and quantum dynamics in an inverse square potential. Journal of Mathematical Physics, 55(10):103509, October 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zarmi:2014:NQM

[2530] Yair Zarmi. Nonlinear quantum-mechanical system associated with Sine– Gordon equation in (1 + 2) dimensions. Journal of Mathematical Physics, 55(10):103510, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon-Cortes:2014:ENO

[2531] L. F. Chacón-Cortes and W. A. Zúñiga-Galindo. Erratum: “Nonlocal operators, parabolic-type equations, and ultrametric random walks” [J. Math. Phys. 54, 113503 (2013)]. Journal of Mathematical Physics, 55 (10):109901, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zafar:2014:PNC

[2532] M. Zafar and V. D. Sharma. Publisher’s note: “Characteristic decomposition of compressible Euler equations for a non-ideal gas in twodimensions” [J. Math. Phys. 55, 093103 (2014)]. Journal of Mathematical Physics, 55(10):109902, October 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2479].

Wu:2014:MCP

[2533] Dan Wu. Mass concentration phenomenon for inhomogeneous fractional Hartree equations. Journal of Mathematical Physics, 55(11):111501, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

You:2014:CNB

[2534] Miman You and Shuanhong Wang. Constructing new braided T - categories over monoidal Hom–Hopf algebras. Journal of Mathematical Physics, 55(11):111701, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2014:NTB

[2535] Xiaohui Zhang and Shuanhong Wang. New Turaev braided group categories and weak (co)quasi-Turaev group coalgebras. Journal of Mathematical Physics, 55(11):111702, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dherin:2014:DMM

[2536] Benoit Dherin and Igor Mencattini. Deformations of momentum maps and G-systems. Journal of Mathematical Physics, 55(11):111703, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anguelova:2014:BFC

[2537] Iana I. Anguelova. Boson-fermion correspondence of type D–A and multilocal Virasoro representations on the Fock space F^{\frac{⊗1}{2}} . Journal of Mathematical Physics, 55(11):111704, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bucher:2014:IPC

[2538] David Bücher and Ingo Runkel. Integrable perturbations of conformal field theories and Yetter–Drinfeld modules. Journal of Mathematical Physics, 55(11):111705, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belletete:2014:PIM

[2539] Jonathan Belletête and Yvan Saint-Aubin. The principal indecomposable modules of the dilute Temperley–Lieb algebra. Journal of Mathematical Physics, 55(11):111706, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakamura:2014:MWO

[2540] Shu Nakamura. Modified wave operators for discrete Schrödinger operators with long-range perturbations. Journal of Mathematical Physics, 55(11):112101, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ganguly:2014:GKV

[2541] A. Ganguly and A. Das. Generalized Korteweg–de Vries equation induced from position-dependent effective mass quantum models and massdeformed soliton solution through inverse scattering transform. Journal of Mathematical Physics, 55(11):112102, November 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marquette:2014:CSA

[2542] Ian Marquette and Christiane Quesne. Combined state-adding and statedeleting approaches to type III multi-step rationally extended potentials: Applications to ladder operators and superintegrability. Journal of Mathematical Physics, 55(11):112103, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Briet:2014:STS

[2543] Ph. Briet, J. Dittrich, and E. Soccorsi. Scattering through a straight quantum waveguide with combined boundary conditions. Journal of Mathematical Physics, 55(11):112104, November 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tusek:2014:ACV

[2544] Matej Tusek. Atoms confined by very thin layers. Journal of Mathematical Physics, 55(11):112105, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agredo:2014:DFS

[2545] Julián Agredo, Franco Fagnola, and Rolando Rebolledo. Decoherence free subspaces of a quantum Markov semigroup. Journal of Mathematical Physics, 55(11):112201, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Audenaert:2014:QSD

[2546] Koenraad M. R. Audenaert. Quantum skew divergence. Journal of Mathematical Physics, 55(11):112202, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jivulescu:2014:RCR

[2547] Maria Anastasia Jivulescu, Nicolae Lupa, and Ion Nechita. On the reduction criterion for random quantum states. Journal of Mathematical Physics, 55(11):112203, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chitambar:2014:ASD

[2548] Eric Chitambar and Min-Hsiu Hsieh. Asymptotic state discrimination and a strict hierarchy in distinguishability norms. Journal of Mathematical Physics, 55(11):112204, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablinger:2014:IBS

[2549] J. Ablinger, J. Blümlein, C. G. Raab, and C. Schneider. Iterated binomial sums and their associated iterated integrals. Journal of Mathematical Physics, 55(11):112301, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knuth:2014:PFE

[2550] Kevin H. Knuth and Newshaw Bahreyni. A potential foundation for emergent space–time. Journal of Mathematical Physics, 55(11):112501, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boonserm:2014:GFM

[2551] Petarpa Boonserm, Auttakit Chatrabhuti, Tritos Ngampitipan, and Matt Visser. Greybody factors for Myers–Perry black holes. Journal of Mathematical Physics, 55(11):112502, November 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klinkhamer:2014:CSD

[2552] F. R. Klinkhamer and F. Sorba. Comparison of spacetime defects which are homeomorphic but not diffeomorphic. Journal of Mathematical Physics, 55(11):112503, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diacu:2014:ALR

[2553] Florin Diacu and Sergiu Popa. All the Lagrangian relative equilibria of the curved 3-body problem have equal masses. Journal of Mathematical Physics, 55(11):112701, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulesza:2014:DCL

[2554] Maité Kulesza and Jaume Llibre. The 3-dimensional cored and logarithm potentials: Periodic orbits. Journal of Mathematical Physics, 55(11): 112702, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gine:2014:SAP

[2555] Jaume Giné. Singularity analysis in planar vector fields. Journal of Mathematical Physics, 55(11):112703, November 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2014:GIS

[2556] Oscar E. Fernandez, Anthony M. Bloch, and Dmitry V. Zenkov. The geometry and integrability of the Suslov problem. Journal of Mathematical Physics, 55(11):112704, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertin:2014:ICS

[2557] M. C. Bertin, B. M. Pimentel, and C. E. Valcárcel. Involutive constrained systems and Hamilton–Jacobi formalism. Journal of Mathematical Physics, 55(11):112901, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fialkovsky:2014:AEL

[2558] I. V. Fialkovsky, M. V. Perel, and A. B. Plachenov. On astigmatic exponentially localized solutions for the wave and the Klein–Gordon–Fock equations. Journal of Mathematical Physics, 55(11):112902, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ho:2014:ECS

[2559] Choon-Lin Ho and Ryu Sasaki. Extensions of a class of similarity solutions of Fokker–Planck equation with time-dependent coefficients and fixed/moving boundaries. Journal of Mathematical Physics, 55(11): 113301, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takemura:2014:MIJ

[2560] Kouichi Takemura. Multi-indexed Jacobi polynomials and Maya diagrams. Journal of Mathematical Physics, 55(11):113501, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:2014:AST

[2561] Slawomir Klimek, Matt McBride, and Sumedha Rathnayake. A p-adic spectral triple. Journal of Mathematical Physics, 55(11):113502, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huh:2014:MLL

[2562] Youngsik Huh, Kyungpyo Hong, Hyoungjun Kim, Sungjong No, and Seungsang Oh. Minimum lattice length and ropelength of 2-bridge knots and links. Journal of Mathematical Physics, 55(11):113503, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrogowska:2014:LBS

[2563] Alina Dobrogowska and Tomasz Goliński. Lie bundle on the space of deformed skew-symmetric matrices. Journal of Mathematical Physics, 55(11):113504, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ikeda:2014:CFD

[2564] Noriaki Ikeda and Xiaomeng Xu. Canonical functions, differential graded symplectic pairs in supergeometry, and Alexandrov–Kontsevich– Schwartz–Zaboronsky sigma models with boundaries. Journal of Mathematical Physics, 55(11):113505, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hohn:2014:CCD

[2565] Philipp A. Höhn. Classification of constraints and degrees of freedom for quadratic discrete actions. Journal of Mathematical Physics, 55(11): 113506, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mikhailov:2014:DTD

[2566] Alexander V. Mikhailov, Georgios Papamikos, and Jing Ping Wang. Darboux transformation with dihedral reduction group. Journal of Mathematical Physics, 55(11):113507, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Czyzycki:2014:GDO

[2567] Tomasz Czyzycki and Jirí Hrivnák. Generalized discrete orbit function transforms of affine Weyl groups. Journal of Mathematical Physics, 55(11):113508, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hakova:2014:IFR

[2568] Lenka Háková and Agnieszka Tereszkiewicz. On immanant functions related to Weyl groups of A_{n}. Journal of Mathematical Physics, 55(11): 113509, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arsie:2014:PNL

[2569] Alessandro Arsie and Paolo Lorenzoni. Purely non-local Hamiltonian formalism, Kohno connections and ∨-systems. Journal of Mathematical Physics, 55(11):113510, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buniy:2014:QHO

[2570] R. V. Buniy, F. Colombo, I. Sabadini, and D. C. Struppa. Quantum harmonic oscillator with superoscillating initial datum. Journal of Mathematical Physics, 55(11):113511, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coffey:2014:ARS

[2571] Mark W. Coffey. Analytic results for spatial coherence and information entropy of an optical vortex field. Journal of Mathematical Physics, 55(11):113512, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amir:2014:CCS

[2572] Naila Amir and Shahid Iqbal. Comment on “Coherent states for the nonlinear harmonic oscillator” [J. Math. Phys. 53, 062104 (2012)]”. Journal of Mathematical Physics, 55(11):114101, November 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1282].

Li:2014:EPS

[2573] Gong-Bao Li and Hong-Yu Ye. The existence of positive solutions with prescribed L^{2}-norm for nonlinear Choquard equations. Journal of Mathematical Physics, 55(12):121501, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrabustovskyi:2014:OCS

[2574] Andrii Khrabustovskyi. Opening up and control of spectral gaps of the Laplacian in periodic domains. Journal of Mathematical Physics, 55(12): 121502, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fang:2014:BSN

[2575] Daoyuan Fang, Zheng Han, and W.-M. Wang. Bounded Sobolev norms for Klein–Gordon equations under non-resonant perturbation. Journal of Mathematical Physics, 55(12):121503, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2014:DRE

[2576] Yuhu Wu and Xiaoping Xue. Decay rate estimates for a class of quasilinear hyperbolic equations with damping terms involving p-Laplacian. Journal of Mathematical Physics, 55(12):121504, December 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zenchuk:2014:IMV

[2577] A. I. Zenchuk. On integration of a multidimensional version of n-wave type equation. Journal of Mathematical Physics, 55(12):121505, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skalak:2014:NRS

[2578] Zdenek Skalák. A note on the regularity of the solutions to the Navier– Stokes equations via the gradient of one velocity component. Journal of Mathematical Physics, 55(12):121506, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2014:NCE

[2579] Lu Ding, Zhaowen Yan, Huiping Zhang, and Wei Zhang. Nontrivial central extensions of 3-Lie algebras. Journal of Mathematical Physics, 55(12):121701, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashton:2014:QBL

[2580] Thomas Ashton and Andrey Mudrov. Quantization of borderline Levi conjugacy classes of orthogonal groups. Journal of Mathematical Physics, 55(12):121702, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kotani:2014:QIC

[2581] Motoko Kotani, Hermann Schulz-Baldes, and Carlos Villegas-Blas. Quantization of interface currents. Journal of Mathematical Physics, 55(12):121901, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2014:CIB

[2582] Pavel Exner and Alexander Minakov. Curvature-induced bound states in Robin waveguides and their asymptotical properties. Journal of Mathematical Physics, 55(12):122101, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molitor:2014:GDJ

[2583] Mathieu Molitor. Gaussian distributions, Jacobi group, and Siegel– Jacobi space. Journal of Mathematical Physics, 55(12):122102, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2014:CBG

[2584] Hong-Biao Zhang, Guang-Yuan Jiang, and San-Xing Guo. Construction of the Barut–Girardello type of coherent states for Pöschl–Teller potential. Journal of Mathematical Physics, 55(12):122103, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sacchetti:2014:ESW

[2585] Andrea Sacchetti. Existence of the Stark–Wannier quantum resonances. Journal of Mathematical Physics, 55(12):122104, December 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sen:2014:SBC

[2586] Surajit Sen and Helal Ahmed. SU(4) based classification of four-level systems and their semiclassical solution. Journal of Mathematical Physics, 55(12):122105, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hatano:2014:TRS

[2587] Naomichi Hatano and Gonzalo Ordonez. Time-reversal symmetric resolution of unity without background integrals in open quantum systems. Journal of Mathematical Physics, 55(12):122106, December 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boche:2014:PDF

[2588] H. Boche and J. Nötzel. Positivity, discontinuity, finite resources, and nonzero error for arbitrarily varying quantum channels. Journal of Mathematical Physics, 55(12):122201, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koenig:2014:HES

[2589] Robert Koenig and John A. Smolin. How to efficiently select an arbitrary Clifford group element. Journal of Mathematical Physics, 55(12):122202, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masanes:2014:ETD

[2590] Ll. Masanes, M. P. Müller, D. Pérez-García, and R. Augusiak. Entanglement and the three-dimensionality of the Bloch ball. Journal of Mathematical Physics, 55(12):122203, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shakeel:2014:HDQ

[2591] Asif Shakeel, David A. Meyer, and Peter J. Love. History dependent quantum random walks as quantum lattice gas automata. Journal of Mathematical Physics, 55(12):122204, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Furrer:2014:PMU

[2592] Fabian Furrer, Mario Berta, Marco Tomamichel, Volkher B. Scholz, and Matthias Christandl. Position-momentum uncertainty relations in the presence of quantum memory. Journal of Mathematical Physics, 55(12): 122205, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amburg:2014:SLS

[2593] Ilya Amburg, Roshan Sharma, Daniel M. Sussman, and William K. Wootters. States that “look the same” with respect to every basis in a mutually unbiased set. Journal of Mathematical Physics, 55(12):122206, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2622].

Mylonas:2014:NGF

[2594] Dionysios Mylonas, Peter Schupp, and Richard J. Szabo. Non-geometric fluxes, quasi-Hopf twist deformations, and nonassociative quantum mechanics. Journal of Mathematical Physics, 55(12):122301, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tlas:2014:ALM

[2595] Tamer Tlas. On the Ashtekar–Lewandowski measure as a restriction of the product one. Journal of Mathematical Physics, 55(12):122302, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dutsch:2014:DRP

[2596] Michael Dütsch, Klaus Fredenhagen, Kai Johannes Keller, and Katarzyna Rejzner. Dimensional regularization in position space and a forest formula for Epstein–Glaser renormalization. Journal of Mathematical Physics, 55(12):122303, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cuenin:2014:DGI

[2597] Jean-Claude Cuenin and Heinz Siedentop. Dipoles in graphene have infinitely many bound states. Journal of Mathematical Physics, 55(12): 122304, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deckert:2014:DEE

[2598] D.-A. Deckert and F. Merkl. Dirac equation with external potential and initial data on Cauchy surfaces. Journal of Mathematical Physics, 55(12):122305, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chanu:2014:TTW

[2599] Claudia Maria Chanu, Luca Degiovanni, and Giovanni Rastelli. The Tremblay–Turbiner–Winternitz system as extended Hamiltonian. Journal of Mathematical Physics, 55(12):122701, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cariglia:2014:CKT

[2600] M. Cariglia, G. W. Gibbons, J.-W. van Holten, P. A. Horvathy, and P.-M. Zhang. Conformal killing tensors and covariant Hamiltonian dynamics. Journal of Mathematical Physics, 55(12):122702, December 2014. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chanu:2014:EHS

[2601] Claudia Maria Chanu, Luca Degiovanni, and Giovanni Rastelli. Extensions of Hamiltonian systems dependent on a rational parameter. Journal of Mathematical Physics, 55(12):122703, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz:2014:LAP

[2602] Bogar Díaz, Daniel Higuita, and Merced Montesinos. Lagrangian approach to the physical degree of freedom count. Journal of Mathematical Physics, 55(12):122901, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lazo:2014:APD

[2603] Matheus J. Lazo and Cesar E. Krumreich. The action principle for dissipative systems. Journal of Mathematical Physics, 55(12):122902, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoang:2014:FST

[2604] Luan T. Hoang, Akif Ibragimov, and Thinh T. Kieu. A family of steady two-phase generalized Forchheimer flows and their linear stability analysis. Journal of Mathematical Physics, 55(12):123101, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2014:GSC

[2605] Yan Guo, Alexandru D. Ionescu, and Benoit Pausader. Global solutions of certain plasma fluid models in three-dimension. Journal of Mathematical Physics, 55(12):123102, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2014:SCP

[2606] Zhongyang Li. Spectral curves of periodic Fisher graphs. Journal of Mathematical Physics, 55(12):123301, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:2014:HTF

[2607] Ivan G. Avramidi. Heat trace and functional determinant in one dimension. Journal of Mathematical Physics, 55(12):123501, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:2014:PTH

[2608] Mahouton Norbert Hounkonnou, Sama Arjika, and Ezinvi Baloïtcha. Pöschl–Teller Hamiltonian: Gazeau–Klauder type coherent states, related statistics, and geometry. Journal of Mathematical Physics, 55(12): 123502, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2014:FOA

[2609] Lin Wang. First order approximation for quadratic dispersive equations by the renormalization group approach. Journal of Mathematical Physics, 55(12):123503, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cross:2014:WCP

[2610] Daniel J. Cross. When the charge on a planar conductor is a function of its curvature. Journal of Mathematical Physics, 55(12):123504, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arcus:2014:WTG

[2611] Constantin M. Arcus, Esmaeil Peyghan, and Esa Sharahi. Weyl’s theory in the generalized Lie algebroids framework. Journal of Mathematical Physics, 55(12):123505, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulze-Halberg:2014:TDP

[2612] Axel Schulze-Halberg and Barnana Roy. Time dependent potentials associated with exceptional orthogonal polynomials. Journal of Mathematical Physics, 55(12):123506, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakabekov:2014:BCO

[2613] Auzhan Sakabekov and Yerkanat Auzhani. Boundary conditions for the one-dimensional nonlinear nonstationary Boltzmann’s moment system equations. Journal of Mathematical Physics, 55(12):123507, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2623].

Fan:2014:LHV

[2614] Guangzhe Fan, Yucai Su, and Henan Wu. Loop Heisenberg–Virasoro Lie conformal algebra. Journal of Mathematical Physics, 55(12):123508, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hardin:2014:PDE

[2615] D. P. Hardin, E. B. Saff, and B. Simanek. Periodic discrete energy for long-range potentials. Journal of Mathematical Physics, 55(12):123509, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirkinis:2014:AMS

[2616] Eleftherios Kirkinis and Robert E. O’Malley, Jr. Amplitude modulation for the Swift–Hohenberg and Kuramoto–Sivashinski equations. Journal of Mathematical Physics, 55(12):123510, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tchokouansi:2014:IST

[2617] Hermann T. Tchokouansi, Victor K. Kuetche, and Timoleon C. Kofane. Inverse scattering transform of a new optical short pulse system. Journal of Mathematical Physics, 55(12):123511, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yao:2014:IST

[2618] Jie Yao, Anne-Cécile Lesage, Bernhard G. Bodmann, Fazle Hussain, and Donald J. Kouri. Inverse scattering theory: Inverse scattering series method for one dimensional non-compact support potential. Journal of Mathematical Physics, 55(12):123512, December 2014. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeier:2014:ESP

[2619] Robert Zeier and Thomas Schulte-Herbrüggen. Erratum: “Symmetry principles in quantum systems theory” [J. Math. Phys. 52, 113510 (2011)]. Journal of Mathematical Physics, 55(12):129901, December 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1044].

Gao:2015:CSB

[2620] Xin-Yi Gao. Comment on “Solitons, Bäcklund transformation, and Lax pair for the (2 + 1)-dimensional Boiti–Leon–Pempinelli equation for the water waves” [J. Math. Phys. 51, 093519 (2010)]. Journal of Mathematical Physics, 56(1):014101, January 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [449].

Han:2015:EGN

[2621] Lei Han and Hai-Jun Wang. Erratum: “General nonlocality in quantum fields” [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 56(2):029901, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [10].

Amburg:2015:ESL

[2622] Ilya Amburg, Roshan Sharma, Daniel M. Sussman, and William K. Wootters. Erratum: “States that ‘look the same’ with respect to every basis in a mutually unbiased set” [J. Math. Phys. 55, 122206 (2014)]. Journal of Mathematical Physics, 56(3):039901, March 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2593].

Sakabekov:2015:EBC

[2623] A. Sakabekov and Y. Auzhani. Erratum: “Boundary conditions for the one-dimensional nonlinear nonstationary Boltzmann’s moment system equations” [J. Math. Phys. 55, 123507 (2014)]. Journal of Mathematical Physics, 56(7):079901, July 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2613].

Zhang:2015:CSE

[2624] Yi Zhang and Ji-Bin Li. Comment on “Superposition of elliptic functions as solutions for a large number of nonlinear equations” [J. Math. Phys.

56, 032104 (2015)]. Journal of Mathematical Physics, 56(8):084101, August 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2236, 2662, 2666].

Bietenholz:2015:CTD

[2625] Wolfgang Bietenholz, Chryssomalis Chryssomalakos, and Marcelo Salgado. Comment on “The two dimensional motion of a particle in an inverse square potential: Classical and quantum aspects” [J. Math. Phys. 54, 053509 (2013)]. Journal of Mathematical Physics, 56(10):104101, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1839].

Kim:2015:ELC

[2626] Do-Hyung Kim. Erratum: “Lorentzian covering space and homotopy classes” [j. math. phys. 52, 122501 (2011)]. Journal of Mathematical Physics, 56(10):109901, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1062].

Harko:2015:ETW

[2627] T. Harko and M. K. Mak. Exact travelling wave solutions of nonlinear reaction-convection-diffusion equations — an Abel equation based approach. Journal of Mathematical Physics, 56(11):111501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marrocos:2015:ENL

[2628] Marcus A. M. Marrocos and Antônio L. Pereira. Eigenvalues of the Neumann Laplacian in symmetric regions. Journal of Mathematical Physics, 56(11):111502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alama:2015:VSW

[2629] Stan Alama, Lia Bronsard, and Xavier Lamy. Vortex structure in pwave superconductors. Journal of Mathematical Physics, 56(11):111503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2015:GSR

[2630] Jianjun Yuan. Global solution of the relativistic Vlasov–Maxwell–Chern– Simons system. Journal of Mathematical Physics, 56(11):111504, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manukure:2015:IES

[2631] Solomon Manukure and Wen-Xiu Ma. An integrable extension of a soliton hierarchy associated with so(3, R) and its bi-Hamiltonian formulation. Journal of Mathematical Physics, 56(11):111505, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:2015:TER

[2632] Jianzhi Han and Chunrui Ai. Three equivalent rationalities of vertex operator superalgebras. Journal of Mathematical Physics, 56(11):111701, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:2015:DBM

[2633] Hongmei Hu and Naihong Hu. Double-bosonization and Majid’s conjecture, (I): Rank-inductions of ABCD. Journal of Mathematical Physics, 56(11):111702, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Perlov:2015:APW

[2634] Leonid Perlov. Analog of the Peter–Weyl expansion for Lorentz group. Journal of Mathematical Physics, 56(11):111703, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lavicka:2015:FDP

[2635] Roman Lávička and Dalibor Šmíd. Fischer decomposition for polynomials on superspace. Journal of Mathematical Physics, 56(11):111704, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dhand:2015:ABR

[2636] Ish Dhand, Barry C. Sanders, and Hubert de Guise. Algorithms for SU(n) boson realizations and D-functions. Journal of Mathematical Physics, 56(11):111705, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fa:2015:LBS

[2637] Huanxia Fa, Junbo Li, and Yelong Zheng. Lie bialgebra structures on the deformative Schrödinger–Virasoro algebras. Journal of Mathematical Physics, 56(11):111706, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruillard:2015:LDR

[2638] Paul Bruillard, Liang Chang, Seung-Moon Hong, Julia Yael Plavnik, Eric C. Rowell, and Michael Yuan Sun. Low-dimensional representations of the three component loop braid group. Journal of Mathematical Physics, 56(11):111707, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qian:2015:SDF

[2639] Zhixin Qian. Static dielectric function with exact exchange contribution in the electron liquid. Journal of Mathematical Physics, 56(11):111901, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedmann:2015:QMD

[2640] Tamar Friedmann and C. R. Hagen. Quantum mechanical derivation of the Wallis formula for π. Journal of Mathematical Physics, 56(11): 112101, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabin:2015:ETB

[2641] Carlos Sabín, Pablo Barberis-Blostein, Cristopher Hernández, Robert B. Mann, and Ivette Fuentes. Effects of three-body collisions in a two-mode Bose–Einstein condensate. Journal of Mathematical Physics, 56(11): 112102, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhatia:2015:SEP

[2642] Rajendra Bhatia and Tanvi Jain. On symplectic eigenvalues of positive definite matrices. Journal of Mathematical Physics, 56(11):112201, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garner:2015:GLC

[2643] David Garner and Sanjaye Ramgoolam. The geometry of the light-cone cell decomposition of moduli space. Journal of Mathematical Physics, 56(11):112301, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

You:2015:HHG

[2644] Miman You, Nan Zhou, and Shuanhong Wang. Hom–Hopf group coalgebras and braided T -categories obtained from Hom–Hopf algebras. Journal of Mathematical Physics, 56(11):112302, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delaFuente:2015:UDM

[2645] Daniel de la Fuente, Alfonso Romero, and Pedro J. Torres. Unchanged direction motion in general relativity: The problems of prescribing acceleration and the extensibility of trajectories. Journal of Mathematical Physics, 56(11):112501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Izumiya:2015:LHA

[2646] Shyuichi Izumiya. Lightlike hypersurfaces along spacelike submanifolds in anti-de Sitter space. Journal of Mathematical Physics, 56(11):112502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khuri:2015:IBS

[2647] Marcus A. Khuri. Inequalities between size and charge for bodies and the existence of black holes due to concentration of charge. Journal of Mathematical Physics, 56(11):112503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bihun:2015:PZG

[2648] Oksana Bihun and Francesco Calogero. Properties of the zeros of generalized basic hypergeometric polynomials. Journal of Mathematical Physics, 56(11):112701, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakhnovich:2015:DSD

[2649] A. L. Sakhnovich. Dynamical and spectral Dirac systems: response function and inverse problems. Journal of Mathematical Physics, 56(11): 112702, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nold:2015:NSM

[2650] Andreas Nold, Martin Oberlack, and Alexei F. Cheviakov. On new stability modes of plane canonical shear flows using symmetry classification. Journal of Mathematical Physics, 56(11):113101, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Adame:2015:EVG

[2651] J. Adame and S. Warzel. Exponential vanishing of the ground-state gap of the quantum random energy model via adiabatic quantum computing. Journal of Mathematical Physics, 56(11):113301, November 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:FCS

[2652] Yongxin Chen, Tryphon T. Georgiou, and Michele Pavon. Fast cooling for a system of stochastic oscillators. Journal of Mathematical Physics, 56(11):113302, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hachem:2015:SMI

[2653] Walid Hachem, Aris Moustakas, and Leonid A. Pastur. The Shannon’s mutual information of a multiple antenna time and frequency dependent channel: An ergodic operator approach. Journal of Mathematical Physics, 56(11):113501, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daletskii:2015:PTC

[2654] Alexei Daletskii, Yuri Kondratiev, and Yuri Kozitsky. Phase transitions in continuum ferromagnets with unbounded spins. Journal of Mathematical Physics, 56(11):113502, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tamizhmani:2015:SRO

[2655] K. M. Tamizhmani, K. Krishnakumar, and P. G. L. Leach. Symmetries and reductions of order for certain nonlinear third- and second-order differential equations with arbitrary nonlinearity. Journal of Mathematical Physics, 56(11):113503, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liemert:2015:FST

[2656] André Liemert and Alwin Kienle. Fundamental solution of the tempered fractional diffusion equation. Journal of Mathematical Physics, 56(11): 113504, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2015:ESC

[2657] Shaowei Liu. Eigenvalues of the string constraints for the Kadomtsev– Petviashvili hierarchy. Journal of Mathematical Physics, 56(11):113505, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sako:2015:GTN

[2658] Akifumi Sako, Toshiya Suzuki, and Hiroshi Umetsu. Gauge theories on noncommutative CP^{N} and Bogomol’nyi–Prasad–Sommerfield-like equations. Journal of Mathematical Physics, 56(11):113506, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baulieu:2015:WLF

[2659] L. Baulieu, N. L. Holanda, and F. Toppan. A world-line framework for 1D topological conformal σ-models. Journal of Mathematical Physics, 56(11):113507, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrogowska:2015:MLP

[2660] Alina Dobrogowska. R-matrix, Lax pair, and multiparameter decompositions of Lie algebras. Journal of Mathematical Physics, 56(11):113508, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strachan:2015:IFA

[2661] Ian A. B. Strachan and Dafeng Zuo. Integrability of the Frobenius algebra-valued Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 56(11):113509, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:2015:RCS

[2662] Avinash Khare and Avadh Saxena. Response to “Comment on ‘Superposition of elliptic functions as solutions for a large number of nonlinear equations’”” [J. Math. Phys. 56, 084101 (2015)]. Journal of Mathematical Physics, 56(11):113510, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2624].

Prodan:2015:GWF

[2663] Emil Prodan. On the generalized Wannier functions. Journal of Mathematical Physics, 56(11):113511, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moskow:2015:NEA

[2664] Shari Moskow. Nonlinear eigenvalue approximation for compact operators. Journal of Mathematical Physics, 56(11):113512, November 2015.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goulart:2015:CAI

[2665] E. Goulart. Characteristic analysis for integrable soliton models on twodimensional target spaces. Journal of Mathematical Physics, 56(11): 113513, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abraham-Shrauner:2015:CCS

[2666] Barbara Abraham-Shrauner. Comment on “Comment on ‘Superposition of elliptic functions as solutions for a large number of nonlinear equations”’ [J. Math. Phys. 56, 084101 (2015)]. Journal of Mathematical Physics, 56(11):114101, November 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2624].

Iacob:2015:ESS

[2667] Felix Iacob and Marina Lute. Exact solution to the Schrödinger’s equation with pseudo-Gaussian potential. Journal of Mathematical Physics, 56(12):121501, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masaki:2015:EUT

[2668] Satoshi Masaki and Takayoshi Ogawa. Existence and uniqueness of two dimensional Euler–Poisson system and WKB approximation to the nonlinear Schrödinger–Poisson system. Journal of Mathematical Physics, 56(12):121502, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2015:VPS

[2669] Zili Chen. The Vlasov–Poisson system with low steady spatial asymptotics. Journal of Mathematical Physics, 56(12):121503, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2015:OSI

[2670] Stephen C. Anco, Abdus Sattar Mia, and Mark R. Willoughby. Oscillatory solitons of U(1)-invariant modified Korteweg–de Vries equations. II. Asymptotic behavior and constants of motion. Journal of Mathematical Physics, 56(12):121504, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheviakov:2015:CCL

[2671] A. F. Cheviakov and S. St. Jean. A comparison of conservation law construction approaches for the two-dimensional incompressible Mooney– Rivlin hyperelasticity model. Journal of Mathematical Physics, 56(12): 121505, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tu:2015:IFR

[2672] Guizhang Tu. An inversion formula for recursion operator of KP hierarchy. Journal of Mathematical Physics, 56(12):121506, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2015:CGS

[2673] Haisheng Li and Jiancai Sun. On certain generalizations of the Schrödinger–Virasoro algebra. Journal of Mathematical Physics, 56(12): 121701, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anguelova:2015:MB

[2674] Iana I. Anguelova. Multilocal bosonization. Journal of Mathematical Physics, 56(12):121702, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Werry:2015:MED

[2675] Jason L. Werry, Mark D. Gould, and Phillip S. Isaac. Matrix elements and duality for type 2 unitary representations of the Lie superalgebra gl(m|n). Journal of Mathematical Physics, 56(12):121703, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdul-Rahman:2015:UAL

[2676] Houssam Abdul-Rahman and Günter Stolz. A uniform area law for the entanglement of eigenstates in the disordered XY chain. Journal of Mathematical Physics, 56(12):121901, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:2015:RSS

[2677] A. G. Smirnov. Reduction by symmetries in singular quantummechanical problems: General scheme and application to Aharonov– Bohm model. Journal of Mathematical Physics, 56(12):122101, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chowdhury:2015:WFN

[2678] S. Hasibul Hassan Chowdhury and S. Twareque Ali. Wigner functions for noncommutative quantum mechanics: A group representation based construction. Journal of Mathematical Physics, 56(12):122102, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landsman:2015:CRT

[2679] Klaas Landsman. On the Colbeck–Renner theorem. Journal of Mathematical Physics, 56(12):122103, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Orsucci:2015:HP

[2680] Davide Orsucci, Daniel Burgarth, Paolo Facchi, Hiromichi Nakazato, Saverio Pascazio, Kazuya Yuasa, and Vittorio Giovannetti. Hamiltonian purification. Journal of Mathematical Physics, 56(12):122104, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lange:2015:DTS

[2681] Rutger-Jan Lange. Distribution theory for schrödinger’s integral equation. Journal of Mathematical Physics, 56(12):122105, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braun:2015:SCS

[2682] Carol Braun, Feifei Li, Anupam Garg, and Michael Stone. The semiclassical coherent state propagator in the Weyl representation. Journal of Mathematical Physics, 56(12):122106, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Valtancoli:2015:CTM

[2683] P. Valtancoli. Canonical transformations and minimal length. Journal of Mathematical Physics, 56(12):122107, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aleixo:2015:NON

[2684] A. N. F. Aleixo and A. B. Balantekin. Normal ordering for nonlinear deformed ladder operators and the f-generalization of the Stirling and Bell numbers. Journal of Mathematical Physics, 56(12):122108, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cariolaro:2015:FEM

[2685] Gianfranco Cariolaro and Gianfranco Pierobon. Fock expansion of multimode pure Gaussian states. Journal of Mathematical Physics, 56(12): 122109, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vubangsi:2015:WPD

[2686] M. Vubangsi, M. Tchoffo, L. C. Fai, and Yu. M. Pisma’k. Wave packet dynamics for a system with position and time-dependent effective mass in an infinite square well. Journal of Mathematical Physics, 56(12):122110, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Band:2015:ANC

[2687] Ram Band, Gregory Berkolaiko, and Tracy Weyand. Anomalous nodal count and singularities in the dispersion relation of honeycomb graphs. Journal of Mathematical Physics, 56(12):122111, December 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blass:2015:OAF

[2688] Andreas Blass, Alex Bocharov, and Yuri Gurevich. Optimal ancilla-free Pauli + V circuits for axial rotations. Journal of Mathematical Physics, 56(12):122201, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carles:2015:HOS

[2689] Rémi Carles, Wolfgang Lucha, and Emmanuel Moulay. Higher-order Schrödinger and Hartree–Fock equations. Journal of Mathematical Physics, 56(12):122301, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimcik:2015:SMP

[2690] Ctirad Klimčík. On supermatrix models, Poisson geometry, and noncommutative supersymmetric gauge theories. Journal of Mathematical Physics, 56(12):122302, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holland:2015:AOP

[2691] Jan Holland and Stefan Hollands. Associativity of the operator product expansion. Journal of Mathematical Physics, 56(12):122303, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2015:EUS

[2692] Xuetao Yang and Quanxin Zhu. Existence, uniqueness, and stability of stochastic neutral functional differential equations of Sobolev-type. Journal of Mathematical Physics, 56(12):122701, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2015:HRG

[2693] Yantao Guo, Ming Wang, and Yanbin Tang. Higher regularity of global attractors of a weakly dissipative fractional Korteweg de Vries equation. Journal of Mathematical Physics, 56(12):122702, December 2015. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:2015:EQP

[2694] Lin Lu and Xuemei Li. Existence of quasi-periodic solutions of fast excited van der Pol–Mathieu–Duffing equation. Journal of Mathematical Physics, 56(12):122703, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2015:BUS

[2695] Jiaoyang Huang, Benjamin Landon, and Horng-Tzer Yau. Bulk universality of sparse random matrices. Journal of Mathematical Physics, 56(12):123301, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouch:2015:EPE

[2696] Gabriel Bouch. The expected perimeter in Eden and related growth processes. Journal of Mathematical Physics, 56(12):123302, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouch:2015:CTS

[2697] Gabriel Bouch. Complex-time singularity and locality estimates for quantum lattice systems. Journal of Mathematical Physics, 56(12):123303, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buzinski:2015:ASC

[2698] David Buzinski and Elizabeth Meckes. Almost sure convergence in quantum spin glasses. Journal of Mathematical Physics, 56(12):123304, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirsten:2015:BFK

[2699] Klaus Kirsten and Yoonweon Lee. The Burghelea–Friedlander– Kappeler–gluing formula for zeta-determinants on a warped product manifold and a product manifold. Journal of Mathematical Physics, 56(12):123501, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zaba:2015:NIF

[2700] Mariusz Żaba and Piotr Garbaczewski. Nonlocally induced (fractional) bound states: Shape analysis in the infinite Cauchy well. Journal of Mathematical Physics, 56(12):123502, December 2015. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Latremoliere:2015:CNT

[2701] Frédéric Latrémolière. Curved noncommutative tori as Leibniz quantum compact metric spaces. Journal of Mathematical Physics, 56(12):123503, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2015:LSG

[2702] Qing Huang and Shoufeng Shen. Lie symmetries and group classification of a class of time fractional evolution systems. Journal of Mathematical Physics, 56(12):123504, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2714].

Nagoya:2015:ICB

[2703] Hajime Nagoya. Irregular conformal blocks, with an application to the fifth and fourth Painlevé equations. Journal of Mathematical Physics, 56(12):123505, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeAngelis:2015:EPI

[2704] Valerio De Angelis and Christophe Vignat. Euler polynomials and identities for non-commutative operators. Journal of Mathematical Physics, 56(12):123506, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dykema:2015:PNS

[2705] Kenneth J. Dykema and Vern Paulsen. Publisher’s note: “Synchronous correlation matrices and Connes’ embedding conjecture” [J. Math. Phys. 57, 015214 (2016)]. Journal of Mathematical Physics, 56(12):129901, December 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Correction to footnoted author in [2706].

Dykema:2016:SCM

[2706] Kenneth J. Dykema and Vern Paulsen. Synchronous correlation matrices and Connes’ embedding conjecture. Journal of Mathematical Physics, 57 (1):015214, January 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paliathanasis:2016:CCL

[2707] A. Paliathanasis and P. G. L. Leach. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)] and its erratum [j. math. phys. 55, 059901 (2014)]. Journal of Mathematical Physics, 57(2):024101, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1836].

Ghoumaid:2016:CAS

[2708] A. Ghoumaid, F. Benamira, and L. Guechi. Comment on “Approximate solutions of the Dirac equation for the Rosen–Morse potential including the spin-orbit centrifugal term” [j. math. phys. 51, 023525 (2010)]. Journal of Mathematical Physics, 57(2):024102, February 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [102].

Magdziarz:2016:CFP

[2709] Marcin Magdziarz and Tomasz Zorawik. Comment on “Fokker–Planck equations for nonlinear dynamical systems driven by non-Gaussian Lévy processes” [J. Math. Phys. 53, 072701 (2012)]. Journal of Mathematical Physics, 57(3):034101, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1330].

Frewer:2016:CAE

[2710] Michael Frewer, George Khujadze, and Holger Foysi. Comment on “Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation” [J. Math. Phys. 54, 072901 (2013)]. Journal of Mathematical Physics, 57(3):034102, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1908] and response [2711].

Waclawczyk:2016:RCA

[2711] Marta Waclawczyk and Martin Oberlack. Response to “Comment on ‘Application of the extended Lie group analysis to the Hopf functional formulation of the Burgers equation”’ [J. Math. Phys. 57, 034102 (2016)]. Journal of Mathematical Physics, 57(3):034103, March 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1908, 2710].

Chuah:2016:ECS

[2712] Meng-Kiat Chuah and Mingjing Zhang. Erratum: “Contragredient supersymmetric pairs” [j. math. phys. 54, 091702 (2013)]. Journal of Mathematical Physics, 57(10):109901, October 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1989].

Carvalho:2016:EUM

[2713] Paulo R. S. Carvalho and Marcelo M. Leite. Erratum: “Unconventional minimal subtraction and Bogoliubov–Parasyuk–Hepp– Zimmermann method: Massive scalar theory and critical exponents” [J. Math. Phys. 54, 093301 (2013)]. Journal of Mathematical Physics, 57(11):119901, November 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2006].

Singla:2017:CLS

[2714] Komal Singla and R. K. Gupta. Comment on “Lie symmetries and group classification of a class of time fractional evolution systems” [J. Math. Phys. 56, 123504 (2015)]. Journal of Mathematical Physics, 58 (5):054101, May 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2702].

Jiao:2017:CEC

[2715] Li Guang Jiao and Li Rong Zan. Comment on “Entropy and complexity analysis of hydrogenic Rydberg atoms” [j. math. phys. 54, 052109 (2013)]. Journal of Mathematical Physics, 58(10):104101, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1813] and response [2716].

Lopez-Rosa:2017:RCE

[2716] S. López-Rosa, I. V. Toranzo, P. Sánchez-Moreno, and J. S. Dehesa. Response to “Comment on ‘Entropy and complexity analysis of hydrogenic Rydberg atoms”’ [j. math. phys. 58, 104101 (2017)]. Journal of Mathematical Physics, 58(10):104102, October 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1813, 2715].

Arrizabalaga:2018:ESA

[2717] Naiara Arrizabalaga, Javier Duoandikoetxea, and Luis Vega. Erratum: “Self-adjoint extensions of Dirac operators with Coulomb type singularity” [J. Math. Phys. 54, 041504 (2013)]. Journal of Mathematical Physics, 59(7):079902, July 2018. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1752].

Iacono:2020:CCL

[2718] Roberto Iacono. Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [j. math. phys. 54, 053506 (2013)]. Journal of Mathematical Physics, 61(4):044101, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1836].

Chandrasekar:2020:RCC

[2719] V. K. Chandrasekar, A. K. Tiwari, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. Response to “Comment on “Classification of Lie point symmetries for quadratic Liénard type equation ẍ + f(x)ẋ^{2} + g(x) = 0” [J. Math. Phys. 54, 053506 (2013)]” [j. math. phys. 61, 044101 (2020)]. Journal of Mathematical Physics, 61(4):044102, April 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moya-Cessa:2020:CTD

[2720] H. M. Moya-Cessa and J. Récamier. Comment on “Time-dependent coupled harmonic oscillators” [j. math. phys. 53, 052101 (2012)]. Journal of Mathematical Physics, 61(11):114101, November 2020. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1253].

DeNittis:2020:EEL

[2721] Giuseppe De Nittis and Max Lein. Erratum: “Exponentially localized Wannier functions in periodic zero flux magnetic fields” [j. math. phys. 52, 112103 (2011)]. Journal of Mathematical Physics, 61(11):119901, November 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1018].